A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2005–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 September 2019 Version 1.04

Title word cross-reference

(0, ∞) [BE07]. (0, m) [dB05b, dB05a]. (A, η) [DF08]. (d + 1) [KV09a]. (d u/d r) [Moh05]. 0 [MVVA06]. 1 [Hsu05, LP07c, MDT09]. 1/9 [SS09e]. 2 [BCH06, CGRGVP08, DL05, Ehr06, GC06, HV06a, KSE09, LN09, LYT07, NGS05, TS06a, WHG07, XXL08]. 3 [BGP07, BFGM06, BCK07, Cim08, GMM09, GF08, KD05, LLL06, LT07, Li06a, LC08c, MMM08, MK09, Scu09, SBO09, VVE08a, VVE08b, ZGV06]. 6 [DS09a, HKA07]. 8 [LW06b]. [−∞, 0] [BGVHN05]. [0, 1)^{s} [LP07b]. [0, 1] [Has09a]. [3/2] [GMFB06]. ] − \frac{π}{2} , \frac{π}{2} [ [Cho05]. ^{1} [LP06b]. ^{2} [HS09a]. ^{X} [TYH08]. _{0} [ZZ09a]. _{10}ϕ_{9} [LV07]. _{2} [AM09a, CZ08b]. _{2}F_{2} [Par05b]. _{3}F_{2} [Ext99, Mil06]. _{3}F_{2}(1) [Mil07]. A [FD09]. A_{T,S}^{(2)} [SCG08]. A^{∗}X + X^{∗}A = B [Djo07].

A^{H}XB = C [yPyHZ07]. A_{T,S}^{(2,3)} [YL09a]. α [FSL09, Jay08, MPR07]. AXB = C [HYG08a]. B [PJ06a]. BaF_{2} [BRVWM07]. C [LP06b, FLS06]. C^{∗} [Par05c]. C^{1} [CHNZ08, DL05, MB06, PS08]. C^{2} [CJ06, Han05, YS06, DS09a]. C^{M} [CD07b]. C^{r} [BFGP08]. O(1) [BG07f]. S [Mor07]. S_{m} [Sid05]. D [QWW08, ZW09a, Ant09, CZ07, CLO09, Kok05]. {D_{n}} [WK07b]. Δ [AGMMB05]. δ^{2} [AGM09]. E [KKW06]. ℓ_{2} [MYY06]. ϵ [BMW08]. η [PZ08]. η_{j} [Cof09b]. F [Abd08, CZ08b]. F (a, b; c; z) [FLS06]. f(x) sin(w/x^{r}) [Has09a]. G [Ant08, HS09a, LW08d]. G^{1} [MW08, MW09c, WM09]. Γ [Yam08]. Γ(1/x)

[Alz06]. Γ(x) [Alz06]. GMRES(m) [BJK09]. H [AU09, BISS05a, BS07, CK05, HR05c, LF08, MPS05, PJ06a, Yun08a, yZXL09, ZHG^{+}09, Din05a, YG09b]. H_{A} [Tur09]. H_{∞} [MZC08]. H′_{ν}(x) + iζH_{ν}(x) = 0 [SA07a]. H_{q,ω} [Pet07b]. hp [ER09, Sur05, GGD08, SV07b, VS08]. I [BFP09, KN06, Tem09]. ijk [Moh09]. ∫_{a}^{b} f(x)e^{iωg(x)}dx [Xia07b]. J [CK09b]. K [LF07b, LF08, ÁNPQ05, FY06, Kil07, LRH07, MW09b, XY05]. K(2, 2) [Waz07b]. K(n, n) [AEET09, LWW08]. K(p, q) [TY07a]. K + 1 [CHJZ09]. K_{n=1}^{∞}a_{n}/1 [LW05a]. L [BBR05, HJR07, tLxLlW07, RS08b, RH09, WtL09, YH09b, Yun08b]. L^{1} [Las09]. L^{2} [GNP09]. l^{2} × l^{2} [LZ05]. L_{1} [CFL05, WFL07]. L_{2} [HW08, PVC06, VC06]. L_{∞} [PVC06]. L_{p} [CW08c]. LU [AKK06, Moh09]. M [CW08d, KBA09, LGYH09, LC09b, MS07d, WS09a, Yun05b, ZM09b, HY09a, HJK09, JG09, LZ08a, LZ08b, LZ09b, LZ09c, LLW09, Pre09, Sal09, SS09b, ZC08a]. R^{1} [RS05c]. R^{2} [Roo06]. R^{N} [AV06, EE06, Koh07, SFV09]. S^{q−1} [Fer07]. N [CKT09, GPTT05, KL08a, CKWZ07, CHCW05, Dub09, FWY09, JG09, Lei08b, LZ09c, LM07b, Par05c, Sto06, WW07]. O(k^{2} + kh^{2} + h^{4}) [MS07c]. ω [Yun05b]. P [DWW06, LW07, LP07c, PZ08, Tsi06, Van06a, Van07a, ZZ09a, BB05, BE07, DF08, Gal09, GJL08, HM09a, He05, HWY09, Hon07, Liu08a, LOA05a, LOA05b, LJ09, Luo07, ML09, WSM^{+}09, XKF06, YG09b, ZFG08]. P_{∗}(κ) [Cho08, WB09]. P_{0} [CP08b, MC08]. P_{m} = (I + S_{m}) [KN09]. ϕ [Lin09a]. q [AÁN07, AGRZ05, CJ07, CZ08b, CS09b, DF08, DO06, IS09, OS07a, PV05, RK08b, YH05]. QR [DV06a, DV06b, Gem05, SEAA08]. QZ [VVV08]. R [CGP09, EH06, HR05b, Yua09, Ant09, Pre09]. sinh [EJ07]. T [CG05, DCF06, Mar07c, WWP09, SO05]. θ [CEJ09, Gan07, KVK08, RS05c, WYL09b]. V A + A^{∗}V = μV [Koh08]. VPAStab(J, L) [GM09c]. W [Ehr06, FL08b, Nov08]. X + A^{∗}X^{−n}A = Q [Iva06]. X + A^{∗}X^{−α}A = Q [yPESlZ07]. X + A^{⋆}X^{−1}A = Q [ESAD05]. X − A\overline{X}F = C [WWD09]. X − AXF = C [WWD09]. X − ∑_{i=1}^{m} A_{i}^{∗}X^{r}A_{i} = Q [DL09c]. x_{n+1} = (α − βx_{n})/(γ + g(x_{n−k})) [ZSZ05]. y′′ = f(x, y) [FT08, Van07b]. y′′ = f(x, y, y′) [VAR06]. Z [CHS08, Kon07, NT06, WHF07]. ζ(2n + 1) [DH06].

-accretive [AU09, DF08, PZ08]. -Adaptive [GGD08]. -adic [HM09a, YH05]. -algebra [Las09]. -algebras [Par05c]. -algorithm [BMW08, DV06a, DV06b]. -analogues [CZ08b]. -asymptotically [Tem09]. -based [Gem05, KKW06]. -Bernoulli [RK08b]. -cardinality [FY06]. -centrohermitian [LF07b]. -centrosymmetric [LF08]. -constant [HKA07]. -convergence [Yam08]. -curve [HJR07, RS08b, RH09]. -curves [PJ06a]. -D [Ehr06, Li06a, LP07c, LC08c, NGS05, XXL08]. -decomposition [AKK06]. -difference [AGRZ05, DO06, IS09]. -dimensional [BFGM06]. -discrepancy [VC06]. -eigenvalues [QWW08]. -error [GNP09]. -expansion [Abd08]. -FEM [SV07b, VS08]. -fold [XY05]. -functions [MS07d]. -gap [ZW09a]. -Ginzburg [Lei08b]. -invexity [MPR07]. -Jacobi [CS09b, PV05]. -Laguerre [OS07a]. -Laplacian [BE07, Gal09, GJL08, He05, HWY09, Hon07, Liu08a, LOA05a, LOA05b, LJ09, ML09, WSM^{+}09, ZFG08]. -like [SEAA08]. -linearly [CGP09]. -matrices [LC09b, Yun08b, CHS08, CK05, HR05c, tLxLlW07, LF08, NT06, WHF07, WtL09, WS09a, Yun08b, yZXL09, ZM09b]. -matrix [CW08d, LGYH09, Mar07c, Yun05b, Yun08a, ZHG^{+}09]. -maximal [Din05a].

-meshes [DCF06]. -method [CEJ09, VVV08]. -methods [Gan07, KVK08, Nov08, RS05c, WYL09b]. -mixing [Lin09a]. -monotone [FD09]. -node [LW06b]. -norms [MYY06]. -order [EH06, HR05b]. -orthogonal [BBR05, Pet07b]. -orthogonality [CLO09, Kok05, CZ07]. -pencil [KV09a]. -periodic [SO05]. -point [CHCW05, DS09a, HY09a, HJK09, JG09, LZ08a, LZ08b, LZ09b, LZ09c, LLW09, Sal09, SS09b, ZC08a, GPTT05]. -policy [CKT09]. -polynomials [AÁN07]. -pre-invex [Ant08]. -preinvex [LW08d]. ’-problem [SS09e]. -proximal [KBA09]. -rational [LP06b]. -regular [BS07]. -series [CJ07, CZ08b, LV07]. -sesquilinear-quadratic [Par05c]. -shaped [YH09b]. -simplex [Sto06]. -Sobolev [AGMMB05]. -species [WW07]. -sphere [LM07b]. -splines [BISS05a, CJ06]. -stability [FL08b]. -stable [MVVA06, Van07a, DWW06, LW07, Tsi06, Van06a]. -step [DF08]. -surfaces [BFGP08]. -symmetric [Yua09]. -th [BB05, CKWZ07, Dub09, FSL09, JG09, MW09b, Luo07]. -th-copy [LRH07]. -th-order [FWY09, LZ09c]. -Toeplitz [ÁNPQ05]. -transformation [Ehr06]. -type [GMFB06, Ant09]. -uniformly [DF08]. -univexity [Jay08]. -version [XKF06]. -waves [LP07c].

/1 [WK07b]. /G/1 [TYH08].

1 [AA09b, ÉT08, TYH08, Ver05, WWP09, WK07b]. 173 [KYF05a]. 1734 [Gau08]. 184 [ByXy06]. 189 [GPPRB07].

204 [AMO08]. 206 [Yun08b]. 214 [yCyZ09]. 217 [SZ09b]. 219 [LC09b]. 224 [LHL11, TSC10, ZJ12]. 231 [Dai11, LB10]. 233 [Asl10, vDvFZ13]. 2nd [Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano07-39, Ano07-40, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano09z].

65th [VAW07].

= [BCH06, Ben07].

ABC [GST06]. Abel

[CC06a, CC08, CJ07, pLT07, SH09]. absolute [BB08a, DZ09a, Dmi08]. absorbing [Tar09]. absorption [Fer09]. abstract [LCA07]. accelerate [Koç08b]. Accelerated [And09b, HR05a, Kou08]. Accelerating [SaLJ08]. acceleration [TF07, VRL07, VR08, ZW09d]. accommodating [BdAR08]. according [Sie07a]. accretive [AU09, DF08, PZ08]. accumulation [ZL09b]. Accuracy [VR08, BG07d, DZ09a, FZL07, Har08a, JF06, JR08, LCZ09, NGS05, YY07a]. Accurate [DAE05, BD07, CH09, DRT09, LT07, LG07, Rad05, RVAD08]. accurately [AOSY05]. acoustic [Dem07a, DDLM07, GLS07, HMPR07, HKM08, Jen07, LmLgF09, RDH07, ZP06]. acoustics [TS05]. activated [KRAH07, ZBB^{+}05]. activation [CG05, Liu07]. activations [Moh07]. active [EM08a, IV08, Liu05]. acute [YY08]. Adams [Gar09, JL06a]. adaptation [GM07b, MS09c]. Adaptive [BN08b, Ban08, FLMR07, FP08, HT06, HO09, JF06, KVK08, Lai09b, Lan05, LT06b, LRT06, Par08b, SWZ05, SW06a, TL09, AAPP07, AS09b, ABG06b, ABP08, BH07a, BH09, BKB06, CGPM08, CMV09, CCCH09, DA09, DVCH06, DDL09, Gan08, GKMS09, HKB^{+}05, HJR07, HAS09b, HRGD08, Hos09, HGS06, LS06a, LY05a, NQDJ06, PP09, SS09c, Sha08, TYH08, WSST05, Wu07, YW08, ZZZ06, ZT06a, GGD08]. add [MS05b]. added [Hat09]. adding [AEET09]. additional [Bno07a]. additive [LZ06, Toc09, Wik08]. adhesion [CG05]. ADI [BLT09, FZL07, HL07a, TG07]. adiabatic [SS08]. adic [HM09a, YH05]. adjacencies [PR05]. Adjoint [MSB05, AGH06, BR07b, Car05b, Cha07, EGFOA09, EGFO09, EMMP07, KN06, KM09c, LP06a, MV08, MS07d]. adjoints [AS09b]. ADM [GJM07]. Adomian [Abb07b, AHC05, BSC07b, Eba09, Has06, Hos06, HLL08a, JDG06, Jan08, Li09a, Swe07a, Waz07a, Zha05]. advances [Bru09, GM09b]. advection [ABR07, CVB07, HPS06, Kan08, MBCV09, RVAD08, Roo06, SK05, SKAW09, Wil06]. advection-reaction [ABR07]. advective [DJ07a]. aeration [AHM09]. aeroacoustic [LN07a]. aerodynamic [FGB07, PLGS09]. aeroelasticity [Svá08a]. Affine [Zhu06, Zhu08a, CKC^{+}07, HYYF05, Lan08b, Lan08a, MK07, SLMW06, Zhu05a, Zhu05b]. affinely [KR07]. after [Sim07a]. against [Zin05]. Age [KS09, LkLkP09, RHQ06, Zha08]. age-dependent [LkLkP09, RHQ06]. age-structured [Zha08]. Age-time [KS09]. aggregation [GW09]. ahead [GM09c]. aided [IOS07]. AIDS [CLGG09]. AINV [RT08]. air [AMO07, AMO08, ZMAZ09]. aircraft [DDL09, KGB^{+}07]. airfoil [DFHS08, Gon07, MT05]. Airy [TV09a]. al. [Mil07]. ALE [SST09]. algebra [GW09, Las09, PPG07, sR07]. Algebraic [ÁNPQ05, AKLT09, BG07b, CZ09, DSW08, GGT07, GL06, Guo06, HS07a, HS09b, Hos06, Kuz08, LZZD09, PS05, SS05, SEAA08, SLMW06, Wan08b, rWlC06, XZS07, XFgS06]. algebraic-logarithmic [HS07a]. algebras [Par05c]. Algorithm [XH08, Yi09, AM09b, AMV06, And09b, Att05, BGRS05, BMW08, Bog05a, BC06, BdAR08, Bru08, BD07, CG08a, CDM07a, CC08, CM05b, CCJ07, CW08b, yCyZ08, yCyZ09, CGP09, CNN07, Cho08, Cof09a, DRT09, DV06a, DV06b, DF08, DC09c, DHS09, EIM^{+}09, EM09, Fan09, FL06c, GXgL05, Gem05, HPS06, HJR07, HM09a, He06a, HHC09, HZW08, JLR06, JTHZ05, bJQlZ07, bJxC09, Kaz06, KBA09, KCLW08, KSE09, LCKK06, LFS08, LH09, LZWQ09, Lin08, LXX08, LW08a, LYW08, LCZ09, LL07b, LHR09, MVV08, MGS08, PP09, QCK09a, SVC09, Sak09, Sha09a, SDP09, SJZ09, SLN06, TGB08a, TH07, TLQ08, VRBF07, VMV07, VB05c, VN09, WZ07, WOK05, Wu07, XZ08, YJGL07, YY07b, YCJ09, ZGSF08, ZW09a, ZSGF09, ZZ09b, ZZ08b, ZCL09, Zhu05a, Zhu05c, Zhu05b, Zhu06, Zhu08c]. algorithmic [ADDdM05, CLS08, Röb08]. Algorithms [CS06b, CW08c, AS09b, AT09c, BMS05, BNP07, BW07a, CMST06, CY08a, CY08c, CZ08a, CHLW08, DP08a, DFZ07, Din05b, Dub09, DS05c, Dus05, FP08, GPS09, KS07b, KK08, Kon07, LM08, Mai08a, MV05a, MVVA06, Mat07a, McC05, NVD08, PLGS09, Qui07, Rom09, SW06b, VAMVC05, WB09, YN08, YSY09]. aliasing [GPV05]. allelopathy [CLCL07, LC06b]. allocation [GBR09, WW09, YW08]. alloys [Ahm07]. Almost [Hsu05, Liu07, PDM08, RS05c, CV09, KN06, LH07a, MLH08, Sch07b, ST07]. along [HeM05]. alternate [Mur07]. alternating [Qin09, Yu06, Zhu07]. alternation [Bla05]. Alternative [JM07a, SdJMH09, YW09a]. amenability [Las09]. American [BS08b, HLJ09, KVK08, LWH09b, NST08, TGB08a, ZDC07]. amplification [CC06b, Van07b]. amplification-fitted [Van07b]. amplitude [CU09a, OW06a, SC07a]. analog [FHtMP06]. analogues [CZ08b, Moh08]. analyses [BG07a, FH09b, GP08]. Analysis [BGJ07, BD06, CFHM06, Chr09, Cui08, DB08, Dol08, ER09, Fai07, GSBB09, HK07a, JW08b, LN09, LTC06, Nag09, NSC07, Tar09, TY07b, TV09b, WNZ09, WK07b, YA08, AHM09, ABD^{+}05, AB09b, BP06, BF08a, Bou07a, BG07e, Bru09, BSKS07, CLGG09, CFV06, CY08c, CHS07, CNN07, Cho05, Civ07, Cop05, CHV^{+}08, CDM07b, DP08a, dCDND06, DF09, Dom09, EKEHR08, EJ07, ÉT08, EPP^{+}09, EKL^{+}07, FFK05, GVC^{+}08, GHT06, HCL08b, HeM05, HyLC06, HGS08, HL09c, IV08, Jia06, JW08a, KJ09, Kar06, KM09a, KG09, KKLY06, KS06, Koi05, KLS^{+}07, Li06a, Li07a, Li07b, LZ09a, LY09, LWMT07, LLZ09, LP09, LL07b, LRB05, MDP08, MSB05, Meh05, MMS^{+}09, Pie09, RK08a, Sen07, SC06a, Tac07, Tan06, Tor05, TM08, VMMM09, Val09, VRL07, VR08, VN09, WPC08, WCW09, WLL09, WZ06, WL09b]. analysis [Won09a, WH09b, XC08, XGM09, YC06, YZ09a, YZL08, YHH05, ZBB^{+}05, ZC08b, vdBHPS07]. Analytic [GS06c, HAS08, Soh05, XFgS06, AS05d, AAK06, FG09, HPS06, JT09a, Lóp07a, MSP08, MSP09, NSS07, Osa08, Rea07, Sch08, Sza05b, VJ09, YA07]. Analytical [HLL08a, LKSC06, LM09a, QW08, RDH07, Tem07, eMH06b, Ahm07, ÁNPQ05, Bat08, FFSS09, HK08, HY09b, LYB09, LB07, PT09b]. Analytical-numerical [LKSC06, LM09a]. and/or [Fin08]. Andersen [GY08, YZ09b]. anergy [CSG05]. Angelesco [BFM09]. angiogenesis [KS07f]. angle [BBB07, Den07a, MS09b]. Anisotropic [AG06, YA07, CYM08, Jai06]. announcement [CSZ09]. Annual [Zhu08d]. annuity [Spr06]. annular [Li09c]. annulus [HC08]. anomalies [LB07]. anomalous [LYB09, LXZG09, Sch06a]. anomaly [yCyZ08, yCyZ09]. antenna [JAE09]. Anti [FWY09, GHZ06, Liu09a, yPyHZ07, Sha09b]. Anti-periodic [FWY09, Liu09a, Sha09b]. anti-reflexive [yPyHZ07]. anti-symmetric [GHZ06]. anticontrol [LY05b]. antigen [ZBB^{+}05]. antigen-activated [ZBB^{+}05].

AOR [Yun08b, tLxLlW07, THC08, WtL09, WS09a, Yun08b]. APC [CSG05]. Apéry [Zud05]. Apéry-like [Zud05]. apparent [MS06b]. Appl [AMO08, Asl10, ByXy06, yCyZ09, Dai11, GPPRB07, KYF05a, LC09b, LB10, LHL11, SZ09b, TGJ19, TSC10, Yun08b, ZJ12, vDvFZ13]. applicability [KM06]. Application [AYK09, AHM08, BMC09, CLQ09, CHS06, CMRS06, DS08, Ehr06, FHM09, GS07a, GJM07, HAS09b, Log08, LW08b, MSB05, QW06, SVC09, SZD09, Sid05, SS08, Svá08a, Abb07b, BLR07, BK09b, Boy09, BSS09, CW08a, Dol08, FHZ09, HO06, HL09c, KAP07, KRAH07, KSE09, Lan08a, LRV^{+}09, LBS09, Mar08, Qi09, SV07a, SAFS07, Tor08, TM08, Val09, XTZ06, YB05, ZLLW08, ZLYW09, ACV05]. applications [Abd08, AZGN06, AABL08, AI06, BRVWM07, ByXy05a, ByXy05b, ByXy06, BFGM06, CMX09, CJCD06, CZ09, Che09, CD07b, DGS06, DS05a, DK06b, Gei08, Gei09, GS08, JAE09, Koh08, KKS^{+}07, KLS^{+}07, Lai09b, Lan06, Las09, LH05b, LHM05, LZ06, MC07, Ma08a, MST05, MJ07, Mon09, NH08, PT09b, PN06, QCK09a, SP09a, SCS08, SLN06, SS09d, TYZ^{+}05, TV09a, TS06b, Tsa09, VMMM09, Wan07a, WPC08, WeR07, Wu08, YZ09d, YL09a, ZL09a, ZWY09, vdBHPS07, ÁNPQ05]. Applied [BBG05, AKW05, BSC07b, CU09b, Fin07, HBS^{+}07, OMC08, PP09, Pul09, Suz05, VN09, WLMS08, WDC^{+}08, ZXZ09]. Applying [TND^{+}09, MK07]. Approach [CU09a, AS05c, AA08, ADLT09, BLR07, BMfX07, BS08b, BBB07, BNDN09, Cak07, CG08a, CS09a, Civ07, DA09, DS06, DVCH06, Den07b, EM06, FL09b, FG07, GGM06, GS06b, HS08a, HHST05, IOO07, JT09a, KL08a, Kho09, KGW09, LZ09d, MLL09, MP06a, MG08, MO07b, MLOP06, NL05, PW08, Pit09, RM07, SL06, Sie07b, SD09, Tok07, VV09b, WLMS08, XGZ09, YYF09, YZ09d, YJY06, YW09a, ZMH09]. approach-application [BLR07]. approaches [ABB07e, BBF^{+}05, JM07a, SEAA08, Wil06, WDC^{+}08]. approximant [GS08]. approximants [DMGVP05, GPTT05, GMFB06, MV05b, Mat07a, Sab08]. Approximate [Bat08, CY08a, GEK09, JKM09, MB07a, MNHA09, SyS08, YG06, And09a, BY09, ED06, Enr06, FC09, HL07b, HYG08a, LXX08, LS06c, MN07b, Smo06, Wil06, WA08]. approximated [Yam08]. Approximating [Lu09a, XW09, ELA09b, HR05a, Rom09, SRB06, SCS08, SH09]. Approximation [AS05c, Ben07, CC07, CGPS08, HV06a, KLLP06, KP08a, KP08b, LNL09, MW09b, PRZW08, Vol08, ZP06, Abb07a, AH07, AAK06, AR09b, And09b, BF05, BGRS05, ByXy05a, ByXy05b, ByXy06, Bla05, BCP07, BT09, CY09, CYZ09, CLN09, CEJ09, Cop05, CD07b, CK08, DH06, DR08, DP08b, EGFOA09, Eng09, Fai07, GJ08, GA09, GHZ06, GSBB09, GNP09, GNS06, GW06b, HS09b, Her09, HMG06, How09b, HV05, KP07, KR08, KSVW07, KKD06, LS07a, LF07b, LLT07, MW07, Mur07, Par05a, Par08a, PR08a, RD08, Roo06, SFSK07, SK09a, SWZ09, SY09, Svá08b, TG09, WT08, Wri08, Xu05a, Xu07b, YWD09, YB05, Yua09, ZW06a, ZLA07, ZLLW08, ZW08, ZYLY09]. Approximations [MYY07, Yüc06, Zud07, ALR05, BLP07, CYB09, Chr09, CCS05, CH09, Dun06, EZ07, Fin07, Ham09, KS07c, LC08a, LY07a, LS05, Mao07, NH08, RVA07a, Spr06, TR09, VCD^{+}08, YG07, YZ09d]. aquifer [Har08b, KY07]. arbitrarily [OTOR07]. arbitrary [EE06, ET06, GH08, KKLM06, Log08, LW06c, LLC08, SF08, SV09, TYZ^{+}05]. arbitrary-order [Log08]. Arc [SLZ06, SFV09]. Arc-length [SLZ06]. ARCH [HL09c]. architecture [Saa06]. area [IT07]. argument [Ext99, Liu08a, LJ09, Mil06, SAD07, SWYZ09, XW07b]. arguments [GX09, MLH08, Sza08, WTL05, WSZ09, WSM^{+}09, Won08, YLN09, ZL07]. arise [Woj06]. arising [CJCD07, Deu07, GZ05, Guo06, LXUK06, NL05, Nou07, TS05, Var07, WF09]. arithmetic [CKC^{+}07, FHL07, MK07, RKAH07, SLMW06, VDL^{+}06]. arithmetic-type [CKC^{+}07]. ARK [Moi05]. ARKN [Fra05]. Arnoldi [LR09, LBS09, VRBF07, Wu07]. array [Oga06, OA06]. arrays [FIR06]. arrival [KL08a]. artery [ICW^{+}09]. Artificial [HZBM06, AEET09, HLY09, Moh08]. Asian [DDV08, FL08a, Hug06, LWH09b, VDL^{+}06]. Askey [CLO09, Koo07, VZ08]. Askey-scheme [CLO09]. aspects [ADF07, ÁNPQ05, BBPR09, BG07b, CDV08, CKWZ07, CLS08, Hui05, PD06, RS05a, Roo06]. Assessing [LWH09a]. assessment [GVC^{+}08, Joh05]. asset [NST08, WWFV07]. assignment [FY06]. assimilation [MDB06, RCR05]. assisted [CBH07, Sch07a, Wat09]. associated [Ass09a, BISS05a, BBR05, Ben06, BFM09, Cho05, Cla05, DMGVP05, GM09a, GQ09b, Kok05, KW07, LZZ09, LF07b, LD06, Skr05, Swa06]. associative [SC07b]. astronomical [ABB^{+}07d]. Asymmetric [DGS06, PRL09, Tay08, CT09b]. Asymptotic [BCGA05b, BDS08, CHBMBR07, Dom09, EPS05, FLS08, IV06, Jia06, KS07c, KTM09, LL01, LS05, Lóp07a, Luo06a, Osa06, Pet05, Sto05b, TV09a, Wan09a, WYO09, YM07, YXJG08, YW09b, AS05a, ADMM07, CdWJ06, Dun06, FL06a, IV08, JM07a, KT09, Lin09a, MM07, Oou08, Par05a, Par08a, RS05c, Sav08, SLH06, SK06, TXZ09, WYL09b, ZZ06b, vD05]. asymptotical [ABÇ06, XZL09]. Asymptotically [Gil07, Tem09, Thi09]. asymptoticity [Ela09a]. Asymptotics [RS05b, AGMMB05, CW05, CW08d, MOZ07, MOC09, MB05, Par09b, Par09a]. Asynchronous [EFS05, MSE08]. atlas [TM08]. atmosphere [BB09a, CdVT07, SM07]. atoms [Sie07a]. attempt [Gau08, IOS07]. attracting [ABP06, ABB^{+}07c]. attractive [LC06a]. attractivity [FL09c, LW05b, LH05a, LF07a, WJS^{+}09]. attractor [ZZL08]. Attractors [LZ05, RW06]. augmented [BG07a, FHM09, GMM09, Sta07, TJ08, WHZ09, ZhL08]. Author [Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05b, Ano05a, Ano05c, Ano05d, Ano05e, Ano05f, Ano06b, Ano06c, Ano06d, Ano06e, Ano06a, Ano06m, Ano06n, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j]. autocorrelation [MPXK06, SJZY09, SJZ09]. autocorrelations [ZSGF09]. automata [Aki08]. autonomous [LC06b, ZCS06]. auxiliary [GH06, ZLYW09, ZA09]. Average [PR05, BH07b, Gu07, Kag09, MLY07, ZZL08]. average-shadowing [Gu07]. averaging [AHL^{+}05, Cao08]. avoid [AMCM05]. axes [rWlC06]. axial [BFT07]. axis [GPTT05, vOP07]. axisymmetric [KS07b, Nke05, Soh05].

B [AM09a, AS05c, Bej06, CMS08, DSI06, GXZ06, KGA08, KRA08, KKD06, LWLL09, Maz08, Müh06a, Ros08, Sab08, WLL06]. B-form [WLL06]. B-spline [AS05c, GXZ06, KGA08, KKD06, LWLL09]. B-splines [Bej06, CMS08, DSI06, KRA08, Maz08, Müh06a, Ros08, Sab08]. backed [CH09]. Bäcklund [WHG07]. backtracking [Zhu05a, Zhu06]. backward [Dao07, HK07b, HR09b, KT07a, RLH09, TGJ19, Wil07, ZGH09]. bacterio [BLR07]. bacterio-phage [BLR07]. Bailey [LV07]. balance [DJ06, DJ07a, Pit09]. ball [HQ06]. Banach [AU09, AABL08, ACQ09, Arg09b, CAY09, DF08, FJG08, FD09, HLH07, Kaz06, KBA09, KA06, yL08, PG07, Par05c, QCK09b, QCKK09, RS05e, Sal09, Thi09, WYL09b, XH08, ZTY07, ZC08a, ZLL08]. band [Fer07, VMV07]. band-limited [Fer07]. banded [NS09a]. barrier [Tar09, WKY^{+}07, YW09a, YSY09]. barycentric [PR07]. based [AAPP07, AA08, AG06, AMV06, ABR07, BP07, BGV^{+}09, BC06, CG08a, CLM06, CGPM08, CK09a, Cas07, CM05b, CHS07, yCyZ08, CP08b, yCyZ09, CGP09, CT09b, CDMR07, DC08, DOY08, EIM^{+}09, FR06, FC09, Gás09, Gem05, GC09, GH08, GDA07, GPS09, GBR09, HPS06, Han08, HTS09, HH07a, HG09, HO07, KKW06, KBA09, Kho09, LT05, LH08a, LW08a, LC08b, LL07b, LYZ09, Ma06, MC08, MDP08, MDT07, MT05, MPP09, MLY07, MNMS05, MS09c, NS09a, NT08, NQDJ06, NaLH07, NMM07, PK09a, Pit09, QLJ09, QZY09, RVA07a, SK05, SJV^{+}09, SS09c, SZ09a, SRW09, Sin09, Śmi07, SLZ06, Sur05, TH07, VVV08, VCD^{+}08, Ver08, WK07a, WYW08, WOK05, WeR07, XGZ09, XC06, YOY07, YLSX06, ZLdST09, ZLLW08, ZW09a, ZSGF09, ZZ09a, ZT06b, vdBHPS07, JKK06, OZL09, PW08]. bases [BBM07, FTT^{+}05, MB06, WC08]. Basic [rWlC06, ABI07, LV06, MJ08a]. basis [BB07, DS09b, DT06b, JAE09, LÅH08, LW08b, OS07a, PLMP08, Pir09, RB05, Sch08, SBO09, STB09, VS08, WZD09, YXYJ06, vOP07]. basket [DDV08, LO08]. batch [CK09b, KL08a, Liu09c, QW08]. batch-arrival [KL08a]. Bayesian [Cal07, Tor05, YYF09]. BDF [MVVA06, RVA07a, VAMVC05]. BDF-type [RVA07a]. beach [Ehr06]. beam [FIR06, lLzS07, WCL08]. beam-induced [FIR06]. beams [AT09d]. bearings [BSKS07]. bed [DP08b]. bed-load [DP08b]. behavior [AA09b, BM08b, CC09, CHL09, EPS05, LMZ08, LJM09, Lin09a, LZZD09, Luo06a, Pet05, PP05, TH07, Wan09a, Yan06c, YM07, YSYX08, YW09b, Zho09b]. behaviors [CLCL07, CHJZ09]. behaviour [BdWG06, CL06a, CP09b, FDFL07, FN09, JM07a, KT09, LLT07]. being [SHC08]. Bell [Dom09, Par09b]. belonging [BBR05]. BEM [Bou07a, Scu09]. Benjamin [OOO^{+}07]. Berger [Tur06]. Bergman [BFGM06]. Berk [AG09]. Bernfeld [ZWF09]. Bernoulli [CDL05, Gau08, RK08b]. Bernstein [Ber05, BB07, BI08b, CEF08, KA08, LW06a, Lin07, LW08b, Lu09b, Sau07, Sza05a, Win06]. Besov [GO07, Kam07]. Besov-type [GO07, Kam07]. Bessel [Sto05b, DMS05, DGS06, DA06, EK07, Har08b, LL01, LD06, Par09c, Xia05]. Best [Xie08, BISS05a, KY06, Qi07, QG08, Qi09]. Bethe [Sie07a]. Between [CK06, AM09a, BCGA05a, COEV05, CFS09, CCW^{+}09, Den07a, DCZ07, FLS08, FT05, HS09a, LW06a, LW08d, Tab07, Waz07a]. Beyond [Jay07, MGV07, CSSL09, JLNP06, Sie07a]. Bézier [CCW^{+}09, HMH08, KI08, KA08, LW08b, Lu09a, Lu09b, WLL06, XW09, HW08]. Bezout [AGRZ06]. BFGS [LNL09, XZ08, XSW09, YLW09a]. BGT2 [RDH07]. bi [CY09, CBDGVN07a, CBDGVN07b, DZLL06, GZLL09]. bi-conjugate [GZLL09]. bi-harmonic [CY09]. bi-orthogonality [CBDGVN07a, CBDGVN07b]. bi-quadratic [DZLL06]. biarcs [MW08]. bias [CV06, Han05]. Bibliography [Ano05-38]. BiCGStab [GM07b]. bidiagonal [TM09]. bidirectional [SC07b]. BIEs [FG09]. Bifurcation [JW08a, Ko07, LSTS06, WZ06, XC08, CCS05, DL09b, FN09, KLO08, LRB05, XGM09, Yan06a, YC06, YHH05].

Bifurcations [ZCH07, LY05b, ZCZ08]. biharmonic [ALR05, Dan06, MDT07]. bilateral [Gon06]. Bilevel [CG08a]. bilinear [KV09b, RS05c, WHG07]. billiards [ET06]. binary [Ahm07, MCD06, Tor05]. Bingham [NT08]. binomial [KG09]. biochemical [SLN08]. Biography [Lor05]. bioinformatics [HKK07, KLS^{+}07]. biological [HSN09, Nag09, RKAH07]. Biorthogonal [CP09c, ARN07, KM05]. biosciences [Car06]. biparametric [HL07a]. Birkhoff [NBSV08, Fin08, WZD09]. birth [Chr05, GLV05, SMT07, YH07]. birthday [VAW07]. bisection [fLxXT08]. Bivariate [CDV08, WL06a, dMFPP09, BRIP09, CG08b, CJ06, DL05, DZLL06, GU08, Liu06, LLC08, WFL07, Win06, ZW06b]. blades [DDL09]. Blasius [AS05d, PT09a]. Blended [BMM06, BM07b, DD08]. blending [ZT06b]. Blind [SJZY09, SJZ09, SC06b, ZSGF09]. Block [KL09, ZT06b, BM08a, Bog05a, BJSS06, DKHNZ08, HR05c, KT07b, Śmi07, TM09, WWM07, Wu07, yZXL09]. Block-based [ZT06b]. block-pulse [BM08a]. block-SOR [DKHNZ08]. block-triangular [Śmi07]. blood [ICW^{+}09, MG08]. blossoms [Maz08]. Blow [Du07b, KV07, Gal09, HO06, MO07a, Rob07]. Blow-up [Du07b, KV07, Gal09, HO06, MO07a, Rob07]. blowflies [LF07a, LD08]. Blowing [CC09, NMKS09]. blowing-up [NMKS09]. blowup [MJX09]. blurring [DESC07]. Board [Ano08m, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. Bochner [VZ08]. bodies [KBSK08]. body [ABF^{+}07, CFS09, TANT09, XZS07]. Boltzmann [DL07a]. boosting [ZZZ08]. Border [SB09]. Born [BCP07]. Bose [CCCH09, Wan07a]. both [SC06a]. bound [AT09c, DFY09, DS05c, LH07b, PIP07, Qi08, QGC08, SDP09, SV07a, XZ08, Zhu05a, Zhu06, Zhu08a]. bound-constrained [AT09c, Zhu05a, Zhu06, Zhu08a]. boundaries [And08, BBSV08, SLN08, XKF06]. Boundary [BN08a, DESC07, Koh07, Sma05, WYS09, eMH08, AABL08, AM09b, AGH06, Att05, BNT08, BV06, CD05, CM05b, CMST06, Cha07, Cha08, CHCW05, CS07d, CDL05, CD07b, CHY05, CG07, Dan06, Dao08, Di 07, DLG05, EE06, EZ07, EGO06, EGFOA09, EJ07, FZ09, FH07, FI07, FJG08, FZ08, Fin07, FOP05, Gao05, Gao06, GC09, Gil07, GJL08, HK07a, HL07b, HY09a, HZBM06, HLX06, He05, Hem06, HCDP08, HVV09b, Hon07, HR09b, HLY09, Hug06, JLNP06, Jan08, JG09, Kar07, Kar09, KMJ06, KV07, Koy09, Lan06, Lan05, LS06b, Li06b, lLzS07, LS09, LZ09a, Li09b, LFHW09, LS07c, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LM09a, LJL06, LT06b, Liu08b, LLW09, Liu09d, Log08, LOA05a, LOA05b, Lu07, xMzY09, MDT07, MD09a, Moh05, MD09b, NM06, NMMS05, NMM07, OYA07, PS07a, PS09a, PK09a, RS08a, Rea07]. boundary [RS07b, Sal09, SFSK07, SA08a, SK09a, SZ08b, SD09, SÁ08b, SO05, SDV07, Sto05a, Sun08, SLZA09, Tat06, TSB09, Tur06, TM08, Wan07c, WWC07, WWW08, WG08, WSZ09, Won09a, WLK06, WKL06, XWZH08, YS06, Yam08, YOA06, YS07, YLN09, Yao09, YL09b, YCJ09, Zha09a, Zha05, ZCL06, ZFG08, ZC08a, ZLL08]. boundary-layer [eMH08]. boundary-only [Gao05]. boundary-value [AABL08, AGH06, FJG08, MDT07, YL09b]. Bounded [MY06, MY08, BCGA05a, DL08a, DL08b, FG08, KI08, Koh07, Luo06b, Roo06, YJJ09]. Boundedness [AHL^{+}05, BGG07]. bounding [MDT09, NSS07]. Bounds [AR06, DDV08, LvGJ06, MS05a, VDL^{+}06, YH09b, ADGH06, ADD^{+}08, BM06, BS07, CL06c, DSW08, DD07, Deu09, ED06, Har06, HYG08b, JM07b, Li08, MS07b, Mor07, MCD06, NH08, Neu08, PK07, Qi07, QG08, Qi09, Shi07, TR09, XDW05, YW08, Zhu08b]. Boussinesq [Gin08, HEOS05, Lai09a, LWLL09, Waz07b, ZT06a]. box [Bej06, CGPS08, FKM05, LMS08b, MRSZ07, BBC07]. box-constrained [FKM05]. box-spline [Bej06]. box-splines [CGPS08]. Bradyrhizobium [MMS^{+}09]. brain [SNV06]. branch [BCeAJ09, Boy09, CS09a, DS05c, SDP09]. branch-and-bound [DS05c]. Branges [KS05a]. Brazilian [MMS^{+}09]. Breakdown [RT09, RT08]. Breakdown-free [RT09, RT08]. breakdowns [WWP09]. Brown [GXgL05]. Brownian [FW09]. Bruno [CHS06]. bubble [LKSC06]. bubble-type [LKSC06]. budgeting [Hua07]. Burger [AS05b]. Burgers [BSC07a, BSC07b, BSC09, CP09b, DA08, DL07a, Gei09, HS08a, KTH08, LLZ09, LYZ09, MNP08, Rad05, Sma05, Waz07b, XHK^{+}08]. Burgers’-type [KTH08]. Burgers/Korteweg [DA08]. Burmeister [CP08b, CGP09]. business [ZL09b]. BVP [HJK09, SS09b, Wei09]. BVPs [AS05e, AKW05, CN07, Has06, Jia09a, Kor06].

C [Yun08b]. Cahn [BN08b]. calcium

[BCeAJ09]. calcium-mediated [BCeAJ09]. calculation [DJ06, EZ06, FPP08a, MVV08, Mot06, SJV^{+}09, Wil06]. Calibration [Pir09]. Can [EFP06]. canal [XFgS06]. candidates [WC07]. Canonical [Ped05, BC06, Wu08, ZTG09]. cantilever [WCL08]. capacitated [KCK09]. capacity [BJ06, KCK09]. capillary [GMFB06, GGP09, KPAT07]. capital [BK09a, Hua07, ZL09b]. Capturing [SRB06, SBK07, WK08]. Carathéodory [BGVHN08, JN09]. cardinal [Bej06, Liu06]. cardinality [FY06]. Carlitz [Dom07]. Carlo [HSN09, Lai09b]. Carrier [OW06b].

Carter [DD07]. case [ADÓ07, BNDN09, CDM07a, CCLS06, CDM07b, DV06a, DESC07, HK08, Kag09, Tup07, Wik08]. cases [Daa09a, DA09, NP07]. Casoratian [WHG07]. Catmull [ZC06]. Cauchy [AS07, DL09a, DS08, ELA09b, HLF09, Kou08, LL08, Mai08b, QFX06, QWS09, RW09]. causal [BBPR09]. cavity [LYT07, RVM08]. CE [QW06]. CE/SE [QW06]. cell [CSG05, CG05, EPP^{+}09]. cell-centered [EPP^{+}09]. cellular [Aki08, LMZ08, Liu07]. censored [LWH09a]. censoring [KG09]. center [DL09b]. center-focus [DL09b]. centered [Ber06a, EPP^{+}09]. centers [CZ09, YJGL07]. central [CD07a]. central-WENO [CD07a]. centro [XZL08]. centro-symmetric [XZL08]. centrohermitian [LF07b]. centrosymmetric [LF08]. cepstrum [SJV^{+}09]. Certain [AUY06, Mil09, VK06, AHL^{+}05, BNP06, CS06a, CS07a, CK07, CS09b, DJ07b, GG07, MN05, MA08b, sR07, YM07, ZWY09]. Certified [Neu08]. Cesáro [AHL^{+}05, CV05a]. CF [Waa05]. CF-coefficients [Waa05]. CFD [PLGS09]. CFSE [DP05b]. CG [ABG06a]. chain [DC09b, MDP08]. challenges [XHW07]. challenging [Lia05]. Chance [Hua07]. Chance-constrained [Hua07]. change [KK09a, KS06]. changes [FN09, MGMTO06]. changing [LLW09, LOA05b, SS09b]. channel [HAS08, LRV^{+}09, LCKK06, MM06]. channels [Khu06]. chaos [CBH07, LY05b, MSA09, SA09, SS09a]. chaotic [NaLH07, Par08b, TH07]. Characteristic [eMH05, eMH06a, HHZ07, KV05, VLR08, YY07a, Zar09]. characteristics [BSKS07, CJ09a, SY09]. characteristics-mixed [CJ09a, SY09]. Characterization [HL08, Ism07, VZ08, AÁN07, BM07a, CLO09, CHS06, HR05b]. characterizations [ÁN06, FPP08b, HCDP08]. Characterizing [BC09b]. Chebyshev [ABGH08, ABBM09, BCGA05b, BG07d, Boy09, KTH08, KK05, KAC05, KW09, LW08b, MDT07, Mat06, Osa08, RVA07a, RVA07b, sR07, SFSK07, Sto08, VRBF07, Xie08, YLS07]. Chebyshev-type [RVA07b, sR07]. Chebyshevian [Maz08]. chemical [Eng09, Tar09]. Chemotactic [SLN06]. chemotaxis [Eps09]. chi [SD06]. chi-squared [SD06]. Chlodowsky [KI08]. choice [AML07, GM06, vOP07]. Choosing [CFGVP07]. Choquet [KM09b]. Christoffel [And08, And09a, BC09b, BD07, CBGV05]. chronic [YY08]. Ciarlet [CY09]. circle [BCGA05a, BR09, BDGV09, CBGV05, CBDGVN07a, CBDGVN07b, Khr05, Mar07b, Pet08, Tem07]. circles [HS09a]. circuit [FHtMP06]. Circular [SACJ09, Ber06b, CMM09, KM09a, Var07]. claim [KG09]. claims [LT06a, ZLY09]. Clark [ZC06]. class [Arg07, Arg09a, AT09c, BT06, BGP09, BF09, CC09, CS09a, CJB05, DL08c, Dio09, EIM^{+}09, FWY09, FJG08, Fra06, GQ09b, HC09, HNS09, HK07b, Hon08, Jan09, Koç08a, LmY08, Liu09a, LGYH09, Liu09b, LW08c, MJ09, MB07a, MJ08a, MVVV06, MLH08, MA08b, MI09, OZY05, PS09a, PJ08, Qi07, QG08, Qi09, SLN08, SCS08, Sha09b, Sid05, Sid08, SK06, TGJ19, TRST09, UHNC09, WWC05, WLL07, WCW09, WF09, WTZ08, XH08, Xie08, XWZH08, XLHZ05, Yao09, Zha05, ZSZ09, ZW09d, dMFPP09]. classes [BBR05, FPP08a, LXSH09, ZLY09]. classical [AÁN07, ÁN06, ABP06, ABB^{+}07c, AGRZ06, Ask05, Car05b, CG06, CSSL09, EKL^{+}07, FPP05a, KMJ06, Kra06, NT06, Pet07b, ZCL09]. classification [eMH05, ADLT09, DC09c, GPS09, IS06, ZZZ08]. Clenshaw [CE09, HS05]. clinical [Ver05]. clock [Nag09]. closed [BGG08]. cluster [MMS^{+}09]. clustering [HKK07, SAFS07].

Clusters [Par07a, NSS07]. CMRH [HS08b]. CMV [Sim07a]. co [BLR07]. co-evolution [BLR07]. Coalescent [TL07]. coalitions [YZ09e]. coated [CCM07]. coatings [CH09]. COCR [SZ07a]. code [Gao05, Gao06, Sak09, SS08]. codes [CGPM08, CM05b, WWW08]. coding [CZ08a, NQDJ06]. coefficient [ASL07, CX07, HV06a, HAS09b, IS06, YYF09, YKS09]. coefficients [Asa05, BM06, BP09, CJCD06, Dem07b, DA06, DI06, ED06, FNW06, Gon09a, HV05, IO06, Jai06, KT09, LMZ08, MP08a, Roe08, SyS08, SZ08a, SZ09b, Woj06, Wri08, XCZ08, YLSX06, YSYX08, ZLYW09, Waa05]. Cohen [SC06a]. coherent [DM05]. cohesive [ADF07]. cohomology [CK07]. coincidence [HL09a]. collage [KLV09]. Collective [HL09a]. collector [Shi07]. collinear [AT09c]. collisionless [Dem07a]. Collocation [CDM07b, Dio09, AM09a, Alh06, ACWL07, AKW05, CD06, CDMR07, DP05a, DL08c, Fin07, KGA08, KRA08, KTH08, KKLY06, LT06b, MDT07, MMM08, MNMS05, NMMS05, PT09a, PT06, RZ07, SZD09, WLK06, WKL06]. collocation-finite [MMM08]. column [Ata07]. Combination [MO08, GNS06]. combinations [BBR05, DJ09, Kho09]. combinatorial [ABÇ06]. combined [Cui07, XGZ09, YXF07]. Combining [TBL09, bJyKyZmT09]. combustion [AR06, ABD^{+}09]. come [DA09]. Comment [Asl10, KGW09, Mil06, Mil07, NK10, TSC10]. commentary [Gau08]. Comments [Jan09, Sto05b, TGJ19, ZJ12, Ter05]. Common [Thi09, CS07e, EEGG08, Jai06, QCK09b]. communal [BNDN09]. communication [EFS05, MGS08]. communication-less [MGS08]. community [SM07]. commutation [BGJ07]. commutative [Kom07a, Kom07b]. Comonotone [FOT07]. Comonotonic [DD07]. Compact [Sut07, ZDC07, CK07, DZ09b, FLS05, GHMPV09, GC06, HRGD08, KCK06, KDN08, LT07, Liu06, MD09b, SS06c, TGB08b, TG07, TM08, WG08, WZ09a]. compacton [TY07a]. companion [BG07d]. comparative [Car07]. comparing [GJM07]. Comparison [BBM07, BB09b, DGR07, Li05c, Rad05, Sav08, WHF07, XTZ06, Eba09, JVVvdZ06, JH08a, NK10, yS05, VB05c, Waz07a]. Comparisons [BDHW07, SD06]. competition [CXS06, Ko07, THS05, Won09a]. competition-diffusion [Won09a]. competitive [LC09a, LOZ08]. complement [Deu09, NL05]. complementarity [BK08, CP08a, CGP09, Cho08, Don09, GKM07, HYYF05, HHC09, JTHZ05, JD05, NB06, Sch07a, WCH05, WB09, ZJW08, ZLL09, Zho09a]. complements [Bac09, LHL06]. complete [GM07c]. completely [Din05a, GQ09b, Qi07, QG08, Qi09]. completeness [Ism05b]. Completions [CK09a]. complex [AMO07, AMO08, BSMT09, CdAHBR06, CFGVP07, FS08, LC08a, LY09, Lui09, MH08, MSB05, MBA09, RS05b, SA07a, SZ07a, TMG09, Wu08]. complex-step [LC08a]. complexity [DFZ07, JF06, SZ09a, YXF07]. complicated [SSW09]. component [BP06, CHS07]. components [GBR09, ZGSF08]. componentwise [XW07a]. Composite [HV05, AMV06, Civ07, LN07b, TV09b, ZWY09]. Composition [GS09, SS06b]. compositions [Dun07]. compound [MS06b]. compressed [WFL07]. compressible [ACV05, Cui07, Cui08, GWG08, Koh07, MS09c, Yan06c]. compression [HV06c, LM08]. Comput [AMO08, Asl10, ByXy06, yCyZ09, Dai11, GPPRB07, KYF05a, LC09b, LB10, LHL11, SZ09b, TGJ19, TSC10, Yun08b, ZJ12, vDvFZ13]. computable [CdAHBR06, TR09]. Computation [ADGH06, AMO07, AMO08, BSS08, CU09a, CDMR07, HS07b, HCL08b, How09a, NMKS09, Röb08, AGM06, BHP07, CHXL06, Daa09b, DB08, DV08a, Dub09, DL07b, ET06, FPO05, Has09a, KW07, LMS08a, Lan05, RK08b, SLN08, SCG08, SV09, Tab07, TM09, UO09, VB05a, VLR08, WK08, YZ09d]. Computational [BBF^{+}05, CDV08, KPAT07, Roo06, Tsa09, ZC06, AT09c, MDP08, PPV05, PD06, RB06, SW06a, TS05]. computations [Deu09, DD08, VV05a, ZFS09]. compute [Ehr06, Gau08]. Computer [GKMS09, CBH07, GW09, IOS07, Sch07a, Wat09]. computer-aided [IOS07]. computer-assisted [CBH07, Sch07a, Wat09]. Computing [CJP08, Cha07, CCW^{+}09, FPP05b, Gau05a, NSS06, PPV05, Tup07, WWW08, Wat06, ZZ07, BSST08, BD07, DRT09, GZLL09, HM09a, KT07b, LSW09, LL06, MVVV06, MN07b, NL05, Scu09, SZ09a, TA09, VT06, ZYLY09]. concave [AMG07, ZTG09]. concentrated [BJ06, Wac05]. concentration [BGGC07]. Concentrators [KBSK08]. concerning [LTC06]. concrete [BM08b]. condensates [CCCH09, Wan07a]. condensation [VS08]. Condition [SZ09a, WD05, CL06a, Dao08, HKA07, Hem06, Koy09, LCH07, LTW07, LS08, LGYH09, LD06, LOA05a, MS09b, NP06, NP07, RDH07, RS07b, SHK^{+}08, SHC08, SZ08b, SD09, Win06, XDW05, XW07a, YS06, YL06, YLS07]. conditional [VCD^{+}08]. conditioned [ABL07, OTOR07]. conditioning [CM05b, CMST06]. Conditions [MO07a, Ant09, AE07, BNT08, BV06, BGG07, COEV05, Cha07, Cha08, CHY05, DESC07, DHS09, EZ07, EGO06, EH06, FT08, FHH08, FJG08, FNW06, FOP05, GK07, GC09, HL07b, HZBM06, Har06, HW07, HR09b, JLNP06, KV07, KLLP06, Li09b, Log08, LP07c, LW08c, MYY07, Rea07, SK09a, SO05, SDV07, WWW08, Won05, Won09a, YOY07, ZFG08, Zho06, ZCL09, Zhu08a]. CONDOR [VB05c]. conducting [IV06]. conduction [CCH06, DA08, HTS09]. Conductor [Maz06]. cone [CP08a, CL09b, GGA08, HYYF05, ZLL09]. cones [ADÓ07, GZX09]. Conference [Ano05n, Ano07o, PRZW08, Ano05o, HH07c]. Confidence [MPP09]. configurations [Mar07b]. confluent [Ell06, YWH05a, YWH05b]. conformal [BDHW07]. Conforming [MK09, CC07, GHT06]. congestion [Kho09]. congruential [Wik08]. conic [HW07, QZZ08, QZY09, Zho09a]. conical [Civ07, Nke05]. conjecture [Bak05, JB09, MK05, ZWF09]. conjugacy [LTW07]. Conjugate [FJT07, And09b, BW06, CXH09, GZLL09, HL07a, HTS09, LLZY06, LTW07, RH06, SG09, SSZ09, TMG09, YLW09b, ZZ08a, Zha09b, ZJ12]. connected [BDHW07, CC09, KS05b, WN08]. Connection [BCGA05a, CC05, GS07c]. Connections [LW06a]. Conservation [AMG07, Ben06, CD07a, DVCH06, JKM09, LZZ09, Mur07, TS08, ZL05]. Conservative [SF08, WCZ07, Mat07b, Mat08]. considering [WC07]. Consistency [KRAH07]. consistent [AB09b, DB08, Kul09, WHG07]. constancy [WYO09]. Constant [ASL07, BS08a, Bra07a, DIKZ07, Gon09a, HKA07, HMG06, HSL08, Kas06, Men09, Pré08b, Roe08, YLN09]. constants [Cof09b, DCZ07, Osa06, Sch06a]. Constrained [Ker09, KA08, SC06b, AT09c, BS09b, Bor07, CR09, FKM05, GBR09, HCL08b, HV06b, Hua07, JTHZ05, bJyKyZmT09, bJxC09, JZW08, Liu05, Lu09b, MZW06, SP09b, TJ08, VRL07, VB05c, WCH05, WZ09b, XZ08, ZW08, ZGSF08, ZXZ09, Zhu05a, Zhu05b, Zhu06, Zhu08a, Zhu08b, Zhu08c, ZA09, BG07e]. constraint [Cim09, GHZ06, Lan08b, PP08, YD08]. constraints [BDL09, CCK06, GGA08, HCDP08, bJQlZ07, KYF05a, KYF05b, Lan08a, MRSZ07, SV07a, TLQ08, WTZ08, YJJ09]. constructed [Chu08]. Constructing [yPlH05, TM08, MJ08b]. Construction [BJ08, CW08a, Sav09, ABP06, Bej06, CMS08, DL08a, Han09, MST05, Nov08, Ros08, VJ09, VW06, ZTY07]. Constructive [LS06c, Sie07b]. consumer [LL07a]. consumer-resource [LL07a]. contact [AB09b, CFHM06, CFV06, CFS09, Cop05, DHS09, FvZ06, GKM07, HV06a, NT08, SL09, Sta07]. containing [AR09c]. containment [Lin07]. context [FGB07]. contexts [CLP08]. contiguous [RI06]. continuation [How09a, Lóp07a]. Continued [Khr05, BHJ05, LW05a, Lut07, VK06]. continuity [HW08, Lu09a]. Continuous [DR08, Zar09, ARG08, ABET08, BJ07a, BS09c, CX08, FR07, Gal06, HQ06, KS09, KM09b, KV07, SP06b, WZ07, Xio07]. Continuous-discontinuous [Zar09]. continuously [SC06a]. continuum [EMMP07]. contracting [BeMBH06, EFS05]. contraction [Bad08, CAHAY09, KLV09]. contractive [ACQ09]. contracts [BBM09, VV09a]. contrasts [FNW06, YA08]. Control [AHM09, SA09, AZGN06, ASL07, BBD06, Bor07, BSV08, BK06, CM07, CY09, CR09, CL08, Chr09, CCK06, FR07, GM06, GL06, HS09a, HWSL05, Hei05, HH07b, HL07c, HGS06, JL07a, KV09b, KS05b, Lan05, LT05, LW05b, LFS08, LFW09, LTC06, LY07b, Liu09c, MZC08, Mar08, MSA09, MLOP06, Par08b, PN06, Pul09, SAHS09, Sma05, Wod05, WTZ08, WWP^{+}08, XXL08, YZ09a]. Controllability [HMPR07, Tat06]. controlled [GK09, KR07]. controllers [XXL08]. controls [GF08, LOZ08]. Convection [SRW09, Alh06, Bog05a, Bog05b, CJ09a, Deu09, DZ09a, Dol08, EKEHR08, GSBB09, Hoo08, IS06, KGA08, Kas06, KHK08, Lin05, LYT07, LS05, MV08, MA09, OS07c, PP09, RK08a, Roi07, SSW09, TG07, YZ09a, YY07a, Zar09, ZT06a]. Convection-diffusion [SRW09]. convection-dominated [CJ09a, YZ09a]. convection-reaction [KHK08]. convective [eMH08, Min05]. conventional [Jay07]. Convergence [BJ06, BJSS06, CJB06, CL06a, CS07e, CHJZ09, CQ09, Cim09, GAC^{+}08, GAR07b, KI08, KKLY06, KD06, LkLkP09, LGYH09, Mai08a, OTOR07, PD06, QCK09a, QCK09b, RHQ06, SST09, SS06a, THC08, VRBF07, Vin09, WLY05, WWM07, WH09b, XCZ08, YL06, YLS07, Yun05a, Yun08a, yZXL09, ZW09c, ZC08b, ABGH08, ALR05, AK09, Arg07, AR09c, Arg09b, BMfX07, Bak05, BJS08, BS07, BR07b, BRW09, BH07b, BHJ05, BS09c, CY08c, CV09, DJ05, DD06, EH06, FH09a, GK09, HR05a, HK07b, Hom05, Hom09, HQ06, Hu05b, HLJ09, JD05, JK07, KYF05a, KYF05b, Kaw07, Koç08a, Koç08b, KLW07a, KLW07b, Kou08, LW05a, LMZ08, LS08, LZ09a, Liu05, LYW08, Lut07, MC08, MZW06, MM09, NS09c, OS07c, OS07b, PS07a, PS09a, PK09b, QCKK09, Sha09a, SX09, Sic08, Sto05a, TGJ19, TD07, Tem09, TFC07, VB05b, VN09]. convergence [WT08, YOY07, Yam08, YN08, YSYX08, ZBL08, ZZ09c, dS07a]. Convergent [PV09b, ZHG^{+}09, BB05, CGP09, Fan09, FHH08, Gau05b, GL07, bJxC09, KGA08, KK08, LP06a, Osa06, Osa07, XSW09]. converse [BI08b]. conversion [WLL06]. Convex [AMG07, ADGH06, ACWL07, BCK07, CL08, DL07b, Fan09, GZ05, KYF05a, KYF05b, KAP07, KK09c, LYW08, MP08b, SCM09, SDP09, XSW09, YXF07, ZW06a].

Convex-concave [AMG07]. convex-convex [SDP09]. Convexity [JT09b, ADD^{+}08, SC06b]. convexity-preserving [ADD^{+}08]. convexized [ZF09]. convolution [AS05c, BS08a, Civ07, CD06, FDFL07, Kok05, MRV09, Swa06]. convolutions [FH07, Lóp07a]. cooling [SS08]. coordinate [LO08]. coordinates [Dmi08, LT07, SBO09]. Copenhagen [PRL09]. copula [HG09]. copula-based [HG09]. copy [LRH07]. core [SM07, YZ09e]. Corner [CO05]. corners [HL07b, HW08]. Cornu [WM09]. corrected [MPP07]. Correction [DS06, Fu08, Han08, Kuz08]. corrections [GGM06]. Corrective [AFT^{+}07]. corrector [Den07b, Din05b, PS05]. correlated [HSN09, WWFV07]. Corresponding [RS05e]. Corrigendum [Dai11, GPPRB07, SZ09b, vDvFZ13]. corrosion [AS07]. cost [CHBMBR07, JRM09]. Cotes [KV09a]. Coulomb [HV06a, LVV05, Sch06a]. countable [BS09c, NS09c, PK09b]. countably [GGG09, GS06c, JG09, LZ08a, LZ08b, LZ09b, LZ09c]. Counter [Bak05, YKS09]. Counter-examples [Bak05]. Coupled [BM08b, BBSV08, AS05b, BV06, BTS07, CYZ09, CX06, CN08, GKM07, HG09, KA07, LRV^{+}09, zLCgS06, MDP08, SM09, Tac07, TRS07, VDV08, Wan07a, WWW08, WNZ09, ZZZ09, ZLDW09]. Coupling [DM07b, BG07c, BGG08, KCLW08, LC08b, RS07b, Sch06a]. coupon [Shi07]. Cournot [GG08, GR09]. Cournot-like [GR09]. covariance [Sen07]. coverage [MPP09]. covering [FH07]. covolume [Bi08a, CJ09a]. Cowell [RVA05]. CP [LVV05]. crack [WDC^{+}08]. cracks [BKT07]. Cramer [DSMY06]. Crank [PP09, WKY^{+}07, YWD09]. Creating [TMOG07]. Credibilistic [Kag09]. credibility [LLS09]. credit [HO09]. crimes [KS07a]. criteria [CS06a, CS07a, CK05, HM09b, LmZbL07, LZG^{+}09, LB09, LM09b, LB10, MS08a, Maz06, PT08b, Sak05, SZ08a, SZ09b, TXZ09, Wan07b, WWC07, WML08, XL07, XX09, YN08, Zhu08d]. criterion [HJR07, MSE08, Sin08]. Critical [Jar09, AR06, MY06, MY08, Xia07a, ZCZ08]. cross [QLJ09, VV05b, WF09, ZW06b]. cross-cut [ZW06b]. cross-diffusion [VV05b]. cross-diffusions [WF09]. cross-entropy [QLJ09]. Crouzeix [BR07b]. crowding [BDHW07]. crystal [BRVWM07]. crystallization [QW08]. Cubature [LM07b, EFP06, GN09, KV09a, Möl07, OS06b, PSV08, SV09, Sto06, Tay08]. cube [PC05]. Cubic [CG08b, GK09, Bej06, CZ09, CFL05, DP05a, GLV05, HS09a, Hom05, Hom09, KRA08, KD06, Swe07b, WFL07, Waz07b, ZBL08, ZCZ08]. cubically [FHH08, Osa06]. cubics [HS08c]. culture [Liu09c]. cumulative [Wil07]. current [FIR06, KKW06, YOO07]. Curtis [CE09, HS05]. curvature [Bra07a, FL06b, ZX06]. curve [And09a, FR06, HMH08, HJR07, JH09, RS08b, RH09, SLZ06]. curved [dCDND06, Her09, Khu06, XKF06]. curves [AKLT09, BC09b, CR09, CCW^{+}09, Han05, Han06, JH08b, KLLP06, KP07, LCZ^{+}08, LW08b, MCD06, PJ06a, SFV09, WZH06, WT08]. curves/surfaces [MCD06]. cuspoid [HCK07]. customer [AA09b]. cut [LXX08, WFW09, ZW06b]. cut-peak [WFW09]. cutting [CG08a, LWYT06]. CVaR [HG09]. cycle [HY09b, ZL09b]. cycles [Cao08, DL09b, GGT07, GG08, Tac07]. cylinder [CMM09, Jon06, KK09c]. cylinders [MOZ07, Nig07]. cylindrical [Civ07, LT07].

D [BCH06, BGP07, BCK07, CGRGVP08, Cim08, Ehr06, GMM09, GC06, GF08, HV06a, Hsu05, KD05, KSE09, LLL06, LT07, Li06a, LN09, LM08, LYT07, LP07c, LC08c, MMM08, MDT09, MK09, NGS05, Scu09, SBO09, TS06a, VVE08a, VVE08b, WHG07, XXL08, ZGV06, JK09]. d-V-type-I [JK09]. d.c [CS09a]. DAE [BMM06, BM07b, CK09a, SW06b]. DAEs [Cam08, Jay07]. damage [CFHM06, CFV06]. Damageable [CFS09]. damped [CBH07]. damping [HM09b, LM09b, YD09b, ZCZ08]. dance [Mer05]. Daniel [Gau08]. Darboux [CBGV05, LZZ09]. Darcy [BH07c, CMX09]. Data [MDB06, SH06, AUY06, ABD^{+}05, AZGN06, BFGP08, BKT07, CR09, CL06b, CP09b, DMM07, DP05b, DD06, EMP08, FL09a, GU08, GNP09, HHZ07, HS08c, KMVWY09, KY07, KP08b, KKS^{+}07, LS05, LOA05b, LP06b, MW08, MMS^{+}09, MR09, RCR05, SV07a, TD06, Val09, Ver05, WWWF09, WLL09, WL09b, WDC07, YA07, ZH08]. data-fitting [CR09]. databases [KLS^{+}07]. datasets [EL06]. Daubechies [CHXL06]. Davey [SJ09]. Davidson [Hoc08]. Dawson [DMGVP05]. DCNNs [Moh07]. DEA [WL09b, Kho09, WC07]. deal [BLR07]. death [Chr05, DP05b, GLV05, YH07]. deblurring [BS09b]. decay [CU09b]. decaying [Gau05a]. Decentralized [HGS06]. decision [Kag09, KGW09, RM07, WY07, WYW08, WLL09]. decision-making [WY07, WYW08]. Decomposition [BR07a, CZ09, Sto08, Tur09, AHC05, AKK06, Bog05a, Bou07a, FL06b, FNW06, FC09, Has06, Hei05, HH07b, Hos06, HLL08a, JDG06, Jan08, KS07b, Li09a, LYZ09, Moh09, PD06, SEAA08, TR07, Ter05, Waz07a, Zha05, Zhu07]. decompositions [KSS07a, Li08, Yi09]. deconvolution [Est07b, RS05a, SC06b]. Dedicated [WS05, VAW07]. Deductibles [DÖ09]. defective [yPlH05]. defects [GXgL05]. deficient [DKHNZ08]. defined [Müh06a]. definite [BW06, DL09c, EKL^{+}07, GT09, Iva06, KMB09b, LWYT06, LHL11, yPESlZ07, RT08, RT09, RH06, TV07, XZS07, XZZ09]. Definitions [Mon09]. deflation [TV07]. deformation [Dmi08, FPS08, LX06, SLZ06, WCL08]. deformations [KBSK08]. deformed [LRV^{+}09, YL06]. Degasperis [ZCH07]. degenerate [Du07b, VN09]. degree [Cao08, EFP06, FH07, FHL07, Han09, HW08, KA08, LW08b, Lu09b, LW06c, LLC08, Sto06, Sza05a, Win06]. degrees [DL09b, OS06b]. Delaunay [DW06b]. Delay [MZC08, PT08b, XLHZ05, AAZ06, ABS07, AK07, BB05, BP05, BJ08, BG09, CS06a, CS07a, CB07, Che05, CLCL07, CHY05, FR07, FL06c, FN09, FL09d, Gan07, GSW07, GW06a, Gug06, HWSL05, HMG06, HGS06, IL05, JZK09, JM07b, Kot08b, LH05a, LH05b, LHM05, LS06b, LF07a, LmZbL07, LZZD09, Liu09b, LB09, LB10, Luo06a, Luo07, Mao07, MY06, MRV09, OZY05, PN06, PP05, RLX08, RW06, Sak06, SFSK07, ST07, TBBC07, TXZ09, VLR08, WLS08, WLS09, WZ06, WW07, WJS^{+}09, Xio07, XZL09, XL07, Xu08, XGM09, XWT09, YG06, ZS05, ZGH09, ZW09c]. Delay-dependent [MZC08, PT08b, XLHZ05, CB07, TXZ09, XZL09]. delay-differential [JZK09, Xio07]. delay-integro-differential [Gan07]. delayed [JW08a, LCH09, LGYH09, LY05b, MLL09, Sin08, SMT07, TYH08, XXL08, Yan06a, YC06, YH07]. delaying [CKT09]. delays [BG09, Bru09, CXS06, FWY09, FG07, FL09c, Har05, JZK09, Jar09, KTM09, Li05b, LZY09, LY09, LOZ08, LC06b, LWMT07, Luo08, LRB05, MZC08, Moh07, Moh08, SC06a, SC07b, TCMS06, TR06, TXZ09, WLL07, WML08, WLF09, WYO09, XLHZ05, XW06, YC05, YZL08, YY07c, ZG09, ZZZ09, ZW09b, ZL09b, Wan09a]. delivery [BNDN09]. delta [Xu07b]. demagnetization [WZH06]. demands [EM09, Siv09]. dendritic [BCeAJ09]. denoising [BP06, BP07, FL06b, LM07a, SRW09]. denominator [DZLL06, Nov08, PR07]. dense [HS08b, NKA09]. densely [Gau05a]. densification [BC06]. density [ASL07, FPP08a, FPP08b, HKB^{+}05, LP09, MJB09, MP06a, NK08, PV09b, Soh05, SMT07, YH07]. density-dependent [MP06a]. dependence [Hat09]. dependences [SC07a]. dependent [BG09, CB07, Cha07, Cha08, CEJ09, FL06a, GSW07, GRW08, Har05, HV06a, IT07, JW08b, JL07b, KGA08, KKW06, Kim05, KC07, LkLkP09, LCH09, LC08c, Ma06, MZC08, MP06a, NPTH09, Nig07, PT08b, PRGPS09, PP09, RHQ06, SMT07, SBO09, TXZ09, TSC10, WPC08, Wan09a, WYO09, XZL09, XLHZ05, XGM09, YC05, YH07, YKS09, ZL05]. depending [Luo06a, SST09, WL07, Xu08]. deposit [TMOG07]. depth [BCK07, ZC06]. Derivation [AT09c, BJ07a, FT08]. Derivative [Zho08, BTI08, BT09, CLA09, CGP09, CXH09, DEMR08, Fin08, HHC09, IO06, Jia09a, LC08a, LLT07, RP09, RP05, SK09a, WeR07, YL06, YLS07, ZJ12, dB05a]. derivative-free [BT09, CGP09, CXH09, DEMR08, ZJ12]. derivative-nodes [dB05a]. derivatives [Abb07a, Daa09b, DJ05, Eig07, HS09b, HSL08, KC07, MW09b, Mat07b, Mat08, NWW08, Röb08, SS09d, SH09, TV09a, WT08]. derived [AMV06, Chi05, Gil07, Las08, SJ09]. Deriving [OW06a]. derogatory [SS07]. Descartes [Eig07]. Descent [KA06, BC06, CP08a, CLQ09, DC09c, DS05b, HLF09, JTHZ05, YLW09b, ZZ08a]. describe [KCS06]. description [KKS^{+}07, LRV^{+}09]. descriptor [HNS09]. descriptors [AA09b]. Design [Sha07, CJCD06, Duf09, GF08, JAE09, KCK09, KGB^{+}07, Kuz08, PLGS09, YJY06, YA08, ZAVS09]. designed [Kaw07]. detect [FL06c]. Detecting [GR09]. detection [AM09b, CB07, yCyZ08, yCyZ09, DMM07, FIR06]. Determinants [Ism05c, DS05a]. Determination [Dao08, DH09, JL07b, TQA06, TANT09, Di 07, Ish07, SAFS07]. determined [GGG09]. Determining [XF07]. deterministic [EMP08, VC06, WO08]. developing [Kro06a]. development [SL06, SLN06, Wod05]. developments [BGRS05, Cak07, Moi05]. deviating [GX09, Liu08a, LJ09, MLH08, SWYZ09, Sza08, WTL05, WSZ09, WSM^{+}09, Won08, XW07b, ZL07]. devices [ACV05]. devising [Gug06]. DFO [VB05c]. di- [Zud07]. diagonalizable [BJS08, SHC08]. Diagonally [Smo06, THC08]. Diagonally-striped [Smo06]. diagram [NSK07]. diagrams [Cof05, KLO08, LSTS06]. dielectric [CCM07, DCZ07]. diffeomorphism [TM08]. difference [AS05e, And09b, AKL05, ABS07, AGRZ05, Bej06, BUG07, BJ06, COEV05, Cen08, CD07a, Daa09b, DH09, DZ09a, DZ09b, DO06, Ela09a, GZL07, GC06, GS06b, Har08a, HLJ09, HZ05, IS09, JL06a, JL06b, KD05, KSE09, KM09c, Li05a, LH05b, LCH07, lLzS07, LZG^{+}09, LC06b, Liu08b, LP06a, LYZ09, MC07, Mag09, MMM08, MY08, MI09, PS07a, PP05, Qin09, RAG07, Roe08, SF08, SCY09, Sut07, TGB08a, TGB08b, Vin09, WG06, WCZ07, WG08, WNZ09, WZ09a, WG09, Won05, WL07, XTZ06, YZ09c, YM07, ZDC07]. differences [ARVP09, FV08, Gal06, HQ06, Müh06a, Sau08, WSST05]. Different [Lai09a, BB09b, DJ09, HZ05, LmLgF09, Müh06b, NaLH07, WWP^{+}08]. differentiability [EH06, Hu05b]. differentiable [ABBM09, AR09c, Jay08]. Differential [DI06, WeR07, AETE07a, Abb07b, AT09a, ADMM07, AAZ06, AMCM05, ADLT09, ABS07, AK07, Ass09a, ADDdM05, BS08a, BB05, BP05, BJ08, BH07a, BG09, BS07, BB07, Bru09, BLP07, BW07b, CS06a, CS07a, Cao08, CJB05, CJB06, Car06, Car05b, CHCW05, CLCL07, CS07d, CL09a, CW05, DG08, DR08, Den07b, DFFW06, DF09, DAE05, DG05, EGO06, EGFO09, Enr06, EM09, EM08c, EK07, EMMP07, EKL^{+}07, FWY09, FWZ09, FPP07, FGRW09, FFSS09, FL06c, FL09b, FN09, FC09, FL08b, FL09d, GG09a, GGD08, Gan07, Gar09, Gei08, GS07b, GL06, GW06a, Gug06, GX09, GS06b, HZBM06, Har05, Has06, HO06, HMG06, Hos06, HLL08a, HR09b, HM09b, IO06, Ish07, IL05, JL06a, JZK09, JDG06, Jan09, JN09, Kaw07, Kom07a, Kom07b, Kom08, Kot08b, KTM09, KLV09, yL08, Lar08, LW05b]. differential [Li05b, LC06a, LLZY06, LS06b, LS07c, LDM09, Liu09a, LZZD09, Liu09b, LZ09d, LB09, LLW09, LM09b, Liu09d, LB10, Lop07b, Luo06a, Luo06b, Luo07, Luo08, MYY05, MYY07, Mao07, MN05, MD09a, MY06, ML09, MA08b, MO07b, MO08, NMKS09, NMMS05, OZY05, PDM08, PT06, PT08a, RB08, RLX08, RLH09, RS05c, RS07b, SZ08a, SZ09b, SFSK07, SDV06, SLZA09, Sza08, TRS07, TRST09, TBBC07, Toc05, TV09b, VLR08, Vil09, Vin09, Vol08, WWC05, Wan07b, WWC07, WML08, WLS08, WPC08, Wan08a, WSZ09, WLS09, Wan09a, WYO09, WYL09b, WS09b, WH09b, Xio07, XL07, Xu08, YS07, YZL08, YG09a, YLN09, YG06, Yüc06, YCJ09, ZS05, ZFG08, ZGH09, ZLL08, ZW09c, Zho09b, Zhu08d, jia09b]. differential-algebraic [GL06, Hos06, LZZD09]. differential-difference [GS06b]. differential-functional [WH09b]. differentiation [ACR06, BXC09, HY09b, Li05a, NWW08, Röb08, Wil07, XC09, ZMH09]. diffraction [BLW06]. diffuse [TL09]. diffusion [eMH06b, ARG08, Alh06, ARVP09, AKL05, ABR07, ABV08, Bog05a, Bog05b, BH07b, BW07a, CVB07, Che05, CWW08, DS08, Deu09, DZ09a, Dol08, Du07b, FL06a, GDM08, GSBB09, GAR07b, GM09b, GC06, GL07, HKV09, HAS09b, HRGD08, HK07b, IS06, KGA08, Ko07, Kan08, KMVWY09, Kot08a, Kro06b, KHK08, KBB05, LL05, LL08, LSTS06, Lew06, LZY09, LWS09, Lin05, LS05, LXZG09, MJX09, MPS05, MP07, Mai08b, MV08, MSZ07, MBCV09, MA09, Mus09, OS07c, PP09, QWW08, RVAD08, Roi07, RKK07, SK05, SSV06, SRW09, Slo08, SM09, TGJ19, TG07, VG09, VV05b, VJ09, WZ09a, WF09, Wil06, Won09a, YZ09a, YY07a, YKS09, Zar09, ZLA07, Zha08, ZC08b]. diffusion-convection [IS06]. diffusion-reaction [ABV08, CVB07, HKV09, SSV06, VJ09]. diffusions [PW08, Pok06, WF09]. diffusive [AT09b, Sea07, SLN06, VVE08a]. diffusivity [CK08]. digamma [Cof05]. digital [AAPP07, AA08, TND^{+}09]. dilute [HS07b]. dimension [CP09c, LW06c, WO08, YG09b]. dimensional [BLR07, Ban05, BS08b, BS07, BFGM06, Bra07b, BG07f, Cof05, CK08, DZ09b, DL07a, EKEHR08, FG09, Gao06, HK07a, HZBM06, HTS09, Har08a, HV06c, Jan09, JVVvdZ06, JL06b, JW08b, KGA08, KCK06, KM06, LÅH08, zLCgS06, LCZ09, LN07b, LOA05a, LLT07, MDT07, MP06a, MG08, MS07c, MD09b, Oga06, OA06, Par08b, PLMP08, Rad05, RB08, RVM08, RCR05, Roe08, SAHS09, ST05, SS06c, SSMB07, TTZ06, TS08, TG09, TRST09, THS05, TG07, TQA06, TANT09, TD06, VRBF07, WGZ08, WLK06, WDC^{+}08, XZS07, XHK^{+}08, XX09, YYF09, ZWF09]. Dimensions [DCF06, Far06, GZL07, WS08, ZKSS06]. dipoles [YOO07]. Dirac [Pie09, Sch06a, Xu07b]. Dirac-like [Pie09]. Direct [BM08a, BI08b, CM07, LDM09, SFPS09, Duf09, FG07, Gás09, MNG09, Wac05, YOO07, YD09b]. direction [BISS05a, Bej06, CGPS08, Qin09, Yu06, YLW09b]. Directional [AB05, Aki09, LBS09, LM08]. directions [GAC^{+}08]. Dirichlet [BNT08, CJCD06, DT06b, GNP09, HL07b, Har06, Jai06, JLNP06, SFFS08, SFPS09, Tur06, Won09a]. disability [Spr06]. disc [CU09a, CU09b, PD06, SB09, Tor08, Wün05]. discontinuities [Wri08]. discontinuity [SL09]. Discontinuous [BHP07, CP09b, Eps09, AMG07, AB07, Ber06a, Cui07, Cui08, DFLP08, ER09, GSS07, Kan08, KLLP06, KP07, LY07a, LN07b, LS05, LC08c, MCR08, Mar09, TRS07, TR09, VJ09, WG06, Zar09, ZLZ09b]. discounted [LXSH09, ZLY09]. discrepancies [PVC06]. Discrepancy [Tab07, PC05, Sin09, VC06, WS08]. Discrete [Civ07, Coj09, EZ07, HKV09, HCDP08, JZK09, Kar07, KK06, PT08a, PT09b, TYH08, AS09b, AÁN07, BR07a, BRIP09, BMS05, BS05, Cal07, CZ08a, CLN09, Che09, CSSL09, DSW08, HJR07, IO06, IV07, KCK06, KM09b, KP07, Kro06b, LOZ08, LXX08, LC06b, LM08, LZZD09, MZC08, Ma08a, Mat07b, Mat08, MJ07, MSA09, Moh08, MS07d, MRSZ07, Nou07, Pie09, RS05b, RAG07, SG05, SV07b, SP06b, ST07, SZ07b, VDL^{+}06, WLL07, WZ08, XC08, XXL08, YJY06, YY07b, YHH05, ZF09]. Discrete-analytical [PT09b]. discrete-complex [RS05b]. Discrete-time [TYH08, MZC08, Moh08, SG05, WLL07, XC08, YJY06, YHH05]. Discretisation [FR07, AK07]. Discretization [CCK06, KS07a, Mai08b, AT09b, Bor07, BK06, Che08, CHJL08, Gao05, MS07c, NP08, SZ08b, VVE08a]. discretization-theory [Gao05]. discretizations [CJB06, Deu07, Han07, Kan08, PDM08, PV09b, Qiu07]. discretized [BP05, CMR06]. Discriminating [WC07]. discrimination [AI06]. disk [EFP06, PSV08]. disks [FT05]. dispersal [CT05, TCMS06]. dispersion [Bat08, CdVT07, MM06, Roo06, SZ07b, TS05]. dispersive [AEET09, FS07, Li06a]. Displacement [YH05, YWH05a, YWH05b, Cui07, Cui08, SY09]. dissipation [AEET09, TS05]. Dissipative [Mat08, FS07, LZ05]. Dissipative/conservative [Mat08]. Dissipativity [Gan07]. distance [CCW^{+}09, DL07b, EE06, Eig07, Iza09]. distributed [AAZ06, ABD^{+}09, Che05, Chr09, DF09, GZLL09, GX09, Hei05, MP07, Moh07, Moh08, Sma05, SC06a, SC07b, UFG09, WTL05, WS08, XW07b, ZLLW08]. distributed-order [DF09]. Distribution [KSE09, AA09a, CLQ09, DA09, DD08, KM09b, LWH09a, LWH09b, LL06, MYY05, MGMTO06, MS06b, NK08, RVAD08, Spr06, VT06, YJGL07, YA08]. Distributional [BR09]. distributions [BN08a, BBR05, CdWJ06, GA09, GK08, KM09b, LP07b, Sen07]. disturbances [Dor09]. divergence [MPP07]. divergences [MPZ09]. divided [DH09, Gal06, HQ06, JL06a, Müh06a]. dividend [GY08, LWS09, LXSH09, YZ09b, YW09a, YSY09]. dividends [FW09, GS06a]. division [DP05b]. Dixon [BE06b]. Does [CMM09]. Domain [FL06b, NSL07, AFT^{+}07, BP06, BC08, Bog05a, Bou07a, BFP09, BDGV09, CH09, Dul08, FHtMP06, FNW06, FG07, Gao05, GZL07, Hei05, HH07b, JW08b, KA08, KW09, LLL06, Lu09b, Lui09, SVC09, SST09, SHK^{+}08, SLN06, YH09b, Zhu07]. domains [AG05, BN08a, BBSV08, BFGM06, CGRGVP08, DL08a, EZ06, FG08, GS06c, HL07b, HZBM06, Her09, KS07b, Koh07, KKS^{+}07, Li09c, Lui09, MSTT08, MM07, Nke05, Roo06, SSW09, SFPS09, SDV07, Var07, YG07, ZKSS06, vOP07]. dominant [MVV08, MA09, THC08]. dominated [CJ09a, MV08, YZ09a]. Doppler [RCR05]. Double [He05, NMMS05, CHM06, He06a, LY05a, Li07a, MRV09, MN07a, MNMS05, Qi07, QG08, Qi08, QGC08, SK06, WLF09]. double-gate [LY05a]. doubling [CHLW08]. Doubly [CMRS06, HR09a, HR09b, RLH09]. DP [CW08a, DHS09]. DQ [Yüc06, LYT07]. Draghicescu [Sak09]. Drazin [WD05]. Drazin-inverse [WD05]. drift [RS05c]. drift-implicit [RS05c]. driven [HR09b, NTS06, RVM08, RLH09, ZZL08]. driving [MD08]. drums [Ban07]. drying [GKMS09]. DSC [Civ07]. DTM [MO08]. DtN [CNN07, Koy07, Koy09]. Dual [BNPR08, How09b, NaLH07, Ber06a, BFP09, Dul08, GXZ06, WFL07, XF07, ZTG09]. Dual-mixed [How09b]. Duality [Ask05, AHS08, Ant09, ACV05, FL08a, GGA08, LTY08, SDP09]. duct [Hoo08]. due [dCDND06, MD08]. Duffing [GC09]. Dunkl [Cho05, Kam07, TX07]. duopoly [GG08]. during [BM08b, VT06]. dust [LmLgF09]. dusty [LmLgF09]. Dynamic [CLCL07, HKB^{+}05, LCZ^{+}08, VV05a, AYK09, ABB^{+}07d, BGV^{+}09, BC06, CHXL06, DaC05, EPS05, EBP09, He05, Hon07, Jac06, LZ05, LXUK06, Mar08, MD08, MSA09, Sak05, Sak06, SL08, XX09, YW09b, YD09a, ZL09b, ZF09]. Dynamical [LZZD09, YTM07, GF08, Hat09, Koh08, KV07, LJM09, LY09, Liu09c, SM07]. Dynamics [BSC07a, BSC09, HZ05, ABD^{+}05, Aue07, BLR07, Ban05, BMC09, BB08b, BB09b, CMRS06, DBE07, Dmi08, FR07, GR09, IOO07, IV07, KS09, LH05a, LH05b, Li09b, Meh05, RW06, SFSK07, SZ07b, Tac07, THS05, VDV08, WF09, YY08]. Dyson [DIKZ07].

each [Ter05]. early [BBM09, Bin07, Mer05, ZBB^{+}05]. Earth [IV08]. easy [CR09]. easy-to-control [CR09]. ecological [WW07]. economy [LCA07]. ECT [Müh06a]. eddy [KKW06]. Edge [LM07a, GG09b, YZ09a]. Edge-preserving [LM07a, GG09b]. edges [ELA09b]. Editorial [GMS08, ZFS09, Ano08m, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Ano09w, Ano09x, Ano09y]. Editors [Ano07u, Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano07-39, Ano07-40, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano09z, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27]. Editors [Ano05-28, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, GD09, VK08]. Educational [Art09]. EEG [CHV^{+}08]. effect [dCDND06, HKA07, Min05, MU08]. Effective [LCH07, CK08, WC08]. Effectiveness [KM06, SS08]. effector [Wod05]. effects [BCP07, FZ08, IV08, LC09a, NPTH09, SBO09, TSC10]. efficiencies [WY07, WYW08]. Efficiency [Kan05, CdAHBR06, CHBMBR07, EHR09, GS09, Pet09b]. Efficient [AKW05, Cas05, EL06, FGB07, FPP08a, GLS07, GPPRRB06, GPPRB07, Guo06, PR08b, SK09b, TR09, VLR08, Xia07a, BBD06, CDM07a, CC06a, CCJ07, Cof09a, DP08a, EGM07, GW09, HC09, HO07, JKM07, KDN08, LMR07, LCKK06, PRP08, SK06, Tan06, VAN06b, WK08, WT07, WC07, WZ07, ZZZ08, Zhu05c]. Ehrenfest [AR09a]. eigenbasis [Ter05]. Eigenfrequencies [Ban07]. eigenfunctions [Gri09, NSS06]. EIGENIND [WWW08]. eigenpairs [yPlH05]. eigenproblem [Yua09]. eigenproblems [Wu07]. eigenspaces [CL06c]. Eigenvalue [CW05, DV08a, Deu07, Deu09, LRV^{+}09, NP06, NP07, ABG06a, AR09b, ALR05, CKWZ07, CCJ07, CLN09, CHLW08, Gei09, Ham09, Hem06, Hoc08, IV08, Ish07, Jar09, Koh08, Li07b, LF07b, LvGJ06, PKS^{+}09, SS07, VRBF07, WMG08, WL09a, WWZY09, YD09a, Zho06]. Eigenvalues [ÅMK06, BE07, Won08, Wu08, AGM06, AV06, BHP07, BSS08, DW06a, HS06, HLL08b, JM07a, Lan06, LHL09, LHL06, QWW08, SHC08, SO05, Toy07, Vol08, WWW08, YXJG08, YH09b, Yüc06, dF07]. eigenvector [EEGG08]. eigenvectors [MVVV06]. eight [LLD09]. eight-node [LLD09]. eighth [BRW09, SS05, Tsi06]. eighth-order [BRW09]. Einstein [CCCH09, Wan07a]. Ekeland [ZL09a]. elastic [ABF^{+}07, CFS09, CCLS06, Dem07b, LCZ09, LP07c, TANT09]. elastically [KS05b]. elasticity [BNPR08, GMM09, Her09, KS07b, SdJMH09, Sta07, XZS07]. elastico [HeM05]. elasto [AYK09, LLM08]. elasto-plastic [AYK09]. elasto-plasticity [LLM08]. elastodynamic [AB09b, BFP09]. elastodynamics [YA07]. electric [CHSW08]. electrically [IV06, NTS06]. electricity [BSMT09]. electrocardiogram [ZSGF09]. electromagnetic [CD05, CCM07, Cak07, EGM07, HCDP08, YOO07]. electromagnetics [CH09, FTT^{+}05, OMC08, AFT^{+}06]. electromagnetism [ABL07, DM07b]. electrophoresis [GMFB06]. Electrostatic [MMFMG07, DRV08, YZ09d]. element [AG06, AB07, ALR05, ABP08, BNT08, BN08b, BG07a, BGP07, BR07b, BNPR08, BFP09, CY09, CYZ09, CYM08, CL08, CWW08, CYB09, Chr09, CCS05, Cop05, DG08, Deu07, dCDND06, Dul08, DBG07, DBR08, EGFO09, EJ07, Fai07, FL08a, GGD08, GMM09, GLOS05, GDA07, GHT06, Ham09, HH07a, HCDP08, Her09, HS06, How09b, HGS08, JW08b, KKW06, KK06, KS07d, KK09c, Koy07, Koy09, LY07a, LLM08, Li06a, LW06b, Li07a, LH08a, xLLhZ08, LLD09, LWLL09, LHL09, LvGJ06, Ma06, MSZ07, MP06a, NH08, NTS06, NS09b, OS07b, OS09, OMV09, PV09b, RTGB09, RKK07, RKK09, SST09, SY09, Svá08a, Svá08b, Tab07, TR08, TSB09, TY05, Won09a, XKF06, XZL08, Xio07, YG07, YZ09d, YG09b, YXJG08, ZP06, Zar09, ZC08b, ZKSS06, vOP07]. element-finite [MP06a]. element-Galerkin [EGFO09]. element/ALE [SST09]. elementary [BXC09, DK06b]. elements [BdAR08, CC07, Dmi08, GLOS05, GRW08, Gil07, HMPR07, HLL08b, Jen07, Kuz08, MS09b, MK09, QCK09b, Sur05, Zha07]. elevated [MM06, Win06]. elevation [Sza05a]. elimination [Saa06]. ellipsoidal [HLY09]. elliptic [AS07, AK09, BNT08, BF05, BR07b, Bi08a, Bor07, BK06, CJCD07, Car05a, DAE05, DBR08, EPP^{+}09, FFSS09, GAC^{+}08, Gan08, Gás09, Gon06, GNP09, Har06, HAS09b, HC08, Jai06, KK06, KK08, KS07c, KSS07a, KKLY06, KV07, Kor06, KV09b, LFHW09, Lui09, Mag09, NH08, Neu08, RS05d, SFFS08, SFPS09, Tem07, TR07, TR08, TR09, TV09b, VG09, Wan07c, WG08, WLK06, YG07, Yan06d, Zhu08d]. elliptical [Nig07]. Elman [CM08]. embedded [FSW09]. embedding [Lui09]. embeddings [GO07]. Emden [LM09a, XL07]. emergent [GPS09]. Empirical [GH06, TFC07]. Enclosing [JT09a]. encoded [DL07b]. encounters [CSG05]. end [LNW06]. endwall [SS08].

Energy [GSS07, AB09b, BFGP08, EZ06, HV06b, KSS07b, LVV05, MDP08, ZGV06]. energy-based [MDP08]. energy-consistent [AB09b]. engagement [CG05]. engine [ABD^{+}09, DDL09]. engineering [GPS09, RTGB09, SLN08]. enhanced [CNN07]. enhancing [Eds09]. ensure [COEV05]. entire [DC09a, Yan06d]. entries [Ism05c]. entropies [MPZ09]. entropy [AMG07, Aki08, Aki09, HCL08b, KL08a, QLJ09]. entry [Mil07, RVK05, RV05]. Enumeration [Cam08, GZ05]. envelopment [Val09, WLL09, WL09b]. environment [CK08, Iza09, PLGS09, YJGL07]. environmental [AMO07, AMO08, CM08, KLS^{+}07, PT09b]. environments [THS05]. enzyme [JSS07]. epidemic [CLGG09, SMT07, WZ06, WJS^{+}09, YH07]. equalities [TS06b]. equality [JZW08, Lan08b, Lan08a, Zhu05b, Zhu08c]. equation [Abb07a, Abb07b, AEBEK09, AAK06, ABL07, ACV05, Asa05, Asl10, Ass09b, BM08a, Ban08, BFT07, BSC07a, BSC07b, BSC09, BY09, BD06, Boy09, BF08b, BR09, Bra07b, Bru08, CRS05, CY09, CLCL07, CHL09, CMRS06, Cla05, CBH07, CP09b, DSI06, Dan06, DP05a, Dao07, Dao08, DSW08, DS09b, DZ09b, DFLV06, Djo07, DJ06, DFY09, DL07a, DL09c, Eba09, EM07, EZ07, ESAD05, Eng09, EPS05, EK07, FSL09, FZ09, FI07, Fer09, FGRW09, FT05, Gal09, Gil07, Gin08, GS07c, Gon07, Gon09a, Guo06, HK07a, HCS05, Har08a, HS08a, HWY09, HBS^{+}07, HLL08a, HYG08a, IO06, IL05, Jen07, JW08b, KCK06, KRA08, KKLY06, Koç08b, KHWF09, KV09b, Kru05, LT07, LWW08, Lai09a, LL05, LRV^{+}09, LT05, LC06a, LLZY06, LS09, LZG^{+}09, LDM09, LM09a, LWLL09, LTM08, Liu08a, LLZ09, LYB09, LD06, hLyHZ08, LOA05a, Lu05, LLT07]. equation [LJ09, LYZ09, Lut07, LXZG09, MDMRP09, Mai08b, MV08, MNP08, MT05, MCA09, MY06, ML09, MRV09, MN07a, MU08, MD09b, Mus09, Nke05, NMMS05, OS07c, OOO^{+}07, OS06a, PT09a, PI06, yPESlZ07, yPyHZ07, PP08, PV09b, QFX06, QWS09, Rad05, RZ07, RW09, RO05, SS05, SWZ05, SyS08, SWYZ09, SJ09, SRW09, uIHA09, SD09, Sma05, SWZ09, SBO09, SCY09, SS06c, Sut07, SSMB07, Swe07b, TY07a, Tur06, VN09, VJ09, Vin09, Vol05, WHG07, WSM^{+}09, WYO09, WG09, WO08, XHK^{+}08, Xio07, XCZ08, YZ09a, Yan06d, YZ09c, YM07, YCJ09, ZCH07, ZLA07, ZLYW09, ZM09a, ZT07, ZL07, ZZ08b, ZL09b]. equations [AETE07a, AEET09, AT09a, eMH05, ARG08, AS05b, ADMM07, AAZ06, AFT^{+}07, AMCM05, ARVP09, ABG06b, AB05, ADLT09, AKL05, ABS07, AK07, ABR07, AGRZ05, AR09c, Arg09a, ABV08, ADDdM05, BS08a, BS09a, BB05, BP05, BBPR09, BN08b, Bar09, BJ08, BH07a, Bat08, Bej06, BG09, BUG07, BLT09, BGP07, BB07, BRW09, BBSV08, BK09b, BF09, BTS07, Bru09, BLP07, BW07b, BSS09, BH07c, COEV05, CS06a, CS07a, CVB07, CM07, CC06a, CC08, CCD09, CJB05, CJB06, CMR06, CMV09, Car05b, CHCW05, CX06, CS07d, CX07, CL09a, CT09c, CYB09, Che08, CXH09, CHJL08, Chr09, CG06, CD06, CMT09, Daa09b, DC08, DSC09, Dai11, DL08b, DL09a, DG08, DR08, DVCH06, Dem07b, Den07b, DFFW06, DF09, DZ09a, DL08c, Dio09, DA08, DO06, DAE05, DFLP08]. equations [DG05, EGO06, EGFO09, Ela09a, Enr06, EBP09, EM08c, FWY09, FWZ09, FH09a, FPP07, FPO05, FLPA09, FFSS09, FL06c, FDFL07, FL09b, FN09, FC09, FZL07, FTT^{+}05, FH09c, GG09a, GGD08, Gan07, GS07a, GJM07, GZL07, GJ08, Gar09, GKM07, Gei08, Gei09, GC09, GRW08, GS07b, GO05, GL06, GL07, GW06a, GSS07, GS08, Gug06, GX09, GS06b, GRL09, HCL08a, HEOS05, HZBM06, Han07, Har05, Has06, He05, HRGD08, HO06, HMG06, Hon07, Hos06, HL09b, Hsi09, Hsu05, HSL08, HR09a, HR09b, HM09b, HMT09, HV06c, Ish07, Iva06, IS06, JL06a, JZK09, Jac06, JDG06, Jan09, JN09, bJyKyZmT09, KKW06, KYF05a, KYF05b, Kaw07, KCS06, KTH08, KA07, Koç08a, KV07, Kom07a, Kom07b, Kom08, KW09, KLO08, Kot08a, Kot08b, Kro06b, KTM09, KLV09, LLL06, LWW08, Lai09b, Lan06, yL08, Lar08, LY07a]. equations [Lew06, LW05b, Li05b, LH05b, Li06a, LS06b, Li06b, lLzS07, LH08b, LH08a, Li09a, LkLkP09, LS07c, Lin05, zLCgS06, LXUK06, pLT07, LY07b, Liu08b, Liu09a, LCZ09, LZ09d, LB09, LLW09, Liu09d, LB10, LRT06, LN07b, Lop07b, LC08c, Luo06a, Luo06b, Luo07, Luo08, Ma06, MW07, MJ09, MJX09, Mag09, MO07a, MB07a, MS09a, MYY05, MYY07, Mao07, MSB05, Mar09, Mat07b, Mat08, MN05, MD09a, MY08, MA08b, MRV09, MS07c, MNHA09, MO07b, MO08, Mot06, MNMS05, MI09, NMKS09, Ned08, Neu08, Nou07, OW06a, OZY05, PDM08, PT06, PT08a, PV09a, PPR07, Pet09a, PP05, Pie09, QW06, Qin09, Qiu07, RA07, RLX08, RLH09, Rob07, RAG07, RS05c, RHQ06, Roo06, SK05, Sak05, Sak06, SST09, SKAW09, SAD07, Sha09a, SZ08a, SZ09b, SFSK07, Sim07b, SRP08, SZ08b, Śmi07, SA07b, ST07, SDV06, SM09, SH09, SF06]. equations [SL08, SLZA09, Sza08, TRS07, TTZ06, TS06a, TRST09, Tat06, TBBC07, Toc05, TV09b, UO09, Uje07, VVE08a, Var07, VLR08, Vil09, Vol08, WWC05, WTL05, WCH05, WCZ07, Wan07b, WWC07, Wan07a, WML08, WLS08, WPC08, WHL08, Wan08a, WZ09a, WSZ09, WH09a, WLS09, Wan09a, Waz07b, WYL09b, WS09b, Wil06, Won05, WWD09, XW07b, XL07, Xu08, YG07, YC05, YS07, YZL08, YWD09, YG09a, YLN09, YW09b, YG06, YLW09a, YXYJ06, ZGV06, ZS05, ZZZ06, ZZ06a, ZFG08, ZGH09, ZJ12, ZW09c, ZLDW09, Zhu06, Zhu08a, ZH09, Zhu07, Zhu08d, ZCZ08, jia09b]. equidistant [Huy09, Mao07]. Equidistributed [SÁ08b]. equilibria [PPV05]. Equilibrium [AKLT09, CY08b, CAHAY09, HyLC06, KA06, LCA07, QCK09b, ZW09a]. equivalence [CLP08]. equivalent [BGRS05]. equivariant [ÅMK06]. ergodicity [Gu07]. Erratum [AMO08, ByXy06, yCyZ09, KYF05a, LC09b, LB10, LHL11, Van07a]. Error [ADD^{+}08, ABR07, Bou07b, Che08, CNN07, Cop05, CDM07b, DSC09, Dai11, DP08a, DZZT07, EJ07, Koy07, Koy09, Li06a, Li07a, Li07b, MS09b, MSP09, VMMM09, Wan07c, YLC09, AT09a, AG06, BRIP09, BG07a, BFP09, CGPM08, CS07c, CL08, CCS05, CN08, DFY09, ED06, EPP^{+}09, FH09b, Fin08, FT05, GPV05, Gon07, GSS07, GNP09, KS07d, Kor06, KV09b, KHK08, LH07b, LY07b, MS05a, MS07b, MvdMV06, MCD06, NH08, Neu08, Osa06, PIP07, Pre09, Sen07, SÁ08b, SF06, TR09, TQA06, YG07, YZ09a, Zhu08a]. errors [BGJ07, CU09b, GPTT05, SB09, VRBF07]. estimate [Che08, FI07, GNP09, HO07, KHK08, SZ09a, TQA06]. estimated [GH06]. estimates [Bou07b, CL08, CCS05, CHBMBR07, DSC09, Dai11, Deu07, FT05, GPTT05, Gon07, GSS07, Koy07, Koy09, KV09b, LY07b, MS09b, MSP09, Moh05, PD06, Vin09]. Estimating [DP05b, MCD06, Sen07, FR06, HO06, MvdMV06, Wat07]. Estimation [ER07, TMG09, AT09a, BBF^{+}05, BTI08, BFP09, CGPM08, CS07c, DMM07, DA09, DZZT07, GG09b, HAS09b, Hu05a, HR05c, KS07d, Kor06, LCKK06, LP09, MV08, MJB09, NGS05, PPR07, SB09, SF06, Ver05, WK07a, YG07, YLC09]. estimations [CN08]. estimators [ABR07, CP09a, CV06, KK09a, MPP09]. Euclidean [Bra07a, KSS07b, Sin09]. Euler [TGJ19, AK07, BB05, Ber06b, CM07, FN09, Gau08, HK07b, JN09, MYY07, Mao07, PDM08, Pré08b, YL06, ZGH09]. Euler-type [BB05]. Eulerian [LN07a]. Evaluating [ABF^{+}07, FH07, SW06b, WYW08, EJ07, HC09, Kho09, MS06a]. Evaluation [Gao05, Vol05, AG09, Asa05, BMS05, BS05, BE06a, CE09, Cof05, DP08a, EJ08, GLS07, Gao06, GM07c, HSN09, HCK07, LVV05, MS08b, PK06, Par07b, Par07a, Par09c, RB05, SLMW06, SK06, Val09, Van05]. evaluations [MNG09]. even [CDL05, GX09, HCK07, KV05, WTL05, YL09b]. even-order [YL09b]. evenoids [HCK07]. events [ZBB^{+}05]. Eventual [ZS05]. everywhere [CV09]. evolution [BLR07, BK09b, EM09, FFSS09, KCS06, Mat08, MA08b, Pie09, SA07b, SACJ09, UO09]. evolutionarily [Lin08]. evolutionary [BBF07, BKB06, KCLW08, Lin08, LW08a, PLGS09, TBL09]. Ewens [Hui07]. Exact [CdWJ06, Eba09, KCS06, Roe08, Tat06, ZGV06, eMH06a, Asl10, Ass09b, CHL09, Gin08, GPS08, Koi05, LWW08, LLZ09, hLyHZ08, Vil09]. exactness [BC08, OS06b]. examples [Bak05]. Exceptional [Daa09a, Sch06a]. excitable [Sha05]. exclusion [Alo07]. exemplified [Wil06]. exercise [BBM09, Gau05b]. exhaustive [Tan06]. exist [CMM09]. Existence [AV06, CHCW05, CXS06, HCS05, HJK09, HLX06, HWY09, LW05b, LF07a, LS07b, LD08, LZ09c, LJL06, LOA05a, LJ09, MK07, NPTH09, RR06, Sun08, TSC10, YC05, Yan06d, Yas07, ZCS06, ZG09, ZC08a, ZL07, ZSZ09, FHL07, JG09, Jia09a, LZ08a, LZ08b, LZ09b, LZW09, MN05, MN07a, NM06, RLX08, SFV09, ZLL08, ZT07]. exit [PW08]. Exotic [Sch06b]. Exp [Asl10, BK09b, Eba09, Ass09b]. Exp-function [Asl10, BK09b, Eba09, Ass09b]. expanded [TMG09]. expanding [BeMBH06, Xu07a]. expansion [Abd08, Boy09, GHZ06, LL01, Mat06, RV05, Sto05b, Tur06]. expansions [AS05a, CV05a, CDL05, DA06, HLL08b, KSS07a, LHL09, Lóp07a, PK06, Par07b, Par07a, TV09a, YXJG08, ZZ06b]. expectations [VCD^{+}08]. expected [CP07, LXSH09]. Experimental [BDL09, VB05c, BGP07, YA08]. Experiments [ARN07, BG07d]. expert [CKC^{+}07]. Explicit [CDL05, Djo07, GG09a, GGT07, JM07b, MVVA06, VV09a, AMCM05, BJ07a, Fra05, Fra06, Gar09, HLJ09, Li05a, LW07, LLZ09, Ma06, WCL08, WG09, WSPJ09, ZP06, ZWZ08, Zhu07]. explicit-implicit [Zhu07]. explicit-magnetic-field [Ma06]. explosion [GEK09]. exponent [ZW09d]. Exponential [BB05, CU09b, Koz08, Koz09, LMZ08, Mao07, MVVA06, Moh08, RTGB09, Zha08, CVB07, DA09, DJ05, FL09c, FL09d, JS08, KSVW07, KT09, LWH09a, LZY09, Luo07, MBCV09, May08, MOC09, Moh07, MNMS05, Nov08, NMMS05, PDM08, Par09a, Rie06, Rom09, Sab08, SÁ08b, SC07b, TYZ^{+}05, VV09b, VAMVC05, XLHZ05]. Exponentially [VV07, CFMR08, CFMR09, Gau05a, HVV09b, Van05, Van06a, Van07a]. Exponentially-fitted [VV07, HVV09b]. exponentials [CJP08]. expression [EKL^{+}07]. expressions [HY09b]. extend [Iza09]. extended [DP05a, HEOS05, Jan08, KKD06, MJ08a, NKA09, SWZ05, WF08]. Extending [SD06]. extension [BM07b, DM05, KK05, Rza09, Sak09, Sid08, SSZ09, SS09e, VB05c]. Extensions [AGRZ06, BE06a, CSSL09, LC08a, Möl07]. exterior [CU09b, CNN07, HLY09, Koy07, Koy09, Qui07, SHK^{+}08, WGZ08]. external [ZZL08]. Extinction [LC09a]. extracting [JK07]. extraction [SK06, ZSGF09]. extragradient [CY08c]. extrapolated [FHZ09, FZL07]. extrapolating [WDC07]. Extrapolation [Mar09, CLA09, CLN09, pLT07]. extrapolations [YXJG08]. extrema [SZ09a]. extremal [Ism07, yPESlZ07]. extreme [FNW06, Kra06, MPZ09, vDvFZ09, vDvFZ13].

Faà [CHS06]. Faber [OS07a]. facilitates [Koç08b]. facility [TFC07]. factor [CP07, HO09, LCYZ08]. Factoring [XC06]. Factorization [GO05, BDY09, DO06, RT09]. Factorizations [KT07a, GSBB09, Yun05a]. factors [And09a, GS07c]. fair [NP08]. fairing [ZX06]. Falkner [Asa05]. families [HHR06]. family [ABBM09, BS09c, CXH09, HHC09, Iva06, JK07, MW08, MW09c, Moh05, Ned08, NS09c, Osa07, PM05, PRP08, PK09b, QCKK09, RS07b, SG09, Tem09, YL06, YLS07, ZJ12]. Farey [IL05]. Farey-type [IL05]. Farlie [GK08]. Fast [BK06, CD06, Cou09, EZ06, LLL06, PLGS09, WDC07, Bru08, CC06a, CMR06, CHLW08, DM07b, DRT09, DFFW06, Est07b, LS09, Mai08b, MVV08, OMC08, PT08c, RTGB09, RO07, Sak09, TGB08a, YH05, YWH05a, YWH05b]. faster [NKM08]. faults [IV08]. Favard [CBGV05, Van08]. FC [HL09a]. FC-spaces [HL09a]. FDM [zLCgS06]. FDTD [FZL07, XTZ06]. FE [BF05, CNN07, HKV09, WDC^{+}08]. feasibility [CS06b, MP08b]. Feasible [bJQlZ07, JTHZ05, bJxC09, ZLZ09a, Zhu08b]. features [BBM09]. February [Gau08]. fed [Liu09c]. fed-batch [Liu09c]. feedback [JL07a, JW08a, LT05, LW05b, lLzS07, LOZ08, MLOP06, SS09a, XXL08]. Feir [OOO^{+}07]. Fejér [SL05]. Feller [ZLA07]. FEM [BCH06, Bou07a, LN09, MK09, MSTT08, MU08, Roo06, SL06, SV07b, UHNC09, VS08, WT07, XCZ08]. FEM/FVM [MSTT08]. fetal [ZSGF09]. FETI [DHS09]. Feynman [Cof05]. FFT [HO07]. Fibonacci [Kil07]. field [AAR09, CCLS06, DCZ07, dCDND06, HPS06, ICW^{+}09, IV06, LvGJ06, Ma06, MD08, NSK07, TTZ06]. fields [AD09, CHSW08, HCDP08, YOO07]. fifth [Cla05, KLW07a, SWZ05]. fifth-order [KLW07a, SWZ05]. filled [Dem07b, LZLH07, LXX08, lSsZ05, WYL09a, YY07b, ZXZ09]. film [BSKS07]. films [VV05a]. Filon [Xia07b]. filter [SP09b, Zho09a]. Filtering [WLL07, JKK06, Pok06]. final [CL06b]. finance [DGS06, HCL08b, VK08]. financial [RTGB09]. find [Tok07]. finders [PRP08]. Finding [GW09, WC08, Alo07, BG07d, Osa06, Osa07, PM05, lSsZ05]. Finetti [YW09a]. Fingering [WZS06]. fingerprint [ZLdST09]. Finite [CGRGVP08, CYZ09, CSZ09, Dmi08, Her09, Ma06, NTS06, NS09b, AG06, AT09b, ARVP09, AB07, AS05e, ALR05, And09b, AKL05, ABS07, ABP08, BNT08, BN08b, BG07a, BUG07, BGP07, BR07b, BG07b, BJ06, BSV08, BNPR08, BFP09, BG07c, BCK07, BdAR08, BE06a, BE06b, CY09, CYM08, CL08, CWW08, CYB09, CC07, Chr09, CCS05, Cop05, CD07a, CK08, Deu07, dCDND06, DZ09b, Dor09, DK05, Dul08, EGFO09, EPP^{+}09, EMP08, Fai07, FPS08, FL08a, GZL07, GMM09, GLOS05, GRW08, GDA07, GC06, GHT06, Ham09, Har08a, HH07a, HK08, HCDP08, How09b, HLJ09, HZ05, HLL08b, JW08b, KD05, KKW06, KK06, KM09a, KS07d, KAC05, KK09c, KSE09, Koy07, Koy09, KM09c, Kuz08, KL09, LLL06, LY07a, LY05a, Li06a, LW06b, Li07a, LCH07, lLzS07, LH08a, xLLhZ08, LLD09, LWLL09, Liu08b, LvGJ06, LP09]. finite [LP06a, LYZ09, MMM08, MSZ07, MS09b, MP06a, Mon09, MS09c, NH08, OS07b, OS09, OMV09, PS07a, PV09b, RTGB09, Roe08, RKK09, SST09, SM07, Siv09, SF08, SY09, Sur05, Sut07, Svá08a, Svá08b, Tab07, TS08, TGB08a, TGB08b, Tem09, TR08, TK07, TY05, UHNC09, VVE08a, VVE08b, WG08, WZ09a, Won09a, WSST05, XKF06, XTZ06, XZL08, Xio07, YG07, YLC09, YZ09d, YG09b, YXJG08, Zar09, Zha07, ZWY09, ZDC07, ZC08b, ZKSS06]. finite-difference [GZL07, Roe08]. finite-differences [ARVP09]. Finite-element [Ma06]. finite-level [BSV08]. finite-part [Mon09, ZWY09]. finite-time [LP09]. finite-volume [SM07]. fire [BM08b, RMKA05]. First [PW08, ADDdM05, BM08a, COEV05, CY09, CMST06, CS07d, CLA09, Che08, Fin08, Gin05, Jia09a, KS07c, KC07, Kom08, LC08a, LS07c, pLT07, Liu09d, Qi07, ST05, SZ08a, SZ09b, WSZ09, WS09b, YLN09, YLS07, ZLL08]. first- [Kom08]. first-order [ADDdM05, Che08, SZ08a, SZ09b, WS09b, YLN09, ZLL08]. first-passage [ST05]. Fischer [CP08b, CGP09]. Fisher [DP05a, DOY08, MDMRP09, OS06a, SRD05]. fishways [AVMVMV08]. fit [BdWG06]. fitted [CFMR08, CFMR09, DWW06, FSW09, HVV09b, PS05, Van05, Van06a, Van07a, Van07b, VV07]. fitting [CR09, CLQ09, JSS07, MVVA06, SV07a, SACJ09, TYZ^{+}05, VV09b, VAMVC05, ZH08]. Five [Sim07a, CMT09]. fixed [CY08b, CAHAY09, CS07e, EST07a, HL09a, HHZ07, Hug06, QCK09b, SVC09, SJZ09, Thi09, Won05, ZTY07, ZGSF08]. fixed-fixed [HHZ07]. fixed-point [SJZ09, ZGSF08]. fixed-sign [Won05]. flame [OW06b]. flame-front [OW06b]. flames [RR06]. flat [FT05, HeM05]. flexibility [AG09]. flexible [Dmi08, EFS05, SAHS09]. Flow [KGB^{+}07, eMH08, BGP07, BTS07, CO05, CEJ09, CMM09, Deu09, DDLM07, DFHS08, ER09, Fai07, FT05, Göt05, GGP09, GS06c, HAS08, HS07b, HeM05, ICW^{+}09, IV06, KPAT07, Kas06, Khu06, Las08, Lin09b, MP06a, MG08, MM06, NTS06, NSK07, Oga06, OA06, RVM08, Svá08b, TSB09, TLQ08, TY07b, WLMS08, Zad09, ZX06, ZZL08]. flows [eMH06a, AB07, AMO07, AMO08, BG07b, BG07c, BGG08, BeMBH06, BdAR08, CK09a, DRT09, Deu07, FZ08, GLOS05, HKA07, How09a, MS09c, OS09, SF08, Soh05, TK07, VVE08b, WT07, Wat09]. fluid [eMH05, BB08b, BB09b, CLA09, CEJ09, Dem07b, GW08, Guo06, Hat09, HKA07, HAS08, How09a, HKM08, IV06, OS09, Svá08b, VDV08, Wat09, Zad09]. fluid-filled [Dem07b]. fluid-solid [HKM08]. fluid-structure [VDV08]. fluids [Yan06c]. flush [SS08]. flush-slot-cooling [SS08]. flux [AMG07, AT09b, Hsu05, Kas06, Kuz08, LNW06, OS09, XF07]. flux-free [OS09]. fluxes [AMG07, ZL05]. FM [Ano05-29, Ano05-30, Ano05-31, Ano05-32, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x]. focal [Won08]. focus [DL09b]. Fokas [PK09a]. Fokker [WO08]. fold [XY05]. following [Sta07]. force [TANT09]. Forced [WWC05, eMH08, CBH07, GC09, Hoo08, HM09b, Li05b, LH08b]. forces [ZZL08]. Forchheimer [Las08]. forcibly [Hat09]. forcing [AML07, ZSZ09]. forecasting [BSMT09, XGZ09]. Foreword [BE06c, GD09, VK08]. form [Fai07, Has09a, JL06a, KA08, LW05a, Lu09b, Sau07, SLMW06, WLL06, WF08, Wu08]. formal [GS08]. formally [LT07]. formation [LNW06]. forms [Lin09a, Moh09]. formula [AOSY05, BXC09, Ber06b, BE06b, CHS06, CCLS06, CN07, CBGV05, ELA09b, HYC05, HS05, KMJ06, MO08, Oou08, Won09b, Xie08, XC09, Yan05, ZC06, vD05]. formulae [DMS05, GN09, Hui07, HV05, JB09, MM07, OS06b, SS09d, Sto06]. Formulas [MP08a, BC08, BP09, BDGV05, BDGV09, CX08, CBDGVN07a, CBDGVN07b, Gon09b, HHST05, Li05a, LM07b, MJ08b, Mil09, Möl07, NSL07, Pet07b, Tay08, Tur09, Wil07, YH05, YWH05a, YWH05b]. Formulation [AB09b, BSV08, BGG08, DM07b, FT05, GDA07, Hug06, LB07, MD08, YXYJ06]. formulations [AFT^{+}07, DRV08, DCZ07, Dul08]. Fornberg [BDHW07]. Fornberg-like [BDHW07]. Fortet [Bru08]. Fortran [Gao05, Gao06]. forward [Dao07]. forward-backward [Dao07]. foundation [CFS09]. Four [RCR05, AOSY05, BBM07, CGPS08, CMT09, DWW06, xMzY09, NBSV08, Par08b, ZCL06]. Four-dimensional [RCR05, Par08b]. four-direction [CGPS08]. four-point [xMzY09, ZCL06]. four-stage [NBSV08]. four-step [DWW06, NBSV08]. Fourier [AGM09, BP09, CZ08a, CLA09, CV09, DD06, DFY09, KK09c, LÅH08, LLT07, LC08c, PT08c, VMMM09, Wri08]. Fourier-finite [KK09c]. Fourth [FOP05, HRGD08, QFX06, WG08, ABG06b, DZ09b, FWZ09, Gal09, Has06, Kar09, Kou08, KBB05, LT07, LJL06, Lu05, LJ09, xMzY09, NMMS05, NMM07, PS09a, PS05, RVA07a, TG07, Wan07c, ZCL06, ZFG08]. Fourth-order [FOP05, HRGD08, QFX06, WG08, ABG06b, DZ09b, FWZ09, Gal09, Has06, Kar09, Kou08, LT07, LJL06, Lu05, LJ09, xMzY09, NMMS05, NMM07, PS09a, RVA07a, TG07, Wan07c, ZCL06, ZFG08]. Fowler [LM09a, XL07]. Fox [MPS05, MP07]. FPM [TK07]. fractal [Ban07, MDT09, SNV06, TH07]. fraction [KSS07a, VK06]. Fractional [Pie09, Abb07a, ARG08, AR09a, AHM08, BGG07, CLA09, CEJ09, Den07b, DFFW06, DA08, EM08c, FGRW09, FC09, GDM08, GG09a, Gar09, GAR07b, GM09b, HS09b, HLL08a, JDG06, Jay08, LW08c, LXZG09, MPS05, MP07, MPR07, MNHA09, MO07b, MO08, PJ06b, PJ08, Roo06, Sal09, SW08, SWZ09, SH09, TV09a]. fractions [BHJ05, Khr05, LW05a, Lut07, Waa05]. fracture [KBSK08]. fractured [LP07c, MSTT08]. fragmentations [Hui05]. frames [CS08]. framework [FPS08, Gon06, LMR07]. Fréchet [ABBM09, Hu05b]. Fréchet-differentiability [Hu05b]. Fredholm [BS09a, BF09, CMR06, CMV09, DL08b, Hsi09, Ma08a, PT08a, RA07]. Fredholm-type [Ma08a]. free [eMH06a, eMH08, AFT^{+}06, BG07c, BGG08, BT09, CGP09, CXH09, Civ07, DEMR08, ESAD05, FZ08, GWG08, GGP09, How09b, HHC09, IV06, Kas06, Koh08, MLL09, MD08, OS09, RT08, RT09, RP09, WCL08, XSP09, Yam08, YB05, ZJ12, ZLZ09a]. free-convection [Kas06]. freeze [GKMS09]. freeze-drying [GKMS09]. frequencies [Gil07]. Frequency [KC07, Van05, ACWL07, AE07, FG07, IV06, Kim05, LN07a, LCKK06, MOZ07, RDH07, SC07a]. frequency-amplitude [SC07a]. Frequency-dependent [KC07]. Frequent [YZ09c]. Freud [MB07b]. friction [HV06a, NT08]. frictional [AB09b, CFHM06, CFV06]. friendly [RMKA05]. Frobenius [BG07d, MV05b, Mat07a]. frog [BD06]. front [OW06b]. fully [Bat08, DDL09, FvZ06]. Function [GS08, eMH05, eMH06a, AMG07, AHL^{+}05, AD09, ADD^{+}08, Asl10, Ass09b, BP05, BK09b, CP08b, CGP09, CCS05, Cof08, DH09, Eba09, EJ08, ELA09b, Ext99, FLS06, FS08, Fin08, FPP05b, Gau05b, GQ09b, GBR09, HKZ09, Hsi09, HS05, JAE09, JT09a, JSS07, KK09a, KKLM06, KSVW07, KD06, LL01, LZLH07, LXX08, Luo06a, LHW08, MC08, MNG09, MPXK06, MCA09, Mil06, MvdMV06, MLY07, PT09a, Par05b, PLMP08, Rui05, Sab08, lSsZ05, SLH06, Sto05b, TJ08, TMG09, Tok07, Tur09, WFW09, WYL09a, XHK^{+}08, Xu07b, Xu08, XSP09, YZ09d, YY07b, ZW09a, ZXZ09, ZZ09a, ZTG09, ZA09]. Function-valued [GS08]. functional [Bru09, CX06, FL09b, FL08b, GO05, GXZ06, Har05, HR09a, JL06a, Lan08b, Lan08a, LW05b, LC06a, LCH09, LS07c, LTC06, LXUK06, Liu08b, Liu09a, LZ09d, Liu09d, MS09a, RLX08, SAD07, Vol05, WWC07, WYL09b, WS09b, WTZ08, WH09b]. functionally [CC07]. functionals [BR09, KW07, Yam08]. functions [AS05a, ADGH06, AHS08, ABET08, Ant08, Ant09, ADGR06, ADÓ07, BM08a, BS08a, BXC09, BNP06, BMW08, Bel06, Ber05, BGJ07, BHJ05, BGVHN05, BGVHN08, Car05a, CZ07, Cof09a, CJ09b, CGP07, CJ08, CDL05, CD07b, DS09b, DFZ07, DC09a, DT06b, Dun06, EIM^{+}09, FKL05, FPP08a, FPP08b, GHMPV09, GG07, Gau05a, GK07, GPTT05, GQ09b, GM07c, Har08b, HC09, HS09b, HCK07, HV05, Ism07, JK07, Jon06, JT09b, KM09a, KI08, KC07, KS05a, LL01, LÅH08, LVV05, Li06b, Liu07, LYW08, LXSH09, LW08d, MPS05, MP07, MDT09, MDSR04, MJ08a, MM07, MC05b, MC05c, MSP08, MSP09, MS07d, NWW08, NSS07, Nou07, Osa08, Par09c, Ped05, Qi07, QG08, Qi09, RA07, RB05, Sau08, Sch08, Sch05, SW08, Skr05, Sto05b, Swa06, TV09a, TFC07, Tur09, VB05a, VB05b]. functions [Var07, VS08, Vid05, WZ07, WGZ08, YXYJ06, ZYD07, ZW06a, ZZ07, ZLY09, dS07a, vOP07, HHC09]. Fundamental [BTS07, CU09a, CU09b, JK07, Li09c, Oga06, OA06]. Further [Sid08, Sza05b]. Fuzzy [BK09a, WL09b, CT09a, CP09a, Dai09, DA09, EM09, FY06, Hua07, Hua08, Iza09, Kaz06, KK09b, KGW09, LLS09, LZWQ09, MM09, QLJ09, RM07, SVC09, SP09a, SRP08, SAFS07, Ver08, WW09, WCW09, XGZ09, YJGL07, YZ09e]. FVE [YY07a]. FVM [MSTT08].

G [AA09b, ÉT08, TYH08, WWP09]. GA [GBR09]. Galerkin [AB07, Ber06a, BHP07, CN08, Cui07, Cui08, DSI06, Deu07, DFLP08, DT06b, EGFO09, ER07, ER09, Eps09, FG09, Gon07, Gon09a, GSS07, GNP09, GRL09, Han07, HL09b, Kan08, KS09, LZ09a, LN07b, LC08c, MCR08, Mar09, Mat08, NMM07, PT08a, Roi07, Scu09, TR09, UO09, VMMM09, Xio07, YZ09a, ZLZ09b, ZH09]. Galerkin/Runge [Han07]. GALI [BMC09]. game [Lin08, LW08a]. games [YZ09e]. Gamma [SLH06, AS05a, AA09a, GQ09b, GM07c, Har08b, MvdMV06]. Gammel [Bak05]. GAOR [ZSWL09]. gap [CSZ09, DIKZ07, Den07a, ZW09a]. gaps [AGM06, Coj09]. gas [Ban05, CMRS06, Las08, Ste07]. gases [SL09]. gate [LY05a]. Gauss [GPPRB07, LC09b, BC09b, CLM06, CFMR09, CEF08, EJ08, FLS06, GPPRRB06, JB09, Kim05, Li05d, LH08a, MS05a, MC05a, MS07b, MSP08, NKM08, Par05a, Par08a, SV09, Vid05, Waa05, ZM09b]. Gauss-fractions [Waa05]. Gaussian [AAR09, BC08, CV05b, FL06b, Gon09b, HSN09, MCA09, MC05b, MSP09, Saa06, ZLLW08, ZX06]. Gaussianity [SJZY09]. Gautschi [JB09, Qi09]. GCDs [WA08]. Gegenbauer [HS05]. Gelfand [MN07a]. gene [HBS^{+}07]. General [CH09, Li05a, Pre09, ZYD07, Zin05, Bno07b, CQ09, DL09b, JTHZ05, JK07, KKS^{+}07, LY09, LC06b, LL07a, LLZ09, Sin09, WCH05, WLS09, WWP09, XH08, ZLDW09, ZL05]. generalised [Par09b]. generalization [HVV09a, HMH08, PJ06b, ZWF09]. Generalizations [BCV^{+}05, OW06a, QGC08]. Generalized [BNT08, BGP09, Dai09, HKZ09, HYC05, LCA07, MPR07, Par05c, PS09b, Skr05, Tor08, VRSS08, VZ08, Wün05, Xu05a, Abd08, ABG06a, AMG07, AHL^{+}05, AU09, ARVP09, ABP08, ByXy05a, ByXy05b, ByXy06, BUG07, Bra07a, BI08b, CY08a, CX06, CP08b, CLA09, CGP09, CG06, CV05a, Din05a, Din05b, DF08, Eba09, EGO06, FD09, GW08, Gin08, GM07c, GK08, Han08, Har08b, HHC09, HR09b, HyLC06, HWL09, Jay08, JK09, bJQlZ07, JC09, JB09, Kaz06, KBA09, Kil07, LWW08, Lai09a, LÅH08, LD08, LH09, LM09a, Lin07, LF07b, LF08, LXZG09, MDMRP09, MT05, Mil09, MO08, Müh06a, Müh06b, PT09a, RP05, SAD07, SW08, SCG08, SFFS08, SFPS09, Śmi07, Sza08, THC08, WGZ08, WL09a, WMN05, WN08, Won08, XZL08, XC09, YL09a, YN08, YA08, YD09a, ZCH07, ZhL08, ZJW08, ZLYW09, ZW09a, yZXL09, ZSGF09]. generalized [ZSZ09, Zhu06]. Generals [RMKA05]. generate [EM07, EM08a]. generated [PS08, VT06]. Generating [PPG07, CZ07, GG07, KKLM06, UFG09, WFL07]. Generation [BMS05, BS05, RB06, BB08b, DW06b, HSN09, JLR06, Swe07a, Van06a, Van07a, Vul07, Zud05]. generative [Gia06]. generator [BS06, Wik08]. generators [Kan05, Tan06]. Generic [Pet07b, KMJ06]. Genetic [McC05, TBBC07, ZZ09b]. genotype [WWWF09]. genus [AKLT09]. Geo [TYH08]. geodesic [Car07]. geodynamo [Che08]. geographical [AMO07, AMO08]. Geometric [CKWZ07, CMS08, Har06, Arp07, Gar07a, MPXK06, SEAA08, SaLJ08, rWlC06]. geometrical [FH07, RD08]. geometries [SV09]. geometry [Koç08b, ZGV06]. geophysical [BB08b, BB09b]. geospatial [KKS^{+}07]. Gerber [ZLY09]. Geronimus [MH08]. Gersgorin [CK06]. Gibbs [BMW08]. Gilbert [Ban08]. Ginzburg [Lei08a, Lei08b, LLT07]. Given [Vil09, Wri08]. Givens [YH09a]. glass [Lan05]. Global [CGPM08, FL06a, FL09c, LH05a, LZY09, Liu05, Liu09b, Moh07, SW08, SC07b, Sto05a, WJS^{+}09, WF09, YH07, YY07c, YY08, ZW06a, ZZ09c, Alo07, ABGH08, BJS08, KMB09a, LW05b, LF07a, LZLH07, LXX08, Liu07, LYW08, MZW06, MK05, RW06, lSsZ05, SaLJ08, Sin08, TF07, Tok07, WZ07, WFW09, WYL09a, WOK05, YY07b, ZW08, ZXZ09, ZA09]. Globally [LC06a, Fan09, XSW09]. GMRES [BJS08, PKS^{+}09, YH09a]. goal [Kor06]. goal-oriented [Kor06]. Goldstein [LmY08]. Göllnitz [VK06]. Good [VC06, HE06b]. goodness [BdWG06]. goodness-of-fit [BdWG06]. Google [LSW09, Wu08]. Gopalsamy [MK05]. Gordon [WCZ07, Bra07b, DS09b, Eba09, OOO^{+}07, VK06]. Gori [Yan05, PS08]. governed [CY09, LY07b, YZ09a]. GP [CGPS08]. grade [CLA09]. graded [CC07, MJ09, ZC08b]. gradient [And09b, BW06, CLQ09, CXH09, DC09c, GZLL09, HL07a, HTS09, HLF09, LNW06, LTW07, PV09b, Pul09, SG09, TBL09, WLY05, YLW09b, ZZ06a, ZZ08a, Zha09b, ZJ12]. gradients [FPP07, How09b]. grain [CLQ09]. grain-size [CLQ09]. Gram [RH06]. graphs [CW05, CW08d, PPG07, TL07]. Grassmannians [GK07]. gravity [eMH06a]. Gray [ZWZ08]. Green [AG09, MM07, MCA09, SV09, YZ09d]. grid [BS08b, BB08b, CYB09, CHJL08, DC08, HK08, LO08, MG08, MS09c, NSS06, Wac05, ZT06a]. grids [ABD^{+}09, Bi08a, BB09b, Car07, EM07, EM08a, GH08, KDN08, MK09, WSST05, YLC09]. Gröbner [STB09]. Gross [Wan07a]. Grossberg [SC06a]. groundwater [KY07]. Group [HeM05, Kas06, RK08a, Val09, ADLT09, BeMBH06, DW05, EKEHR08, IS06, NP08, TM08]. groups [CK07, LV06, LV07]. growing [SLN06]. Growth [CP07, BRVWM07]. Grünwald [GAR07b]. GS [Li05d, NKA09]. Guaranteed [NH08, WK07a, KRAH07]. guarantees [HLF09]. Guest [GD09, VK08]. guild [YTM07]. Gumbel [GK08]. gyroscopic [YD09b].

Haar [BS09a, BSS08, BSS09, RA07, ZYLY09]. Hadamard [FFK05, PK06, Par07b, Par07a, Par09c]. Haddock [ZWF09]. Hahn [BR09, FLS08]. Hahn-type [FLS08]. half [BGVHN05, CMV09, LZ08b, LZW09, OYA07, Sak05, YA07, ZM09a]. half-line [LZ08b, LZW09, OYA07]. half-linear [Sak05]. Halley [ABBM09, Osa08, YL06, YLS07]. Halton [VC06]. Hamilton [Qiu07, TS06a]. Hamiltonian [FS07, MS07d, Tup07]. Hammerstein [AEBEK09, Lan06]. Hamming [Iza09]. Han [JH09]. hand [CHSW08, GQ09a]. handling [CKC^{+}07]. Hankel [AH07, DS05a, EJ08, EK07, KM09a, NP07]. haptotaxis [Eps09]. hard [KS07e]. harder [PS09b]. hardware [KDH08]. Hardy [EL07, Gau05b]. Harmonic [Cho05, Swe07a, BS06, CY09, DDLM07, DFLP08, LN07b, WGZ08, Wu07]. harmonics [DLROY05]. Harten [ABT06]. harvested [Kar06]. harvesting [KOS07, LZZD09]. having [SHC08]. hazard [KK09a, Men09]. head [KY07]. health [DÖ09]. heart [VDV08]. heat [eMH06b, AZGN06, BF08b, CCH06, Dao07, Dao08, DA08, DFY09, EM07, Fer09, GS07a, GKMS09, HTS09, Har08a, HeM05, JL07b, Kas06, Lan05, MMM08, RR06, TQA06, WLMS08, XF07]. heated [CLA09, KD05]. heating [VT06]. Hedging [WWFV07]. Hedstrom [YS06]. helices [BM07a]. Hellinger [Sur05]. Helmert [Wat06]. Helmholtz [BFT07, EZ07, Gil07, GDA07, HK07a, Koy07, Koy09, Li06b, LS09, Lu05, QWS09, RS08a, RW09, SHK^{+}08, Sut07, ZKSS06]. Helmholtz-type [SHK^{+}08]. Hemivariational [Noo05]. Hemivariational-like [Noo05]. Hermite [Boy09, CV05a, DSZ05, DH09, Fin08, GJ08, Han09, HYC05, KW07, KSVW07, LMS08a, MST05, MW08, MW09c, Müh06b, NBSV08, Par09b, PS08, Sab08, XY05]. Hermitian [CL06c, DHM07, Du07a, DL09c, HWL09, JC09, KN06, KMB09b, LHL06, LvGJ06, Toy07]. Hertz [GKM07]. Hessenberg [Gem05]. Hessian [And09b]. Hessian/vector [And09b]. heteroclinic [Tac07]. heuristic [KCK09, Tok07]. hexahedral [Zha07]. HHDMR [TD06]. hidden [ABF^{+}07]. hierarchical [MB06]. hierarchies [LZZ09]. High [AS05e, BFT07, Bor07, CCD09, CH09, CD07a, Dus05, GGG09, Har08a, LN07a, LYT07, MOZ07, MV05a, Par09c, SS06c, WG06, YLSX06, ABB^{+}07d, ACWL07, AE07, BS08b, CS07a, CM07, DZ09a, Dor09, Dub09, Gil07, Huy09, IV06, JKM07, Kan08, KHWF09, LW07, LCZ09, OS06b, Pet09b, RVAD08, Rom09, TGB08a, TGB08b, TD06, VVE08a, VMMM09, WWC07, WS08, YSYX08, ZG09, ZW09b]. high-dimensional [BS08b]. High-frequency [MOZ07, ACWL07, AE07, IV06]. High-order [AS05e, BFT07, Bor07, CH09, Dus05, GGG09, SS06c, WG06, Dub09, Huy09, JKM07, Kan08, RVAD08, TGB08a, TGB08b, VMMM09, WWC07, YSYX08, ZG09, ZW09b]. High-performance [MV05a, Rom09]. High-precision [Par09c]. Higher [AGH06, Fin07, LTY08, WZ09a, ALR05, DLG05, Eig07, FH09a, FH09c, Gei08, HCL08a, KCK06, KL09, LWH09b, PM05, PRP08, Rad05, RP05, SS09d, VMV07, WKY^{+}07, WT08, YM07, YL09b]. Higher-order [AGH06, LTY08, WZ09a, DLG05, FH09a, Gei08, HCL08a, PM05, Rad05]. highest [KK09b, SK09a, WeR07]. highly [AT09b, AT09d, BJ08, GLS07, HC09, HV06c, Xia07a, Xia07b]. Hilbert [AB09a, Bac09, BFM09, CQ09, FKL05, KM05, Pie08, WMN05, WN08, ZYLY09]. Hilliard [BN08b]. Hirota [CX06].

Histopolating [Sie08]. histopolation [FO05, FOT07]. HIV [ABD^{+}05, CLGG09, IOS07, IOO07, Mer05, RMKA05, Ver05]. HIV-1 [Ver05]. HIV/AIDS [CLGG09]. HJB [ZZ08b]. hodograph [KL08b]. Hoff [FZ08]. Hölder [HQ06, LS08, YL06, YLS07]. hole [LLM08]. Holling [LCH09, LTC06]. home [BNDN09]. homeomorphism [HJK09, SS09b, Wei09]. homoclinic [KW07]. homogeneous [Cao08, LS07a]. Homogenization [CJCD06, CJCD07, Dul08, FIR06, Las08]. homomorphism [HJK09, SS09b, Wei09]. homotopy [AHM08, FHM09, GS07a, Li09a, YXF07]. homotopy-perturbation [AHM08, GS07a]. Honeycomb [XY05, LCZ09]. Hook [Dun07]. Hook-lengths [Dun07]. Hopf [Yan06a]. Hopfield [FL09c, YSYX08, ZG09, ZW09b]. horizontal [GF08, LFS08, LFW09, WB09]. Householder [SEAA08]. HPM [GC09]. HSS [LHL11, GT09]. huge [BLR07]. hull [DL07b]. Humbert [AGRZ06]. Hurwitz [Cof08, Cof09a]. Hybrid [CD07a, Hsi09, LLM08, SS06b, TD06, AOSY05, BDL09, CY08c, CY08b, DVCH06, DDL09, EM09, FT08, Fra06, GDA07, GZLL09, KCLW08, KS07e, LZWQ09, NS09a, PDM08, Van07b, VT06, WT08, YW08, ZZ08a]. Hybrid-Trefftz [LLM08]. hydro [NT08]. hydro-mechanical [NT08]. hydrodynamic [HS06]. hydrodynamics [eMH05]. hydrogenic [DLROY05]. hydrostatic [BB08a]. hydrothermal [BGRS05, BGRS06, BSMT09]. Hyers [Par05c]. hygro [BM08b]. hygro-thermal [BM08b]. Hyman [GH09]. hyper [GST06]. Hyperasymptotic [PK06, Par07b, Par07a]. Hyperasymptotics [Daa09a]. Hyperbolic [ZL05, BUG07, CHY05, DG08, DZ09b, EMP08, GGD08, HK08, zLCgS06, LP07c, MCR08, PH09, TS08, WWC05, XCZ08, YA08, ZLZ09b]. hypercube [PVC06].

Hypergeometric

[AGRZ05, BE06a, BE06b, CJ09b, DSMY06, DJ07b, Ell06, Ext99, FLS06, LL01, LV06, Mil06, Mil09, Par05b, RI06, RVK05, RAG07, RS05d, Rui05, SRD05, Sto05b, Swa06, Tur09, Vid05, ZYD07, Zud09]. hypergeometric-type [ZYD07]. Hypergroups [FLS05, Ass09a, Las09]. Hyperinterpolation [CDV07]. hypersingular [FG09]. hyperterminants [Daa09a].

I.V.P. [RVA07b]. ideals [Sau08]. idempotent [TS06b]. identifiability [Jai06]. Identification [BP05, FZ08, AS07, Gon06, LC08b, MDT09, TBL09, UHNC09, YOO07]. identify [YYF09]. identifying [CK05, DFY09]. identities [CJ07, GZ05, VK06]. identity [AGRZ06]. IFC [Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano07-39, Ano07-40, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano09z]. II [LW08b, PPR07, AMG07, BS05, Ber07, GA09, KG09, KSVW07, LWH09a, LTC06, Par08a, Ped09, Slo08, TR08, ZL05]. III [CKWZ07, Par07a]. IIPG [Dol08]. ill [BR07a, Cal07, Gon09b, HL08, LHW07, MRSZ06, MRSZ07, OTOR07, RS08b, WHL08, WH09a]. ill-conditioned [OTOR07]. ill-posed [BR07a, Cal07, LHW07, MRSZ06, MRSZ07, RS08b, WHL08, WH09a]. ill-posedness [HL08]. ill-scaled [Gon09b]. ILU [KT07b, RT09, RH06, VS08, Yun05a]. Image [BP06, DMM07, NQDJ06, AAPP07, AA08, BS09b, BJ07b, DESC07, Est07b, FL06b, GG09b, HT06, LM07a, LNL09, SRW09, TM08, Yi09, ZLdST09]. images [AAPP07, ABB^{+}07d, TND^{+}09]. Imaging [BKT07, BBM07, FIR06]. imbedding [CMM09]. imbeddings [Maz06]. IMEX [Kot08a, Kot08b, SSV06]. immiscible [OS09]. immune [RMKA05, Wod05]. immunity [Zin05]. Immunological [CPM^{+}05, LRB05]. Immunology [BBG05, Zin05, ABB07e, BBF^{+}05]. Impact [SZ07b, CSG05, DK06a]. impact-induced [DK06a]. impedance [BKT07]. Implementation [MSA09, WWZY09, HS08b, PT09b, SM07, SFFS08, SK09b, VAR06, WSST05]. implementing [AG09]. Implicit [AZGN06, ACQ09, Bno07b, BS09c, BB08a, BB09a, BMM06, BM07b, Chr09, Din05a, EMP08, FN09, Gug06, HL09b, HyLC06, KBA09, Kom08, yL08, MVVA06, MS07c, NSC07, PJ08, Qin09, RS05c, TYZ^{+}05, Tem09, VVE08a, VVE08b, WH09a, rWlC06, ZWZ08, Zhu07]. implicit-explicit [ZWZ08]. Imposing [SK09a, WLL09]. Improved [AMV06, BW07b, Kaw07, hLyHZ08, MNG09, OS06b, PLMP08, Abd08, BIR09, GZLL09, LmY08, LWLL09, NS09b, RP09]. Improvement [SO05, ZHG^{+}09]. Improvements [tLxLlW07, Yun08b, NT06]. Improving [EHR09, GWG08, GS09]. impulse [Kar09, Liu09d]. impulses [LLZY06, LS06b, LS07b, LZG^{+}09, ZG09, ZW09b]. Impulsive [JL07a, LFW09, AAZ06, CS07d, CHY05, DC09b, FL08b, FL09c, FL09d, yL08, Li05b, LH05a, LF07a, LS07c, LTC06, LZ09d, MS08a, MD09a, NPTH09, SZ08a, SZ09b, TSC10, WPC08, YS07, YZL08, ZCS06, ZS05, ZLL08]. imputation [MR09, WWWF09]. IMT [Oou08]. IMT-type [Oou08]. inclined [EKEHR08]. Inclusion [Swa06, ABF^{+}07, Di 07, Din05b, HC08, PM05, PIP07, TFC07]. inclusions [AU09, Din05a, DF08, FD09, HLH07, Hon08, Kaz06, KBA09, Kon07, PZ08]. Incomplete [Har08b, STB09, BDY09, Dun06, GSBB09, GM07c, KS07c, KSS07a, Par05a, Par08a, YH09a]. incompressible [Deu07, Deu09, ER09, GLOS05, GRW08, SY09, TY05, TY07b, Wat09, ZZL08]. incorporating [Kar06]. increasing [Eds09, HJK09, SS09b, Wei09]. indefinite [BR07b, Du07a, Gan08, HS07a, LvGJ06]. independence [GK08]. independent [AK09, CHS07, DA09, Nig07, ZGSF08]. Indeterminate [Chr05, BGVHN08, Ped09]. Index [Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano08k, Ano08l, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano05b, Ano05a, Ano05c, Ano05d, Ano05e, Ano05f, Ano05-41, Ano06b, Ano06c, Ano06d, Ano06e, Ano06a, Ano06m, Ano06n, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06-38, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07-41, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Cam08, CM07, CK09a, CV06, DL09a, Den09, DW05, EHR09, HY09b, LWH09a, LRH07, LN09]. indicator [HAS09b, KKLM06]. indices [WWW08]. induced [DK06a, DFHS08, FIR06, FN09]. inducing [Bra07a]. industry [FGRW09]. Inequalities [AA09a, Alz06, JS08, Bad08, BT06, Bno07a, Bno07b, CY08a, CAY09, Che09, CV05a, Fu08, GG07, He06a, HZW08, JM07b, LHM05, LmY08, LHL06, MC07, Ma08a, Maz06, MJ07, Noo05, QCK09a, XH08]. Inequality [Ple09, CLP08, Dai09, Gon06, Han08, bJQlZ07, lLpZ08, LH09, Liu05, MZW06, Qi07, QG08, Qi08, QGC08, Qi09, ZZ08b, Zhu05b, Zhu08c]. inertial [JKM07, Mai08a]. Inexact [Don09, AUY06, AML07, Arg09b, Bno07b, BS09b, CY08c, CL06a, LS08, WA08, Zhu06]. infection [IOS07, IOO07, IV07, MSB05, Mer05, Zin05]. infections [Wod05]. inference [YYF09]. Infinite [BF08a, BS07, BeMBH06, CW08c, CW08b, FLMR07, FL09d, GLdS08, HL08, HeM05, LLL06, LWYT06, Luo08, MW09a, MS05b, QCKK09, RLX08, TR06, TSB09, TLQ08, VRSS08, Ver08, WW07, XWT09, YOA06, ZLL08]. infinite-dimensional [BS07]. infinitely [LP08]. infinitesimal [BS06]. infinity [Mor07]. Influence [BSMT09, Ham09]. Information [DLROY05, DSMY06, DOY08, GA09, GH06, SRD05, WWWF09]. infrequent [yCyZ08, yCyZ09]. inhibitory [LMZ08, Liu07]. inhomogeneous [CWMT07, Di 07, EGM07, LD06, SHK^{+}08]. initial [AUY06, AOSY05, BP05, DWW06, GGG09, LS07b, Mar07a, PK09a, Pre08a, WZ08]. initial-value [AOSY05, DWW06, GGG09, Pre08a, WZ08]. Initializers [CLM06]. injection [BeMBH06]. Inner [PK07, AK09, RS05b, TG09]. inner-outer [AK09]. inpainting [DMM07]. Input [HL07c, Kho09]. Input-to-state [HL07c]. inputs [Kho09]. inserting [Won09b]. instability [BB08a, CB07, OOO^{+}07, WZS06]. instructions [CGRGVP08]. instrument [GM06]. insurance [BBM09, DÖ09, VV09a]. insurer [BC09a]. integer [HM09a, MS05b, lSsZ05, ZF09]. integrable [LZZ09]. Integral [FT05, GPS08, WSZ09, AEBEK09, ABL07, BM08a, BS08a, BS09a, Bar09, BNP06, BF09, CRS05, CC06a, CC08, CMR06, CMV09, Car05a, CT09c, CD06, DM07b, DL08b, DL09a, DMGVP05, DFLV06, DL08c, Dio09, ELA09b, FJG08, FDFL07, FTT^{+}05, GJ08, GKM07, GC09, Gon09a, GS08, HC09, HS07a, Hon08, Hsi09, HV06c, Jan09, JM07b, KS07c, KSS07a, Lan06, LL01, LHM05, LS09, pLT07, MJ09, MO07a, MB07a, MRV09, MS08b, MNMS05, PK06, PV09a, RS08a, RZ07, RA07, Sto05b, TRST09, Vol05, Xia07b, ZFG08, ZWY09]. integral-functional [Vol05]. integrals [BBC07, CJP08, EJ07, EJ08, GLS07, Gao05, Gao06, Gau05a, Gin05, Has09a, HCK07, Kál05, KAC05, KM09b, MS06a, Mon09, Scu09, Sin09, SK06, Xia07a, XW08, Par07b, Par07a]. integrals-III [Par07a]. integrals-IV [Par07b]. integrands [Gon09b, Has09a, HS07a]. integrated [DA06, MDT07]. Integrating [ADDdM05]. Integration [GH08, ADMM07, Arp07, BG07f, CE09, FG09, Gar07a, Gei08, HPS06, Ham09, HO09, Jay07, Koz08, Koz09, Kul09, LRH07, LLD09, Mon07, Pet07a, RTGB09, RVA07b, Rza09, Sid05, Sid08, Tay08, Zha07]. integrations [LH08a]. integrator [CVB07, GOT06, KHWF09, MBCV09]. integrators [CJB05, Cas07, Cas05, LG07, VRL07]. integro [AT09a, BS08a, CS07d, Gan07, HZBM06, Has06, yL08, PT06, PT08a, WSZ09, YS07, YCJ09, jia09b]. integro-differential [AT09a, BS08a, CS07d, HZBM06, Has06, yL08, PT06, PT08a, WSZ09, YS07, YCJ09, jia09b]. intelligence [PPV05]. intelligent [LZWQ09]. interaction [BC09a, Bla05, CC06b, HKM08, MD08, VDV08]. Interactions [Dem07a]. interchanging [dB05a]. interconnected [HGS06]. interest [FW09, LCYZ08]. interesting [HK08]. interface [ADF07, AK09, DHS09, FNW06, OS07b, OW06b, TL09, TV09b]. interfaces [LLL06, Soh05]. Interior [GSS07, BS09b, Cho08, ER07, Fan09, FKM05, GNP09, WB09, Wu07, YXF07, Zhu05a, Zhu05b, Zhu06, Zhu08a]. interior-point [FKM05, WB09]. Interlacing [BV06]. internal [SC07a, SBK07]. Internalization [LNW06]. International [PRZW08]. interpolant [Bar09]. interpolants [BRIP09, BISS05a, BISS05b, BSST08, DL08a, Han09, LMS08a, MST05, PS08, PR07]. interpolated [XCZ08]. interpolating [Bou07b, DL05, DS09a, HKZ09, JH08b, KP08a, KM09c, MW09c, Rom09]. Interpolation [Bel06, Car07, ADD^{+}08, ABET08, BRIP09, Bej06, CDV08, CG08b, CHW09, CHNZ08, DJ05, DH09, DK05, DZLL06, FL09a, Fin07, Fin08, Gás09, Gau08, GM07a, HYC05, HT06, HS08c, JAE09, KC07, KY06, LWC^{+}06, LS06c, LP06b, LMS08b, MW09b, Mag09, MDT09, MJ08b, MT05, Müh06b, SH06, Sch08, WM09, WL06a, WZD09, XZS07, XC09, ZT06b, ZW06b, ZZ06c, dB05c, dB05b, dB05a]. interpolatory [BDGV05]. interpretation [Ber06b, ZLA07]. interpretations [He07]. interproximation [PJ06a]. intersection [GLdS08]. Intertwining [Xu05b]. Interval [Aue07, CKC^{+}07, KMB09a, KLS^{+}07, Mar07b, Wan07b, WML08, AV06, BCGA05a, FDC08, FPO05, FHL07, GVC^{+}08, GBR09, Mar07a, May07, MC05a, PT08b, RKAH07, Sch07b, Sin08, VB05a, WY07, WYW08, YOA06]. Interval-type [CKC^{+}07]. intervals [Boy09, DL08b, KKS^{+}07, PV09a]. intra [YTM07]. intra-guild [YTM07]. intracellular [ZBB^{+}05]. intracellular-signaling [ZBB^{+}05]. Introduction [GPS09, ZFS09]. intuitionistic [MM09, SVC09]. Invariance [LV06, KS06]. Invariant [Sto06, MvdMV06, RD08]. invariants [Arp07]. inventory [Siv09]. Inverse [CM05a, KLV09, Yua09, BMfX07, BGP07, BW07a, CD05, CCM07, Cak07, CL06b, CHS06, CDMR07, DW05, FI07, FKL05, GPP07, HTS09, HM09a, KS07a, LDM09, LF07b, LHW08, Mat06, Mor07, MN07b, SCG08, SZD09, Smo06, TA09, WD05, XZZ09, YM08, YL09a, Yi09, YD09a, Zho06]. inverses [Car05a]. inversion [Cal07, CDM07a, ESAD05, KSE09, PS09b, RS05d, YH05, YWH05a, YWH05b, vdBHPS07]. inverted [Ata07]. invertible [Aki08]. inverting [CDM07b, OTOR07]. investigation [LKSC06, LM09a, RDH07, SBK07]. investigations [QW08]. invex [Ant08]. invexity [MPR07]. inviscid [TY07b]. involving [AHS08, Alz06, Ant09, BXC09, CJP08, CJ09b, Din05a, EE06, EJ08, Fin08, Hu05b, LLL06, LD06, VG09, WYS09, Xia07a]. ion [Dem07a]. ion-acoustic [Dem07a]. IRKC [SSV06]. Irrationality [CS09b]. irreducible [HYG08b]. irregular [BS08b, KW09]. isolation [LL07b]. Isoparametric [BCH06]. Isospectrality [JLNP06]. isotropic [TANT09]. Issue [BBG05, GD09, WS05, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano06-28, Ano06-29, Ano06-30, Ano06-31, Ano06-32, Ano06-33, Ano06-34, Ano06-35, Ano06-36, Ano06-37, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano07-38, Ano07-39, Ano07-40, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano08-33, Ano08-34, Ano08-35, Ano09z, BBF07, GPS09, VAW07, VK08]. issues [Gug06, SC06b]. iterated [CLN09, Dio09, HP07, PRGPS09]. Iterates [BS06, Bog05b]. iteration [AETE07a, AETE07b, Abb07a, Abb07b, AS05b, BS09c, DC08, DS08, ESAD05, GS07a, GJM07, GT09, HVV09a, He07, JC09, JK07, KMB09b, Lu07, LHL11, MP08b, Pet09a, SA07b, Swe07a, Swe07b, TD07, Tem09, TY07a, Waz07a, Waz07b, XHW07, YLS07]. iterations [EFS05, GPRB05, GPPRRB06, GPPRB07, MSE08, SC06b, Thi09, YL06]. Iterative [ABB^{+}07d, CMT09, Dan06, DRV08, Gei08, HLH07, Kaz06, Kon07, MW07, MRSZ06, yPESlZ07, PRZW08, ZTY07, AT09b, AKK06, ABG06b, ABP06, ABGH08, ABBM09, AK09, ACQ09, BMfX07, BGP07, BRW09, BK06, Cal07, CMR06, CdVT07, CY08b, CAHAY09, CAY09, CL09a, CHS08, CQ09, Din05b, DF08, EM08b, EHR09, FP08, GM09c, Hei05, HYG08a, JH08a, KBA09, Koç08a, Koç08b, KL09, tLxLlW07, MRSZ07, Ned08, NK10, NT06, QCK09a, Sha09a, SFPS09, yS05, THC08, Uje07, WHF07, WZ08, WtL09, WS09a, WWM07, WHZ09, XW08, Yao09, Yun08b, ZZ07, yZXL09, ZZ08b]. IV [Par07b]. IVPs [AHM08, Fra06, LW07, PS05].

J [AMO08, Asl10, ByXy06, yCyZ09, Dai11, GPPRB07, KYF05a, LC09b, LB10, LHL11, SZ09b, Sto05b, TGJ19, TSC10, Yun08b, ZJ12, vDvFZ13]. Jackson [Ple09]. Jacobi [AMV06, ADGR06, Cho05, CSZ09, Coj09, CS09b, EKL^{+}07, FLS08, Hoc08, JM07a, LW06a, yPlH05, PV05, Qiu07, Sau07, Sch05, TS06a]. Jacobian [Car05a, HMT09, Śmi07]. Jaulent [GJM07]. Java [OOM^{+}07]. Jensen [DC09a]. Joint [KKLM06]. Jordan [Wu08]. July [VGT09]. jump [HK07b, Kru05, LWS09, LP07c, Pok06, TGJ19]. jump-diffusion [HK07b, LWS09, TGJ19]. jump-diffusions [Pok06]. jumps [BMW08, BLP07].

Kantorovich [HSL08]. Karlsson [RS05d]. Kármán [CG06]. Kawahara [Asl10, Ass09b]. KdV [CX06, Dem07b, Waz07b, ZM09a]. Keith [WS05]. Kepler [Boy09]. Kernel [HM08, MS09a, vdBHPS07, BFGM06, CG07, EIM^{+}09, GJ08, HV06c, LS07a, Mus09, SL05, WMN05, WN08, XHK^{+}08, Zho08]. Kernel-based [vdBHPS07]. kernels [CCD09, PT06, PT08a, PV09a]. Kershaw [Qi07, QG08, Qi08, QGC08, Qi09]. key [SD06]. kind [AGRZ05, ADGR06, BM08a, BS09a, BF09, DK06a, DL08b, DZZT07, GJ08, Hsi09, pLT07, LTM08, MB07a, SWYZ09, Skr05, WPC08, ZLZ09a, ZL07]. Kinetic [CG05, BGG08, JSS07]. Kirchhoff [HGS08]. Klein [DS09b, Eba09, WCZ07]. KM [Mai08a]. KM-type [Mai08a]. knots [Rie06]. knowledge [WK07a]. knowledge-based [WK07a]. Koch [NSS06]. Kolmogorov [MDMRP09, DP05a, Wat09]. Koornwinder [vD05]. Kopel [GG08]. Korteweg [BBSV08, DA08, SWZ05]. KP [LWW08]. Krawtchouk [FLS08]. Kronecker [XDW05]. Krylov [BJSS06, DGR07, FLMR07, Mil05, Sea07, vdESvG05]. Kummer [BE06a, CZ08b, Par05b]. Kummer-type [CZ08b, Par05b]. Kuramoto [LT05, WG09]. kurtosis [QWW08]. Kutta [AMCM05, AMCM06, Van06a, Van07a, AM09a, AS05f, BJ07a, BJ08, BH07a, BH09, CGPM08, CFMR08, CFMR09, Cam08, HVV09a, Han07, HL07c, Jay07, KR07, Koi05, Kom07a, Kom07b, Kom08, Kot08a, Kot08b, LZ06, NSC07, RVA07a, RB08, SKAW09, Sun09, Toc05, TS05, VMMM09, Van05, Van06a, Ver06, WLS08, WHL08, Wan08a, YLN09]. Kuznetsov [Bat08, MNHA09].

L [Sto05b]. L. [Dom07]. lacunary [MM09]. ladder [ZYD07]. Ladyzhenskaya [Fai07]. lag [HMG06, LW07, SS05]. Lagrange [ABET08, BG07a, BNPR08, CDV08, CM07, CHNZ08, DJ05, GZX09, HYC05, LWC^{+}06, MB06, XC09, ZW06b, ZH09]. Lagrangian [BXC09, BB09a, HPS06, Sta07, TJ08]. Lagrangians [JKM09]. Laguerre [ByXy06, ADÓ07, Ass09a, ByXy05a, ByXy05b, BR09, DLROY05, DC09a, DR09, DJ09, GW06b, MC05a, OS07a, PT09a, WGZ08]. Lambert [CS09b]. Lamé [OOO^{+}07]. laminar [eMH08]. laminated [Civ07]. lamination [Dul08]. Lanczos [RP05, WL09a, YB05]. Landau [Cim08, Sto05b, Ban08, Lei08a, Lei08b, LLT07]. Landau-type [Lei08b]. Landweber [LHW07, WHL08, WH09a]. Langevin [CdVT07]. Laplace [CDM07a, CDMR07, CDM07b, FZ09, FI07, Kru05, Li07b, Par07b, Par07a, QFX06, Woj06]. Laplace-type [Par07b, Par07a]. Laplacian [BE07, Gal09, GJL08, He05, HWY09, Hem06, Hon07, Liu08a, LOA05a, LOA05b, LJ09, ML09, Rea07, SNV06, WSM^{+}09, YH09b, ZFG08]. Large [ZFS09, BR07a, BY09, CC08, CXH09, Cho08, CMM09, Dmi08, FLS06, GKMS09, GM09c, Hsu05, Jon06, LT06a, PKS^{+}09, SCY09, VL06, WZ07, WCL08, Wu07, XZ08, YH09a, ZJ12]. large-deformation [Dmi08]. large-scale [CXH09, GKMS09, PKS^{+}09, VL06, XZ08, ZJ12]. large-update [Cho08]. LaSalle [Liu09a]. laser [CC06b, FIR06]. laser-plasma [CC06b]. lasers [KD05]. Lattice [DL07a, Ell06, HE06b, LZ05, ST05, Sin09, WHG07]. lattices [CK07, KV09a, LRH07, Mag09]. Laudatum [BE06d]. Laurent [CBGV05, CBDGVN07a, CBDGVN07b, DMGVP05, UHNC09]. law [AMG07, BIR09, GW08, Las08]. laws [Ben06, CD07a, JKM09, LZZ09, Mur07, TS08, ZL05]. Lax [Qiu07, WHG07]. layer [eMH08, CH09, LT06b, Vul07, YZ09b, ZKSS06]. layer-resolving [Vul07]. layers [CWMT07, SÁ08b, Zar09]. LBB [GLOS05]. LBB-stable [GLOS05]. LBT [DESC07, Est07b]. leader [CFGVP07]. leaky [Har08b]. leap [BD06]. leap-frog [BD06]. Learning [DC09c, AMV06, EL06, XC08, Zho08]. Least [DG08, DJ07a, BTI08, CR09, Cou09, DKHNZ08, DB08, GGD08, GRL09, HR05b, JZW08, JSS07, KA07, Log08, MJB09, MV05a, RKK07, RKK09, SV07a, TY05, WC07, WA08, XZL08, XF07, YH09a, ZH08, ZSWL09, ZLDW09]. Least-squares [DG08, DJ07a, CR09, Cou09, DB08, GRL09, KA07, RKK07, RKK09, TY05, XZL08, YH09a]. Lee [DD07]. left [EKL^{+}07]. left-definite [EKL^{+}07]. Legendre [EJ08, GW06b, HS06, Pré08b]. LEM [CVB07]. length [FR06, MS06b, SFV09, SLZ06]. lengths [Dun07]. less [MGS08]. Lessons [CPM^{+}05]. Letnikov [GAR07b]. letter [Gau08]. level [BSV08, DMZ08, Gás09, KD05, lLpZ08, MS07c, XDW05, ZH09, CKWZ07]. level-2 [XDW05]. levels [LLS09]. Levenberg [KYF05a, KYF05b, MW09a, Wat07, Zhu08a]. Levin [Xia07b, XW08]. Levinson [VMV07]. Levinson-like [VMV07]. Lévy [HR09b, RLH09, Sch06b, YW09a, ZLA07]. Li [Yun08b, Cof09b]. Liao [Bno07a]. Lie [BG07b, BeMBH06, CK07, EKEHR08, LLZ09, Röb08]. Lie-group [BeMBH06]. Liénard [LTM08, Liu08a, MLH08, ML09, SWYZ09, WSM^{+}09, ZL07, ZSZ09, ZCZ08]. Liénard-type [MLH08]. life [BBM09, VV09a]. lifetime [LWH09a]. Lifshitz [Ban08, Cim08]. lifting [CHXL06]. light [THS05]. light-limited [THS05]. like [Arg09a, BDHW07, CX07, Dai09, Din05b, FD09, FTT^{+}05, GR09, HR05b, Kaz06, KLO08, MP08a, Noo05, PG07, Pie09, SEAA08, Sch07b, VMV07, YH05, YWH05a, YWH05b, ZT06b, Zud05]. likelihood [FvZ06]. Limit [Cao08, Sch08, ADD^{+}08, AGMMB05, DL09b, GGT07]. limited [Fer07, THS05, VL06, Wri08, XZ08]. limited-memory [VL06]. Limiting [JR08, Lei08a]. limits [CJCD07]. line [BGVHN05, CMV09, DEMR08, bJQlZ07, Kam07, LZ08b, LZW09, hLyHZ08, OYA07, Pet08, QZY09, SS09c, SS06a, TJ08, XSW09, XSP09, YP08, ZM09a]. line-search [DEMR08]. Linear [LWLL09, RAG07, AUY06, AT07, AAZ06, AKK06, AMCM05, AK07, ASL07, BR07a, BW06, Bar09, BF05, BBR05, BM06, BY09, BJSS06, BdAR08, BW07a, BCP07, BKT07, BW07b, Cal07, CJCD07, CLP08, CJB06, CMST06, CHW09, CHS08, Cho08, CCLS06, DaC05, Daa09b, DG08, DW05, Don09, DJ09, EM06, EIM^{+}09, EMMP07, FLMR07, FJT07, FS07, FL06c, FG07, GMM09, GKM07, GRW08, GHZ06, GHT06, GM09c, GT09, GQ09a, GNP09, HL08, HTS09, HNS09, Hei05, HS08b, HM08, HLL08a, Jan09, bJyKyZmT09, JD05, JH08a, KGA08, KT07a, Koh08, Koi05, KMB09b, Li05c, Li05d, LmZbL07, LW07, LM09b, LHR09, LHL11, MJ08b, MSE08, Moh05, MS07c, MLOP06, MRSZ07, MN07b, NS09a, NKA09, OOM^{+}07, PKS^{+}09, PN06, PP05, PK07, RB08, Roe08, RKK09, RO07, Sak05, SS09a, Sch07a, Sch07b, SAD07]. linear [SFFS08, SV07a, SZ08b, SZ07a, SSZ09, Sta07, SF06, Swe07a, TRST09, Ter05, THC08, Var07, WCH05, WHF07, WLS09, WCW09, WB09, WD05, WWM07, XDW05, XZS07, XW05, Yan06d, YA08, ZW06a, ZW08, yZXL09, ZGH09, ZM09b, ZSWL09, Zhu05b]. linearised [Ehr06]. linearization [CS07c, OS07a]. Linearized [Har05]. Linearly [ZW08, CGP09, PJ08]. Lines [WS05, Bra07b, CG08b, Cas05, HKB^{+}05, HEOS05, Mil05, SWZ05, WSST05, rWlC06]. linked [VV09a]. linking [DVCH06]. Liouville [Abb07a, AGM06, ABI07, BV06, BBB07, BSS08, CKWZ07, Cha07, Cha08, CW08d, FPP05b, FPP08a, FPP08b, GGM06, Liu08b, MJ08a, Sim07b, SO05, WWW08, Won05, Yüc06]. Liouvillian [ADDdM05]. Lipschitz [Bel06, CL06a, Liu07, MYY07]. Lipschitzian [NS09c]. liquid [WT07]. little [CS09b, OS07a, PV05, Sza05b]. Littlewood [Gau05b]. Liu [MK05, Zha09b]. Living [Gia06]. LLL [HE06b]. LLL-spectral [HE06b]. LM [ZL09b]. LMI [MLOP06, YJY06]. LMS [VLR08]. load [DP08b, WCL08]. Lobatto [CEF08, JR07, JB09]. Local [LS08, LY05b, Qui07, Wri08, YOY07, ZCZ08, ABR07, BMfX07, BGG07, CJB06, DMM07, DL05, DL08a, Dao08, Di 07, EZ07, Han05, Han06, Jay07, JC09, Ko07, KYF05a, KYF05b, KCLW08, LH08a, MN07a, RW06, SZ09a, VVE08b, Zhu08a]. localization [BMC09]. localized [Fer07]. Locally [Rie06, Osa07]. Locating [BCK07]. location [TFC07, YJGL07]. locking [Tac07]. Log [GZX09, ZLLW08]. Log-det [ZLLW08]. Log-Sigmoid [GZX09].

Logarithmic

[Ped09, HS07a, KS07c, MOC09].

logarithmically

[GQ09b, Qi07, QG08, Qi09]. logarithms [Gau08, Zud07]. logistic [MPP09]. Logistics [YJGL07, NSK09]. lognormals [VCD^{+}08]. Long [SF06, Ben06, CP09b, LLT07]. Long-time [SF06, CP09b, LLT07]. longevity [Den09]. look [GM09c]. look-ahead [GM09c]. Lorentzian [MP08a]. Lorentzian-like [MP08a]. Lorenz [NaLH07]. loss [HO09, MvdMV06]. losses [RR06]. lossy [VT06]. Lotka [TSC10, Ko07, LC09a, LJM09, LOZ08, LL07b, NPTH09, PI06, Won09a, XWT09, YC06]. Low [DFZ07, WS08, BGJ07, CS09a, GGP09, HKB^{+}05, JF06, KM06, PC05, RDH07, VVE08b]. low-density [HKB^{+}05]. low-dimensional [KM06]. low-discrepancy [PC05]. low-frequency [RDH07]. lower [COEV05, Har06, HYG08b, LOA05b, LW06c, LLC08, xMzY09, Qi08, Shi07, WYS09, YOA06, YG09a, ZCL06]. lowering [CC05]. lowest [Hem06]. LULU [CdWJ06]. lumping [LSW09]. Lur’e [HWSL05]. Luswili [Sto05b]. Lyapunov [SG05]. lying [COEV05]. lymphocyte [Meh05, ZBB^{+}05]. Lyskova [dMFPP09].

M [AA09b, ÉT08, WWP09, WK07b]. M/ [WK07b]. M/G/1 [AA09b, ÉT08, WWP09]. MAC [FLPA09]. Macdonald [FS08]. Mach [NVD08, VVE08b]. Mach-uniform [NVD08]. machine [EL06, WCY09]. machines [TMOG07]. Maclaurin [Ber06b]. magnet [WZH06]. magnetic [ACV05, CHSW08, ICW^{+}09, IV06, KM06, Ma06, Sch06a, SV07a, VV05a]. magnetized [LmLgF09]. magneto [eMH08, HeM05]. magneto-elastico-viscous [HeM05]. magneto-forced-unsteady [eMH08]. magnetoencephalography [FP08, Suz05]. magnetohydrodynamic [BTS07, NTS06, QW06, TSB09]. magnetohydrodynamics [MDB06]. magnetostatic [DRV08]. magnetostriction [Ban08]. Magnus [DS06, GOT06, Lu05]. Magnus-type [GOT06]. Main [GA09]. Main-mass [GA09]. making [KGW09, RM07, WY07, WYW08, WLL09]. management [AVMVMV08, IOS07, Lan08a, Mar08, WCY09]. Managing [MGMTO06]. manifold [TM08]. manifolds [JLR06, LS07a, Lop07b]. Mann [MP08b]. Mann-type [MP08b]. many [JG09, KTM09, LZ08a, LZ08b, LZ09b, LZ09c]. map [HL07b, Jai06, SFFS08, SFPS09, TMOG07]. mapping [And08, BDHW07, CHV^{+}08, LH09, Pet08, SNV06]. mappings [And09a, ACQ09, BS09c, CAHAY09, CQ09, DT06a, Dai09, Din05a, HyLC06, Kaz06, KBA09, KL08b, Kon07, KLV09, NS09c, Par05c, PK09b, QCK09a, QCKK09, Tem09, Thi09, ZTY07, ZL09a]. Marchenko [CCLS06]. Marder [WWZY09]. Marguerre [CG06]. Marker [Lew06, NSK07]. marker-particle [NSK07]. Markov [BS06, EMP08, GEK09, Kag09, KR08]. Markov-modulated [GEK09]. Markovian [LkLkP09, MYY05, MYY07, PPG07, YG06, ZW09c]. Marquardt [KYF05a, KYF05b, MW09a, Wat07, Zhu08a]. Maruyama [MYY07, Mao07, PDM08]. mask [LP08]. Mass [BGGC07, EZ06, GA09, GKMS09, HHZ07]. mass-spring [HHZ07]. masses [Won09b]. massively [MBCV09]. master [Eng09, HBS^{+}07]. matched [ZKSS06]. matches [MW08]. matching [NGS05]. materials [DK06a, LN07b, MPXK06]. Math [AMO08, Asl10, ByXy06, yCyZ09, Dai11, GPPRB07, KYF05a, LC09b, LB10, LHL11, SZ09b, TGJ19, TSC10, Yun08b, ZJ12, vDvFZ13]. Mathematical [AVMVMV08, GDM08, KS07f, NSK09, Wod05, ABB07e, Ant08, CPM^{+}05, IOO07, Kro06a, WWP^{+}08, ZMAZ09]. Mathematics [BBG05, Gia06, Lig09, CWMT07, Art09]. MATLAB [GS07b, Sha08, WSST05, BCH06, Gau05c]. matrices [ÅMK06, ÁNPQ05, AB09a, BS08a, BJS08, BGGC07, CL06c, CHS08, CSZ09, Coj09, CK05, DHM07, DV08a, DV08b, Du07a, Duf09, EST07a, Gem05, GLV05, HM09a, HR05c, HYG08b, JM07a, tLxLlW07, LHL06, LF07b, LF08, LC09b, LvGJ06, MLL09, MVVV06, Mor07, NL05, NT06, OTOR07, yPlH05, RT08, RT09, RH06, Sim07a, Smo06, SS07, TV07, TS06b, THC08, TA09, TM09, VVV08, VMV07, WHF07, WtL09, Wan09b, WS09a, XW07a, XZL08, YH05, YWH05a, YWH05b, Yi09, YD09b, Yua09, Yun08b, yZXL09, ZM09b, Zho06, dF07]. Matrix [KS07b, KN06, AEBEK09, AH07, BM08a, BG07d, CJP08, CW08d, DSW08, DH09, DIKZ07, DS05a, Deu07, DL09c, DG05, DI06, FKL05, GH08, GHZ06, HYYF05, HYG08a, HV06c, Iva06, KM06, Kil07, KA07, Koh08, KW09, KV05, Lai09b, LLZ08, LSW09, LGYH09, LM09b, hLyHZ08, Lop07b, Mar07c, yPESlZ07, yPyHZ07, PS09b, RS05d, SHC08, TG09, Toy07, VG09, Wu08, WWD09, YD08, Yun05b, Yun08a, ZHG^{+}09, ZLDW09]. matrix-splitting [HYYF05]. matrix-valued [FKL05]. max [LXX08, fLxXT08]. max-bisection [fLxXT08]. max-cut [LXX08]. maximal [BC08, BDGV09, BGG07, CY08c, Din05a]. maximization [DS05a]. Maximum [FvZ06, KL08a, HKV09, KK06, SV07b]. Maxwell [AFT^{+}07, BHP07, Cim08, DFLP08, GZL07, GSS07, Ham09, Li06a, LN07b, Ma06, SHB05, SZ08b]. MBFGS [LYW08, ZZ09c]. McKendrick [PI06]. MDO [FGB07]. meal [BNDN09]. Mean [Hua08, LS06a, RB08, Sic08, Toc05, YZL08, BB05, BH09, Bra07a, CT09a, Luo07]. Mean-semivariance [Hua08].

Mean-square [LS06a, RB08, Sic08, Toc05, BH09]. means [ABÇ06, CV05a, CDL05, CD07b, IO06, Lóp07a, MPZ09, MNMS05]. measure [DS07b, EMP08, NL05]. measurement [CL06b]. measurements [Di 07, FZ09, ZS05]. measures [AS09a, CE09, CT09b, DLROY05, EST07a, GD09, LZ09d, MH08, VCD^{+}08]. Measuring [HG09, WY07]. Mechanical [pLT07, AHM09, NT08]. mechanics [ABÇ06, SRP08, Tsa09]. mechanism [VJ09]. mechanisms [TF07]. media [CMX09, Cui07, DCZ07, EGM07, FS07, GS07a, Li06a, MP06a, Sea07, Sha05, SY09, WG06]. median [AA09a]. mediated [BCeAJ09]. medium [eMH06b, EKEHR08, VT06]. MEG [FP08]. Meixner [AGMMB05, FLS08]. Mellin [Lóp07a]. membrane [KS05b]. memory [BDL09, BF08b, Den07b, SC07b, VL06, XZ08]. MEMS [FGRW09]. Menten [JSS07]. merit [CGP09, LRH07]. Merton [May08]. Mesh [AK09, AG06, AFT^{+}06, AKW05, BISS05a, BKB06, CM05b, CMST06, CGPS08, DB08, DVCH06, DW06b, FLPA09, KGA08, MJ09, MJX09, Moh05, RO05, SÁ08b, TTZ06, Vul07, ZX06]. mesh-free [AFT^{+}06]. meshes [AT09b, BFGP08, BHP07, DCF06, Gar07a, GWG08, HKV09, Koz09, LX06, LN07b, RB06, TS06a, ZC08b]. meshfree [Kro06b, uIHA09]. Meshless [PSV08, YXYJ06, AFT^{+}07, DT06b, Gás09, HS08a]. meta [BLR07, Meh05]. meta-dynamics [BLR07]. meta-population [Meh05]. metaheuristic [fLxXT08]. metal [CH09]. metallic [dCDND06]. metamaterials [Li07a, OMC08]. metapopulation [SZ07b]. metastable [OW06b]. meteorological [KKS^{+}07]. Method [HEOS05, WS05, AETE07b, Abb07a, Abb07b, AT09a, eMH05, eMH06a, eMH06b, AYK09, AS05b, Abd08, AEBEK09, AHM08, ABG06a, AOSY05, Ahm07, AHC05, AABL08, Alh06, ABG06b, ABB^{+}07c, ABGH08, AB07, AB05, AML07, AS05f, ABS07, AK09, ABP08, ACWL07, Arg07, ACV05, Asl10, Ass09b, AT09d, ADDdM05, BM08a, BNT08, BSST08, BG07a, BH07a, BH09, BSC07b, BJS08, BDHW07, BUG07, BLT09, Ber06a, BR07b, Bi08a, Bno07a, Bno07b, BFGM06, BJ06, BS09b, BIR09, BK09b, Bou07a, BFP09, BMC09, BeMBH06, BG07d, Boy09, Bra07b, BCP07, BK08, BH07c, Cao08, CC06a, CMR06, CMV09, CR09, Cas05, CJ09a, CHXL06, CHM06, CP08a, CP08b, CWW08, CLQ09, CLA09, CHS08, CL09b, CU09a, CU09b, CQ09, CEJ09, CJ07, Chu08, CN08, Cui07, CDMR07, CDM07b, DSI06, DWW06, DC08, DSC09, Dai11, Dan06]. method [DP05a, Dao07, DM07b, DB08, DL09a, De 08, DG08, DS08, DH09, DVCH06, dCDND06, DO06, DAE05, Dol08, Dom09, DJ07a, Dor09, DS05b, DT06b, DMZ08, Dub09, Eba09, EGO06, EGFO09, EEGG08, EM08c, EPP^{+}09, EH06, FHH08, FH09a, FHM09, FLPA09, FFSS09, FL08a, FKM05, Fu08, GH09, GAC^{+}08, Gal06, GGD08, GJM07, Gan08, GPP07, GMM09, GG07, GC09, GK09, GH08, GO05, Gon07, GL06, GS09, GM09c, GSS07, GZLL09, GZX09, GG09b, GW06b, HKB^{+}05, HCL08a, Han08, HTS09, HC09, Has06, HLF09, HYYF05, HH07a, He07, HMPR07, HK08, Hei05, HH07b, HeM05, HR05a, HS08b, HS06, HO07, HO06, Hoc08, HVV09b, HQ06, Hos06, Hos09, HM08, Hsi09, Hsu05, Hu05b, HLL08a, HYG08a, HGS08, HMT09, HKM08, IO06, Jan08, Jan09, bJyKyZmT09, JW08b, JH08a, JYX07, KGA08]. method [KKW06, Kaw07, KRA08, KTH08, KKLY06, KK09c, KW09, KW07, KSE09, KLW07a, KLW07b, Kou08, Koy07, Koy09, KS07e, Kro06b, KKD06, KM09c, KLV09, yL08, Lar08, Lew06, Li05d, LY05a, LWYT06, Li07b, LCH07, tLxLlW07, LHW07, LW07, LmY08, LH08a, xLLhZ08, LLZ08, lLpZ08, Li09a, LS09, LZ09a, Li09c, LY09, LZLH07, LX06, LDM09, LWLL09, LH07b, LC09b, LYB09, LYT07, Log08, LN07b, hLyHZ08, Lu07, LC08c, LHW08, Ma06, MC08, MW09b, MJ09, MW09a, MJX09, xMzY09, MDP08, MDT07, MMM08, MV08, MNG09, MS09a, MNP08, MYY05, MCR08, Mar09, MT05, Mat08, Mil05, MN07a, MZW06, MLY07, MD09b, MO08, MRSZ07, Mot06, MNMS05, MS09c, MA09, NTS06, NKM08, NKA09, NSK07, NSS07, NB06, NMMS05, NMM07, Oga06, OA06, OS09, OTOR07, OS06a, OMC08, OZL09, OOM^{+}07, PS09a, PT09a, PG07, PKS^{+}09, PT06, PT08a]. method [PRGPS09, Pet09b, PJ06b, Pre08a, Pre09, Pul09, QW06, QFX06, QWS09, QZZ08, QZY09, RT08, RVA07a, RZ07, RD08, Rea07, RW09, RA07, Sak09, SA09, SST09, SS09c, SWZ05, Sav09, SHK^{+}08, Sea07, SAD07, lSsZ05, SaLJ08, SJ09, SS06a, SG08, SZD09, uIHA09, SK09a, Śmi07, SZ07a, SSZ09, SA07b, SM09, Sto05a, SP09b, yS05, SY09, Svá08a, Svá08b, Swe07a, Swe07b, TRS07, TTZ06, TS06a, TJ08, TRST09, TD07, TG07, TY07a, THC08, TK07, TA09, Toy07, TY05, TY07b, TV09b, Uje07, UFG09, VVV08, VDV08, VVE08b, VAN06b, Vin09, WK08, WT07, WCH05, WFL07, WHF07, Wan07c, Wan07a, WG08, WGZ08, WHL08, WYL08, WFW09, WZ09a, WH09a, WWWF09, WYL09a, WtL09, WCY09, WL09a, Wat07, Waz07a, Waz07b, WL09b, Woj06, WO08, Won09a, WSST05, WLK06, WKL06, WeR07, WTZ08]. method [WWZY09, WHZ09, XKF06, XW08, XZS07, XSW09, XZZ09, XWZH08, XZL08, XF07, Xio07, XHW07, XSP09, YA07, YOO07, YZ09a, YY07a, YLC09, YG09a, Yao09, YG09b, YH09a, YXF07, Yua08, YLW09a, YLW09b, Yüc06, Yun08a, Yun08b, Zar09, Zha05, ZZ06a, ZCL06, ZZZ08, ZhL08, ZJW08, ZHG^{+}09, ZXZ09, ZLZ09a, Zha09b, ZM09a, ZLL09, ZGH09, ZZ09a, ZDC07, ZBL08, ZH08, ZZ09c, ZW09d, ZYLY09, Zho09a, Zhu08a, ZF09, ZA09, ZH09, Zhu07, ZKSS06, dS07a, NK10]. Methods [May08, AUY06, AT07, AM09a, AG06, AZGN06, AMCM05, AMCM06, AR09b, ABP06, ABB^{+}07d, ADLT09, ASL07, AR09c, Arg09a, Arg09b, AKW05, AB09b, Bad08, BR07a, BW06, BN08b, BJ07a, BJ08, BDL09, BY09, BRW09, BGP09, BJ07b, BNPR08, BJSS06, BMM06, BM07b, BW07b, Cal07, CLM06, CFMR08, CFMR09, Cam08, CCD09, CGRGVP08, CAY09, CL06a, CL09a, CLN09, CT09c, CYB09, CHJZ09, CXH09, CCK06, CX08, CD06, CMT09, DT06a, DKHNZ08, DMS05, DGR07, DK06b, DFLV06, DL08c, Dio09, DFLP08, Don09, DBG07, DBR08, Eds09, ED06, ER07, ER09, Eps09, FT08, FLMR07, FJT07, FL06b, FNW06, FR06, FN09, FZL07, Fra05, Fra06, GG09a, Gan07, GS07a, GZL07, Gar09, Gás09, Gei08, Gei09, GPPRRB06, GPPRB07, GT09, GQ09a, Guo06, HL07a, HLH07, HR05b, Hom05, Hom09]. methods [How09a, HL07c, IT07, Ish07, JL06a, JZK09, JKM07, Jay07, JD05, JC09, JR08, KYF05a, KYF05b, KVK08, Kil07, KS09, KR07, Koç08a, Koi05, Kom07a, Kom07b, Kom08, KA06, Kot08b, Koz08, Koz09, KMB09b, Kul09, KL09, LN07a, Lai09b, LVV05, LR09, Li06a, Li07a, LTW07, LS08, LZ06, pLT07, Lop07b, LNL09, LP06a, LHL11, Mar07a, Mar07b, MVVA06, Moh05, Moi05, MRSZ06, MR09, Ned08, NSC07, NST08, Nov08, OMV09, Osa06, Osa07, Osa08, PS07a, PPV05, PI06, PT09b, yPESlZ07, PM05, PIP07, PPR07, PLMP08, PJ08, Qin09, RVA05, RVA07b, RS08a, RB08, RS05c, Roi07, RKK07, RKK09, SS05, Sch07b, SX09, SG09, Sic08, SW06b, SS06b, Sta07, Tem07, Toc05, TS05, Tsi06, UO09, VRBF07, Van05, Van06a, Van07a, Van07b, Van09, VV07, VLR08, Ver06]. methods [Ver05, VAR06, VL06, VK08, WG06, WLY05, WCZ07, WLS08, Wan08a, WLS09, WS09a, WSPJ09, WWM07, WYL09b, WH09b, Xia07b, Xu07a, YOY07, YLN09, YXJG08, YJJ09, YH09a, Yu06, YH09b, YKS09, Yun05a, ZP06, ZT06a, ZZ08a, yZXL09, ZLZ09b, ZJ12, ZM09b, Zho08, ZSWL09, ZC08b, jia09b, vdBHPS07, vdESvG05, PRZW08]. metric [SVC09, VL06]. metrology [SD06]. MFS [KS07b, SK09b]. MHD [EKEHR08, HAS08]. MIC [MK09]. Micchelli [Yan05]. Michaelis [JSS07]. micro [KD05]. micro-sphere [KD05]. microbial [Liu09c]. microlithography [BRVWM07]. Micromagnetic [WZH06]. micromagnetism [Cim09]. microreactor [KPAT07]. microscale [Har08a]. microscopic [MMM08]. microwave [VT06]. mid [Cim09]. mid-point [Cim09]. migration [SLN06]. mild [EH06]. mildly [GM07b]. Miller [WS05]. Milstein [HMG06]. MIMD [Saa06]. Mindlin [XKF06]. Minimal [BFGP08, CK06, EFP06, KSS07b, SS05, TS05]. minimally [BISS05b]. minimax [Jay08, bJQlZ07, LW08c, MPR07]. minimisation [dS07a]. Minimization [Lan08b, Lan08a, MYY06, DFZ07, XSW09, ZZ09c]. minimizer [Lei08b, lSsZ05]. minimizers [Lei08a]. Minimizing [BRIP09, PVC06, GEK09]. minimum [Alo07, CCW^{+}09, Den07a, WLL09]. Minkowski [FH07]. minmax [AHS08]. Minton [RS05d]. Minton-type [RS05d]. Miodek [GJM07]. miscible [Cui07, Cui08, SY09]. missing [DMM07, MR09, WWWF09]. MITC [XKF06]. Mixed [FL09b, RS08a, WGZ08, ALR05, BG07a, Bi08a, Bno07b, BNPR08, BFP09, CMX09, CD05, CY09, CY08b, CJ09a, CL08, CHS08, CN08, Cui07, Dai09, DRV08, Deu07, Din05a, Din05b, DF08, GMM09, GDA07, Ham09, How09b, HyLC06, JLNP06, KK06, yL08, Li07a, LWMT07, LD06, MMM08, MP06a, Rea07, RKK07, SY09, Sur05, XH08, XW07a, YXJG08]. Mixed-type [FL09b, CHS08]. mixing [Lin09a]. mKdV [ZLYW09]. Möbius [dB05b]. mode [SA09]. Model [HNS09, WWWF09, WL09b, AR09a, AG09, AT09d, BCeAJ09, BLR07, BBF^{+}05, BK09a, BM08b, BC09a, BB09a, BG07c, BGG08, BNP07, CMX09, CLGG09, CK09b, Che05, CXS06, CLCL07, CC06b, DD07, DWY09, DL07a, ET06, Fai07, FIR06, FW09, GY08, GSBB09, GG08, HG09, HSN09, HKS07, Hu05a, HGS06, HL09c, IV06, JRM09, JL07a, Joh05, JSS07, KD05, KS07a, Ko07, Kho09, KS09, KN06, LH05a, LF07a, LD08, LC09a, LFW09, LCH09, LBS09, LTC06, LZZD09, LL06, MG08, MK05, NPTH09, QW08, QZZ08, RW06, RR06, SS09c, SM07, Sen07, SP06b, SMT07, SZ07b, TCMS06, TL09, Thi05, TSC10, TD06, VDV08, WZ06, WJS^{+}09, XC08, XZL08, XGM09, YTM07, Yan06a, YZ09b, YH07, YY08, Yua08, YD09a, ZWZ08, ZLY09, ZZZ09, ZZ09b, ZL09b]. model-moving [MG08]. Modeling [ABD^{+}05, LLS09, LP07c, Meh05, TK07, ZBB^{+}05, AVMVMV08, Aue07, BRVWM07, CPM^{+}05, GDM08, MD08, RMKA05, rWlC06]. Modelling [Kar06, MSTT08, WW09, ABB07e, BBPR09, FDFL07, KPAT07, MSB05, McC05, NL05]. Models [Ver05, AKL05, BB08a, BB08b, CB07, DK06b, Eps09, FPS08, GR09, Guo06, Hua07, Hua08, HO09, IV07, JVVvdZ06, Jay07, Joh05, KOS07, KS07f, LSTS06, LCYZ08, LL07a, LRB05, MMFMG07, MBCV09, May08, MPP09, Men09, NGS05, RB06, Sch06b, WK07a, WC08, Wil07, Wod05, YC05, ZMAZ09]. modes [XZS07]. modification [JYX07, Zha05]. modifications [Chu08, FH09a, HCL08a, KLW07a]. Modified [And09a, GW06b, GN09, HMT09, KK09a, NB06, QWS09, SLMW06, XZ08, AETE07a, AETE07b, And08, AR09c, BSC07b, FHM09, GXgL05, GS09, GNS06, HMG06, Koy09, Kru05, Li06b, LN09, fLxXT08, LYB09, Par08b, Pré08b, QFX06, WHL08, Wu07, WWZY09, WHZ09, YOY07, YLSX06, ZZZ08, ZM09b]. modulated [GEK09]. modules [FKL05, Par05c]. Molcanov [BBB07]. molecular [Li09b]. Moment [GA09, Ber07, BGVHN08, Chr05, EST07a, FSL09, Ism05a, MPP07, PDM08, Ped09, RW09, Sch06a, SV09]. moment-corrected [MPP07]. moments [DJ06, KM09b, LWH09b, Pré08b, SP09a]. Mond [LTY08]. monic [KMJ06, Pet07b]. monitoring [CM08]. monodomain [HRGD08]. Monotone [Bog05b, yL08, WZ08, Bog05a, CY08c, CS07c, CL09a, Din05a, FD09, Fu08, HyLC06, bJQlZ07, KA06, PK09b, Wan07c, YG09a, ZZ06a]. monotonic [CRS05, GQ09b, Qi07, QG08, Qi09]. Monotonicity [DR09, FO05, PJ06a, Waa05, AD09, CHW09]. Monotonicity-preserving [PJ06a]. monotony [KS05a]. Monte [HSN09, Lai09b]. Moore [TA09]. Morgenstern [GK08]. Morley [HGS08]. morphogen [LNW06]. Morrey [BGG07]. Morrey-type [BGG07]. Morrison [EM06]. mortality [DD07]. mortar [DHS09]. MOSFETs [LY05a]. motion [AMO07, AMO08, DMM07, FW09, JKK06, NP08]. Moving [LX06, Alh06, AG05, BBSV08, Kas06, Koz09, MJX09, MG08, Mil05, TTZ06]. MR2333703 [Yun08b]. MRLW [KRA08]. MSOR [MYY06]. Multi [AG09, yCyZ08, yCyZ09, Gás09, JRM09, LÅH08, SS09a, Zad09, BW07b, DT06a, DLG05, FJT07, FC09, GJL08, HTS09, HW08, HO09, Jia09a, KMB09a, Kon07, LLS09, LZW09, zLCgS06, MD09a, Ned08, NST08, NM06, PLMP08, RB08, SyS08, SL09, Sic08, WF09, YZ09b, ZAVS09, ZZ09b]. multi-asset [NST08]. multi-degree [HW08]. Multi-dimensional [LÅH08, HTS09, zLCgS06, PLMP08, RB08]. multi-elliptic [Gás09]. multi-factor [HO09]. multi-gases [SL09]. multi-layer [YZ09b]. Multi-level [Gás09]. Multi-objective [JRM09, Zad09]. multi-pantograph [SyS08]. multi-period [LLS09, ZZ09b]. multi-point [DLG05, GJL08, Jia09a, LZW09, MD09a, Ned08]. multi-points [NM06]. multi-response [ZAVS09]. Multi-scale [yCyZ08, yCyZ09]. multi-section [KMB09a]. multi-species [WF09]. Multi-stage [AG09]. multi-step [BW07b, DT06a, FJT07, Sic08]. Multi-synchronization [SS09a]. multi-term [FC09]. multi-valued [Kon07]. multiband [NQDJ06]. multibit [TND^{+}09]. multibody [Aue07, BBD06, Dmi08]. multicommodity [KCK09]. Multiderivative [SS05]. Multidimensional [MMS^{+}09, CZ08a, DD08, MSTT08, PW08]. multidomain [GW06b, JKM07]. Multifilters [CLS08]. multifunctions [CY08a]. Multigrid [KDH08, MA09, Sea07, Bor07, BSKS07, CCCH09, De 08, VN09, XZS07, WT07]. multigrid-FEM [WT07]. multilevel [MS09c]. multimedia [CGRGVP08]. multiobjective [Ant09, JK09, UFG09]. Multiple [CHSW08, GJL08, Hon08, Lan06, MD09a, PV05, Won05, BJSS06, BKT07, CJP08, CSG05, CMRS06, CV05b, DESC07, Eig07, Est07b, GK09, GQ09a, HWSL05, Hom09, LZW09, MLH08, MS06a, Röb08, Tan06, TYH08, Toy07, WYO09, Wik08, YB05, YW08, ZC08a]. multiple-image [DESC07, Est07b]. multiple-objective [YW08]. multiplication [CT09a, PS09b]. Multiplicative [Li08, CG08a, ZW09d]. Multiplicity [HC08, YOA06]. multipliers [BG07a, BNPR08, Tan06, ZH09]. multiply [BDHW07, WN08]. multipole [DM07b, OMC08]. multiquadric [CHW09, LT06b, MW09b]. Multirate [SKAW09, Sav08, Sav09]. Multiresolution [GP08, ABT06, CP09c, KJ09, TS08, ZYLY09]. Multiscale [LCZ09, CHXL06, LG07, MS09c, VR08]. multiscale-based [MS09c]. multisplitting [CHJZ09, Don09, Yun05a, Yun08a, ZHG^{+}09]. multistable [DBE07]. multistep [AT07, ASL07, Gar09, Mar07a, TYZ^{+}05, VAR06]. Multivariate [GNS06, Mat07a, PR08a, Ple09, Sau08, BI08b, KR08, LW06c, WLL06, Wan08b, WZD09, dMFPP08]. multiwave [PS07b]. multiwavelet [BP07]. Müntz [MC05c, MDSR04]. mutual [WWWF09]. Myshkis [CCH06].

N [Sto05b]. Naik [AG09]. nanocrystalline [WZH06]. nanoscale [LY05a]. Nash [PPV05, PLGS09]. natural [BBM07, EKEHR08, LS09, LYT07, RK08a]. nature [TH07]. Navier [Dai11, DC08, DSC09, DMZ08, FLPA09, FH09c, GRW08, HH07a, HL09b, KDN08, Mot06, NHN07, SS08, VVE08a, YWD09, YXYJ06, ZH09]. NCP [HHC09, MC08, ZZ09a]. NCP-function [MC08]. NCP-functions [HHC09]. NCPs [CP08b]. Near [BISS05a, BISS05b, KS07c]. Near-best [BISS05a]. nearly [EJ07, Scu09, Toy07]. nearness [YD08]. Necessary [BGG07, HW07, LW08c]. negative [BM06, DL09a, LVV05, Li07a, LZZ09, SZ08a, SZ09b, YW09a]. negatively [BS09b]. Nernst [Hsu05]. nested [RV05, Wac05, vdESvG05]. nested-grid [Wac05]. network [CFGVP07, DA09, HKS07, KCK09, Lin09b, SC06a, Yan06a, YHH05]. Networks [CM08, BNP07, CdAHBR06, CHBMBR07, FL09c, GEK09, HBS^{+}07, LMZ08, LZY09, LY09, Liu07, LH07a, LGYH09, LS06c, LC08b, MLL09, Moh08, Sha09b, Sin08, SC07b, Ste07, TBBC07, TXZ09, XLHZ05, YSYX08, YY07c, ZG09, ZW09b]. Neumann [CV09, Dan06, HL07b, Hem06, HR09b, Jai06, JLNP06, Pie09, SFFS08, SFPS09, WMN05, WN08]. Neumann-type [Pie09]. Neural [CM08, FL09c, LMZ08, LZY09, Liu07, LH07a, LGYH09, LS06c, LC08b, MLL09, Moh08, Sha09b, Sin08, SC06a, SC07b, TXZ09, XLHZ05, YSYX08, YY07c, YHH05, ZG09, ZZZ09, ZW09b]. neuron [Yan06a]. neuronic [HCS05]. Neurospora [Nag09]. neutral [BBPR09, BG09, HR09a, JL06a, KTM09, LmZbL07, LWMT07, LB09, LB10, MLL09, MY08, PP05, Sak06, WWC05, WTL05, WLS08, WLS09, WL07, WH09b, XZL09, XW07b, XL07, YC05, ZW09c]. neutral-type [MLL09]. Nevanlinna [GLV05]. Newmark [ZP06]. Newton [AML07, Arg07, AR09c, Arg09a, Arg09b, BK08, CR09, CL06a, CP08b, CL09b, Chu08, DC08, DH09, DFZ07, Dub09, Dus05, EEGG08, EH06, FHH08, FH09a, GK09, GPRB05, GS09, HCL08a, HR05a, HR05b, Hom05, Hom09, Hsu05, Hu05b, HMT09, JYX07, Kaw07, KLW07a, KV09a, hLyHZ08, MC08, MNG09, PG07, Sch07b, Sea07, Śmi07, WZD09, WCY09, XZZ09, YOY07, ZJW08, ZZ09a, ZBL08, Zhu06]. Newton-like [Arg09a, HR05b, PG07, Sch07b]. Newton-penalty [Dus05]. Newton-secant [GK09]. Newton-type [CR09, FHH08, GPRB05, Hom05, Hom09, ZJW08]. Newtonian [GW08, ICW^{+}09, Svá08b, ZZL08]. Nicholson [LF07a, LD08]. Nicolson [PP09, WKY^{+}07, YWD09]. nine [DL09b, GSBB09]. nine-point [GSBB09]. Njåstad [Ano05-38, Lor05]. NLS [CX07]. no [Yun08b]. Nodal [BC08, Dmi08, Sie07b]. node [LW06b, LZ09a, LLD09]. nodes [BDGV09, Mil05, MS06a, dB05a]. Noether [JKM09]. Noether-type [JKM09]. Noise [FN09, BS07, CX06, Toc09]. Noise-induced [FN09]. noisy [BFGP08, KP08b]. Non [ARVP09, BS09b, Jay08, SFV09, UO09, ABBM09, AKL05, AR09c, BR07b, Bou07a, Cha07, CC07, DS09a, Dao08, DG08, DFLV06, DHS09, EZ07, Fan09, FIR06, GGT07, Gia06, GHT06, HSN09, HyLC06, ICW^{+}09, KK09c, Kom07b, KMB09b, LC06b, LvGJ06, LP06a, MYY07, MK09, MY08, Moh05, MS07c, MS07d, MS05b, RW06, SJZY09, SZ08b, SDV07, Svá08b, Swe07a, XSW09, Yan06d, YG09a, ZZL08, ZCL09]. non-algebraic [GGT07]. non-autonomous [LC06b]. non-classical [ZCL09]. non-commutative [Kom07b]. non-conforming [CC07, GHT06, MK09]. non-convex [Fan09, KK09c, XSW09]. non-critical [MY08]. Non-differentiable [Jay08, AR09c]. Non-existence [SFV09]. non-Fréchet [ABBM09]. non-Gaussian [HSN09]. non-Gaussianity [SJZY09]. non-Hermitian [KMB09b, LvGJ06]. non-integer [MS05b]. non-linear [DG08, Moh05, MS07c, SZ08b]. non-Lipschitz [MYY07]. non-living [Gia06]. non-local [Dao08, EZ07, RW06]. non-monotone [HyLC06]. Non-negatively [BS09b]. non-Newtonian [ICW^{+}09, Svá08b, ZZL08]. non-overlapping [Bou07a]. non-penetration [DHS09]. non-polygonal [SDV07]. non-self-adjoint [BR07b, Cha07, MS07d]. Non-separable [UO09]. non-singular [Yan06d]. non-smooth [DFLV06]. Non-standard [ARVP09, AKL05, LP06a]. non-stationary [DS09a]. non-uniformities [FIR06]. nonautonomous [Che05, GOT06, LZY09, LOZ08, SP06b, TR06, WLF09, XW06]. noncoercive [NH08]. nonconforming [AR09b, CYM08]. nonconstant [Bra07a]. nonconvex [DFZ07, YG07, ZZ09c, Zhu08b]. nondifferentiable [AHS08, Ant09, JK09, LW08c, MPR07]. nonequispaced [PT08c]. nonexpansive [CS07e, CQ09, PK09b, QCK09a, QCKK09, Tem09, Thi09]. nonextensive [Tsa09]. nonhomogeneous [AS05d, GNP09]. nonincreasing [MLY07]. nonisolated [SCM09]. Nonlinear [ABT06, CS07d, GF08, MJB09, Mil05, Pok06, Slo08, WLS08, WLS09, Xu07b, AETE07a, AEET09, Abb07a, AT09a, eMH06a, eMH06b, AEBEK09, ADMM07, AS07, AB05, AMV06, Arg09a, AB09b, BS09a, BLW06, BFT07, Bat08, BRW09, BH07b, BK09b, BF09, Bor07, BMC09, BK08, BSS09, CC09, CK09a, CS07c, CYZ09, Che05, CXS06, CGP09, CL09a, CYB09, CXH09, CHJL08, CMT09, DS09b, Dem07a, DA08, Du07b, EGO06, EGFOA09, EPS05, FWY09, Fan09, FWZ09, FPS08, FH09a, FI07, FD09, FKM05, FOP05, FL08b, GGG09, Gan07, GS07a, GJM07, GPP07, GXgL05, Gei09, GM07b, GZX09, HCL08a, HY09a, HJK09, Han07, HK07b, How09b, HL07c, HR09b, HyLC06, HM09b, HMT09, IO06, IS06, JDG06, Jan09, JG09, bJQlZ07, JM07b, KYF05a, KYF05b, KK06, KCS06, KRAH07, KR07, Koç08a, Koç08b, KHWF09, LWW08, yL08]. nonlinear [LL05, LL08, LSTS06, Lew06, LHW07, lLpZ08, Li09a, LZ08a, LZ08b, LZ09b, Lia05, LKSC06, zLCgS06, Liu08b, LYB09, LB09, LLW09, Liu09c, LB10, LRT06, LD06, LOA05a, LOA05b, LNL09, Luo06b, MC07, Ma08a, MV08, MO07a, Mat07b, Mat08, MY06, MJ07, MY08, MA08b, Min05, MK07, MO08, Ned08, NB06, OW06a, Pet09a, PR08b, Rob07, RKK09, SK05, Sak06, SS09b, Sha09a, SJ09, SJZY09, SJZ09, SV07a, SA07b, ST07, SP09b, SL08, Swe07b, TGJ19, TRST09, Uje07, UFG09, VW06, WMG08, WK07a, WLL07, Wan07b, WG08, WZ08, WHL08, WNZ09, WH09a, WYL09b, WWFV07, WWZY09, XZL09, Xu07a, XSP09, YY07a, YZ09c, YM07, Yas07, YW09b, YLW09a, Zha05, ZZ06a, ZJW08, Zha09b, ZJ12, Zho09a, Zhu05a, Zhu05c, Zhu05b, ZF09, Zhu08d]. nonlinearities [Cao08, HWSL05, LOA05b]. nonlinearity [IL05, SS09b]. nonlinearly [bJxC09, ZA09]. Nonlocal [CJCD07, Du07b, MJX09, SD09]. nonmonotone [DEMR08, LYW08, MLY07, QZZ08, QZY09, SS06a, SP09b, WLY05, XSW09, YP08, ZZ09c]. Nonnegative [OS07a, HYG08b]. Nonoscillation [EBP09, JL06b, Xu08]. nonparametric [MJB09]. nonperiodic [BG07b]. nonreflecting [YS06]. Nonself [KN06, Thi09, ZTY07]. Nonself-adjoint [KN06]. nonself-mappings [Thi09, ZTY07]. nonsingular [VMV07]. Nonsmooth [BGRS06, AZGN06, Arg07, BGP09, KMVWY09, MW09a, PV09a, WTZ08, ZCH07, ZXZ09]. nonstandard [DK06b]. Nonstationary [CL09a, CGP07, Dol08, GP08, Yun05a]. nonsymmetric [Gan08, GQ09a, GNP09, Guo06, JR08, KK08, RT08, RT09, RH06, SHC08, SSZ09]. nonsymmetrical [ZGV06]. Nontrivial [LLW09]. nonuniform [DL05, KGA08, KDN08]. nonuniformly [SLN06]. Nordsieck [Kul09]. norm [CV05a, GSS07, HW08, bJyKyZmT09, Mor07, WA08]. norm-relaxed [bJyKyZmT09]. normalized [Swa06]. normed [BISS05b, MM09]. norms [MYY06]. Norton [FZ08]. Note [AOSY05, CHS08, Du07a, SA07a, AD05, BE06b, CL06c, DT06a, DS05a, Gra06, GZ05, HP07, KN09, KP07, LP07b, Li05d, LC09b, Lu09b, Luo07, MS07b, MB07b, RVA05, RP09, SG05, SHB05, VV06, Yun05b, Yun08b, ZZ06c]. Notebooks [RVK05, RV05]. notion [SCS08]. Novel [PK09a, TXZ09, XZL09, HMH08, MO08, SRW09, SK06, XWZH08, ZYLY09, dS07a]. NPV [Hua07]. nuclear [LRV^{+}09]. nullity [VV06]. number [Kan05, LCH07, MS05b, SZ09a, SL08, Tan06, VVE08b, WD05, Wik08]. numbers [Ben07, CT09a, CMM09, EKL^{+}07, GGP09, KG09, NP06, NP07, SP09a, WD05, XDW05, XW07a]. Numeric [SL09, BBD06]. Numerical [AT09a, AGM06, ADF07, AM09b, AVMMS^{+}06, BS08a, BS09a, BCeAJ09, BDY09, BLW06, BSC07b, BG07d, BNP07, CFV06, Car06, CB07, Civ07, CP09b, DH06, DL08b, DS09b, dCDND06, DF09, DFLV06, DJ06, DFHS08, Eds09, ÉT08, FZ09, Fer09, FDFL07, Gao06, GRW08, GHT06, HO06, IO06, Ish07, JL06a, Jen07, Ko07, LL05, LL08, LLD09, LYB09, Lop07b, LRB05, MJX09, MDMRP09, MYY05, MD08, MCA09, MN07a, MO07b, MNMS05, Mus09, NHN07, NWW08, Nig07, PI06, PT08c, RVM08, Sha05, SA07b, SBO09, SDV06, Sun09, SSMB07, TGB08b, VRL07, VK08, WTZ08, XHK^{+}08, YJJ09, ZZZ06, ZLA07, Zha07, ZCL09, AEET09, ACR06, Arp07, ABV08, Att05, BXC09, BS05, BFT07, BSC09, BGP07, BF09, BSV08, BB08b, BB09b, BMM06, Bru08, Bru09, CCD09, CJB06, Cas07, Cas05, CMM09]. numerical [CHSW08, CC06b, DMM07, DSI06, DFFW06, DK06b, FFSS09, FL09b, Gug06, GGP09, Ham09, Has09a, HV06a, Hat09, HV06b, HY09b, HCK07, HLL08b, IT07, JVVvdZ06, Jay07, Kaw07, KAP07, KTM09, LN07a, Li05a, LkLkP09, LS09, Lia05, LDM09, LKSC06, LM09a, LHL09, MV05a, MT05, MG08, Mon07, MS08b, NST08, PK09a, Pet07a, Pre08a, QW08, RS05a, RDH07, RHQ06, RK08b, SK05, SS05, SHK^{+}08, SAD07, Sid05, Sid08, SFFS08, SLN06, uIHA09, SD09, SW06b, TRS07, UO09, VAN06b, Vil09, Wan07c, Win06, XC09, ZL05, Zha08, ZMH09, ZW09c]. Numerically [LCKK06, BKB06, CDM07b]. numerics [HZ05]. Numerov [AOSY05, FT08, HVV09b, Tsi06, VV07]. Numerov-type [FT08]. nutrient [Thi05]. Nyström [Van07a, AMCM05, AMCM06, Bar09, DL09a, Van06a].

objective [FY06, JRM09, LYW08, YW08, Zad09, ZW06a]. objects [DL07b]. Obrechkoff [DWW06, NBSV08]. observed [FvZ06, SB09]. observer [YJY06]. obstacle [BCK07, Hem06, KS07e, PS07b]. obstacles [Mar07c, Oga06, OA06]. obtain [LB07]. obtained [AG09]. obtaining [BCP07, OOM^{+}07]. occasion [VAW07]. ocean [BCK07]. octree [DL07b]. octree-encoded [DL07b]. odd [CS07a, HCK07, KV05, OZY05]. odd-order [OZY05]. oddoids [HCK07]. ODE [BM07b, HKS07, NBSV08, NMM07, OZL09, SW06b, Wil07]. ODE-based [OZL09]. ODE/DAE [SW06b]. ODEs [Arp07, DS06, ED06, Hos09, HZ05, Jay07, Sav09, Sha07]. off [NSK09, VRBF07]. offset [LCKK06]. Oharu [VAW07]. oil [FGRW09]. Olav [Ano05-38, Lor05]. Old [AT07, MMFMG07]. Oldroyd [HKA07]. Olever [SWZ09]. One [Lar08, Müh06b, Roi07, CL09b, Cof05, DW05, EZ07, EHR09, GK07, HK07a, Hu05b, JVVvdZ06, JW08b, KGA08, Kho09, LOA05a, MC08, MG08, Osa07, ST05, Sen07, TTZ06, Vil09, WO08, Wu08, XHK^{+}08, YYF09, YG09b, ZCS06]. one- [TTZ06]. one-dimensional [Cof05, HK07a, JVVvdZ06, JW08b, KGA08, LOA05a, MG08, ST05, XHK^{+}08, YYF09]. one-model [Kho09]. one-point [EHR09]. One-step [Lar08, Roi07, CL09b, MC08, Vil09]. one-way [EZ07, Sen07]. only [CDMR07, Fin08, Gao05, Hu05b]. Onsager [sR07]. Open [Ano05-39, Gug06, CJCD06, EE06, GEK09, MMFMG07, MM06, PV09a]. open-channel [MM06]. operational [AT09a, BM08a, BS08a, DMS05, Hos09]. Operator [Gei09, GGP09, WLMS08, BJ06, Cho05, DSC09, Dai11, Djo07, FL09a, Gei08, HHST05, Kam07, Koz08, LN09, LD06, MJ08b, MS07d, Sch06a, Sha09a, SBO09, Vin09, WHL08, WH09a, Xu05a]. operator-differential [Vin09]. Operator-splitting [Gei09, DSC09, Dai11, Gei08]. Operators [EL07, AGM06, AS05c, AHL^{+}05, AU09, ABBM09, Arg07, Ass09a, BNT08, Bad08, BGG07, CY08c, CS07c, CC05, CN08, CW05, DF08, EFS05, EMMP07, FD09, GP07, GPS08, IS09, KI08, KN06, MJ08b, MYY06, PZ08, RS07b, Swa06, WL06a, Xu05b]. opposite [DCZ07]. OPSFA09 [VGT09]. optics [Tor08]. Optimal [Ata07, COEV05, CHNZ08, CCS05, DFY09, FW09, HW08, HLJ09, KS05b, Liu09c, SAHS09, VCD^{+}08, WCY09, YW08, ABD^{+}05, AZGN06, AVMVMV08, BBD06, Bor07, BSV08, BG07e, CJCD06, CM07, CY09, CL08, Chr09, CCK06, Gan08, GS06a, GHZ06, GL06, GF08, Hei05, HH07b, HVV09b, HYG08a, KV09b, LP07b, LFS08, LFW09, LHW08, Mar08, NT06, Pul09, Rie06, Saa06, SCM09, TLQ08, WTZ08, YZ09a]. Optimality [Ant09, YW09a, CG08a, Mar07b]. Optimally [Fer07]. optimisation [McC05]. Optimization [BC06, WWP09, Wan09b, ABÇ06, AVMMS^{+}06, And09b, AT09c, BRVWM07, BGRS05, BGRS06, BSMT09, BGP09, BG07e, BNDN09, BT09, CL06b, DEMR08, DS05b, DMZ08, DDL09, EL06, EIM^{+}09, FGB07, GAC^{+}08, GZX09, HL08, HCL08b, JTHZ05, bJyKyZmT09, bJxC09, KMB09a, KCLW08, KGB^{+}07, LMR07, LTW07, LTY08, LZLH07, Liu05, LW08a, MZW06, OZL09, QZZ08, QZY09, lSsZ05, SaLJ08, SW08, SG08, Ste07, SP09b, Suz05, TF07, TH07, UFG09, VL06, WCH05, WZ07, WFW09, WYL09a, WOK05, XZ08, YW08, YSY09, YY07b, YP08, YLW09b, Zad09, ZW06a, ZW08, ZZ08a, ZXZ09, Zha09b, ZZ09b, Zhu05b, Zhu08c, ZTG09, ZA09]. Optimized [FNW06, AS05f, KM09c, LM08]. optimum [Tok07]. option [HLJ09, Hug06, LÅH08, NST08, PLMP08, VV09a, ZDC07]. options [BF08a, BS08b, CP09a, DDV08, FL08a, KVK08, LO08, LWH09b, Pir09, Sch06b, TGB08a, TGB08b, VDL^{+}06, WKY^{+}07]. orbit [SAFS07]. orbital [AS05f]. orbits [ABP06, ABB^{+}07c, KW07, PRL09]. Order [AMCM05, LZG^{+}09, AOSY05, AHS08, ABG06b, ABP06, ABB^{+}07c, AS05e, AGH06, ALR05, ADDdM05, BNT08, BJ07a, BFT07, BM06, BRW09, Bor07, BM07b, COEV05, CS06a, CS07a, CFMR09, CMST06, Cen08, CHCW05, CS07d, CP08a, Che08, CL09b, CCS05, Chu08, CH09, CMT09, CD07a, CDL05, CN08, DSC09, Dai11, DF09, DZ09b, DD08, DO06, Dor09, DLG05, Dub09, DG05, Dus05, ED06, EGO06, EGFOA09, EPS05, EBP09, EH06, FT08, FWY09, FWZ09, FH09a, FJT07, FPP07, FS08, Fin07, FOP05, Fra06, FH09c, Gal09, GGG09, Gei08, GO05, GPPRRB06, GPPRB07, GOT06, GS09, GZX09, GX09, GGA08, HCL08a, Has06, HYYF05, HAS08, HRGD08, HR05b, HM09b, Huy09, Ish07, JKM07, Jai06, JG09, Jia09a, JYX07, Jon06, KCK06, Kan08, Kar07, Kar09, KT07a, Kil07, KKLY06, Koç08a, Koç08b, Kom07a, Kom07b]. order [Kom08, KHWF09, KLW07a, KLW07b, Kou08, KL09, KBB05, LT07, LT05, LG07, Lei08a, LLZY06, LS07b, LW07, LTY08, LBS09, LS07c, LZ09c, LJL06, Liu08b, LB09, LLW09, LM09b, Liu09d, LB10, Log08, Lu05, LJ09, Luo06a, xMzY09, MP07, MD09a, MY06, MY08, Min05, MO07b, NWW08, NMKS09, NBSV08, NM06, NMMS05, NMM07, OZY05, PS09a, Ped05, Ped09, PM05, PRP08, PJ08, PS05, QFX06, Rad05, RVA07a, RVA07b, RB08, RVAD08, SS05, Sak05, Sak06, Sal09, ST05, SS09b, SWZ05, SZ08a, SZ09b, SA08a, SF08, SS09d, Sun08, SLZA09, SS06c, Sut07, Sza08, TRS07, TYZ^{+}05, TGB08a, TGB08b, TG07, Toc05, Toc09, Tsi06, VVE08a, VMMM09, VMV07, Ver06, WKY^{+}07, WG06, WTL05, Wan07c, Wan07b, WWC07, WG08, WML08, WYS09, WZ09a, WSZ09, Wei09, WS09b, Wri08, XW07b, Xu08, YLSX06, Yan06b, YS07]. order [YY07a, YLN09, YM07, YL09b, YSYX08, YW09b, YCJ09, ZCS06, ZCL06, ZWZ08, ZFG08, ZG09, ZLL09, ZW09b, ZLL08, Zhu08d]. order- [Kil07]. ordered [yPlH05, YG09a]. Ordering [MPZ09, ZGSF08]. orders [RP05]. ordinary [BW07b, CHCW05, HO06, Ish07, Lop07b, NMMS05, RS05c, TRS07, Vil09]. oriented [KMB09a, Kor06]. originated [GP08]. Orr [MCA09]. Orthogonal [BGVHN05, DP05a, DMGVP05, Ell06, Gau05c, MC05b, Pet08, RAG07, AÁN07, AB09a, ARN07, AGMMB05, AGRZ06, Ask05, BS08a, BDY09, BMS05, BS05, BBR05, BCGA05b, BCGA05a, BCV^{+}05, BDGV05, CJB05, Car05b, CSSL09, CV05b, CBGV05, CBDGVN07a, CBDGVN07b, DSMY06, DOY08, DM05, DG05, DI06, FPP05a, FPP07, Gau05a, GLV05, Gri09, GS08, HHR06, Ism05c, IS09, JT09b, Khr05, KMJ06, Kra06, KT09, LYZ09, MP08a, MS07a, MJ08a, MOC09, MB05, MB07b, Pet07b, Pie08, RS05b, SRD05, VB05a, VB05b, Van08, Vol08, dMFPP07, dMFPP08, dMFPP09, vDvFZ09, vDvFZ13]. Orthogonality [DSZ05, MDSR04, MC05c, Sza05b, CLO09, CSSL09, CBDGVN07a, CBDGVN07b, DS07b, Kok05, CZ07, Ism05b]. orthogonalization [ARN07, VS08, YH09a]. orthonormal [Liu06]. oscillated [GDA07]. oscillating [Gau05a, PS05, ZZL08]. Oscillation [GX09, HM09b, JL06b, LLZY06, LM09b, OZY05, Sak05, Sak06, SZ08a, SZ09b, WTL05, WWC07, XW07b, Xu08, XX09, AK07, BV06, CHY05, Li05b, Liu09a, LB09, LB10, MY06, MY08, WWC05, Wan07b, WML08, XL07, ZW09b, Zhu08d]. oscillator [BSS09, JW08a]. oscillators [FSW09, Fra05, MDP08, OW06a, Van09, Xu07a, ZZZ09]. Oscillatory [HCK07, LZG^{+}09, DS06, GLS07, HC09, HV06c, KC07, MC05b, WLMS08, Xia07a, Xia07b, XW08, YZ09c]. osculatory [HT06]. Oseen [TY05]. Ostrowski [KLW07b]. other [Art09, FH07, Ter05]. outer [AK09, PK07]. outlines [SRB06]. output [CX08, LC08b, SS09a, XXL08]. overcomplete [WC08]. overestimation [RKAH07]. Overlapping [Dao07, Bou07a]. Oversampling [GHMPV09]. overview [Sch06b].

P. [AGRZ06]. P2Q2Iso2D [BCH06]. package [GS07b]. packet [HV06c]. packing [Mar07b]. Padé [AETE07a, Boy09, DMGVP05, EZ07, GPTT05, GS08, KR08, KSVW07, MV05b, Mat07a, Sab08, TG09, YB05]. Padé-type [GS08, TG09]. Padua [CDV08]. PageRank [DGR07, LSW09, Wu08]. Painlevé [Cla05, Nou07]. pair [FSW09, GM07a, WHG07]. pairing [Cou09]. pairs [DM05, Dun07]. Pál [dB05c, dB05b, dB05a]. Pál-type [dB05c, dB05b, dB05a]. palindromic [CHLW08]. pantograph [FSL09, SAD07, SyS08]. paper [Dom07, Mil07]. para [BDGV05]. para-orthogonal [BDGV05]. Parabolic [Jon06, AZGN06, ABS07, BUG07, Bog05b, BJ06, CC09, CJCD06, CL06b, CYB09, Chr09, CP09b, DAE05, DBG07, DBR08, Gei08, GKMS09, GOT06, HZBM06, Han07, HH07b, KD05, KV07, Li05b, LDM09, zLCgS06, LD06, MW07, MS07c, OZY05, RKK09, Sch07b, SD09, SF06, YYF09, ZCL09, Zhu07]. paraboloid [PH09]. Parallel [Sch07b, SCY09, Suz05, BJ08, BGP07, CC06a, CC08, CCD09, GAC^{+}08, GZLL09, HE06b, LY05a, MBCV09, MGS08, NS09a, PKS^{+}09, VB05c, WW09, Zhu07]. parallel-series [WW09]. Parameter [DOY08, GSW07, MV08, MDT09, NGS05, Wil07, ABS07, BBF^{+}05, BJK09, Cha07, Cha08, Daa09b, DA09, Dao08, GK09, GG09b, Han06, JKM09, LCKK06, LC08b, MJB09, Min05, MvdMV06, OS07c, Osa07, RS07b, SM09, WK07a, WS09b, Xu07a, ZAVS09]. Parameter-based [DOY08]. parameter-controlled [GK09]. Parameter-dependent [GSW07, Cha07]. parameter-expanding [Xu07a]. parameter-robust [ABS07]. parameter-uniform [OS07c, SM09]. Parameterisation [GMFB06]. parameterization [GL06, HAS09b, KY06]. parameterizations [SFV09]. parameterized [AD05, Cen08, Wan09b, XWZH08]. parameters [AR06, CB07, CHJZ09, EL06, ER07, GVC^{+}08, GH06, HSN09, Hua07, HHZ07, KY07, Par08b, Rie06, Sen07, TR07, TR08, WL07]. Parametric [Din05a, PH09, CLP08, CMS08, FR06, KY06, LBS09, LCZ^{+}08, PK07]. parametrization [JH08b]. paraxial [AT09d]. Pareto [BdWG06, UFG09]. Pareto-type [BdWG06]. Parshin [CCLS06]. part [AT09a, Har06, KAC05, Mon09, ZWY09, CH09]. Partial [Jac06, AETE07a, AMCM05, ABS07, CL09a, DAE05, EMMP07, FPP07, FFSS09, GGD08, Gei08, GX09, HO06, IO06, JKM09, KSS07a, LDM09, Luo08, Mat08, MA08b, MO08, RAG07, SDV06, SCY09, VS08, WWC07, WML08, YZ09c]. partially [CCM07]. Participants [Ano05-40, Ano05n]. Particle [AFT^{+}06, AFT^{+}07, LW08a, NSK07, WO08, WZS06, YW08]. particles [Göt05, JKM07, LmLgF09, MM06]. partition [NQDJ06]. partitioned [TD06, VDV08]. partitioning [WOK05]. partitions [BISS05b, PR05, ZW06b]. partly [LZ05]. parts [CJ07]. Pascal [LHR09]. passage [ST05, SS08]. patchwork [Bak05]. Path [Sta07, KK09b, Lin09b]. Path-following [Sta07]. pathology [Wod05]. pathways [ZBB^{+}05]. Pattern [BGP09, yCyZ08, yCyZ09]. patterned [NP07]. payments [ADGH06]. PC [CX08]. PCG [KK08]. PCR [MMS^{+}09]. PDE [BG07e, BKB06, JKK06, LCZ^{+}08, Min05, PR08a, PW08, RW06, SRW09, Ste07, VK08, WWFV07]. PDE-based [JKK06, PW08]. PDE-constrained [BG07e]. PDEs [BF05, CC09, Har06, JKM09, Lui09, PRGPS09, SSV06, Sha05, SFFS08, SFPS09, VMMM09]. PDIEs [HR09b, RLH09]. Peaceman [PJ06b]. peak [WFW09]. peakon [TY07a]. peaks [GMFB06]. Pearcey [PK06].

Pearson [NK08]. Pearson-type [NK08]. peer [WSPJ09]. Pell [Kil07]. Penalty [NST08, BIR09, CG08a, Dus05, ER07, FLPA09, GSS07, GNP09, GBR09, LXSH09, LHW08, XSP09, ZLY09]. penalty-function-free [XSP09]. penalty-projection [FLPA09]. pencil [KV09a]. pencils [CG08b]. pendulum [CBH07]. penetration [DHS09]. Penrose [TA09]. perceptions [AS09a]. perfect [eMH05]. perfectly [ZKSS06]. performance [CCD09, Eds09, LWH09a, MV05a, Oou08, RDH07, Rom09, Val09]. performances [DL05, WY07, WYW08]. Period [AD09, LLS09, ZZ09b]. Periodic [CX07, FWZ09, IL05, LS06b, LS07c, LTM08, Liu08a, Liu09d, ML09, SWYZ09, SP06b, ST07, WSM^{+}09, XW06, YS07, ABP06, ABB^{+}07c, BK09b, COEV05, CJCD07, CCCH09, CXS06, CMRS06, Coj09, DWW06, DC09b, FWY09, HCS05, Har05, HWY09, LW05b, LH05a, LH05b, LC06a, LF07a, LW07, LD08, LH08b, LC09a, LC06b, Liu07, LH07a, Liu09a, LJ09, MN05, MLH08, NPTH09, Oga06, OA06, PRL09, PP05, RR06, SG05, Sha09b, SO05, TSC10, WW07, WS09b, WL07, YC05, YG09a, YLN09, ZCS06, ZG09, ZT07, ZL07, ZSZ09, Zho09b]. Periodic-like [CX07]. Periodical [XW05]. periodicity [LCH09]. periodizing [Sid08]. periods [ZCZ08]. perishable [Siv09]. Permanence [CS07b, CT05, LCH09, LOZ08, LL07a, SMT07, TCMS06, WJS^{+}09]. Permutation [KS06]. permutations [VC06]. Perron [AAZ06, HYG08b]. Persistence [KBB05, TR06]. perspective [Cal07, dMFPP08]. Perturbation [WCW09, AHM08, AD05, BF08b, Cen08, CL06c, DC09b, ED06, FHM09, GS07a, Hoo08, Jia06, Li08, Li09a, LP06a, Tem07, VAN06b, Vul07, XDW05, Yan06b]. perturbations [Coj09, DW05, Du07a, WZ09b, XZL09]. Perturbed [LS07a, ABS07, Att05, FSW09, Fra05, GY08, GC06, HR09a, KGA08, KM09c, LC06a, Lin05, LS05, MM07, MSE08, MN07a, MS07c, OS07c, RO05, Sto05a, Sto08, TRS07, TR07, TR08, Van09, XWZH08, YZ09b, ZC08b]. Petrovsky [MDMRP09]. phage [BLR07]. Phase [Van07b, Ahm07, DK06a, ÉT08, ER09, HS07b, HCK07, LW07, Min05, SS05, TTZ06, TK07, WT07, WZH06]. phase-field [TTZ06]. Phase-fitted [Van07b]. phase-lag [LW07, SS05]. phase-transforming [DK06a]. phases [CKT09]. phenomenon [BMW08]. phi [MPP07]. phi-divergence [MPP07]. Philos [XL07]. Philos-type [XL07]. photogravitational [PRL09]. Photos [Ano07o]. Physical [HHZ07, Lai09a, PS07b]. phytoplankton [Thi05, THS05]. phytoplankton-nutrient [Thi05]. PIC [AT09d]. Pick [Ped05]. PIDEs [May08]. Piecewise [Han06, ABET08, And08, BS08a, BdAR08, ED06, EMP08, GRW08, HV05, KD06, LY09, Müh06b, Wan08b, YLN09, ZW06a, ZW08]. piecewise-linear [ZW06a, ZW08]. pile [MD08]. pipe [BeMBH06]. pipes [BG07c, BGG08]. PIRK [CX08]. PIRK-type [CX08]. Piscounov [MDMRP09]. Pitaevskii [Wan07a]. Pitfalls [DFFW06]. Pitolli [PS08]. pivoting [CP07]. pixel [BBM07]. Planar [GM07a, Rho05, AD09, DL09b, EM07, FH07, MW08, Oga06, SRB06, ZCS06]. Planck [Hsu05, WO08]. Plane [PD06, CG08a, DSMY06, LWYT06, Li06b, Lu09a, MH08]. planetary [LB07]. plankton [CLCL07, LC06b]. planning [LLS09]. plant [KOS07]. plasma [CC06b, Dem07a, LmLgF09]. plastic [AYK09, KBSK08, NT08]. plasticity [LLM08]. plate [Bou07b, CYM08, HeM05, Kas06, KS05b, PSV08, RK08a]. plate-membrane [KS05b]. plates [CC07, HGS08, XKF06]. plotting [SLMW06]. plus [VMV07]. PML [HK07a].

Poincaré [Gin05, SSMB07]. Point [FR06, PPR07, AHC05, Bac09, BS09b, CM05b, CMST06, CY08c, CY08b, CAHAY09, CHCW05, CS07e, CWW08, Cho08, Cim09, CN07, CG07, DS09a, DMGVP05, DLG05, EST07a, EHR09, Fan09, Fin07, FKM05, GPTT05, GSBB09, GJL08, HY09a, Han08, HJK09, He05, HL09a, HVV09b, Hon07, Hu05b, HWL09, Jan08, JG09, Jia09a, JC09, JR08, Kar07, KK09a, LP07b, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LLW09, Lu07, xMzY09, MW08, MW09c, MD09a, MN07a, Moh05, Ned08, NB06, PVC06, QCK09b, SVC09, Sal09, SS09b, SHC08, SJZ09, Sun08, SLZA09, WCL08, Wan09b, WB09, Wei09, Won09b, WKL06, YN08, Yas07, Zha09a, ZCL06, ZGSF08, ZC08a]. Point-based [FR06]. points [Bla05, Boy09, CDV08, EFP06, HE06b, Huy09, KKLY06, KMJ06, KSS07b, NM06, Nke05, Par07a, Pet07b, Thi09, TMG09, WC08, Xia07a, Yan06b, ZTY07]. pointset [TK07]. Pointwise [DJ05, DD06, Pul09]. poised [CJ07, Zud05]. Poisson [CCJ07, CMRS06, EZ06, Hsu05, KK09c, KW09, LLL06, LT07, LC08c, MGV07, Nke05, SS06c, Wac05]. Poisson-type [KW09]. Pol [JW08a]. Pole [SHK^{+}08, AM09b, CMRS06]. Poles [Par07b, Bla05, BGVHN05, SHB05]. policy [CKT09, KL08a, WWP09]. Pollaczek [FLS08, ZZ06b]. pollutants [AMO07, AMO08]. polluted [AVMMS^{+}06]. pollution [GM06]. Pólya [DC09a]. polycrystalline [MPXK06]. polyethylene [HKB^{+}05]. polygamma [BHJ05, Cof05]. polygonal [BN08a, BFP09, GS06c, SDV07, Wat07, YG07, ZZ07, ZKSS06]. polygons [ACWL07, LLD09]. polyharmonic [Ros08]. polyhedra [ZW06a]. polyhedral [KK09c]. polyhexes [Rho05]. polyiamonds [Rho05]. Polynomial [CD07b, Gin05, WB09, BB07, Cao08, CDL05, EIM^{+}09, FL09a, FPO05, GGT07, GGM06, GS07c, Gri09, GS06b, Han05, Han06, HCK07, Ism05c, Jar09, KLO08, Las09, PIP07, PPR07, Pet09b, Pet08, RAG07, SLMW06, STB09, Sza05a, Tay08, VV09b, WT08, XW09, YA07, YWH05a, YWH05b]. polynomial-like [KLO08]. polynomial-time [EIM^{+}09]. polynomials [Abb07b, AÁN07, ÁN06, AB09a, AGMMB05, AGRZ06, Ask05, BS06, BMS05, BS05, BBR05, BCGA05b, BCGA05a, BG07d, BCV^{+}05, BDGV05, BI08b, Car05b, Cla05, CEF08, CDL05, CSSL09, CV05b, CS09b, CBGV05, CBDGVN07a, CBDGVN07b, DJ05, DSZ05, DLROY05, DSMY06, DOY08, DM05, DMGVP05, DC09a, DR09, DA06, Dom09, DJ07b, DJ09, DG05, DI06, Ell06, EKL^{+}07, FDC08, FPP05a, FPP07, FLS08, GM09a, Gau05a, Gau05c, GU08, GLV05, GDA07, GPS08, GS08, GNS06, HHR06, IS09, JT09b, Khr05, KK05, KA08, KMJ06, Koo07, KV05, Kra06, KM05, KT09, LW06a, Lin07, Lu09b, MP08a, MDT07, MMFMG07, MS07a, MDSR04, Mat06, MBA09, MOC09, MC05b, MC05c, MB05, MB07b, Müh06b, OS07a, Par09b, Pet05, Pet07b, Pet08, Pie08, PV05]. polynomials [sR07, RS05b, RK08b, SRD05, SL05, Sau07, Skr05, Sto08, Sza05b, Tor08, VZ08, Vol08, Win06, WA08, Wün05, ZZ06b, dMFPP07, dMFPP08, dMFPP09, vD05, vDvFZ09, vDvFZ13]. polyominoes [GZ05, Rho05]. Population [LJM09, DJ06, DJ07a, IV07, KOS07, KS09, LSTS06, LH05b, LkLkP09, Lin08, Meh05, MvdMV06, MK05, RW06, RHQ06, SFSK07, Siv09, SP06b, TCMS06, WZ06, WF09, Zha08]. populations [BLR07, MPZ09, Tac07]. poroelastic [BSKS07]. porous [BeMBH06, CMX09, CO05, Cui07, EKEHR08, GS07a, HAS08, Hoo08, MP06a, SY09]. porous-saturated [Hoo08]. Portfolio [CHS07, QLJ09, CT09b, Hua08, HO09, JRM09, Lan08a, LZWQ09, Mar08, VV09a, YJJ09]. Portfolios [Ver08]. posed [AR09b, BR07a, Cal07, LHW07, MRSZ06, MRSZ07, RS08b, WHL08, WH09a]. posedness [HL08]. posets [DK05]. position [YKS09]. Positive [BDGV05, DK06b, FJG08, HY09a, LH05b, LZZ09, LC06b, MLH08, SS09b, SLZA09, WW07, WS09b, XXL08, Zha09a, BW06, CXS06, CGP07, DL09c, GT09, GJL08, HJK09, He05, Hon07, HC08, Hon08, HS08c, Iva06, JG09, Jia09a, Ko07, Kar09, KMB09b, LW05b, LH05a, LF07a, LS07b, LD08, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LHL11, MD09a, Möl07, NM06, yPESlZ07, RT08, RT09, RH06, SHC08, SZ08a, SZ09b, Sun08, SL08, TV07, Thi05, Wei09, Yan06d, Yas07, ZFG08, ZC08a]. Positivity [HS08c, KS05a]. Positivity-preserving [HS08c]. positone [Zha09a]. possibilistic [CT09a, JRM09, SP09a]. posteriori [BG07a, CS07c, CL08, CN08, ED06, EPP^{+}09, FH09b, KS07d, Kor06, KHK08, LY07b, PIP07, YZ09a]. Postprocessing [ALR05]. potential [BE07, Cha08, CCCH09, FG09, GS07c, KKS^{+}07, XHW07, YZ09d]. potentially [BLR07]. potentials [AKLT09, BBB07, Ehr06, LC08c, TX07]. Poussin [MT05]. Powell [BFGP08, MB06, SDV06, SDV07, VB05c]. power [BIR09, GW08, HHST05, TLQ08, VT06]. power-law [BIR09, GW08]. powers [Ber07]. practical [LZ09d, ZS05]. pre [AA08, Ant08]. pre-processing [AA08]. precise [SAFS07]. precision [Par09c]. Preconditioned [Cal07, WS09a, WL09a, YH09a, Yun08b, ZSWL09, Deu07, Li05c, Li05d, tLxLlW07, LC09b, NKM08, WHF07, WtL09]. preconditioner [Deu09, GH08, HWL09, KN09, KT07b, MN07b, RH06, STB09]. Preconditioners [DBR08, BW06, BDY09, HL07a, LRT06, Smo06, SCY09, Wan09b]. Preconditioning [CMX09, EM08b, BGP07, CNN07, MK09, VS08, VVE08b]. predation [YTM07]. predator [CS07b, Che05, CT05, DK06b, DC09b, FL06a, JL07a, Kar06, LCH09, LTC06, LZZD09, NPTH09, TR06, TSC10, WLF09, XW06, XGM09, YC06]. prediction [Fu08, Han08, MGMTO06, SS08]. Predictor [Din05b, Den07b, PS05, VW06]. Predictor-corrector [Din05b]. Preface [AHS07, ANR07, BRS09, BDF^{+}09, BCS^{+}06, BCP05, BLL07, CDZ06, CDG07, DM07a, DVD^{+}08, GHK^{+}09, GSL07, GVW06, Goo08, KMS09, LQW08, LRW05, MP06b, PRV09, RS07a, SP05, SP06a, Sol08, TMMS07, VA06, WL06b, zCEF07]. prefilters [CLS08]. preinvex [LW08d]. preliminary [MPP09]. prescribed [BDGV09, yPlH05]. presence [DDLM07, IV07, SdJMH09]. present [ADGH06, Spr06]. preservation [CFMR08, Moh08]. preserved [DV06a, DV06b]. preserving [ADD^{+}08, CHW09, CFL05, CHLW08, Cim09, EZ06, FL09a, FO05, FOT07, GG09b, HS08c, IT07, KL08b, LM07a, PJ06a, PS08, SLZ06]. pressure [Deu09, dCDND06, LL06, Mot06, TY05]. pressurised [BG07c, BGG08]. prey [CS07b, Che05, CXS06, DK06b, DC09b, FL06a, JL07a, Kar06, LCH09, LTC06, LZZD09, NPTH09, TR06, TSC10, WLF09, XW06, XGM09, YC06]. prey-chain [DC09b]. prey-competition [CXS06]. prey-predator [Kar06, LZZD09]. PrFeB [WZH06]. price [VDL^{+}06]. Prices [GM06, BSMT09]. Pricing [BBM09, BF08a, LWH09b, BS08b, CP09a, KVK08, LÅH08, LO08, May08, PLMP08, TGB08a, TGB08b, WKY^{+}07, WWFV07, ZDC07]. primal [WFL07]. primal-dual [WFL07]. principal [AV06, BP06, yPlH05]. principle [Den07b, HKV09, SV07b, Sur05, ZL09a]. Principles [Pet07a, KK06, KS06, PT09b]. priori [BG07a, BFP09, FH09b, KV09b, YZ09a]. prism [YLC09]. prismatic [HKV09]. privacy [KLS^{+}07]. Probabilistic [DS05c, AS05c]. probabilities [DD07, LP09, MPP09]. probability [Bin07, DIKZ07, GA09, KKLM06, PW08, ZLA07]. Problem [Ber05, Chi05, Ism05a, Ism05b, Xu05b, dB05c, AYK09, AT07, AOSY05, AH07, AR06, AS05d, AD05, ABP08, Bog05a, Bog05b, BH07b, BJ06, Bou07a, BNPR08, BFP09, BFM09, BGVHN08, CD05, CCM07, CFV06, CY09, Cen08, CS06b, CCJ07, CYZ09, CJ09a, CHCW05, CL06b, CCH06, CS07d, CP08a, CLA09, CU09b, Cop05, Cui07, CG07, Cui08, Dan06, DS08, DGR07, DLG05, DL09b, DMZ08, Ehr06, EEGG08, EM09, EMP08, FI07, FY06, FJG08, GKM07, GA09, GEK09, GHZ06, Gon06, GSBB09, GBR09, Ham09, HKV09, HLX06, HTS09, HAS09b, HH07a, HCL08b, HV06b, Her09, Hoc08, HLY09, Jia06, JZW08, KGA08, Kar07, Kar09, KCK09, KK09c, Koh08, Koy07, Koy09, Kru05, KM05, KKD06, LP07a, LLS09, LL08, LRV^{+}09, LS07b, xLLhZ08, LZWQ09, LS07c, Lia05, LDM09, LM09a, Lin09b, LF07b, LS05, LHW08]. problem [Mai08b, Mar07a, MP08b, MD09a, NT08, NMMS05, NMM07, OS07b, OW06b, PRL09, PP09, Pie08, Pul09, QFX06, QWS09, RW09, Sal09, SDP09, SZD09, SFSK07, Shi07, SO05, SS09e, Sun08, SS07, Tem07, TR07, TR08, TY05, Tur06, VRBF07, WYL08, WCY09, WB09, WMN05, WN08, Wu08, XZZ09, YM08, YZ09a, Yan06b, YJGL07, YXJG08, YW09a, YD09a, YKS09, ZJW08, ZLL09, Zho09a, ZA09, Zud09]. problems [eMH06b, AUY06, ABG06a, AV06, ADF07, AHC05, AZGN06, AABL08, ABÇ06, AS07, Alh06, AM09b, AR09b, AS05f, AGH06, ALR05, Ano05-39, AK09, Ant09, ABR07, Att05, AB09b, BN08a, BR07a, BMfX07, BBF07, BLW06, BGRS06, BTI08, BBD06, BR07b, Bi08a, BV06, BBB07, BGP09, BIR09, Bor07, BSV08, BMM06, BM07b, BG07f, BK08, BSS08, BK06, CG08a, Cal07, CJCD06, CM07, CFHM06, CKWZ07, Car06, CMST06, CY08b, CAHAY09, Cha07, Cha08, CC05, CW08b, CL08, CWW08, CGP09, Che09, CNN07, Cho08, Chr05, Chr09, CCK06, CHLW08, CW08d, DWW06, DKHNZ08, Dol08, Don09, DJ07a, DT06b, DBG07, DBR08, EE06, EM08b, EGFOA09, EPP^{+}09, FG09, FH09b, Fin07, FR07, FNW06, FOP05, FKL05, FPP05b, FPP08a, FPP08b, GAC^{+}08, GGG09, Gan08, GPP07, GMM09, Gei09, GKMS09]. problems [GGM06, GDA07, GPRB05, GPPRRB06, GPPRB07, GOT06, GL06, GC06, GZX09, GNP09, GJL08, HY09a, Han08, HV06a, HYYF05, He05, HK08, Hei05, HH07b, HCDP08, HK07b, HO06, HVV09b, Hon07, How09b, HHC09, HyLC06, HWL09, JLNP06, Jan08, Jar09, JVVvdZ06, JG09, bJQlZ07, JD05, JC09, JR08, KS07a, Kan08, KS07b, KK06, KOS07, KAP07, Koh07, KGW09, KA06, KV09b, KM09c, KLV09, KHK08, LN07a, Lan06, LMR07, LG07, LS06b, Li06b, Li07b, LHW07, LFHW09, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LKSC06, LJL06, LCA07, LT06b, LXX08, fLxXT08, LY07b, Liu08b, LLW09, Liu09d, Log08, Lu07, LP06a, MW09a, xMzY09, MDT07, MMM08, MV08, MSZ07, MMFMG07, MCR08, MK09, Moh05, Mon07, MRSZ06, MRSZ07, MA09, NHN07, NH08, NST08, NM06, NB06, OYA07, Oga06, OA06, OZL09, PS07a]. problems [PS09a, PK09a, PKS^{+}09, Ped09, PN06, Pre08a, QCK09b, RTGB09, RS05a, RS08a, Rea07, RS08b, RDH07, Roi07, RM07, RKK07, RKK09, Sch08, Sch07a, SHK^{+}08, Sch07b, SCM09, SA08a, SÁ08b, Sta07, Sto05a, SLZA09, TGJ19, TG07, TR09, TFC07, TY07b, TS05, VR08, VV05b, VAN06b, Vul07, WMG08, Wan07c, WWC07, WG08, WZ08, WGZ08, WYS09, WSZ09, WL09a, Won08, WLK06, WKL06, WTZ08, WWZY09, XZS07, XWZH08, YOA06, YYF09, YS07, YY07a, Yao09, YL09b, Yas07, YH09a, Yu06, YD08, YCJ09, ZAVS09, Zar09, Zha09a, Zha05, ZCL06, ZZZ08, ZW09a, Zha09b, ZC08a, ZLL08, Zho06, ZW09d, ZCL09, ZSWL09, ZC08b, ZKSS06]. procedure [UHNC09]. procedures [GRL09]. Proceedings [VGT09]. Procesi [ZCH07]. process [AHM09, AGM09, AMV06, ACQ09, BRVWM07, BS06, FvZ06, KRAH07, Kom08, LWS09, LBS09, MGV07, PPG07, RH06, SACJ09, Tem09, ZLLW08]. processes [AR09a, AI06, ACV05, Chr05, DGS06, EMP08, EHR09, GH06, Gia06, GLV05, GM09b, Gri09, HR09b, Kag09, KS06, LL08, LXSH09, Mar08, MSTT08, RKAH07, RLH09, YW09a]. processing [AA08, HT06]. produce [MS05b]. produced [CU09a]. Product [KMB09b, And09b, DP08a, Gar07a, GN09, Hoc08, KJ09, Lu09a, NK08, RS05b, Rza09, TG09, XDW05]. Product-type [KMB09b]. production [HKB^{+}05, LLS09, ZMAZ09]. products [FFK05, LWH09a, Ped05]. Professor [VAW07]. profiles [SBK07]. programming [AHS08, AH07, Ant08, Ant09, CW08c, CL09b, Fan09, Hua07, Jay08, JK09, LWYT06, LXUK06, LW08c, MW09a, MSA09, MPR07, PR08b, SaLJ08, SCM09, TJ08, TLQ08, VRSS08, Ver08, XZZ09, XSP09, Yu06, YXF07, ZMAZ09, ZW09d, Zhu05c, Zhu08b, Zhu08c, ZF09]. programs [CS09a]. progress [DW06b, Ela09a]. progressive [KG09]. progressively [LWH09a]. project [AG09]. Projected [BY09, Pul09, WLY05]. projection [Bno07a, CLN09, DD07, FLPA09, He06a, Vin09, ZZ06a]. projection-difference [Vin09]. projections [GU08, WS08]. projective [LG07, Par08b, Zhu05b]. projectors [ARN07, LMS08b]. prolate [KM09a]. proof [CS09b, CBH07, Wat09, ZWF09]. proofreading [CG05]. proofs [Sch07a]. propagation [DK06a, DDLM07, Li07a, Pre09, WDC^{+}08]. proper [LYZ09]. Properties [HHC09, HL09c, Mat07a, BH09, BCGA05b, BDS08, BJSS06, CFL05, Cim08, Cof09b, DSZ05, GG09a, HWL09, KYF05a, KYF05b, KS05a, LS06a, Las09, Lei08a, LF08, Mon09, MS06b, NS09b, Pie09, PS07b, Sav08, Sch05, WLY05, WTL05, WWWF09, XFgS06, YTM07, Zho08]. property [FL09a, Gu07]. proportional [Men09]. proportionality [CGPM08]. proposed [WYL08]. protocols [ABD^{+}05]. prove [OS07c]. Proving [FHL07]. proximal [CY08a, CY08c, Han08, KBA09, NB06, YN08]. proximal-point [NB06]. proximal-point-based [Han08]. proxy [XC06]. Prüfer [BBB07]. pruning [HJR07]. Pseudo [Ass09a, ACQ09, CLM06, CAHAY09, CP09b, KCS06]. pseudo-contraction [CAHAY09]. pseudo-contractive [ACQ09]. Pseudo-differential [Ass09a]. pseudo-parabolic [CP09b]. pseudo-spherical [KCS06]. pseudo-symplecticity [CLM06]. pseudocontractive [BS09c]. pseudohyperbolic [GRL09]. pseudomonotone [CY08a, LH09]. pseudorandom [Kan05]. Pseudospectra [GW06a, DW06a, Gra06, WMG08]. pseudospectral [BD06, GH08, GW06b, Koz09, OS06a, TL09]. pulse [BM08a, WJS^{+}09]. pulsed [KD05]. pump [HS07b]. purification [AVMMS^{+}06]. pyramid [ABT06]. Pythagorean [KL08b]. Pythagorean-hodograph [KL08b].

qd [AM09b]. qd-algorithm [AM09b]. QP [bJyKyZmT09, ZLZ09a]. QP-free [ZLZ09a]. quad [NQDJ06]. quad-tree [NQDJ06]. quadrangular [YLC09]. quadratic [Abb07b, ABG06b, CW08c, DL05, DL08a, DZLL06, FL09a, Han05, Hei05, Kuz08, Lan08b, Lan08a, Lin09a, hLyHZ08, Par05c, TJ08, XZZ09, YLC09, Zhu05c, Zhu08b, Zhu08c, dS07a]. quadratically [TJ08]. Quadrature [BP09, BDGV09, ELA09b, HS07a, MS06a, SH09, BC08, BDGV05, CE09, CV05b, CBDGVN07a, CBDGVN07b, EJ08, Gon09b, HS05, HV05, Huy09, JR07, KAC05, Kim05, pLT07, MJ08b, MCA09, MC05a, MC05b, NSL07, OMV09, Oou08, SL05, Sha08, WeR07, Xia05, Xia07a, Xie08, Yan05, Yüc06]. quadratures [ADDdM05, KK06, MS05a, MS07b, MSP08, MSP09]. quadrilateral [CJ06, EM08a, LW06b, LLD09, Zha07]. quadrilaterals [GHT06]. Qualitative [CHL09, WPC08, Cak07, FDFL07, Zho09b]. quality [DW06b, TMOG07]. qualocation [MNP08]. quantification [SV07a]. quantiles [MR09]. Quantitative [BG07e, NS09b]. quantities [GM06, Kor06]. quantized [NBSV08]. Quantum [AB09a, BSV08, ET06, PV09b, Tor08]. quartic [Chr05, CJ06, CHNZ08, Han06, HW07, JH09, WL06a]. quartically [Osa07]. Quasi [LMS08b, Sza08, BRIP09, BISS05a, BISS05b, BK08, CHW09, DL05, DL08a, DFZ07, Din05a, Din05b, DF08, FD09, FL09a, HLF09, Kaz06, Kul09, LCA07, MW09b, SVC09, SZ08b, WL06a, YOY07, Yan06d]. quasi-Cauchy [HLF09]. quasi-consistent [Kul09]. quasi-equilibrium [LCA07]. quasi-interpolants [BRIP09, BISS05a, BISS05b, DL08a]. quasi-interpolating [DL05]. Quasi-interpolation [LMS08b, BRIP09, CHW09, FL09a, MW09b, WL06a]. quasi-linear [Yan06d]. quasi-metric [SVC09]. quasi-Newton [BK08, DFZ07, YOY07]. Quasi-solutions [Sza08]. quasi-static [SZ08b]. quasi-variational [Din05a, DF08]. quasi-variational-like [Din05b, FD09, Kaz06]. quasicommutators [BGGC07]. quasiconvex [CS06b]. quasilinear [Dol08, Luo06a, Tat06, Vul07, Yao09]. quasilinearization [EGO06]. quasistatic [CFV06, CHSW08]. quasivariational [ZZ08b]. quaternion [JZW08]. quenching [Fer09, Rob07]. queue [AA09b, KL08a, TYH08, WWP09, WK07b]. queueing [ÉT08]. queues [GEK09]. Quicksort [SVC09]. quincunx [ABT06]. quintic [BM07a, CJ06, CMS08, DSI06].

R [LHW07, LM08]. Rachford [PJ06b].

radar [RCR05]. Radau

[JB09, MSP08, PRGPS09]. radial [BE07, CP09c, DRT09, DS09b, DT06b, FT05, HC08, JAE09, LÅH08, PLMP08, RB05, SS05, Sch08, YXYJ06]. radially [AV06, MN07a]. radiation [AE07, Sea07, ZGV06]. radiative [Eds09, Lan05]. radicals [JK07]. radius [HR05c]. Radon [GU08, vdBHPS07]. Raman [CC06b]. Ramanujan [Mil07, RVK05, RV05, VK06]. Random [HSN09, ARG08, ADGH06, AAR09, CHBMBR07, CK08, Dai09, DIKZ07, Hui05, JN09, Kan05, KLV09, Lin09a, SB09, ST05, Tan06, Tar09, TL07, WW09, Wik08]. randomness [Hat09]. range [ABB^{+}07d, LvGJ06, LP06b, PS07b, Wil07]. Rank [DV06a, TS06b, CS09a, DKHNZ08, DV08a, DV08b, GXgL05, GK07, HSL08, LLZ08, TM09, Wu08]. rank-deficient [DKHNZ08]. rank-one [Wu08]. rank-revealing [LLZ08]. Ranking [WLL09, WL09b]. RANS [GWG08]. Rao [DSMY06, Men09, Mil07]. Rapid [RB05, May08]. rapidly [ZZL08]. Rashba [MU08]. Rassias [Par05c]. rate [BS07, HO06, HLJ09, KI08, Lut07, SMT07]. rates [Chr05, DP05b, DC09c, HK07b, LCYZ08, TGJ19, YH07]. ratio [AS05a, FL06a, LCH09, XGM09]. ratio-dependent [FL06a, LCH09, XGM09]. Rational [JAE09, Mag09, PT09a, SWZ09, Bla05, BGVHN05, BGVHN08, CR09, Cla05, DZLL06, DZZT07, FO05, FKL05, GM09a, Gon09b, GM07a, GMFB06, HT06, HW07, HS08c, Joh05, LP06b, MDSR04, MC05c, PR07, SFV09, SH06, VB05a, VB05b, WT08, Waz07b, XW09, ZM09a, ZT06b, ZZ06c]. rationality [AS09a]. Rationalized [RA07]. rationing [BK09a]. ratios [SDP09, Soh05]. Raviart [BR07b, CY09, Jen07]. ray [Dom09]. Rayleigh [HWY09, LH08b, ZT07]. Rayleigh-type [LH08b]. Razumikhin [FL09d]. Razumikhin-type [FL09d]. RCMS [DS06]. re [DESC07, NMKS09]. re-blurring [DESC07]. re-scaling [NMKS09]. reaction [Alh06, ARVP09, AKL05, ABR07, ABV08, BH07b, CVB07, DS08, Du07b, GDM08, GC06, GL07, HKV09, HRGD08, KMVWY09, Kot08a, KHK08, LL05, LL08, LZY09, MJX09, MSZ07, RKK07, SSV06, SM09, VJ09, WZ09a, WF09, ZC08b]. reactor [HKB^{+}05, NGS05]. read [ACV05]. read/write [ACV05]. real [ADÓ07, BGVHN05, CDM07a, CDL05, CDMR07, CDM07b, GPTT05, Kam07, LL07b, Pet08, SHC08, Tab07, XXL08]. Reality [XHW07, Lig09]. realization [HV06a, ZKSS06]. reciprocal [CMRS06]. reconfigurable [KDH08]. reconstruction [ABB^{+}07d, GHMPV09, HHZ07, KY07, KS07e, PLGS09, YM08]. reconstructions [PS07b]. Recovering [FI07]. Recovery [LY07b, FZ09, YZ09d]. rectangular [BFGM06, DL08a, KS05b]. Recurrence [KV05, PG07, HM08, KT09, MP08a, Maz08, ZYD07]. recurrences [BM06, FFK05, KT07a]. recurrent [LH07a, LGYH09, Sha09b]. recursion [ADÓ07]. recursions [GW09, GST06, Zud05]. Recursive [LMS08a, Asa05, BSST08, Han09, Lar08, MVV08, MST05, Tan06, VW06, Wik08, ZSZ05]. reduce [CS09a]. Reduced [LT05, CU09b, LYZ09, MN07b, Pir09]. Reduced-order-based [LT05]. Reduction [BMW08, RKAH07, AG06, AMCM05, CK09a, GP07, HNS09, HW08, KS07a, KA08, LBS09, LW08b, Lu09b, Tsi06, Wil07]. redundancy [GBR09, WW09]. reference [HGS06, ZSGF09]. Refinable [CJ06, GPS08, HKZ09, Sau08]. refined [GW06b, LN07b]. refinement [AG06, AKK06, BKB06, DVCH06, HAS09b, HK08, Ker09]. reflection [GS06a]. Reflectionless [Sim07b]. reflexive [yPyHZ07, Sal09, Zho06]. reformulation [CP08a]. refraction [GS06a]. refuge [Kar06]. regime [FPS08, LJM09]. region [CU09b, MLY07, OZL09, QZZ08, QZY09, SS09c, SG08, SX09, SP09b, TSB09, YLW09a, Zho09a, Zhu05a]. regional [BB09a]. regions [And08, BDHW07, CB07, Lin07, Wat07, WN08]. registrations [TM08]. regression [Hu05a, MPP09, ZLLW08]. Regular [Bin07, AM09b, BS07, BSS08, CM07, CDL05, Gao05, May07, PS07a, PR05]. Regularity [Cim08, LN09, BGJ07, GS06c, PV09a]. regularization [BJ07b, CDM07a, CP08b, DD06, DFY09, GG09b, LR09, QWS09, RH09, TQA06]. Regularized [Est07b, Yam08]. regularly [Gal06]. regulator [BC09a]. regulator-insurer [BC09a]. regulatory [HBS^{+}07, TBBC07]. Reinsurance [LT06a]. Reissner [Sur05, XKF06]. related [Art09, Chr05, Cof05, Cof09a, DSZ05, EMP08, FKL05, Gau05a, GLV05, JT09b, LTW07, MP08a, MC05b, Ped05, ZZ08b]. relation [AGMMB05, Ext99, HLF09, Koo07, MS07a, Mil06]. relations [AM09a, ADÓ07, CHS06, FLS08, KV05, Maz08, NSK09, PG07, RI06, ZYD07]. relationships [LW08d]. relative [WC07]. relativistic [AT09d, Rui05, Sie07a]. Relaxation [DP08b, vdESvG05, Dao07, MV08, Mur07, TV09b, WH09b]. relaxed [Ban05, CAY09, CHJZ09, bJyKyZmT09, Kho09, LWYT06]. reliability [GBR09, KK09b]. reliable [KL08a]. Rellich [BT06]. reload [BF08a]. Relocation [DBE07]. remainder [MSP08, XC09]. remark [GRL09, Pet09a, RKK07].

Remarks [KS07d, MC05c, RLX08, Dom07, OS09]. removal [DMM07]. removals [Yi09]. renewal [AI06, DWY09, LWS09, MGV07]. repair [CKT09, WCY09]. repeated [GM06]. repertoires [Meh05]. report [Ela09a]. Representation [BXC09, WLL06, AGRZ05, DM07b, HW07, sR07, SCG08, Sza05a, TD06, YL09a]. Representations [vDvFZ09, vDvFZ13, ARN07, Cof08, CJ09b, KT07a]. representing [Lut07]. reproducing [CHW09, CG07, XZS07, XHK^{+}08, Zho08]. reproduction [FL09a]. Research [WWP^{+}08]. researches [LWC^{+}06, Wan08b]. residual [DD08, RVAD08, SSZ09]. resilience [YJY06]. Resolution [PP08, BCP07, LYT07]. Resolving [ZT06a, AG09, Vul07]. resonance [DLG05, SV07a]. resource [LL07a, YW08]. respect [Gau05a, Gei09, MM09, RS05b, Ter05]. respiratory [LH05a]. response [FS07, ICW^{+}09, LCH09, LTC06, RMKA05, ZAVS09]. responses [Wod05]. restart [BJK09]. restoration [BJ07b, GG09b, LNL09, SRW09]. restricted [LP06b, Nov08]. restricted-denominator [Nov08]. restriction [WLL09]. restrictions [Gon06]. Restrictive [BW06]. result [Cim09, VB05b]. resulting [WZ06]. results [AT07, AGRZ06, AT09c, BJS08, BI08b, Cam08, CHCW05, CCH06, CJ09b, CdAHBR06, HLX06, He07, JH08a, Lei08b, Li05c, LH08b, LmLgF09, LH07a, Rob07, VB05c, WHF07]. retarded [ZT07]. RETRACTED [Wan09a, WYO09]. retrial [AA09b, CK09b, Siv09, WK07b]. retrieval [ZLdST09]. returns [CT09b, LZWQ09, SdJMH09, Ver08]. revealing [LLZ08]. reverse [SCM09, WYS09]. revisited [GH09]. Reynolds [ACV05, CMM09]. RFLP [MMS^{+}09]. RFLP/PCR [MMS^{+}09]. rheology [NT08]. Riccati [Abb07b, DSW08, Guo06, Lut07, Mag09, RO05]. Richardson [CLN09, FHZ09]. Richardson-extrapolated [FHZ09]. Ridge [Hu05a, Ism07]. ridges [MOZ07]. Riemann [Abb07a, BFM09, Dun06, KM05, Pie08, WMN05, WN08]. Riemannian [JLR06]. Riesz [KSS07b, MB06, TV09a, TX07]. right [CHSW08, GQ09a, LWH09a, Won08, YLC09, DS06]. right-hand [CHSW08, GQ09a]. rigid [CFS09, XZS07]. rigorous [JT09a]. Risk [AS09a, GD09, KG09, CHS07, Den09, DWY09, FW09, GW09, LXSH09, VCD^{+}08, YZ09b, ZLY09]. risks [HG09]. Rival [ABB07e]. RK [CLM06, Pre09]. RKGL [Pre08a]. RKHSM [GC09]. RLW [DSI06, uIHA09]. road [BNP07, NL05]. Robin [CHY05, FZ09, YYF09]. Robust [CT09b, HWSL05, Kan08, TR08, AT09b, ABS07, DW06b, FSW09, GSW07, HGS06, MZC08, PT08b, Sin08, TND^{+}09, WG09, YY07c, ZAVS09, dS07a]. rock [LP07c, MSTT08]. RODAS [Sav09]. Rodrigues [AGRZ05]. role [CMST06, CSG05, MP07]. roofs [dCDND06]. root [Boy09, Dub09, HYG08b, Lan08b, Lan08a, LL07b, PRP08, RP09]. root-solvers [RP09]. Rooted [Koi05]. roots [Eig07, HR05a, Hom09, RV05, RK08b, VLR08, ZZ07, dB05b]. Rössler [NaLH07]. rotating [Wan07a]. rotational [TY07b]. rough [CWMT07]. round [VRBF07]. round-off [VRBF07]. routing [EM09]. ruin [LP09]. Rule [Eig07, CE09, EFP06, HS07a, KAC05, MC05a, SH09, Tar09, ZWY09]. rules [Gar07a, Huy09, JR07, Kim05, KC07, KV09a, MJ08b, MC05b, MS06a, OMV09, Rza09, Sin09, Zin05]. Rump [OTOR07]. Runge [Van07a, AM09a, AMCM05, AMCM06, AS05f, BJ07a, BJ08, BH07a, BH09, CGPM08, CFMR08, CFMR09, Cam08, HVV09a, Han07, HL07c, Jay07, KR07, Koi05, Kom07a, Kom07b, Kom08, Kot08a, Kot08b, LZ06, NSC07, RVA07a, RB08, SKAW09, Sun09, Toc05, TS05, VMMM09, Van05, Van06a, Ver06, WLS08, WHL08, Wan08a, YLN09]. Runs [AI06].

S [Yun08b]. Sabin [BFGP08, MB06, SDV06, SDV07]. Sabin-type [BFGP08]. saddle [Bac09, HWL09, JC09, JR08, Par07a, SHC08, Wan09b]. safe [LMR07]. safety [NL05]. Saint [LP07a]. Saint-Venant [LP07a]. Salpeter [Sie07a]. same [Oou08]. sample [PR08b]. samples [CDM07b, LWH09a]. Sampling [FG08, ABI07, BCP07, Hui07]. Saran [Tur09]. satisfying [BC09b, DG05, FL09a]. Satsuma [CX06]. saturated [EKEHR08, Hoo08, Zad09]. scalable [DHS09]. scalar [EMP08, Kom08, TS06a]. Scale [SdJMH09, yCyZ08, yCyZ09, CXH09, GKMS09, HLX06, Joh05, MvdMV06, PKS^{+}09, SCY09, VL06, WZ07, XZ08, ZT06a, ZJ12, ZFS09]. Scaled [Lin07, Gon09b, PT09a]. scales [AGH06, DaC05, EPS05, EBP09, HY09a, HJK09, He05, Hon07, Jac06, LZ09b, MS08a, RS05e, Sak05, Sak06, SL08, XX09, Yas07, YW09b]. scaling [AT09c, KCK09, NMKS09, Zhu05a, Zhu05b, Zhu06, Zhu08a]. scanning [SAHS09]. scattered [FL09a, KKLY06, LP06b, ZH08]. scattering [ABL07, ACWL07, AE07, Bou07a, BCK07, BG07f, CD05, CCM07, Cak07, CWMT07, CU09a, CNN07, EGM07, EM08b, GLS07, HMPR07, Mar07c, Nig07, RDH07, SHK^{+}08, ZKSS06]. scenarios [SK09b]. schedule [Saa06]. scheduling [BDL09, ZMAZ09]. Scheifele [Van09]. scheme [AT09b, ACR06, BH07b, BF09, BD06, BB09a, BG07c, BG07f, Cen08, CY08b, CAHAY09, CCCH09, CLO09, Che08, Cim09, CD07a, Cui08, DS09a, DZ09a, DZ09b, DD08, EMP08, GAR07b, GC06, GL07, Har08a, HMG06, HL09b, HLJ09, JN09, KD05, KCK06, LT07, lLzS07, Lu05, LYZ09, Ma06, MN07b, PS08, PP09, PS05, SK05, SZ08b, TS08, VVE08a, VRL07, Vil09, WKY^{+}07, WNZ09, WG09, YWD09, ZZZ06, ZWZ08]. schemes [ARVP09, ABBM09, AS05e, Ban05, BB08a, BB09b, CJB06, CHJL08, CGP07, CP09b, DSC09, Dai11, FC09, GLOS05, GP08, HPS06, HRGD08, HZ05, HP07, KMVWY09, Kot08a, Kuz08, Mat07b, NP08, NT06, Qiu07, Rad05, RVAD08, Roe08, Rom09, SKAW09, SBK07, SF08, Sun09, Sut07, Tab07, TYZ^{+}05, TGB08b, Toc09, VMMM09, VR08, WKY^{+}07, XY05, XC06, YLSX06, YXYJ06, ZP06]. Schmidt [RH06]. Schrödinger [BD06, CCJ07, CHJL08, EZ06, GS07c, JW08b, KCK06, KHWF09, LRV^{+}09, LN09, LC08c, MU08, MD09b, SS05, SBO09, Swe07b, WNZ09, ZZZ06]. Schur [Bac09, Deu09, LHL06, NL05]. Schwarz [And08, And09a, Bad08, Bog05b, Dao07, lLpZ08, Sch07b, SM09]. science [GPS09]. scientific [ZFS09]. scores [WC07]. Scott [ZWZ08]. scrambled [VC06]. screens [CD05]. SDD [Mor07]. SDE [LS06a]. SDELab [GS07b]. SE [QW06]. Search [BGP09, DEMR08, HE06b, bJQlZ07, KCLW08, QZY09, SS09c, SS06a, TJ08, XSW09, XSP09, YP08]. searches [hLyHZ08]. Secant [Gal06, HQ06, AB05, DFZ07, GK09, YOY07, YA08]. Second [AHS08, BM06, DSC09, Dai11, DD08, DO06, FPP07, GGA08, LG07, MS07a, ZWZ08, AS05e, AGRZ05, ADGR06, BS09a, BNT08, BF09, BM07b, Cen08, CHCW05, CP08a, CLA09, CL09b, CN08, DL08b, DZ09b, DG05, ED06, EGO06, EGFOA09, EBP09, FJT07, Fra06, GJ08, GO05, GPPRRB06, GPPRB07, GOT06, GZX09, HYYF05, Hsi09, HM09b, Ish07, Jai06, Jia09a, KT07a, KKLY06, Kom07b, Kom08, LC08a, Lei08a, LLZY06, LS07b, Liu08b, LB09, LLW09, LM09b, LB10, Luo06a, MB07a, MD09a, MY06, MY08, NWW08, NMKS09, NM06, PJ08, QG08, Qi08, QGC08, Qi09, RVA07b, RB08, Sak05, Sak06, Skr05, SLZA09, Sza08, TRS07, Toc05, Wan07b, WML08, XW07b, Xu08, Yan06b, YS07, YY07a, YL06, YCJ09, ZLL09, Zhu08d]. Second-order [AHS08, BM06, DSC09, Dai11, GGA08, LG07, ZWZ08, AS05e, BNT08, BM07b, Cen08, CP08a, CL09b, DZ09b, ED06, EGO06, FJT07, Fra06, GPPRRB06, GPPRB07, GOT06, GZX09, HYYF05, HM09b, Ish07, Jia09a, KT07a, KKLY06, Kom07b, Kom08, LS07b, Liu08b, LB09, LLW09, LB10, MD09a, MY06, MY08, NMKS09, NM06, RVA07b, RB08, Sak05, Sak06, SLZA09, Toc05, Wan07b, Yan06b, YS07, YCJ09, Zhu08d]. section [KMB09a, UHNC09]. sections [HW07]. sector [LS05]. sediment [DP08b]. sedimentary [CLQ09]. seeds [ZLLW08]. segment [WM09]. segregated [JR08]. SEI [YY08]. Seidel [Li05d, LC09b, NKM08, ZM09b]. selected [MvdMV06]. selecting [MPZ09]. Selection [Ver08, AKW05, BBF^{+}05, CM05b, CMST06, CT09b, DS05c, Hua08, Hui07, JRM09, LZWQ09, QLJ09, RVA05, YJJ09]. selective [RW06]. Self [Car05b, AGH06, BR07b, CPM^{+}05, Cha07, dCDND06, EGFOA09, EGFO09, EST07a, EMMP07, Gal09, KM09c, LLZY06, LP06a, MS07d, Pet09b, SS09c, WHG07]. self-adaptive [SS09c]. Self-adjoint [Car05b, AGH06, EGFOA09, EGFO09, EMMP07, KM09c, LP06a]. self-conjugate [LLZY06]. self-consistent [WHG07]. self-similar [EST07a, Gal09]. self-tolerance [CPM^{+}05]. self-validated [Pet09b]. self-weighted [dCDND06]. semantics [PPG07]. Semi [eMH06b, Bej06, BB09a, FG09, ZSGF09, Ahm07, ADDdM05, BB08a, BeMBH06, CW08b, FFSS09, FN09, GLdS08, HPS06, HL08, HeM05, HL09b, KCK06, LWYT06, LB07, MW09a, MLL09, SF06, TV07, TLQ08, Var07, VRSS08, Ver08, WG09, XZS07, XZZ09, YOA06, Zha09a]. semi-algorithmic [ADDdM05]. Semi-analytic [FG09]. Semi-analytical [eMH06b, Ahm07, FFSS09, LB07]. Semi-blind [ZSGF09]. Semi-cardinal [Bej06]. semi-definite [LWYT06, TV07, XZS07, XZZ09]. semi-discrete [KCK06]. semi-explicit [WG09]. semi-free [MLL09]. semi-implicit [BB08a, BB09a, FN09, HL09b]. semi-infinite [BeMBH06, CW08b, GLdS08, HL08, HeM05, LWYT06, MW09a, TLQ08, VRSS08, Ver08, YOA06]. Semi-Lagrangian [BB09a, HPS06]. semi-linear [SF06, Var07]. semi-positone [Zha09a]. Semiclassical [dMFPP07, MS07a, dMFPP08]. semiconductor [DBE07, FIR06]. semiconductors [PV09b]. semiconvergent [MRSZ06]. semidefinite [AH07, Yu06]. semidiscrete [BW07a, Cui08, KV07]. semigroup [Kok05]. semigroups [CS07e]. semilinear [AK09, Bog05a, Bog05b, EPP^{+}09, Fer09, KV07, MW07, SD09, Wan07c, XCZ08]. semilocal [Arg09b, Hu05b, ZBL08]. semiparametric [Hu05a]. semiseparable [Gem05, VVV08, VMV07]. semisingularly [ZC08b]. semismooth [CP08b, Śmi07, Zhu06, Zhu08a]. semistrictly [LW08d]. semitransparent [Sea07]. semivariance [Hua08]. Sensitivity [HyLC06, LP09, AHM09, DW06a]. Separable [SV07a, UO09, ZW06a]. separated [BV06, Log08]. separating [Toy07]. Separation [DJ07b, KSS07b, SJ09, SJZY09, SJZ09]. Septic [SA08a]. sequence [DMM07, Kil07, ZSZ05]. Sequences [CBDGVN07a, CBDGVN07b, Ber07, DJ09, JKK06, MDSR04, MC05c, MRSZ06, PC05, SCS08, Tor05, VC06, WS08]. Sequential [WCH05, FHZ09, TJ08, Zhu05c, Zhu08c]. serial [CG05]. Series [KSS07a, AGM09, BG07d, BE06a, BE06b, BSS09, yCyZ08, yCyZ09, CJ07, CZ08b, CV09, Cof05, Cof08, CJ09b, CS09b, Gau05b, Gin05, GBR09, HHST05, LV06, LV07, Mat06, MGMTO06, Mil09, Par09c, RI06, RVK05, Rea07, RS05d, SLH06, TMG09, UHNC09, Van08, WW09, DS06]. server [CKT09, KL08a, TYH08, WWP09]. service [BNDN09, CKT09, LLS09]. sesquilinear [Par05c]. set [Chi05, DMZ08, EM08a, HyLC06, bJxC09, KAP07, KGW09, LTY08, Lin05, Liu05, MMS^{+}09, PK07, RM07, UFG09, ZL09a]. set-valued [HyLC06, LTY08, ZL09a]. sets [CJCD06, CZ09, CK06, GLdS08, MPP09, XGZ09]. settling [MM06]. seventh [KLW07b]. seventh-order [KLW07b]. several [HHST05, Li05b, Nig07, Xio07, Xu05a]. SGBEM [ADF07]. Shabat [CHL09]. shadowing [Gu07]. shallow [AVMMS^{+}06, AB07, CGRGVP08, QW06, SF08]. shallow-water [CGRGVP08, SF08]. Shape [CFL05, DMZ08, PS08, YM08, Ata07, FGB07, FL09a, FOT07, GA09, HS09a, Han06, KGB^{+}07]. Shape-preserving [CFL05, PS08, FL09a, FOT07]. Shape-topology [DMZ08]. shaped [Li09c, YH09b]. shapes [ET06, PLGS09, SRB06]. Sharma [SWZ09]. Sharp [Che09, GO07]. shear [DDLM07]. sheared [eMH06a]. sheet [Pit09]. Sheffer [Kok05]. shells [Civ07]. Sherman [EM06]. Shifted [Sin09, VL06]. shifts [Wu07]. Shinnosuke [VAW07]. Shishkin [RO05]. Shiu [ZLY09]. Shock [TY07a, AAK06, GWG08, SBK07]. shock-compacton [TY07a]. shock-free [GWG08]. Shock-peakon [TY07a]. Shohat [Van08]. Short [Den07b, Sen07]. short-tailed [Sen07]. shortest [KK09b, Lin09b]. shunting [LMZ08, Liu07]. side [CHSW08]. sided [GPTT05, HYC05, Müh06b]. sides [GQ09a]. Sigmoid [GZX09]. sign [AMV06, DCZ07, LLW09, LOA05b, SS09b, Won05]. sign-based [AMV06]. sign-changing [LLW09, SS09b]. signaling [ZBB^{+}05]. signals [ZSGF09]. signature [XC06]. signomial [SaLJ08]. Signorini [xLLhZ08]. similar [EST07a, Gal09]. Similarity [eMH08, DA08, GW08]. Simple [Koç08b, BJK09, GR09, HR05c, LmZbL07, SS09c, Sau08, TTZ06, ZZZ09]. simplex [KA08, Lu09b, PC05, Sto06]. simplicial [BHP07, SDP09]. Simplified [HY09b, Toc09, BDHW07]. simply [Chu08]. Simpson [Gar07a, ZWY09]. Simulating [Göt05, BdAR08]. Simulation [ARG08, Dor09, SSW09, THS05, ACV05, BRVWM07, CHSW08, DFHS08, FHtMP06, GKMS09, GM09b, HKB^{+}05, IO06, KM06, KRAH07, KSE09, KGB^{+}07, LY05a, MJX09, OMC08, RKAH07, Sha05, WT07]. simulations [ABD^{+}09, BNP07, CFV06, Hat09, WZH06]. simulator [SL06]. Simultaneous [MV05b, PKS^{+}09, GAC^{+}08, PPR07]. simultaneously [CJCD06]. Sinc [AYK09, MNMS05, NMMS05, RZ07, SZD09, SK09a, WLK06, WKL06, NMM07]. Sinc-collocation [NMMS05, RZ07, SZD09, WLK06]. Sine [OOO^{+}07, Bra07b]. Sine-Gordon [OOO^{+}07, Bra07b]. single [BSS09, CT05, KL08a, LmZbL07, Lin08, PVC06, WM09, WZH06]. single-phase [WZH06]. single-population [Lin08]. single-species [CT05]. single-term [BSS09]. Singular [CK07, Yan06c, AM09b, AD05, ABV08, AKW05, BMfX07, Bar09, Ben07, BBB07, BF08b, CE09, CCD09, Cen08, CT09c, Civ07, CG07, DL09a, DV06a, DW05, DL08c, Dio09, EGFOA09, EGFO09, EJ07, ELA09b, Fer09, FPP05b, FPP08a, FPP08b, Gao05, Gao06, GJ08, GXgL05, Gon09a, HS07a, HMT09, HV05, IL05, KA07, Lan06, LS07b, Li08, LZ08a, LZ08b, LZ09b, LKSC06, LM09a, LJL06, LLW09, LOA05a, LOA05b, LP06a, LYZ09, MZC08, MJ09, MB07a, MVV08, May07, OYA07, PS07a, PS09a, PT06, PT08a, Pet07b, Scu09, SLZA09, TV07, Tem07, VG09, VAN06b, Vul07, WZ08, WD05, YOA06, Yan06d, Yan06b, YL09b, Zha05]. singularities [CO05, Kru05, MS06a]. Singularity [FFK05, HS09b, KS07c, Moh05]. singularly [ABS07, Att05, GC06, KGA08, KM09c, Lin05, LS05, MM07, MS07c, OS07c, RO05, Sto05a, TRS07, TR07, TR08, XWZH08, ZC08b]. sinh [CMRS06]. sinh-Poisson [CMRS06]. SIR [SMT07, YH07]. SIS [WZ06]. Sivashinsky [LT05]. Sixth [CFMR09, SA08a, Sut07]. Sixth-order [CFMR09, SA08a]. Size [KOS07, CLQ09, NBSV08, RVA05, Sic08, VRL07]. Size-structured [KOS07]. sizes [PR08b]. Skan [Asa05]. skeleton [PR05]. skeleton-regular [PR05]. Skew [AAR09, GQ09a, DGS06, Duf09, HYG08a, HWL09, JC09, KMB09b, Pie08]. Skew-Gaussian [AAR09]. skew-Hermitian [HWL09, JC09, KMB09b]. skew-orthogonal [Pie08]. Skew-symmetric [GQ09a, HYG08a]. skewed [AT09b]. slate [TMOG07]. sliding [SA09]. slip [HKA07]. slot [AVMVMV08, SS08]. slowly [Gau05b]. SLP [WWW08]. sludge [KRAH07]. slug [KPAT07]. small [CV06, FL06c, Hem06, JKM09, Mar07c, Ped05, ZT06a]. small-scale [ZT06a]. Smallest [MB05]. smooth [AU09, ABET08, And08, BN08a, BMW08, CT09c, DF08, DFLV06, HV05, JLR06, Mus09, ZCH07, ZZL08]. Smoothed [AFT^{+}06, CJ08]. smoothers [CdWJ06, Kan08]. Smoothing [GU08, KMVWY09, BK08, CL09b, DD06, HZW08, KP08b, MC08, Scu09, Sun09, WKY^{+}07, XZZ09, ZJW08, ZLL09, ZZ09a, Zho09a]. smoothing-type [HZW08]. Snowflake [NSS06]. SNP [WWWF09]. Sobolev [AGMMB05, BMS05, BS05, BCGA05b, Bou07b, Cim08, GS06c, Maz06, MOC09, MB07b, RS05b, Xie08]. soft [KGW09, RM07, XGZ09]. softening [ADF07]. software [CN07, DMM07, Gug06, Kro06a, MV05a, Sha07, SW06b]. soil [MD08]. soil-structure [MD08]. solid [HKM08, WT07]. solid-liquid [WT07]. soliton [BK09b]. Solution [Asa05, DFLP08, HLY09, Koh08, RZ07, TR07, VV05b, WWFV07, YKS09, Abb07a, AT09a, ADF07, Alh06, AAK06, AM09b, AS05e, AS05f, ABR07, AGRZ05, ABV08, BS08a, BS09a, BCeAJ09, BLW06, BF05, BFT07, Bat08, BTI08, BGP07, BBD06, BY09, Bor07, BSV08, BTS07, Bra07b, BMM06, BSS09, BK06, CM07, Car06, CdVT07, CXS06, CU09a, CU09b, CMM09, DSI06, Dao07, DSW08, DS09b, Djo07, DL09c, Eds09, EM06, EGM07, EM08b, ED06, FLMR07, FC09, GAC^{+}08, GEK09, GRW08, GLdS08, Gon09a, GM09c, Gug06, GGP09, HS08a, HV06a, HAS08, HK08, Hei05, HV06b, HO06, Hoo08, Hos06, HLL08a, HYG08a, JL06a, Jen07, JR08, Kas06, KAP07, Koç08b, LP07a, LL05, LL08, Lew06, LC06a, LS07b, LFHW09, LC06b, xMzY09, MB07a, MSZ07, May08, MP06a, MCA09]. solution [MG08, MN07a, MK07, Moh05, MS07c, MD09b, MNMS05, Mus09, NTS06, NPTH09, NST08, NM06, Oga06, OA06, OS06a, PK09a, PKS^{+}09, yPESlZ07, PK07, Pre08a, RO07, SK05, SS05, SAD07, SyS08, Sha09b, SCM09, SFPS09, uIHA09, SWZ09, SBO09, SDV06, Ste07, SSMB07, TL09, TR08, TSB09, TSC10, Tur06, VVE08a, VVE08b, Vol05, WCL08, WYL08, WD05, XHK^{+}08, ZCS06, ZCL06, Zhu07]. solution-dependent [HV06a]. Solutions [BB07, LCYZ08, LXUK06, LXZG09, OYA07, eMH06a, eMH08, AMG07, AS05d, Asl10, Ass09b, ADDdM05, BB05, BSC07a, BSC09, BK09b, BG07e, Bru08, BSS08, COEV05, CRS05, CX06, CX07, CL08, CHL09, CT09c, CMRS06, Cim08, Cla05, Daa09b, DFLV06, DA08, DAE05, Eba09, EZ06, Eng09, Enr06, EPS05, FWY09, FWZ09, FJG08, FPO05, Fin07, FL06c, FDFL07, GW08, Gin08, GJL08, HY09a, HEOS05, HCS05, HJK09, HKV09, Hat09, He05, HWY09, Hon07, HC08, Hon08, IL05, Iva06, JG09, Jia09a, Ko07, Kar09, KK06, Kaw07, KCS06, KS07d, KA07, KV07, KBB05, LWW08, Lai09a, LW05b, Li05b, LH05a, LH05b, LF07a, LD08, LH08b, LkLkP09, Li09c, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LKSC06, LJL06, LXX08, Liu07, LH07a, LTM08, Liu08a, LLZ09, LLW09, LOA05b, LJ09, Luo06a, Luo06b, Luo07, Lut07, MW07, MDMRP09, Mag09]. solutions [MO07a, MN05, MD09a, MLH08, ML09, MNHA09, NHN07, NMKS09, OS07b, OOM^{+}07, PV09a, yPyHZ07, PP05, PS05, RTGB09, RLX08, RAG07, RHQ06, RR06, SS09b, SWZ05, Sch07a, SWYZ09, SA08a, Soh05, SdJMH09, SA07b, SP06b, ST07, Sun08, SL08, SLZA09, Sza08, Thi05, TY07a, VJ09, Wan07c, WHG07, WYS09, WSM^{+}09, Wan09a, Waz07b, WW07, Wei09, WF09, WS09b, Won05, WL07, XW06, YS06, YOA06, YC05, Yan06d, YZ09c, YG09a, YM07, Yas07, YW09b, YG06, ZGV06, ZSZ05, Zha09a, ZCH07, Zha08, ZFG08, ZG09, ZC08a, ZLL08, ZT07, ZL07, ZW09c, ZSZ09, Zho09b, ZLDW09]. solvability [Cas07, NT08, Zho06]. solvable [DO06]. solve [AMCM05, BM08a, Koç08a, hLyHZ08, ZZ09b]. solved [SB09]. Solvency [Pit09, GD09]. solver [AFT^{+}06, ABD^{+}09, Duf09, HBS^{+}07, KDN08, LS06a, NS09a, NBSV08, SSV06, Wac05]. solvers [DRV08, DFFW06, GWG08, JF06, KDH08, KL09, LLL06, RP09, SS06b, SS06c]. Solving [AETE07a, AH07, AS07, BG09, BUG07, CGRGVP08, CL06b, CN07, CG07, EM08c, JDG06, Jan09, TRST09, YL09b, Yu06, ZA09, AEET09, eMH06b, AS05b, AKK06, AR09c, Arg09a, AT09d, BRW09, Bno07a, BM07b, BKB06, CS09a, Cou09, Dan06, DS08, DVCH06, Din05b, DZ09a, DJ07a, ESAD05, FT08, FH09a, FHM09, FFSS09, FZL07, FKM05, Fu08, GA09, GXgL05, GC09, GDA07, GS07b, GS08, GS06b, Guo06, GBR09, HCL08a, Has06, HLH07, HS08b, HVV09b, Hsi09, HZW08, JH08a, KD05, KYF05a, KYF05b, KRA08, KTH08, Kro06b, Lar08, LMR07, Li05c, Li06b, Li07b, lLzS07, lLpZ08, Li09a, pLT07, Lu07, LHW08, MW09a, Mar07a, MA09, Ned08, PT09a, PPR07, Pet09a, QZY09, Rad05, RA07, Sha09a, SZD09, SZ07a, Swe07b, TG07, TLQ08, Uje07, WWZY09, XH08, ZZ06a, ZZZ08, ZSWL09, Zhu05a, Zhu06]. Some [BBPR09, CCH06, CJ09b, CHY05, DL05, Dom07, DLG05, FH09a, Gin08, GM09b, HCL08a, He07, JK09, KLW07a, KLW07b, Kou08, Lei08b, LHM05, LmZbL07, LWC^{+}06, LHL06, LZ06, LF08, MC07, Ma08a, Mon07, OS09, Pie09, Sch05, Shi07, ÁNPQ05, ABP06, AGRZ06, Att05, Bar09, BDS08, BJSS06, CGPM08, Cof08, CP07, FWZ09, Gar09, GJL08, HR05b, HL09c, JM07a, JG09, JM07b, JT09b, KK06, KK08, KCS06, MO07a, MMFMG07, MDSR04, MJ07, MC05c, Moi05, MB05, MS06b, Pet07b, RTGB09, Tab07, Vid05, Vol08, WZ09b, XW07a, dF07]. Sommerfeld [MCA09, SK06]. SOR [NK10, DKHNZ08, MYY06, NKM08, yS05, WHF07, Yun05b, ZhL08]. SOR-type [WHF07]. sound [KS07e]. sound-hard [KS07e]. source [ABV08, CHV^{+}08, Dao08, DFY09, Gal09, HCDP08, JL07b, lLpZ08, LYB09, MM06, SJZY09, SJZ09, TRS07, TQA06, Wac05, ZSGF09]. sources [Du07b, SJ09, WHG07]. sourcing [LLS09]. sp. [MMS^{+}09]. Space [AB07, ARG08, AMG07, BLR07, BS06, Bra07a, CHV^{+}08, CG07, Fer07, GEK09, KDN08, LW06c, LLC08, MS07c, MM09, QW06, SAHS09, SLN08, Sin09, TG09, Thi09, YA07, ZL05, ZZZ06, Zhu08b]. space-dependent [ZL05]. space-fractional [ARG08]. space-localized [Fer07].

Space-time [AB07, BS06, QW06, ZZZ06]. spaces [AHL^{+}05, AU09, AABL08, ACQ09, Arg09b, BN08a, Bac09, BMS05, BS05, Bou07b, BGG07, CAY09, CW08c, CQ09, Cim08, DCF06, DF08, Far06, FJG08, FD09, GS06c, GO07, HLH07, HL09a, Kam07, Kaz06, KBA09, KA06, KSS07b, yL08, OYA07, PG07, QCK09b, QCKK09, RS05e, SVC09, Sal09, WL06a, WYL09b, XH08, ZTY07, ZC08a, ZLL08]. spacewise [JL07b]. Sparre [GY08, YZ09b]. sparse [AKK06, Duf09, GM09c, LO08, YH09a]. Spatial [MCR08, BB09b, HH07b]. Special [BBF07, BBG05, Cla05, GD09, Nou07, Var07, VAW07, VK08, WS05, GPS09, Koç08b, Pet07b, SCM09, Wik08]. species [CT05, LOZ08, WW07, WF09, Won09a, YTM07]. Specific [Tac07, FPP08a]. SPECT [BBM07]. Spectral [AR09b, CT09c, DHM07, DBG07, EE06, Eng09, HKK07, JLNP06, KM09a, Lui09, SBO09, ZZ06a, AGM06, BSC07a, ByXy05a, ByXy05b, ByXy06, DG08, DAE05, DJ07a, DBR08, FPP05b, FPP08a, FPP08b, GGD08, GM09a, GGM06, HMPR07, Her09, HS06, HR05c, HWL09, KTH08, KW07, Koz08, Li08, LLT07, LC08c, MDT07, MH08, MCR08, VRBF07, WLY05, Wan07a, WGZ08, ZP06, ZM09a, jia09b, vOP07, HE06b]. spectral-discontinuous [LC08c]. spectrally [YW09a]. spectroscopy [SV07a]. spectrum [Cha07, Chi05, Coj09, Har06, KN06, OOO^{+}07, Sie07a]. speed [BHJ05, Eds09, TFC07]. Speeding [GPRB05]. speeds [Dor09]. sphere [BDS08, KD05, LM07b, Sak09]. spheres [KSS07b, Sid05]. spherical [eMH06b, AG05, Car07, CM05a, DLROY05, KCS06, LMS08a, NP08, RR06, SBO09, WGZ08]. spherically [MDMRP09, ZGV06]. spheroidal [KM09a]. spin [KM06]. Spiral [AG05, WM09]. spirals [GM07a, MW09c]. Spline [PT06, AS05c, BISS05b, Bej06, DL08a, DP05a, DCF06, DZZT07, Far06, FO05, FOT07, GXZ06, Han09, JLR06, KGA08, KJ09, KKD06, LMS08a, LW06b, LLD09, LWLL09, LP06b, LW06c, LLC08, MST05, SH06, SA08a, SACJ09, WL06a, Wan08b]. splines [BISS05a, Bej06, Bou07b, CFL05, CJ06, CHNZ08, CGPS08, CMS08, DSI06, DL05, KJ09, KRA08, KP08a, KP08b, Liu06, LMS08b, Maz08, Müh06a, PR08a, PH09, PSV08, Rie06, Rom09, Ros08, Sab08, Sie07b, Sie08, SDV06, SDV07, UO09, WFL07, WF08, ZW06b]. Splinoid [DK05]. Split [RKK09, EZ07, MCA09, SBO09, Ter05, ZGH09]. split-Gaussian [MCA09]. split-operator [SBO09]. split-step [EZ07, ZGH09]. Splitting [HTS09, BB09a, Cas07, CHS08, DSC09, Dai11, FHZ09, FH09c, GZL07, Gei08, Gei09, GGP09, HYYF05, HWL09, HP07, JD05, JC09, KKW06, Koz08, Koz09, KMB09b, SS06b, TY07b, Wan07a, WWM07, WLMS08, Yun05b, ZZZ06]. Splitting-based [HTS09]. spring [HHZ07]. Spurious [AK07]. SQP [JTHZ05, bJQlZ07, Liu05, MZW06, XSP09]. square [BH09, CDV07, HR05a, LS06a, Log08, Mar07b, Möl07, OS06b, RB08, Sic08, SFPS09, Tay08, Toc05, YZL08]. squared [MvdMV06, SD06]. squares [BTI08, CR09, Cou09, DKHNZ08, DB08, DG08, DJ07a, GGD08, GRL09, JZW08, JSS07, KA07, MJB09, MV05a, RKK07, RKK09, SV07a, TY05, XZL08, XF07, YH09a, ZH08, ZSWL09, ZLDW09]. squeeze [BSKS07]. Srivastava [Tur09]. SSE [CGRGVP08]. SSLE [bJxC09]. SSLE-type [bJxC09]. SSOR [WHZ09, Yun08a, ZHG^{+}09]. stabilisation [DRV08]. stabilised [BW07a]. Stability [AMCM06, CS06a, CS07a, CLGG09, CWW08, DaC05, Ela09a, FZL07, Fra05, FG07, GLdS08, Kot08b, LWMT07, LL07b, Luo08, MS08a, MI09, PN06, SCS08, SC06a, WYL09b, XGM09, YC06, YHH05, ZSZ05, AT07, BB05, BS05, BH09, BG07e, COEV05, CS07b, CXS06, CdAHBR06, DZ09a, Eds09, FL06a, FSL09, FI07, FL08b, FL09c, FL09d, GG09a, GSW07, Har05, HWSL05, HS06, HL07c, KTM09, LS06a, LmZbL07, LZY09, Li09b, LY09, Liu09b, LZ09d, Luo07, Mao07, MMS^{+}09, Moh07, Moh08, MK05, NPTH09, PDM08, Par05c, PT08b, PT08c, RB08, RS05c, RR06, SG05, Sav08, SLN08, Sin08, SW06a, SC07b, Tab07, TXZ09, TSC10, Toc05, TLQ08, VRL07, VAMVC05, Wan07c, WLS08, WLS09, XZL09, XLHZ05, Yan06a, YZL08, YH07, YY07c, ZS05, Zha08, ZG09]. stabilization [CK09a, XW05, YZ09a]. Stabilized [GLOS05, LY07a, ABP08, BH07c, HH07a, HL09b, LH08a]. Stable [Di 07, GG08, Huy09, NSK07, Sur05, ACR06, ABL07, BJ08, DWW06, DK06b, KD05, LCKK06, LW07, Lin08, MVVA06, Tsi06, Van06a, Van07a, GLOS05]. stacks [Dul08]. Stage [Tsi06, AG09, CS07b, Fu08, GPP07, GEK09, JD05, LC09a, LCH09, LL07a, LZZD09, NBSV08, Ver06, Wan08a, WLF09, XW06, YTM07]. stage-order [Ver06]. stage-structured [LC09a, LL07a, LZZD09]. stages [YY08]. Staircasing [BW07a]. standard [ARVP09, AKL05, CN08, LP06a]. standing [SC07a]. star [PVC06, VB05b]. start [CO05, MJB09]. start-up [CO05]. Starting [Ver06, AOSY05]. startup [WWP09]. State [IT07, MLOP06, BLR07, BG09, CCK06, GEK09, Har05, Hsu05, HL07c, JL07a, KV09b, LC08b, MY06, NPTH09, RW06, SLN08, TSC10, WLL07, WPC08, Wan09a, WYO09, YC05]. State-dependent [IT07, BG09, Har05, WPC08, Wan09a, WYO09, YC05]. state-selective [RW06]. states [MY08, Yan06c]. Static [VS08, AG05, SZ08b]. Stationary [HL07a, ZW09b, CMM09, DS09a, HH07a, Ko07, KKD06, MYY05, Mot06, NHN07, TY05, ZH09]. statistic [BdWG06, Men09]. Statistical [Hui05, KS07a, ABÇ06, KLS^{+}07, MM09, Tsa09]. Statistically [BP07]. statistics [HL09c, MPP07, SD06, Tup07]. stator [SS08]. steady [DD08, Hsu05, MS09c, WK08]. steady-state [Hsu05]. steepest [DS05b]. Stefan [JVVvdZ06, VV05b]. Steffensen [AABL08]. stenosed [ICW^{+}09]. stenosis [Dem07b]. step [ASL07, BJ07a, BJ08, BRW09, Bno07a, BW07b, CL09b, CEJ09, CCLS06, DWW06, DT06a, DF08, EZ07, FJT07, Fra06, HRGD08, JLR06, KD05, LC08a, Lar08, LW07, MC08, NBSV08, PJ06b, PJ08, RVA05, Roi07, Sic08, Sza05b, Van07b, Van09, VRL07, Ver06, Vil09, WSPJ09, ZGH09]. step-size [RVA05, Sic08]. step-transition [DT06a]. stepping [AS09b, GKMS09, LS06a, SW06a]. steps [Eds09, SB09]. stepsize [RVA07b, VAR06]. Stewartson [SJ09]. Stieltjes [Ber07, BHJ05, GPTT05, TMG09]. stiff [GPRB05, JF06, Sav09, SSV06, VMMM09, VR08, WYL09b]. Stirling [EKL^{+}07]. STL [RB06]. Stochastic [Gri09, HR09b, LFS08, Mar08, RLH09, TBBC07, XWT09, YSY09, AR09a, AK07, BB05, BH07a, BH09, BS07, BLP07, BW07b, CJB05, CJB06, Car06, CX06, CX07, DR08, FSL09, FR07, FN09, GS07b, Gin08, Guo06, HBS^{+}07, HMG06, HR09a, KS06, Kom07a, Kom07b, Kom08, LkLkP09, Lin09b, Luo06b, Luo07, Luo08, MSZ07, MYY05, MYY07, Mao07, Mer05, PDM08, PR08b, RB08, RLX08, RS05c, RHQ06, SA09, Sha05, Sic08, TBL09, Toc05, WLL07, Wan08a, YZL08, YWD09, YJY06, YG06, Zha08, ZGH09, ZZ09b, ZW09c]. stochastic-flow [Lin09b]. Stochastically [Luo06b]. Stock [GM06, CT09b]. Stokes [Dai11, ABP08, BIR09, BH07c, CMX09, CLA09, DC08, DSC09, DMZ08, FLPA09, FH09c, GRW08, Göt05, GS06c, HH07a, HL09b, How09b, KDN08, Khu06, Koh07, KKD06, LY07a, LH08a, LY07b, Mot06, NHN07, Oga06, OS07b, SST09, SS08, VVE08a, YM08, YWD09, YXJG08, YXYJ06, ZH09]. Stokes/Darcy [CMX09]. stopping [MSE08]. storage [ACV05].

Storey [Zha09b]. Störmer [RVA05]. strategies [CP07, LP07a, Lin08, NGS05, PLGS09, Sav08, Ver08, YSY09, vdESvG05]. strategy [BJK09, BKB06, DWY09, EM08a, GY08, LWS09, LTC06, LXSH09, SS09a, Scu09, YZ09b, YW09a]. streamfunction [Fai07]. streamline [CWW08, HPS06]. stretching [EKEHR08, LO08]. strict [CAHAY09, JTHZ05, WCH05]. strictly [ACQ09, THC08]. strike [Hug06]. striped [Smo06]. Strong [BLP07, HK07b, QCKK09, TGJ19, BS09c, KYF05a, KYF05b, KS06, MS06a]. strongest [Ata07]. strongly [CS07c, bJxC09, KMB09b, VMV07, VDV08, Xu07a]. Structural [GVC^{+}08, BC09a, Mar07b, SRP08, Yua08, YD09a]. Structure [CFMR08, LW06c, sR07, BGP09, CS07b, CK07, CHLW08, ÉT08, Hat09, Ko07, Koo07, LCH09, LLC08, MS07a, MD08, RK08b, SBK07, VDV08, WLF09, XW06]. structure-preserving [CHLW08]. Structured [DW05, DW06a, WMG08, WA08, XW07a, DV08a, DV08b, Du07a, Gra06, KOS07, LLZ08, LC09a, LL07a, LM08, LZZD09, MV05a, MVVV06, NP06, YTM07, Zha08]. Structures [DV06b, DV06a, Jay07, Lai09a, LCZ09, SAHS09, YH05, YWH05a, YWH05b, ZT06a]. studies [Nig07]. study [AEET09, ABÇ06, BDY09, BSC07b, BGP07, BDL09, BNDN09, CJCD07, Car07, FH09b, Koh08, NSK09, RVM08, Tup07, Var07, ZL05, ZZZ06, ZTG09]. Sturm [AGM06, ABI07, BV06, BBB07, BSS08, CKWZ07, Cha07, Cha08, CW08d, FPP05b, FPP08a, FPP08b, GGM06, Liu08b, MJ08a, Sim07b, SO05, WWW08, Won05, Yüc06]. style [ABG06a]. sub [bJxC09, MP07, Mus09]. sub-diffusion [MP07, Mus09]. sub-feasible [bJxC09]. subcovariance [NL05]. subdiffusion [LYB09]. Subdivision [NP08, CS08, CGP07, CJ08, DS09a, GP08, MCD06, PS08, Rom09, SLZ06, XY05, ZC06]. Subdivisions [LP08, CJ06]. subgrid [Joh05]. Subject [Ano05-41, Ano06-38, Ano07-41, PP08, WTZ08]. subjected [Kas06]. sublinear [EBP09]. submatrix [yPlH05, YD08]. submerged [MOZ07]. subproblem [DS05c, SS09c]. subproblems [bJyKyZmT09]. subsets [EE06]. subsonic [CMRS06]. subspace [BJSS06, DGR07, FLMR07, SX09, XZ08]. subspaces [Den07a, GPV05, MVV08, WZ09b]. substances [LC09a]. Substructuring [WDC^{+}08]. successive [BT09, MLY07]. successively [Wu08]. suction [BeMBH06]. sufficient [BGG07, CHY05, HW07]. suitable [GZLL09]. suited [OOM^{+}07]. sum [Ism07, Par09a, SW08, SDP09]. Summation [Cof09b, CJ07, DMS05, HHST05, Mil09]. summations [CHM06, MS05b]. sums [FH07, Par05a, Par08a, VCD^{+}08]. Super [ABB^{+}07c, RO07, BCP07]. Super-attracting [ABB^{+}07c]. Super-fast [RO07]. super-resolution [BCP07]. superconductors [Slo08]. Superconvergence [Bi08a, DL08c, xLLhZ08, Xio07, YG09b, ZWY09, ZLZ09b, LY07b, MCR08, ZC08b]. Superconvergent [WSPJ09, BF05, CYM08]. superlattices [DBE07]. superlinear [AK09, LYW08, YOY07]. Superlinearly [KK08, bJxC09]. superpositions [AI06]. superquadratic [Sha09a]. supervision [Pit09]. support [BCK07, EL06, GHMPV09, Jai06, TMOG07]. supported [LP08]. supports [Liu06]. supremum [GP07]. sure [PDM08, RS05c]. Surface [SNV06, eMH06a, AE07, Ben06, BG07c, BGG08, BG07f, CW08a, CCLS06, EKEHR08, GLS07, GGP09, HMH08, IV06, IV08, RB06, SC07a, SLMW06, TK07, WK08, XF07, ZX06]. surfaces [BFGP08, Bra07a, CKWZ07, CM05a, DP08a, HW08, KCS06, KLLP06, KP07, Lu09a, MCD06, Sid05, WLL06, Wri08, rWlC06, XFgS06, XW09, ZC06]. surprising [CV09]. surrender [VV09a]. surrogate [WC08]. survey [CD07b, DG05, MGV07, Nou07, PPR07]. survival [CT05, DD07]. susceptibility [VV05a]. SVEIR [WJS^{+}09]. swarm [LW08a, YW08]. switch [VJ09]. switched [MLOP06, XW05]. switching [DBE07, LkLkP09, LJM09, MYY05, MYY07, YG06, ZW09c]. Sylvester [BLT09, BJ07b, HYC05, LLZ08, ZLDW09]. Symbolic [BBD06, FT08, CN07, HHST05]. Symbolic-numeric [BBD06]. Symmetric [DKHNZ08, ZFG08, ABG06a, AV06, ADÓ07, BW06, BXC09, BBR05, CFMR09, CK08, DM05, Duf09, ER07, GHZ06, GQ09a, GGA08, HYYF05, HYG08a, KV05, MDMRP09, MJ08a, MN07a, Sen07, SZ07a, TV07, VMV07, XZL08, YY07a, YD08, YLW09a, Yua09, ZhL08]. symmetrical [GMFB06]. symmetries [JKM09]. Symmetry [LV07, Tac07, BFT07, LLZ09, MPP07, NSS06, Wil06, vOP07]. Symplectic [KHWF09, SEAA08, CFMR09, WNZ09]. symplecticity [CLM06, DT06a, FJT07, IT07]. Synchronization [LY09, LC08b, BSC09, NaLH07, Par08b, SS09a]. Synchronizing [ZMAZ09]. syntactic [PPG07]. System [AU09, AT09a, BF05, BBSV08, Bra07a, CS07b, Cim08, CT05, DW05, DF08, DC09b, Du07b, DL09b, FL06a, FD09, GL07, GBR09, HC08, HM08, HHZ07, HZW08, JDG06, Jai06, Kar06, KBA09, Koh07, KS05b, KBB05, yL08, Lan08a, LJM09, LOZ08, zLCgS06, LC06b, Liu09c, MW07, MS09a, Min05, NS09a, OS07b, Par08b, PZ08, Siv09, Śmi07, SM09, TRS07, TBL09, VG09, WNZ09, WW09, WLF09, Wan09a, WD05, WW07, Wil06, Won05, Won08, XW06, XWT09, YC06, ZGV06, ZCS06, Zha08, ZZZ09, ZM09b, ZZ08b]. systematic [ADLT09]. Systems [Art09, FC09, SRP08, AKK06, ABI07, Aue07, Bac09, BW06, BSMT09, BC08, BBD06, BY09, BS07, BMC09, BJSS06, BFM09, CJCD07, CLP08, Cao08, CC08, CMST06, CGRGVP08, CKC^{+}07, CZ09, CHS08, CXH09, CMT09, Cou09, CHY05, DaC05, DL08b, DLROY05, EM06, EZ06, ÉT08, EM08c, FLMR07, FHM09, FPO05, FKM05, FG07, FL08b, FL09d, GDM08, GSW07, GGT07, GXgL05, Gia06, GLdS08, GM07b, GM09c, GT09, GQ09a, GF08, HWSL05, HNS09, HS08b, Hos09, HL07c, HSL08, HGS06, HMT09, bJyKyZmT09, JL06b, JH08a, KM06, KK08, KMVWY09, KRAH07, KR07, Koh08, Kok05, KT07b, KMB09b, KTM09, Lar08, LRV^{+}09, LZ05, Li05c, Li05b, Li05d, LmZbL07, LCA07, LWMT07, Liu09b, LM09b, LRT06, LY05b, LL07b, LHR09, LHL11, MZC08, MS08a, MSB05, May07, MGS08, MDB06, MLH08, MSA09, MK07, MS07d, MLOP06]. systems [MN07b, NSK09, Ned08, NKA09, NaLH07, OOM^{+}07, PKS^{+}09, PT08b, PN06, PK07, QCK09a, RB08, Roe08, RO07, SG05, SLN08, SZ07a, SSZ09, SL08, STB09, SCY09, TR06, THC08, Tup07, WCH05, WHF07, WLL07, WZ08, WYL08, WZ09a, WCW09, WWM07, WF09, WL07, WH09b, WHZ09, WZS06, XDW05, XW05, XZL09, XXL08, XX09, Yua08, ZhL08, yZXL09, ZLZ09b, ZJ12, ZLL08, ZSZ09, Zho09b, Zhu05a]. Systems-based [FC09]. Szego [GM09a, JR07, NSL07, Pet05, SL05].

T [CSG05, ZBB^{+}05]. T-lymphocyte

[ZBB^{+}05]. tabu [XC08]. tail [CV06]. tailed [Sen07]. tails [GA09]. talks [Ano05o]. tangent [KLLP06, Lu09a]. Tasso [SWZ09]. tau [DAE05, KW09, AT09a, GK07, Hos09]. Tau-functions [GK07]. Taylor [AZGN06, Asa05, CN07, GNS06, GS06b, MO08, Roi07, SAD07, Toc09, TMG09, XTZ06, Zha07]. tearing [NS09a]. tearing-based [NS09a]. technical [HL09c]. technique [AETE07a, CLA09, DH09, DEMR08, Dul08, FH09c, GBR09, HK08, JZW08, KM06, KCLW08, LYB09, MLY07, OS07c, Pet09a, SBO09, SK06, VT06, VVE08b, WZ08, YW08, YP08, ZLYW09, Zhu05a, Zhu06]. techniques [Ban08, CMX09, CK09a, CKC^{+}07, CHSW08, CHV^{+}08, EM08b, FL08a, GPS09, KMB09a, KLS^{+}07, Lia05, NMKS09, PK09a, TBL09]. telescoping [CHM06]. temperature [KSE09, XF07]. temporal [LB07, ST05]. tensile [DK06a]. tension [Bou07b, Rie06, TK07]. Tensor [Arp07, DP08a, Joh05, KJ09, Lu09a, SLMW06]. tensor-product [KJ09]. tensors [QWW08]. term [AEET09, AR09c, BSS09, Ext99, Fin08, FC09, KHWF09, LV06, LYB09, LLW09, Mil06, MSP08, TRS07, Tan06, ZSZ09]. terminating [CJ07]. terms [AML07, ABV08, FI07, Koi05, lLpZ08, LZY09, LZ09d, MS05b, VVE08a, VC06, Wac05, Xio07, ZS05]. Test [GK08, Alo07, HK08, Kan05, MPP07, MRV09, MK07, PS07b]. Testing [CV06]. tests [HE06b, KRAH07]. tetrahedral [PR05]. texture [vdBHPS07]. th [BB05, CKWZ07, Dub09, FWY09, FSL09, JG09, LRH07, LZ09c, MW09b, Luo07]. their [BM07b, CK09a, CZ08b, DGS06, Eig07, GG09a, GF08, LHM05, LZ06, pLT07, MC07, Ma08a, MJ07, NH08, OW06a, SP09a, SRB06, TYZ^{+}05, WC07, WS08, WD05, XDW05, XTZ06, Yua09, vdBHPS07]. theorem [AAZ06, BC09b, CLO09, CBGV05, EST07a, GP07, HL09a, Hu05b, LW05a, Liu09a, MJ08a, NK10, PK09b, SVC09, yS05, Van08, VV06, VZ08]. Theorems [ZZL08, BS09c, DJ07b, FG08, FL08b, FL09d, HL09a, HSL08, Kaw07, LOA05a, Mai08a, NS09c, QCK09b, QCKK09, Swa06, Tem09, WWM07]. theoretic [KGW09, RM07]. Theoretical [Gon06, LmLgF09, ADLT09, CP09a]. theories [LB07, WWP^{+}08]. Theory [Gao06, HCK07, XZL08, ABI07, Cak07, DIKZ07, FS07, Gao05, GM09b, MS07d, RW06, SHB05, Tab07, Zho08]. thermal [BGRS05, BM08b, CYZ09, Dor09, FZ08, SSW09]. thermally [RVM08]. thermo [KS07b]. thermo-elasticity [KS07b]. thermoelasticity [Swe07a]. thermoviscoelastic [Cop05]. theta [Sch05]. Thiele [ZT06b]. Thiele-like [ZT06b]. thin [Bou07b, CH09, FTT^{+}05, PSV08, VV05a]. thin-plate [PSV08]. thin-wire [FTT^{+}05]. Third [JYX07, LZG^{+}09, ABP06, ABB^{+}07c, CS06a, Chu08, EPS05, HAS08, Kar07, KSS07a, Koç08a, Koç08b, SS09b, Sun08, Wei09, YW09b]. Third-Order [LZG^{+}09, JYX07, ABP06, CS06a, Chu08, EPS05, Kar07, Koç08a, SS09b, Sun08, YW09b]. Thomas [Jen07]. Three [BRW09, Gal09, MP06a, NVD08, Wan08a, BISS05a, BJ07a, Bej06, Bra07a, BG07f, CP09c, DCZ07, EKEHR08, FG09, Har08a, He05, HR05b, Hon07, KD05, Kar07, LV06, LCZ09, LLT07, MJB09, Oga06, OA06, RVM08, SK09b, Sun08, SLZA09, SS06c, THS05, WGZ08, WF08, Wei09, Won09a, WDC^{+}08, YTM07, Yas07, Zha09a, ZWF09]. Three-dimensional [MP06a, BG07f, FG09, Har08a, LCZ09, LLT07, Oga06, OA06, RVM08, SS06c, THS05, WGZ08, WDC^{+}08, ZWF09]. three-direction [BISS05a, Bej06]. three-field [DCZ07]. three-level [KD05]. three-parameter [MJB09]. three-point [He05, Hon07, Kar07, Sun08, SLZA09, Wei09, Yas07, Zha09a]. three-species [Won09a, YTM07]. Three-stage [Wan08a]. Three-step [BRW09]. threshold [DWY09, GY08, JZK09, LWS09, LXSH09]. threshold-type [JZK09]. thresholding [FP08, LM07a]. Tight [CS08, LL06]. Tikhonov [BJ07b, LR09, QWS09, RH09]. Tikhonov-regularization [BJ07b]. tilings [Rho05]. Time [DDLM07, FHtMP06, Lin09b, SZ08b, ARG08, AR09a, AFT^{+}07, AS09b, AB07, AGH06, BGV^{+}09, BS06, BB09a, Cas05, yCyZ08, yCyZ09, Che08, CEJ09, CH09, CP09b, DaC05, DFLP08, Dul08, EIM^{+}09, EPS05, EBP09, FS07, FR07, FZL07, FG07, FL09c, GDM08, GSW07, GZL07, Gei08, GKMS09, GRW08, GAR07b, HY09a, HJK09, HLX06, He05, HRGD08, Hei05, HO06, HMG06, Hon07, HGS06, IV07, Jac06, JW08b, KGA08, KKW06, KS09, Koz08, Koz09, Kro06b, LS06a, LMZ08, LZY09, LY09, LZ09b, LZZD09, LN07b, LP09, LLT07, LC08c, LRB05, Ma06, MZC08, MS08a, MGMTO06, Moh08, Nig07, PW08, PT08b, PRGPS09, PN06, PP09, QW06, Qiu07, RTGB09, RR06, RS05e, SG05, Sak05, Sak06, SST09, SW06a, SP06b, SC06a, SBO09, SF06, SL08, TCMS06, TYH08, TXZ09, VRL07, WLL07, Wan07a, WLF09, WJS^{+}09]. time [XC08, XLHZ05, XW06, XX09, XGM09, Yas07, YJY06, YSYX08, YY07c, YW09b, YHH05, ZS05, ZZZ06, ZW09b, ZZL08, ZL09b]. time-delay [GSW07, HGS06]. time-delays [MZC08]. time-dependent [CEJ09, GRW08, JW08b, KGA08, KKW06, LC08c, Ma06, Nig07, PRGPS09, PP09]. time-discrete [Kro06b]. Time-discretization [SZ08b]. time-domain [CH09, Dul08, GZL07, Hei05]. time-extrapolated [FZL07]. time-fractional [AR09a, GAR07b]. Time-harmonic [DDLM07, DFLP08, LN07b]. time-independent [Nig07]. time-periodic [RR06]. time-splitting [BB09a, Wan07a, ZZZ06]. time-stepping [LS06a, SW06a]. time-varying [FG07, LMZ08, LZY09, LY09, PT08b, PN06, SC06a, TXZ09, YSYX08]. times [ST05, WWP09]. Timoshenko [lLzS07]. tip [WCL08]. tips [Kru05]. Toda [BG07b, CK07, WHG07]. Toeplitz [LHL11, AEBEK09, ÁNPQ05, BGGC07, DHM07, GT09, KV05, MGS08, NP07, TM09, WL09a]. tolerance [CGPM08, CPM^{+}05]. tomography [Qui07]. topics [Art09]. topographies [AMO07, AMO08]. topological [AHM09, Aki08, Aki09, FHL07, Gu07, HO07]. topologies [Sal09]. topology [DMZ08]. TOPSIS [Iza09]. tori [RD08]. Total [JSS07, HL08, MV05a, Pit09, WC07, WA08]. Totally [CGP07]. toxic [LC09a]. trace [How09b]. trace-free [How09b]. tracking [HPS06, JKM07]. trade [NSK09]. trade-off [NSK09]. traditional [NP06]. traffic [BNP07, HKS07]. trains [CHLW08]. trajectory [GF08, LFS08, LFW09]. transaction [JRM09]. transcription [CM07]. Transfer [KM06, Den09, Eds09, GS07a, GKMS09, HeM05, Lan05, WLMS08, ZGV06]. transform [CDM07a, CZ08a, CDMR07, CDM07b, EM08c, EK07, HV06c, Jan09, KM09a, KR08, LÅH08, LM08, TRST09, TX07, ZYLY09, dB05b, vdBHPS07]. transformation [BD07, CEF08, Eba09, Ehr06, EJ07, Koç08b, LO08, LZZ09, Mil09, MNMS05, NMMS05, NMM07, Par05b, Vul07, WHG07]. Transformations [CZ08b, Vid05, Ahm07, CMRS06, DHM07, GM09a, LV06, LV07, LW08b, SEAA08, Sid05, Sid08, Ter05, Wat06, Xia05]. Transforming [PC05, DK06a]. transforms [BNP06, MH08, MS08b, PK09a, PT08c]. transient [Alh06, KDN08, Ste07, TLQ08]. transients [AFT^{+}06]. transition [DT06a, ÉT08, HS09a]. transitions [Min05]. translation [Xu05a]. transmission [DCZ07, MDP08, RS08a]. transonic [DRT09]. Transport [PP08, CJ09a, DP08b, MMM08]. transportation [CW08b, ZMAZ09]. Transpose [YB05]. Transpose-free [YB05]. trapezoidal [Rza09]. trapped [KHWF09, MOZ07]. traveling [ZCH07]. travelling [CHL09, KBB05, LWW08]. treatment [ABD^{+}05, CLGG09, DL08b, FS07, GM07c, Koi05, MDMRP09, RKAH07, SL09, WLK06, WKL06]. tree [FFK05, Koi05, LM08, NQDJ06, Sak09]. tree-code [Sak09]. tree-structured [LM08]. Trefftz [LLM08, Li07b]. tri [Yan06a, Zud07]. tri-logarithms [Zud07]. tri-neuron [Yan06a]. triangle [LW06a, Tay08]. Triangular [ZX06, Bi08a, CW08a, HW08, KMB09b, MS09b, Śmi07, XW09]. triangulation [LW06c, LLC08, WOK05]. triangulation-based [WOK05]. triangulations [BDS08, CHNZ08, DL05, Far06, MB06]. tridiagonal [ÁNPQ05, KT07b, MGS08, Ter05, dF07]. Trigonometric [GJ08, VV09b, DWW06, Han05, NSC07, Xu07b]. trigonometrically [FSW09, PS05]. TRIOPT [WOK05]. Triple [Hon07, Wei09, Sun08]. trivariate [LWC^{+}06, SLMW06]. true [LB07]. truncated [BS07, BG07d, Moh09, ABG06a]. truncated-CG [ABG06a]. truncations [Zha07]. trust [MLY07, OZL09, QZZ08, QZY09, SS09c, SG08, SX09, SP09b, YLW09a, Zho09a, Zhu05a]. trust-region [QZZ08, YLW09a, Zhu05a]. Tsuzuki [WG09]. tube [WLMS08]. tubes [Dem07b]. tubular [HKB^{+}05]. tumour [KS07f]. Turán [MS05a, MS07b]. Turán-type [MS05a, MS07b]. turbomachines [Las08]. turbulent [CdVT07, DFHS08, SSW09]. turning [AHC05, CWW08, MN07a, Yan06b]. tutorial [MGV07, Sau08, VRSS08]. TVD [TS06a]. TVD-type [TS06a]. twice [Hu05b]. Two [AS05a, CYB09, CHJL08, CV05a, DCZ07, Jan08, Kim05, lLpZ08, Spr06, TF07, Ter05, Wil06, WL07, ZZ08a, ZM09b, ZH09, AB09b, Ban05, BJ07a, BJ08, Bra07b, CFHM06, CM05b, CMST06, CCW^{+}09, CKT09, CN07, CG07, DC08, Den07a, DMGVP05, DZ09b, DL07a, ER09, Ext99, FY06, FPP05a, FPP07, Fin07, Fra06, FT05, Fu08, Gao06, GZL07, GPP07, GEK09, GPTT05, GZ05, HS09a, HZBM06, HS07b, HK08, HVV09b, HYC05, HZ05, HV06c, JLR06, Jan09, JD05, JL06b, Joh05, KD05, KCK06, LWW08, LW06a, LW07, LH08a, LCYZ08, LBS09, LXUK06, LZ09d, LN07b, Lu07, LXSH09, MDT07, Maz06, MW08, MW09c, Mil06, Moh05, MS07c, MD09b, MS08b, NWW08, NaLH07, OS09, Rad05, RAG07, Roe08, SB09, SAHS09, Sav08, SSMB07, TRS07, TTZ06, Tan06, TS08, TG09, TRST09, TR07, TR08, Thi09, TG07]. two [TK07, TQA06, TANT09, VRBF07, Van07b, Van09, Ver06, WT07, WSM^{+}09, WSPJ09, WLK06, WKL06, WWP^{+}08, XZS07, XX09, Yan06b, ZS05, ZLY09, ZBL08, ZL07, ZL09b, ZKSS06, dMFPP07]. Two- [DCZ07]. two-dimensional [Ban05, Bra07b, DZ09b, DL07a, Gao06, HZBM06, HV06c, Jan09, JL06b, KCK06, LN07b, MDT07, MD09b, Rad05, Roe08, SAHS09, SSMB07, TTZ06, TS08, TG09, TRST09, TG07, TQA06, TANT09, VRBF07, WLK06, XZS07, XX09]. two-directional [LBS09]. two-factor [LCYZ08]. two-fluid [OS09]. Two-frequency-dependent [Kim05]. Two-grid [CYB09, CHJL08, DC08]. Two-level [lLpZ08, ZH09, MS07c]. two-objective [FY06]. two-phase [ER09, HS07b, TK07]. Two-point [Jan08, CM05b, CN07, CG07, DMGVP05, Fin07, HVV09b, Lu07, MW08, MW09c, WKL06]. two-sided [GPTT05, HYC05]. two-space [MS07c]. two-stage [Fu08, GPP07, JD05]. two-step [BJ07a, BJ08, Fra06, JLR06, KD05, LW07, Van07b, Van09, Ver06, WSPJ09]. two-term [Ext99, Mil06, Tan06]. two-variable [LW06a]. Two-weight [CV05a, Maz06]. Twostep [CX08]. Twostep-by-twostep [CX08]. tying [DB08]. type [AEBEK09, AMG07, AABL08, AGRZ05, AGMMB05, BS08a, BB05, BFGP08, BdWG06, BBSV08, BGG07, CRS05, CFMR09, CC06a, CC08, CR09, CKC^{+}07, CX06, CX07, CHS08, CZ08b, CX08, CD06, DL05, Dan06, EL07, ÉT08, ELA09b, FT08, FHH08, FLS08, FLS05, FL09b, FL09d, GU08, Gin08, GPRB05, GOT06, GMFB06, GS08, GO07, HWY09, Hoc08, Hom05, Hom09, HSL08, HZW08, HS05, IL05, JZK09, JK09, bJxC09, JKM09, Kam07, KG09, KTH08, KV07, KW09, Kor06, KMB09b, yL08, Lei08b, LHW07, LmY08, LH08b, LCH09, LKSC06, LY07b, Liu08a, LHR09, Ma08a, MB06, MLL09, Mai08a, MYY07, MP08b, MJ08b, Maz06, MLH08, ML09, MS05a, MS07b, NK08, OOO^{+}07, Oou08, Par05b, Par07b, Par07a, Ped09, Pie09, Qiu07, RVA07a, RVA07b, sR07, RAG07, RS05d, Rza09, SHK^{+}08, TS06a, TG09]. type [Tsi06, Van08, WHF07, WHG07, WHL08, WSM^{+}09, XZZ09, XL07, ZYD07, ZJW08, Zha09b, ZT07, dB05c, dB05b, dB05a, jia09b, AHS08, Ant09, KSVW07, LWH09a, Slo08]. type- [DL05]. Type-I [AHS08]. type-II [Slo08]. types [Gal09, LWW08, MB05, WF08].

Ulam [Par05c]. ultra [DM07b, HKM08]. ultra-weak [DM07b, HKM08]. ultradiffusion [Mar08]. ultraparabolic [Mar09]. ultrashort [KD05]. ultrashort-pulsed [KD05]. ultraspherical [DAE05]. un-reliable [KL08a]. unbalanced [EL06]. Unbounded [YOA06, CWMT07, EZ06, ELA09b, HZBM06, JM07a, JW08b, Moh07, ST07]. uncertain [GSW07, GVC^{+}08, KRAH07, MZC08, Neu08, Par08b, PT08b]. uncertainty [CKC^{+}07, KKS^{+}07]. unconfined [KY07]. Unconstrained [BT09, And09b, DEMR08, LTW07, OZL09, QZZ08, QZY09, SG08, VL06, WYL09a, YP08, YLW09b, Zha09b, ZZ09c]. unconstricted [Far06]. undamped [Yua08]. underdetermined [FKM05]. undergoing [FN09]. underlying [LWH09b]. Understanding [Maz08]. underwater [MOZ07]. underwear [LL06]. Unified [GM07c, WF08, BH07c]. Uniform [BR07b, BH07b, Dun06, HS09b, MM07, ZZ06b, AK07, BISS05b, BFGP08, NVD08, OS07c, SM09]. uniformities [FIR06]. uniformization [AKLT09]. Uniformly [LP06a, SH09, AU09, DF08, GL07, KGA08, RVAD08, ZLLW08]. union [PV09a]. Unique [Jai06]. Uniqueness [CLS08, TQA06, CXS06, DS07b, Fin08, HWY09, LFHW09, LJL06, MN07a, RLX08, ZL07, ZSZ09]. unit [BCGA05a, BR09, BDGV09, CBGV05, CBDGVN07a, CBDGVN07b, Ext99, Khr05, Mar07b, Mil06, PD06, Pet08, PVC06, VV09a]. unit-linked [VV09a]. Unitary [DV08b, DV08a, Gem05]. units [WY07, WYW08, WLL09]. unity [dB05b]. Univalency [BNP06]. Univariate [Liu06, CFL05, FDC08, GXZ06, Win06, dS07a]. univexity [Jay08, JK09]. unknown [DFY09, Luo06a, PS07b, Xu08]. unreliable [CKT09, TYH08]. unstable [Soh05]. Unsteady [ICW^{+}09, eMH08, BeMBH06, DD08, DL07a, EKEHR08, KCK06, Kas06, Rad05, TG07, YXYJ06]. unstructured [GWG08, MK09, SS08, TS06a]. unusual [EMMP07, MS05b]. UOBYQA [VB05c]. update [Cho08]. updated [Wu08]. updating [XZL08, Yua08, YD09b, YD09a]. Upper [BS07, LOA05b, xMzY09, Mor07, YOA06, ZCL06, COEV05, DSW08, HYG08b, QGC08, Shi07, WYS09, YG09a]. upwind [DD08, WSST05, YLSX06]. USAOR [LH07b]. use [BCP07, Duf09, Gar07a, HKK07, Jay07, MD09b, Par07b, Par07a]. used [AMCM05, BBPR09, EJ07, FGRW09]. uses [CDMR07]. Using [Iza09, ZZ09b, AETE07a, Abb07b, AH07, And09a, Asl10, Ass09b, BM08a, BS08a, BS09a, BP06, BDL09, BCK07, Bra07b, BKT07, BSS08, CGRGVP08, CHXL06, CN07, CS09b, CHV^{+}08, DSI06, DM07b, DH09, DS09b, Dmi08, DFLP08, Dor09, DT06b, DMZ08, EM08c, Fin07, FL06b, FL08a, FHL07, GMFB06, HKB^{+}05, HE06b, How09a, How09b, HLY09, JDG06, Jen07, Kan05, KJ09, KK06, KKLY06, KC07, KM09c, KLV09, LL01, LÅH08, LLD09, LB07, MJX09, MDT09, MSTT08, MJ08a, MJ08b, MD08, Mat08, MCA09, MR09, NMKS09, PLGS09, PSV08, RH06, SA09, SH06, SK09a, SA07b, SACJ09, Sto05b, TMOG07, TGB08b, VRBF07, VLR08, VVE08b, WY07, WGZ08, Wil07, Wil06, XHK^{+}08, YW08, Yüc06, YKS09, YXYJ06, Yun05a, ZMAZ09]. utilizing [WWWF09]. Uzawa [Wan09b].

V [JK09]. vacation [KL08a, WCY09]. vacations [CK09b, TYH08]. vaccination [WJS^{+}09]. vacuum [GKMS09, Yan06c]. validated [KKS^{+}07, Pet09b, RO07]. validity [BDGV09, NSL07]. Vallée [MT05]. valuation [HLJ09]. value [AG09, AOSY05, ADGH06, AABL08, AM09b, AGH06, Att05, BN08a, BMfX07, CD05, CM05b, CMST06, CHCW05, CHS07, CT09a, CG07, DWW06, Dan06, DLG05, EGFOA09, FI07, FJG08, Fin07, GGG09, GJL08, HY09a, HLX06, He05, HVV09b, Hon07, Hug06, Jan08, JG09, Kar07, Kar09, Koh07, Lan06, LS06b, Li06b, LS07b, Li08, LFHW09, LS07c, LZ08a, LZ08b, LZ09b, LZ09c, LZW09, LM09a, LJL06, Liu08b, LLW09, Liu09d, Lu07, LYZ09, LHW08, xMzY09, MDT07, Mar07a, MD09a, Moh05, MD09b, NM06, NMMS05, NMM07, OYA07, PS07a, PS09a, PK09a, Pre08a, Sal09, SFSK07, SA08a, Spr06, Sto05a, Sun08, SLZA09, Tur06, Wan07c, WWC07, WG08, WZ08, WYS09, WSZ09, WLK06, WKL06, XWZH08, YOA06, YS07, YLN09, Yao09, YL09b, YCJ09, Zha09a, Zha05, ZCL06, ZC08a, ZLL08, dB05a]. value- [dB05a]. valued [DS05a, FKL05, GS08, GBR09, HyLC06, Kon07, LTY08, ZL09a, ZZ06c]. values [Ben07, CDL05, CD07b, CDMR07, Fin08, FV08, KMJ06, LvGJ06, MLY07, Pet07b, PR07]. Valuing [FL08a]. valve [VDV08]. Vandermonde [YH05, YWH05a, YWH05b]. Vandermonde-like [YH05, YWH05a, YWH05b]. vane [SS08]. vanishing [AK07]. variability [CT09b]. Variable [Dem07b, RVA07b, VAR06, BJ06, Bru09, CJCD06, CX07, IO06, IS06, JZK09, LW06a, LCYZ08, Moh05, SyS08, SJ09, Sic08, Sid05, Sid08, VL06, ZLYW09, AMG07]. variable-coefficient [IS06]. variable-II [AMG07]. Variable-stepsize [RVA07b]. variables [CM08, DA09, FPP05a, FPP07, KKLM06, Lin09a, LB07, NWW08, RAG07, Xu05a, dMFPP07]. variably [Zad09]. variance [CT09a, YA08]. variant [KI08, LmY08, MZW06, RT08, RH09, ZZ09a]. Variants [Ban05, KLW07b, Kou08, YLS07, ZBL08]. variation [Bin07, HSN09, KI08]. Variational [AS05b, He07, Lu07, Swe07b, XHW07, AETE07a, AETE07b, Abb07a, Abb07b, AU09, Bno07a, Bno07b, CY08a, CAY09, Che09, CMM09, Dai09, DM07b, DS08, Din05a, Din05b, DF08, DMZ08, FPS08, FD09, Fu08, GS07a, GJM07, Gon06, Han08, He06a, HLH07, IO06, Kaz06, KBA09, Ker09, KP08a, KP08b, Kro06b, LmY08, lLpZ08, LH09, Mat07b, PT09b, PZ08, Pet09a, QCK09a, RCR05, SA07b, Swe07a, TD07, TY07a, Waz07a, Waz07b, XH08, YH09b, YKS09, ZL09a]. variational-like [Dai09]. variationally [AR09b]. variationally-posed [AR09b]. variety [Wan08b]. Various [Las09]. varying [BJK09, DaC05, FG07, KT09, LMZ08, LZY09, LY09, PT08b, PN06, SC06a, TXZ09, YSYX08, YY07c]. Vector [CZ08a, KA07, AD09, And09b, AKLT09, DS05b, EL06, KKLM06, KP08b, LCKK06, LCA07, Min05, TMOG07, VV05b, ZZ06c]. vector-valued [ZZ06c]. Vectorized [Sha08]. vectors [HKZ09]. vehicle [EM09]. velocity [HPS06, How09b, TY05]. velocity-vorticity-pressure [TY05]. Venant [LP07a]. Verification [KLO08, HO07, Ko07, Kaw07, NHN07]. Verified [DL07b, NSS07, OOM^{+}07, Pet07a, TF07]. Verifying [Enr06, Mar07b, MN07a]. version [ABG06b, IV07, Lin09b, RT09, XKF06, YG09b]. versions [KLS^{+}07]. versus [GM06, NP06, Wod05]. vertex [Ber06a]. vertex-centered [Ber06a]. Vertical [AVMVMV08, Kas06, RK08a]. via [AHM09, ADDdM05, Cao08, CC05, CZ07, CMM09, DMM07, Eba09, GVC^{+}08, Gem05, Gin05, GS08, HLF09, HY09b, JK07, JKM09, KSS07a, Koç08b, LSW09, LC08b, Maz08, PK06, Par09c, Pet08, SS09a, XXL08, Yam08, Yu06, ZW06a, ZW08, ZX06, ZTG09, ZA09]. Vibration [CHLW08, HGS08]. vibrations [Civ07, DFHS08, MD08]. VIEs [CC09]. VII [NK08]. VIM [MNHA09]. viral [Wod05]. virtual [PLGS09]. Virus [Wod05, MSB05]. visco [NT08]. visco-plastic [NT08]. viscoelastic [BdAR08, CEJ09, FPS08, FH09b, How09a]. viscoelasticity [CFV06]. viscometer [FGRW09]. viscoplastic [CFHM06]. viscosity [CAY09]. viscous [CMM09, Deu07, Fai07, HeM05, MCA09, OA06, Sma05, VVE08b, Wat09]. visualization [SH06]. Vlasov [WO08]. Vleck [MBA09]. VMs [AM09a]. VNS [fLxXT08]. Volterra [Jan09, TSC10, AT09a, BM08a, BS08a, Bar09, Bru09, CRS05, CC06a, CC08, CCD09, CMR06, CMV09, CT09c, CD06, DFLV06, DL08c, Dio09, Ela09a, Gan07, GKM07, HZBM06, Hsi09, Ko07, LC09a, LJM09, LOZ08, LL07b, Ma08a, MJ09, MO07a, NPTH09, RZ07, RA07, Rob07, ST07, TRST09, WYL09b, Won09a, XWT09, YC06, jia09b].

Volterra-type [jia09b]. Volume [Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano08k, Ano08l, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, CGRGVP08, AT09b, ABF^{+}07, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06b, Ano06c, Ano06d, Ano06e, Ano06a, Ano06m, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, BR07b, BG07c, CYB09, EPP^{+}09, EMP08, HH07a, HK08, HCDP08, KL09, LY05a, MP06a, MS09c, SM07, TS08, VVE08a, VVE08b, YLC09]. volumes [Ano05a, Ano05-41, Ano06n, Ano06-38, Ano07n, Ano07-41, MDT09]. Voronoi [NSK07]. vortex [Sak09]. vorticity [TY05, YXYJ06]. voting [WYL08]. VPAStab [GM07b]. Vries [BBSV08, DA08, SWZ05]. vulnerability [CHBMBR07].

w.r.t [CE09]. walk [ARG08, CK08]. walks [Kan05, ST05, Tar09, TL07]. wall [BeMBH06]. Wang [WYL08]. Warnaar [RS05d]. warping [BGV^{+}09]. wastewater [RKAH07]. water [AB07, CGRGVP08, QW06, SF08, WK08]. watermark [AA08]. watermarking [AAPP07]. watermarks [TND^{+}09]. waters [AVMMS^{+}06]. wave [AAK06, AE07, CHL09, CU09a, CU09b, DCZ07, Dor09, FZL07, Jen07, KM09a, KBB05, LWW08, Lai09a, Li06b, Li07a, Mat07b, PD06, Qin09, Swe07a, Tat06, ZCH07]. waveform [Dao07, WH09b]. Wavelet [BGV^{+}09, BSKS07, FTT^{+}05, KM09c, RW09, SJV^{+}09, ZLdST09, AAPP07, AA08, BP06, BSS09, BK06, CS08, CHXL06, De 08, Gan08, GJ08, GPV05, GPS09, HV05, HV06c, LM07a, LM08, NQDJ06, XF07, ZZZ06].

Wavelet-based [BGV^{+}09, SJV^{+}09, ZLdST09, AAPP07, AA08, GPS09]. Wavelet-like [FTT^{+}05]. Wavelet-multigrid [BSKS07]. wavelet-packet [HV06c]. Wavelets [CM08, BS09a, BSS08, Fer07, KJ09, KM09c, NS09b]. Waves [Dem07b, AG05, Ben06, DK06a, Dem07a, LmLgF09, LP07c, MOZ07, SC07a, WK08, WG06, ZP06, HH07c]. way [EZ07, Sen07]. Weak [BS09c, CJB06, FPP05a, Kom07a, Kom07b, Kom08, NS09c, PK09b, CL06a, CZ09, DM07b, DR08, FHH08, HKM08, MS09b, Sal09, SV07b, Toc09, VB05b]. weak-star [VB05b]. weakly [Bar09, CCD09, CT09c, DL08c, Dio09, GJ08, MJ09, OW06a, PS07a, PT06, PT08a, TRS07]. weakly-nonlinear [OW06a]. Weibull [MJB09]. Weierstrass [Bra07a]. weight [CV05a, ELA09b, Gau05a, HS05, Maz06, MC05b, WLL09, Xie08]. weight-function [HS05]. Weighted [KKD06, OMV09, ZLDW09, BNP06, BH07b, Cim08, Cou09, dCDND06, GS06c, KJ09, MJB09, MLY07, OYA07, Sin09, ZH08, ZSWL09]. weighting [MLL09]. weights [DJ05, JS08, KT09, MP08a, MOC09, PR07]. Weir [LTY08]. Weitzner [WWZY09]. Well [Zud05, ABL07, CJ07, GF08, UFG09, WS08, YG09a]. well-conditioned [ABL07]. well-distributed [UFG09]. well-ordered [YG09a]. Well-poised [Zud05, CJ07]. Wellposedness [Hug06]. wells [LFS08, LFW09]. Wendroff [Qiu07]. WENO [CD07a, Qiu07]. where [DA09]. which [CDMR07, HLF09, KCS06, Vil09]. White [CX06]. Whittaker [LVV05]. Wick [CX06, CX07, Gin08, MSZ07]. Wick-stochastic [MSZ07]. Wick-type [CX06, CX07]. Widom [DIKZ07]. Wiener [Kom08]. Wills [Bak05]. Wilson [Koo07, LHL09, VZ08]. wind [dCDND06, RCR05]. window [XTZ06]. wing [UHNC09]. wire [FTT^{+}05].

WITHDRAWN [KYF05b]. within

[FS07, UHNC09]. without

[GA09, Hui07, JTHZ05, Liu07, LGYH09, MNG09, WCH05, Yi09]. words [SVC09]. working [bJxC09, WCY09]. Wright [MP07]. write [ACV05]. Wu [Yun08b].

XFE [WDC^{+}08].

years [Bin07, Sim07a].

Zakharov [Bat08, MNHA09, WCZ07]. Zernike [Tor08, Wün05]. Zero [NP06, Osa06, Osa07, PM05]. zerofinding [Osa07, PM05]. Zero-structured [NP06]. Zeros [ADGR06, DJ09, FDC08, FS08, MOC09, BG07d, DR09, DJ07b, FHL07, GG07, GK09, JT09a, JT09b, Kra06, Lin07, MMFMG07, MBA09, MB05, MB07b, NSS07, PIP07, Pet09b, SA07a, Vol08, vDvFZ09, vDvFZ13]. Zeta [Dun06, Cof08, Cof09a, FV08]. Zhiber [CHL09]. Zones [Kál05].

References

Agreste:2008:NAP

[AA08] Santa Agreste and Guido Andaloro. A new approach to pre-processing digital image for wavelet-based watermark. Journal of Computational and Applied Mathematics, 221(2):274–283, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                                      //www.sciencedirect.com/
                                                      science/article/pii/S0377042707005602.

Adell:2009:IMG

[AA09a] José A. Adell and Horst Alzer. Inequalities for the median of the gamma distribution. Journal of Computational and Applied Mathematics, 232(2):481–495, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003811.

Amador:2009:MGR

[AA09b] J. Amador and J. R. Artalejo. The M/G/1 retrial queue: new descriptors of the customer’s behavior. Journal of Computational and Applied Mathematics, 223(1):15–26, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006541.

Alarcon:2008:STM

[AABL08] Virginia Alarcón, Sergio Amat, Sonia Busquier, and David J. López. A Steffensen’s type method in Banach spaces with applications on boundary-value problems. Journal of Computational and Applied Mathematics, 216(1):243–250, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002336.

Allan:2006:AAS

[AAK06] Fathi M. Allan and Kamel Al-Khaled. An approximation of the analytic solution of the shock wave equation. Journal of Computational and Applied Mathematics, 192(2):301–309, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003493.

Alfaro:2007:CCO

[AÁN07] M. Alfaro and R. Álvarez-Nodarse. A characterization of the classical orthogonal discrete and q-polynomials. Journal of Computational and Applied Mathematics, 201(1):48–54, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600077X.

Agreste:2007:IAW

[AAPP07] Santa Agreste, Guido Andaloro, Daniela Prestipino, and Luigia Puccio. An image adaptive, wavelet-based watermarking of digital images. Journal of Computational and Applied Mathematics, 210(1–2):13–21, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006431.

Alodat:2009:SGR

[AAR09] M. T. Alodat and M. Y. Al-Rawwash. Skew-Gaussian random field. Journal of Computational and Applied Mathematics, 232(2): 496–504, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003823.

Akhmet:2006:PTL

[AAZ06] M. U. Akhmet, J. Alzabut, and A. Zafer. Perron’s theorem for linear impulsive differential equations with distributed delay. Journal of Computational and Applied Mathematics, 193(1): 204–218, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003882.

An:2005:DSM

[AB05] Heng-Bin An and Zhong-Zhi Bai. Directional secant method for nonlinear equations. Journal of Computational and Applied Mathematics, 175(2):291–304, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400233X.

Ambati:2007:STD

[AB07] V. R. Ambati and O. Bokhove.

Space-time discontinuous Galerkin finite element method for shallow water flows. Journal of Computational and Applied Mathematics, 204 (2):452–462, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003724.

Andersen:2009:QHM

[AB09a] Jørgen Ellegaard Andersen and Christian Berg. Quantum Hilbert matrices and orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(3): 723–729, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001186.

Ayyad:2009:FAT

[AB09b] Y. Ayyad and M. Barboteu. Formulation and analysis of two energy-consistent methods for nonlinear elastodynamic frictional contact problems. Journal of Computational and Applied Mathematics, 228(1):254–269, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004731.

Abbasbandy:2007:ASN

[Abb07a] S. Abbasbandy. An approximation solution of a nonlinear equation with Riemann– Liouville’s fractional derivatives by He’s variational iteration method. Journal of Computational and Applied Mathematics, 207(1): 53–58, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004602.

Abbasbandy:2007:NAH

[Abb07b] S. Abbasbandy. A new application of He’s variational iteration method for quadratic Riccati differential equation by using Adomian’s polynomials. Journal of Computational and Applied Mathematics, 207(1):59–63, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004614.

Amat:2007:SAP

[ABB^{+}07c] S. Amat, C. Bermúdez, S. Busquier, J. Carrasco, and S. Plaza. Super-attracting periodic orbits for a classical third order method. Journal of Computational and Applied Mathematics, 206(1): 599–602, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004869.

Anconelli:2007:IMR

[ABB^{+}07d] B. Anconelli, M. Bertero, P. Boccacci, M. Carbillet, and H. Lanteri. Iterative methods for the reconstruction of astronomical images with high dynamic range. Journal of Computational and Applied Mathematics, 198(2): 321–331, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500717X.

Andrew:2007:RAM

[ABB07e] Sarah M. Andrew, Christopher T. H. Baker, and Gennady A. Bocharov. Rival approaches to mathematical modelling in immunology. Journal of Computational and Applied Mathematics, 205(2):669–686, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600392X.

Amat:2009:FHC

[ABBM09] Sergio Amat, Concepción Bermúdez, Sonia Busquier, and Driss Mestiri. A family of Halley–Chebyshev iterative schemes for non-Fréchet differentiable operators. Journal of Computational and Applied Mathematics, 228 (1):486–493, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004627.

Albrecher:2006:ASC

[ABÇ06] Hansjörg Albrecher, Rainer E.

Burkard, and Eranda Çela. An asymptotical study of combinatorial optimization problems by means of statistical mechanics. Journal of Computational and Applied Mathematics, 186(1): 148–162, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001949.

Adams:2005:HDM

[ABD^{+}05] B. M. Adams, H. T. Banks, M. Davidian, Hee-Dae Kwon, H. T. Tran, S. N. Wynne, and E. S. Rosenberg. HIV dynamics: Modeling, data analysis, and optimal treatment protocols. Journal of Computational and Applied Mathematics, 184(1):10–49, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000592.

Abdou:2008:IGE

[Abd08] M. A. Abdou. An improved generalized F -expansion method and its applications. Journal of Computational and Applied Mathematics, 214 (1):202–208, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001082.

Antonelli:2009:DCS

[ABD^{+}09] Laura Antonelli, Paola Belardini, Pasqua D’Ambra, Francesco Gregoretti, and Gennaro Oliva. A distributed combustion solver for engine simulations on grids. Journal of Computational and Applied Mathematics, 226 (2):197–204, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004007.

Amat:2008:LIC

[ABET08] Sergio Amat, Sonia Busquier, Antonio Escudero, and J. Carlos Trillo. Lagrange interpolation for continuous piecewise smooth functions. Journal of Computational and Applied Mathematics, 221(1): 47–51, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005237.

Alessandrini:2007:EVH

[ABF^{+}07] G. Alessandrini, A. Bilotta, G. Formica, A. Morassi, E. Rosset, and E. Turco. Evaluating the volume of a hidden inclusion in an elastic body. Journal of Computational and Applied Mathematics, 198(2):288–306, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007156.

Absil:2006:TCS

[ABG06a] P.-A. Absil, C. G. Baker, and K. A. Gallivan. A truncated-CG style method for symmetric generalized eigenvalue problems. Journal of Computational and Applied Mathematics, 189(1– 2):274–285, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006163.

Amat:2006:AVF

[ABG06b] Sergio Amat, Sonia Busquier, and José M. Gutiérrez. An adaptive version of a fourthorder iterative method for quadratic equations. Journal of Computational and Applied Mathematics, 191 (2):259–268, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005546.

Amat:2008:GCC

[ABGH08] S. Amat, S. Busquier, J. M.

Gutiérrez, and M. A. Hernández. On the global convergence of Chebyshev’s iterative method. Journal of Computational and Applied Mathematics, 220(1–2):17–21, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003913.

Annaby:2007:STB

[ABI07] M. H. Annaby, J. Bustoz, and M. E. H. Ismail. On sampling theory and basic Sturm– Liouville systems. Journal of Computational and Applied Mathematics, 206(1): 73–85, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003827.

Alouges:2007:SWC

[ABL07] F. Alouges, S. Borel, and D. P. Levadoux. A stable wellconditioned integral equation for electromagnetism scattering. Journal of Computational and Applied Mathematics, 204(2):440–451, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003712.

Amat:2006:CAP

[ABP06] S. Amat, Sonia Busquier, and S. Plaza. A construction of attracting periodic orbits for some classical third-order iterative methods. Journal of Computational and Applied Mathematics, 189 (1–2):22–33, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001342.

Araya:2008:ASF

[ABP08] Rodolfo Araya, Gabriel R.

Barrenechea, and Abner Poza. An adaptive stabilized finite element method for the generalized Stokes problem. Journal of Computational and Applied Mathematics, 214(2):457–479, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001409.

Araya:2007:EEA

[ABR07] Rodolfo Araya, Edwin Behrens, and Rodolfo Rodríguez. Error estimators for advectionreaction–diffusion equations based on the solution of local problems. Journal of Computational and Applied Mathematics, 206(1): 440–453, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004900.

Ansari:2007:PRF

[ABS07] A. R. Ansari, S. A. Bakr, and G. I. Shishkin. A parameter-robust finite difference method for singularly perturbed delay parabolic partial differential equations. Journal of Computational and Applied Mathematics, 205(1):552–566, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003384.

Amat:2006:NHM

[ABT06] Sergio Amat, S. Busquier, and J. C. Trillo. Nonlinear Harten’s multiresolution on the quincunx pyramid. Journal of Computational and Applied Mathematics, 189(1–2):555–567, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001433.

Ashyraliyev:2008:NSD

[ABV08] M. Ashyraliyev, J. G. Blom, and J. G. Verwer. On the numerical solution of diffusionreaction equations with singular source terms. Journal of Computational and Applied Mathematics, 216 (1):20–38, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002051.

Argyros:2009:IIP

[ACQ09] Ioannis K. Argyros, Yeol Je Cho, and Xiaolong Qin.

On the implicit iterative process for strictly pseudocontractive mappings in Banach spaces. Journal of Computational and Applied Mathematics, 233(2):208– 216, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004117.

Ahn:2006:SSN

[ACR06] Soyoung Ahn, U. Jin Choi, and Alexander G. Ramm. A scheme for stable numerical differentiation. Journal of Computational and Applied Mathematics, 186(2): 325–334, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000610.

Arregui:2005:DMC

[ACV05] Iñigo Arregui, J. Jesús Cendán, and Carlos Vázquez. A duality method for the compressible Reynolds equation. Application to simulation of read/write processes in magnetic storage devices. Journal of Computational and Applied Mathematics, 175(1):31–40, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002481.

Arden:2007:CMH

[ACWL07] S. Arden, S. N. Chandler-Wilde, and S. Langdon. A collocation method for high-frequency scattering by convex polygons. Journal of Computational and Applied Mathematics, 204 (2):334–343, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600361X.

Amiraliyev:2005:NPS

[AD05] G. M. Amiraliyev and Hakki Duru. A note on a parameterized singular perturbation problem. Journal of Computational and Applied Mathematics, 182(1): 233–242, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006132.

Al-Dosary:2009:PFM

[AD09] Khalil I. T. Al-Dosary. Period function monotonicity of planar vector fields. Journal of Computational and Applied Mathematics, 224(1): 405–414, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002471.

Amat:2008:EBC

[ADD^{+}08] S. Amat, K. Dadourian, R. Donat, J. Liandrat, and J. C. Trillo. Error bounds for a convexity-preserving interpolation and its limit function. Journal of Computational and Applied Mathematics, 211(1):36–44, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006935.

Avellar:2005:IFO

[ADDdM05] J. Avellar, L. G. S. Duarte, S. E. S. Duarte, and L. A. C. P. da Mota. Integrating first-order differential equations with Liouvillian solutions via quadratures: a semialgorithmic method. Journal of Computational and Applied Mathematics, 182(2): 327–332, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006326.

Aimi:2007:NAS

[ADF07] A. Aimi, M. Diligenti, and F. Freddi. Numerical aspects in the SGBEM solution of softening cohesive interface problems. Journal of Computational and Applied Mathematics, 210(1–2): 22–33, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006443.

Ahcan:2006:CCB

[ADGH06] Ales Ahcan, Grzegorz Darkiewicz, Marc Goovaerts, and Tom Hoedemakers. Computation of convex bounds for present value functions with random payments. Journal of Computational and Applied Mathematics, 186(1): 23–42, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001871.

Area:2006:ZJF

[ADGR06] Iván Area, Dimitar K. Dimitrov, Eduardo Godoy, and André Ronveaux. Zeros of Jacobi functions of second kind. Journal of Computational and Applied Mathematics, 188(1):65–76, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001615.

Andriopoulos:2009:SAC

[ADLT09] K. Andriopoulos, S. Dimas, P. G. L. Leach, and D. Tsoubelis. On the systematic approach to the classification of differential equations by group theoretical methods. Journal of Computational and Applied Mathematics, 230(1): 224–232, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800602X.

Agarwal:2007:AIN

[ADMM07] Ravi P. Agarwal, Smail Djebali, Toufik Moussaoui, and Octavian G. Mustafa. On the asymptotic integration of nonlinear differential equations. Journal of Computational and Applied Mathematics, 202(2):352–376, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001452.

Aristidou:2007:LFS

[ADÓ07] Michael Aristidou, Mark Davidson, and Gestur Ólafsson. Laguerre functions on symmetric cones and recursion relations in the real case. Journal of Computational and Applied Mathematics, 199(1): 95–112, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007429.

Atle:2007:SRC

[AE07] Andreas Atle and Bjorn Engquist. On surface radiation conditions for high-frequency wave scattering. Journal of Computational and Applied Mathematics, 204(2): 306–316, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003591.

Abdou:2009:TMM

[AEBEK09] M. A. Abdou, M. M. El-Borai, and M. M. El-Kojok. Toeplitz matrix method and nonlinear integral equation of Hammerstein type. Journal of Computational and Applied Mathematics, 223(2): 765–776, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000733.

Abassy:2009:NSA

[AEET09] Tamer A. Abassy, H. El Zoheiry, and Magdy A. El-Tawil. A numerical study of adding an artificial dissipation term for solving the nonlinear dispersive equations K(n, n). Journal of Computational and Applied Mathematics, 232(2):388–401, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003653.

Abassy:2007:SNP

[AETE07a] Tamer A. Abassy, Magdy A. El-Tawil, and H. El Zoheiry.

Solving nonlinear partial differential equations using the modified variational iteration Padé technique. Journal of Computational and Applied Mathematics, 207 (1):73–91, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004638.

Abassy:2007:TMV

[AETE07b] Tamer A. Abassy, Magdy A.

El-Tawil, and H. El Zoheiry. Toward a modified variational iteration method. Journal of Computational and Applied Mathematics, 207(1): 137–147, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004699.

Ala:2006:SPE

[AFT^{+}06] G. Ala, E. Francomano, A. Tortorici, E. Toscano, and F. Viola. Smoothed particle ElectroMagnetics: a mesh-free solver for transients. Journal of Computational and Applied Mathematics, 191(2):194–205, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005480.

Ala:2007:CMP

[AFT^{+}07] G. Ala, E. Francomano, A. Tortorici, E. Toscano, and F. Viola. Corrective meshless particle formulations for time domain Maxwell’s equations. Journal of Computational and Applied Mathematics, 210(1–2):34–46, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006455.

Amdjadi:2005:SWS

[AG05] Faridon Amdjadi and Jagnnathan Gomatam. Spiral waves on static and moving spherical domains. Journal of Computational and Applied Mathematics, 182(2): 472–486, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006430.

Aguilar:2006:AMR

[AG06] Juan C. Aguilar and Jonathan B. Goodman. Anisotropic mesh refinement for finite element methods based on error reduction. Journal of Computational and Applied Mathematics, 193(2): 497–515, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004280.

Abid:2009:MSI

[AG09] Fathi Abid and Dorra Guermazi. Multi-stage IT project evaluation: the flexibility value obtained by implementing and resolving Berk, Green and Naik (2004) model. Journal of Computational and Applied Mathematics, 233(1): 73–82, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002118.

Anderson:2006:HOS

[AGH06] Douglas R. Anderson, Gusein Sh. Guseinov, and Joan Hoffacker. Higher-order selfadjoint boundary-value problems on time scales. Journal of Computational and Applied Mathematics, 194(2): 309–342, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004760.

Aceto:2006:NCE

[AGM06] L. Aceto, P. Ghelardoni, and M. Marletta. Numerical computation of eigenvalues in spectral gaps of Sturm– Liouville operators. Journal of Computational and Applied Mathematics, 189(1– 2):453–470, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500021X.

Abebe:2009:FSP

[AGM09] Eyuel Abebe, Jameson Graber, and Charles N. Moore. Fourier series and the δ^{2} process. Journal of Computational and Applied Mathematics, 224(1):146–151, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001854.

Area:2005:SOP

[AGMMB05] I. Area, E. Godoy, F. Marcellán, and J. J. Moreno-Balcázar. Δ-Sobolev orthogonal polynomials of Meixner type: asymptotics and limit relation. Journal of Computational and Applied Mathematics, 178(1–2):21–36, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400384X.

Area:2005:HTD

[AGRZ05] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Hypergeometric type q-difference equations: Rodrigues type representation for the second kind solution. Journal of Computational and Applied Mathematics, 173 (1):81–92, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001359.

Area:2006:ESR

[AGRZ06] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Extensions of some results of P. Humbert on Bezout’s identity for classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 196(1): 212–228, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005595.

Al-Homidan:2007:SHM

[AH07] Suliman Al-Homidan. Solving Hankel matrix approximation problem using semidefinite programming. Journal of Computational and Applied Mathematics, 202 (2):304–314, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001415.

Al-Hayani:2005:ADM

[AHC05] Waleed Al-Hayani and Luis Casasús. The Adomian decomposition method in turning point problems. Journal of Computational and Applied Mathematics, 177 (1):187–203, May 1, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004224.

Agrawal:2005:BGC

[AHL^{+}05] M. R. Agrawal, P. G. Howlett, S. K. Lucas, S. Naik, and S. Ponnusamy. Boundedness of generalized Cesáro averaging operators on certain function spaces. Journal of Computational and Applied Mathematics, 180(2): 333–344, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005266.

Ahmed:2007:NSA

[Ahm07]  S. G. Ahmed.  A new

semi-analytical method for phase transformations in binary alloys. Journal of Computational and Applied Mathematics, 206(1):409– 419, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004870.

Abdulaziz:2008:AHP

[AHM08] O. Abdulaziz, I. Hashim, and S. Momani. Application of homotopy-perturbation method to fractional IVPs. Journal of Computational and Applied Mathematics, 216(2):574–584, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003081.

Abdelwahed:2009:CMA

[AHM09] M. Abdelwahed, M. Hassine, and M. Masmoudi. Control of a mechanical aeration process via topological sensitivity analysis. Journal of Computational and Applied Mathematics, 228(1):480–485, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004226.

Adimurthi:2007:P

[AHS07] A. Adimurthi, S. B. Hazra, and V. Schulz. Preface. Journal of Computational and Applied Mathematics, 203(2):309, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002147.

Ahmad:2008:SOD

[AHS08] I. Ahmad, Z. Husain, and Sarita Sharma. Secondorder duality in nondifferentiable minmax programming involving Type-I functions. Journal of Computational and Applied Mathematics, 215(1):91–102, May 15, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001768.

Alsmeyer:2006:RSR

[AI06] Gerold Alsmeyer and Albrecht Irle. Runs in superpositions of renewal processes with applications to discrimination. Journal of Computational and Applied Mathematics, 186(1):283–299, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002013.

Appleby:2007:SOU

[AK07] John A. D. Appleby and Cónall Kelly. Spurious oscillation in a uniform Euler discretisation of linear stochastic differential equations with vanishing delay. Journal of Computational and Applied Mathematics, 205(2): 923–935, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004134.

Antal:2009:MIS

[AK09] István Antal and János Karátson. Mesh independent superlinear convergence of an inner-outer iterative method for semilinear elliptic interface problems. Journal of Computational and Applied Mathematics, 226 (2):190–196, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003993.

Akin:2008:TEI

[Aki08] Hasan Akin. The topological entropy of invertible cellular automata. Journal of Computational and Applied Mathematics, 213(2):501–508, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000684.

Akin:2009:TDE

[Aki09] Hasan Akin. On the topological directional entropy. Journal of Computational and Applied Mathematics, 225(2): 459–466, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004184.

Al-Kurdi:2006:DIR

[AKK06] Ahmad Al-Kurdi and David R.

Kincaid. LU-decomposition with iterative refinement for solving sparse linear systems. Journal of Computational and Applied Mathematics, 185(2):391–403, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001226.

Anguelov:2005:NSF

[AKL05] R. Anguelov, P. Kama, and J. M.-S. Lubuma. On non-standard finite difference models of reaction–diffusion equations. Journal of Computational and Applied Mathematics, 175(1):11–29, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002419.

Aptekarev:2009:EVP

[AKLT09] A. I. Aptekarev, V. A. Kalyagin, V. G. Lysov, and D. N. Toulyakov. Equilibrium of vector potentials and uniformization of the algebraic curves of genus 0. Journal of Computational and Applied Mathematics, 233(3): 602–616, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001022.

Auzinger:2005:EMS

[AKW05] W. Auzinger, O. Koch, and E. Weinmüller. Efficient mesh selection for collocation methods applied to singular BVPs. Journal of Computational and Applied Mathematics, 180(1):213–227, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005084.

Alhumaizi:2006:MCM

[Alh06] Khalid Alhumaizi. A moving collocation method for the solution of the transient convection–diffusion– reaction problems. Journal of Computational and Applied Mathematics, 193(2): 484–496, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004279.

Alolyan:2007:NET

[Alo07] Ibraheem Alolyan. A new exclusion test for finding the global minimum. Journal of Computational and Applied Mathematics, 200(2): 491–502, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000306.

Andreev:2005:PHO

[ALR05] A. B. Andreev, R. D.

Lazarov, and M. R. Racheva. Postprocessing and higher order convergence of the mixed finite element approximations of biharmonic eigenvalue problems. Journal of Computational and Applied Mathematics, 182(2): 333–349, October 15, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006338.

Alzer:2006:II

[Alz06]  Horst Alzer.  Inequalities

involving Γ(x) and Γ(1/x). Journal of Computational and Applied Mathematics, 192(2):460–480, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003742.

Aceto:2009:RBB

[AM09a] L. Aceto and C. Magherini. On the relations between B_{2}VMs and Runge–Kutta collocation methods. Journal of Computational and Applied Mathematics, 231(1): 11–23, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000363.

Allouche:2009:NSS

[AM09b] H. Allouche and N. Marhraoui. Numerical solution of singular regular boundary value problems by pole detection with qd-algorithm. Journal of Computational and Applied Mathematics, 233(2): 420–436, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004403.

Alonso-Mallo:2005:ORH

[AMCM05] I. Alonso-Mallo, B. Cano, and M. J. Moreta. Order reduction and how to avoid it when explicit Runge–Kutta– Nyström methods are used to solve linear partial differential equations. Journal of Computational and Applied Mathematics, 176(2):293–318, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003395.

Alonso-Mallo:2006:SRK

[AMCM06] I. Alonso-Mallo, B. Cano, and M. J. Moreta. Stability of Runge–Kutta–Nyström methods. Journal of Computational and Applied Mathematics, 189(1–2):120–131, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042705000178.

Adimurthi:2007:CLF

[AMG07] Adimurthi, Siddhartha Mishra, and G. D. Veerappa Gowda. Conservation law with the flux function discontinuous in the space variable-II: Convex-concave type fluxes and generalized entropy solutions. Journal of Computational and Applied Mathematics, 203(2):310–344, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002172.

Aahlander:2006:EEM

[ÅMK06] Krister Åhlander and Hans Munthe-Kaas. Eigenvalues for equivariant matrices. Journal of Computational and Applied Mathematics, 192(1):89–99, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003109.

An:2007:CFT

[AML07] Heng-Bin An, Ze-Yao Mo, and Xing-Ping Liu. A choice of forcing terms in inexact Newton method. Journal of Computational and Applied Mathematics, 200 (1):47–60, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007958.

Arima:2007:CAF

[AMO07] T. Arima, Y. Matsuura, and S. Oharu. Computation of air flows and motion of environmental pollutants over complex geographical topographies. Journal of Computational and Applied Mathematics, 204(1):187–196, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002512.

See erratum [AMO08].

Arima:2008:ECA

[AMO08] T. Arima, Y. Matsuura, and S. Oharu. Erratum to “Computation of air flows and motion of environmental pollutants over complex geographical topographies”: [J. Comput. Appl. Math. 204(1) 187–196]. Journal of Computational and Applied Mathematics, 220(1–2): 759–761, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002567.

See [AMO07].

Anastasiadis:2006:ISB

[AMV06] Aristoklis D. Anastasiadis, George D. Magoulas, and Michael N. Vrahatis. Improved sign-based learning algorithm derived by the composite nonlinear Jacobi process. Journal of Computational and Applied Mathematics, 191(2):166–178, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005467.

Alvarez-Nodarse:2006:CCP

[ÁN06]  R. Álvarez-Nodarse.  On

characterizations of classical polynomials. Journal of Computational and Applied Mathematics, 196(1): 320–337, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005662.

Andersson:2008:MSC

[And08]  Anders Andersson.  A

modified Schwarz–Christoffel mapping for regions with piecewise smooth boundaries. Journal of Computational and Applied Mathematics, 213(1):56–70, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000052.

Andersson:2009:MSC

[And09a] Anders Andersson. Modified Schwarz–Christoffel mappings using approximate curve factors. Journal of Computational and Applied Mathematics, 233(4):1117– 1127, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709006281.

Andrei:2009:ACG

[And09b] Neculai Andrei. Accelerated conjugate gradient algorithm with finite difference Hessian/ vector product approximation for unconstrained optimization. Journal of Computational and Applied Mathematics, 230(2):570–582, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006523.

Anonymous:2005:AIVi

[Ano05a]  Anonymous.  Author in-

dex 10 volumes. Journal of Computational and Applied Mathematics, 180(2): 481–492, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002736.

Anonymous:2005:AIVh

[Ano05b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 180(2):479–480, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002724.

Anonymous:2005:AIVj

[Ano05c]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 181(2):473– 474, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003262.

Anonymous:2005:AIVk

[Ano05d] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 182(2):487–488, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003468.

Anonymous:2005:AIVl

[Ano05e]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 183(2):358– 359, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004450.

Anonymous:2005:AIVm

[Ano05f]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 184(2):538– 539, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500508X.

Anonymous:2005:AIVa

[Ano05g] Anonymous. Author index volume 173 (2005). Journal of Computational and Applied Mathematics, 173(2): 397–398, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004868.

Anonymous:2005:AIVb

[Ano05h] Anonymous. Author index volume 174 (2005). Journal of Computational and Applied Mathematics, 174(2): 437–438, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400514X.

Anonymous:2005:AIVc

[Ano05i] Anonymous. Author index volume 175 (2005). Journal of Computational and Applied Mathematics, 175(2): 465–466, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005680.

Anonymous:2005:AIVd

[Ano05j] Anonymous. Author index volume 176 (2005). Journal of Computational and Applied Mathematics, 176 (2):467–468, April 15, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006272.

Anonymous:2005:AIVe

[Ano05k] Anonymous. Author index volume 177 (2005). Journal of Computational and Applied Mathematics, 177 (2):475–476, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000348.

Anonymous:2005:AIVf

[Ano05l] Anonymous. Author index volume 178 (2005). Journal of Computational and Applied Mathematics, 178(1– 2):537–538, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000920.

Anonymous:2005:AIVg

[Ano05m] Anonymous. Author index volume 179 (2005). Journal of Computational and Applied Mathematics, 179(1– 2):381–382, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001767.

Anonymous:2005:CP

[Ano05n] Anonymous. Conference participants. Journal of Computational and Applied Mathematics, 179(1–2):ix–x, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001792.

Anonymous:2005:CT

[Ano05o] Anonymous. The conference talks. Journal of Computational and Applied Mathematics, 179(1–2):xi, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001809.

Anonymous:2005:Ea

[Ano05p] Anonymous. Editors. Journal of Computational and Applied Mathematics, 173 (1):ii, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004698.

Anonymous:2005:Eb

[Ano05q] Anonymous. Editors. Journal of Computational and Applied Mathematics, 173 (2):ifc, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004820.

Anonymous:2005:Ec

[Ano05r] Anonymous. Editors. Journal of Computational and Applied Mathematics, 174 (1):ii, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004959.

Anonymous:2005:Ed

[Ano05s] Anonymous. Editors. Journal of Computational and Applied Mathematics, 174 (2):ifc, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005114.

Anonymous:2005:Ee

[Ano05t] Anonymous. Editors. Journal of Computational and Applied Mathematics, 175(1): ii, March 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005321.

Anonymous:2005:Ef

[Ano05u] Anonymous. Editors. Journal of Computational and Applied Mathematics, 175 (2):ifc, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005643.

Anonymous:2005:Eg

[Ano05v] Anonymous. Editors. Journal of Computational and Applied Mathematics, 176(1):ii, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005758.

Anonymous:2005:Eh

[Ano05w] Anonymous. Editors. Journal of Computational and Applied Mathematics, 176(2): ifc, April 15, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006235.

Anonymous:2005:Ei

[Ano05x] Anonymous. Editors. Journal of Computational and Applied Mathematics, 177(1):ii, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000075.

Anonymous:2005:Ej

[Ano05y] Anonymous. Editors. Journal of Computational and Applied Mathematics, 177 (2):CO2, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000300.

Anonymous:2005:Ek

[Ano05z] Anonymous. Editors. Journal of Computational and Applied Mathematics, 178(1– 2):ii, June 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000877.

Anonymous:2005:El

[Ano05-27] Anonymous. Editors. Journal of Computational and Applied Mathematics, 179(1– 2):ii, July 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001718.

Anonymous:2005:Em

[Ano05-28] Anonymous. Editors. Journal of Computational and Applied Mathematics, 180(1): ii, August 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002050.

Anonymous:2005:FEa

[Ano05-29] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 181(1):ii, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002888.

Anonymous:2005:FEb

[Ano05-30] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 182(1):ii, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003353.

Anonymous:2005:FEc

[Ano05-31] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 183(1):ii, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004140.

Anonymous:2005:FEd

[Ano05-32] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 184(1):ii, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004589.

Anonymous:2005:IEIa

[Ano05-33] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 180(2):CO2, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002682.

Anonymous:2005:IEIb

[Ano05-34] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 181(2):CO2, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003225.

Anonymous:2005:IEIc

[Ano05-35] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 182(2):CO2, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003420.

Anonymous:2005:IEId

[Ano05-36] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 183(2):CO2, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004413.

Anonymous:2005:IEIe

[Ano05-37] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 184(2):CO2, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005042.

Anonymous:2005:ONB

[Ano05-38] Anonymous. Olav Njåstad:

Bibliography. Journal of Computational and Applied Mathematics, 179(1–2):xxi– xxviii, July 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004388.

Anonymous:2005:OP

[Ano05-39] Anonymous. Open problems. Journal of Computational and Applied Mathematics, 178(1–2):523, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004005.

Anonymous:2005:P

[Ano05-40] Anonymous. Participants.

Journal of Computational and Applied Mathematics, 178(1–2):xi–xiv, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004078.

Anonymous:2005:SIV

[Ano05-41]  Anonymous.  Subject in-

dex 10 volumes. Journal of Computational and Applied Mathematics, 180(2): 493–501, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002748.

Anonymous:2006:AIVe

[Ano06a] Anonymous. Author index to volume 189. Journal of Computational and Applied Mathematics, 189(1– 2):719–721, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000112.

Anonymous:2006:AIVa

[Ano06b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 185(2):482–483, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005418.

Anonymous:2006:AIVb

[Ano06c] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 186(2):542–543, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005947.

Anonymous:2006:AIVc

[Ano06d] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 187(2):261, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006394.

Anonymous:2006:AIVd

[Ano06e] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 188(2):336–337, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006813.

Anonymous:2006:AIVh

[Ano06f] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 191(2):306–307, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001701.

Anonymous:2006:AIVi

[Ano06g] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 192(2):481–482, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600207X.

Anonymous:2006:AIVj

[Ano06h]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 193(2):665– 666, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002792.

Anonymous:2006:AIVk

[Ano06i] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 194(2):460–461, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003074.

Anonymous:2006:AIVl

[Ano06j]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 195(1–2):351– 352, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003013.

Anonymous:2006:AIVm

[Ano06k]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 196(2):664– 665, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004444.

Anonymous:2006:AIVn

[Ano06l]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 197(2):612– 613, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004961.

Anonymous:2006:AIVf

[Ano06m]  Anonymous.  Author in-

dex volume 190. Journal of Computational and Applied Mathematics, 190(1– 2):549–550, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000604.

Anonymous:2006:AIVg

[Ano06n] Anonymous. Author index volumes 181–190. Journal of Computational and Applied Mathematics, 190(1– 2):551–562, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000616.

Anonymous:2006:FEa

[Ano06o] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 185(1):ii, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005273.

Anonymous:2006:FEb

[Ano06p] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 186(1):ii, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005753.

Anonymous:2006:FEc

[Ano06q] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 187(1):ii, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006060.

Anonymous:2006:FEd

[Ano06r] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 188(1):ii, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006667.

Anonymous:2006:FEe

[Ano06s] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 189(1–2):ii, May 1, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000070.

Anonymous:2006:FEf

[Ano06t] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 190(1–2):ii, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000550.

Anonymous:2006:FEg

[Ano06u] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 191(1):ii, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001336.

Anonymous:2006:FEh

[Ano06v] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 192(1):iv, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001920.

Anonymous:2006:FEi

[Ano06w] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 193(1):ii, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002676.

Anonymous:2006:FEj

[Ano06x] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 194(1):ii, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002883.

Anonymous:2006:FEk

[Ano06y] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 195(1–2):ii, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002974.

Anonymous:2006:FEl

[Ano06z] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 196(1):ii, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004304.

Anonymous:2006:FEm

[Ano06-27] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 197(1):ii, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004791.

Anonymous:2006:IEIa

[Ano06-28] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 185(2):CO2, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005376.

Anonymous:2006:IEIb

[Ano06-29] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 186(2):CO2, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500590X.

Anonymous:2006:IEIc

[Ano06-30] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 187(2):CO2, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006369.

Anonymous:2006:IEId

[Ano06-31] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 188(2):CO2, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006771.

Anonymous:2006:IEIe

[Ano06-32] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 191(2):CO2, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001658.

Anonymous:2006:IEIf

[Ano06-33] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 192(2):CO2, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002020.

Anonymous:2006:IEIg

[Ano06-34] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 193(2):CO2, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002755.

Anonymous:2006:IEIh

[Ano06-35] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 194(2):CO2, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003050.

Anonymous:2006:IEIi

[Ano06-36] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 196(2):CO2, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004407.

Anonymous:2006:IEIj

[Ano06-37] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 197(2):CO2, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004924.

Anonymous:2006:SIV

[Ano06-38] Anonymous. Subject index volumes 181–190. Journal of Computational and Applied Mathematics, 190(1– 2):563–572, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000628.

Anonymous:2007:AIVa

[Ano07a] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 198(2):521–522, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005486.

Anonymous:2007:AIVb

[Ano07b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 199(2):452–453, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005899.

Anonymous:2007:AIVc

[Ano07c] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 200(2):780–781, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007813.

Anonymous:2007:AIVd

[Ano07d] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 201(2):395–396, April 15, 2007. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000209.

Anonymous:2007:AIVe

[Ano07e] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 202(2):573–574, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000647.

Anonymous:2007:AIVf

[Ano07f] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 203(2):581–582, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001471.

Anonymous:2007:AIVg

[Ano07g] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 204(2):560–561, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001902.

Anonymous:2007:AIVh

[Ano07h]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 205(2):1019– 1021, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002427.

Anonymous:2007:AIVi

[Ano07i]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 206(2):1137– 1139, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002816.

Anonymous:2007:AIVj

[Ano07j] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 207(2):371–372, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003792.

Anonymous:2007:AIVk

[Ano07k]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 208(2):440– 441, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004256.

Anonymous:2007:AIVl

[Ano07l] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 209(2):267, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004773.

Anonymous:2007:AIVm

[Ano07m]  Anonymous.  Author in-

dex volume 210. Journal of Computational and Applied Mathematics, 210(1– 2):254, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005006.

Anonymous:2007:AIVn

[Ano07n] Anonymous. Author index volumes 201 to 210. Journal of Computational and Applied Mathematics, 210(1–2): 255–272, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005018.

Anonymous:2007:CP

[Ano07o] Anonymous. Conference photos. Journal of Computational and Applied Mathematics, 202(1):I–VIII, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000581.

Anonymous:2007:Ea

[Ano07p] Anonymous. Editors. Journal of Computational and Applied Mathematics, 201 (2):CO2, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000155.

Anonymous:2007:Eb

[Ano07q] Anonymous. Editors. Journal of Computational and Applied Mathematics, 202(1):ii, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000490.

Anonymous:2007:Ed

[Ano07r] Anonymous. Editors. Journal of Computational and Applied Mathematics, 203(1):ii, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001148.

Anonymous:2007:Ee

[Ano07s] Anonymous. Editors. Journal of Computational and Applied Mathematics, 204(2):ifc, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001859.

Anonymous:2007:Ef

[Ano07t] Anonymous. Editors. Journal of Computational and Applied Mathematics, 205(1): ii, August 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002117.

Anonymous:2007:Ec

[Ano07u] Anonymous. Editors 2. Journal of Computational and Applied Mathematics, 202 (2):CO2, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700060X.

Anonymous:2007:FEa

[Ano07v] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 198(1):ii, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005322.

Anonymous:2007:FEb

[Ano07w] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 199(1):ii, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005541.

Anonymous:2007:FEc

[Ano07x] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 200(1):ii, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006819.

Anonymous:2007:FEd

[Ano07y] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 201(1):ii, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600793X.

Anonymous:2007:FEe

[Ano07z]  Anonymous.  FM edi-

tors. Journal of Computational and Applied Mathematics, 204(1):ii, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700163X.

Anonymous:2007:FEf

[Ano07-27] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 206(1):ii, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002610.

Anonymous:2007:FEg

[Ano07-28] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 207(1):ii, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003007.

Anonymous:2007:FEh

[Ano07-29] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 208(1):ii, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003986.

Anonymous:2007:FEi

[Ano07-30] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 209(1):ii, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004426.

Anonymous:2007:FEj

[Ano07-31] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 210(1–2):ii, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004955.

Anonymous:2007:IEIa

[Ano07-32] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 198(2):CO2, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005437.

Anonymous:2007:IEIb

[Ano07-33] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 199(2):CO2, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005656.

Anonymous:2007:IEIc

[Ano07-34] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 200(2):CO2, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007771.

Anonymous:2007:IEId

[Ano07-35] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 203(2):ifc, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001422.

Anonymous:2007:IEIe

[Ano07-36] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 205(2):ifc, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002373.

Anonymous:2007:IEIf

[Ano07-37] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 206(2):ifc, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002774.

Anonymous:2007:IEIg

[Ano07-38] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 207(2):ifc, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003743.

Anonymous:2007:IEIh

[Ano07-39] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 208(2):ifc, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004219.

Anonymous:2007:IEIi

[Ano07-40] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 209(2):ifc, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004682.

Anonymous:2007:SIV

[Ano07-41] Anonymous. subject index volumes 201 to 210. Journal of Computational and Applied Mathematics, 210(1–2): 273–286, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700502X.

Anonymous:2008:AIVa

[Ano08a] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 211(2):232–233, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005870.

Anonymous:2008:AIVb

[Ano08b] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 212(2):469–470, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006425.

Anonymous:2008:AIVc

[Ano08c] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 213(2):616–617, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000228.

Anonymous:2008:AIVd

[Ano08d] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 214(2):637–638, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000630.

Anonymous:2008:AIVe

[Ano08e] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 215(2):639–641, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000988.

Anonymous:2008:AIVf

[Ano08f] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 216(2):600–601, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001738.

Anonymous:2008:AIVg

[Ano08g] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 217(2):457–458, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002331.

Anonymous:2008:AIVh

[Ano08h]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 218(2):616– 618, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002872.

Anonymous:2008:AIVi

[Ano08i] Anonymous. Author index volume. Journal of Computational and Applied Mathematics, 219(2):551–552, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003439.

Anonymous:2008:AIVj

[Ano08j]  Anonymous.  Author in-

dex volume. Journal of Computational and Applied Mathematics, 220(1–2):762– 764, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003920.

Anonymous:2008:AIVk

[Ano08k] Anonymous. Author index volume 221 (2008). Journal of Computational and Applied Mathematics, 221(2): 472–473, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004822.

Anonymous:2008:AIVl

[Ano08l] Anonymous. Author index volume 222 (2008). Journal of Computational and Applied Mathematics, 222(2): 732–734, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005311.

Anonymous:2008:EB

[Ano08m] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 222(1):ii, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004925.

Anonymous:2008:FEa

[Ano08n] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 211(1):ii, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005389.

Anonymous:2008:FEb

[Ano08o] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 212(1):ii, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006127.

Anonymous:2008:FEc

[Ano08p] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 213(1):ii, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006620.

Anonymous:2008:FEd

[Ano08q] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 214(1):ii, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000447.

Anonymous:2008:FEe

[Ano08r] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 215(1):ii, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000770.

Anonymous:2008:FEf

[Ano08s] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 216(1):ii, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001246.

Anonymous:2008:FEg

[Ano08t] Anonymous. FM editors.

Journal of Computational and Applied Mathematics, 217(1):ii, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001970.

Anonymous:2008:FEh

[Ano08u] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 218(1):ii, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002689.

Anonymous:2008:FEi

[Ano08v] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 219(1):ii, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800304X.

Anonymous:2008:FEj

[Ano08w] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 220(1–2):ii, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003750.

Anonymous:2008:FEk

[Ano08x] Anonymous. FM editors. Journal of Computational and Applied Mathematics, 221(1):ii, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004470.

Anonymous:2008:IEIa

[Ano08y] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 211(2):ifc, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005833.

Anonymous:2008:IEIb

[Ano08z] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 212(2):CO2, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006383.

Anonymous:2008:IEIc

[Ano08-27] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 213(2):ifc, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000198.

Anonymous:2008:IEId

[Ano08-28] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 214(2):ifc, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000599.

Anonymous:2008:IEIe

[Ano08-29] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 215(2):ifc, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000939.

Anonymous:2008:IEIf

[Ano08-30] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 216(2):ifc, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001696.

Anonymous:2008:IEIg

[Ano08-31] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 217(2):ifc, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002288.

Anonymous:2008:IEIh

[Ano08-32] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 218(2):ifc, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002823.

Anonymous:2008:IEIi

[Ano08-33] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 219(2):ifc, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003373.

Anonymous:2008:IEIj

[Ano08-34] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 221(2):ifc, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004779.

Anonymous:2008:IEIk

[Ano08-35] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 222(2):ifc, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800527X.

Anonymous:2009:AIVa

[Ano09a] Anonymous. Author index volume 223 (2009). Journal of Computational and Applied Mathematics, 223(2): 1087–1090, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006080.

Anonymous:2009:AIVb

[Ano09b] Anonymous. Author index volume 224 (2009). Journal of Computational and Applied Mathematics, 224(2): 779–781, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006614.

Anonymous:2009:AIVc

[Ano09c] Anonymous. Author index volume 225 (2009). Journal of Computational and Applied Mathematics, 225(2): 621–622, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000326.

Anonymous:2009:AIVd

[Ano09d] Anonymous. Author index volume 226 (2009). Journal of Computational and Applied Mathematics, 226 (2):388–389, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000806.

Anonymous:2009:AIVe

[Ano09e] Anonymous. Author index volume 227 (2009). Journal of Computational and Applied Mathematics, 227 (2):320–321, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002027.

Anonymous:2009:AIVf

[Ano09f] Anonymous. Author index volume 228 (2009). Journal of Computational and Applied Mathematics, 228 (2):600–602, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002611.

Anonymous:2009:AIVg

[Ano09g] Anonymous. Author index volume 229 (2009). Journal of Computational and Applied Mathematics, 229 (2):488–490, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002842.

Anonymous:2009:AIVh

[Ano09h] Anonymous. Author index volume 230 (2009). Journal of Computational and Applied Mathematics, 230(2): 823–825, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003288.

Anonymous:2009:AIVi

[Ano09i] Anonymous. Author index volume 231 (2009). Journal of Computational and Applied Mathematics, 231(2): 1005–1008, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003719.

Anonymous:2009:AIVj

[Ano09j] Anonymous. Author index volume 232 (2009). Journal of Computational and Applied Mathematics, 232(2): 655–657, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004786.

Anonymous:2009:EBa

[Ano09k] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 223(1):ii, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005694.

Anonymous:2009:EBb

[Ano09l] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 223(2):ifc, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006055.

Anonymous:2009:EBc

[Ano09m] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 224(1):ii, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006158.

Anonymous:2009:EBd

[Ano09n] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 224(2):ifc, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006572.

Anonymous:2009:EBe

[Ano09o] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 225(1):ii, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000193.

Anonymous:2009:EBf

[Ano09p] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 225(2):ifc, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000296.

Anonymous:2009:EBg

[Ano09q] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 226(1):ii, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000673.

Anonymous:2009:EBh

[Ano09r] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 226(2):ifc, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000752.

Anonymous:2009:EBi

[Ano09s] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 227(1):ii, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001885.

Anonymous:2009:EBj

[Ano09t] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 227(2):ifc, May 15, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001976.

Anonymous:2009:EBk

[Ano09u] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 228(2):ifc, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002362.

Anonymous:2009:EBl

[Ano09v] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 230(2):ifc, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003094.

Anonymous:2009:EBm

[Ano09w] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 231(2):ifc, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003677.

Anonymous:2009:EBn

[Ano09x] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 232(2):ifc, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004725.

Anonymous:2009:EBo

[Ano09y] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 233(3):ifc, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006116.

Anonymous:2009:IEI

[Ano09z] Anonymous. IFC editors 2nd issue. Journal of Computational and Applied Mathematics, 229(2):ifc, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002775.

Alvarez-Nodarse:2005:STT

[ÁNPQ05] R. Álvarez-Nodarse, J. Petronilho, and N. R. Quintero. On some tridiagonal k-Toeplitz matrices: Algebraic and analytical aspects. Applications. Journal of Computational and Applied Mathematics, 184(2):518–537, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000464.

Alefeld:2007:P

[ANR07] Götz Alefeld, Mitsuhiro T.

Nakao, and Siegfried M. Rump. Preface. Journal of Computational and Applied Mathematics, 199(2): 197–198, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007508.

Antczak:2008:PIF

[Ant08] Tadeusz Antczak. G-pre-invex functions in mathematical programming. Journal of Computational and Applied Mathematics, 217 (1):212–226, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003500.

Antczak:2009:OCD

[Ant09] Tadeusz Antczak. Optimality conditions and duality for nondifferentiable multiobjective programming problems involving d–r-Type I functions. Journal of Computational and Applied Mathematics, 225(1):236–250, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800366X.

Adee:2005:NSN

[AOSY05] S. O. Adee, P. Onumanyi, U. W. Sirisena, and Y. A. Yahaya. Note on starting the Numerov method more accurately by a hybrid formula of order four for an initial-value problem. Journal of Computational and Applied Mathematics, 175(2): 369–373, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002559.

Al-Refai:2006:BCP

[AR06] Mohammed Al-Refai. Bounds and critical parameters for a combustion problem. Journal of Computational and Applied Mathematics, 188 (1):33–43, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001585.

Abdel-Rehim:2009:EMT

[AR09a] E. A. Abdel-Rehim. From the Ehrenfest model to timefractional stochastic processes. Journal of Computational and Applied Mathematics, 233(2):197–207, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004105.

Alonso:2009:SAV

[AR09b] Ana Alonso and Anahí Dello Russo. Spectral approximation of variationally-posed eigenvalue problems by nonconforming methods. Journal of Computational and Applied Mathematics, 223(1): 177–197, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800006X.

Argyros:2009:CMN

[AR09c] Ioannis K. Argyros and Hongmin Ren. On the convergence of modified Newton methods for solving equations containing a non-differentiable term. Journal of Computational and Applied Mathematics, 231(2):897–906, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003185.

Argyros:2007:CNM

[Arg07] Ioannis K. Argyros. On the convergence of Newton’s method for a class of nonsmooth operators. Journal of Computational and Applied Mathematics, 205(1): 584–593, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003414.

Abdel-Rehim:2008:SCT

[ARG08] E. A. Abdel-Rehim and R. Gorenflo. Simulation of the continuous time random walk of the space-fractional diffusion equations. Journal of Computational and Applied Mathematics, 222(2): 274–283, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005961.

Argyros:2009:CNL

[Arg09a] Ioannis K. Argyros. On a class of Newton-like methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 228(1):115–122, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004433.

Argyros:2009:SCI

[Arg09b] Ioannis K. Argyros. On the semilocal convergence of inexact Newton methods in Banach spaces. Journal of Computational and Applied Mathematics, 228 (1):434–443, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005098.

Andrle:2007:EOB

[ARN07] Miroslav Andrle and Laura Rebollo-Neira. Experiments on orthogonalization by biorthogonal representations of orthogonal projectors. Journal of Computational and Applied Mathematics, 205(1):545–551, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003372.

Arponen:2007:TIN

[Arp07] Teijo Arponen. Tensor invariants in numerical geometric integration of ODEs. Journal of Computational and Applied Mathematics, 205(2): 791–801, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600402X.

Artemiadis:2009:ESM

[Art09] Nicolaos K. Artemiadis. Educational Systems (ES) in Mathematics and other related topics. Journal of Computational and Applied Mathematics, 227(1):5–9, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003245.

Alvarez-Ramirez:2009:NSF

[ARVP09] Jose Alvarez-Ramirez and Francisco J. Valdes-Parada. Non-standard finite-differences schemes for generalized reaction– diffusion equations. Journal of Computational and Applied Mathematics, 228 (1):334–343, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005013.

Abad:2005:TNA

[AS05a] Julio Abad and Javier Sesma. Two new asymptotic expansions of the ratio of two gamma functions. Journal of Computational and Applied Mathematics, 173(2): 359–363, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001669.

Abdou:2005:VIM

[AS05b] M. A. Abdou and A. A. Soliman. Variational iteration method for solving Burger’s and coupled Burger’s equations. Journal of Computational and Applied Mathematics, 181(2): 245–251, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005825.

Adell:2005:ABS

[AS05c] J. A. Adell and C. Sangüesa.

Approximation by B-spline convolution operators. A probabilistic approach. Journal of Computational and Applied Mathematics, 174(1): 79–99, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001736.

Allan:2005:ASN

[AS05d] Fathi M. Allan and Muhammed I.

Syam. On the analytic solutions of the nonhomogeneous Blasius problem. Journal of Computational and Applied Mathematics, 182(2): 362–371, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006351.

Amodio:2005:HOF

[AS05e] Pierluigi Amodio and Ivonne Sgura. High-order finite difference schemes for the solution of second-order BVPs. Journal of Computational and Applied Mathematics, 176(1):59–76, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003206.

Anastassi:2005:ORK

[AS05f] Z. A. Anastassi and T. E. Simos. An optimized Runge– Kutta method for the solution of orbital problems. Journal of Computational and Applied Mathematics, 175(1):1–9, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002432.

Alessandrini:2007:SEC

[AS07] Giovanni Alessandrini and Eva Sincich. Solving elliptic Cauchy problems and the identification of nonlinear corrosion. Journal of Computational and Applied Mathematics, 198(2): 307–320, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007168.

Aalabaf-Sabaghi:2009:RPR

[AS09a] Morteza Aalabaf-Sabaghi. Risk perceptions and rationality in measures of risk. Journal of Computational and Applied Mathematics, 233(1):46–50, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004664.

Alexe:2009:DAA

[AS09b] Mihai Alexe and Adrian Sandu. On the discrete adjoints of adaptive time stepping algorithms. Journal of Computational and Applied Mathematics, 233(4):1005– 1020, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006062.

Asaithambi:2005:SFS

[Asa05] Asai Asaithambi. Solution of the Falkner–Skan equation by recursive evaluation of Taylor coefficients. Journal of Computational and Applied Mathematics, 176 (1):203–214, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003279.

Askey:2005:DCO

[Ask05] Richard Askey. Duality for classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 178(1–2):37–43, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003644.

Arevalo:2007:CCL

[ASL07] Carmen Arévalo, Gustaf Söderlind, and José Diaz López. Constant coefficient linear multistep methods with step density control. Journal of Computational and Applied Mathematics, 205(2): 891–900, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004109.

Aslan:2010:CNE

[Asl10]  Ismail Aslan.  Comment

on: “New exact solutions for the Kawahara equation using Exp-function method” [J. Comput. Appl. Math. 233 (2009) 97–102]. Journal of Computational and Applied Mathematics, 234(12): 3213–3215, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042710002347.

See [Ass09b].

Assal:2009:PDO

[Ass09a]  Miloud Assal.  Pseudo-

differential operators associated with Laguerre hypergroups. Journal of Computational and Applied Mathematics, 233(3):617–620, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001034.

Assas:2009:NES

[Ass09b] Laila M. B. Assas. New exact solutions for the Kawahara equation using Exp-function method. Journal of Computational and Applied Mathematics, 233(2):97–102, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004038.

See comment [Asl10].

Aceto:2007:SPL

[AT07] L. Aceto and D. Trigiante.

The stability problem for linear multistep methods: Old and new results. Journal of Computational and Applied Mathematics, 210(1–2): 2–12, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600642X.

Abbasbandy:2009:NSS

[AT09a] S. Abbasbandy and A. Taati.

Numerical solution of the system of nonlinear Volterra integro-differential equations with nonlinear differential part by the operational Tau method and error estimation. Journal of Computational and Applied Mathematics, 231(1):106–113, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000442.

Ahipo:2009:RIS

[AT09b] Yves Marcel Ahipo and Philippe Traore. A robust iterative scheme for finite volume discretization of diffusive flux on highly skewed meshes. Journal of Computational and Applied Mathematics, 231(1):478–491, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002131.

Ariyawansa:2009:CCS

[AT09c] K. A. Ariyawansa and Wayne L. Tabor. A class of collinear scaling algorithms for boundconstrained optimization: Derivation and computational results. Journal of Computational and Applied Mathematics, 230(1): 143–163, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005955.

Assous:2009:PMS

[AT09d] F. Assous and F. Tsipis. A PIC method for solving a paraxial model of highly relativistic beams. Journal of Computational and Applied Mathematics, 227 (1):136–146, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003476.

Atanackovic:2007:OSS

[Ata07] Teodor M. Atanackovic. Optimal shape of a strongest inverted column. Journal of Computational and Applied Mathematics, 203 (1):209–218, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001956.

Attili:2005:NAS

[Att05] Basem S. Attili. A numerical algorithm for some singularly perturbed boundary value problems. Journal of Computational and Applied Mathematics, 184(2): 464–474, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000427.

Ahmad:2009:SGV

[AU09] Rais Ahmad and Farhat Usman. System of generalized variational inclusions with H-accretive operators in uniformly smooth Banach spaces. Journal of Computational and Applied Mathematics, 230(2):424–432, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006390.

Auer:2007:IMD

[Aue07]  Ekaterina Auer.  Interval

modeling of dynamics for multibody systems. Journal of Computational and Applied Mathematics, 199(2): 251–256, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007582.

Abramov:2006:CML

[AUY06] A. A. Abramov, V. I. Ul’yanova, and L. F. Yukhno. Certain methods for linear problems with inexact initial data. Journal of Computational and Applied Mathematics, 192(1):2–10, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500302X.

Afrouzi:2006:EIP

[AV06] G. A. Afrouzi and Javad Vahidi. Existence of an interval of principal eigenvalues for radially symmetric problems on all of R^{N}. Journal of Computational and Applied Mathematics, 187(2): 253–259, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001603.

Alvarez-Vazquez:2006:NOP

[AVMMS^{+}06] L. J. Alvarez-Vázquez, A. Martínez, R. Muñoz-Sola, C. Rodríguez, and M. E. Vázquez-Méndez. Numerical optimization for the purification of polluted shallow waters. Journal of Computational and Applied Mathematics, 189(1– 2):191–206, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003638.

Alvarez-Vazquez:2008:VSF

[AVMVMV08] L. J. Alvarez-Vázquez, A. Martínez, M. E. Vázquez-Méndez, and M. A. Vilar. Vertical slot fishways: Mathematical modeling and optimal management. Journal of Computational and Applied Mathematics, 218(2):395–403, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700204X.

Abdella:2009:ASM

[AYK09] K. Abdella, X. Yu, and I. Kucuk. Application of the sinc method to a dynamic elastoplastic problem. Journal of Computational and Applied Mathematics, 223(2): 626–645, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000551.

Al-Zanaidi:2006:ITM

[AZGN06] M. A. Al-Zanaidi, C. Grossmann, and A. Noack. Implicit Taylor methods for parabolic problems with nonsmooth data and applications to optimal heat control. Journal of Computational and Applied Mathematics, 188(1):121–149, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001834.

Bacuta:2009:SCH

[Bac09] Constantin Bacuta. Schur complements on Hilbert spaces and saddle point systems. Journal of Computational and Applied Mathematics, 225(2):581–593, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004305.

Badea:2008:SMI

[Bad08] Lori Badea. Schwarz methods for inequalities with contraction operators. Journal of Computational and Applied Mathematics, 215 (1):196–219, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001963.

Baker:2005:CEB

[Bak05] George A. Baker, Jr. Counterexamples to the Baker– Gammel–Wills conjecture and patchwork convergence. Journal of Computational and Applied Mathematics, 179(1–2):1–14, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400439X.

Banda:2005:VRS

[Ban05] Mapundi Kondwani Banda.

Variants of relaxed schemes and two-dimensional gas dynamics. Journal of Computational and Applied Mathematics, 175(1):41–62, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400247X.

Banjai:2007:EFD

[Ban07] Lehel Banjai. Eigenfrequencies of fractal drums. Journal of Computational and Applied Mathematics, 198 (1):1–18, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500703X.

Banas:2008:ATL

[Ban08]  L’ubomír Banas.  Adap-

tive techniques for Landau– Lifshitz–Gilbert equation with magnetostriction. Journal of Computational and Applied Mathematics, 215 (2):304–310, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007400.

Baratella:2009:NIS

[Bar09] Paola Baratella. A Nyström interpolant for some weakly singular linear Volterra integral equations. Journal of Computational and Applied Mathematics, 231(2): 725–734, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002726.

Batiha:2008:AAS

[Bat08]  Khaldoun Batiha.  Ap-

proximate analytical solution for the Zakharov–Kuznetsov equations with fully nonlinear dispersion. Journal of Computational and Applied Mathematics, 216(1): 157–163, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002257.

Baker:2005:ESM

[BB05] Christopher T. H. Baker and Evelyn Buckwar. Exponential stability in p-th mean of solutions, and of convergent Euler-type solutions, of stochastic delay differential equations. Journal of Computational and Applied Mathematics, 184(2): 404–427, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000397.

Bhatti:2007:SDE

[BB07] M. Idrees Bhatti and P. Bracken.

Solutions of differential equations in a Bernstein polynomial basis. Journal of Computational and Applied Mathematics, 205(1): 272–280, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003153.

Bourchtein:2008:AIS

[BB08a] Andrei Bourchtein and Ludmila Bourchtein. On the absolute instability of semiimplicit schemes for hydrostatic models. Journal of Computational and Applied Mathematics, 218(2): 404–420, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002075.

Bourchtein:2008:GGN

[BB08b] Ludmila Bourchtein and Andrei Bourchtein. On grid generation for numerical models of geophysical fluid dynamics. Journal of Computational and Applied Mathematics, 218(2):317–328, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000854.

Bourchtein:2009:SLS

[BB09a] Andrei Bourchtein and Ludmila Bourchtein. Semi-Lagrangian semi-implicit timesplitting scheme for a regional model of the atmosphere. Journal of Computational and Applied Mathematics, 227(1):115–125, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003348.

Bourchtein:2009:CDS

[BB09b] Ludmila Bourchtein and Andrei Bourchtein. Comparison of different spatial grids for numerical schemes of geophysical fluid dynamics. Journal of Computational and Applied Mathematics, 227(1):161–170, May 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800349X.

Binding:2007:PAA

[BBB07] Paul Binding, Lyonell Boulton, and Patrick J. Browne. A Prüfer angle approach to singular Sturm–Liouville problems with molcanov potentials. Journal of Computational and Applied Mathematics, 208(1):226–234, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042706006376.

Bailey:2007:BI

[BBC07] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Box integrals. Journal of Computational and Applied Mathematics, 206(1): 196–208, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004250.

Bertolazzi:2006:SNE

[BBD06] Enrico Bertolazzi, Francesco Biral, and Mauro Da Lio. Symbolic-numeric efficient solution of optimal control problems for multibody systems. Journal of Computational and Applied Mathematics, 185(2):404–421, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001238.

Baker:2005:CAP

[BBF^{+}05] C. T. H. Baker, G. A. Bocharov, J. M. Ford, P. M. Lumb, S. J. Norton, C. A. H. Paul, T. Junt, P. Krebs, and B. Ludewig. Computational approaches to parameter estimation and model selection in immunology. Journal of Computational and Applied Mathematics, 184(1): 50–76, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000609.

Baker:2007:SIE

[BBF07] Christopher T. H. Baker, Kevin Burrage, and Neville J. Ford. Special issue on evolutionary problems. Journal of Computational and Applied Mathematics, 205(2): 667–668, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003918.

Bocharov:2005:SIM

[BBG05] Gennady Bocharov, Hermann Brunner, and Zvi Grossman. Special issue on mathematics applied to immunology. Journal of Computational and Applied Mathematics, 184(1):1–3, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000579.

Bevilacqua:2007:CFN

[BBM07] R. Bevilacqua, E. Bozzo, and O. Menchi. Comparison of four natural pixel bases for SPECT imaging. Journal of Computational and Applied Mathematics, 198(2): 361–377, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500720X.

Bacinello:2009:PLI

[BBM09] Anna Rita Bacinello, Enrico Biffis, and Pietro Millossovich. Pricing life insurance contracts with early exercise features. Journal of Computational and Applied Mathematics, 233(1): 27–35, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002458.

Baker:2009:SAC

[BBPR09] Christopher T. H. Baker, Gennady Bocharov, Eugene Parmuzin, and Fathalla Rihan. Some aspects of causal & neutral equations used in modelling. Journal of Computational and Applied Mathematics, 229(2):335–349, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001556.

Batelo:2005:LCO

[BBR05] M. A. Batelo, C. F. Bracciali, and A. Sri Ranga. On linear combinations of L-orthogonal polynomials associated with distributions belonging to symmetric classes. Journal of Computational and Applied Mathematics, 179 (1–2):15–29, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004406.

Bisognin:2008:CSK

[BBSV08] Eleni Bisognin, Vanilde Bisognin, Mauricio Sepúlveda, and Octavio Vera. Coupled system of Korteweg–de Vries equations type in domains with moving boundaries. Journal of Computational and Applied Mathematics, 220(1–2): 290–321, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004529.

Bousson:2006:OAB

[BC06] K. Bousson and S. D. Correia. Optimization algorithm based on densification and dynamic canonical descent. Journal of Computational and Applied Mathematics, 191(2):269–279, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005558.

Berriochoa:2008:NSM

[BC08] E. Berriochoa and A. Cachafeiro.

Nodal systems with maximal domain of exactness for Gaussian quadrature formulas. Journal of Computational and Applied Mathematics, 212(2):272–281, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007291.

Bernard:2009:RII

[BC09a] Carole Bernard and An Chen.

On the regulator-insurer interaction in a structural model. Journal of Computational and Applied Mathematics, 233(1):3–15, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001908.

Berriochoa:2009:CCS

[BC09b] E. Berriochoa and A. Cachafeiro. Characterizing curves satisfying the Gauss–Christoffel theorem. Journal of Computational and Applied Mathematics, 233(3):630–633, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001058.

Baer:2009:NSC

[BCeAJ09] S. M. Baer, S. Crook, M. Dur e Ahmad, and Z. Jackiewicz. Numerical solution of calcium-mediated dendritic branch model. Journal of Computational and Applied Mathematics, 229 (2):416–424, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001623.

Berriochoa:2005:CBO

[BCGA05a] E. Berriochoa, A. Cachafeiro, and J. García-Amor. Connection between orthogonal polynomials on the unit circle and bounded interval. Journal of Computational and Applied Mathematics, 177 (1):205–223, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004248.

Berriochoa:2005:APC

[BCGA05b] E. Berriochoa, A. Cachafeiro, and J. M. Garcia-Amor. Asymptotic properties of Chebyshev–Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 178 (1–2):63–74, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003668.

Bartels:2006:PIF

[BCH06] S. Bartels, C. Carstensen, and A. Hecht. P2Q2Iso2D = 2D isoparametric FEM in Matlab. Journal of Computational and Applied Mathematics, 192(2):219–250, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002839.

Bourgeois:2007:LOF

[BCK07] L. Bourgeois, C. Chambeyron, and S. Kusiak. Locating an obstacle in a 3D finite depth ocean using the convex scattering support. Journal of Computational and Applied Mathematics, 204 (2):387–399, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003669.

Berg:2005:P

[BCP05] C. Berg, J. S. Christiansen, and H. L. Pedersen. Preface. Journal of Computational and Applied Mathematics, 178(1–2):xv, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003991.

Brignone:2007:ULS

[BCP07] M. Brignone, J. Coyle, and M. Piana. The use of the linear sampling method for obtaining super-resolution effects in Born approximation. Journal of Computational and Applied Mathematics, 203(1):145–158, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001798.

Beirlant:2006:P

[BCS^{+}06] Jan Beirlant, An Carbonez, Wim Schoutens, Walter Van Assche, and Noel Veraverbeke. Preface. Journal of Computational and Applied Mathematics, 186 (1):1–3, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003511.

Bultheel:2005:GOP

[BCV^{+}05] A. Bultheel, A. Cuyt, W. Van Assche, M. Van Barel, and B. Verdonk. Generalizations of orthogonal polynomials. Journal of Computational and Applied Mathematics, 179(1–2):57–95, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004443.

Borzi:2006:ALF

[BD06] A. Borzì and E. Decker.

Analysis of a leap-frog pseudospectral scheme for the Schrödinger equation. Journal of Computational and Applied Mathematics, 193 (1):65–88, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003808.

Bueno:2007:MAA

[BD07] María I. Bueno and Froilán M.

Dopico. A more accurate algorithm for computing the Christoffel transformation. Journal of Computational and Applied Mathematics, 205(1):567–582, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003396.

Brasil:2008:NAS

[BdAR08] A. P. Brasil, Jr., J. H. Carneiro de Araujo, and V. Ruas. A new algorithm for simulating viscoelastic flows accommodating piecewise linear finite elements. Journal of Computational and Applied Mathematics, 215(2):311–319, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007412.

Baker:2009:P

[BDF^{+}09] Christopher T. H. Baker, Teresa Diogo, George Flessas, Pedro Lima, Athena Makroglou, and Jie Shen. Preface. Journal of Computational and Applied Mathematics, 229 (2):333–334, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001544.

Bultheel:2005:PIQ

[BDGV05] Adhemar Bultheel, Leyla Daruis, and Pablo González-Vera. Positive interpolatory quadrature formulas and para-orthogonal polynomials. Journal of Computational and Applied Mathematics, 179(1–2):97–119, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004455.

Bultheel:2009:QFU

[BDGV09] Adhemar Bultheel, Leyla Daruis, and Pablo González-Vera. Quadrature formulas on the unit circle with prescribed nodes and maximal domain of validity. Journal of Computational and Applied Mathematics, 231(2): 948–963, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003240.

Benchama:2007:SFL

[BDHW07] N. Benchama, T. K. DeLillo, T. Hrycak, and L. Wang. A simplified Fornberg-like method for the conformal mapping of multiply connected regions — comparisons and crowding. Journal of Computational and Applied Mathematics, 209 (1):1–21, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006157.

Berlinska:2009:ESS

[BDL09] J. Berlińska, M. Drozdowski, and M. Lawenda. Experimental study of scheduling with memory constraints using hybrid methods. Journal of Computational and Applied Mathematics, 232(2): 638–654, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003999.

Boal:2008:APS

[BDS08] N. Boal, V. Domínguez, and F.-J. Sayas. Asymptotic properties of some triangulations of the sphere. Journal of Computational and Applied Mathematics, 211(1): 11–22, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006911.

Beirlant:2006:GFS

[BdWG06] Jan Beirlant, Tertius de Wet, and Yuri Goegebeur. A goodness-of-fit statistic for Pareto-type behaviour. Journal of Computational and Applied Mathematics, 186(1): 99–116, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001913.

Bai:2009:NSI

[BDY09] Zhong-Zhi Bai, Iain S. Duff, and Jun-Feng Yin. Numerical study on incomplete orthogonal factorization preconditioners. Journal of Computational and Applied Mathematics, 226(1):22–41, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002185.

Brown:2006:EKE

[BE06a] B. M. Brown and M. S. P.

Eastham. Extensions of the Kummer evaluation of a finite hypergeometric series. Journal of Computational and Applied Mathematics, 194(1): 131–140, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004073.

Brown:2006:NDF

[BE06b] B. M. Brown and M. S. P.

Eastham. A note on the Dixon formula for a finite hypergeometric series. Journal of Computational and Applied Mathematics, 194(1): 173–175, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004103.

Brown:2006:F

[BE06c] B. M. Brown and W. D.

Evans. Foreword. Journal of Computational and Applied Mathematics, 194 (1):1, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003961.

Brown:2006:L

[BE06d] B. M. Brown and W. D. Evans. Laudatum. Journal of Computational and Applied Mathematics, 194 (1):2, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003973.

Brown:2007:ERL

[BE07] B. M. Brown and M. S. P. Eastham. Eigenvalues of the radial p-Laplacian with a potential on (0, ∞). Journal of Computational and Applied Mathematics, 208(1): 111–119, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006303.

Bejancu:2006:SCI

[Bej06]  Aurelian Bejancu.  Semi-

cardinal interpolation and difference equations: From cubic B-splines to a threedirection box-spline construction. Journal of Computational and Applied Mathematics, 197(1):62–77, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006448.

Beliakov:2006:ILF

[Bel06] Gleb Beliakov. Interpolation of Lipschitz functions. Journal of Computational and Applied Mathematics, 196(1): 20–44, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005157.

Boutros:2006:LGM

[BeMBH06] Youssef Z. Boutros, Mina B. Abd el Malek, Nagwa A. Badran, and Hossam S. Hassan. Lie-group method for unsteady flows in a semi-infinite expanding or contracting pipe with injection or suction through a porous wall. Journal of Computational and Applied Mathematics, 197(2): 465–494, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006941.

Benney:2006:CLA

[Ben06] D. J. Benney. Conservation laws associated with long surface waves. Journal of Computational and Applied Mathematics, 190(1– 2):136–141, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002268.

Bennewitz:2007:ANS

[Ben07] Christer Bennewitz. Approximation numbers = singular values. Journal of Computational and Applied Mathematics, 208(1):102– 110, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006297.

Berg:2005:PBF

[Ber05] Christian Berg. Problem 1. Bernstein functions. Journal of Computational and Applied Mathematics, 178(1– 2):525–526, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004017.

Berggren:2006:VCD

[Ber06a] Martin Berggren. A vertexcentered, dual discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 192 (1):175–181, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003195.

Berrut:2006:CIE

[Ber06b] Jean-Paul Berrut. A circular interpretation of the Euler– Maclaurin formula. Journal of Computational and Applied Mathematics, 189(1– 2):375–386, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001408.

Berg:2007:PSM

[Ber07] Christian Berg. On powers of Stieltjes moment sequences, II. Journal of Computational and Applied Mathematics, 199(1):23–38, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007351.

Barbeiro:2005:SLF

[BF05] S. Barbeiro and J. A. Ferreira. A superconvergent linear FE approximation for the solution of an elliptic system of PDEs. Journal of Computational and Applied Mathematics, 177(2):287–300, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004261.

Belanger:2008:IRO

[BF08a] A. C. Bélanger and P. A.

Forsyth. Infinite reload options: Pricing and analysis. Journal of Computational and Applied Mathematics, 222(1):54–81, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005493.

Branco:2008:SPH

[BF08b] J. R. Branco and J. A. Ferreira. A singular perturbation of the heat equation with memory. Journal of Computational and Applied Mathematics, 218(2): 376–394, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001975.

Borzabadi:2009:NSC

[BF09] Akbar H. Borzabadi and Omid S. Fard. A numerical scheme for a class of nonlinear Fredholm integral equations of the second kind. Journal of Computational and Applied Mathematics, 232(2): 449–454, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900377X.

Bock:2006:DBK

[BFGM06] S. Bock, M. I. Falcão, K. Gürlebeck, and H. Malonek. A 3-dimensional Bergman kernel method with applications to rectangular domains. Journal of Computational and Applied Mathematics, 189(1–2):67–79, May 1, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003614.

Barrera:2008:MES

[BFGP08] D. Barrera, M. A. Fortes, P. González, and M. Pasadas. Minimal energy C^{r}-surfaces on uniform Powell–Sabintype meshes for noisy data. Journal of Computational and Applied Mathematics, 218(2):592–602, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000149.

Branquinho:2009:RHP

[BFM09] A. Branquinho, U. Fidalgo, and A. Foulquié Moreno. Riemann–Hilbert problem associated with Angelesco systems. Journal of Computational and Applied Mathematics, 233(3):643–651, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001071.

Boulaajine:2009:PEE

[BFP09] L. Boulaajine, M. Farhloul, and L. Paquet. A priori error estimation for the dual mixed finite element method of the elastodynamic problem in a polygonal domain, I. Journal of Computational and Applied Mathematics, 231(1): 447–472, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002210.

Baruch:2007:HON

[BFT07] G. Baruch, G. Fibich, and S. Tsynkov. High-order numerical solution of the nonlinear Helmholtz equation with axial symmetry. Journal of Computational and Applied Mathematics, 204 (2):477–492, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003748.

Barrios:2007:AMF

[BG07a] Tomás P. Barrios and Gabriel N. Gatica. An augmented mixed finite element method with Lagrange multipliers: a priori and a posteriori error analyses. Journal of Computational and Applied Mathematics, 200(2):653–676, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000422.

Bloch:2007:LAA

[BG07b] Anthony M. Bloch and Michael I. Gekhtman. Lie algebraic aspects of the finite nonperiodic Toda flows.

Journal of Computational and Applied Mathematics, 202(1):3–25, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001154.

Bourdarias:2007:FVS

[BG07c] C. Bourdarias and S. Gerbi.

A finite volume scheme for a model coupling free surface and pressurised flows in pipes. Journal of Computational and Applied Mathematics, 209(1):109–131, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006686.

Boyd:2007:NEA

[BG07d] John P. Boyd and Daniel H.

Gally. Numerical experiments on the accuracy of the Chebyshev–Frobenius companion matrix method for finding the zeros of a truncated series of Chebyshev polynomials. Journal of Computational and Applied Mathematics, 205(1): 281–295, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003177.

Brandes:2007:QSA

[BG07e] Kerstin Brandes and Roland Griesse. Quantitative stability analysis of optimal solutions in PDE-constrained optimization. Journal of Computational and Applied Mathematics, 206(2):908– 926, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005401.

Bruno:2007:IST

[BG07f] Oscar P. Bruno and Christophe A. Geuzaine. An O(1) integration scheme for threedimensional surface scattering problems. Journal of Computational and Applied Mathematics, 204(2): 463–476, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003736.

Bellen:2009:SND

[BG09] Alfredo Bellen and Nicola Guglielmi. Solving neutral delay differential equations with state-dependent delays. Journal of Computational and Applied Mathematics, 229(2):350–362, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001568.

Burenkov:2007:NSC

[BGG07] Viktor I. Burenkov, Huseyn V.

Guliyev, and Vagif S. Guliyev. Necessary and sufficient conditions for the boundedness of fractional maximal operators in local Morrey-type spaces. Journal of Computational and Applied Mathematics, 208(1):280–301, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006406.

Bourdarias:2008:KFM

[BGG08] C. Bourdarias, S. Gerbi, and M. Gisclon. A kinetic formulation for a model coupling free surface and pressurised flows in closed pipes. Journal of Computational and Applied Mathematics, 218(2): 522–531, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004815.

Bottcher:2007:MCQ

[BGGC07] Albrecht Böttcher, Jesús Gutiérrez-Gutiérrez, and Pedro M. Crespo. Mass concentration in quasicommutators of Toeplitz matrices. Journal of Computational and Applied Mathematics, 205(1): 129–148, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002603.

Berselli:2007:ACE

[BGJ07] Luigi C. Berselli, Carlo R. Grisanti, and Volker John. Analysis of commutation errors for functions with low regularity. Journal of Computational and Applied Mathematics, 206(2):1027– 1045, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005760.

Bergamaschi:2007:NES

[BGP07] L. Bergamaschi, G. Gambolati, and G. Pini. A numerical experimental study of inverse preconditioning for the parallel iterative solution to 3D finite element flow equations. Journal of Computational and Applied Mathematics, 210(1–2): 64–70, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006480.

Bogani:2009:GPS

[BGP09] C. Bogani, M. G. Gasparo, and A. Papini. Generalized pattern search methods for a class of nonsmooth optimization problems with structure. Journal of Computational and Applied Mathematics, 229(1):283–293, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005773.

Bayon:2005:NDE

[BGRS05] L. Bayón, J. M. Grau, M. M.

Ruiz, and P. M. Suárez. New developments on equivalent thermal in hydrothermal optimization: an algorithm of approximation. Journal of Computational and Applied Mathematics, 175(1):63–75, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002468.

Bayon:2006:NOH

[BGRS06] L. Bayón, J. M. Grau, M. M.

Ruiz, and P. M. Suárez. Nonsmooth optimization of hydrothermal problems. Journal of Computational and Applied Mathematics, 192 (1):11–19, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003031.

Barbon:2009:WBD

[BGV^{+}09] Sylvio Barbon, Jr., Rodrigo Capobianco Guido, Lucimar Sasso Vieira, Everthon Silva Fonseca, Fabrício Lopes Sanchez, Paulo Rogério Scalassara, Carlos Dias Maciel, José Carlos Pereira, and Shi-Huang Chen. Waveletbased dynamic time warping. Journal of Computational and Applied Mathematics, 227(2):271–287, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001106.

Bultheel:2005:ORF

[BGVHN05] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Orthogonal rational functions on the real half line with poles in [−∞, 0]. Journal of Computational and Applied Mathematics, 179(1–2):121–155, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004467.

Bultheel:2008:IRM

[BGVHN08] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. An indeterminate rational moment problem and Carathéodory functions. Journal of Computational and Applied Mathematics, 219(2):359–369, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042707002300.

Bastani:2007:NAR

[BH07a] A. Foroush Bastani and S. Mohammad Hosseini. A new adaptive Runge–Kutta method for stochastic differential equations. Journal of Computational and Applied Mathematics, 206(2): 631–644, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005103.

Boglaev:2007:UCW

[BH07b] Igor Boglaev and Matthew Hardy. Uniform convergence of a weighted average scheme for a nonlinear reaction–diffusion problem. Journal of Computational and Applied Mathematics, 200(2):705–721, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000689.

Burman:2007:USM

[BH07c] Erik Burman and Peter Hansbo. A unified stabilized method for Stokes’ and Darcy’s equations. Journal of Computational and Applied Mathematics, 198 (1):35–51, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007053.

Bastani:2009:MSS

[BH09] A. Foroush Bastani and S. Mohammad Hosseini. On mean-square stability properties of a new adaptive stochastic Runge–Kutta method. Journal of Computational and Applied Mathematics, 224(2):556–564, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002628.

Bonan-Hamada:2005:SCF

[BHJ05] Catherine M. Bonan-Hamada and William B. Jones. Stieltjes continued fractions for polygamma functions; speed of convergence. Journal of Computational and Applied Mathematics, 179(1– 2):47–55, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400442X.

Buffa:2007:DGC

[BHP07] Annalisa Buffa, Paul Houston, and Ilaria Perugia. Discontinuous Galerkin computation of the Maxwell eigenvalues on simplicial meshes. Journal of Computational and Applied Mathematics, 204(2):317–333, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003608.

Bi:2008:SMC

[Bi08a] Chunjia Bi. Superconvergence of mixed covolume method for elliptic problems on triangular grids. Journal of Computational and Applied Mathematics, 216 (2):534–544, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002932.

Buyukyazici:2008:DCR

[BI08b] Ibrahim Büyükyazici and Ertan Ibikli. Direct and converse results for multivariate generalized Bernstein polynomials. Journal of Computational and Applied Mathematics, 219(1): 145–155, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700372X.

Bingham:2007:RVP

[Bin07]  N. H. Bingham.  Regu-

lar variation and probability: the early years. Journal of Computational and Applied Mathematics, 200 (1):357–363, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000203.

Borggaard:2009:IPM

[BIR09] Jeff Borggaard, Traian Iliescu, and John Paul Roop. An improved penalty method for power-law Stokes problems. Journal of Computational and Applied Mathematics, 223(2):646–658, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000563.

Barrera:2005:NBQ

[BISS05a] D. Barrera, M. J. Ibáñez, P. Sablonnière, and D. Sbibih. Near-best quasi-interpolants associated with H-splines on a three-direction mesh. Journal of Computational and Applied Mathematics, 183(1): 133–152, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000208.

Barrera:2005:NMN

[BISS05b] D. Barrera, M. J. Ibáñez, P. Sablonnière, and D. Sbibih. Near minimally normed spline quasi-interpolants on uniform partitions. Journal of Computational and Applied Mathematics, 181(1): 211–233, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005813.

Bojovic:2006:CFD

[BJ06] Dejan Bojović and Bosko S.

Jovanović. Convergence of finite difference method for the parabolic problem with concentrated capacity and variable operator. Journal of Computational and Applied Mathematics, 189(1– 2):286–303, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500138X.

Bartoszewski:2007:DCE

[BJ07a] Z. Bartoszewski and Z. Jackiewicz. Derivation of continuous explicit two-step Runge– Kutta methods of order three. Journal of Computational and Applied Mathematics, 205(2):764–776, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004006.

Bouhamidi:2007:STR

[BJ07b] A. Bouhamidi and K. Jbilou.

Sylvester Tikhonov-regularization methods in image restoration. Journal of Computational and Applied Mathematics, 206(1):86–98, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003839.

Bartoszewski:2008:CHS

[BJ08] Z. Bartoszewski and Z. Jackiewicz. Construction of highly stable parallel two-step Runge–Kutta methods for delay differential equations. Journal of Computational and Applied Mathematics, 220(1–2):257–270, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004505.

Baker:2009:SSV

[BJK09] A. H. Baker, E. R. Jessup, and Tz. V. Kolev. A simple strategy for varying the restart parameter in GMRES(m). Journal of Computational and Applied Mathematics, 230(2): 751–761, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000132.

Jian:2007:FGM

[bJQlZ07] Jin bao Jian, Ran Quan, and Xue lu Zhang. Feasible generalized monotone line search SQP algorithm for nonlinear minimax problems with inequality constraints. Journal of Computational and Applied Mathematics, 205(1):406–429, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600327X.

Bellalij:2008:NCR

[BJS08] M. Bellalij, K. Jbilou, and H. Sadok. New convergence results on the global GMRES method for diagonalizable matrices. Journal of Computational and Applied Mathematics, 219(2): 350–358, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005146.

Bouyouli:2006:CPS

[BJSS06] R. Bouyouli, K. Jbilou, R. Sadaka, and H. Sadok. Convergence properties of some block Krylov subspace methods for multiple linear systems. Journal of Computational and Applied Mathematics, 196(2): 498–511, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006011.

Jian:2009:SCS

[bJxC09] Jin bao Jian and Wei xin Cheng. A superlinearly convergent strongly subfeasible SSLE-type algorithm with working set for nonlinearly constrained optimization. Journal of Computational and Applied Mathematics, 225(1):172–186, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003610.

Jian:2009:MCN

[bJyKyZmT09] Jin bao Jian, Xiao yan Ke, Hai yan Zheng, and Chun ming Tang. A method combining norm-relaxed QP subproblems with systems of linear equations for constrained optimization. Journal of Computational and Applied Mathematics, 223(2): 1013–1027, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001428.

Burstedde:2006:FIS

[BK06] Carsten Burstedde and Angela Kunoth. Fast iterative solution of elliptic control problems in wavelet discretization. Journal of Computational and Applied Mathematics, 196(1):299– 319, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005650.

Buhmiler:2008:NSQ

[BK08] Sandra Buhmiler and Natasa Krejić. A new smoothing quasi-Newton method for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 211(2): 141–155, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007011.

Bas:2009:FCR

[BK09a] E. Bas and C. Kahraman. Fuzzy capital rationing model. Journal of Computational and Applied Mathematics, 224(2):628–645, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800280X.

Borhanifar:2009:NPS

[BK09b] A. Borhanifar and M. M.

Kabir. New periodic and soliton solutions by application of exp-function method for nonlinear evolution equations. Journal of Computational and Applied Mathematics, 229(1):158–167, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005384.

Burgarelli:2006:NAM

[BKB06] Denise Burgarelli, Mauricio Kischinhevsky, and Rodney Josué Biezuner. A new adaptive mesh refinement strategy for numerically solving evolutionary PDE’s. Journal of Computational and Applied Mathematics, 196(1):115–131, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005212.

Bryan:2007:IML

[BKT07] Kurt Bryan, Rachel Krieger, and Nic Trainor. Imaging of multiple linear cracks using impedance data. Journal of Computational and Applied Mathematics, 200 (1):388–407, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000239.

Blatt:2005:IAP

[Bla05] Hans-Peter Blatt. The interaction of alternation points and poles in rational approximation. Journal of Computational and Applied Mathematics, 179(1–2):31–46, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004418.

Brown:2007:P

[BLL07] B. M. Brown, J. Lang, and R. T. Lewis. Preface. Journal of Computational and Applied Mathematics, 208 (1):1–2, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006224.

Bruti-Liberati:2007:SAS

[BLP07] Nicola Bruti-Liberati and Eckhard Platen. Strong approximations of stochastic differential equations with jumps. Journal of Computational and Applied Mathematics, 205(2):982–1001, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004183.

Baguelin:2007:HDP

[BLR07] Marc Baguelin, Jacques LeFèvre, and Jean-Pierre Richard. How to deal with potentially huge dimensional state space: the meta-dynamics approachapplication to a model of the co-evolution of bacteriophage populations. Journal of Computational and Applied Mathematics, 205(2): 687–695, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003931.

Benner:2009:AMS

[BLT09] Peter Benner, Ren-Cang Li, and Ninoslav Truhar. On the ADI method for Sylvester equations. Journal of Computational and Applied Mathematics, 233 (4):1035–1045, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709006050.

Bao:2006:NSN

[BLW06] Gang Bao, Ying Li, and Haijun Wu. Numerical solution of nonlinear diffraction problems. Journal of Computational and Applied Mathematics, 190(1–2):170–189, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002293.

Berenhaut:2006:SOB

[BM06] Kenneth S. Berenhaut and Daniel C. Morton. Secondorder bounds for linear recurrences with negative coefficients. Journal of Computational and Applied Mathematics, 186(2):504–522, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001299.

Beltran:2007:CQH

[BM07a] J. V. Beltran and J. Monterde. A characterization of quintic helices. Journal of Computational and Applied Mathematics, 206(1): 116–121, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003852.

Brugnano:2007:BIM

[BM07b] Luigi Brugnano and Cecilia Magherini. Blended implicit methods for solving ODE and DAE problems, and their extension for second-order problems. Journal of Computational and Applied Mathematics, 205(2):777–790, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004018.

Babolian:2008:DMS

[BM08a] E. Babolian and Z. Masouri. Direct method to solve Volterra integral equation of the first kind using operational matrix with block-pulse functions. Journal of Computational and Applied Mathematics, 220(1– 2):51–57, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004153.

Benes:2008:CMH

[BM08b] Michal Benes and Petr Mayer. Coupled model of hygrothermal behavior of concrete during fire. Journal of Computational and Applied Mathematics, 218(1): 12–20, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003226.

Bountis:2009:AGM

[BMC09] T. Bountis, T. Manos, and H. Christodoulidi. Application of the GALI method to localization dynamics in nonlinear systems. Journal of Computational and Applied Mathematics, 227 (1):17–26, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003269.

Bai:2007:LCI

[BMfX07] Zheng-Jian Bai, Benedetta Morini, and Shu fang Xu. On the local convergence of an iterative approach for inverse singular value problems. Journal of Computational and Applied Mathematics, 198(2):344–360, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007193.

Brugnano:2006:BIM

[BMM06] Luigi Brugnano, Cecilia Magherini, and Filippo Mugnai. Blended implicit methods for the numerical solution of DAE problems. Journal of Computational and Applied Mathematics, 189 (1–2):34–50, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003316.

Barrio:2005:GEOa

[BMS05] R. Barrio, B. Melendo, and S. Serrano. Generation and evaluation of orthogonal polynomials in discrete Sobolev spaces. I: algorithms. Journal of Computational and Applied Mathematics, 181(2): 280–298, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005850.

Beckermann:2008:RGP

[BMW08] Bernhard Beckermann, Ana C.

Matos, and Franck Wielonsky. Reduction of the Gibbs phenomenon for smooth functions with jumps by the ϵ-algorithm. Journal of Computational and Applied Mathematics, 219(2): 329–349, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006206.

Babuska:2008:BVP

[BN08a] Ivo Babuska and Victor Nistor. Boundary value problems in spaces of distributions on smooth and polygonal domains. Journal of Computational and Applied Mathematics, 218(1):137–148, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003354.

Banas:2008:AFE

[BN08b] L’ubomír Banas and Robert Nürnberg. Adaptive finite element methods for Cahn– Hilliard equations. Journal of Computational and Applied Mathematics, 218 (1):2–11, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003214.

Braysy:2009:OAC

[BNDN09] Olli Bräysy, Pentti Nakari, Wout Dullaert, and Pekka Neittaanmäki. An optimization approach for communal home meal delivery service: a case study. Journal of Computational and Applied Mathematics, 232 (1):46–53, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005438.

Bnouhachem:2007:APS

[Bno07a] Abdellah Bnouhachem. An additional projection step to He and Liao’s method for solving variational inequalities. Journal of Computational and Applied Mathematics, 206(1):238– 250, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004389.

Bnouhachem:2007:IIM

[Bno07b] Abdellah Bnouhachem. An inexact implicit method for general mixed variational inequalities. Journal of Computational and Applied Mathematics, 200(1):377– 387, March 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000227.

Barnard:2006:UWI

[BNP06] R. W. Barnard, S. Naik, and S. Ponnusamy. Univalency of weighted integral transforms of certain functions. Journal of Computational and Applied Mathematics, 193(2): 638–651, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500436X.

Bretti:2007:NAS

[BNP07] G. Bretti, R. Natalini, and B. Piccoli. Numerical algorithms for simulations of a traffic model on road networks. Journal of Computational and Applied Mathematics, 210(1–2):71–77, December 31, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006492.

Boulaajine:2008:DMF

[BNPR08] L. Boulaajine, S. Nicaise, L. Paquet, and Rafilipojaona. Dual mixed finite element methods for the elasticity problem with Lagrange multipliers. Journal of Computational and Applied Mathematics, 221(1): 234–260, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005936.

Babuska:2008:GFE

[BNT08] Ivo Babuska, Victor Nistor, and Nicolae Tarfulea. Generalized finite element method for second-order elliptic operators with Dirichlet boundary conditions. Journal of Computational and Applied Mathematics, 218(1): 175–183, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003391.

Boglaev:2005:BMD

[Bog05a] Igor Boglaev. A block monotone domain decomposition algorithm for a semilinear convection–diffusion problem. Journal of Computational and Applied Mathematics, 173(2):259–277, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001566.

Boglaev:2005:MSI

[Bog05b]  Igor Boglaev.  Monotone

Schwarz iterates for a semilinear parabolic convection– diffusion problem. Journal of Computational and Applied Mathematics, 183(1): 191–209, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000257.

Borzi:2007:HOD

[Bor07] A. Borzi‘. High-order discretization and multigrid solution of elliptic nonlinear constrained optimal control problems. Journal of Computational and Applied Mathematics, 200(1):67–85, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007971.

Boubendir:2007:ABF

[Bou07a] Yassine Boubendir. An analysis of the BEM–FEM nonoverlapping domain decomposition method for a scattering problem. Journal of Computational and Applied Mathematics, 204(2):282–291, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003578.

Bouhamidi:2007:EES

[Bou07b] A. Bouhamidi. Error estimates in Sobolev spaces for interpolating thin plate splines under tension. Journal of Computational and Applied Mathematics, 200 (1):208–216, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500806X.

Boyd:2009:CEI

[Boy09] John P. Boyd. Chebyshev expansion on intervals with branch points with application to the root of Kepler’s equation: a Chebyshev– Hermite–Padé method. Journal of Computational and Applied Mathematics, 223(2): 693–702, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000678.

Baker:2005:IIF

[BP05] C. T. H. Baker and E. I. Parmuzin. Identification of the initial function for discretized delay differential equations. Journal of Computational and Applied Mathematics, 181(2):420–441, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005941.

Bacchelli:2006:IDU

[BP06] Silvia Bacchelli and Serena Papi. Image denoising using principal component analysis in the wavelet domain. Journal of Computational and Applied Mathematics, 189(1– 2):606–621, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002815.

Bacchelli:2007:SBM

[BP07] S. Bacchelli and S. Papi.

Statistically based multiwavelet denoising. Journal of Computational and Applied Mathematics, 210(1–2):

47–55, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006467.

Bojanov:2009:QFF

[BP09] Borislav Bojanov and Guergana Petrova. Quadrature formulas for Fourier coefficients. Journal of Computational and Applied Mathematics, 231(1):378– 391, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001836.

Baglama:2007:DML

[BR07a] James Baglama and Lothar Reichel. Decomposition methods for large linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 198(2):332–343, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007181.

Bi:2007:UCF

[BR07b] Chunjia Bi and Hongxing Rui. Uniform convergence of finite volume element method with Crouzeix–Raviart element for non-self-adjoint and indefinite elliptic problems. Journal of Computational and Applied Mathematics, 200(2):555–565, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000355.

Branquinho:2009:DEL

[BR09] A. Branquinho and M. N. Rebocho. Distributional equation for Laguerre–Hahn functionals on the unit circle. Journal of Computational and Applied Mathematics, 233(3):634–642, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900106X.

Bracken:2007:GWS

[Bra07a] Paul Bracken. The generalized Weierstrass system inducing surfaces of constant and nonconstant mean curvature in Euclidean three space. Journal of Computational and Applied Mathematics, 202(1):122–132, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001105.

Bratsos:2007:STD

[Bra07b] A. G. Bratsos. The solution of the two-dimensional sine-Gordon equation using the method of lines. Journal of Computational and Applied Mathematics, 206(1): 251–277, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600447X.

Barrera-Rosillo:2009:MQI

[BRIP09] Domingo Barrera-Rosillo and María José Ibáñez-Pérez. Minimizing the quasi-interpolation error for bivariate discrete quasi-interpolants. Journal of Computational and Applied Mathematics, 224(1): 250–268, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002069.

Bai:2009:P

[BRS09] Zhong-Zhi Bai, Lothar Reichel, and Zhong-Ci Shi. Preface. Journal of Computational and Applied Mathematics, 226(1):1, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800215X.

Brumen:2008:FAN

[Bru08] Gorazd Brumen. A fast algorithm for numerical solutions to Fortet’s equation. Journal of Computational and Applied Mathematics, 220(1–2): 574–587, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004888.

Brunner:2009:RAN

[Bru09] Hermann Brunner. Recent advances in the numerical analysis of Volterra functional differential equations with variable delays. Journal of Computational and Applied Mathematics, 228 (2):524–537, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001349.

Backofen:2007:MAM

[BRVWM07] Rainer Backofen, Angel Ribalta, Axel Voigt, and Dirk Wulff-Molder. BaF_{2} for microlithography applications: modeling, simulation and optimization of the crystal growth process. Journal of Computational and Applied Mathematics, 203 (2):362–375, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002299.

Bi:2009:TSI

[BRW09] Weihong Bi, Hongmin Ren, and Qingbiao Wu. Threestep iterative methods with eighth-order convergence for solving nonlinear equations.

Journal of Computational and Applied Mathematics, 225(1):105–112, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003555.

Barrio:2005:GEOb

[BS05] R. Barrio and S. Serrano. Generation and evaluation of orthogonal polynomials in discrete Sobolev spaces II: numerical stability. Journal of Computational and Applied Mathematics, 181(2): 299–320, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005862.

Barrieu:2006:IIG

[BS06] Pauline Barrieu and Wim Schoutens. Iterates of the infinitesimal generator and space-time harmonic polynomials of a Markov process. Journal of Computational and Applied Mathematics, 186(1):300–323, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002517.

Bessaih:2007:UBR

[BS07] H. Bessaih and H. Schurz. Upper bounds on the rate of convergence of truncated stochastic infinitedimensional differential systems with H-regular noise. Journal of Computational and Applied Mathematics, 208(2):354–361, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005887.

Babolian:2008:NSV

[BS08a] E. Babolian and A. Salimi Shamloo. Numerical solution of Volterra integral and integro-differential equations of convolution type by using operational matrices of piecewise constant orthogonal functions. Journal of Computational and Applied Mathematics, 214(2):495–508, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001513.

Berridge:2008:IGA

[BS08b] S. J. Berridge and J. M. Schumacher. An irregular grid approach for pricing highdimensional American options. Journal of Computational and Applied Mathematics, 222(1):94–111, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005511.

Babolian:2009:NSN

[BS09a] E. Babolian and A. Shahsavaran. Numerical solution of nonlinear Fredholm integral equations of the second kind using Haar wavelets. Journal of Computational and Applied Mathematics, 225(1):87–95, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003221.

Bonettini:2009:NNC

[BS09b] Silvia Bonettini and Thomas Serafini. Non-negatively constrained image deblurring with an inexact interior point method. Journal of Computational and Applied Mathematics, 231(1): 236–248, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000570.

Boonchari:2009:WSC

[BS09c] Daruni Boonchari and Satit Saejung. Weak and strong convergence theorems of an implicit iteration for a countable family of continuous pseudocontractive mappings. Journal of Computational and Applied Mathematics, 233(4):1108–1116, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006293.

Basto:2007:DSS

[BSC07a] Mário Basto, Viriato Semiao, and Francisco Calheiros. Dynamics in spectral solutions of Burgers equation. Journal of Computational and Applied Mathematics, 205(1): 296–304, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003189.

Basto:2007:NSM

[BSC07b] Mário Basto, Viriato Semiao, and Francisco L. Calheiros. Numerical study of modified Adomian’s method applied to Burgers equation. Journal of Computational and Applied Mathematics, 206(2): 927–949, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005413.

Basto:2009:DSN

[BSC09] Mário Basto, Viriato Semiao, and Francisco Calheiros. Dynamics and synchronization of numerical solutions of the Burgers equation. Journal of Computational and Applied Mathematics, 231(2): 793–806, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002921.

Bujurke:2007:WMA

[BSKS07] N. M. Bujurke, C. S. Salimath, Ramesh B. Kudenatti, and S. C. Shiralashetti. Wavelet-multigrid analysis of squeeze film characteristics of poroelastic bearings. Journal of Computational and Applied Mathematics, 203 (1):237–248, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001993.

Bayon:2009:IFE

[BSMT09] L. Bayón, P. Suárez, J. M. Matías, and J. Taboada. Influence of forecasting electricity prices in the optimization of complex hydrothermal systems. Journal of Computational and Applied Mathematics, 232(2):262–274, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003537.

Bujurke:2008:CES

[BSS08] N. M. Bujurke, C. S. Salimath, and S. C. Shiralashetti. Computation of eigenvalues and solutions of regular Sturm–Liouville problems using Haar wavelets. Journal of Computational and Applied Mathematics, 219(1): 90–101, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003676.

Bujurke:2009:AST

[BSS09] N. M. Bujurke, S. C. Shiralashetti, and C. S. Salimath. An application of single-term Haar wavelet series in the solution of nonlinear oscillator equations. Journal of Computational and Applied Mathematics, 227 (2):234–244, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001076.

Barrera:2008:RMC

[BSST08] D. Barrera, D. Sbibih, A. Serghini, and A. Tijini. A recursive method for computing interpolants. Journal of Computational and Applied Mathematics, 216 (2):435–450, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002749.

Borzi:2008:FNS

[BSV08] A. Borzi‘, J. Salomon, and S. Volkwein. Formulation and numerical solution of finitelevel quantum optimal control problems. Journal of Computational and Applied Mathematics, 216(1):170–197, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002270.

Barbatis:2006:CRI

[BT06] G. Barbatis and A. Tertikas. On a class of rellich inequalities. Journal of Computational and Applied Mathematics, 194(1): 156–172, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004097.

Burmen:2009:UDF

[BT09] Árpád Bűrmen and Tadej Tuma. Unconstrained derivativefree optimization by successive approximation. Journal of Computational and Applied Mathematics, 223 (1):62–74, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006681.

Belward:2008:DES

[BTI08] John A. Belward, Ian W. Turner, and Milos Ilić. On derivative estimation and the solution of least squares problems. Journal of Computational and Applied Mathematics, 222(2):511–523, December 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006280.

Bozkaya:2007:FSC

[BTS07] C. Bozkaya and M. Tezer-Sezgin. Fundamental solution for coupled magnetohydrodynamic flow equations. Journal of Computational and Applied Mathematics, 203(1):125–144, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001786.

Benito:2007:SPH

[BUG07] J. J. Benito, F. Ureña, and L. Gavete. Solving parabolic and hyperbolic equations by the generalized finite difference method. Journal of Computational and Applied Mathematics, 209(2): 208–233, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600687X.

Binding:2006:IOS

[BV06] P. A. Binding and H. Volkmer. Interlacing and oscillation for Sturm–Liouville problems with separated and coupled boundary conditions.

Journal of Computational and Applied Mathematics, 194(1):75–93, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004048.

Bai:2006:RPC

[BW06] Zhong-Zhi Bai and Zeng-Qi Wang. Restrictive preconditioners for conjugate gradient methods for symmetric positive definite linear systems. Journal of Computational and Applied Mathematics, 187(2):202–226, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001536.

Breuss:2007:SSS

[BW07a] M. Breuß and M. Welk. Staircasing in semidiscrete stabilised inverse linear diffusion algorithms. Journal of Computational and Applied Mathematics, 206(1): 520–533, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005024.

Buckwar:2007:ILM

[BW07b] Evelyn Buckwar and Renate Winkler. Improved linear multi-step methods for stochastic ordinary differential equations. Journal of Computational and Applied Mathematics, 205(2): 912–922, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004122.

Bai:2009:RLN

[BXC09] Honghuan Bai, Aimin Xu, and Feng Cui. Representation for the Lagrangian numerical differentiation formula involving elementary symmetric functions. Journal of Computational and Applied Mathematics, 231(2): 907–913, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003203.

Bertsekas:2009:PEM

[BY09] Dimitri P. Bertsekas and Huizhen Yu. Projected equation methods for approximate solution of large linear systems. Journal of Computational and Applied Mathematics, 227(1):27–50, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003270.

Ben-yu:2005:NGLa

[ByXy05a] Guo Ben-yu and Zhang Xiaoyong. A new generalized Laguerre spectral approximation and its applications. Journal of Computational and Applied Mathematics, 181(2):342–363, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005904.

Ben-yu:2005:NGLb

[ByXy05b] Guo Ben-yu and Zhang Xiaoyong. A new generalized Laguerre spectral approximation and its applications. Journal of Computational and Applied Mathematics, 184(2):382–403, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000385.

See erratum [ByXy06].

Ben-yu:2006:ENG

[ByXy06] Guo Ben-yu and Zhang Xiaoyong. Erratum to “A new generalized Laguerre spectral approximation and its applications” [J. Comput. Appl. Math. 184 (2005) 382– 403]. Journal of Computational and Applied Mathematics, 196(1):211, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005583.

See [ByXy05b].

Ceng:2009:ISE

[CAHAY09] L.-C. Ceng, S. Al-Homidan, Q. H. Ansari, and J.-C. Yao. An iterative scheme for equilibrium problems and fixed point problems of strict pseudo-contraction mappings. Journal of Computational and Applied Mathematics, 223(2):967–974, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042708001179.

Cakoni:2007:RDQ

[Cak07] Fioralba Cakoni. Recent developments in the qualitative approach to inverse electromagnetic scattering theory. Journal of Computational and Applied Mathematics, 204(2):242–255, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003542.

Calvetti:2007:PIM

[Cal07] Daniela Calvetti. Preconditioned iterative methods for linear discrete ill-posed problems from a Bayesian inversion perspective. Journal of Computational and Applied Mathematics, 198(2): 378–395, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007211.

Cameron:2008:ERR

[Cam08] Frank Cameron. Enumeration results for Runge– Kutta methods for index 2 DAEs. Journal of Computational and Applied Mathematics, 213(1):294–299, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007874.

Cao:2008:LCP

[Cao08] Jinlong Cao. Limit cycles of polynomial differential systems with homogeneous nonlinearities of degree 4 via the averaging method. Journal of Computational and Applied Mathematics, 220(1–2): 624–631, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004918.

Carlson:2005:JEF

[Car05a] B. C. Carlson. Jacobian elliptic functions as inverses of an integral. Journal of Computational and Applied Mathematics, 174(2):355–359, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002201.

Caruntu:2005:SAD

[Car05b] Dumitru I. Caruntu. Selfadjoint differential equations for classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 180(1):107–118, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004807.

Carletti:2006:NSS

[Car06] M. Carletti. Numerical solution of stochastic differential problems in the biosciences. Journal of Computational and Applied Mathematics, 185(2):422–440, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500124X.

Carfora:2007:ISG

[Car07] Maria Francesca Carfora.

Interpolation on spherical geodesic grids: a comparative study. Journal of Computational and Applied Mathematics, 210(1–2):99– 105, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006522.

Cash:2005:ETI

[Cas05] J. R. Cash. Efficient time integrators in the numerical method of lines. Journal of Computational and Applied Mathematics, 183(2): 259–274, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000646.

Casas:2007:NNI

[Cas07] Fernando Casas. New numerical integrators based on solvability and splitting. Journal of Computational and Applied Mathematics, 205(2): 802–813, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004031.

Ceng:2009:RVI

[CAY09] L.-C. Ceng, Q. H. Ansari, and J. C. Yao. On relaxed viscosity iterative methods for variational inequalities in Banach spaces. Journal of Computational and Applied Mathematics, 230(2):813–822, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000351.

Carletti:2007:NDI

[CB07] Margherita Carletti and Edoardo Beretta. Numerical detection of instability regions for delay models with delay-dependent parameters. Journal of Computational and Applied Mathematics, 205(2):835–848, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004067.

Cruz-Barroso:2007:SOLa

[CBDGVN07a] Ruymán Cruz-Barroso, Leyla Daruis, Pablo González-Vera, and Olav Njåstad. Sequences of orthogonal Laurent polynomials, bi-orthogonality and quadrature formulas on the unit circle. Journal of Computational and Applied Mathematics, 200(1): 424–440, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000252.

Cruz-Barroso:2007:SOLb

[CBDGVN07b] Ruymán Cruz-Barroso, Leyla Daruis, Pablo González-Vera, and Olav Njåstad. Sequences of orthogonal Laurent polynomials, bi-orthogonality and quadrature formulas on the unit circle. Journal of Computational and Applied Mathematics, 206(2):950– 966, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005607.

Cruz-Barroso:2005:CDF

[CBGV05] Ruymán Cruz-Barroso and Pablo González-Vera. A Christoffel–Darboux formula and a Favard’s theorem for orthogonal Laurent polynomials on the unit circle. Journal of Computational and Applied Mathematics, 179(1– 2):157–173, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004479.

Csendes:2007:TCA

[CBH07] Tibor Csendes, Balázs Bánhelyi, and László Hatvani. Towards a computer-assisted proof for chaos in a forced damped pendulum equation. Journal of Computational and Applied Mathematics, 199(2): 378–383, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007752.

Cheikh:2005:CPL

[CC05] Y. Ben Cheikh and H. Chaggara. Connection problems via lowering operators. Journal of Computational and Applied Mathematics, 178 (1–2):45–61, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003656.

Capobianco:2006:EFP

[CC06a] Giovanni Capobianco and Dajana Conte. An efficient and fast parallel method for Volterra integral equations of Abel type. Journal of Computational and Applied Mathematics, 189(1– 2):481–493, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001627.

Colin:2006:NMR

[CC06b] M. Colin and T. Colin. A numerical model for the Raman amplification for laserplasma interaction. Journal of Computational and Applied Mathematics, 193(2): 535–562, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004309.

Chinosi:2007:AFG

[CC07] Claudia Chinosi and Lucia Della Croce. Approximation of functionally graded plates with non-conforming finite elements. Journal of Computational and Applied Mathematics, 210(1–2):106– 115, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006534.

Capobianco:2008:PAL

[CC08] G. Capobianco and A. Cardone. A parallel algorithm for large systems of Volterra integral equations of Abel type. Journal of Computational and Applied Mathematics, 220(1–2):749–758, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002550.

Calabro:2009:BBC

[CC09] F. Calabrò and G. Capobianco. Blowing up behavior for a class of nonlinear VIEs connected with parabolic PDEs. Journal of Computational and Applied Mathematics, 228(2):580–588, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001362.

Chang:2009:AMS

[CCCH09] S.-L. Chang, H.-S. Chen, C.- S. Chien, and D. J. Han. An adaptive multigrid scheme for Bose–Einstein condensates in a periodic potential. Journal of Computational and Applied Mathematics, 231(1): 268–287, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000612.

Capobianco:2009:HPP

[CCD09] Giovanni Capobianco, Dajana Conte, and Ida Del Prete. High performance parallel numerical methods for Volterra equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 228 (2):571–579, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001374.

Chen:2006:SRH

[CCH06] Jong-Yi Chen, Yunshyong Chow, and June Hsieh. Some results on a heat conduction problem by myshkis. Journal of Computational and Applied Mathematics, 190(1– 2):190–199, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500230X.

Chang:2007:EAS

[CCJ07] S.-L. Chang, C.-S. Chien, and B.-W. Jeng. An efficient algorithm for the Schrödinger– Poisson eigenvalue problem. Journal of Computational and Applied Mathematics, 205(1):509–532, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003359.

Chryssoverghi:2006:DMO

[CCK06] I. Chryssoverghi, I. Coletsos, and B. Kokkinis. Discretization methods for optimal control problems with state constraints. Journal of Computational and Applied Mathematics, 191(1):1–31, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002591.

Connell:2006:EFS

[CCLS06] Cameron R. Connell, Russel E. Caflisch, Erding Luo, and Geoff Simms. The elastic field of a surface step: the Marchenko–Parshin formula in the linear case. Journal of Computational and Applied Mathematics, 196(2): 368–386, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005704.

Cakoni:2007:IES

[CCM07] Fioralba Cakoni, David Colton, and Peter Monk. The inverse electromagnetic scattering problem for a partially coated dielectric. Journal of Computational and Applied Mathematics, 204 (2):256–267, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003554.

Chun:2005:OOE

[CCS05] Changbum Chun, Q-Heung Choi, and Michael W. Smiley. Optimal order error estimates for finite element approximations of a bifurcation function. Journal of Computational and Applied Mathematics, 181(2): 404–419, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400593X.

Chen:2009:CMD

[CCW^{+}09] Xiao-Diao Chen, Linqiang Chen, Yigang Wang, Gang Xu, Jun-Hai Yong, and Jean-Claude Paul. Computing the minimum distance between two Bézier curves. Journal of Computational and Applied Mathematics, 229 (1):294–301, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005785.

Cakoni:2005:IES

[CD05] Fioralba Cakoni and Eric Darrigrand. The inverse electromagnetic scattering problem for a mixed boundary value problem for screens. Journal of Computational and Applied Mathematics, 174(2):251–269, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400216X.

Conte:2006:FCM

[CD06] Dajana Conte and Ida Del Prete. Fast collocation methods for Volterra integral equations of convolution type. Journal of Computational and Applied Mathematics, 196(2):652–663, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500628X.

Costa:2007:HOH

[CD07a] Bruno Costa and Wai Sun Don. High order hybrid central-WENO finite difference scheme for conservation laws. Journal of Computational and Applied Mathematics, 204(2):209–218, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003517.

Costabile:2007:PAF

[CD07b] F. A. Costabile and F. Dell’Accio. Polynomial approximation of C^{M} functions by means of boundary values and applications: a survey. Journal of Computational and Applied Mathematics, 210(1–2):

116–135, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006546.

Criado:2006:NRC

[CdAHBR06] R. Criado, A. García del Amo, B. Hernández-Bermejo, and M. Romance. New results on computable efficiency and its stability for complex networks. Journal of Computational and Applied Mathematics, 192(1):59–74, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003080.

Costabile:2007:P

[CDG07] F. A. Costabile, F. Dell’Accio, and M. I. Gualtieri. Preface. Journal of Computational and Applied Mathematics, 210(1–2):1, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006418.

Costabile:2005:EPE

[CDL05] F. A. Costabile, F. Dell’Accio, and R. Luceri. Explicit polynomial expansions of regular real functions by means of even order Bernoulli polynomials and boundary values. Journal of Computational and Applied Mathematics, 176(1):77–90, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003218.

Campagna:2007:EAR

[CDM07a] R. Campagna, L. D’Amore, and A. Murli. An efficient algorithm for regularization of Laplace transform inversion in real case. Journal of Computational and Applied Mathematics, 210(1–2): 84–98, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006510.

Cuomo:2007:EAC

[CDM07b] S. Cuomo, L. D’Amore, and A. Murli. Error analysis of a collocation method for numerically inverting a Laplace transform in case of real samples. Journal of Computational and Applied Mathematics, 210(1–2): 149–158, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600656X.

Cuomo:2007:CIL

[CDMR07] S. Cuomo, L. D’Amore, A. Murli, and M. Rizzardi. Computation of the inverse Laplace transform based on a collocation method which uses only real values. Journal of Computational and Applied Mathematics, 198(1): 98–115, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007089.

Caliari:2007:HS

[CDV07] Marco Caliari, Stefano De Marchi, and Marco Vianello. Hyperinterpolation on the square. Journal of Computational and Applied Mathematics, 210(1–2):78–83, December 31, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006509.

Caliari:2008:BLI

[CDV08] Marco Caliari, Stefano De Marchi, and Marco Vianello. Bivariate Lagrange interpolation at the Padua points: Computational aspects. Journal of Computational and Applied Mathematics, 221(2): 284–292, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005614.

Carvalho:2007:ILS

[CdVT07] Jonas C. Carvalho, Marco Túllio M. B. de Vilhena, and Mark Thompson. An iterative Langevin solution for turbulent dispersion in the atmosphere. Journal of Computational and Applied Mathematics, 206(1): 534–548, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005036.

Conradie:2006:EAD

[CdWJ06] W. J. Conradie, T. de Wet, and M. Jankowitz. Exact and asymptotic distributions of LULU smoothers. Journal of Computational and Applied Mathematics, 186(1): 253–267, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001998.

Ciarlet:2006:P

[CDZ06] Philippe G. Ciarlet, Hui-Hui Dai, and Qiang Zhang. Preface. Journal of Computational and Applied Mathematics, 190(1–2):1–2, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002177.

Calabro:2009:ECC

[CE09] F. Calabrò and A. Corbo Esposito. An evaluation of Clenshaw–Curtis quadrature rule for integration w.r.t. singular measures. Journal of Computational and Applied Mathematics, 229 (1):120–128, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005256.

Coluccio:2008:GLB

[CEF08] Loredana Coluccio, Alfredo Eisinberg, and Giuseppe Fedele. Gauss–Lobatto to Bernstein polynomials transformation. Journal of Computational and Applied Mathematics, 222(2):690–700, December 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006498.

Chrispell:2009:FSM

[CEJ09] J. C. Chrispell, V. J. Ervin, and E. W. Jenkins. A fractional step θ-method approximation of time-dependent viscoelastic fluid flow. Journal of Computational and Applied Mathematics, 232(2): 159–175, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003306.

Cen:2008:SOD

[Cen08] Zhongdi Cen. A secondorder difference scheme for a parameterized singular perturbation problem. Journal of Computational and Applied Mathematics, 221(1): 174–182, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005341.

Criado:2007:CLC

[CFGVP07] R. Criado, J. Flores, M. I. González-Vasco, and J. Pello. Choosing a leader on a complex network. Journal of Computational and Applied Mathematics, 204 (1):10–17, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002342.

Campo:2006:ATF

[CFHM06] M. Campo, J. R. Fernández, and T.-V. Hoarau-Mantel. Analysis of two frictional viscoplastic contact problems with damage. Journal of Computational and Applied Mathematics, 196(1): 180–197, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005352.

Cheng:2005:SPP

[CFL05] Hao Cheng, Shu-Cherng Fang, and John E. Lavery. Shape-preserving properties of univariate cubic L_{1} splines. Journal of Computational and Applied Mathematics, 174(2):361–382, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002213.

Calvo:2008:SPE

[CFMR08] M. Calvo, J. M. Franco, J. I.

Montijano, and L. Rández. Structure preservation of exponentially fitted Runge– Kutta methods. Journal of Computational and Applied Mathematics, 218(2): 421–434, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700249X.

Calvo:2009:SOS

[CFMR09] M. Calvo, J. M. Franco, J. I.

Montijano, and L. Rández. Sixth-order symmetric and symplectic exponentially fitted Runge–Kutta methods of the Gauss type. Journal of Computational and Applied Mathematics, 223(1): 387–398, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000307.

Campo:2009:DCB

[CFS09] M. Campo, J. R. Fernández, and A. Silva. Damageable contact between an elastic body and a rigid foundation. Journal of Computational and Applied Mathematics, 224(2):646–657, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002902.

Campo:2006:NAS

[CFV06] M. Campo, J. R. Fernández, and J. M. Viaño. Numerical analysis and simulations of a quasistatic frictional contact problem with damage in viscoelasticity. Journal of Computational and Applied Mathematics, 192(1):30–39, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003055.

Coombs:2005:CAK

[CG05] Daniel Coombs and Byron Goldstein. T cell activation: Kinetic proofreading, serial engagement and cell adhesion. Journal of Computational and Applied Mathematics, 184(1): 121–139, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000737.

Ciarlet:2006:CGK

[CG06] Philippe G. Ciarlet and Liliana Gratie. From the classical to the generalized von Kármán and Marguerre–von Kármán equations. Journal of Computational and Applied Mathematics, 190(1– 2):470–486, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002475.

Cui:2007:SST

[CG07] Minggen Cui and Fazhan Geng. Solving singular twopoint boundary value problem in reproducing kernel space. Journal of Computational and Applied Mathematics, 205(1):6–15, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002524.

Calvete:2008:BMP

[CG08a] Herminia I. Calvete and Carmen Galé. Bilevel multiplicative problems: a penalty approach to optimality and a cutting plane based algorithm. Journal of Computational and Applied Mathematics, 218(2): 259–269, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000258.

Carnicer:2008:CPL

[CG08b] J. M. Carnicer and M. Gasca. Cubic pencils of lines and bivariate interpolation. Journal of Computational and Applied Mathematics, 219(2): 370–382, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002713.

Conti:2007:TPF

[CGP07] Costanza Conti, Laura Gori, and Francesca Pitolli. Totally positive functions through nonstationary subdivision schemes. Journal of Computational and Applied Mathematics, 200(1):255–265, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008101.

Chen:2009:LCD

[CGP09] Jein-Shan Chen, Hung-Ta Gao, and Shaohua Pan. An R-linearly convergent derivative-free algorithm for nonlinear complementarity problems based on the generalized Fischer–Burmeister merit function. Journal of Computational and Applied Mathematics, 232(2): 455–471, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003781.

Calvo:2008:GEE

[CGPM08] M. Calvo, S. González-Pinto, and J. I. Montijano. Global error estimation based on the tolerance proportionality for some adaptive Runge–Kutta codes. Journal of Computational and Applied Mathematics, 218(2): 329–341, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001240.

Conti:2008:AGB

[CGPS08] Costanza Conti, Laura Gori, Francesca Pitolli, and Paul Sablonnière. Approximation by GP box-splines on a four-direction mesh. Journal of Computational and Applied Mathematics, 221(2): 310–329, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700564X.

Castro:2008:SSW

[CGRGVP08] M. J. Castro, J. A. García-Rodríguez, J. M. González-Vida, and C. Parés. Solving shallow-water systems in 2D domains using finite volume methods and multimedia SSE instructions. Journal of Computational and Applied Mathematics, 221(1): 16–32, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005201.

Chun:2009:HOA

[CH09] S. Chun and J. S. Hesthaven. High-order accurate thin layer approximations for time-domain electromagnetics. Part I: General metal backed coatings. Journal of Computational and Applied Mathematics, 231(2): 598–611, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002520.

Chanane:2007:CSN

[Cha07] B. Chanane. Computing the spectrum of non-self-adjoint Sturm–Liouville problems with parameter-dependent boundary conditions. Journal of Computational and Applied Mathematics, 206(1): 229–237, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004377.

Chanane:2008:SLP

[Cha08] B. Chanane. Sturm–Liouville problems with parameter dependent potential and boundary conditions. Journal of Computational and Applied Mathematics, 212(2): 282–290, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007308.

Criado:2007:AEE

[CHBMBR07] R. Criado, B. Hernández-Bermejo, J. Marco-Blanco, and M. Romance. Asymptotic estimates for efficiency, vulnerability and cost for random networks. Journal of Computational and Applied Mathematics, 204 (1):166–171, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002482.

Chen:2005:ERP

[CHCW05] Shihua Chen, Jia Hu, Li Chen, and Changping Wang. Existence results for n-point boundary value problem of second order ordinary differential equations. Journal of Computational and Applied Mathematics, 180(2): 425–432, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005424.

Chen:2005:NNP

[Che05] Fengde Chen. On a nonlinear nonautonomous predator– prey model with diffusion and distributed delay. Journal of Computational and Applied Mathematics, 180 (1):33–49, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004777.

Cheng:2008:EEF

[Che08] Ting Cheng. Error estimate of a first-order time discretization scheme for the geodynamo equations. Journal of Computational and Applied Mathematics, 219(1): 35–50, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003615.

Cheung:2009:SDI

[Che09] Wing-Sum Cheung. Sharp discrete inequalities and applications to discrete variational problems. Journal of Computational and Applied Mathematics, 232(2): 176–186, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900332X.

Chihara:2005:PDS

[Chi05] T. S. Chihara. Problem 2. On the derived set of the spectrum. Journal of Computational and Applied Mathematics, 178(1–2):527–528, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004029.

Chien:2008:TGD

[CHJL08] C.-S. Chien, H.-T. Huang, B.- W. Jeng, and Z.-C. Li. Twogrid discretization schemes for nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 214(2):549–571, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001549.

Cheng:2009:CBM

[CHJZ09] Guang-Hui Cheng, Ting-Zhu Huang, Yan-Fei Jing, and Li-Tao Zhang. Convergence behaviors of multisplitting methods with K + 1 relaxed parameters. Journal of Computational and Applied Mathematics, 229(1):61–69, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005207.

Chen:2009:QBE

[CHL09] Aiyong Chen, Wentao Huang, and Jibin Li. Qualitative behavior and exact travelling wave solutions of the Zhiber– Shabat equation. Journal of Computational and Applied Mathematics, 230(2): 559–569, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006511.

Chu:2008:VFT

[CHLW08] Eric King-Wah Chu, Tsung-Min Hwang, Wen-Wei Lin, and Chin-Tien Wu. Vibration of fast trains, palindromic eigenvalue problems and structure-preserving doubling algorithms. Journal of Computational and Applied Mathematics, 219(1): 237–252, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003925.

Chen:2006:TMD

[CHM06] William Y. C. Chen, Qing-Hu Hou, and Yan-Ping Mu. A telescoping method for double summations. Journal of Computational and Applied Mathematics, 196(2): 553–566, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006205.

Chui:2008:OLI

[CHNZ08] C. K. Chui, G. Hecklin, G. Nürnberger, and F. Zeilfelder. Optimal Lagrange interpolation by quartic C^{1} splines on triangulations. Journal of Computational and Applied Mathematics, 216(2):344–363, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002555.

Chouchene:2005:HAA

[Cho05] Frej Chouchene. Harmonic analysis associated with the Jacobi–Dunkl operator on ] − \frac{π}{2} , \frac{π}{2} [. Journal of Computational and Applied Mathematics, 178(1–2):75–89, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400367X.

Cho:2008:NLU

[Cho08] Gyeong-Mi Cho. A new largeupdate interior point algorithm for P_{∗}(κ) linear complementarity problems. Journal of Computational and Applied Mathematics, 216 (1):265–278, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002452.

Christiansen:2005:IMP

[Chr05] Jacob S. Christiansen. Indeterminate moment problems related to birth and death processes with quartic rates. Journal of Computational and Applied Mathematics, 178(1–2):91–98, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003681.

Chrysafinos:2009:AFE

[Chr09] Konstantinos Chrysafinos. Analysis and finite element approximations for distributed optimal control problems for implicit parabolic equations. Journal of Computational and Applied Mathematics, 231(1): 327–348, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000995.

Chou:2006:AFB

[CHS06] W.-S. Chou, Leetsch C. Hsu, and Peter J.-S. Shiue. Application of Faà di Bruno’s formula in characterization of inverse relations. Journal of Computational and Applied Mathematics, 190(1– 2):151–169, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002281.

Chen:2007:PVR

[CHS07] Ying Chen, Wolfgang Härdle, and Vladimir Spokoiny. Portfolio value at risk based on independent component analysis. Journal of Computational and Applied Mathematics, 205(1):594–607, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003426.

Cheng:2008:NMT

[CHS08] Guang-Hui Cheng, Ting-Zhu Huang, and Shu-Qian Shen. Note to the mixed-type splitting iterative method for Zmatrices linear systems. Journal of Computational and Applied Mathematics, 220(1– 2):1–7, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003573.

Clemens:2008:MRH

[CHSW08] Markus Clemens, Moritz Helias, Thorsten Steinmetz, and Georg Wimmer. Multiple right-hand side techniques for the numerical simulation of quasistatic electric and magnetic fields. Journal of Computational and Applied Mathematics, 215 (2):328–338, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007436.

Chun:2008:SCT

[Chu08] Changbum Chun. A simply constructed third-order modifications of Newton’s method. Journal of Computational and Applied Mathematics, 219(1):81–89, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003664.

Crevecoeur:2008:ESA

[CHV^{+}08] G. Crevecoeur, H. Hallez, P. Van Hese, Y. D’Asseler, L. Dupré, and R. Van de Walle. EEG source analysis using space mapping techniques. Journal of Computational and Applied Mathematics, 215(2):339–347, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007448.

Chen:2009:MQI

[CHW09] Ronghua Chen, Xuli Han, and Zongmin Wu. A multiquadric quasi-interpolation with linear reproducing and preserving monotonicity. Journal of Computational and Applied Mathematics, 231(2): 517–525, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002337.

Chen:2006:DML

[CHXL06] Xuefeng Chen, Zhengjia He, Jiawei Xiang, and Bing Li. A dynamic multiscale lifting computation method using Daubechies wavelet. Journal of Computational and Applied Mathematics, 188 (2):228–245, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002529.

Cui:2005:SSC

[CHY05] Bao Tong Cui, Maoan Han, and Huizhong Yang. Some sufficient conditions for oscillation of impulsive delay hyperbolic systems with Robin boundary conditions. Journal of Computational and Applied Mathematics, 180(2): 365–375, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400528X.

Cimrak:2008:RPS

[Cim08]  Ivan Cimrák.  Regularity

properties of the solutions to the 3D Maxwell–Landau– Lifshitz system in weighted Sobolev spaces. Journal of Computational and Applied Mathematics, 215(2): 320–327, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007424.

Cimrak:2009:CRC

[Cim09] Ivan Cimrák. Convergence result for the constraint preserving mid-point scheme for micromagnetism. Journal of Computational and Applied Mathematics, 228 (1):238–246, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004718.

Civalek:2007:NAF

[Civ07] Ömer Civalek. Numerical analysis of free vibrations of laminated composite conical and cylindrical shells: Discrete singular convolution (DSC) approach. Journal of Computational and Applied Mathematics, 205(1): 251–271, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003141.

Chui:2006:RBQ

[CJ06] Charles K. Chui and Qingtang Jiang. Refinable bivariate quartic and quintic C^{2}-splines for quadrilateral subdivisions. Journal of Computational and Applied Mathematics, 196(2): 402–424, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005844.

Chu:2007:AMS

[CJ07] Wenchang Chu and Cangzhi Jia. Abel’s method on summation by parts and terminating well-poised qseries identities. Journal of Computational and Applied Mathematics, 207(2): 360–370, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600598X.

Conti:2008:SFS

[CJ08] C. Conti and K. Jetter.

Smoothed hat functions in subdivision. Journal of Computational and Applied Mathematics, 221(2):330– 345, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005651.

Che:2009:CMC

[CJ09a] Haitao Che and Ziwen Jiang. A characteristicsmixed covolume method for a convection-dominated transport problem. Journal of Computational and Applied Mathematics, 231(2): 760–770, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002891.

Coffey:2009:SRI

[CJ09b] M. W. Coffey and S. J. Johnston. Some results involving series representations of hypergeometric functions. Journal of Computational and Applied Mathematics, 233(3): 674–679, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001113.

Carbonell:2005:COI

[CJB05] F. Carbonell, J. C. Jimenez, and R. J. Biscay. A class of orthogonal integrators for stochastic differential equations. Journal of Computational and Applied Mathematics, 182(2):350–361, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400634X.

Carbonell:2006:WLL

[CJB06] F. Carbonell, J. C. Jimenez, and R. J. Biscay. Weak local linear discretizations for stochastic differential equations: Convergence and numerical schemes. Journal of Computational and Applied Mathematics, 197(2): 578–596, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007016.

Calvo-Jurado:2006:HDP

[CJCD06] Carmen Calvo-Jurado and Juan Casado-Díaz. Homogenization of Dirichlet parabolic problems for coefficients and open sets simultaneously variable and applications to optimal design. Journal of Computational and Applied Mathematics, 192(1):20–29, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003043.

Calvo-Jurado:2007:NLS

[CJCD07] Carmen Calvo-Jurado and Juan Casado-Díaz. Nonlocal limits in the study of linear elliptic systems arising in periodic homogenization. Journal of Computational and Applied Mathematics, 204(1):3–9, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002330.

Carbonell:2008:CMI

[CJP08] F. Carbonell, J. C. Jímenez, and L. M. Pedroso. Computing multiple integrals involving matrix exponentials. Journal of Computational and Applied Mathematics, 213(1):300–305, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000283.

Cvetkovic:2005:NCI

[CK05] Ljiljana Cvetković and Vladimir Kostić. New criteria for identifying H-matrices. Journal of Computational and Applied Mathematics, 180(2): 265–278, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005229.

Cvetkovic:2006:BGM

[CK06] Ljiljana Cvetković and Vladimir Kostić. Between Gersgorin and minimal Gersgorin sets. Journal of Computational and Applied Mathematics, 196(2):452–458, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005881.

Casian:2007:SST

[CK07] Luis Casian and Yuji Kodama. Singular structure of Toda lattices and cohomology of certain compact Lie groups. Journal of Computational and Applied Mathematics, 202(1):56–79, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001166.

Cudna:2008:FDA

[CK08] Małgorzata Cudna and Tomasz Komorowski. A finite dimensional approximation of the effective diffusivity for a symmetric random walk in a random environment. Journal of Computational and Applied Mathematics, 213(1): 186–204, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000234.

Campbell:2009:CND

[CK09a] Stephen L. Campbell and Peter Kunkel. Completions of nonlinear DAE flows based on index reduction techniques and their stabilization. Journal of Computational and Applied Mathematics, 233 (4):1021–1034, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709006049.

Chang:2009:BRM

[CK09b] Fu-Min Chang and Jau-Chuan Ke. On a batch retrial model with J vacations. Journal of Computational and Applied Mathematics, 232(2):402–414, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003744.

Ceberio:2007:ITA

[CKC^{+}07] Martine Ceberio, Vladik Kreinovich, Sanjeev Chopra, Luc Longpré, Hung T. Nguyen, Bertram Ludäscher, and Chitta Baral. Intervaltype and affine arithmetictype techniques for handling uncertainty in expert systems. Journal of Computational and Applied Mathematics, 199(2):403–410, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500779X.

Choudhury:2009:PUS

[CKT09] Gautam Choudhury, Jau-Chuan Ke, and Lotfi Tadj. The N-policy for an unreliable server with delaying repair and two phases of service. Journal of Computational and Applied Mathematics, 231(1):349–364, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900137X.

Cao:2007:GAS

[CKWZ07] Xifang Cao, Qingkai Kong, Hongyou Wu, and Anton Zettl. Geometric aspects of Sturm–Liouville problems. III. Level surfaces of the n-th eigenvalue. Journal of Computational and Applied Mathematics, 208(1): 176–193, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006340.

Chen:2006:CBI

[CL06a] Jinhai Chen and Weiguo Li.

Convergence behaviour of inexact Newton methods under weak Lipschitz condition. Journal of Computational and Applied Mathematics, 191(1):143–164, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002773.

Chen:2006:SIP

[CL06b] Qun Chen and Jijun Liu.

Solving an inverse parabolic problem by optimization from final measurement data. Journal of Computational and Applied Mathematics, 193(1):183–203, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003870.

Chen:2006:NPB

[CL06c] Xiaoshan Chen and Wen Li. A note on the perturbation bounds of eigenspaces for Hermitian matrices. Journal of Computational and Applied Mathematics, 196(1): 338–346, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005674.

Chen:2008:PEE

[CL08] Yanping Chen and Wenbin Liu. A posteriori error estimates for mixed finite element solutions of convex optimal control problems. Journal of Computational and Applied Mathematics, 211(1): 76–89, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006972.

Chen:2009:NMI

[CL09a] Ren-Chuen Chen and Jinn-Liang Liu. Nonstationary monotone iterative methods for nonlinear partial differential equations. Journal of Computational and Applied Mathematics, 233(4): 859–877, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004506.

Chi:2009:OSS

[CL09b] Xiaoni Chi and Sanyang Liu.

A one-step smoothing Newton method for second-order cone programming. Journal of Computational and Applied Mathematics, 223(1): 114–123, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006723.

Clarkson:2005:SPA

[Cla05] Peter A. Clarkson. Special polynomials associated with rational solutions of the fifth Painlevé equation. Journal of Computational and Applied Mathematics, 178(1– 2):111–129, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400370X.

Chen:2009:FME

[CLA09] Chang-Ming Chen, F. Liu, and V. Anh. A Fourier method and an extrapolation technique for Stokes’ first problem for a heated generalized second grade fluid with fractional derivative. Journal of Computational and Applied Mathematics, 223(2): 777–789, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000836.

Chen:2007:DBD

[CLCL07] Fengde Chen, Zhong Li, Xiaoxing Chen, and Jitka Laitochová. Dynamic behaviors of a delay differential equation model of plankton allelopathy. Journal of Computational and Applied Mathematics, 206(2): 733–754, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600519X.

Cai:2009:SAH

[CLGG09] Liming Cai, Xuezhi Li, Mini Ghosh, and Baozhu Guo. Stability analysis of an HIV/ AIDS epidemic model with treatment. Journal of Computational and Applied Mathematics, 229(1):313–323, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005815.

Calvo:2006:IRG

[CLM06] M. Calvo, M. P. Laburta, and J. I. Montijano. Initializers for RK–Gauss methods based on pseudo-symplecticity. Journal of Computational and Applied Mathematics, 189(1– 2):228–241, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002803.

Chen:2009:REI

[CLN09] Zhongying Chen, Guangqing Long, and Gnaneshwar Nelakanti. Richardson extrapolation of iterated discrete projection methods for eigenvalue approximation. Journal of Computational and Applied Mathematics, 223(1):48–61, January 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006577.

Cheikh:2009:ASO

[CLO09] Y. Ben Cheikh, I. Lamiri, and A. Ouni. On Askeyscheme and d-orthogonality. I. A characterization theorem. Journal of Computational and Applied Mathematics, 233(3):621–629, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001046.

Canovas:2008:EPC

[CLP08] M. J. Cánovas, M. A. López, and J. Parra. On the equivalence of parametric contexts for linear inequality systems. Journal of Computational and Applied Mathematics, 217(2):448–456, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700101X.

Chen:2009:AGD

[CLQ09] Haibo Chen, Xianqing Lv, and Yansong Qiao. Application of gradient descent method to the sedimentary grain-size distribution fitting. Journal of Computational and Applied Mathematics, 233(4):1128–1138, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900627X.

Cotronei:2008:MPU

[CLS08] Mariantonia Cotronei, Maria Laura Lo Cascio, and Tomas Sauer. Multifilters and prefilters: Uniqueness and algorithmic aspects. Journal of Computational and Applied Mathematics, 221(2):346– 354, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005663.

Casciola:2005:ISS

[CM05a] G. Casciola and S. Morigi. Inverse spherical surfaces. Journal of Computational and Applied Mathematics, 176(2):411–424, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003450.

Cash:2005:NMS

[CM05b] J. R. Cash and F. Mazzia.

A new mesh selection algorithm, based on conditioning, for two-point boundary value codes. Journal of Computational and Applied Mathematics, 184(2): 362–381, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000373.

Campbell:2007:DTS

[CM07] Stephen L. Campbell and Roswitha März. Direct transcription solution of high index optimal control problems and regular Euler–Lagrange equations. Journal of Computational and Applied Mathematics, 202(2):186–202, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600121X.

Ciarlini:2008:WEN

[CM08] Patrizia Ciarlini and Umberto Maniscalco. Wavelets and Elman neural networks for monitoring environmental variables. Journal of Computational and Applied Mathematics, 221(2):302– 309, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005638.

Christov:2009:DSV

[CMM09] Christo I. Christov, Rossitza S. Marinova, and Tchavdar T. Marinov. Does the stationary viscous flow around a circular cylinder exist for large Reynolds numbers? a numerical solution via variational imbedding. Journal of Computational and Applied Mathematics, 226(2): 205–217, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004019.

Cardone:2006:FIM

[CMR06] A. Cardone, E. Messina, and E. Russo. A fast iterative method for discretized Volterra–Fredholm integral equations. Journal of Computational and Applied Mathematics, 189(1– 2):568–579, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003651.

Chow:2006:DPM

[CMRS06] K. W. Chow, C. C. Mak, C. Rogers, and W. K. Schief. Doubly periodic and multiple pole solutions of the sinh-Poisson equation: Application of reciprocal transformations in subsonic gas dynamics. Journal of Computational and Applied Mathematics, 190(1–2):114–126, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002244.

Cravero:2008:GCQ

[CMS08] Isabella Cravero, Carla Manni, and M. Lucia Sampoli. Geometric construction of quintic parametric B-splines. Journal of Computational and Applied Mathematics, 221(2): 355–366, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005675.

Cash:2006:RCM

[CMST06] J. R. Cash, F. Mazzia, N. Sumarti, and D. Trigiante. The role of conditioning in mesh selection algorithms for first order systems of linear two point boundary value problems. Journal of Computational and Applied Mathematics, 185(2): 212–224, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001111.

Cordero:2009:IMO

[CMT09] Alicia Cordero, Eulalia Martínez, and Juan R. Torregrosa. Iterative methods of order four and five for systems of nonlinear equations. Journal of Computational and Applied Mathematics, 231(2): 541–551, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002453.

Cardone:2009:AMV

[CMV09] A. Cardone, E. Messina, and A. Vecchio. An adaptive method for Volterra– Fredholm integral equations on the half line. Journal of Computational and Applied Mathematics, 228 (2):538–547, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001398.

Cai:2009:PTM

[CMX09] Mingchao Cai, Mo Mu, and Jinchao Xu. Preconditioning techniques for a mixed Stokes/Darcy model in porous media applications. Journal of Computational and Applied Mathematics, 233(2):346–355, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004269.

Costabile:2007:SBU

[CN07] F. Costabile and A. Napoli.

Solving BVPs using twopoint Taylor formula by a symbolic software. Journal of Computational and Applied Mathematics, 210(1–2): 136–148, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006558.

Creuse:2008:PEE

[CN08] Emmanuel Creusé and Serge Nicaise. A posteriori error estimations of a coupled mixed and standard Galerkin method for second order operators. Journal of Computational and Applied Mathematics, 213(1):35–55, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000040.

Chindelevitch:2007:EAP

[CNN07] Leonid Chindelevitch, David P.

Nicholls, and Nilima Nigam. Error analysis and preconditioning for an enhanced DtN– FE algorithm for exterior scattering problems. Journal of Computational and Applied Mathematics, 204 (2):493–504, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600375X.

Chapwanya:2005:CSS

[CO05] M. Chapwanya and S. B. G. O’Brien. Corner and startup singularities in porous flow. Journal of Computational and Applied Mathematics, 176(1):163–177, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003255.

Cabada:2005:OCE

[COEV05] Alberto Cabada, Victoria Otero-Espinar, and Dolores R. Vivero. Optimal conditions to ensure the stability of periodic solutions of first order difference equations lying between lower and upper solutions. Journal of Computational and Applied Mathematics, 176(1):45–57, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400319X.

Coffey:2005:ODD

[Cof05] Mark W. Coffey. On onedimensional digamma and polygamma series related to the evaluation of Feynman diagrams. Journal of Computational and Applied Mathematics, 183(1):84–100, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000154.

Coffey:2008:SSR

[Cof08] Mark W. Coffey. On some series representations of the Hurwitz zeta function. Journal of Computational and Applied Mathematics, 216 (1):297–305, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002476.

Coffey:2009:EAH

[Cof09a] Mark W. Coffey. An efficient algorithm for the Hurwitz zeta and related functions. Journal of Computational and Applied Mathematics, 225(2):338–346, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003713.

Coffey:2009:SPL

[Cof09b] Mark W. Coffey. Summation properties of the η_{j} and Li constants. Journal of Computational and Applied Mathematics, 233(3): 667–673, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001101.

Cojuhari:2009:DSG

[Coj09] P. A. Cojuhari. Discrete spectrum in the gaps for perturbations of periodic Jacobi matrices. Journal of Computational and Applied Mathematics, 225(2):374–386, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003865.

Copetti:2005:EAF

[Cop05] M. I. M. Copetti. Error analysis for a finite element approximation of a thermoviscoelastic contact problem. Journal of Computational and Applied Mathematics, 180(1): 181–190, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005059.

Courrieu:2009:FSW

[Cou09] Pierre Courrieu. Fast solving of weighted pairing leastsquares systems. Journal of Computational and Applied Mathematics, 231(1):

39–48, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000387.

Cortes:2007:GFE

[CP07] V. Cortés and J. M. Peña.

Growth factor and expected growth factor of some pivoting strategies. Journal of Computational and Applied Mathematics, 202(2): 292–303, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001403.

Chen:2008:DMR

[CP08a] Jein-Shan Chen and Shaohua Pan. A descent method for a reformulation of the secondorder cone complementarity problem. Journal of Computational and Applied Mathematics, 213(2):547–558, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000714.

Chen:2008:RSN

[CP08b] Jein-Shan Chen and Shaohua Pan. A regularization semismooth Newton method based on the generalized Fischer–Burmeister function for P_{0}–NCPs. Journal of Computational and Applied Mathematics, 220(1–2): 464–479, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004669.

Chrysafis:2009:TPO

[CP09a] Konstantinos A. Chrysafis and Basil K. Papadopoulos. On theoretical pricing of options with fuzzy estimators. Journal of Computational and Applied Mathematics, 223(2):552–566, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700636X.

Cuesta:2009:NSP

[CP09b] C. M. Cuesta and I. S. Pop. Numerical schemes for a pseudo-parabolic Burgers equation: Discontinuous data and long-time behaviour. Journal of Computational and Applied Mathematics, 224(1):269–283, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002070.

Cui:2009:BRM

[CP09c] Li Cui and Lizhong Peng. Biorthogonal radial multiresolution in dimension three. Journal of Computational and Applied Mathematics, 224(2):581–591, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002756.

Carneiro:2005:IST

[CPM^{+}05] Jorge Carneiro, Tiago Paixão, Dejan Milutinovic, João Sousa, Kalet Leon, Rui Gardner, and Jose Faro. Immunological self-tolerance: Lessons from mathematical modeling. Journal of Computational and Applied Mathematics, 184(1):77–100, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000713.

Cho:2009:CGI

[CQ09] Yeol Je Cho and Xiaolong Qin. Convergence of a general iterative method for nonexpansive mappings in Hilbert spaces. Journal of Computational and Applied Mathematics, 228(1):458–465, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005116.

Casciola:2009:NTM

[CR09] G. Casciola and L. Romani. A Newton-type method for constrained least-squares datafitting with easy-to-control rational curves. Journal of Computational and Applied Mathematics, 223(2): 672–692, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000666.

Caballero:2005:MSI

[CRS05] J. Caballero, J. Rocha, and K. Sadarangani. On monotonic solutions of an integral equation of Volterra type. Journal of Computational and Applied Mathematics, 174(1):119–133, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001840.

Cahlon:2006:SCC

[CS06a] Baruch Cahlon and Darrell Schmidt. Stability criteria for certain third-order delay differential equations. Journal of Computational and Applied Mathematics, 188 (2):319–335, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500258X.

Censor:2006:AQF

[CS06b] Yair Censor and Alexander Segal. Algorithms for the quasiconvex feasibility problem. Journal of Computational and Applied Mathematics, 185(1):34–50, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500049X.

Cahlon:2007:SCC

[CS07a] Baruch Cahlon and Darrell Schmidt. Stability criteria for certain high odd order delay differential equations. Journal of Computational and Applied Mathematics, 200(1):408–423, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000240.

Cai:2007:PSP

[CS07b] Liming Cai and Xinyu Song.

Permanence and stability of a predator–prey system with stage structure for predator. Journal of Computational and Applied Mathematics, 201(2):356–366, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000884.

Chaillou:2007:PEL

[CS07c] Alexandra Chaillou and Manil Suri. A posteriori estimation of the linearization error for strongly monotone nonlinear operators. Journal of Computational and Applied Mathematics, 205 (1):72–87, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002561.

Chen:2007:NBP

[CS07d] Lijing Chen and Jitao Sun. Nonlinear boundary problem of first order impulsive integro-differential equations. Journal of Computational and Applied Mathematics, 202(2):392–401, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001476.

Chen:2007:CCF

[CS07e] Rudong Chen and Yunyan Song. Convergence to common fixed point of nonexpansive semigroups. Journal of Computational and Applied Mathematics, 200(2): 566–575, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000367.

Charina:2008:TWF

[CS08] Maria Charina and Joachim Stöckler. Tight wavelet frames for subdivision. Journal of Computational and Applied Mathematics, 221(2): 293–301, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005626.

Cambini:2009:BRA

[CS09a] Riccardo Cambini and Francesca Salvi. A branch and reduce approach for solving a class of low rank d.c. programs. Journal of Computational and Applied Mathematics, 233(2): 492–501, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004580.

Coussement:2009:IPC

[CS09b] J. Coussement and C. Smet.

Irrationality proof of certain Lambert series using little q-Jacobi polynomials. Journal of Computational and Applied Mathematics, 233(3): 680–690, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001125.

Chan:2005:IMC

[CSG05] Cliburn Chan, Jaroslav Stark, and Andrew J. T. George. The impact of multiple T cell–APC encounters and the role of anergy. Journal of Computational and Applied Mathematics, 184(1): 101–120, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000725.

Costas-Santos:2009:EDC

[CSSL09] R. S. Costas-Santos and J. F. Sánchez-Lara. Extensions of discrete classical orthogonal polynomials beyond the orthogonality. Journal of Computational and Applied Mathematics, 225(2): 440–451, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003968.

Christiansen:2009:FGJ

[CSZ09] Jacob S. Christiansen, Barry Simon, and Maxim Zinchenko. Finite gap Jacobi matrices: an announcement. Journal of Computational and Applied Mathematics, 233(3): 652–662, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001083.

Cui:2005:PSS

[CT05] Jing’an Cui and Yasuhiro Takeuchi. Permanence of a single-species dispersal system and predator survival. Journal of Computational and Applied Mathematics, 175(2):375–394, March 15, 2005. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003115.

Chen:2009:PMV

[CT09a] Wei Chen and Shaohua Tan.

On the possibilistic mean value and variance of multiplication of fuzzy numbers. Journal of Computational and Applied Mathematics, 232(2):327–334, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003604.

Chen:2009:RPS

[CT09b] Wei Chen and Shaohua Tan.

Robust portfolio selection based on asymmetric measures of variability of stock returns. Journal of Computational and Applied Mathematics, 232(2):295–304, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003562.

Chen:2009:SMW

[CT09c] Yanping Chen and Tao Tang.

Spectral methods for weakly singular Volterra integral equations with smooth solutions. Journal of Computational and Applied Mathematics, 233(4):938–950, December 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709005342.

Chiba:2009:CSA

[CU09a] Fumihiro Chiba and Teruo Ushijima. Computation of the scattering amplitude for a scattering wave produced by a disc — approach by a fundamental solution method. Journal of Computational and Applied Mathematics, 233(4):1155–1174, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709006256.

Chiba:2009:EDE

[CU09b] Fumihiro Chiba and Teruo Ushijima. Exponential decay of errors of a fundamental solution method applied to a reduced wave problem in the exterior region of a disc. Journal of Computational and Applied Mathematics, 231(2): 869–885, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003057.

Cui:2007:CMD

[Cui07]  Mingrong Cui.  A com-

bined mixed and discontinuous Galerkin method for compressible miscible displacement problem in porous media. Journal of Computational and Applied Mathematics, 198(1):19–34, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007041.

Cui:2008:ASD

[Cui08]  Mingrong Cui.  Analysis

of a semidiscrete discontinuous Galerkin scheme for compressible miscible displacement problem. Journal of Computational and Applied Mathematics, 214 (2):617–636, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001707.

Ciaurri:2005:TWN

[CV05a] Óscar Ciaurri and Juan L.

Varona. Two-weight norm inequalities for the Cesàro means of generalized Hermite expansions. Journal of Computational and Applied Mathematics, 178(1– 2):99–110, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003693.

Coussement:2005:GQM

[CV05b] Jonathan Coussement and Walter Van Assche. Gaussian quadrature for multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 178(1– 2):131–145, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003711.

Csorgo:2006:TSB

[CV06] Sándor Csörgő and László Viharos. Testing for small bias of tail index estimators. Journal of Computational and Applied Mathematics, 186(1):232–252, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001986.

Ciaurri:2009:SAE

[CV09] Óscar Ciaurri and Juan Luis Varona. The surprising almost everywhere convergence of Fourier–Neumann series. Journal of Computational and Applied Mathematics, 233(3):663–666, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001095.

Caliari:2007:LEI

[CVB07] Marco Caliari, Marco Vianello, and Luca Bergamaschi. The LEM exponential integrator for advection–diffusionreaction equations. Journal of Computational and Applied Mathematics, 210(1–2): 56–63, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006479.

Currie:2005:EAD

[CW05] Sonja Currie and Bruce A.

Watson. Eigenvalue asymptotics for differential operators on graphs. Journal of Computational and Applied Mathematics, 182(1): 13–31, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005977.

Chen:2008:CTD

[CW08a] Jie Chen and Guo-Jin Wang.

Construction of triangular DP surface and its application. Journal of Computational and Applied Mathematics, 219(1):312–326, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042707004128.

Chen:2008:ASI

[CW08b] Shen-Yu Chen and Soon-Yi Wu. An algorithm for semiinfinite transportation problems. Journal of Computational and Applied Mathematics, 217(2):365–380, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000969.

Chen:2008:AIQ

[CW08c] Shen-Yu Chen and Soon-Yi Wu. Algorithms for infinite quadratic programming in L_{p} spaces. Journal of Computational and Applied Mathematics, 213(2):408–422, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000404.

Currie:2008:MAS

[CW08d] Sonja Currie and Bruce A. Watson. M-matrix asymptotics for Sturm–Liouville problems on graphs. Journal of Computational and Applied Mathematics, 218(2): 568–578, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006279.

Chandler-Wilde:2007:MSU

[CWMT07] Simon N. Chandler-Wilde, Peter Monk, and Martin Thomas. The mathematics of scattering by unbounded, rough, inhomogeneous layers. Journal of Computational and Applied Mathematics, 204(2):549–559, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003803.

Chen:2008:SSD

[CWW08] Long Chen, Yonggang Wang, and Jinbiao Wu. Stability of a streamline diffusion finite element method for turning point problems. Journal of Computational and Applied Mathematics, 220(1–2): 712–724, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005109.

Chen:2006:WNF

[CX06] Bin Chen and Yingchao Xie.

White noise functional solutions of Wick-type stochastic generalized Hirota–Satsuma coupled KdV equations. Journal of Computational and Applied Mathematics, 197(2):345–354, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006734.

Chen:2007:PLS

[CX07] Bin Chen and Yingchao Xie.

Periodic-like solutions of variable coefficient and Wicktype stochastic NLS equations. Journal of Computational and Applied Mathematics, 203(1):249–263, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002007.

Cong:2008:TTP

[CX08] Nguyen Huu Cong and Le Ngoc Xuan. Twostepby-twostep PIRK-type PC methods with continuous output formulas. Journal of Computational and Applied Mathematics, 221(1): 165–173, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700533X.

Cheng:2009:FDF

[CXH09] Wanyou Cheng, Yunhai Xiao, and Qing-Jie Hu. A family of derivative-free conjugate gradient methods for large-scale nonlinear systems of equations. Journal of Computational and Applied Mathematics, 224(1):11–19, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001490.

See comments [ZJ12].

Chen:2006:EUS

[CXS06] Fengde Chen, Xiangdong Xie, and Jinlin Shi. Existence, uniqueness and stability of positive periodic solution for a nonlinear prey-competition model with delays. Journal of Computational and Applied Mathematics, 194(2): 368–387, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005005.

Ceng:2008:APA

[CY08a] L. C. Ceng and J. C. Yao.

Approximate proximal algorithms for generalized variational inequalities with pseudomonotone multifunctions. Journal of Computational and Applied Mathematics, 213(2):423–438, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000416.

Ceng:2008:HIS

[CY08b] Lu-Chuan Ceng and Jen-Chih Yao. A hybrid iterative scheme for mixed equilibrium problems and fixed point problems. Journal of Computational and Applied Mathematics, 214(1):186–201, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001069.

Ceng:2008:CAI

[CY08c] Lu-Chuan Ceng and Jen-Chih Yao. On the convergence analysis of inexact hybrid extragradient proximal point algorithms for maximal monotone operators. Journal of Computational and Applied Mathematics, 217(2): 326–338, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000933.

Cao:2009:CRM

[CY09] Weidong Cao and Danping Yang. Ciarlet–Raviart mixed finite element approximation for an optimal control problem governed by the first biharmonic equation. Journal of Computational and Applied Mathematics, 233(2): 372–388, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900435X.

Chen:2009:TGM

[CYB09] Chuanjun Chen, Min Yang, and Chunjia Bi. Twogrid methods for finite volume element approximations of nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 228(1):123–132, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004548.

Chen:2008:ASN

[CYM08] Shaochun Chen, Li Yin, and Shipeng Mao. An anisotropic, superconvergent nonconforming plate finite element. Journal of Computational and Applied Mathematics, 220(1–2): 96–110, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004190.

Chang:2009:FEA

[CYZ09] Yanzhen Chang, Danping Yang, and Jiang Zhu. Finite element approximation to nonlinear coupled thermal problem. Journal of Computational and Applied Mathematics, 225(2): 467–477, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004196.

Cheikh:2007:OGF

[CZ07] Y. Ben Cheikh and A. Zaghouani. d-Orthogonality via generating functions. Journal of Computational and Applied Mathematics, 199 (1):2–22, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500734X.

Chen:2008:VCA

[CZ08a] Zhaodou Chen and Lijing Zhang. Vector coding algorithms for multidimensional discrete Fourier transform. Journal of Computational and Applied Mathematics, 212(1):63–74, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007126.

Chu:2008:TKT

[CZ08b] Wenchang Chu and Wenlong Zhang. Transformations of Kummer-type for 2 F _{2}-series and their qanalogues. Journal of Computational and Applied Mathematics, 216(2):467–473, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002865.

Chen:2009:DAS

[CZ09] Xingwu Chen and Weinian Zhang. Decomposition of algebraic sets and applications to weak centers of cubic systems. Journal of Computational and Applied Mathematics, 232(2):565–581, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003884.

Doha:2006:CIE

[DA06] E. H. Doha and H. M.

Ahmed. On the coefficients of integrated expansions of Bessel polynomials. Journal of Computational and Applied Mathematics, 187 (1):58–71, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001457.

Djordjevic:2008:SSN

[DA08] Vladan D. Djordjevic and Teodor M. Atanackovic. Similarity solutions to nonlinear heat conduction and Burgers/Korteweg–de Vries fractional equations. Journal of Computational and Applied Mathematics, 222(2): 701–714, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006516.

Dalkilic:2009:FAN

[DA09] Turkan Erbay Dalkilic and Aysen Apaydin. A fuzzy adaptive network approach to parameter estimation in cases where independent variables come from an exponential distribution. Journal of Computational and Applied Mathematics, 233(1): 36–45, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800397X.

Daalhuis:2009:HHE

[Daa09a] A. B. Olde Daalhuis. Hyperasymptotics and hyperterminants: Exceptional cases. Journal of Computational and Applied Mathematics, 233(2):555–563, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004683.

Daalhuis:2009:CPD

[Daa09b] A. B. Olde Daalhuis. On the computation of parameter derivatives of solutions of linear difference equations. Journal of Computational and Applied Mathematics, 230(1):128–134, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800592X.

DaCunha:2005:STV

[DaC05] Jeffrey J. DaCunha. Stability for time varying linear dynamic systems on time scales. Journal of Computational and Applied Mathematics, 176(2):381–410, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003449.

Doha:2005:ASS

[DAE05] E. H. Doha and W. M. Abd-Elhameed. Accurate spectral solutions for the parabolic and elliptic partial differential equations by the ultraspherical tau method. Journal of Computational and Applied Mathematics, 181(1): 24–45, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005473.

Dai:2009:GMV

[Dai09]  Hong-Xia Dai.  Gener-

alized mixed variationallike inequality for random fuzzy mappings. Journal of Computational and Applied Mathematics, 224(1): 20–28, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001507.

Dai:2011:CEE

[Dai11]  Xiaoxia Dai.  Corrigen-

dum to “Error estimates for an operator-splitting method for Navier–Stokes equations: Second-order schemes” [J. Comput. Appl. Math. 231

(2) (2009) 696–704]. Journal of Computational and Applied Mathematics, 235(9): 3110–3111, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042710006448.

See [DSC09].

Dang:2006:IMS

[Dan06]  Quang A. Dang.  It-

erative method for solving the Neumann boundary value problem for biharmonic type equation. Journal of Computational and Applied Mathematics, 196(2): 634–643, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006266.

Daoud:2007:OSW

[Dao07] Daoud S. Daoud. Overlapping Schwarz waveform relaxation method for the solution of the forward-backward heat equation. Journal of Computational and Applied Mathematics, 208(2): 380–390, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005930.

Daoud:2008:DSP

[Dao08] Daoud S. Daoud. Determination of the source parameter in a heat equation with a nonlocal boundary condition. Journal of Computational and Applied Mathematics, 221(1):261–272, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005948.

deBruin:2005:PTIb

[dB05a] Marcel G. de Bruin. (0, m)

Pál-type interpolation: interchanging value- and derivativenodes. Journal of Computational and Applied Mathematics, 179(1–2):175–184, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004431.

deBruin:2005:PTIa

[dB05b] Marcel G. de Bruin. (0, m)

Pál-type interpolation on the Möbius transform of roots of unity. Journal of Computational and Applied Mathematics, 178(1–2):147–153, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003723.

deBruin:2005:PPT

[dB05c] Marcel G. de Bruin. Problem 3. Pál-type interpolation. Journal of Computational and Applied Mathematics, 178(1–2):529–530, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004030.

Day:2008:ACL

[DB08] David Day and Pavel Bochev. Analysis and computation of a least-squares method for consistent mesh tying. Journal of Computational and Applied Mathematics, 218 (1):21–33, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003238.

DellAcqua:2007:RDM

[DBE07] Guido Dell’Acqua, Luis L. Bonilla, and Ramón Escobedo. Relocation dynamics and multistable switching in semiconductor superlattices. Journal of Computational and Applied Mathematics, 204(1):18–24, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002354.

Dutt:2007:SEM

[DBG07] P. Dutt, P. Biswas, and S. Ghorai. Spectral element methods for parabolic problems. Journal of Computational and Applied Mathematics, 203(2):461–486, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002226.

Dutt:2008:PSE

[DBR08] P. Dutt, P. Biswas, and G. Naga Raju. Preconditioners for spectral element methods for elliptic and parabolic problems. Journal of Computational and Applied Mathematics, 215(1):152–166, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001823.

Dai:2008:TGM

[DC08] Xiaoxia Dai and Xiaoliang Cheng. A two-grid method based on Newton iteration for the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 220(1–2):566–573, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004876.

Dimitrov:2009:LPJ

[DC09a] Dimitar K. Dimitrov and Youssèf Ben Cheikh. Laguerre polynomials as Jensen polynomials of Laguerre– Pólya entire functions. Journal of Computational and Applied Mathematics, 233(3): 703–707, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001150.

Dong:2009:PPP

[DC09b] Lingzhen Dong and Lansun Chen. A periodic predator– prey-chain system with impulsive perturbation. Journal of Computational and Applied Mathematics, 223(2): 578–584, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000514.

Dong:2009:LRG

[DC09c] Xue-Mei Dong and Di-Rong Chen. Learning rates of gradient descent algorithm for classification. Journal of Computational and Applied Mathematics, 224(1): 182–192, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800188X.

Diaz:2006:NAP

[dCDND06] J. J. del Coz Díaz, P. J. García Nieto, and F. J. Suárez Domínguez. Numerical analysis of pressure field on curved self-weighted metallic roofs due to the wind effect by the finite element method. Journal of Computational and Applied Mathematics, 192(1):40–50, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003067.

Deng:2006:DSS

[DCF06] Jiansong Deng, Falai Chen, and Yuyu Feng. Dimensions of spline spaces over T - meshes. Journal of Computational and Applied Mathematics, 194(2):267–283, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004644.

Dhia:2007:TTF

[DCZ07] A. S. Bonnet-Ben Dhia, P. Ciarlet, Jr., and C. M. Zwölf. Two- and three-field formulations for wave transmission between media with opposite sign dielectric constants. Journal of Computational and Applied Mathematics, 204(2):408–417, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003682.

DeCanditiis:2006:PCF

[DD06] Daniela De Canditiis and Italia De Feis. Pointwise convergence of Fourier regularization for smoothing data. Journal of Computational and Applied Mathematics, 196(2):540–552, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006199.

Denuit:2007:CBS

[DD07] Michel Denuit and Jan Dhaene. Comonotonic bounds on the survival probabilities in the Lee–Carter model for mortality projection. Journal of Computational and Applied Mathematics, 203(1):169–176, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001816.

Dobes:2008:SOB

[DD08] Jirí Dobes and Herman Deconinck. Second order blended multidimensional upwind residual distribution scheme for steady and unsteady computations. Journal of Computational and Applied Mathematics, 215 (2):378–389, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007485.

Dumas:2009:FAH

[DDL09] L. Dumas, B. Druez, and N. Lecerf. A fully adaptive hybrid optimization of aircraft engine blades. Journal of Computational and Applied Mathematics, 232 (1):54–60, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800544X.

Dhia:2007:THA

[DDLM07]  Anne-Sophie  Bonnet-Ben

Dhia, Ève-Marie Duclairoir, Guillaume Legendre, and Jean-François Mercier. Timeharmonic acoustic propagation in the presence of a shear flow. Journal of Computational and Applied Mathematics, 204(2):428–439, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003700.

Deelstra:2008:BAB

[DDV08] Griselda Deelstra, Ibrahima Diallo, and Michèle Vanmaele. Bounds for Asian basket options. Journal of Computational and Applied Mathematics, 218(2): 215–228, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007886.

DeLeon:2008:NWM

[De 08] Doreen De Leon. A new wavelet multigrid method. Journal of Computational and Applied Mathematics, 220(1–2):674–685, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005080.

Demiray:2007:INI

[Dem07a]  Hilmi Demiray.  Inter-

actions of nonlinear ionacoustic waves in a collisionless plasma. Journal of Computational and Applied Mathematics, 206(2): 826–831, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005255.

Demiray:2007:WFF

[Dem07b] Hilmi Demiray. Waves in fluid-filled elastic tubes with a stenosis: Variable coefficients KdV equations. Journal of Computational and Applied Mathematics, 202 (2):328–338, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001439.

Diniz-Ehrhardt:2008:DFN

[DEMR08] M. A. Diniz-Ehrhardt, J. M. Martínez, and M. Raydan. A derivative-free nonmonotone line-search technique for unconstrained optimization. Journal of Computational and Applied Mathematics, 219(2):383–397, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003834.

Deng:2007:MGA

[Den07a] Chun-Yuan Deng. On the minimum gap and the angle between two subspaces. Journal of Computational and Applied Mathematics, 206(2): 625–630, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005097.

Deng:2007:SMP

[Den07b] Weihua Deng. Short memory principle and a predictor– corrector approach for fractional differential equations. Journal of Computational and Applied Mathematics, 206(1):174–188, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004201.

Denuit:2009:ILR

[Den09] Michel M. Denuit. An index for longevity risk transfer. Journal of Computational and Applied Mathematics, 230(2):411–417, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006377.

Donatelli:2007:BCM

[DESC07] Marco Donatelli, Claudio Estatico, and Stefano Serra-Capizzano. Boundary conditions and multiple-image reblurring: the LBT case. Journal of Computational and Applied Mathematics, 198(2): 426–442, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007247.

Deuring:2007:EEP

[Deu07] Paul Deuring. Eigenvalue estimates for a preconditioned Galerkin matrix arising from mixed finite element discretizations of viscous incompressible flows. Journal of Computational and Applied Mathematics, 205(1): 453–457, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003293.

Deuring:2009:EBS

[Deu09] Paul Deuring. Eigenvalue bounds for the Schur complement with a pressure convection–diffusion preconditioner in incompressible flow computations. Journal of Computational and Applied Mathematics, 228 (1):444–457, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005104.

daFonseca:2007:EST

[dF07] C. M. da Fonseca. On the eigenvalues of some tridiagonal matrices. Journal of Computational and Applied Mathematics, 200(1): 283–286, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600015X.

Ding:2008:SIA

[DF08] Xie Ping Ding and Hai Rong Feng. The p-step iterative algorithm for a system of generalized mixed quasi-variational inclusions with (A, η)-accretive operators in q-uniformly smooth Banach spaces. Journal of Computational and Applied Mathematics, 220(1–2): 163–174, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004335.

Diethelm:2009:NAD

[DF09] Kai Diethelm and Neville J.

Ford. Numerical analysis for distributed-order differential equations. Journal of Computational and Applied Mathematics, 225(1):96–104, March 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003464.

Diethelm:2006:PFN

[DFFW06] Kai Diethelm, Judith M. Ford, Neville J. Ford, and Marc Weilbeer. Pitfalls in fast numerical solvers for fractional differential equations. Journal of Computational and Applied Mathematics, 186(2):482–503, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001287.

Dubcova:2008:NSA

[DFHS08] Lenka Dubcová, Miloslav Feistauer, Jaroslav Horácek, and Petr Svácek. Numerical simulation of airfoil vibrations induced by turbulent flow. Journal of Computational and Applied Mathematics, 218(1):34–42, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700324X.

Dolean:2008:STH

[DFLP08] V. Dolean, H. Fol, S. Lanteri, and R. Perrussel. Solution of the time-harmonic Maxwell equations using discontinuous Galerkin methods. Journal of Computational and Applied Mathematics, 218(2):

435–445, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002683.

Diogo:2006:NMV

[DFLV06] Teresa Diogo, Neville J.

Ford, Pedro Lima, and Svilen Valtchev. Numerical methods for a Volterra integral equation with non-smooth solutions. Journal of Computational and Applied Mathematics, 189(1–2):412–423, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500614X.

Dou:2009:OEB

[DFY09] Fang-Fang Dou, Chu-Li Fu, and Feng-Lian Yang. Optimal error bound and Fourier regularization for identifying an unknown source in the heat equation. Journal of Computational and Applied Mathematics, 230(2): 728–737, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000119.

DiFiore:2007:LCS

[DFZ07] Carmine Di Fiore, Stefano Fanelli, and Paolo Zellini. Low complexity secant quasi-Newton minimization algorithms for nonconvex functions. Journal of Computational and Applied Mathematics, 210(1–2):167–174, December 31, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006583.

Duran:2005:SOM

[DG05] Antonio J. Durán and F. Alberto Grünbaum. A survey on orthogonal matrix polynomials satisfying second order differential equations. Journal of Computational and Applied Mathematics, 178(1–2):169–190, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003747.

DeMaerschalck:2008:LSS

[DG08] B. De Maerschalck and M. I. Gerritsma. Least-squares spectral element method for non-linear hyperbolic differential equations. Journal of Computational and Applied Mathematics, 215 (2):357–367, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007461.

DelCorso:2007:CKS

[DGR07] Gianna M. Del Corso, Antonio Gullí, and Francesco Romani. Comparison of Krylov subspace methods on the PageRank problem. Journal of Computational and Applied Mathematics, 210(1–2): 159–166, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006571.

Decamps:2006:ASB

[DGS06] Marc Decamps, Marc Goovaerts, and Wim Schoutens. Asymmetric skew Bessel processes and their applications to finance. Journal of Computational and Applied Mathematics, 186(1):130–147, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001937.

Dancs:2006:NA

[DH06] Michael J. Dancs and Tian-Xiao He. Numerical approximation to ζ(2n + 1). Journal of Computational and Applied Mathematics, 196(1): 150–154, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005236.

Dehghan:2009:DMF

[DH09] Mehdi Dehghan and Masoud Hajarian. Determination of a matrix function using the divided difference method of Newton and the interpolation technique of Hermite. Journal of Computational and Applied Mathematics, 231(1): 67–81, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000405.

Daruis:2007:STH

[DHM07] Leyla Daruis, Javier Hernández, and Francisco Marcellán. Spectral transformations for Hermitian Toeplitz matrices. Journal of Computational and Applied Mathematics, 202(2):155–176, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001087.

Dostal:2009:SFD

[DHS09] Zdenek Dostál, David Horák, and Dan Stefanica. A scalable FETI–DP algorithm with non-penetration mortar conditions on contact interface. Journal of Computational and Applied Mathematics, 231(2):577–591, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002489.

DiCristo:2007:SDI

[Di 07] Michele Di Cristo. Stable determination of an inhomogeneous inclusion by local boundary measurements. Journal of Computational and Applied Mathematics, 198(2):414–425, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007235.

Duran:2006:DCO

[DI06] Antonio J. Duran and Mourad E. H. Ismail. Differential coefficients of orthogonal matrix polynomials. Journal of Computational and Applied Mathematics, 190(1– 2):424–436, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500244X.

Deift:2007:WDC

[DIKZ07] P. Deift, A. Its, I. Krasovsky, and X. Zhou. The Widom– Dyson constant for the gap probability in random matrix theory. Journal of Computational and Applied Mathematics, 202(1):26–47, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001130.

Ding:2005:PCG

[Din05a] Xie Ping Ding. Parametric completely generalized mixed implicit quasi-variational inclusions involving h-maximal monotone mappings. Journal of Computational and Applied Mathematics, 182(2): 252–269, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006144.

Ding:2005:PCI

[Din05b] Xie Ping Ding. Predictorcorrector iterative algorithms for solving generalized mixed quasi-variational-like inclusion. Journal of Computational and Applied Mathematics, 182(1):1–12, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005953.

Diogo:2009:CIC

[Dio09]  Teresa Diogo.  Colloca-

tion and iterated collocation methods for a class of weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 229(2):363–372, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800157X.

Damelin:2005:PCD

[DJ05] S. B. Damelin and H. S.

Jung. Pointwise convergence of derivatives of Lagrange interpolation polynomials for exponential weights. Journal of Computational and Applied Mathematics, 173(2): 303–319, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001621.

Dorao:2006:NCM

[DJ06] C. A. Dorao and H. A. Jakobsen. Numerical calculation of the moments of the population balance equation. Journal of Computational and Applied Mathematics, 196(2): 619–633, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006254.

Dorao:2007:LSS

[DJ07a] C. A. Dorao and H. A. Jakobsen. Least-squares spectral method for solving advective population balance problems. Journal of Computational and Applied Mathematics, 201(1):247–257, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001038.

Driver:2007:STZ

[DJ07b] K. Driver and K. Jordaan. Separation theorems for the zeros of certain hypergeometric polynomials. Journal of Computational and Applied Mathematics, 199(1): 48–55, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007375.

Driver:2009:ZLC

[DJ09] Kathy Driver and Kerstin Jordaan. Zeros of linear combinations of Laguerre polynomials from different sequences. Journal of Computational and Applied Mathematics, 233(3): 719–722, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001174.

Djordjevic:2007:ESO

[Djo07] Dragan S. Djordjević. Explicit solution of the operator equation A^{∗}X + X^{∗}A = B. Journal of Computational and Applied Mathematics, 200(2):701–704, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000677.

Doslic:2005:SIF

[DK05] Tomislav Doslić and Douglas J. Klein. Splinoid interpolation on finite posets. Journal of Computational and Applied Mathematics, 177(1):175–185, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004212.

Dai:2006:PII

[DK06a] Hui-Hui Dai and De-Xing Kong. The propagation of impact-induced tensile waves in a kind of phasetransforming materials. Journal of Computational and Applied Mathematics, 190 (1–2):57–73, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002219.

Dimitrov:2006:PES

[DK06b] Dobromir T. Dimitrov and Hristo V. Kojouharov. Positive and elementary stable nonstandard numerical methods with applications to predator–prey models. Journal of Computational and Applied Mathematics, 189 (1–2):98–108, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002141.

Darvishi:2008:SBS

[DKHNZ08] M. T. Darvishi, F. Khani, S. Hamedi-Nezhad, and B. Zheng. Symmetric block-SOR methods for rankdeficient least squares problems. Journal of Computational and Applied Mathematics, 215(1):14–27, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001719.

Dagnino:2005:SPL

[DL05] Catterina Dagnino and Paola Lamberti. Some performances of local bivariate quadratic C^{1} quasiinterpolating splines on nonuniform type-2 triangulations. Journal of Computational and Applied Mathematics, 173(1):21–37, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001311.

Duan:2007:LBM

[DL07a] YaLi Duan and RuXun Liu. Lattice Boltzmann model for two-dimensional unsteady Burgers’ equation. Journal of Computational and Applied Mathematics, 206(1): 432–439, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004894.

Dyllong:2007:VCH

[DL07b] Eva Dyllong and Wolfram Luther. Verified convex hull and distance computation for octree-encoded objects. Journal of Computational and Applied Mathematics, 199(2): 358–364, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007727.

Dagnino:2008:CLQ

[DL08a] C. Dagnino and P. Lamberti.

On the construction of local quadratic spline quasiinterpolants on bounded rectangular domains. Journal of Computational and Applied Mathematics, 221(2): 367–375, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005687.

DeBonis:2008:NTS

[DL08b] M. C. De Bonis and C. Laurita. Numerical treatment of second kind Fredholm integral equations systems on bounded intervals. Journal of Computational and Applied Mathematics, 217 (1):64–87, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003160.

Diogo:2008:SCM

[DL08c] Teresa Diogo and Pedro Lima. Superconvergence of collocation methods for a class of weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 218(2):307–316, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000830.

DeBonis:2009:NMC

[DL09a] M. C. De Bonis and C. Laurita. Nyström method for Cauchy singular integral equations with negative index. Journal of Computational and Applied Mathematics, 232(2):523–538, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003859.

Du:2009:PGC

[DL09b] Chaoxiong Du and Yirong Liu. The problem of general center-focus and bifurcation of limit cycles for a planar system of nine degrees. Journal of Computational and Applied Mathematics, 223(2): 1043–1057, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001441.

Duan:2009:HPD

[DL09c] Xuefeng Duan and Anping Liao. On Hermitian positive definite solution of the matrix equation X − ∑_{i=1}^{m} A_{i}^{∗}X^{r}A_{i} = Q. Journal of Computational and Applied Mathematics, 229 (1):27–36, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005153.

Du:2005:SHO

[DLG05] Zengji Du, Xiaojie Lin, and Weigao Ge. Some higherorder multi-point boundary value problem at resonance. Journal of Computational and Applied Mathematics, 177(1):55–65, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003620.

Dehesa:2005:IMH

[DLROY05] J. S. Dehesa, S. López-Rosa, B. Olmos, and R. J. Yáñez. Information measures of hydrogenic systems, Laguerre polynomials and spherical harmonics. Journal of Computational and Applied Mathematics, 179(1– 2):185–194, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004480.

Delgado:2005:ESC

[DM05] Antonia M. Delgado and Francisco Marcellán. On an extension of symmetric coherent pairs of orthogonal polynomials. Journal of Computational and Applied Mathematics, 178(1–2):155–168, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003735.

Damelin:2007:P

[DM07a] Steven B. Damelin and Wen-Xiu Ma. Preface. Journal of Computational and Applied Mathematics, 202(1): 1–2, May 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001099.

Darrigrand:2007:CUW

[DM07b] E. Darrigrand and P. Monk. Coupling of the ultra-weak variational formulation and an integral representation using a fast multipole method in electromagnetism. Journal of Computational and Applied Mathematics, 204 (2):400–407, July 15, 2007.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003670.

deMorales:2007:SOP

[dMFPP07] María Álvarez de Morales, Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. Semiclassical orthogonal polynomials in two variables. Journal of Computational and Applied Mathematics, 207(2):323–330, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600608X.

deMorales:2008:SPM

[dMFPP08] María Álvarez de Morales, Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. A semiclassical perspective on multivariate orthogonal polynomials. Journal of Computational and Applied Mathematics, 214(2):447–456, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001392.

deMorales:2009:BOP

[dMFPP09] María Álvarez de Morales, Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. Bivariate orthogonal polynomials in the lyskova class. Journal of Computational and Applied Mathematics, 233(3):597–601, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001010.

Diaz-Mendoza:2005:OLP

[DMGVP05] C. Díaz-Mendoza, P. González-Vera, and M. Jiménez Paiz. Orthogonal Laurent polynomials and two-point Padé approximants associated with Dawson’s integral. Journal of Computational and Applied Mathematics, 179(1– 2):195–213, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004492.

Dmitrochenko:2008:FEU

[Dmi08]  Oleg Dmitrochenko.  Fi-

nite elements using absolute nodal coordinates for largedeformation flexible multibody dynamics. Journal of Computational and Applied Mathematics, 215 (2):368–377, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007473.

DAmore:2007:ISI

[DMM07] L. D’Amore, L. Marcellino, and A. Murli. Image sequence inpainting: Towards numerical software for detection and removal of local missing data via motion estimation. Journal of Computational and Applied Mathematics, 198(2):396–413, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007223.

Dattoli:2005:BSF

[DMS05] G. Dattoli, M. Migliorati, and H. M. Srivastava. Bessel summation formulae and operational methods. Journal of Computational and Applied Mathematics, 173(1): 149–154, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400144X.

Duan:2008:STO

[DMZ08] Xian-Bao Duan, Yi-Chen Ma, and Rui Zhang. Shapetopology optimization for Navier–Stokes problem using variational level set method. Journal of Computational and Applied Mathematics, 222(2):487–499, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006255.

Dobrogowska:2006:SOD

[DO06] Alina Dobrogowska and Anatol Odzijewicz. Second order q-difference equations solvable by factorization method. Journal of Computational and Applied Mathematics, 193(1):319–346, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003948.

Dimitriyadis:2009:DHI

[DÖ09] I. Dimitriyadis and Ü. N. Öney. Deductibles in health insurance. Journal of Computational and Applied Mathematics, 233(1): 51–60, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004640.

Dolejsi:2008:AAI

[Dol08] Vít Dolejsí. Analysis and application of the IIPG method to quasilinear nonstationary convection–diffusion problems. Journal of Computational and Applied Mathematics, 222(2):251–273, December 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700595X.

Dominici:2007:SRP

[Dom07] Diego Dominici. Some remarks on a paper by L. Carlitz. Journal of Computational and Applied Mathematics, 198(1):129–142, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007107.

Dominici:2009:AAB

[Dom09] Diego Dominici. Asymptotic analysis of the Bell polynomials by the ray method. Journal of Computational and Applied Mathematics, 233(3): 708–718, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001162.

Dong:2009:IMM

[Don09] Jun-Liang Dong. Inexact multisplitting methods for linear complementarity problems. Journal of Computational and Applied Mathematics, 223(2):714–724, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000691.

Dorao:2009:STD

[Dor09] C. A. Dorao. Simulation of thermal disturbances with finite wave speeds using a high order method. Journal of Computational and Applied Mathematics, 231(2): 637–647, September 15, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002635.

Dehesa:2008:PBF

[DOY08] J. S. Dehesa, B. Olmos, and R. J. Yáñez. Parameter-based Fisher’s information of orthogonal polynomials. Journal of Computational and Applied Mathematics, 214 (1):136–147, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001021.

Danumjaya:2005:OCS

[DP05a] P. Danumjaya and Amiya K. Pani. Orthogonal cubic spline collocation method for the extended Fisher– Kolmogorov equation. Journal of Computational and Applied Mathematics, 174(1): 101–117, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001748.

DeBoer:2005:EDD

[DP05b] Rob J. De Boer and Alan S. Perelson. Estimating division and death rates from CFSE data. Journal of Computational and Applied Mathematics, 184(1):140– 164, December 1, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000749.

Delgado:2008:EAE

[DP08a] J. Delgado and J. M. Peña.

Error analysis of efficient evaluation algorithms for tensor product surfaces. Journal of Computational and Applied Mathematics, 219(1): 156–169, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003858.

Delis:2008:RAB

[DP08b] A. I. Delis and I. Papoglou.

Relaxation approximation to bed-load sediment transport. Journal of Computational and Applied Mathematics, 213(2):521–546, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000702.

Debrabant:2008:CWA

[DR08] Kristian Debrabant and Andreas Rößler. Continuous weak approximation for stochastic differential equations. Journal of Computational and Applied Mathematics, 214(1):259–273, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001215.

Dimitrov:2009:MZL

[DR09] Dimitar K. Dimitrov and Fernando R. Rafaeli. Monotonicity of zeros of Laguerre polynomials. Journal of Computational and Applied Mathematics, 233(3): 699–702, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001149.

DeSterck:2009:FAA

[DRT09] Hans De Sterck, Scott Rostrup, and Feng Tian. A fast and accurate algorithm for computing radial transonic flows. Journal of Computational and Applied Mathematics, 223(2):916–928, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001040.

Deliege:2008:ISS

[DRV08] Geoffrey Deliége, Eveline Rosseel, and Stefan Vandewalle. Iterative solvers and stabilisation for mixed electrostatic and magnetostatic formulations. Journal of Computational and Applied Mathematics, 215 (2):348–356, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600745X.

Dette:2005:NMM

[DS05a] Holger Dette and W. J. Studden. A note on the maximization of matrix valued Hankel determinants with applications. Journal of Computational and Applied Mathematics, 177(1):129–140, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004182.

Drummond:2005:SDM

[DS05b] L. M. Graña Drummond and B. F. Svaiter. A steepest descent method for vector optimization. Journal of Computational and Applied Mathematics, 175(2): 395–414, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003127.

Dur:2005:PSS

[DS05c] Mirjam Dür and Volker Stix. Probabilistic subproblem selection in branch-and-bound algorithms. Journal of Computational and Applied Mathematics, 182 (1):67–80, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006016.

Degani:2006:RRC

[DS06] Ilan Degani and Jeremy Schiff. RCMS: Right Correction Magnus Series approach for oscillatory ODEs. Journal of Computational and Applied Mathematics, 193(2): 413–436, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500422X.

daSilva:2007:NMR

[dS07a] M. A. Salgueiro da Silva. A novel method for robust minimisation of univariate functions with quadratic convergence. Journal of Computational and Applied Mathematics, 200(1):168–177, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008034.

Driver:2007:UOM

[DS07b] Kathy Driver and Herbert Stahl. The uniqueness of an orthogonality measure. Journal of Computational and Applied Mathematics, 207(2): 180–185, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006005.

Dehghan:2008:AHV

[DS08] Mehdi Dehghan and Fatemeh Shakeri. Application of He’s variational iteration method for solving the Cauchy reaction–diffusion problem. Journal of Computational and Applied Mathematics, 214(2):435–446, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001380.

Daniel:2009:IPN

[DS09a] Sunita Daniel and P. Shunmugaraj. An interpolating 6point C^{2} non-stationary subdivision scheme. Journal of Computational and Applied Mathematics, 230(1): 164–172, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005967.

Dehghan:2009:NSN

[DS09b] Mehdi Dehghan and Ali Shokri. Numerical solution of the nonlinear Klein– Gordon equation using radial basis functions. Journal of Computational and Applied Mathematics, 230(2): 400–410, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006365.

Dai:2009:EEO

[DSC09] Xiaoxia Dai, Jie Sun, and Xiaoliang Cheng. Error estimates for an operatorsplitting method for Navier– Stokes equations: Secondorder schemes. Journal of Computational and Applied Mathematics, 231(2): 696–704, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002696.

See corrigendum [Dai11].

Dag:2006:GMN

[DSI06] Idris Dag, Bülent Saka, and Dursun Irk. Galerkin method for the numerical solution of the RLW equation using quintic B-splines. Journal of Computational and Applied Mathematics, 190(1– 2):532–547, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002785.

Dehesa:2006:CRI

[DSMY06] J. S. Dehesa, P. Sánchez-Moreno, and R. J. Yáñez. Cramer–Rao information plane of orthogonal hypergeometric polynomials. Journal of Computational and Applied Mathematics, 186(2): 523–541, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001305.

Davies:2008:NUS

[DSW08] Richard Davies, Peng Shi, and Ron Wiltshire. New upper solution bounds of the discrete algebraic Riccati matrix equation. Journal of Computational and Applied Mathematics, 213 (2):307–315, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000349.

Dattoli:2005:OPH

[DSZ05] G. Dattoli, H. M. Srivastava, and K. Zhukovsky. Orthogonality properties of the Hermite and related polynomials. Journal of Computational and Applied Mathematics, 182(1):165–172, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006089.

Dai:2006:NSS

[DT06a] Gui-Dong Dai and Yi-Fa Tang. A note on symplecticity of step-transition mappings for multi-step methods. Journal of Computational and Applied Mathematics, 196(2): 474–477, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005984.

Duan:2006:MGM

[DT06b] Yong Duan and Yong-Ji Tan. A meshless Galerkin method for Dirichlet problems using radial basis functions. Journal of Computational and Applied Mathematics, 196(2): 394–401, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005832.

Du:2007:NSI

[Du07a] Kui Du. Note on structured indefinite perturbations to Hermitian matrices. Journal of Computational and Applied Mathematics, 202 (2):258–265, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001270.

Du:2007:BDR

[Du07b] Lili Du. Blow-up for a degenerate reaction–diffusion system with nonlinear nonlocal sources. Journal of Computational and Applied Mathematics, 202(2):237–247, May 15, 2007. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001257.

Dubeau:2009:NMH

[Dub09]  François Dubeau.  New-

ton’s method and highorder algorithms for the nth root computation. Journal of Computational and Applied Mathematics, 224(1): 66–76, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001775.

Duff:2009:DUS

[Duf09] Iain S. Duff. The design and use of a sparse direct solver for skew symmetric matrices. Journal of Computational and Applied Mathematics, 226(1):50–54, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002203.

Dular:2008:TDH

[Dul08]  Patrick Dular.  A time-

domain homogenization technique for lamination stacks in dual finite element formulations. Journal of Computational and Applied Mathematics, 215(2):390–399, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007497.

Dunster:2006:UAA

[Dun06]  T. M. Dunster.  Uni-

form asymptotic approximations for incomplete Riemann zeta functions. Journal of Computational and Applied Mathematics, 190(1– 2):339–353, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002396.

Dunkl:2007:HLP

[Dun07] Charles F. Dunkl. Hooklengths and pairs of compositions. Journal of Computational and Applied Mathematics, 199(1):39–47, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007363.

Dussault:2005:HON

[Dus05] Jean-Pierre Dussault. Highorder Newton-penalty algorithms. Journal of Computational and Applied Mathematics, 182(1):117–133, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006053.

Delvaux:2006:RSP

[DV06a] Steven Delvaux and Marc Van Barel. Rank structures preserved by the QRalgorithm: the singular case. Journal of Computational and Applied Mathematics, 189(1–2):157–178, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001329.

Delvaux:2006:SPA

[DV06b] Steven Delvaux and Marc Van Barel. Structures preserved by the QR-algorithm. Journal of Computational and Applied Mathematics, 187(1):29–40, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001330.

Delvaux:2008:ECU

[DV08a] Steven Delvaux and Marc Van Barel. Eigenvalue computation for unitary rank structured matrices. Journal of Computational and Applied Mathematics, 213(1): 268–287, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000350.

Delvaux:2008:URS

[DV08b] Steven Delvaux and Marc Van Barel. Unitary rank structured matrices. Journal of Computational and Applied Mathematics, 215 (1):49–78, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001744.

Delage:2006:HLA

[DVCH06] Stephanie Delage, Stephane Vincent, Jean-Paul Caltagirone, and Jean-Philippe Heliot. A hybrid linking approach for solving the conservation equations with an adaptive mesh refinement method. Journal of Computational and Applied Mathematics, 191(2):280–296, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500556X.

Dick:2008:P

[DVD^{+}08] Erik Dick, Jan Vierendeels, Luc Dupré, Lieven Vandevelde, Roger Van Keer, and Marián Slodicka. Preface. Journal of Computational and Applied Mathematics, 215(2):303, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007394.

Diao:2005:SPG

[DW05] Huaian Diao and Yimin Wei.

Structured perturbations of group inverse and singular linear system with index one. Journal of Computational and Applied Mathematics, 173(1):93–113, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001360.

Du:2006:SPS

[DW06a] Kui Du and Yimin Wei.

Structured pseudospectra and structured sensitivity of eigenvalues. Journal of Computational and Applied Mathematics, 197(2): 502–519, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006965.

Du:2006:RPR

[DW06b] Qiang Du and Desheng Wang.

Recent progress in robust and quality Delaunay mesh generation. Journal of Computational and Applied Mathematics, 195(1–2):8–23, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004681.

Dai:2006:FST

[DWW06] Yongming Dai, Zhongcheng Wang, and Dongmei Wu. A four-step trigonometric fitted P -stable Obrechkoff method for periodic initial-value problems. Journal of Computational and Applied Mathematics, 187(2):192–201, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001524.

Dong:2009:RRM

[DWY09] Yinghui Dong, Guojing Wang, and Kam C. Yuen. On the renewal risk model under a threshold strategy. Journal of Computational and Applied Mathematics, 230 (1):22–33, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005839.

Ding:2009:NDS

[DZ09a] Hengfei Ding and Yuxin Zhang. A new difference scheme with high accuracy and absolute stability for solving convection– diffusion equations. Journal of Computational and Applied Mathematics, 230(2): 600–606, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006547.

Ding:2009:NFO

[DZ09b] Hengfei Ding and Yuxin Zhang. A new fourthorder compact finite difference scheme for the twodimensional second-order hyperbolic equation. Journal of Computational and Applied Mathematics, 230(2): 626–632, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900003X.

Duan:2006:BRI

[DZLL06] Qi Duan, Huanling Zhang, Aikui Liu, and Huaigu Li. A bivariate rational interpolation with a bi-quadratic denominator. Journal of Computational and Applied Mathematics, 195(1–2):24– 33, October 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004693.

Duan:2007:EEK

[DZZT07] Qi Duan, Huanling Zhang, Yunfeng Zhang, and E. H. Twizell. Error estimation of a kind of rational spline. Journal of Computational and Applied Mathematics, 200 (1):1–11, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007910.

Ebaid:2009:ESG

[Eba09] A. Ebaid. Exact solutions for the generalized Klein–Gordon equation via a transformation and exp-function method and comparison with Adomian’s method. Journal of Computational and Applied Mathematics, 223(1):278–290, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000125.

Erbe:2009:NSO

[EBP09] Lynn Erbe, Jia Baoguo, and Allan Peterson. Nonoscillation for second order sublinear dynamic equations on time scales. Journal of Computational and Applied Mathematics, 232(2): 594–599, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003902.

El-Daou:2006:PEB

[ED06] Mohamed K. El-Daou. A posteriori error bounds for the approximate solution of second-order ODEs by piecewise coefficients perturbation methods. Journal of Computational and Applied Mathematics, 189(1–2):51–66, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500018X.

Edstrom:2009:NPS

[Eds09] Per Edström. Numerical performance of stability enhancing and speed increasing steps in radiative transfer solution methods. Journal of Computational and Applied Mathematics, 228(1):104–114, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004421.

Edmunds:2006:SPA

[EE06] D. E. Edmunds and W. D.

Evans. Spectral problems on arbitrary open subsets of R^{n} involving the distance to the boundary. Journal of Computational and Applied Mathematics, 194(1): 36–53, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004024.

ElGhazi:2008:NMC

[EEGG08] Abdellatif El Ghazi, Said El Hajji, Luc Giraud, and Serge Gratton. Newton’s method for the common eigenvector problem. Journal of Computational and Applied Mathematics, 219(2): 398–407, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006000.

Easwaran:2006:CMD

[EFP06] Chirakkal Easwaran, Lawrence Fialkow, and Srdjan Petrovic. Can a minimal degree 6 cubature rule for the disk have all points inside? Journal of Computational and Applied Mathematics, 185(1): 144–165, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000543.

ElBaz:2005:AIF

[EFS05] Didier El Baz, Andreas Frommer, and Pierre Spiteri. Asynchronous iterations with flexible communication: contracting operators. Journal of Computational and Applied Mathematics, 176 (1):91–103, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400322X.

El-Gebeily:2009:FEG

[EGFO09] Mohamed A. El-Gebeily, Khaled M. Furati, and Donal O’Regan. The finite element-Galerkin method for singular self-adjoint differential equations. Journal of Computational and Applied Mathematics, 223(2):735–752, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800071X.

El-Gebeily:2009:ANS

[EGFOA09] M. A. El-Gebeily, K. M. Furati, Donal O’Regan, and Ravi Agarwal. On the approximation of nonlinear singular self-adjoint second order boundary value problems. Journal of Computational and Applied Mathematics, 224(1):360–372, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002148.

Egidi:2007:ESE

[EGM07] Nadaniela Egidi, Romina Gobbi, and Pierluigi Maponi. The efficient solution of electromagnetic scattering for inhomogeneous media. Journal of Computational and Applied Mathematics, 210(1–2): 175–182, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006595.

El-Gebeily:2006:GQM

[EGO06] Mohamed El-Gebeily and Donal O’Regan. A generalized quasilinearization method for second-order nonlinear differential equations with nonlinear boundary conditions. Journal of Computational and Applied Mathematics, 192(2):270–281, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003201.

Ezquerro:2006:OCN

[EH06] J. A. Ezquerro and M. A. Hernández. On the R-order of convergence of Newton’s method under mild differentiability conditions. Journal of Computational and Applied Mathematics, 197(1): 53–61, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006345.

Ehrenmark:2006:ATC

[Ehr06] Ulf Ehrenmark. Application of the W -transformation to compute the linearised 2-D beach problem potentials. Journal of Computational and Applied Mathematics, 197(2):457–464, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500693X.

Ezquerro:2009:IEI

[EHR09] J. A. Ezquerro, M. A.

Hernández, and N. Romero. Improving the efficiency index of one-point iterative processes. Journal of Computational and Applied Mathematics, 223(2):879–892, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000915.

Eigenwillig:2007:MRD

[Eig07] Arno Eigenwillig. On multiple roots in Descartes’ rule and their distance to roots of higher derivatives. Journal of Computational and Applied Mathematics, 200 (1):226–230, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008083.

ElGhami:2009:PTA

[EIM^{+}09] M. El Ghami, I. Ivanov, J. B. M. Melissen, C. Roos, and T. Steihaug. A polynomial-time algorithm for linear optimization based on a new class of kernel functions. Journal of Computational and Applied Mathematics, 224(2):500–513, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002574.

Elliott:2007:EAT

[EJ07] David Elliott and Peter R. Johnston. Error analysis for sinh transformation used in evaluating nearly singular boundary element integrals. Journal of Computational and Applied Mathematics, 203(1):103–124, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001774.

Elliott:2008:GLQ

[EJ08] David Elliott and Peter R. Johnston. Gauss–Legendre quadrature for the evaluation of integrals involving the Hankel function. Journal of Computational and Applied Mathematics, 211(1): 23–35, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006923.

Everitt:2007:BDE

[EK07] W. N. Everitt and H. Kalf. The Bessel differential equation and the Hankel transform. Journal of Computational and Applied Mathematics, 208(1):3–19, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006236.

El-Kabeir:2008:LGA

[EKEHR08] S. M. M. El-Kabeir, M. A. El-Hakiem, and A. M. Rashad. Lie group analysis of unsteady MHD three dimensional by natural convection from an inclined stretching surface saturated porous medium. Journal of Computational and Applied Mathematics, 213(2):582–603, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700074X.

Everitt:2007:JSN

[EKL^{+}07] W. N. Everitt, K. H. Kwon, L. L. Littlejohn, R. Wellman, and G. J. Yoon. Jacobi– Stirling numbers, Jacobi polynomials, and the leftdefinite analysis of the classical Jacobi differential expression. Journal of Computational and Applied Mathematics, 208(1):29–56, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600625X.

Eitrich:2006:EOS

[EL06] Tatjana Eitrich and Bruno Lang. Efficient optimization of support vector machine learning parameters for unbalanced datasets. Journal of Computational and Applied Mathematics, 196(2): 425–436, November 15, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005856.

Edmunds:2007:OHT

[EL07] D. E. Edmunds and J. Lang. Operators of Hardy type. Journal of Computational and Applied Mathematics, 208(1):20–28, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006248.

Elaydi:2009:SAV

[Ela09a] Saber Elaydi. Stability and asymptoticity of Volterra difference equations: a progress report. Journal of Computational and Applied Mathematics, 228(2):504–513, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001337.

Eshkuvatov:2009:QFA

[ELA09b] Z. K. Eshkuvatov, N. M. A. Nik Long, and M. Abdulkawi. Quadrature formula for approximating the singular integral of Cauchy type with unbounded weight function on the edges. Journal of Computational and Applied Mathematics, 233(2): 334–345, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004257.

Elliott:2006:OCH

[Ell06] C. James Elliott. Orthogonal confluent hypergeometric lattice polynomials. Journal of Computational and Applied Mathematics, 193(1): 89–108, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500381X.

Egidi:2006:SMA

[EM06] N. Egidi and P. Maponi. A Sherman–Morrison approach to the solution of linear systems. Journal of Computational and Applied Mathematics, 189(1–2):703–718, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001032.

Egidi:2007:HEG

[EM07] N. Egidi and P. Maponi.

The heat equation to generate planar grids. Journal of Computational and Applied Mathematics, 210(1–2): 183–190, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006601.

Egidi:2008:ASS

[EM08a] N. Egidi and P. Maponi. An active set strategy to generate quadrilateral grids. Journal of Computational and Applied Mathematics, 218(2): 492–505, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700413X.

Egidi:2008:PTI

[EM08b] Nadaniela Egidi and Pierluigi Maponi. Preconditioning techniques for the iterative solution of scattering problems. Journal of Computational and Applied Mathematics, 218(2): 229–237, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706008004.

Erturk:2008:SSF

[EM08c] Vedat Suat Ertürk and Shaher Momani. Solving systems of fractional differential equations using differential transform method. Journal of Computational and Applied Mathematics, 215 (1):142–151, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001811.

Erbao:2009:HDE

[EM09] Cao Erbao and Lai Mingyong. A hybrid differential evolution algorithm to vehicle routing problem with fuzzy demands. Journal of Computational and Applied Mathematics, 231(1): 302–310, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000636.

Abd-el-Malek:2005:CFM

[eMH05] M. B. Abd el Malek and M. M. Helal. Characteristic function method for classification of equations of hydrodynamics of a perfect fluid. Journal of Computational and Applied Mathematics, 182(1):105–116, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006041.

Abd-el-Malek:2006:CFM

[eMH06a] Mina B. Abd el Malek and Medhat M. Helal. The characteristic function method and exact solutions of nonlinear sheared flows with free surface under gravity. Journal of Computational and Applied Mathematics, 189 (1–2):2–21, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002979.

Abd-el-Malek:2006:SAM

[eMH06b] Mina B. Abd el Malek and Medhat M. Helal. Semianalytical method for solving nonlinear heat diffusion problems in spherical medium. Journal of Computational and Applied Mathematics, 193(1):10–21, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003766.

Abd-el-Malek:2008:SSM

[eMH08] Mina B. Abd el Malek and Medhat M. Helal. Similarity solutions for magneto-forced-unsteady free convective laminar boundary-layer flow. Journal of Computational and Applied Mathematics, 218(2):202–214, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007163.

Everitt:2007:CUS

[EMMP07] W. N. Everitt, L. Markus, M. Muzzulini, and M. Plum. A continuum of unusual selfadjoint linear partial differential operators. Journal of Computational and Applied Mathematics, 208(1): 164–175, November 1, 2007.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006339.

Eymard:2008:IFV

[EMP08] Robert Eymard, Sophie Mercier, and Alain Prignet. An implicit finite volume scheme for a scalar hyperbolic problem with measure data related to piecewise deterministic Markov processes. Journal of Computational and Applied Mathematics, 222(2): 293–323, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005985.

Engblom:2009:SAS

[Eng09] Stefan Engblom. Spectral approximation of solutions to the chemical master equation. Journal of Computational and Applied Mathematics, 229(1):208–221, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005578.

Enright:2006:VAS

[Enr06] W. H. Enright. Verifying approximate solutions to differential equations. Journal of Computational and Applied Mathematics, 185(2): 203–211, January 15, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500110X.

Estep:2009:PEA

[EPP^{+}09] Don Estep, Michael Pernice, Du Pham, Simon Tavener, and Haiying Wang. A posteriori error analysis of a cell-centered finite volume method for semilinear elliptic problems. Journal of Computational and Applied Mathematics, 233(2): 459–472, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900452X.

Erbe:2005:ABS

[EPS05] L. Erbe, A. Peterson, and S. H. Saker. Asymptotic behavior of solutions of a thirdorder nonlinear dynamic equation on time scales. Journal of Computational and Applied Mathematics, 181(1): 92–102, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005552.

Epshteyn:2009:DGM

[Eps09] Yekaterina Epshteyn. Discontinuous Galerkin methods for the chemotaxis and haptotaxis models. Journal of Computational and Applied Mathematics, 224(1): 168–181, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001878.

Epshteyn:2007:EPP

[ER07] Yekaterina Epshteyn and Béatrice Rivière. Estimation of penalty parameters for symmetric interior penalty Galerkin methods. Journal of Computational and Applied Mathematics, 206(2): 843–872, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005279.

Epshteyn:2009:ADG

[ER09] Yekaterina Epshteyn and Beatrice Rivière. Analysis of hp discontinuous Galerkin methods for incompressible two-phase flow. Journal of Computational and Applied Mathematics, 225(2): 487–509, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004214.

El-Sayed:2005:NIF

[ESAD05] Salah M. El-Sayed and Asmaa M. Al-Dbiban. A new inversion free iteration for solving the equation X + A^{⋆}X^{−1}A = Q. Journal of Computational and Applied Mathematics, 181(1): 148–156, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005588.

Escribano:2007:FPT

[EST07a] C. Escribano, M. A. Sastre, and E. Torrano. A fixed point theorem for moment matrices of self-similar measures. Journal of Computational and Applied Mathematics, 207(2):352–359, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006091.

Estatico:2007:RFM

[Est07b] Claudio Estatico. Regularized fast multiple-image deconvolution for LBT. Journal of Computational and Applied Mathematics, 198(2): 443–459, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007259.

Erhan:2006:MCQ

[ET06] Inci M. Erhan and H. Taseli. A model for the computation of quantum billiards with arbitrary shapes. Journal of Computational and Applied Mathematics, 194(2): 227–244, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004528.

Elteto:2008:NAM

[ÉT08] T. Éltető and M. Telek. Numerical analysis of M/G/1 type queueing systems with phase type transition structure. Journal of Computational and Applied Mathematics, 212(2):331–340, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007345.

Exton:1999:NTT

[Ext99] H. Exton. A new two-term relation for the _{3}F_{2} hypergeometric function of unit argument. Journal of Computational and Applied Mathematics, 106(2):395–397, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042799000771.

See comment [Mil06].

Ehrhardt:2006:FCE

[EZ06] Matthias Ehrhardt and Andrea Zisowsky. Fast calculation of energy and mass preserving solutions of Schrödinger–Poisson systems on unbounded domains. Journal of Computational and Applied Mathematics, 187 (1):1–28, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001317.

Ehrhardt:2007:DNL

[EZ07] Matthias Ehrhardt and Andrea Zisowsky. Discrete non-local boundary conditions for split-step Padé approximations of the one-way Helmholtz equation. Journal of Computational and Applied Mathematics, 200(2): 471–490, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600029X.

Fairag:2007:AFE

[Fai07] Faisal A. Fairag. Analysis and finite element approximation of a Ladyzhenskaya model for viscous flow in streamfunction form. Journal of Computational and Applied Mathematics, 206(1): 374–391, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004754.

Fan:2009:GCI

[Fan09] Xiaona Fan. A globally convergent interior point algorithm for non-convex nonlinear programming. Journal of Computational and Applied Mathematics, 224(2): 622–627, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002793.

Farin:2006:DSS

[Far06] Gerald Farin. Dimensions of spline spaces over unconstricted triangulations. Journal of Computational and Applied Mathematics, 192(2): 320–327, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003535.

Ford:2009:SBD

[FC09] Neville J. Ford and Joseph A.

Connolly. Systems-based decomposition schemes for the approximate solution of multi-term fractional differential equations. Journal of Computational and Applied Mathematics, 229(2): 382–391, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001593.

Feng:2009:NSG

[FD09] Hai Rong Feng and Xie Ping Ding. A new system of generalized nonlinear quasivariational-like inclusions with A-monotone operators in Banach spaces. Journal of Computational and Applied Mathematics, 225(2): 365–373, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003841.

Fan:2008:ZUI

[FDC08] Xuchuan Fan, Jiansong Deng, and Falai Chen. Zeros of univariate interval polynomials. Journal of Computational and Applied Mathematics, 216(2):563–573, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002968.

Ford:2007:NMQ

[FDFL07] Neville J. Ford, Teresa Diogo, Judith M. Ford, and Pedro Lima. Numerical modelling of qualitative behaviour of solutions to convolution integral equations. Journal of Computational and Applied Mathematics, 205(2): 849–858, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004079.

Fernandez:2007:OSL

[Fer07] Noemí Laín Fernández. Optimally space-localized bandlimited wavelets on S^{q−1}. Journal of Computational and Applied Mathematics, 199(1):68–79, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007399.

Ferreira:2009:NQS

[Fer09] Raúl Ferreira. Numerical quenching for the semilinear heat equation with a singular absorption. Journal of Computational and Applied Mathematics, 228(1):92–103, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800441X.

Fill:2005:SAH

[FFK05] James Allen Fill, Philippe Flajolet, and Nevin Kapur. Singularity analysis, Hadamard products, and tree recurrences. Journal of Computational and Applied Mathematics, 174(2): 271–313, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002171.

Fokas:2009:SAN

[FFSS09] A. S. Fokas, N. Flyer, S. A. Smitheman, and E. A. Spence. A semi-analytical numerical method for solving evolution and elliptic partial differential equations. Journal of Computational and Applied Mathematics, 227(1):59–74, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003294.

Fridman:2007:SLS

[FG07] Emilia Fridman and Michael Gil’. Stability of linear systems with time-varying delays: a direct frequency domain approach. Journal of Computational and Applied Mathematics, 200 (1):61–66, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500796X.

Fornasier:2008:STB

[FG08] Massimo Fornasier and Laura Gori. Sampling theorems on bounded domains. Journal of Computational and Applied Mathematics, 221(2): 376–385, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005699.

Fata:2009:SAI

[FG09] S. Nintcheu Fata and L. J.

Gray. Semi-analytic integration of hypersingular Galerkin BIEs for threedimensional potential problems. Journal of Computational and Applied Mathematics, 231(2):561–576, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709002477.

Fazzolari:2007:EAS

[FGB07] Antonio Fazzolari, Nicolas R.

Gauger, and Joël Brezillon. Efficient aerodynamic shape optimization in MDO context. Journal of Computational and Applied Mathematics, 203(2):548–560, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002214.

Fitt:2009:FDE

[FGRW09] A. D. Fitt, A. R. H. Goodwin, K. A. Ronaldson, and W. A. Wakeham. A fractional differential equation for a MEMS viscometer used in the oil industry. Journal of Computational and Applied Mathematics, 229(2):373–381, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001581.

Farouki:2007:EBC

[FH07] Rida T. Farouki and Joel Hass. Evaluating the boundary and covering degree of planar Minkowski sums and other geometrical convolutions. Journal of Computational and Applied Mathematics, 209(2):246–266, December 15, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600690X.

Fang:2009:SMN

[FH09a] Liang Fang and Guoping He. Some modifications of Newton’s method with higherorder convergence for solving nonlinear equations. Journal of Computational and Applied Mathematics, 228 (1):296–303, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004871.

Fernandez:2009:PPE

[FH09b] J. R. Fernández and P. Hild. A priori and a posteriori error analyses in the study of viscoelastic problems. Journal of Computational and Applied Mathematics, 225(2): 569–580, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004299.

Frochte:2009:STH

[FH09c] Jörg Frochte and Wilhelm Heinrichs. A splitting technique of higher order for the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 228(1):373–390, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005049.

Fang:2008:CCN

[FHH08] Liang Fang, Guoping He, and Zhongyong Hu. A cubically convergent Newtontype method under weak conditions. Journal of Computational and Applied Mathematics, 220(1–2):409–412, October 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700461X.

Frommer:2007:PEZ

[FHL07] A. Frommer, F. Hoxha, and B. Lang. Proving the existence of zeros using the topological degree and interval arithmetic. Journal of Computational and Applied Mathematics, 199(2): 397–402, February 15, 2007.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007788.

Feng:2009:AMH

[FHM09] Xinlong Feng, Yinnian He, and Jixiang Meng. Application of modified homotopy perturbation method for solving the augmented systems. Journal of Computational and Applied Mathematics, 231(1):288–301, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000624.

Fijnvandraat:2006:TDA

[FHtMP06] J. G. Fijnvandraat, S. H. M. J. Houben, E. J. W. ter Maten, and J. M. F. Peters. Time domain analog circuit simulation. Journal of Computational and Applied Mathematics, 185(2): 441–459, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001251.

Farago:2009:RES

[FHZ09] István Faragó, Ágnes Havasi, and Zahari Zlatev. Richardsonextrapolated sequential splitting and its application. Journal of Computational and Applied Mathematics, 226 (2):218–227, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004020.

Fasino:2007:RNT

[FI07] Dario Fasino and Gabriele Inglese. Recovering nonlinear terms in an inverse boundary value problem for Laplace’s equation: a stability estimate. Journal of Computational and Applied Mathematics, 198(2):460–470, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007260.

Finden:2007:HOA

[Fin07] W. F. Finden. Higher order approximations using interpolation applied to collocation solutions of two-point boundary value problems. Journal of Computational and Applied Mathematics, 206(1):99–115, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003840.

Finden:2008:ETU

[Fin08]  W. F. Finden.  An er-

ror term and uniqueness for Hermite–Birkhoff interpolation involving only function values and/or first derivative values. Journal of Computational and Applied Mathematics, 212(1):1–15, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007072.

Fang:2006:HML

[FIR06] Weifu Fang, Kazufumi Ito, and David A. Redfern. A homogenization model for laser beam-induced current imaging and detection of nonuniformities in semiconductor arrays. Journal of Computational and Applied Mathematics, 194(2):395–408, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005029.

Feng:2008:PSC

[FJG08] Meiqiang Feng, Dehong Ji, and Weigao Ge. Positive solutions for a class of boundary-value problem with integral boundary conditions in Banach spaces. Journal of Computational and Applied Mathematics, 222(2): 351–363, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006036.

Feng:2007:CSS

[FJT07] Quan-Dong Feng, Yan-Dong Jiao, and Yi-Fa Tang. Conjugate symplecticity of secondorder linear multi-step methods. Journal of Computational and Applied Mathematics, 203(1):6–14, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001555.

Fritzsche:2005:HMR

[FKL05] Bernd Fritzsche, Bernd Kirstein, and Andreas Lasarow. On Hilbert modules of rational matrix-valued functions and related inverse problems. Journal of Computational and Applied Mathematics, 179(1–2):215–248, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004509.

Francisco:2005:IPM

[FKM05] J. B. Francisco, N. Krejić, and J. M. Martínez. An interior-point method for solving box-constrained underdetermined nonlinear systems. Journal of Computational and Applied Mathematics, 177(1):67–88, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400398X.

Fan:2006:GAS

[FL06a] Yong-Hong Fan and Wan-Tong Li. Global asymptotic stability of a ratio-dependent predator–prey system with diffusion. Journal of Computational and Applied Mathematics, 188(2):205–227, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002505.

Firsov:2006:DDM

[FL06b] D. Firsov and S. H. Lui. Domain decomposition methods in image denoising using Gaussian curvature. Journal of Computational and Applied Mathematics, 193(2): 460–473, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004255.

Ford:2006:ADS

[FL06c] Neville J. Ford and Patricia M. Lumb. An algorithm to detect small solutions in linear delay differential equations. Journal of Computational and Applied Mathematics, 193(1):121–139, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003833.

Foufas:2008:VAO

[FL08a] Georgios Foufas and Mats G.

Larson. Valuing Asian options using the finite element method and duality techniques. Journal of Computational and Applied Mathematics, 222(1): 144–158, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005547.

Fu:2008:STN

[FL08b] Xilin Fu and Xiaodi Li. W - stability theorems of nonlinear impulsive functional differential systems. Journal of Computational and Applied Mathematics, 221(1): 33–46, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005213.

Feng:2009:SPQ

[FL09a] Renzhong Feng and Feng Li. A shape-preserving quasi-interpolation operator satisfying quadratic polynomial reproduction property to scattered data. Journal of Computational and Applied Mathematics, 225(2): 594–601, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004317.

Ford:2009:MTF

[FL09b] Neville J. Ford and Patricia M. Lumb. Mixedtype functional differential equations: a numerical approach. Journal of Computational and Applied Mathematics, 229(2):471–479, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001672.

Fu:2009:GES

[FL09c] Xilin Fu and Xiaodi Li. Global exponential stability and global attractivity of impulsive Hopfield neural networks with time delays. Journal of Computational and Applied Mathematics, 231(1): 187–199, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000521.

Fu:2009:RTT

[FL09d] Xilin Fu and Xiaodi Li. Razumikhin-type theorems on exponential stability of impulsive infinite delay differential systems. Journal of Computational and Applied Mathematics, 224 (1):1–10, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001489.

Favati:2007:ASI

[FLMR07] P. Favati, G. Lotti, O. Menchi, and F. Romani. Adaptive solution of infinite linear systems by Krylov subspace methods. Journal of Computational and Applied Mathematics, 210(1–2):191– 199, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006613.

Fevriere:2009:PPM

[FLPA09] C. Févrière, J. Laminie, P. Poullet, and Ph. Angot. On the penalty-projection method for the Navier–Stokes equations with the MAC mesh. Journal of Computational and Applied Mathematics, 226(2):228–245, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004032.

Filbir:2005:HCT

[FLS05] Frank Filbir, Rupert Lasser, and Ryszard Szwarc. Hypergroups of compact type. Journal of Computational and Applied Mathematics, 178(1– 2):205–214, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003759.

Ferreira:2006:GHF

[FLS06] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. The Gauss hypergeometric function F (a, b; c; z) for large c. Journal of Computational and Applied Mathematics, 197(2):568–577, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007004.

Ferreira:2008:ARB

[FLS08] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. Asymptotic relations between the Hahn-type polynomials and Meixner– Pollaczek, Jacobi, Meixner and Krawtchouk polynomials. Journal of Computational and Applied Mathematics, 217(1):88–109, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003172.

Ling:2008:MVM

[fLxXT08] Ai fan Ling, Cheng xian Xu, and Le Tang. A modified VNS metaheuristic for maxbisection problems. Journal of Computational and Applied Mathematics, 220(1–2): 413–421, October 15, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004621.

Ford:2009:NIC

[FN09] Neville J. Ford and Stewart J. Norton. Noise-induced changes to the behaviour of semi-implicit Euler methods for stochastic delay differential equations undergoing bifurcation. Journal of Computational and Applied Mathematics, 229(2):462–470, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001660.

Flauraud:2006:OIC

[FNW06] E. Flauraud, F. Nataf, and F. Willien. Optimized interface conditions in domain decomposition methods for problems with extreme contrasts in the coefficients. Journal of Computational and Applied Mathematics, 189(1–2):539–554, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003675.

Fischer:2005:MPR

[FO05] M. Fischer and P. Oja. Monotonicity preserving rational spline histopolation. Journal of Computational and Applied Mathematics, 175(2): 195–208, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002286.

Franco:2005:FOP

[FOP05] Daniel Franco, Donal O’Regan, and Juan Perán. Fourthorder problems with nonlinear boundary conditions. Journal of Computational and Applied Mathematics, 174(2):315–327, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002183.

Fischer:2007:CSP

[FOT07] Malle Fischer, Peeter Oja, and Helle Trossmann. Comonotone shape-preserving spline histopolation. Journal of Computational and Applied Mathematics, 200(1):127– 139, March 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008009.

Fornasier:2008:AIT

[FP08] Massimo Fornasier and Francesca Pitolli. Adaptive iterative thresholding algorithms for magnetoencephalography (MEG). Journal of Computational and Applied Mathematics, 221(2):386– 395, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005705.

Ferreira:2005:CSS

[FPO05] J. A. Ferreira, F. Patrício, and F. Oliveira. On the computation of solutions of systems of interval polynomial equations. Journal of Computational and Applied Mathematics, 173(2): 295–302, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400158X.

Fernandez:2005:WCO

[FPP05a] Lidia Fernández, Teresa E.

Pérez, and Miguel A. Piñar. Weak classical orthogonal polynomials in two variables. Journal of Computational and Applied Mathematics, 178(1–2):191–203, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003863.

Fulton:2005:CSF

[FPP05b] Charles Fulton, David Pearson, and Steven Pruess. Computing the spectral function for singular Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 176(1):131–162, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003243.

Fernandez:2007:SOP

[FPP07] Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. Second order partial differential equations for gradients of orthogonal polynomials in two variables. Journal of Computational and Applied Mathematics, 199(1): 113–121, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007430.

Fulton:2008:ECS

[FPP08a] Charles Fulton, David Pearson, and Steven Pruess. Efficient calculation of spectral density functions for specific classes of singular Sturm– Liouville problems. Journal of Computational and Applied Mathematics, 212 (2):150–178, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007205.

Fulton:2008:NCS

[FPP08b] Charles Fulton, David Pearson, and Steven Pruess. New characterizations of spectral density functions for singular Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 212(2):194–213, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007230.

Fancello:2008:VFN

[FPS08] E. A. Fancello, J. P. Ponthot, and L. Stainier. A variational framework for nonlinear viscoelastic models in finite deformation regime. Journal of Computational and Applied Mathematics, 215 (2):400–408, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007503.

Floater:2006:PBM

[FR06] Michael S. Floater and Atgeirr F. Rasmussen. Pointbased methods for estimating the length of a parametric curve. Journal of Computational and Applied Mathematics, 196(2):512– 522, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006023.

Fischer:2007:DSC

[FR07] Markus Fischer and Markus Reiß. Discretisation of stochastic control problems for continuous time dynamics with delay. Journal of Computational and Applied Mathematics, 205(2): 969–981, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004171.

Franco:2005:SEA

[Fra05] J. M. Franco. Stability of explicit ARKN methods for perturbed oscillators. Journal of Computational and Applied Mathematics, 173(2): 389–396, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002377.

Franco:2006:CET

[Fra06] J. M. Franco. A class of explicit two-step hybrid methods for second-order IVPs. Journal of Computational and Applied Mathematics, 187(1):41–57, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001445.

Figotin:2007:HTT

[FS07] Alexander Figotin and Jeffrey Schenker. Hamiltonian treatment of time dispersive and dissipative media within the linear response theory. Journal of Computational and Applied Mathematics, 204(2):199–208, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003505.

Ferreira:2008:ZMF

[FS08] Erasmo M. Ferreira and Javier Sesma. Zeros of the Macdonald function of complex order. Journal of Computational and Applied Mathematics, 211(2): 223–231, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007084.

Fan:2009:MSS

[FSL09] Zhencheng Fan, Minghui Song, and Mingzhu Liu. The α-th moment stability for the stochastic pantograph equation. Journal of Computational and Applied Mathematics, 233(2): 109–120, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002751.

Fang:2009:RTF

[FSW09] Yonglei Fang, Yongzhong Song, and Xinyuan Wu. A robust trigonometrically fitted embedded pair for perturbed oscillators. Journal of Computational and Applied Mathematics, 225(2): 347–355, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003816.

Frankel:2005:IEF

[FT05] J. I. Frankel and K. Taira. Integral equation formulation and error estimates for radial flow between two flat disks. Journal of Computational and Applied Mathematics, 181(1):103–124, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005564.

Famelis:2008:SDO

[FT08] I. Th. Famelis and Ch. Tsitouras. Symbolic derivation of order conditions for hybrid Numerov-type methods solving y′′ = f(x, y). Journal of Computational and Applied Mathematics, 218(2): 543–555, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005171.

Francomano:2005:WLB

[FTT^{+}05] E. Francomano, A. Tortorici, E. Toscano, G. Ala, and F. Viola. Wavelet-like bases for thin-wire integral equations in electromagnetics. Journal of Computational and Applied Mathematics, 175(1):77–86, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002456.

Fu:2008:TSP

[Fu08] Xiaoling Fu. A two-stage prediction–correction method for solving monotone variational inequalities. Journal of Computational and Applied Mathematics, 214 (2):345–355, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001306.

Flajolet:2008:DZV

[FV08] Philippe Flajolet and Linas Vepstas. On differences of zeta values. Journal of Computational and Applied Mathematics, 220(1– 2):58–73, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004165.

Fiocco:2006:MLF

[FvZ06] Marta Fiocco and Willem R. van Zwet. Maximum likelihood for the fully observed contact process. Journal of Computational and Applied Mathematics, 186(1): 117–129, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001925.

Fang:2009:ODB

[FW09] Ying Fang and Rong Wu. Optimal dividends in the Brownian motion risk model with interest. Journal of Computational and Applied Mathematics, 229(1):145–151, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005359.

Fan:2009:APS

[FWY09] Qiyi Fan, Wentao Wang, and Xuejun Yi. Anti-periodic solutions for a class of nonlinear n-th-order differential equations with delays. Journal of Computational and Applied Mathematics, 230(2): 762–769, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000144.

Fan:2009:PSS

[FWZ09] Qiyi Fan, Wentao Wang, and Jinglei Zhou. Periodic solutions of some fourth-order nonlinear differential equations. Journal of Computational and Applied Mathematics, 233(2): 121–126, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003987.

Feng:2006:TOF

[FY06] Yuan Feng and Lixing Yang.

A two-objective fuzzy kcardinality assignment problem. Journal of Computational and Applied Mathematics, 197(1):233–244, December 1, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006564.

Ferchichi:2008:IFB

[FZ08] J. Ferchichi and J. P. Zolésio.

Identification of a free boundary in Norton–Hoff flows with thermal effects. Journal of Computational and Applied Mathematics, 220(1–2): 181–197, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004359.

Fang:2009:NRR

[FZ09] Weifu Fang and Suxing Zeng. Numerical recovery of Robin boundary from boundary measurements for the Laplace equation. Journal of Computational and Applied Mathematics, 224(2): 573–580, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002641.

Fornberg:2007:SAT

[FZL07] Bengt Fornberg, Julia Zuev, and Jongwoo Lee. Stability and accuracy of timeextrapolated ADI–FDTD methods for solving wave equations. Journal of Computational and Applied Mathematics, 200(1):178– 192, March 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008046.

Gavriliadis:2009:MIP

[GA09] P. N. Gavriliadis and G. A. Athanassoulis. Moment information for probability distributions, without solving the moment problem, II: Mainmass, tails and shape approximation. Journal of Computational and Applied Mathematics, 229(1):7–15, July 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800513X.

Galo:2008:COP

[GAC^{+}08] J. R. Galo, I. I. Albarreal, M. C. Calzada, J. L. Cruz, E. Fernández-Cara, and M. Marín. Convergence and optimization of the parallel method of simultaneous directions for the solution of elliptic problems. Journal of Computational and Applied Mathematics, 222(2): 458–476, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006231.

Galperin:2006:SMR

[Gal06] A. Galperin. Secant method with regularly continuous divided differences. Journal of Computational and Applied Mathematics, 193(2): 574–595, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004334.

Galaktionov:2009:TTS

[Gal09] V. A. Galaktionov. Three types of self-similar blowup for the fourth-order p-Laplacian equation with source. Journal of Computational and Applied Mathematics, 223(1):326–355, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000265.

Gan:2007:DMN

[Gan07]  Siqing Gan.  Dissipativ-

ity of θ-methods for nonlinear Volterra delay-integro-differential equations. Journal of Computational and Applied Mathematics, 206(2): 898–907, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005395.

Gantumur:2008:OAW

[Gan08] Tsogtgerel Gantumur. An optimal adaptive wavelet method for nonsymmetric and indefinite elliptic problems. Journal of Computational and Applied Mathematics, 211(1):90–102, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006984.

Gao:2005:ERS

[Gao05]  Xiao-Wei Gao.  Evalua-

tion of regular and singular domain integrals with boundary-only discretizationtheory and Fortran code. Journal of Computational and Applied Mathematics, 175(2):265–290, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002328.

Gao:2006:NET

[Gao06] Xiao-Wei Gao. Numerical evaluation of two-dimensional singular boundary integrals

— theory and Fortran code. Journal of Computational and Applied Mathematics, 188(1):44–64, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001597.

Garrappa:2007:UGM

[Gar07a] Roberto Garrappa. The use of geometric meshes in product integration Simpson’s rules. Journal of Computational and Applied Mathematics, 210(1–2):200–209, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006625.

Gorenflo:2007:CGL

[GAR07b] R. Gorenflo and E. A. Abdel-Rehim. Convergence of the Grünwald–Letnikov scheme for time-fractional diffusion. Journal of Computational and Applied Mathematics, 205(2):871–881, August 15, 2007. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004092.

Garrappa:2009:SEA

[Gar09]  Roberto Garrappa.  On

some explicit Adams multistep methods for fractional differential equations. Journal of Computational and Applied Mathematics, 229 (2):392–399, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800160X.

Gaspar:2009:MLM

[Gás09] C. Gáspár. Multi-level meshless methods based on direct multi-elliptic interpolation. Journal of Computational and Applied Mathematics, 226(2):259–267, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004056.

Gautschi:2005:CPO

[Gau05a] Walter Gautschi. Computing polynomials orthogonal with respect to densely oscillating and exponentially decaying weight functions and related integrals. Journal of Computational and Applied Mathematics, 184(2): 493–504, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000440.

Gautschi:2005:HLF

[Gau05b] Walter Gautschi. The Hardy– Littlewood function: an exercise in slowly convergent series. Journal of Computational and Applied Mathematics, 179(1–2):249–254, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004510.

Gautschi:2005:OPM

[Gau05c] Walter Gautschi. Orthogonal polynomials (in Matlab). Journal of Computational and Applied Mathematics, 178(1–2):215–234, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003760.

Gautschi:2008:EAC

[Gau08] Walter Gautschi. On Euler’s attempt to compute logarithms by interpolation: a commentary to his letter of February 16, 1734 to Daniel Bernoulli. Journal of Computational and Applied Mathematics, 219(2): 408–415, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600714X.

Gupta:2009:GBP

[GBR09] R. K. Gupta, A. K. Bhunia, and D. Roy. A GA based penalty function technique for solving constrained redundancy allocation problem of series system with interval valued reliability of components. Journal of Computational and Applied Mathematics, 232(2):275–284, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003549.

Gracia:2006:CFD

[GC06] J. L. Gracia and C. Clavero. A compact finite difference scheme for 2D reaction– diffusion singularly perturbed problems. Journal of Computational and Applied Mathematics, 192(1):152–167, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003171.

Geng:2009:NMB

[GC09] Fazhan Geng and Minggen Cui. New method based on the HPM and RKHSM for solving forced Duffing equations with integral boundary conditions. Journal of Computational and Applied Mathematics, 233(2):

165–172, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004051.

Gebizlioglu:2009:RMS

[GD09] Omer L. Gebizlioğlu and Jan Dhaene. Risk measures and solvency — special issue: Guest Editors’ foreword. Journal of Computational and Applied Mathematics, 233(1):1–2, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004324.

Gillman:2007:MHF

[GDA07] Adrianna Gillman, Rabia Djellouli, and Mohamed Amara. A mixed hybrid formulation based on oscillated finite element polynomials for solving Helmholtz problems. Journal of Computational and Applied Mathematics, 204(2):515–525, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003773.

Gafiychuk:2008:MMT

[GDM08] V. Gafiychuk, B. Datsko, and V. Meleshko. Mathematical modeling of time fractional reaction–diffusion systems. Journal of Computational and Applied Mathematics, 220(1–2):215–225, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004372.

Geiser:2008:IOS

[Gei08]  Jürgen Geiser.  Iterative

operator-splitting methods with higher-order time integration methods and applications for parabolic partial differential equations. Journal of Computational and Applied Mathematics, 217 (1):227–242, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003512.

Geiser:2009:OSM

[Gei09] Jürgen Geiser. Operatorsplitting methods in respect of eigenvalue problems for nonlinear equations and applications for Burgers equations. Journal of Computational and Applied Mathematics, 231(2): 815–827, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900301X.

Gemikonakli:2009:AST

[GEK09] Orhan Gemikonakli, Enver Ever, and Altan Kocyigit.

Approximate solution for two stage open networks with Markov-modulated queues minimizing the state space explosion problem. Journal of Computational and Applied Mathematics, 223(1): 519–533, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000502.

Gemignani:2005:UHB

[Gem05] Luca Gemignani. A unitary Hessenberg QR-based algorithm via semiseparable matrices. Journal of Computational and Applied Mathematics, 184(2):505– 517, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000452.

Guo:2008:NDS

[GF08] Yuzhen Guo and Enmin Feng.

Nonlinear dynamical systems of trajectory design for 3D horizontal well and their optimal controls. Journal of Computational and Applied Mathematics, 212(2): 179–186, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007217.

Gatteschi:2007:MGI

[GG07] Luigi Gatteschi and Carla Giordano. On a method for generating inequalities for the zeros of certain functions. Journal of Computational and Applied Mathematics, 207(2):186–191, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006030.

Govaerts:2008:SCC

[GG08] W. Govaerts and R. Khoshsiar Ghaziani. Stable cycles in a Cournot duopoly model of kopel. Journal of Computational and Applied Mathematics, 218(2): 247–258, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000246.

Galeone:2009:EMF

[GG09a] Luciano Galeone and Roberto Garrappa. Explicit methods for fractional differential equations and their stability properties. Journal of Computational and Applied Mathematics, 228(2): 548–560, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001350.

Gu:2009:NMP

[GG09b] Xiaojuan Gu and Li Gao.

A new method for parameter estimation of edgepreserving regularization in image restoration. Journal of Computational and Applied Mathematics, 225(2): 478–486, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004202.

Gulati:2008:SOS

[GGA08] T. R. Gulati, S. K. Gupta, and I. Ahmad. Secondorder symmetric duality with cone constraints. Journal of Computational and Applied Mathematics, 220(1–2): 347–354, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004554.

Galvao:2008:ALS

[GGD08] Árpád Galvão, Marc Gerritsma, and Bart De Maerschalck. hp-Adaptive least squares spectral element method for hyperbolic partial differential equations. Journal of Computational and Applied Mathematics, 215 (2):409–418, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007515.

Gamez:2009:HON

[GGG09] D. Gámez, A. I. Garralda Guillem, and M. Ruiz Galán. High-order nonlinear initialvalue problems countably determined. Journal of Computational and Applied Mathematics, 228(1):77–82, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004391.

Ghelardoni:2006:PAS

[GGM06] P. Ghelardoni, G. Gheri, and M. Marletta. A polynomial approach to the spectral corrections for Sturm– Liouville problems. Journal of Computational and Applied Mathematics, 185(2): 360–376, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001202.

Guidoboni:2009:OSN

[GGP09] Giovanna Guidoboni, Roland Glowinski, and Matteo Pasquali. Operator splitting for the numerical solution of free surface flow at low capillary numbers. Journal of Computational and Applied Mathematics, 232(1): 72–81, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005463.

Gasull:2007:ENA

[GGT07] A. Gasull, H. Giacomini, and J. Torregrosa. Explicit nonalgebraic limit cycles for polynomial systems. Journal of Computational and Applied Mathematics, 200(1): 448–457, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000276.

Genz:2006:EPE

[GH06] Michael Genz and Erich Haeusler. Empirical processes with estimated parameters under auxiliary information. Journal of Computational and Applied Mathematics, 186(1):191–216, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001962.

Ghoreishi:2008:IMB

[GH08] F. Ghoreishi and S. Mohammad Hosseini. Integration matrix based on arbitrary grids with a preconditioner for pseudospectral method. Journal of Computational and Applied Mathematics, 214(1):274–287, April 15, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001227.

Galantai:2009:HMR

[GH09] A. Galántai and C. J. Hegedűs. Hyman’s method revisited. Journal of Computational and Applied Mathematics, 226(2):246–258, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004044.

Gallopoulos:2009:P

[GHK^{+}09] Efstratios Gallopoulos, Apostolos Hadjidimos, Ilias S. Kotsireas, Dimitrios Noutsos, and Michael N. Vrahatis. Preface. Journal of Computational and Applied Mathematics, 227(1):1–4, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003233.

Garcia:2009:ORF

[GHMPV09] A. G. García, M. A. Hernández-Medina, and G. Pérez-Villalón. Oversampling and reconstruction functions with compact support. Journal of Computational and Applied Mathematics, 227 (2):245–253, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001088.

Grajewski:2006:NAN

[GHT06] Matthias Grajewski, Jaroslav Hron, and Stefan Turek. Numerical analysis for a new non-conforming linear finite element on quadrilaterals. Journal of Computational and Applied Mathematics, 193(1):38–50, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500378X.

Gong:2006:EPA

[GHZ06] Lisha Gong, Xiyan Hu, and Lei Zhang. The expansion problem of antisymmetric matrix under a linear constraint and the optimal approximation. Journal of Computational and Applied Mathematics, 197(1): 44–52, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006321.

Giannantoni:2006:MGP

[Gia06] Corrado Giannantoni. Mathematics for generative processes: Living and non-living systems. Journal of Computational and Applied Mathematics, 189(1–2):324–340, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001391.

Giladi:2007:ADB

[Gil07] Eldar Giladi. Asymptotically derived boundary elements for the Helmholtz equation in high frequencies. Journal of Computational and Applied Mathematics, 198(1):52–74, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007065.

Gine:2005:PFI

[Gin05] Jaume Giné. Polynomial first integrals via the Poincaré series. Journal of Computational and Applied Mathematics, 184(2):428–441, December 15, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000403.

Ginovart:2008:SEW

[Gin08]  Frédéric Ginovart.  Some

exact Wick type stochastic generalized Boussinesq equation solutions. Journal of Computational and Applied Mathematics, 220(1–2): 559–565, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004864.

Gao:2008:THW

[GJ08] Jing Gao and Yao-Lin Jiang.

Trigonometric Hermite wavelet approximation for the integral equations of second kind with weakly singular kernel. Journal of Computational and Applied Mathematics, 215(1):242–259, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002002.

Guo:2008:MPS

[GJL08] Yanping Guo, Yude Ji, and Xiujun Liu. Multiple positive solutions for some multipoint boundary value problems with p-Laplacian. Journal of Computational and Applied Mathematics, 216 (1):144–156, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002245.

Ganji:2007:AHV

[GJM07] D. D. Ganji, M. Jannatabadi, and E. Mohseni. Application of He’s variational iteration method to nonlinear Jaulent–Miodek equations and comparing it with ADM. Journal of Computational and Applied Mathematics, 207(1):35–45, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004584.

Gekhtman:2007:TFG

[GK07] Michael Gekhtman and Alex Kasman. Tau-functions, Grassmannians and rank one conditions. Journal of Computational and Applied Mathematics, 202(1):80–87, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001117.

Guven:2008:TIG

[GK08] Bilgehan Güven and Samual Kotz. Test of independence for generalized Farlie– Gumbel–Morgenstern distributions. Journal of Computational and Applied Mathematics, 212(1):102–111, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007187.

Geum:2009:CCP

[GK09] Young Hee Geum and Young Ik Kim. Cubic convergence of parameter-controlled Newtonsecant method for multiple zeros. Journal of Computational and Applied Mathematics, 233(4):931–937, December 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709005354.

Gauthier:2007:HCP

[GKM07] A. Gauthier, P. A. Knight, and S. McKee. The Hertz contact problem, coupled Volterra integral equations and a linear complementarity problem. Journal of Computational and Applied Mathematics, 206(1): 322–340, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004523.

Georgiev:2009:ATS

[GKMS09] K. Georgiev, N. Kosturski, S. Margenov, and J. Starý. On adaptive time stepping for large-scale parabolic problems: Computer simulation of heat and mass transfer in vacuum freeze-drying. Journal of Computational and Applied Mathematics, 226 (2):268–274, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004068.

Gorbunov:2006:PMO

[GL06] Vladimir K. Gorbunov and Igor V. Lutoshkin. The parameterization method in optimal control problems and differential-algebraic equations. Journal of Computational and Applied Mathematics, 185(2):377–390, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001214.

Gracia:2007:UCS

[GL07] J. L. Gracia and F. J. Lisbona. A uniformly convergent scheme for a system of reaction–diffusion equations. Journal of Computational and Applied Mathematics, 206(1):1–16, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600389X.

Goberna:2008:SIS

[GLdS08] Miguel A. Goberna, Mercedes Larriqueta, and Virginia N. Vera de Serio. Stability of the intersection of solution sets of semi-infinite systems. Journal of Computational and Applied Mathematics, 217(2):420–431, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000994.

Gelhard:2005:SFE

[GLOS05] Tobias Gelhard, Gert Lube, Maxim A. Olshanskii, and Jan-Hendrik Starcke. Stabilized finite element schemes with LBB-stable elements for incompressible flows. Journal of Computational and Applied Mathematics, 177 (2):243–267, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004236.

Ganesh:2007:EEH

[GLS07] M. Ganesh, S. Langdon, and I. H. Sloan. Efficient evaluation of highly oscillatory acoustic scattering surface integrals. Journal of Computational and Applied Mathematics, 204(2):363–374, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003645.

Gilewicz:2005:NNM

[GLV05] Jacek Gilewicz, Elie Leopold, and Galliano Valent. New Nevanlinna matrices for orthogonal polynomials related to cubic birth and death processes. Journal of Computational and Applied Mathematics, 178(1–2):235–245, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003917.

Germain:2006:PVQ

[GM06] Marc Germain and Alphonse Magnus. Prices versus quantities: Stock pollution control with repeated choice of the instrument. Journal of Computational and Applied Mathematics, 197(2): 437–445, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006916.

Goodman:2007:PIP

[GM07a] T. N. T. Goodman and D. S. Meek. Planar interpolation with a pair of rational spirals. Journal of Computational and Applied Mathematics, 201(1):112–127, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000811.

Graves-Morris:2007:BVA

[GM07b] P. R. Graves-Morris. BiCGStab, VPAStab and an adaptation to mildly nonlinear systems. Journal of Computational and Applied Mathematics, 201(1):284–299, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001063.

Guseinov:2007:UTE

[GM07c] I. I. Guseinov and B. A. Mamedov. Unified treatment for the evaluation of generalized complete and incomplete gamma functions. Journal of Computational and Applied Mathematics, 202 (2):435–439, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001506.

Garza:2009:STR

[GM09a] L. Garza and F. Marcellán.

Szegő transformations and rational spectral transformations for associated polynomials. Journal of Computational and Applied Mathematics, 233(3):730–738, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001198.

Gorenflo:2009:SRA

[GM09b] Rudolf Gorenflo and Francesco Mainardi. Some recent advances in theory and simulation of fractional diffusion processes. Journal of Computational and Applied Mathematics, 229(2):400–415, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001611.

Graves-Morris:2009:IML

[GM09c] P. R. Graves-Morris. VPAStab(J, L):

an iterative method with look-ahead for the solution of large sparse linear systems. Journal of Computational and Applied Mathematics, 230(2):674–689, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000077.

Graves-Morris:2006:PSP

[GMFB06] P. R. Graves-Morris, A. F. Fell, and M. Bensalem. Parameterisation of symmetrical peaks in capillary electrophoresis using [3/2]-type rational approximants. Journal of Computational and Applied Mathematics, 189(1– 2):220–227, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001366.

Gatica:2009:AMF

[GMM09] Gabriel N. Gatica, Antonio Márquez, and Salim Meddahi. An augmented mixed finite element method for 3D linear elasticity problems. Journal of Computational and Applied Mathematics, 231(2):526–540, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002349.

Gori:2008:E

[GMS08] Laura Gori, Carla Manni, and Elisabetta Santi. Editorial. Journal of Computational and Applied Mathematics, 221(2):273, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005596.

Gushev:2009:MPC

[GN09] Vesselin Gushev and Geno Nikolov. Modified product cubature formulae. Journal of Computational and Applied Mathematics, 224(2): 465–475, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002525.

Gudi:2009:EEN

[GNP09] Thirupathi Gudi, Neela Nataraj, and Amiya K. Pani. On L^{2}-error estimate for nonsymmetric interior penalty Galerkin approximation to linear elliptic problems with nonhomogeneous Dirichlet data. Journal of Computational and Applied Mathematics, 228(1):30–40, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004354.

Guessab:2006:MAC

[GNS06] Allal Guessab, Otheman Nouisser, and Gerhard Schmeisser.

Multivariate approximation by a combination of modified Taylor polynomials. Journal of Computational and Applied Mathematics, 196(1): 162–179, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005340.

Golinski:2005:FMS

[GO05] Tomasz Goliński and Anatol Odzijewicz. Factorization method for second order functional equations. Journal of Computational and Applied Mathematics, 176 (2):331–355, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003413.

Gurka:2007:SEB

[GO07] Petr Gurka and Bohumír Opic. Sharp embeddings of Besov-type spaces. Journal of Computational and Applied Mathematics, 208(1): 235–269, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006388.

Gonzalez:2006:TFI

[Gon06] Graciela Adriana González. Theoretical framework of an identification problem for an elliptic variational inequality with bilateral restrictions. Journal of Computational and Applied Mathematics, 197(1):245–252, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006576.

Gong:2007:NEE

[Gon07] Yafang Gong. New error estimates for Galerkin method to an airfoil equation. Journal of Computational and Applied Mathematics, 206(1): 278–287, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004481.

Gong:2009:GSS

[Gon09a] Yafang Gong. Galerkin solution of a singular integral equation with constant coefficients. Journal of Computational and Applied Mathematics, 230(2):393–399, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006341.

Gonzalez:2009:GRQ

[Gon09b] J. R. Illán González. Gaussian rational quadrature formulas for ill-scaled integrands. Journal of Computational and Applied Mathematics, 233(3):745–748, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001216.

Goovaerts:2008:P

[Goo08]  M. J. Goovaerts.  Pref-

ace. Journal of Computational and Applied Mathematics, 218(2):201, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002537.

Gotz:2005:SPS

[Göt05] Thomas Götz. Simulating particles in Stokes flow. Journal of Computational and Applied Mathematics, 175(2): 415–427, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003139.

Gonzalez:2006:SOM

[GOT06] C. González, A. Ostermann, and M. Thalhammer. A second-order Magnus-type integrator for nonautonomous parabolic problems. Journal of Computational and Applied Mathematics, 189(1– 2):142–156, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002955.

Gogatishvili:2007:RTS

[GP07] Amiran Gogatishvili and Lubos Pick. A reduction theorem for supremum operators. Journal of Computational and Applied Mathematics, 208(1):270–279, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600639X.

Gori:2008:MAO

[GP08] Laura Gori and Francesca Pitolli. Multiresolution analyses originated from nonstationary subdivision schemes. Journal of Computational and Applied Mathematics, 221(2):406–415, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005729.

Gasparo:2007:TSM

[GPP07] Maria Grazia Gasparo, Alessandra Papini, and Aldo Pasquali. A two-stage method for nonlinear inverse problems. Journal of Computational and Applied Mathematics, 198(2): 471–482, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007272.

Gonzalez-Pinto:2007:CEI

[GPPRB07] S. González-Pinto, S. Pérez-Rodríguez, and R. Rojas Bello. Corrigendum to “Efficient iterations for Gauss methods on second-order problems”: [J. Comput. Appl. Math. 189 (2006) 80– 97]. Journal of Computational and Applied Mathematics, 205(1):583, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003402.

See [GPPRRB06].

Gonzalez-Pinto:2006:EIG

[GPPRRB06] S. González-Pinto, S. Pérez-Rodríguez, and R. Rojas-Bello. Efficient iterations for Gauss methods on secondorder problems. Journal of Computational and Applied Mathematics, 189(1– 2):80–97, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003626.

See corrigendum [GPPRB07].

Gonzalez-Pinto:2005:SNT

[GPRB05] S. González-Pinto and R. Rojas-Bello. Speeding up Newtontype iterations for stiff problems. Journal of Computational and Applied Mathematics, 181(2):266–279, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042704005849.

Gori:2008:IRO

[GPS08] L. Gori, E. Pellegrino, and E. Santi. Integral refinable operators exact on polynomials. Journal of Computational and Applied Mathematics, 221(2):396–405, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005717.

Guido:2009:ISI

[GPS09] Rodrigo Capobianco Guido, José Carlos Pereira, and Jan Frans Willem Slaets. Introduction to the special issue on wavelet-based algorithms and emergent classification techniques in science and engineering. Journal of Computational and Applied Mathematics, 227(2):221–222, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001052.

Gilewicz:2005:PPA

[GPTT05] Jacek Gilewicz, Maciej Pindor, J. Joachim Telega, and Stanislaw Tokarzewski. Npoint Padé approximants and two-sided estimates of errors on the real axis for Stieltjes functions. Journal of Computational and Applied Mathematics, 178(1– 2):247–253, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003772.

Garcia:2005:AEW

[GPV05] A. G. García and G. Pérez-Villalón. On the aliasing error in wavelet subspaces. Journal of Computational and Applied Mathematics, 183(1): 153–167, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000221.

Gu:2009:SSM

[GQ09a] Chuanqing Gu and Hongjun Qian. Skew-symmetric methods for nonsymmetric linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 223(2): 567–577, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000022.

Guo:2009:CLC

[GQ09b] Senlin Guo and Feng Qi. A class of logarithmically completely monotonic functions associated with the gamma function. Journal of Computational and Applied Mathematics, 224(1): 127–132, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001829.

Guirao:2009:DSD

[GR09] Juan Luis García Guirao and Raquel García Rubio. Detecting simple dynamics in Cournot-like models. Journal of Computational and Applied Mathematics, 233 (4):1091–1095, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006220.

Graillat:2006:NSP

[Gra06] Stef Graillat. A note on structured pseudospectra. Journal of Computational and Applied Mathematics, 191 (1):68–76, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002621.

Griffiths:2009:SPO

[Gri09] Bob Griffiths. Stochastic processes with orthogonal polynomial eigenfunctions. Journal of Computational and Applied Mathematics, 233(3): 739–744, December 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001204.

Guo:2009:RLS

[GRL09] Hui Guo, Hongxing Rui, and Chao Lin. A remark on leastsquares Galerkin procedures for pseudohyperbolic equations. Journal of Computational and Applied Mathematics, 229(1):108–119, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005244.

Ghadi:2008:NST

[GRW08] Fattehallah Ghadi, Vitoriano Ruas, and Mohamed Wakrim. Numerical solution of the time-dependent incompressible Navier–Stokes equations by piecewise linear finite elements. Journal of Computational and Applied Mathematics, 215(2):429–437, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007539.

Gerber:2006:ODR

[GS06a] Hans U. Gerber and Elias S. W. Shiu. On optimal dividends: From reflection to refraction. Journal of Computational and Applied Mathematics, 186(1):4–22, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500186X.

Gulsu:2006:TPA

[GS06b] Mustafa Gülsu and Mehmet Sezer. A Taylor polynomial approach for solving differential-difference equations. Journal of Computational and Applied Mathematics, 186(2):349–364, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000993.

Guo:2006:ARS

[GS06c] Benqi Guo and Christoph Schwab. Analytic regularity of Stokes flow on polygonal domains in countably weighted Sobolev spaces. Journal of Computational and Applied Mathematics, 190(1–2):487–519, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002487.

Ganji:2007:AHP

[GS07a] D. D. Ganji and A. Sadighi.

Application of homotopyperturbation and variational iteration methods to nonlinear heat transfer and porous media equations. Journal of Computational and Applied Mathematics, 207(1): 24–34, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004572.

Gilsing:2007:SPS

[GS07b] Hagen Gilsing and Tony Shardlow. SDELab: a package for solving stochastic differential equations in MAT-LAB. Journal of Computational and Applied Mathematics, 205(2):1002–1018, August 15, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004195.

Gomez:2007:CFS

[GS07c] Francisco J. Gómez and Javier Sesma. Connection factors in the Schrödinger equation with a polynomial potential. Journal of Computational and Applied Mathematics, 207(2):291–300, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005954.

Gu:2008:FVP

[GS08] Chuanqing Gu and Jindong Shen. Function-valued Padétype approximant via the formal orthogonal polynomials and its applications in solving integral equations. Journal of Computational and Applied Mathematics, 221(1): 114–131, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005298.

Grau-Sanchez:2009:IOE

[GS09] Miquel Grau-Sánchez. Improving order and efficiency: Composition with a modified Newton’s method. Journal of Computational and Applied Mathematics, 231(2): 592–597, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002507.

Gopaul:2009:AIF

[GSBB09] A. Gopaul, M. S. Sunhaloo, R. Boojhawon, and M. Bhuruth. Analysis of incomplete factorizations for a nine-point approximation to a convection–diffusion model problem. Journal of Computational and Applied Mathematics, 224(2): 719–733, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002999.

Gil:2007:P

[GSL07] Amparo Gil, Javier Segura, and José Luis López. Preface. Journal of Computational and Applied Mathematics, 207(2):165, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005942.

Grote:2007:IPD

[GSS07] Marcus J. Grote, Anna Schneebeli, and Dominik Schötzau. Interior penalty discontinuous Galerkin method for Maxwell’s equations: Energy norm error estimates. Journal of Computational and Applied Mathematics, 204(2):375–386, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003657.

Gil:2006:AHR

[GST06] Amparo Gil, Javier Segura, and Nico M. Temme. The ABC of hyper recursions. Journal of Computational and Applied Mathematics, 190(1–2):270–286, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002359.

Gao:2007:PDR

[GSW07] Huijun Gao, Peng Shi, and Junling Wang. Parameterdependent robust stability of uncertain time-delay systems. Journal of Computational and Applied Mathematics, 206(1):366–373, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004742.

Gu:2009:HIM

[GT09] Chuanqing Gu and Zhaolu Tian. On the HSS iteration methods for positive definite Toeplitz linear systems. Journal of Computational and Applied Mathematics, 224(2):709–718, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002975.

See erratum [LHL11].

Gu:2007:ASP

[Gu07] Rongbao Gu. The averageshadowing property and topological ergodicity. Journal of Computational and Applied Mathematics, 206(2): 796–800, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600522X.

Georgieva:2008:SRP

[GU08] Irina Georgieva and Rumen Uluchev. Smoothing of Radon projections type of data by bivariate polynomials. Journal of Computational and Applied Mathematics, 215 (1):167–181, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001938.

Guglielmi:2006:OID

[Gug06] Nicola Guglielmi. Open issues in devising software for the numerical solution of implicit delay differential equations. Journal of Computational and Applied Mathematics, 185(2):261–277, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001147.

Guo:2006:EMS

[Guo06] Chun-Hua Guo. Efficient methods for solving a nonsymmetric algebraic Riccati equation arising in stochastic fluid models. Journal of Computational and Applied Mathematics, 192(2): 353–373, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003560.

Garcia:2008:SAU

[GVC^{+}08] Oscar García, Josep Vehí, Jose Campos e Matos, António Abel Henriques, and Joan Ramon Casas. Structural assessment under uncertain parameters via interval analysis. Journal of Computational and Applied Mathematics, 218(1):43–52, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003251.

Goovaerts:2006:P

[GVW06] M. J. Goovaerts, S. Vandewalle, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 189(1–2):1, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000033.

Green:2006:PDD

[GW06a] Kirk Green and Thomas Wagenknecht. Pseudospectra and delay differential equations. Journal of Computational and Applied Mathematics, 196(2):567– 578, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006217.

Guo:2006:MLP

[GW06b] Ben-Yu Guo and Li-Lian Wang. Modified Laguerre pseudospectral method refined by multidomain Legendre pseudospectral approximation. Journal of Computational and Applied Mathematics, 190(1–2):304–324, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002372.

Gao:2008:SSP

[GW08] Wenjie Gao and Junyu Wang. Similarity solutions to the power-law generalized Newtonian fluid. Journal of Computational and Applied Mathematics, 222(2):381– 391, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006061.

Gerhold:2009:FER

[GW09] Stefan Gerhold and Richard Warnung. Finding efficient recursions for risk aggregation by computer algebra. Journal of Computational and Applied Mathematics, 223(1): 499–507, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000393.

Georges:2008:ISF

[GWG08] Laurent Georges, Grégoire Winckelmans, and Philippe Geuzaine. Improving shockfree compressible RANS solvers for LES on unstructured meshes. Journal of Computational and Applied Mathematics, 215 (2):419–428, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007527.

Gui:2009:OEO

[GX09] Gaihua Gui and Zhiting Xu.

Oscillation of even order partial differential equations with distributed deviating arguments. Journal of Computational and Applied Mathematics, 228(1):20–29, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004342.

Ge:2005:MBA

[GXgL05] Ren-Dong Ge, Zun-Quan Xia, and Jian guo Liu. A modified Brown algorithm for solving singular nonlinear systems with rank defects. Journal of Computational and Applied Mathematics, 181(2): 252–265, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005837.

Guohui:2006:DFU

[GXZ06] Zhao Guohui, Liu Xiuping, and Su Zhixun. A dual functional to the univariate B-spline. Journal of Computational and Applied Mathematics, 195(1–2): 292–299, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004930.

Gao:2008:PSA

[GY08] Heli Gao and Chuancun Yin. The perturbed Sparre Andersen model with a threshold dividend strategy. Journal of Computational and Applied Mathematics, 220(1–2): 394–408, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004591.

Guo:2005:NTI

[GZ05] Victor J. W. Guo and Jiang Zeng. A note on two identities arising from enumeration of convex polyominoes. Journal of Computational and Applied Mathematics, 180(2): 413–423, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005412.

Gao:2007:SFD

[GZL07] Liping Gao, Bo Zhang, and Dong Liang. The splitting finite-difference time-domain methods for Maxwell’s equations in two dimensions. Journal of Computational and Applied Mathematics, 205(1): 207–230, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002834.

Gu:2009:IPH

[GZLL09] Tong-Xiang Gu, Xian-Yu Zuo, Xing-Ping Liu, and Pei-Lu Li. An improved parallel hybrid bi-conjugate gradient method suitable for distributed parallel computing. Journal of Computational and Applied Mathematics, 226(1):55–65, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002215.

Gu:2009:LSN

[GZX09] Jian Gu, Liwei Zhang, and Xiantao Xiao. Log-sigmoid nonlinear Lagrange method for nonlinear optimization problems over second-order cones. Journal of Computational and Applied Mathematics, 229(1):129–144, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005347.

Hamelinck:2009:INI

[Ham09] Wouter Hamelinck. On the influence of numerical integration on mixed finite element approximations of a Maxwell eigenvalue problem. Journal of Computational and Applied Mathematics, 223(2):929–937, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001143.

Han:2005:QTP

[Han05] Xuli Han. C^{2} quadratic trigonometric polynomial curves with local bias. Journal of Computational and Applied Mathematics, 180(1): 161–172, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005035.

Han:2006:PQP

[Han06] Xuli Han. Piecewise quartic polynomial curves with a local shape parameter. Journal of Computational and Applied Mathematics, 195(1– 2):34–45, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500470X.

Hansen:2007:GRK

[Han07]  Eskil Hansen.  Galerkin/

Runge–Kutta discretizations of nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 205(2):882–890, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004237.

Han:2008:GPP

[Han08]  Deren Han.  A gener-

alized proximal-point-based prediction–correction method for variational inequality problems. Journal of Computational and Applied Mathematics, 221(1):183– 193, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700578X.

Han:2009:DDR

[Han09] Xuli Han. A degree by degree recursive construction of Hermite spline interpolants. Journal of Computational and Applied Mathematics, 225(1):113–123, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003567.

Hartung:2005:LSP

[Har05] Ferenc Hartung. Linearized stability in periodic functional differential equations with state-dependent delays. Journal of Computational and Applied Mathematics, 174(2):201–211, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001876.

Harrell:2006:GLB

[Har06] Evans M. Harrell. Geometric lower bounds for the spectrum of elliptic PDEs with Dirichlet conditions in part. Journal of Computational and Applied Mathematics, 194(1):26–35, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004012.

Harfash:2008:HAF

[Har08a] Akil J. Harfash. High accuracy finite difference scheme for three-dimensional microscale heat equation. Journal of Computational and Applied Mathematics, 220(1–2): 335–346, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004542.

Harris:2008:IBG

[Har08b] Frank E. Harris. Incomplete Bessel, generalized incomplete gamma, or leaky aquifer functions. Journal of Computational and Applied Mathematics, 215(1):260–269, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002014.

Hashim:2006:ADM

[Has06] Ishak Hashim. Adomian decomposition method for solving BVPs for fourth-order integro-differential equations. Journal of Computational and Applied Mathematics, 193(2):658–664, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004383.

Hayat:2008:ASM

[HAS08] T. Hayat, Naveed Ahmed, and M. Sajid. Analytic solution for MHD flow of a third order fluid in a porous channel. Journal of Computational and Applied Mathematics, 214(2):572–582, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001550.

Hascelik:2009:NCI

[Has09a] A. Ihsan Hascelik. On numerical computation of integrals with integrands of the form f(x) sin(w/x^{r}) on [0, 1]. Journal of Computational and Applied Mathematics, 223(1):399–408, January 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000319.

Hayek:2009:NRI

[HAS09b] Mohamed Hayek, Philippe Ackerer, and Éric Sonnendrücker. A new refinement indicator for adaptive parameterization: Application to the estimation of the diffusion coefficient in an elliptic problem. Journal of Computational and Applied Mathematics, 224(1):307–319, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002100.

Hataue:2009:DDS

[Hat09] Itaru Hataue. On dependence of dynamical structure of numerical solutions of fluid simulations on forcibly added randomness. Journal of Computational and Applied Mathematics, 232 (1):82–89, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005475.

Hegland:2007:SSM

[HBS^{+}07] Markus Hegland, Conrad Burden, Lucia Santoso, Shev MacNamara, and Hilary Booth. A solver for the stochastic master equation applied to gene regulatory networks. Journal of Computational and Applied Mathematics, 205(2): 708–724, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003955.

Hong:2008:MPR

[HC08] Shihuang Hong and Ji Chen.

Multiplicity of positive radial solutions for an elliptic inclusion system on an annulus. Journal of Computational and Applied Mathematics, 221(1):66–75, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005250.

Harris:2009:EME

[HC09] Paul J. Harris and Ke Chen.

An efficient method for evaluating the integral of a class of highly oscillatory functions. Journal of Computational and Applied Mathematics, 230(2):433–442, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006407.

Henneron:2008:DFE

[HCDP08] Thomas Henneron, Stéphane Clénet, Patrick Dular, and Francis Piriou. Discrete finite element characterizations of source fields for volume and boundary constraints in electromagnetic problems. Journal of Computational and Applied Mathematics, 215 (2):438–447, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007540.

Hobbs:2007:TNE

[HCK07] C. A. Hobbs, J. N. L. Connor, and N. P. Kirk. Theory and numerical evaluation of oddoids and evenoids: Oscillatory cuspoid integrals with odd and even polynomial phase functions. Journal of Computational and Applied Mathematics, 207(2): 192–213, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006042.

Ham:2008:SHO

[HCL08a] YoonMee Ham, Changbum Chun, and Sang-Gu Lee. Some higher-order modifications of Newton’s method for solving nonlinear equations. Journal of Computational and Applied Mathematics, 222(2):477–486, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006243.

He:2008:CAC

[HCL08b] Changhong He, Thomas F.

Coleman, and Yuying Li. Computation and analysis for a constrained entropy optimization problem in finance. Journal of Computational and Applied Mathematics, 222(1):159–174, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005559.

Han:2005:EPS

[HCS05] Maoan Han, Hua Chen, and Desheng Shang. Existence of periodic solutions of a neuronic equation. Journal of Computational and Applied Mathematics, 181(2): 235–244, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005618.

He:2005:DPS

[He05] Zhimin He. Double positive solutions of three-point boundary value problems for p-Laplacian dynamic equations on time scales. Journal of Computational and Applied Mathematics, 182(2): 304–315, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400620X.

He:2006:NDP

[He06a] Yiran He. A new double projection algorithm for variational inequalities. Journal of Computational and Applied Mathematics, 185(1): 166–173, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000555.

Hechenleitner:2006:PSG

[HE06b] Bernhard Hechenleitner and Karl Entacher. A parallel search for good lattice points using LLL-spectral tests. Journal of Computational and Applied Mathematics, 189(1–2):424–441, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001652.

He:2007:VIM

[He07] Ji-Huan He. Variational iteration method — some recent results and new interpretations. Journal of Computational and Applied Mathematics, 207(1):3–17, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/

science/article/pii/S0377042706004559.

See remark [Pet09a].

Heinkenschloss:2005:TDD

[Hei05] Matthias Heinkenschloss. A time-domain decomposition iterative method for the solution of distributed linear quadratic optimal control problems. Journal of Computational and Applied Mathematics, 173(1): 169–198, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001505.

Helal:2005:GMA

[HeM05] M. M. Helal and M. B. Abd el Malek. Group method analysis of magneto-elastico-viscous flow along a semiinfinite flat plate with heat transfer. Journal of Computational and Applied Mathematics, 173(2):199–210, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001529.

Hempel:2006:LEL

[Hem06] Rainer Hempel. On the lowest eigenvalue of the Laplacian with Neumann boundary condition at a small obstacle. Journal of Computational and Applied Mathematics, 194(1):54–74, September 15, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004036.

Hamdi:2005:MLS

[HEOS05] S. Hamdi, W. H. Enright, Y. Ouellet, and W. E. Schiesser. Method of lines solutions of the extended Boussinesq equations. Journal of Computational and Applied Mathematics, 183(2): 327–342, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000695.

Hernandez:2009:FEA

[Her09] Erwin Hernández. Finite element approximation of the elasticity spectral problem on curved domains. Journal of Computational and Applied Mathematics, 225(2): 452–458, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004172.

He:2009:MCR

[HG09] Xubiao He and Pu Gong. Measuring the coupled risks: a copula-based CVaR model. Journal of Computational and Applied Mathematics, 223(2):1066–1080, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001465.

Hua:2006:DRM

[HGS06] Changchun Hua, Xinping Guan, and Peng Shi. Decentralized robust model reference adaptive control for interconnected timedelay systems. Journal of Computational and Applied Mathematics, 193(2): 383–396, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004206.

Huang:2008:VAK

[HGS08] Jianguo Huang, Ling Guo, and Zhongci Shi. Vibration analysis of Kirchhoff plates by the Morley element method. Journal of Computational and Applied Mathematics, 213(1):14–34, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000039.

He:2007:FVM

[HH07a] Guoliang He and Yinnian He. The finite volume method based on stabilized finite element for the stationary Navier–Stokes problem. Journal of Computational and Applied Mathematics, 205(1):651–665, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004535.

Heinkenschloss:2007:SDD

[HH07b] Matthias Heinkenschloss and Michael Herty. A spatial domain decomposition method for parabolic optimal control problems. Journal of Computational and Applied Mathematics, 201(1):88–111, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600080X.

Hesthaven:2007:WC

[HH07c] Jan S. Hesthaven and Houssem Haddar. Waves 2005 Conference. Journal of Computational and Applied Mathematics, 204(2):197–198, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003499.

Hu:2009:PFG

[HHC09] Sheng-Long Hu, Zheng-Hai Huang, and Jein-Shan Chen. Properties of a family of generalized NCP-functions and a derivative free algorithm for complementarity problems. Journal of Computational and Applied Mathematics, 230(1):69–82, August 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005876.

Hounga:2006:NFO

[HHR06] C. Hounga, M. N. Hounkonnou, and A. Ronveaux. New families of orthogonal polynomials. Journal of Computational and Applied Mathematics, 193(2): 474–483, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004267.

He:2005:SOA

[HHST05] T. X. He, L. C. Hsu, P. J.- S. Shiue, and D. C. Torney. A symbolic operator approach to several summation formulas for power series. Journal of Computational and Applied Mathematics, 177(1):17–33, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003607.

Huang:2007:PPR

[HHZ07] Xiantong Huang, Xiyan Hu, and Lei Zhang. Physical parameters reconstruction of a fixed-fixed massspring system from its characteristic data. Journal of Computational and Applied Mathematics, 206(2):

645–655, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005115.

Han:2009:EPS

[HJK09] Wei Han, Zhen Jin, and Shugui Kang. Existence of positive solutions of nonlinear m-point BVP for an increasing homeomorphism and positive homomorphism on time scales. Journal of Computational and Applied Mathematics, 233(2): 188–196, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004075.

Hansen:2007:APA

[HJR07] Per Christian Hansen, Toke Koldborg Jensen, and Giuseppe Rodriguez. An adaptive pruning algorithm for the discrete L-curve criterion. Journal of Computational and Applied Mathematics, 198(2): 483–492, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007284.

Ha:2007:AOD

[HK07a] Taeyoung Ha and Imbunm Kim. Analysis of onedimensional Helmholtz equation with PML boundary. Journal of Computational and Applied Mathematics, 206(1):586–598, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006170.

Higham:2007:SCR

[HK07b] Desmond J. Higham and Peter E. Kloeden. Strong convergence rates for backward Euler on a class of nonlinear jump-diffusion problems. Journal of Computational and Applied Mathematics, 205(2):949–956, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004158.

See comments [TGJ19].

Heineken:2008:AST

[HK08] Wolfram Heineken and Matthias Kunik. The analytical solution of two interesting hyperbolic problems as a test case for a finite volume method with a new grid refinement technique. Journal of Computational and Applied Mathematics, 214(2):509–532, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001525.

Hayat:2007:ESC

[HKA07] T. Hayat, Masood Khan, and M. Ayub. The effect of the slip condition on flows of an Oldroyd 6-constant fluid. Journal of Computational and Applied Mathematics, 202(2):402–413, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001488.

Hafele:2005:DST

[HKB^{+}05] M. Häfele, A. Kienle, M. Boll, C.-U. Schmidt, and M. Schwibach. Dynamic simulation of a tubular reactor for the production of low-density polyethylene using adaptive method of lines. Journal of Computational and Applied Mathematics, 183(2): 288–300, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500066X.

Higham:2007:SCU

[HKK07] Desmond J. Higham, Gabriela Kalna, and Milla Kibble. Spectral clustering and its use in bioinformatics. Journal of Computational and Applied Mathematics, 204 (1):25–37, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002366.

Huttunen:2008:UWM

[HKM08] T. Huttunen, J. P. Kaipio, and P. Monk. An ultraweak method for acoustic fluid-solid interaction. Journal of Computational and Applied Mathematics, 213(1): 166–185, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000131.

Herty:2007:OTN

[HKS07] M. Herty, A. Klar, and A. K. Singh. An ODE traffic network model. Journal of Computational and Applied Mathematics, 203 (2):419–436, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002159.

Hannukainen:2009:DMP

[HKV09] Antti Hannukainen, Sergey Korotov, and Tomás Vejchodský. Discrete maximum principle for FE solutions of the diffusion-reaction problem on prismatic meshes. Journal of Computational and Applied Mathematics, 226(2):275–287, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800407X.

Han:2009:GIR

[HKZ09] Bin Han, Soon-Geol Kwon, and Xiaosheng Zhuang. Generalized interpolating refinable function vectors. Journal of Computational and Applied Mathematics, 227 (2):254–270, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800109X.

Hadjidimos:2007:SBA

[HL07a] A. Hadjidimos and M. Lapidakis. Stationary biparametric ADI preconditioners for conjugate gradient methods. Journal of Computational and Applied Mathematics, 205(1):364–381, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003244.

Halpern:2007:DNM

[HL07b] Laurence Halpern and Olivier Lafitte. Dirichlet to Neumann map for domains with corners and approximate boundary conditions. Journal of Computational and Applied Mathematics, 204(2):505–514, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003761.

Hu:2007:ISS

[HL07c] Guang-Da Hu and Mingzhu Liu. Input-to-state stability of Runge–Kutta methods for nonlinear control systems. Journal of Computational and Applied Mathematics, 205(1):633–639, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003451.

Hantoute:2008:CTI

[HL08] A. Hantoute and M. A.

López. Characterization of total ill-posedness in linear semi-infinite optimization. Journal of Computational and Applied Mathematics, 217(2):350–364, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000957.

He:2009:CFP

[HL09a] Rong-Hua He and Hong-Xu Li. Collective fixed point theorem and coincidence theorems in FC-spaces. Journal of Computational and Applied Mathematics, 225 (1):227–235, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003658.

Hou:2009:SSI

[HL09b] Yanren Hou and Qingfang Liu. A stabilized semiimplicit Galerkin scheme for Navier–Stokes equations. Journal of Computational and Applied Mathematics, 231(2):552–560, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002465.

Huang:2009:PSS

[HL09c] Xudong Huang and Wei Liu. Properties of some statistics for AR–ARCH model with application to technical analysis. Journal of Computational and Applied Mathematics, 225(2):522–530, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800424X.

Hassan:2009:NGM

[HLF09] Malik Abu Hassan, Wah June Leong, and Mahboubeh Farid. A new gradient method via quasi-Cauchy relation which guarantees descent. Journal of Computational and Applied Mathematics, 230(1):300–305, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006250.

He:2007:IMS

[HLH07] Xin-Feng He, Jian Lou, and Zhen He. Iterative methods for solving variational inclusions in Banach spaces. Journal of Computational and Applied Mathematics, 203(1):80–86, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001749.

Hu:2009:OCR

[HLJ09] Bei Hu, Jin Liang, and Lishang Jiang. Optimal convergence rate of the explicit finite difference scheme for American option valuation. Journal of Computational and Applied Mathematics, 230(2): 583–599, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006535.

Hu:2008:ASL

[HLL08a] Yizheng Hu, Yong Luo, and Zhengyi Lu. Analytical solution of the linear fractional differential equation by Adomian decomposition method. Journal of Computational and Applied Mathematics, 215(1):220–229, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001987.

Huang:2008:NEN

[HLL08b] Hung-Tsai Huang, Zi-Cai Li, and Qun Lin. New expansions of numerical eigenvalues by finite elements. Journal of Computational and Applied Mathematics, 217(1):9–27, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003111.

Hao:2006:ERT

[HLX06] Zhao-Cai Hao, Jin Liang, and Ti-Jun Xiao. Existence results for time scale boundary value problem. Journal of Computational and Applied Mathematics, 197(1): 156–168, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006503.

Huang:2009:SEP

[HLY09] Hongying Huang, Dongjie Liu, and Dehao Yu. Solution of exterior problem using ellipsoidal artificial boundary. Journal of Computational and Applied Mathematics, 231(1):434–446, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002143.

Long:2008:INM

[hLyHZ08] Jian hui Long, Xi yan Hu, and Lei Zhang. Improved Newton’s method with exact line searches to solve quadratic matrix equation. Journal of Computational and Applied Mathematics, 222(2): 645–654, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006462.

Hou:2008:KML

[HM08] Qing-Hu Hou and Toufik Mansour. Kernel method and linear recurrence system. Journal of Computational and Applied Mathematics, 216(1):227–242, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002324.

Haramoto:2009:AAC

[HM09a] H. Haramoto and M. Matsumoto. A p-adic algorithm for computing the inverse of integer matrices. Journal of Computational and Applied Mathematics, 225 (1):320–322, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003907.

Huang:2009:OCF

[HM09b] Yan Huang and Fanwei Meng. Oscillation criteria for forced second-order nonlinear differential equations with damping. Journal of Computational and Applied Mathematics, 224(1):339–345, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002124.

Hofmann:2006:MMS

[HMG06] Norbert Hofmann and Thomas Müller-Gronbach. A modified Milstein scheme for approximation of stochastic delay differential equations with constant time lag. Journal of Computational and Applied Mathematics, 197(1): 89–121, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006461.

Han:2008:NGB

[HMH08] Xi-An Han, YiChen Ma, and XiLi Huang. A novel generalization of Bézier curve and surface. Journal of Computational and Applied Mathematics, 217(1):180–193, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003482.

Heikkola:2007:CMA

[HMPR07] Erkki Heikkola, Sanna Mönkölä, Anssi Pennanen, and Tuomo Rossi. Controllability method for acoustic scattering with spectral elements. Journal of Computational and Applied Mathematics, 204 (2):344–355, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003621.

Hueso:2009:MNM

[HMT09] José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Modified Newton’s method for systems of nonlinear equations with singular Jacobian. Journal of Computational and Applied Mathematics, 224(1): 77–83, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001787.

Hechme:2009:MRC

[HNS09] G. Hechme, Yu. M. Nechepurenko, and M. Sadkane. Model reduction for a class of linear descriptor systems. Journal of Computational and Applied Mathematics, 229(1):54–60, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005189.

Hirota:2006:NME

[HO06] Chiaki Hirota and Kazufumi Ozawa. Numerical method of estimating the blow-up time and rate of the solution of ordinary differential equations

— an application to the blowup problems of partial differential equations. Journal of Computational and Applied Mathematics, 193(2): 614–637, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004358.

Hiraoka:2007:EEB

[HO07] Yasuaki Hiraoka and Toshiyuki Ogawa. An efficient estimate based on FFT in topological verification method. Journal of Computational and Applied Mathematics, 199(2): 238–244, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007569.

Huang:2009:AIM

[HO09] Xinzheng Huang and Cornelis W. Oosterlee. Adaptive integration for multi-factor portfolio credit loss models. Journal of Computational and Applied Mathematics, 231(2):506–516, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002325.

Hochstenbach:2008:JDT

[Hoc08] Michiel E. Hochstenbach.

A Jacobi–Davidson type method for the product eigenvalue problem. Journal of Computational and Applied Mathematics, 212(1): 46–62, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007114.

Homeier:2005:NTM

[Hom05] H. H. H. Homeier. On Newton-type methods with cubic convergence. Journal of Computational and Applied Mathematics, 176 (2):425–432, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003462.

Homeier:2009:NTM

[Hom09] H. H. H. Homeier. On Newton-type methods for multiple roots with cubic convergence. Journal of Computational and Applied Mathematics, 231(1): 249–254, September 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000582.

Hong:2007:TPS

[Hon07] Shihuang Hong. Triple positive solutions of three-point boundary value problems for p-Laplacian dynamic equations on time scales. Journal of Computational and Applied Mathematics, 206(2): 967–976, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005619.

Hong:2008:MPS

[Hon08] Shihuang Hong. Multiple positive solutions for a class of integral inclusions. Journal of Computational and Applied Mathematics, 214 (1):19–29, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000787.

Hooman:2008:PSF

[Hoo08] K. Hooman. A perturbation solution for forced convection in a porous-saturated duct. Journal of Computational and Applied Mathematics, 211(1):57–66, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006959.

Hosseini:2006:ADM

[Hos06]  M. M. Hosseini.  Ado-

mian decomposition method for solution of differentialalgebraic equations. Journal of Computational and Applied Mathematics, 197(2): 495–501, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006953.

Hosseini:2009:AOT

[Hos09] S. Mohammad Hosseini. The adaptive operational Tau method for systems of ODEs. Journal of Computational and Applied Mathematics, 231(1):24–38, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000375.

Howell:2009:CVF

[How09a] Jason S. Howell. Computation of viscoelastic fluid flows using continuation methods. Journal of Computational and Applied Mathematics, 225(1):187–201, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003622.

Howell:2009:DMF

[How09b] Jason S. Howell. Dual-mixed finite element approximation of Stokes and nonlinear Stokes problems using tracefree velocity gradients. Journal of Computational and Applied Mathematics, 231(2): 780–792, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900291X.

Hundsdorfer:2007:NIS

[HP07] Willem Hundsdorfer and Laura Portero. A note on iterated splitting schemes. Journal of Computational and Applied Mathematics, 201 (1):146–152, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000938.

Ham:2006:STA

[HPS06] David A. Ham, Julie Pietrzak, and Guus S. Stelling. A streamline tracking algorithm for semi-Lagrangian advection schemes based on the analytic integration of the velocity field. Journal of Computational and Applied Mathematics, 192(1):168–174, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003183.

Hongmin:2006:CBS

[HQ06] Ren Hongmin and Wu Qingbiao. The convergence ball of the secant method under Hölder continuous divided differences. Journal of Computational and Applied Mathematics, 194(2): 284–293, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004656.

Hernandez:2005:ACN

[HR05a] M. A. Hernández and N. Romero.

Accelerated convergence in Newton’s method for approximating square roots. Journal of Computational and Applied Mathematics, 177 (1):225–229, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004315.

Hernandez:2005:CSN

[HR05b] M. A. Hernández and N. Romero.

On a characterization of some Newton-like methods of Rorder at least three. Journal of Computational and Applied Mathematics, 183(1): 53–66, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000026.

Huang:2005:SES

[HR05c] Ting-Zhu Huang and Rui-Sheng Ran. A simple estimation for the spectral radius of (block) H-matrices. Journal of Computational and Applied Mathematics, 177 (2):455–459, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004765.

Hu:2009:DPN

[HR09a] Lanying Hu and Yong Ren. Doubly perturbed neutral stochastic functional equations. Journal of Computational and Applied Mathematics, 231(1):319–326, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042709000983.

Hu:2009:SPN

[HR09b] Lanying Hu and Yong Ren. Stochastic PDIEs with nonlinear Neumann boundary conditions and generalized backward doubly stochastic differential equations driven by Lévy processes. Journal of Computational and Applied Mathematics, 229 (1):230–239, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005608.

Heidenreich:2008:FOC

[HRGD08] E. A. Heidenreich, J. F.

Rodríguez, F. J. Gaspar, and M. Doblaré. Fourthorder compact schemes with adaptive time step for monodomain reaction–diffusion equations. Journal of Computational and Applied Mathematics, 216(1):39–55, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002063.

Hunter:2005:QFC

[HS05] D. B. Hunter and H. V.

Smith. A quadrature formula of Clenshaw–Curtis type for the Gegenbauer weightfunction. Journal of Computational and Applied Mathematics, 177(2):389–400, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004352.

Hill:2006:LSE

[HS06] A. A. Hill and B. Straughan.

A Legendre spectral element method for eigenvalues in hydrodynamic stability. Journal of Computational and Applied Mathematics, 193(1): 363–381, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004000.

Hasegawa:2007:QRI

[HS07a] Takemitsu Hasegawa and Hiroshi Sugiura. Quadrature rule for indefinite integral of algebraic-logarithmic singular integrands. Journal of Computational and Applied Mathematics, 205(1): 487–496, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003335.

Hazra:2007:CDT

[HS07b] S. B. Hazra and K. Steiner. Computation of dilute twophase flow in a pump. Journal of Computational and Applied Mathematics, 203 (2):444–460, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002196.

Hashemian:2008:MAS

[HS08a] Alireza Hashemian and Hossein M. Shodja. A meshless approach for solution of Burgers’ equation. Journal of Computational and Applied Mathematics, 220(1–2): 226–239, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004384.

Heyouni:2008:NIC

[HS08b] M. Heyouni and H. Sadok.

A new implementation of the CMRH method for solving dense linear systems. Journal of Computational and Applied Mathematics, 213 (2):387–399, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000386.

Hussain:2008:PPI

[HS08c] Malik Zawwar Hussain and Muhammad Sarfraz. Positivitypreserving interpolation of positive data by rational cubics. Journal of Computational and Applied Mathematics, 218(2):446–458, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042707002841.

Habib:2009:CTB

[HS09a] Zulfiqar Habib and Manabu Sakai. G^{2} cubic transition between two circles with shape control. Journal of Computational and Applied Mathematics, 223(1):133–144, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006747.

Hasegawa:2009:UAF

[HS09b] Takemitsu Hasegawa and Hiroshi Sugiura. Uniform approximation to fractional derivatives of functions of algebraic singularity. Journal of Computational and Applied Mathematics, 228 (1):247–253, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800472X.

Hsiao:2009:HFM

[Hsi09] Chun-Hui Hsiao. Hybrid function method for solving Fredholm and Volterra integral equations of the second kind. Journal of Computational and Applied Mathematics, 230(1):59–68, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005864.

Hu:2008:KTT

[HSL08] Nuchun Hu, Weiping Shen, and Chong Li. Kantorovich’s type theorems for systems of equations with constant rank derivatives. Journal of Computational and Applied Mathematics, 219(1): 110–122, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700369X.

Hertog:2009:MCE

[HSN09] Maarten L. A. T. M. Hertog, Nico Scheerlinck, and Bart M. Nicolaï. Monte Carlo evaluation of biological variation: Random generation of correlated non-Gaussian model parameters. Journal of Computational and Applied Mathematics, 223(1):1–14, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700653X.

Hsu:2005:ANM

[Hsu05] Viktoria R. T. Hsu. Almost Newton method for large flux steady-state of 1D Poisson–Nernst–Planck equations. Journal of Computational and Applied Mathematics, 183(1):1–15, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006405.

Hu:2006:AOR

[HT06] Min Hu and Jieqing Tan.

Adaptive osculatory rational interpolation for image processing. Journal of Computational and Applied Mathematics, 195(1–2):46–53, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004711.

Hao:2009:SBC

[HTS09] Dinh Nho Hào, Nguyen Trung Thành, and Hichem Sahli. Splitting-based conjugate gradient method for a multidimensional linear inverse heat conduction problem. Journal of Computational and Applied Mathematics, 232(2):361–377, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900363X.

Hu:2005:RES

[Hu05a] Hongchang Hu. Ridge estimation of a semiparametric regression model. Journal of Computational and Applied Mathematics, 176 (1):215–222, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003292.

Hu:2005:NSC

[Hu05b] Zhongyong Hu. A new semilocal convergence theorem for Newton’s method involving twice Fréchet-differentiability at only one point. Journal of Computational and Applied Mathematics, 181(2): 321–325, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005874.

Huang:2007:CCP

[Hua07]  Xiaoxia Huang.  Chance-

constrained programming models for capital budgeting with NPV as fuzzy parameters. Journal of Computational and Applied Mathematics, 198(1):149–159, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007120.

Huang:2008:MSM

[Hua08]  Xiaoxia Huang.  Mean-

semivariance models for fuzzy portfolio selection. Journal of Computational and Applied Mathematics, 217(1): 1–8, July 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700310X.

Hugger:2006:WBV

[Hug06] Jens Hugger. Wellposedness of the boundary value formulation of a fixed strike Asian option. Journal of Computational and Applied Mathematics, 185(2):460–481, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001263.

Huillet:2005:SAR

[Hui05] Thierry Huillet. Statistical aspects of random fragmentations. Journal of Computational and Applied Mathematics, 181(2):364– 387, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005916.

Huillet:2007:ESF

[Hui07]  Thierry Huillet.  Ewens

sampling formulae with and without selection. Journal of Computational and Applied Mathematics, 206(2): 755–773, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005206.

Huybrechs:2009:SHO

[Huy09] Daan Huybrechs. Stable highorder quadrature rules with equidistant points. Journal of Computational and Applied Mathematics, 231(2): 933–947, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003239.

Huybrechs:2005:CQF

[HV05] Daan Huybrechs and Stefan Vandewalle. Composite quadrature formulae for the approximation of wavelet coefficients of piecewise smooth and singular functions. Journal of Computational and Applied Mathematics, 180(1): 119–135, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004893.

Haslinger:2006:ANR

[HV06a] J. Haslinger and O. Vlach.

Approximation and numerical realization of 2D contact problems with Coulomb friction and a solution-dependent coefficient of friction. Journal of Computational and Applied Mathematics, 197(2): 421–436, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006898.

Helsen:2006:NSC

[HV06b] S. Helsen and M. Van Barel.

A numerical solution of the constrained energy problem. Journal of Computational and Applied Mathematics, 189(1–2):442–452, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001664.

Huybrechs:2006:TDW

[HV06c] Daan Huybrechs and Stefan Vandewalle. A twodimensional wavelet-packet transform for matrix compression of integral equations with highly oscillatory kernel. Journal of Computational and Applied Mathematics, 197(1):218–232, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006552.

Haelterman:2009:GRK

[HVV09a] R. Haelterman, J. Vierendeels, and D. Van Heule. A generalization of the Runge– Kutta iteration. Journal of Computational and Applied Mathematics, 224(1): 152–167, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001866.

Hollevoet:2009:OEF

[HVV09b] D. Hollevoet, M. Van Daele, and G. Vanden Berghe. The optimal exponentiallyfitted Numerov method for solving two-point boundary value problems. Journal of Computational and Applied Mathematics, 230(1): 260–269, August 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006225.

Hu:2007:NSC

[HW07] Qian-Qian Hu and Guo-Jin Wang. Necessary and sufficient conditions for rational quartic representation of conic sections. Journal of Computational and Applied Mathematics, 203 (1):190–208, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001944.

Hu:2008:OMD

[HW08] Qian-Qian Hu and Guo-Jin Wang. Optimal multi-degree reduction of triangular Bézier surfaces with corners continuity in the norm L_{2}. Journal of Computational and Applied Mathematics, 215 (1):114–126, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002178.

Huang:2009:SPH

[HWL09] Ting-Zhu Huang, Shi-Liang Wu, and Cui-Xia Li. The spectral properties of the Hermitian and skew-Hermitian splitting preconditioner for generalized saddle point problems. Journal of Computational and Applied Mathematics, 229(1):37–46, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005165.

He:2005:RSD

[HWSL05] Yong He, Min Wu, Jin-Hua She, and Guo-Ping Liu. Robust stability for delay Lur’e control systems with multiple nonlinearities. Journal of Computational and Applied Mathematics, 176 (2):371–380, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003437.

He:2009:EUP

[HWY09] Zhanbing He, Wentao Wang, and Xuejun Yi. Existence and uniqueness of periodic solutions for Rayleigh type p-Laplacian equation. Journal of Computational and Applied Mathematics, 232(2): 558–564, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003872.

Hamal:2009:PSN

[HY09a] N. Aykut Hamal and Fulya Yoruk. Positive solutions of nonlinear m-point boundary value problems on time scales. Journal of Computational and Applied Mathematics, 231(1):92–105, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000430.

Hickernell:2009:SAE

[HY09b] Fred J. Hickernell and Shijun Yang. Simplified analytical expressions for numerical differentiation via cycle index. Journal of Computational and Applied Mathematics, 224(1):433–443, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002501.

Hu:2005:GHF

[HYC05] Yong-Jian Hu, Zheng-Hong Yang, and Gong-Ning Chen. Generalized Hermite formula for the two-sided Lagrange– Sylvester interpolation. Journal of Computational and Applied Mathematics, 173(2): 345–358, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001670.

Huang:2008:IMS

[HYG08a] Guang-Xin Huang, Feng Yin, and Ke Guo. An iterative method for the skewsymmetric solution and the optimal approximate solution of the matrix equation AXB = C. Journal of Computational and Applied Mathematics, 212(2): 231–244, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007266.

Huang:2008:LUB

[HYG08b] Guang-Xin Huang, Feng Yin, and Ke Guo. The lower and upper bounds on Perron root of nonnegative irreducible matrices. Journal of Computational and Applied Mathematics, 217(1):259–267, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700355X.

Huang:2006:SAN

[HyLC06] Nan-Jing Huang, Heng you Lan, and Yeol Je Cho. Sensitivity analysis for nonlinear generalized mixed implicit equilibrium problems with non-monotone setvalued mappings. Journal of Computational and Applied Mathematics, 196(2): 608–618, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006242.

Hayashi:2005:MSM

[HYYF05] Shunsuke Hayashi, Takahiro Yamaguchi, Nobuo Yamashita, and Masao Fukushima. A matrix-splitting method for symmetric affine second-order cone complementarity problems. Journal of Computational and Applied Mathematics, 175(2):335–353, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002389.

Huang:2005:DNT

[HZ05] Yu Huang and Xingfu Zou.

Dynamics in numerics: on two different finite difference schemes for ODEs. Journal of Computational and Applied Mathematics, 181(2): 388–403, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005928.

Han:2006:ABC

[HZBM06] Houde Han, Liang Zhu, Hermann Brunner, and Jingtang Ma. Artificial boundary conditions for parabolic Volterra integro-differential equations on unbounded twodimensional domains. Journal of Computational and Applied Mathematics, 197(2): 406–420, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006886.

Huang:2008:STA

[HZW08] Zheng-Hai Huang, Ying Zhang, and Wei Wu. A smoothing-type algorithm for solving system of inequalities. Journal of Computational and Applied Mathematics, 220(1–2):355–363, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004566.

Ikbal:2009:URN

[ICW^{+}09] Md. A. Ikbal, S. Chakravarty, Kelvin K. L. Wong, J. Mazumdar, and P. K. Mandal. Unsteady response of non-Newtonian blood flow through a stenosed artery in magnetic field. Journal of Computational and Applied Mathematics, 230(1):243–259, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006031.

Ivanov:2005:PSS

[IL05] Anatoli Ivanov and Eduardo Liz. Periodic solutions of a singular differential delay equation with the Fareytype nonlinearity. Journal of Computational and Applied Mathematics, 180(1): 137–145, August 1, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400490X.

Ide:2006:NSN

[IO06] Takanori Ide and Masami Okada. Numerical simulation for a nonlinear partial differential equation with variable coefficients by means of the discrete variational derivative method. Journal of Computational and Applied Mathematics, 194(2):425–459, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005133.

Ida:2007:MAH

[IOO07] A. Ida, S. Oharu, and Y. Oharu. A mathematical approach to HIV infection dynamics. Journal of Computational and Applied Mathematics, 204(1):172–186, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002500.

Ida:2007:ACA

[IOS07] A. Ida, Y. Oharu, and O. Sankey. An attempt at the computer-aided management of HIV infection. Journal of Computational and Applied Mathematics, 204 (1):159–165, July 1, 2007.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002470.

Ivanova:2006:GCV

[IS06] N. M. Ivanova and C. Sophocleous. On the group classification of variablecoefficient nonlinear diffusionconvection equations. Journal of Computational and Applied Mathematics, 197(2): 322–344, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006722.

Ismail:2009:DOO

[IS09] Mourad E. H. Ismail and Plamen Simeonov. qdifference operators for orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(3): 749–761, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001228.

Ishikawa:2007:NME

[Ish07]  Hideaki Ishikawa.  Nu-

merical methods for the eigenvalue determination of second-order ordinary differential equations. Journal of Computational and Applied Mathematics, 208(2):

404–424, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006133.

Ismail:2005:PMP

[Ism05a] Mourad Ismail. Problem 4.

A moment problem. Journal of Computational and Applied Mathematics, 178(1– 2):531–532, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004042.

Ismail:2005:POC

[Ism05b] Mourad Ismail. Problem 5.

Orthogonality and completeness. Journal of Computational and Applied Mathematics, 178(1–2):533–534, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004054.

Ismail:2005:DOP

[Ism05c] Mourad E. H. Ismail. Determinants with orthogonal polynomial entries. Journal of Computational and Applied Mathematics, 178(1– 2):255–266, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003784.

Ismailov:2007:CES

[Ism07] Vugar E. Ismailov. Characterization of an extremal sum of ridge functions. Journal of Computational and Applied Mathematics, 205(1): 105–115, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002585.

Iavernaro:2007:SDS

[IT07] Felice Iavernaro and Donato Trigiante. State-dependent symplecticity and area preserving numerical methods. Journal of Computational and Applied Mathematics, 205(2):814–825, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004043.

Ilin:2006:AMF

[IV06] K. I. Ilin and V. A. Vladimirov. Asymptotic model for free surface flow of an electrically conducting fluid in a high-frequency magnetic field. Journal of Computational and Applied Mathematics, 190(1– 2):520–531, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002499.

Izzo:2007:DTV

[IV07] Giuseppe Izzo and Antonia Vecchio. A discrete time version for models of population dynamics in the presence of an infection. Journal of Computational and Applied Mathematics, 210(1–2): 210–221, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006637.

Ionescu:2008:ESE

[IV08] Ioan R. Ionescu and Darko Volkov. Earth surface effects on active faults: an eigenvalue asymptotic analysis. Journal of Computational and Applied Mathematics, 220(1–2): 143–162, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004323.

Ivanov:2006:PDS

[Iva06] Ivan G. Ivanov. On positive definite solutions of the family of matrix equations X + A^{∗}X^{−n}A = Q. Journal of Computational and Applied Mathematics, 193(1): 277–301, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003924.

Izadikhah:2009:UHD

[Iza09] Mohammad Izadikhah. Using the Hamming distance to extend TOPSIS in a fuzzy environment. Journal of Computational and Applied Mathematics, 231(1): 200–207, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000533.

Jackson:2006:PDE

[Jac06]  B. Jackson.  Partial dy-

namic equations on time scales. Journal of Computational and Applied Mathematics, 186(2):391–415, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001019.

Jakobsson:2009:RRB

[JAE09] Stefan Jakobsson, Björn Andersson, and Fredrik Edelvik. Rational radial basis function interpolation with applications to antenna design. Journal of Computational and Applied Mathematics, 233(4): 889–904, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004634.

Jais:2006:UIC

[Jai06] Mathias Jais. Unique identifiability of the common support of coefficients of a second order anisotropic elliptic system by the Dirichlet– Neumann map. Journal of Computational and Applied Mathematics, 194(1): 3–16, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003985.

Jang:2008:TPB

[Jan08] Bongsoo Jang. Two-point boundary value problems by the extended Adomian decomposition method. Journal of Computational and Applied Mathematics, 219(1): 253–262, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004062.

Jang:2009:CSC

[Jan09] Bongsoo Jang. Comments on “Solving a class of twodimensional linear and nonlinear Volterra integral equations by the differential transform method”. Journal of Computational and Applied Mathematics, 233(2): 224–230, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004130.

See [TRST09].

Jarlebring:2009:CDP

[Jar09] Elias Jarlebring. Critical delays and polynomial eigenvalue problems. Journal of Computational and Applied Mathematics, 224(1): 296–306, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002094.

Jay:2007:BCR

[Jay07] Laurent O. Jay. Beyond conventional Runge–Kutta methods in numerical integration of ODEs and DAEs by use of structures and local models. Journal of Computational and Applied Mathematics, 204(1):56–76, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002391.

Jayswal:2008:NDM

[Jay08]  Anurag Jayswal.  Non-

differentiable minimax fractional programming with generalized α-univexity. Journal of Computational and Applied Mathematics, 214 (1):121–135, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000891.

Joulak:2009:GCG

[JB09] Hédi Joulak and Bernhard Beckermann. On Gautschi’s conjecture for generalized Gauss–Radau and Gauss– Lobatto formulae. Journal of Computational and Applied Mathematics, 233(3): 768–774, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001241.

Jiang:2009:LHS

[JC09] Mei-Qun Jiang and Yang Cao. On local Hermitian and skew-Hermitian splitting iteration methods for generalized saddle point problems. Journal of Computational and Applied Mathematics, 231(2): 973–982, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003276.

Jiang:2005:CTS

[JD05] Meiqun Jiang and Junliang Dong. On convergence of twostage splitting methods for linear complementarity problems. Journal of Computational and Applied Mathematics, 181(1):58–69, September 1, 2005. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005497.

Jafari:2006:SSN

[JDG06] Hossein Jafari and Varsha Daftardar-Gejji. Solving a system of nonlinear fractional differential equations using Adomian decomposition. Journal of Computational and Applied Mathematics, 196(2):644–651, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006278.

Jenkins:2007:NSA

[Jen07] Eleanor W. Jenkins. Numerical solution of the acoustic wave equation using Raviart– Thomas elements. Journal of Computational and Applied Mathematics, 206(1): 420–431, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004882.

Jannelli:2006:ASS

[JF06] Alessandra Jannelli and Riccardo Fazio. Adaptive stiff solvers at low accuracy and complexity. Journal of Computational and Applied Mathematics, 191(2):246–258, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005534.

Ji:2009:ECM

[JG09] Yude Ji and Yanping Guo.

The existence of countably many positive solutions for some nonlinear n-th order mpoint boundary value problems. Journal of Computational and Applied Mathematics, 232(2):187–200, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003331.

Jing:2008:NIM

[JH08a] Yan-Fei Jing and Ting-Zhu Huang. On a new iterative method for solving linear systems and comparison results. Journal of Computational and Applied Mathematics, 220(1–2):74–84, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004177.

Juhasz:2008:PIC

[JH08b] Imre Juhász and Miklós Hoffmann. On parametrization of interpolating curves. Journal of Computational and Applied Mathematics, 216 (2):413–424, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002725.

Juhasz:2009:QCH

[JH09] Imre Juhász and Miklós Hoffmann. On the quartic curve of han. Journal of Computational and Applied Mathematics, 223(1):124–132, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006735.

Jiang:2006:AAP

[Jia06] X. H. Jiang. Asymptotic analysis of a perturbation problem. Journal of Computational and Applied Mathematics, 190(1–2):22–36, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002190.

Jiang:2009:EPS

[Jia09a] Weihua Jiang. The existence of positive solutions for second-order multi-point BVPs with the first derivative. Journal of Computational and Applied Mathematics, 225(2):387–392, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003877.

jiang:2009:SMV

[jia09b] Ying-Jun jiang. On spectral methods for Volterratype integro-differential equations. Journal of Computational and Applied Mathematics, 230(2):333–340, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006298.

Jin:2007:GCE

[JK07] Yi Jin and Bahman Kalantari. On general convergence in extracting radicals via a fundamental family of iteration functions. Journal of Computational and Applied Mathematics, 206(2): 832–842, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005267.

Jayswal:2009:SNM

[JK09] Anurag Jayswal and Rajnish Kumar. Some nondifferentiable multiobjective programming under generalized d-V-type-I univexity. Journal of Computational and Applied Mathematics, 229 (1):175–182, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005542.

Jablonski:2006:PBF

[JKK06] Bartosz Jablonski, Ryszard Klempous, and Marek Kulbacki. PDE-based filtering of motion sequences. Journal of Computational and Applied Mathematics, 189(1– 2):660–675, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001421.

Jacobs:2007:TET

[JKM07] G. B. Jacobs, D. A. Kopriva, and F. Mashayek. Towards efficient tracking of inertial particles with high-order multidomain methods. Journal of Computational and Applied Mathematics, 206(1): 392–408, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004857.

Johnpillai:2009:ANT

[JKM09] A. G. Johnpillai, A. H. Kara, and F. M. Mahomed. Approximate Noether-type symmetries and conservation laws via partial Lagrangians for PDEs with a small parameter. Journal of Computational and Applied Mathematics, 223(1):508–518, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800040X.

Jackiewicz:2006:NSN

[JL06a] Z. Jackiewicz and E. Lo. Numerical solution of neutral functional differential equations by Adams methods in divided difference form. Journal of Computational and Applied Mathematics, 189(1– 2):592–605, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500141X.

Jiang:2006:ONT

[JL06b] Jianchu Jiang and Xiaoping Li. Oscillation and nonoscillation of two-dimensional difference systems. Journal of Computational and Applied Mathematics, 188 (1):77–88, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001639.

Jiang:2007:ISF

[JL07a] Guirong Jiang and Qishao Lu. Impulsive state feedback control of a predator–prey model. Journal of Computational and Applied Mathematics, 200(1):193–207, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008058.

Johansson:2007:DSD

[JL07b] Tomas Johansson and Daniel Lesnic. Determination of a spacewise dependent heat source. Journal of Computational and Applied Mathematics, 209(1):66–80, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006200.

Jakobson:2006:SPM

[JLNP06] Dmitry Jakobson, Michael Levitin, Nikolai Nadirashvili, and Iosif Polterovich. Spectral problems with mixed Dirichlet–Neumann boundary conditions: Isospectrality and beyond. Journal of Computational and Applied Mathematics, 194(1): 141–155, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004085.

Jakubiak:2006:TSA

[JLR06] Janusz Jakubiak, Fátima Silva Leite, and Rui C. Rodrigues. A two-step algorithm of smooth spline generation on Riemannian manifolds. Journal of Computational and Applied Mathematics, 194(2): 177–191, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004395.

Janas:2007:AAA

[JM07a] Jan Janas and Maria Malejki. Alternative approaches to asymptotic behaviour of eigenvalues of some unbounded Jacobi matrices. Journal of Computational and Applied Mathematics, 200(1):342–356, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000197.

Jiang:2007:EBS

[JM07b] Fangcui Jiang and Fanwei Meng. Explicit bounds on some new nonlinear integral inequalities with delay. Journal of Computational and Applied Mathematics, 205(1): 479–486, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003323.

Jentzen:2009:RES

[JN09] A. Jentzen and A. Neuenkirch.

A random Euler scheme for Carathéodory differential equations. Journal of Computational and Applied Mathematics, 224(1): 346–359, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002136.

John:2005:ATM

[Joh05] Volker John. An assessment of two models for the subgrid scale tensor in the rational LES model. Journal of Computational and Applied Mathematics, 173 (1):57–80, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001347.

Jones:2006:PCF

[Jon06] D. S. Jones. Parabolic cylinder functions of large order. Journal of Computational and Applied Mathematics, 190(1–2):453–469, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002463.

Jagels:2007:SLQ

[JR07] Carl Jagels and Lothar Reichel. Szegő–Lobatto quadrature rules. Journal of Computational and Applied Mathematics, 200(1): 116–126, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007995.

Jiranek:2008:LAS

[JR08] Pavel Jiránek and Miroslav Rozlozník. Limiting accuracy of segregated solution methods for nonsymmetric saddle point problems. Journal of Computational and Applied Mathematics, 215(1):28–37, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001720.

Jana:2009:MOP

[JRM09] P. Jana, T. K. Roy, and S. K.

Mazumder. Multi-objective possibilistic model for portfolio selection with transaction cost. Journal of Computational and Applied Mathematics, 228(1):188–196, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004639.

Jung:2008:IEW

[JS08] H. S. Jung and R. Sakai.

Inequalities with exponential weights. Journal of Computational and Applied Mathematics, 212(2):359–373, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007369.

Jukic:2007:TLS

[JSS07] Dragan Jukić, Kristian Sabo, and Rudolf Scitovski. Total least squares fitting Michaelis–Menten enzyme kinetic model function. Journal of Computational and Applied Mathematics, 201 (1):230–246, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001026.

Johnson:2009:EAZ

[JT09a] Tomas Johnson and Warwick Tucker. Enclosing all zeros of an analytic function — a rigorous approach. Journal of Computational and Applied Mathematics, 228 (1):418–423, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005074.

Jordaan:2009:CZS

[JT09b] Kerstin Jordaan and Ferenc Tookos. Convexity of the zeros of some orthogonal polynomials and related functions. Journal of Computational and Applied Mathematics, 233(3): 762–767, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900123X.

Jian:2005:FDS

[JTHZ05] Jin-Bao Jian, Chun-Ming Tang, Qing-Jie Hu, and Hai-Yan Zheng. A feasible descent SQP algorithm for general constrained optimization without strict complementarity. Journal of Computational and Applied Mathematics, 180(2):391–412, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005400.

Javierre:2006:CNM

[JVVvdZ06] E. Javierre, C. Vuik, F. J.

Vermolen, and S. van der Zwaag. A comparison of numerical models for onedimensional Stefan problems. Journal of Computational and Applied Mathematics, 192(2):445–459, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003730.

Jiang:2008:BAV

[JW08a] Weihua Jiang and Junjie Wei. Bifurcation analysis in van der Pol’s oscillator with delayed feedback. Journal of Computational and Applied Mathematics, 213 (2):604–615, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000751.

Jin:2008:AFE

[JW08b] Jicheng Jin and Xiaonan Wu. Analysis of finite element method for onedimensional time-dependent Schrödinger equation on unbounded domain. Journal of Computational and Applied Mathematics, 220(1–2): 240–256, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004499.

Jisheng:2007:TOM

[JYX07] Kou Jisheng, Li Yitian, and Wang Xiuhua. Third-order modification of Newton’s method. Journal of Computational and Applied Mathematics, 205(1):1–5, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600197X.

Jackiewicz:2009:DVM

[JZK09] Z. Jackiewicz and B. Zubik-Kowal. Discrete variable methods for delay-differential equations with threshold-type delays. Journal of Computational and Applied Mathematics, 228(2):514–523, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001325.

Jiang:2008:NTQ

[JZW08] Tongsong Jiang, Jianli Zhao, and Musheng Wei. A new technique of quaternion equality constrained least squares problem. Journal of Computational and Applied Mathematics, 216 (2):509–513, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002907.

Konnov:2006:DMM

[KA06] I. V. Konnov and M. S. S.

Ali. Descent methods for monotone equilibrium problems in Banach spaces. Journal of Computational and Applied Mathematics, 188 (2):165–179, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500213X.

Kilicman:2007:VLS

[KA07] Adem Kiliçman and Zeyad Abdel Aziz Al Zhour. Vector least-squares solutions for coupled singular matrix equations. Journal of Computational and Applied Mathematics, 206(2):1051– 1069, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005784.

Kim:2008:CDR

[KA08] Hoi Sub Kim and Young Joon Ahn. Constrained degree reduction of polynomials in Bernstein–Bézier form over simplex domain. Journal of Computational and Applied Mathematics, 216 (1):14–19, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001835.

Kim:2005:CQR

[KAC05] Philsu Kim, Soyoung Ahn, and U. Jin Choi. On the Chebyshev quadrature rule of finite part integrals. Journal of Computational and Applied Mathematics, 180(1): 147–159, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004911.

Kageyama:2009:CMD

[Kag09] Masayuki Kageyama. Credibilistic Markov decision processes: the average case. Journal of Computational and Applied Mathematics, 224(1):140–145, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001842.

Kalovics:2005:ZI

[Kál05]  Ferenc Kálovics.  Zones

and integrals. Journal of Computational and Applied Mathematics, 182(2): 243–251, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006119.

Kamoun:2007:BTS

[Kam07] Lotfi Kamoun. Besov-type spaces for the Dunkl operator on the real line. Journal of Computational and Applied Mathematics, 199(1): 56–67, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007387.

Kang:2005:ETP

[Kan05] Mihyun Kang. Efficiency test of pseudorandom number generators using random walks. Journal of Computational and Applied Mathematics, 174(1):165–177, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001864.

Kanschat:2008:RSH

[Kan08] Guido Kanschat. Robust smoothers for high-order discontinuous Galerkin discretizations of advection– diffusion problems. Journal of Computational and Applied Mathematics, 218 (1):53–60, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003263.

Ketabchi:2007:SSC

[KAP07] S. Ketabchi and E. Ansari-Piri. On the solution set of convex problems and its numerical application. Journal of Computational and Applied Mathematics, 206(1): 288–292, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004493.

Kar:2006:MAH

[Kar06] Tapan Kumar Kar. Modelling and analysis of a harvested prey-predator system incorporating a prey refuge. Journal of Computational and Applied Mathematics, 185 (1):19–33, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000488.

Karaca:2007:DTO

[Kar07] Ilkay Yaslan Karaca. Discrete third-order three-point boundary value problem. Journal of Computational and Applied Mathematics, 205(1):458–468, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600330X.

Karaca:2009:PSF

[Kar09] Ilkay Yaslan Karaca. On positive solutions for fourth-order boundary value problem with impulse. Journal of Computational and Applied Mathematics, 225(2):356–364, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800383X.

Kassem:2006:GSU

[Kas06] M. Kassem. Group solution for unsteady free-convection flow from a vertical moving plate subjected to constant heat flux. Journal of Computational and Applied Mathematics, 187(1):72–86, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001469.

Kawanago:2007:ICT

[Kaw07] Tadashi Kawanago. Improved convergence theorems of Newton’s method designed for the numerical verification for solutions of differential equations. Journal of Computational and Applied Mathematics, 199(2):365–371, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007739.

Kazmi:2006:IAG

[Kaz06] Kaleem Raza Kazmi. Iterative algorithm for generalized quasi-variational-like inclusions with fuzzy mappings in Banach spaces. Journal of Computational and Applied Mathematics, 188 (1):1–11, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500155X.

Kazmi:2009:IAB

[KBA09] K. R. Kazmi, M. I. Bhat, and Naeem Ahmad. An iterative algorithm based on M-proximal mappings for a system of generalized implicit variational inclusions in Banach spaces. Journal of Computational and Applied Mathematics, 233(2): 361–371, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004312.

Kyrychko:2005:PTW

[KBB05] Y. N. Kyrychko, M. V. Bartuccelli, and K. B. Blyuss. Persistence of travelling wave solutions of a fourth order diffusion system. Journal of Computational and Applied Mathematics, 176(2): 433–443, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003474.

Khromov:2008:CDF

[KBSK08] A. I. Khromov, A. A.

Bukhanko, S. L. Stepanov, and E. P. Kocherov. Concentrators of deformations and fracture of plastic bodies. Journal of Computational and Applied Mathematics, 215(2):457–466, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007564.

Kim:2007:FDI

[KC07] Kyung Joong Kim and Seung Hoe Choi. Frequencydependent interpolation rules using first derivatives for oscillatory functions. Journal of Computational and Applied Mathematics, 205(1): 149–160, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002615.

Kalita:2006:SDH

[KCK06] Jiten C. Kalita, Puneet Chhabra, and Sudhanshu Kumar. A semi-discrete higher order compact scheme for the unsteady two-dimensional Schrödinger equation. Journal of Computational and Applied Mathematics, 197(1): 141–149, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006485.

Katayama:2009:CSH

[KCK09] N. Katayama, M. Chen, and M. Kubo. A capacity scaling heuristic for the multicommodity capacitated network design problem. Journal of Computational and Applied Mathematics, 232(1): 90–101, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005487.

Kelner:2008:HOT

[KCLW08] Vincent Kelner, Florin Capitanescu, Olivier Léonard, and Louis Wehenkel. A hybrid optimization technique coupling an evolutionary and a local search algorithm. Journal of Computational and Applied Mathematics, 215 (2):448–456, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007552.

Khater:2006:ESS

[KCS06] A. H. Khater, D. K. Callebaut, and S. M. Sayed. Exact solutions for some nonlinear evolution equations which describe pseudo-spherical surfaces. Journal of Computational and Applied Mathematics, 189(1–2):387–411, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006175.

Kaba:2005:STL

[KD05] Ibrahima K. Kaba and Weizhong Dai. A stable three-level finite difference scheme for solving the parabolic two-step model in a 3D micro-sphere heated by ultrashort-pulsed lasers. Journal of Computational and Applied Mathematics, 181(1):125–147, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005576.

Kumar:2006:CCP

[KD06] Arun Kumar and V. B.

Das. Convergence of cubic piecewise function. Journal of Computational and Applied Mathematics, 194(2): 388–394, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005017.

Kasbah:2008:MSR

[KDH08] Safaa J. Kasbah, Issam W. Damaj, and Ramzi A. Haraty. Multigrid solvers in reconfigurable hardware. Journal of Computational and Applied Mathematics, 213 (1):79–94, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000076.

Kalita:2008:ETN

[KDN08] Jiten C. Kalita, Anoop K. Dass, and Nimisha Nidhi. An efficient transient Navier– Stokes solver on compact nonuniform space grids. Journal of Computational and Applied Mathematics, 214 (1):148–162, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001033.

Kersey:2009:CVR

[Ker09] Scott N. Kersey. Constrained variational refinement. Journal of Computational and Applied Mathematics, 223(2):

983–996, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001209.

Kemaloglu:2009:RAU

[KG09] Sibel Acik Kemaloğlu and Omer L. Gebizlioğlu. Risk analysis under progressive type II censoring with binomial claim numbers. Journal of Computational and Applied Mathematics, 233(1): 61–72, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005633.

Kadalbajoo:2008:UCB

[KGA08] Mohan K. Kadalbajoo, Vikas Gupta, and Ashish Awasthi. A uniformly convergent Bspline collocation method on a nonuniform mesh for singularly perturbed onedimensional time-dependent linear convection–diffusion problem. Journal of Computational and Applied Mathematics, 220(1–2):271–289, October 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004517.

Kroll:2007:FSS

[KGB^{+}07] Norbert Kroll, Nicolas R.

Gauger, Joël Brezillon, Richard Dwight, Antonio Fazzolari, Daniel Vollmer, Klaus Becker, Holger Barnewitz, Volker Schulz, and Subhendu Hazra. Flow simulation and shape optimization for aircraft design. Journal of Computational and Applied Mathematics, 203(2):397–411, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002202.

Kong:2009:CFS

[KGW09] Zhi Kong, Liqun Gao, and Lifu Wang. Comment on “A fuzzy soft set theoretic approach to decision making problems”. Journal of Computational and Applied Mathematics, 223(2): 540–542, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000162.

See [RM07].

Kuzmin:2008:NPE

[KHK08] Dmitri Kuzmin, Antti Hannukainen, and Sergey Korotov. A new a posteriori error estimate for convectionreaction–diffusion problems. Journal of Computational and Applied Mathematics, 218(1):70–78, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003287.

Khodabakhshi:2009:OMA

[Kho09] Mohammad Khodabakhshi.

A one-model approach based on relaxed combinations of inputs for evaluating input congestion in DEA. Journal of Computational and Applied Mathematics, 230(2): 443–450, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006419.

Khrushchev:2005:CFO

[Khr05] Sergey Khrushchev. Continued fractions and orthogonal polynomials on the unit circle. Journal of Computational and Applied Mathematics, 178(1–2):267–303, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003796.

Khuri:2006:SFC

[Khu06] S. A. Khuri. Stokes flow in curved channels. Journal of Computational and Applied Mathematics, 187(2): 171–191, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001512.

Kong:2009:SIN

[KHWF09] Linghua Kong, Jialin Hong, Lan Wang, and Fangfang Fu. Symplectic integrator for nonlinear high order Schrödinger equation with a trapped term. Journal of Computational and Applied Mathematics, 231(2):664–679, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002660.

Karsli:2008:CRN

[KI08] Harun Karsli and Ertan Ibikli. Convergence rate of a new Bézier variant of Chlodowsky operators to bounded variation functions. Journal of Computational and Applied Mathematics, 212(2):431–443, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007898.

Kilic:2007:GOF

[Kil07] Emrah Kilic. The generalized order-k Fibonacci–Pell sequence by matrix methods. Journal of Computational and Applied Mathematics, 209(2):133–145, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006698.

Kim:2005:TFD

[Kim05] Kyung Joong Kim. Twofrequency-dependent Gauss quadrature rules. Journal of Computational and Applied Mathematics, 174(1): 43–55, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001700.

Kapl:2009:MAT

[KJ09] Mario Kapl and Bert Jüttler.

A multiresolution analysis for tensor-product splines using weighted spline wavelets. Journal of Computational and Applied Mathematics, 231(2):828–839, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003021.

Kiepiela:2005:ECP

[KK05] Katarzyna Kiepiela and Dominika Klimek. An extension of the Chebyshev polynomials. Journal of Computational and Applied Mathematics, 178(1–2):305–312, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003802.

Karatson:2006:DMP

[KK06] J. Karátson and S. Korotov.

Discrete maximum principles for finite element solutions of some mixed nonlinear elliptic problems using quadratures. Journal of Computational and Applied Mathematics, 192(1):75–88, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003092.

Karatson:2008:SCP

[KK08] J. Karátson and T. Kurics. Superlinearly convergent PCG algorithms for some nonsymmetric elliptic systems. Journal of Computational and Applied Mathematics, 212 (2):214–230, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007254.

Karasoy:2009:MEC

[KK09a] Durdu Karasoy and Cem Kadilar. Modified estimators for the change point in hazard function. Journal of Computational and Applied Mathematics, 229(1):152–157, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005372.

Keshavarz:2009:FSP

[KK09b] Esmaile Keshavarz and Esmaile Khorram. A fuzzy shortest path with the highest reliability. Journal of Computational and Applied Mathematics, 230(1): 204–212, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005992.

Kim:2009:FFE

[KK09c] Young Pyo Kim and Jae Ryong Kweon. The Fourierfinite element method for the Poisson problem on a nonconvex polyhedral cylinder. Journal of Computational and Applied Mathematics, 233(4):951–968, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709005755.

Kumar:2006:WEB

[KKD06] V. V. K. Srinivas Kumar, B. V. Rathish Kumar, and P. C. Das. Weighted extended B-spline method for the approximation of the stationary Stokes problem. Journal of Computational and Applied Mathematics, 186(2): 335–348, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000981.

Kolev:2006:JPG

[KKLM06] Nikolai Kolev, Ekaterina T.

Kolkovska, and José Alfredo López-Mimbela. Joint probability generating function for a vector of arbitrary indicator variables. Journal of Computational and Applied Mathematics, 186(1): 89–98, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001901.

Kim:2006:CAS

[KKLY06] Changho Kim, Sangdong Kim, Yong Hun Lee, and Jungho Yoon. Convergence analysis for a second-order elliptic equation by a collocation method using scattered points. Journal of Computational and Applied Mathematics, 186(2):450–465, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001068.

Kreinovich:2007:IDT

[KKS^{+}07] Vladik Kreinovich, Olga Kosheleva, Scott A. Starks, Kavitha Tupelly, Graçaliz P. Dimuro, Antônio Carlos da Rocha Costa, and Karen Villaverde. From intervals to domains: Towards a general description of validated uncertainty, with potential applications to geospatial and meteorological data. Journal of Computational and Applied Mathematics, 199(2): 411–417, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007806.

Kang:2006:BSF

[KKW06] Tong Kang, Kwang Ik Kim, and Zhengpeng Wu. An Ebased splitting finite element method for time-dependent eddy current equations. Journal of Computational and Applied Mathematics, 196(2): 358–367, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005698.

Ke:2008:MEA

[KL08a] Jau-Chuan Ke and Chuen-Horng Lin. Maximum entropy approach for batcharrival queue under N policy with an un-reliable server and single vacation. Journal of Computational and Applied Mathematics, 221(1): 1–15, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005195.

Kim:2008:PHP

[KL08b] Gwang-Il Kim and Sunhong Lee. Pythagorean-hodograph preserving mappings. Journal of Computational and Applied Mathematics, 216 (1):217–226, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002312.

Kwak:2009:BIS

[KL09] Do Y. Kwak and Hijin Lee. Block iterative solvers for higher order finite volume methods. Journal of Computational and Applied Mathematics, 232(2): 378–387, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003641.

Kouibia:2006:ADC

[KLLP06] A. Kouibia, A. J. López-Linares, and M. Pasadas. Approximation of discontinuous curves and surfaces with tangent conditions. Journal of Computational and Applied Mathematics, 193 (1):51–64, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003791.

Korman:2008:VBD

[KLO08] Philip Korman, Yi Li, and Tiancheng Ouyang. Verification of bifurcation diagrams for polynomial-like equations. Journal of Computational and Applied Mathematics, 212(2):187–193, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007229.

Kreinovich:2007:IVS

[KLS^{+}07] Vladik Kreinovich, Luc Longpré, Scott A. Starks, Gang Xiang, Jan Beck, Raj Kandathi, Asis Nayak, Scott Ferson, and Janos Hajagos. Interval versions of statistical techniques with applications to environmental analysis, bioinformatics, and privacy in statistical databases. Journal of Computational and Applied Mathematics, 199(2): 418–423, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007818.

Kunze:2009:IPR

[KLV09] H. E. Kunze, D. La Torre, and E. R. Vrscay. Inverse problems for random differential equations using the collage method for random contraction mappings. Journal of Computational and Applied Mathematics, 223(2): 853–861, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000885.

Kou:2007:SMN

[KLW07a] Jisheng Kou, Yitian Li, and Xiuhua Wang. Some modifications of Newton’s method with fifth-order convergence. Journal of Computational and Applied Mathematics, 209(2):146–152, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006704.

Kou:2007:SVO

[KLW07b] Jisheng Kou, Yitian Li, and Xiuhua Wang. Some variants of Ostrowski’s method with seventh-order convergence. Journal of Computational and Applied Mathematics, 209(2):153–159, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006716.

Kuijlaars:2005:RHP

[KM05] A. B. J. Kuijlaars and K. T.- R. McLaughlin. A Riemann– Hilbert problem for biorthogonal polynomials. Journal of Computational and Applied Mathematics, 178(1– 2):313–320, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003814.

Kamieniarz:2006:TMS

[KM06] G. Kamieniarz and R. Matysiak.

Transfer matrix simulation technique: Effectiveness and applicability to the lowdimensional magnetic spin systems. Journal of Computational and Applied Mathematics, 189(1–2):471–480, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006187.

Karoui:2009:SAF

[KM09a] Abderrazek Karoui and Taher Moumni. Spectral analysis of the finite Hankel transform and circular prolate spheroidal wave functions. Journal of Computational and Applied Mathematics, 233(2):315–333, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004245.

Kojadinovic:2009:MDD

[KM09b] Ivan Kojadinovic and Jean-Luc Marichal. On the moments and distribution of discrete Choquet integrals from continuous distributions. Journal of Computational and Applied Mathematics, 230(1):83–94, August 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005888.

Kumar:2009:WOF

[KM09c] Vivek Kumar and Mani Mehra. Wavelet optimized finite difference method using interpolating wavelets for self-adjoint singularly perturbed problems. Journal of Computational and Applied Mathematics, 230(2): 803–812, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000272.

Karmakar:2009:IOM

[KMB09a] S. Karmakar, S. K. Mahato, and A. K. Bhunia. Interval oriented multi-section techniques for global optimization. Journal of Computational and Applied Mathematics, 224(2):476–491, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002549.

Krukier:2009:PTS

[KMB09b] Lev A. Krukier, Tatiana S. Martynova, and Zhong-Zhi Bai. Product-type skew-Hermitian triangular splitting iteration methods for strongly non-Hermitian positive definite linear systems. Journal of Computational and Applied Mathematics, 232 (1):3–16, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005414.

Koepf:2006:GFV

[KMJ06] Wolfram Koepf and Mohammad Masjed-Jamei. A generic formula for the values at the boundary points of monic classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 191(1):98–105, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002645.

Kako:2009:P

[KMS09] Takashi Kako, Frédéric Magoulès, and Hiroshi Suito. Preface. Journal of Computational and Applied Mathematics, 232(1):1–2, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005402.

Khaliq:2009:SSR

[KMVWY09] A. Q. M. Khaliq, J. Martín-Vaquero, B. A. Wade, and M. Yousuf. Smoothing schemes for reaction–diffusion systems with nonsmooth data. Journal of Computational and Applied Mathematics, 223(1):374–386, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000290.

Kiselev:2006:NAO

[KN06] Alexander V. Kiselev and Serguei N. Naboko. Nonselfadjoint operators with almost Hermitian spectrum: Matrix model. I. Journal of Computational and Applied Mathematics, 194(1): 115–130, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004061.

Kohno:2009:NP

[KN09] Toshiyuki Kohno and Hiroshi Niki. A note on the preconditioner P_{m} = (I + S_{m}). Journal of Computational and Applied Mathematics, 225(1):316–319, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003853.

Kan-on:2007:BSP

[Ko07] Yukio Kan-on. Bifurcation structure of positive stationary solutions for a Lotka– Volterra competition model with diffusion. I: Numerical verification of local structure. Journal of Computational and Applied Mathematics, 201(2):317–326, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000847.

Kocak:2008:CIM

[Koç08a] M. Çetin Koçak. A class of iterative methods with thirdorder convergence to solve nonlinear equations. Journal of Computational and Applied Mathematics, 218(2): 290–306, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000738.

Kocak:2008:SGF

[Koç08b] M. Çetin Koçak. Simple geometry facilitates iterative solution of a nonlinear equation via a special transformation to accelerate convergence to third order. Journal of Computational and Applied Mathematics, 218(2): 350–363, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001318.

Kohr:2007:BVP

[Koh07] Mirela Kohr. Boundary value problems for a compressible Stokes system in bounded domains in R^{n}. Journal of Computational and Applied Mathematics, 201 (1):128–145, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000823.

Kohaupt:2008:SME

[Koh08] L. Kohaupt. Solution of the matrix eigenvalue problem V A + A^{∗}V = μV with applications to the study of free linear dynamical systems. Journal of Computational and Applied Mathematics, 213(1):142–165, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700012X.

Koikari:2005:RTA

[Koi05] S. Koikari. Rooted tree analysis of Runge–Kutta methods with exact treatment of linear terms. Journal of Computational and Applied Mathematics, 177(2):427–453, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004650.

Kokonendji:2005:OSS

[Kok05] Célestin C. Kokonendji. On d-orthogonality of the Sheffer systems associated to a convolution semigroup. Journal of Computational and Applied Mathematics, 181(1): 83–91, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005539.

Komori:2007:WOS

[Kom07a] Yoshio Komori. Weak order stochastic Runge–Kutta methods for commutative stochastic differential equations. Journal of Computational and Applied Mathematics, 203(1):57–79, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001737.

Komori:2007:WSO

[Kom07b] Yoshio Komori. Weak secondorder stochastic Runge– Kutta methods for noncommutative stochastic differential equations. Journal of Computational and Applied Mathematics, 206(1): 158–173, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003906.

Komori:2008:WFS

[Kom08] Yoshio Komori. Weak first- or second-order implicit Runge– Kutta methods for stochastic differential equations with a scalar Wiener process. Journal of Computational and Applied Mathematics, 217 (1):166–179, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003469.

Konnov:2007:IAM

[Kon07] Igor V. Konnov. Iterative algorithms for multi-valued inclusions with Z mappings. Journal of Computational and Applied Mathematics, 206(1):358–365, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004730.

Koornwinder:2007:SRA

[Koo07] Tom H. Koornwinder. The structure relation for Askey– Wilson polynomials. Journal of Computational and Applied Mathematics, 207(2): 214–226, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006029.

Korotov:2006:PEE

[Kor06] S. Korotov. A posteriori error estimation of goal-oriented quantities for elliptic type BVPs. Journal of Computational and Applied Mathematics, 191(2):216–227, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005509.

Kato:2007:SSP

[KOS07] Nobuyuki Kato, Shinnosuke Oharu, and Koichi Shitaoka. Size-structured plant population models and harvesting problems. Journal of Computational and Applied Mathematics, 204(1):114–123, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002494.

Koto:2008:IRK

[Kot08a]  Toshiyuki Koto.  IMEX

Runge–Kutta schemes for reaction–diffusion equations. Journal of Computational and Applied Mathematics, 215(1):182–195, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001951.

Koto:2008:SIR

[Kot08b] Toshiyuki Koto. Stability of IMEX Runge–Kutta methods for delay differential equations. Journal of Computational and Applied Mathematics, 211(2):201–212, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007059.

Kou:2008:SVC

[Kou08] Jisheng Kou. Some variants of Cauchy’s method with accelerated fourth-order convergence. Journal of Computational and Applied Mathematics, 213(1):71–78, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000064.

Koyama:2007:EED

[Koy07] Daisuke Koyama. Error estimates of the DtN finite element method for the exterior Helmholtz problem. Journal of Computational and Applied Mathematics, 200 (1):21–31, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007934.

Koyama:2009:EEF

[Koy09] Daisuke Koyama. Error estimates of the finite element method for the exterior Helmholtz problem with a modified DtN boundary condition. Journal of Computational and Applied Mathematics, 232(1):109–121, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005505.

Kozlov:2008:EOS

[Koz08] Roman Kozlov. Exponential operator splitting time integration for spectral methods. Journal of Computational and Applied Mathematics, 222(2):592–607, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006346.

Kozlov:2009:EST

[Koz09] Roman Kozlov. Exponential splitting time integration for pseudospectral methods on moving meshes. Journal of Computational and Applied Mathematics, 228 (1):56–69, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004378.

Kouibia:2007:NDA

[KP07] A. Kouibia and M. Pasadas.

A note on the discrete approximation of discontinuous curves and surfaces. Journal of Computational and Applied Mathematics, 208(2): 373–379, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005917.

Kouibia:2008:AIV

[KP08a] A. Kouibia and M. Pasadas. Approximation by interpolating variational splines. Journal of Computational and Applied Mathematics, 218(2): 342–349, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001276.

Kouibia:2008:ASV

[KP08b] A. Kouibia and M. Pasadas. Approximation by smoothing variational vector splines for noisy data. Journal of Computational and Applied Mathematics, 211(2): 213–222, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007060.

Kashid:2007:CMS

[KPAT07] M. N. Kashid, F. Platte, D. W. Agar, and S. Turek. Computational modelling of slug flow in a capillary microreactor. Journal of Computational and Applied Mathematics, 203 (2):487–497, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002184.

Kloeden:2007:RKM

[KR07] Peter E. Kloeden and Andreas Rößler. Runge–Kutta methods for affinely controlled nonlinear systems. Journal of Computational and Applied Mathematics, 205(2):957–968, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600416X.

Kounchev:2008:PAM

[KR08] Ognyan Kounchev and Hermann Render. Padé approximation for a multivariate Markov transform. Journal of Computational and Applied Mathematics, 219(2): 416–430, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700235X.

Krasikov:2006:EZC

[Kra06] Ilia Krasikov. On extreme zeros of classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 193(1): 168–182, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003869.

Khalifa:2008:CMC

[KRA08] A. K. Khalifa, K. R. Raslan, and H. M. Alzubaidi. A collocation method with cubic B-splines for solving the MRLW equation. Journal of Computational and Applied Mathematics, 212(2): 406–418, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007850.

Kletting:2007:CTG

[KRAH07] Marco Kletting, Andreas Rauh, Harald Aschemann, and Eberhard P. Hofer. Consistency tests in guaranteed simulation of nonlinear uncertain systems with application to an activated sludge process. Journal of Computational and Applied Mathematics, 199(2):213–219, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007533.

Krogh:2006:DMS

[Kro06a] Fred T. Krogh. On developing mathematical software. Journal of Computational and Applied Mathematics, 185(2):196–202, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001081.

Krottje:2006:VMM

[Kro06b] Johannes K. Krottje. A variational meshfree method for solving time-discrete diffusion equations. Journal of Computational and Applied Mathematics, 192(2): 183–204, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002797.

Krutitskii:2005:MJP

[Kru05] P. A. Krutitskii. The modified jump problem for the Laplace equation and singularities at the tips. Journal of Computational and Applied Mathematics, 183(1): 232–240, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000270.

Koepf:2005:PMP

[KS05a] Wolfram Koepf and Dieter Schmersau. Positivity and monotony properties of the de Branges functions. Journal of Computational and Applied Mathematics, 173(2): 279–294, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001578.

Kucuk:2005:OCE

[KS05b] Ismail Kucuk and Ibrahim Sadek. Optimal control of an elastically connected rectangular plate-membrane system. Journal of Computational and Applied Mathematics, 180(2):345–363, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005278.

Kirch:2006:PPC

[KS06] Claudia Kirch and Josef Steinebach. Permutation principles for the change analysis of stochastic processes under strong invariance. Journal of Computational and Applied Mathematics, 186(1):64–88, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001895.

Kaipio:2007:SIP

[KS07a] Jari Kaipio and Erkki Somersalo. Statistical inverse problems: Discretization, model reduction and inverse crimes. Journal of Computational and Applied Mathematics, 198(2):493–504, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007296.

Karageorghis:2007:MDM

[KS07b] Andreas Karageorghis and Yiorgos-Sokratis Smyrlis. Matrix decomposition MFS algorithms for elasticity and thermo-elasticity problems in axisymmetric domains. Journal of Computational and Applied Mathematics, 206(2): 774–795, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005218.

Karp:2007:AAF

[KS07c] D. Karp and S. M. Sitnik.

Asymptotic approximations for the first incomplete elliptic integral near logarithmic singularity. Journal of Computational and Applied Mathematics, 205(1):186–206, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002822.

Kikuchi:2007:RPE

[KS07d] Fumio Kikuchi and Hironobu Saito. Remarks on a posteriori error estimation for finite element solutions. Journal of Computational and Applied Mathematics, 199(2): 329–336, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007685.

Kress:2007:HMS

[KS07e] Rainer Kress and Pedro Serranho. A hybrid method for sound-hard obstacle reconstruction. Journal of Computational and Applied Mathematics, 204(2):418–427, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003694.

Kubo:2007:MMT

[KS07f] Akisato Kubo and Takashi Suzuki. Mathematical models of tumour angiogenesis. Journal of Computational and Applied Mathematics, 204(1):48–55, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600238X.

Kim:2009:ATC

[KS09] Mi-Young Kim and Tsendauysh Selenge. Age-time continuous Galerkin methods for a model of population dynamics. Journal of Computational and Applied Mathematics, 223(2):659–671, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000575.

Koshigoe:2009:DAF

[KSE09] H. Koshigoe, T. Shiraishi, and M. Ehara. Distribution algorithm in finite difference method and its application to a 2D simulation of temperature inversion. Journal of Computational and Applied Mathematics, 232(1):102–108, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005499.

Karp:2007:SET

[KSS07a] D. Karp, A. Savenkova, and S. M. Sitnik. Series expansions for the third incomplete elliptic integral via partial fraction decompositions. Journal of Computational and Applied Mathematics, 207(2):331–337, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006108.

Kuijlaars:2007:SMR

[KSS07b] A. B. J. Kuijlaars, E. B.

Saff, and X. Sun. On separation of minimal Riesz energy points on spheres in Euclidean spaces. Journal of Computational and Applied Mathematics, 199(1): 172–180, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500748X.

Kuijlaars:2007:TIH

[KSVW07] A. B. J. Kuijlaars, H. Stahl, W. Van Assche, and F. Wielonsky. Type II Hermite–Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 207(2): 227–244, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005978.

Kilic:2007:FRB

[KT07a] Emrah Kilic and Dursun Tasci. Factorizations and representations of the backward second-order linear recurrences. Journal of Computational and Applied Mathematics, 201(1):182–197, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000975.

Koulaei:2007:CBI

[KT07b] M. H. Koulaei and F. Toutounian. On computing of block ILU preconditioner for block tridiagonal systems. Journal of Computational and Applied Mathematics, 202 (2):248–257, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001269.

Kuijlaars:2009:ABR

[KT09] A. B. J. Kuijlaars and P. M. J.

Tibboel. The asymptotic behaviour of recurrence coefficients for orthogonal polynomials with varying exponential weights. Journal of Computational and Applied Mathematics, 233(3): 775–785, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001253.

Khater:2008:CSC

[KTH08] A. H. Khater, R. S. Temsah, and M. M. Hassan. A Chebyshev spectral collocation method for solving Burgers’-type equations. Journal of Computational and Applied Mathematics, 222(2):333–350, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006024.

Kuang:2009:ANS

[KTM09] Jiaoxun Kuang, Hongjiong Tian, and Taketomo Mitsui. Asymptotic and numerical stability of systems of neutral differential equations with many delays. Journal of Computational and Applied Mathematics, 223(2): 614–625, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800054X.

Kulikov:2009:QCI

[Kul09] G. Yu. Kulikov. On quasiconsistent integration by Nordsieck methods. Journal of Computational and Applied Mathematics, 225 (1):268–287, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003683.

Kuzmin:2008:DAF

[Kuz08] Dmitri Kuzmin. On the design of algebraic flux correction schemes for quadratic finite elements. Journal of Computational and Applied Mathematics, 218(1): 79–87, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003299.

Kostic:2005:RRE

[KV05] Aleksandra Kostić and Heinrich Voss. Recurrence relations for the even and odd characteristic polynomials of a symmetric Toeplitz matrix. Journal of Computational and Applied Mathematics, 173(2):365–369, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001724.

Koleva:2007:BCS

[KV07] Miglena Koleva and Lubin Vulkov. Blow-up of continuous and semidiscrete solutions to elliptic equations with semilinear dynamical boundary conditions of parabolic type. Journal of Computational and Applied Mathematics, 202(2):414–434, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600149X.

Kozak:2009:NCC

[KV09a] Jernej Kozak and Vito Vitrih.

Newton–Cotes cubature rules over (d + 1)-pencil lattices. Journal of Computational and Applied Mathematics, 231(1):392–402, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001848.

Kroner:2009:PEE

[KV09b] Axel Kröner and Boris Vexler.

A priori error estimates for elliptic optimal control problems with a bilinear state equation. Journal of Computational and Applied Mathematics, 230(2): 781–802, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000260.

Khaliq:2008:AMP

[KVK08] Abdul Q. M. Khaliq, David A. Voss, and Kamran Kazmi. Adaptive θ-methods for pricing American options. Journal of Computational and Applied Mathematics, 222(1): 210–227, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005584.

Korostyshevskiy:2007:HSM

[KW07] Valeriy R. Korostyshevskiy and Thomas Wanner. A Hermite spectral method for the computation of homoclinic orbits and associated functionals. Journal of Computational and Applied Mathematics, 206 (2):986–1006, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005632.

Kong:2009:CTM

[KW09] Weibin Kong and Xionghua Wu. Chebyshev tau matrix method for Poisson-type equations in irregular domain. Journal of Computational and Applied Mathematics, 228(1):158–167, June 1, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004585.

Kuznetsov:2006:BPP

[KY06] E. B. Kuznetsov and A. Yu.

Yakimovich. The best parameterization for parametric interpolation. Journal of Computational and Applied Mathematics, 191(2):239–245, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005522.

Knowles:2007:RGP

[KY07] Ian Knowles and Aimin Yan.

The reconstruction of groundwater parameters from head data in an unconfined aquifer. Journal of Computational and Applied Mathematics, 208(1):72–81, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006273.

Kanzow:2005:ELM

[KYF05a] Christian Kanzow, Nobuo Yamashita, and Masao Fukushima. Erratum to “Levenberg– Marquardt methods with strong local convergence properties for solving nonlinear equations with convex constraints”: [J. Comput. Appl. Math. 173 (2005) 321–343]. Journal of Computational and Applied Mathematics, 177(2):241, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006417.

See [KYF05b].

Kanzow:2005:WLM

[KYF05b] Christian Kanzow, Nobuo Yamashita, and Masao Fukushima. WITHDRAWN: Levenberg– Marquardt methods with strong local convergence properties for solving nonlinear equations with convex constraints. Journal of Computational and Applied Mathematics, 173(2): 321–343, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001645.

See [KYF05a].

Larsson:2008:MDO

[LÅH08] Elisabeth Larsson, Krister Åhlander, and Andreas Hall. Multi-dimensional option pricing using radial basis functions and the generalized Fourier transform. Journal of Computational and Applied Mathematics, 222(1): 175–192, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005560.

Lai:2009:DPS

[Lai09a]  Shaoyong Lai.  Different

physical structures of solutions for a generalized Boussinesq wave equation. Journal of Computational and Applied Mathematics, 231(1): 311–318, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900096X.

Lai:2009:AMC

[Lai09b]  Yongzeng Lai.  Adap-

tive Monte Carlo methods for matrix equations with applications. Journal of Computational and Applied Mathematics, 231(2): 705–714, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002702.

Lang:2005:ACB

[Lan05] Jens Lang. Adaptive computation for boundary control of radiative heat transfer in glass. Journal of Computational and Applied Mathematics, 183(2):312– 326, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000683.

Lan:2006:MES

[Lan06] K. Q. Lan. Multiple eigenvalues for singular Hammerstein integral equations with applications to boundary value problems. Journal of Computational and Applied Mathematics, 189(1– 2):109–119, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001354.

Landsman:2008:MRQb

[Lan08a] Zinoviy Landsman. Minimization of the root of a quadratic functional under a system of affine equality constraints with application to portfolio management. Journal of Computational and Applied Mathematics, 220(1–2): 739–748, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005122.

Landsman:2008:MRQa

[Lan08b] Zinoviy Landsman. Minimization of the root of a quadratic functional under an affine equality constraint. Journal of Computational and Applied Mathematics, 216(2):319–327, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002506.

Lara:2008:OSR

[Lar08] Luis Lara. One-step recursive method for solving systems of differential equations. Journal of Computational and Applied Mathematics, 216(1):210–216, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002294.

Laschet:2008:FLD

[Las08] Gottfried Laschet. Forchheimer law derived by homogenization of gas flow in turbomachines. Journal of Computational and Applied Mathematics, 215(2):467–476, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007576.

Lasser:2009:VAP

[Las09] R. Lasser. Various amenability properties of the L^{1}- algebra of polynomial hypergroups and applications. Journal of Computational and Applied Mathematics, 233(3):786–792, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001265.

Lopez:2007:FOS

[LB07] J. A. López and M. Barreda. A formulation to obtain semianalytical planetary theories using true anomalies as temporal variables. Journal of Computational and Applied Mathematics, 204 (1):77–83, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002408.

Liu:2009:NOC

[LB09] Lihua Liu and Yuzhen Bai. New oscillation criteria for second-order nonlinear neutral delay differential equations. Journal of Computational and Applied Mathematics, 231(2): 657–663, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002659.

See erratum [LB10].

Liu:2010:ENO

[LB10] Lihua Liu and Yuzhen Bai. Erratum to: “New oscillation criteria for second-order nonlinear neutral delay differential equations” [J. Comput. Appl. Math. 231 (2009) 657–663]. Journal of Computational and Applied Mathematics, 233(10):2755, March 15, 2010. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.math.utah.edu/pub/
tex/bib/jcomputapplmath2005.
bib;   http://www.sciencedirect.
com/science/article/pii/
S0377042709007717.  See

[LB09].

Li:2009:TDA

[LBS09] Yung-Ta Li, Zhaojun Bai, and Yangfeng Su. A twodirectional Arnoldi process and its application to parametric model order reduction. Journal of Computational and Applied Mathematics, 226(1):10–21, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002173.

Li:2006:GAP

[LC06a] Jing-Wen Li and Sui Sun Cheng. Globally attractive periodic solution of a perturbed functional differential equation. Journal of Computational and Applied Mathematics, 193(2): 652–657, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004371.

Liu:2006:PPS

[LC06b] Zhijun Liu and Lansun Chen.

Positive periodic solution of a general discrete nonautonomous difference system of plankton allelopathy with delays. Journal of Computational and Applied Mathematics, 197(2): 446–456, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006928.

Lai:2008:EFS

[LC08a] K.-L. Lai and J. L. Crassidis. Extensions of the first and second complex-step derivative approximations. Journal of Computational and Applied Mathematics, 219(1): 276–293, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004086.

Lou:2008:SNN

[LC08b] Xuyang Lou and Baotong Cui. Synchronization of neural networks based on parameter identification and via output or state coupling. Journal of Computational and Applied Mathematics, 222(2):440–457, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006218.

Lu:2008:FSD

[LC08c] Tiao Lu and Wei Cai. A Fourier spectral-discontinuous Galerkin method for timedependent 3-D Schrödinger– Poisson equations with discontinuous potentials. Journal of Computational and Applied Mathematics, 220(1–2): 588–614, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700489X.

Li:2009:EPC

[LC09a] Zhong Li and Fengde Chen.

Extinction in periodic competitive stage-structured Lotka– Volterra model with the effects of toxic substances. Journal of Computational and Applied Mathematics, 231(1):143–153, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900048X.

Liu:2009:ENP

[LC09b] Qingbing Liu and Guoliang Chen. Erratum to: “A note on the preconditioned Gauss–Seidel method for Mmatrices” [J. Comput. Appl. Math. 219 (1) (2008) 59– 71]. Journal of Computational and Applied Mathematics, 228(1):498–502, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/

                                        science/article/pii/S0377042708003890.

See [NKM08].

Lin:2007:GAE

[LCA07] Lai-Jiu Lin, Li-Fang Chen, and Qamrul Hasan Ansari. Generalized abstract economy and systems of generalized vector quasi-equilibrium problems. Journal of Computational and Applied Mathematics, 208(2):341– 353, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005875.

Li:2007:ECN

[LCH07] Zi-Cai Li, Cheng-Sheng Chien, and Hung-Tsai Huang. Effective condition number for finite difference method. Journal of Computational and Applied Mathematics, 198(1):208–235, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007879.

Li:2009:PPD

[LCH09] Zuxiong Li, Lansun Chen, and Jianmin Huang. Permanence and periodicity of a delayed ratio-dependent predator–prey model with Holling type functional response and stage structure. Journal of Computational and Applied Mathematics, 233(2):173–187, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004063.

Lee:2006:NSE

[LCKK06] Chong Hyun Lee, Jae Sang Cha, Kyungsup Kwak, and Jae Moung Kim. Numerically stable and efficient algorithm for vector channel parameter and frequency offset estimation. Journal of Computational and Applied Mathematics, 191(2):179–193, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005479.

Li:2008:STF

[LCYZ08] Jinglu Li, C. B. Clemons, G. W. Young, and J. Zhu. Solutions of two-factor models with variable interest rates. Journal of Computational and Applied Mathematics, 222(1):30–41, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700547X.

Liu:2008:DPP

[LCZ^{+}08] Xu-Zheng Liu, Xia Cui, Guo-Qin Zheng, Jun-Hai Yong, and Jia-Guang Sun. Dynamic PDE parametric curves. Journal of Computational and Applied Mathematics, 220(1–2):322–334, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004530.

Liu:2009:MAH

[LCZ09] Xiao-Qi Liu, Li-Qun Cao, and Qi-Ding Zhu. Multiscale algorithm with high accuracy for the elastic equations in three-dimensional honeycomb structures. Journal of Computational and Applied Mathematics, 233(4): 905–921, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004701.

Long:2006:NPE

[LD06] Nguyen Thanh Long and Alain Pham Ngoc Dinh. On a nonlinear parabolic equation involving Bessel’s operator associated with a mixed inhomogeneous condition. Journal of Computational and Applied Mathematics, 196(1): 267–284, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005637.

Li:2008:EPP

[LD08] Jingwen Li and Chaoxiong Du. Existence of positive periodic solutions for a generalized Nicholson’s blowflies model. Journal of Computational and Applied Mathematics, 221(1): 226–233, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005912.

Liao:2009:DNM

[LDM09] Wenyuan Liao, Mehdi Dehghan, and Akbar Mohebbi. Direct numerical method for an inverse problem of a parabolic partial differential equation. Journal of Computational and Applied Mathematics, 232(2): 351–360, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003628.

Lei:2008:LPS

[Lei08a] Yutian Lei. Limiting properties of the second order Ginzburg–Landau minimizers. Journal of Computational and Applied Mathematics, 214(2):596–609, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001574.

Lei:2008:SRG

[Lei08b] Yutian Lei. Some results on an n-Ginzburg–Landau-type minimizer. Journal of Computational and Applied Mathematics, 217(1):123–136, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003421.

Lewandowski:2006:MMS

[Lew06] Jerome L. V. Lewandowski. Marker method for the solution of nonlinear diffusion equations. Journal of Computational and Applied Mathematics, 196(2): 523–539, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006126.

Li:2007:EGA

[LF07a] Wan-Tong Li and Yong-Hong Fan. Existence and global attractivity of positive periodic solutions for the impulsive delay Nicholson’s blowflies model. Journal of Computational and Applied Mathematics, 201(1):55–68, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000781.

Liu:2007:IEP

[LF07b] Zhongyun Liu and Heike Faßbender. An inverse eigenvalue problem and an associated approximation problem for generalized Kcentrohermitian matrices. Journal of Computational and Applied Mathematics, 206(1):578–585, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005061.

Liu:2008:SPG

[LF08] Zhongyun Liu and H. Faßbender. Some properties of generalized K-centrosymmetric Hmatrices. Journal of Computational and Applied Mathematics, 215(1):38–48, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001732.

Li:2009:SUE

[LFHW09] Zi-Cai Li, Qing Fang, Hung-Tsai Huang, and Yimin Wei. On solution uniqueness of elliptic boundary value problems. Journal of Computational and Applied Mathematics, 233(2): 293–307, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004191.

Li:2008:SOC

[LFS08] An Li, Enmin Feng, and Xuelian Sun. Stochastic optimal control and algorithm of the trajectory of horizontal wells. Journal of Computational and Applied Mathematics, 212(2):419–430, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007862.

Li:2009:IOC

[LFW09] An Li, Enmin Feng, and Lei Wang. Impulsive optimal control model for the trajectory of horizontal wells. Journal of Computational and Applied Mathematics, 223(2): 893–900, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001015.

Lee:2007:SOA

[LG07] Steven L. Lee and C. William Gear. Second-order accurate projective integrators for multiscale problems. Journal of Computational and Applied Mathematics, 201 (1):258–274, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600104X.

Liu:2009:CCD

[LGYH09] Bingwen Liu, Shuhua Gong, Xuejun Yi, and Lihong Huang. Convergence for a class of delayed recurrent neural networks without Mmatrix condition. Journal of Computational and Applied Mathematics, 233(2): 137–141, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004014.

Li:2005:GAP

[LH05a] Wan-Tong Li and Hai-Feng Huo. Global attractivity of positive periodic solutions for an impulsive delay periodic model of respiratory dynamics. Journal of Computational and Applied Mathematics, 174(2):227–238, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001918.

Li:2005:PPS

[LH05b] Wan-Tong Li and Hai-Feng Huo. Positive periodic solutions of delay difference equations and applications in population dynamics. Journal of Computational and Applied Mathematics, 176(2): 357–369, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003425.

Liu:2007:NRA

[LH07a] Bingwen Liu and Lihong Huang. New results of almost periodic solutions for recurrent neural networks. Journal of Computational and Applied Mathematics, 206(1): 293–305, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600450X.

Liu:2007:EBU

[LH07b] Fu-Ti Liu and Ting-Zhu Huang. An error bound for the USAOR method. Journal of Computational and Applied Mathematics, 205(1): 608–616, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003438.

Li:2008:SFEa

[LH08a] Jian Li and Yinnian He. A stabilized finite element method based on two local Gauss integrations for the Stokes equations. Journal of Computational and Applied Mathematics, 214 (1):58–65, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000829.

Li:2008:NRP

[LH08b] Yaqiong Li and Lihong Huang. New results of periodic solutions for forced Rayleigh-type equations. Journal of Computational and Applied Mathematics, 221(1): 98–105, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005274.

Li:2009:AGV

[LH09] Fenglian Li and Yiran He.

An algorithm for generalized variational inequality with pseudomonotone mapping. Journal of Computational and Applied Mathematics, 228(1):212–218, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004676.

Liu:2006:SIE

[LHL06] Jianzhou Liu, Yunqing Huang, and Anping Liao. Some inequalities for eigenvalues of Schur complements of Hermitian matrices. Journal of Computational and Applied Mathematics, 196(2): 439–451, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500587X.

Lin:2009:NEN

[LHL09] Qun Lin, Hung-Tsai Huang, and Zi-Cai Li. New expansions of numerical eigenvalues by Wilson’s element. Journal of Computational and Applied Mathematics, 225 (1):213–226, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003646.

Lv:2011:EHI

[LHL11] Xiao-Guang Lv, T. Z. Huang, and Liang Li. Erratum to: “On the HSS iteration methods for positive definite Toeplitz linear systems” [J. Comput. Appl. Math. 224 (2009) 709–718]. Journal of Computational and Applied Mathematics, 235(9): 3112–3114, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042710006369.

See [GT09].

Li:2005:SND

[LHM05] Wei Nian Li, Maoan Han, and Fan Wei Meng. Some new delay integral inequalities and their applications. Journal of Computational and Applied Mathematics, 180(1): 191–200, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005060.

Lv:2009:NAL

[LHR09] Xiao-Guang Lv, Ting-Zhu Huang, and Zhi-Gang Ren. A new algorithm for linear systems of the Pascal type. Journal of Computational and Applied Mathematics, 225(1):309–315, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003828.

Li:2007:RKT

[LHW07] L. Li, B. Han, and W. Wang.

R–K type Landweber method for nonlinear ill-posed problems. Journal of Computational and Applied Mathematics, 206(1):341–357, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042706004729.

Lv:2008:PFM

[LHW08] Yibing Lv, Tiesong Hu, and Zhongping Wan. A penalty function method for solving inverse optimal value problem. Journal of Computational and Applied Mathematics, 220(1–2):175–180, October 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004347.

Li:2005:GED

[Li05a] Jianping Li. General explicit difference formulas for numerical differentiation. Journal of Computational and Applied Mathematics, 183(1): 29–52, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006454.

Li:2005:FOS

[Li05b] Wei Nian Li. On the forced oscillation of solutions for systems of impulsive parabolic differential equations with several delays. Journal of Computational and Applied Mathematics, 181(1): 46–57, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005485.

Li:2005:CRS

[Li05c] Wen Li. Comparison results for solving preconditioned linear systems. Journal of Computational and Applied Mathematics, 176 (2):319–329, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003401.

Li:2005:NPG

[Li05d] Wen Li. A note on the preconditioned Gauss–Seidel (GS) method for linear systems. Journal of Computational and Applied Mathematics, 182(1):81–90, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006028.

Li:2006:EAF

[Li06a] Jichun Li. Error analysis of finite element methods for 3-D Maxwell’s equations in dispersive media. Journal of Computational and Applied Mathematics, 188 (1):107–120, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001676.

Li:2006:SBV

[Li06b] Xin Li. On solving boundary value problems of modified Helmholtz equations by plane wave functions. Journal of Computational and Applied Mathematics, 195(1– 2):66–82, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004735.

Li:2007:EAM

[Li07a] Jichun Li. Error analysis of mixed finite element methods for wave propagation in double negative metamaterials. Journal of Computational and Applied Mathematics, 209(1):81–96, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006212.

Li:2007:EAT

[Li07b] Zi-Cai Li. Error analysis of the Trefftz method for solving Laplace’s eigenvalue problems. Journal of Computational and Applied Mathematics, 200(1):231–254, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008095.

Li:2008:MPB

[Li08] Wen Li. Multiplicative perturbation bounds for spectral and singular value decompositions. Journal of Computational and Applied Mathematics, 217(1):243–251, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003536.

Li:2009:ADM

[Li09a] Jian-Lin Li. Adomian’s decomposition method and homotopy perturbation method in solving nonlinear equations. Journal of Computational and Applied Mathematics, 228(1):168–173, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004597.

Li:2009:SBC

[Li09b] Xiantao Li. On the stability of boundary conditions for molecular dynamics. Journal of Computational and Applied Mathematics, 231(2): 493–505, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002209.

Li:2009:MFS

[Li09c] Zi-Cai Li. The method of fundamental solutions for annular shaped domains. Journal of Computational and Applied Mathematics, 228 (1):355–372, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005037.

Liao:2005:CNP

[Lia05] Shi-Jun Liao. A challenging nonlinear problem for numerical techniques. Journal of Computational and Applied Mathematics, 181(2):

467–472, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005989.

Ligomenides:2009:RM

[Lig09] Panos A. Ligomenides. The reality of mathematics. Journal of Computational and Applied Mathematics, 227 (1):10–16, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003257.

Linss:2005:SSP

[Lin05]  Torsten Linß.  On a

set of singularly perturbed convection–diffusion equations. Journal of Computational and Applied Mathematics, 180(1):173–179, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005047.

Linden:2007:SGB

[Lin07] Hansjörg Linden. Scaled generalized Bernstein polynomials and containment regions for the zeros of polynomials. Journal of Computational and Applied Mathematics, 206(1):216–228, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004365.

Lin:2008:AES

[Lin08] Zhi Lin. An algorithm of evolutionarily stable strategies for the single-population evolutionary game. Journal of Computational and Applied Mathematics, 217 (1):157–165, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003457.

Lin:2009:ABQ

[Lin09a] Fuming Lin. The asymptotic behavior of quadratic forms in ϕ-mixing random variables. Journal of Computational and Applied Mathematics, 233(2):437–448, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004476.

Lin:2009:TVS

[Lin09b] Yi-Kuei Lin. Time version of the shortest path problem in a stochastic-flow network. Journal of Computational and Applied Mathematics, 228(1):150–157, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004573.

Liu:2005:GCA

[Liu05] Xin-Wei Liu. Global convergence on an active set SQP for inequality constrained optimization. Journal of Computational and Applied Mathematics, 180(1): 201–211, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005072.

Liu:2006:UBO

[Liu06]  Xiaoyan Liu.  Univariate

and bivariate orthonormal splines and cardinal splines on compact supports. Journal of Computational and Applied Mathematics, 195(1–2): 93–105, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004772.

Liu:2007:APS

[Liu07] Bingwen Liu. Almost periodic solutions for shunting inhibitory cellular neural networks without global Lipschitz activation functions. Journal of Computational and Applied Mathematics, 203(1):159–168, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001804.

Liu:2008:PSL

[Liu08a] Bingwen Liu. Periodic solutions for Liénard type p-Laplacian equation with a deviating argument. Journal of Computational and Applied Mathematics, 214 (1):13–18, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000775.

Liu:2008:SLB

[Liu08b] Yuji Liu. On Sturm–Liouville boundary value problems for second-order nonlinear functional finite difference equations. Journal of Computational and Applied Mathematics, 216(2):523–533, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002920.

Liu:2009:APL

[Liu09a]  Bingwen Liu.  An anti-

periodic LaSalle oscillation theorem for a class of functional differential equations. Journal of Computational and Applied Mathematics, 223(2):1081–1086, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001477.

Liu:2009:GSC

[Liu09b] Bingwen Liu. Global stability of a class of delay differential systems. Journal of Computational and Applied Mathematics, 233(2): 217–223, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004129.

Liu:2009:OCN

[Liu09c] Chongyang Liu. Optimal control for nonlinear dynamical system of microbial fed-batch culture. Journal of Computational and Applied Mathematics, 232(2):252–261, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003483.

Liu:2009:PBV

[Liu09d]  Yansheng Liu.  Periodic

boundary value problems for first order functional differential equations with impulse. Journal of Computational and Applied Mathematics, 223(1):27–39, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006553.

Lu:2009:EPS

[LJ09] Shiping Lu and Shan Jin. Existence of periodic solutions for a fourth-order p-Laplacian equation with a deviating argument. Journal of Computational and Applied Mathematics, 230(2):513–520, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006481.

Lin:2006:EUS

[LJL06] Xiaoning Lin, Daqing Jiang, and Xiaoyue Li. Existence and uniqueness of solutions for singular fourth-order boundary value problems. Journal of Computational and Applied Mathematics, 196(1):155–161, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005339.

Li:2009:PDB

[LJM09] Xiaoyue Li, Daqing Jiang, and Xuerong Mao. Population dynamical behavior of Lotka–Volterra system under regime switching. Journal of Computational and Applied Mathematics, 232(2): 427–448, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003768.

Li:2009:CNS

[LkLkP09] Ronghua Li, Ping kei Leung, and Wan kai Pang. Convergence of numerical solutions to stochastic age-dependent population equations with Markovian switching. Journal of Computational and Applied Mathematics, 233 (4):1046–1055, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006128.

Lima:2006:ANI

[LKSC06] P. M. Lima, N. B. Konyukhova, A. I. Sukov, and N. V. Chemetov. Analyticalnumerical investigation of bubble-type solutions of nonlinear singular problems. Journal of Computational and Applied Mathematics, 189(1–2):260–273, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003304.

Landau:2001:AEB

[LL01] L. J. Landau and N. J. Luswili. Asymptotic expansion of a Bessel function integral using hypergeometric functions. Journal of Computational and Applied Mathematics, 132(2):387–397, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/

science/article/pii/S0377042700004416.

See comments [Sto05b].

LeRoux:2005:NSN

[LL05] Alain-Yves Le Roux and Marie-Noelle Le Roux. Numerical solution of a nonlinear reaction diffusion equation. Journal of Computational and Applied Mathematics, 173(2):211–237, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001530.

Luo:2006:CMP

[LL06] Xiaonan Luo and Huimin Luo. A computing model of pressure distribution from tight underwear. Journal of Computational and Applied Mathematics, 195(1–2): 106–112, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004784.

Liu:2007:PGS

[LL07a] Shengqiang Liu and Zhuojun Liu. Permanence of general stage-structured consumerresource models. Journal of Computational and Applied Mathematics, 201(2): 381–388, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000914.

Luo:2007:SAL

[LL07b] Yong Luo and Zhengyi Lu. Stability analysis for Lotka– Volterra systems based on an algorithm of real root isolation. Journal of Computational and Applied Mathematics, 201(2):367–373, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000896.

LeRoux:2008:NSC

[LL08] Alain-Yves Le Roux and Marie-Noelle Le Roux. Numerical solution of a Cauchy problem for nonlinear reaction diffusion processes. Journal of Computational and Applied Mathematics, 214 (1):90–110, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000878.

Luo:2008:SBS

[LLC08] Zhongxuan Luo, Yu Liu, and Lijuan Chen. On structure of bivariate spline space of lower degree over arbitrary triangulation. Journal of Computational and Applied Mathematics, 220(1–2):34–50, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004141.

Li:2009:NIP

[LLD09] Chong-Jun Li, Paola Lamberti, and Catterina Dagnino. Numerical integration over polygons using an eightnode quadrilateral spline finite element. Journal of Computational and Applied Mathematics, 233(2):279– 292, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900418X.

Lai:2006:FSP

[LLL06] Ming-Chih Lai, Zhilin Li, and Xiaobiao Lin. Fast solvers for 3D Poisson equations involving interfaces in a finite or the infinite domain. Journal of Computational and Applied Mathematics, 191(1):106–125, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002657.

Leconte:2008:THT

[LLM08] Nicolas Leconte, Bertrand Langrand, and Eric Markiewicz. Toward a hybrid-Trefftz element with a hole for elastoplasticity? Journal of Computational and Applied Mathematics, 218(1):88–95, August 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003305.

Li:2008:TLS

[lLpZ08] Chen liang Li and Jin ping Zeng. Two-level Schwarz method for solving variational inequality with nonlinear source terms. Journal of Computational and Applied Mathematics, 211(1): 67–75, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006960.

Lan:2009:MFM

[LLS09] Yan-Fei Lan, Yan-Kui Liu, and Gao-Ji Sun. Modeling fuzzy multi-period production planning and sourcing problem with credibility service levels. Journal of Computational and Applied Mathematics, 231(1): 208–221, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000545.

Lu:2007:FSA

[LLT07] Shujuan Lü, Qishao Lu, and E. H. Twizell. Fourier spectral approximation to long-time behaviour of the derivative three-dimensional Ginzburg–Landau equation. Journal of Computational and Applied Mathematics, 198(1):167–186, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007326.

Liu:2009:NSP

[LLW09] Lishan Liu, Bingmei Liu, and Yonghong Wu. Nontrivial solutions of m-point boundary value problems for singular second-order differential equations with a sign-changing nonlinear term. Journal of Computational and Applied Mathematics, 224(1):373–382, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002434.

Li:2008:SRR

[LLZ08] Bingyu Li, Zhuojun Liu, and Lihong Zhi. A structured rank-revealing method for Sylvester matrix. Journal of Computational and Applied Mathematics, 213(1): 212–223, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000271.

Liu:2009:LSA

[LLZ09] Hanze Liu, Jibin Li, and Quanxin Zhang. Lie symmetry analysis and exact explicit solutions for general Burgers’ equation. Journal of Computational and Applied Mathematics, 228(1):1–9, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003130.

Li:2007:FDS

[lLzS07] Fu le Li and Zhi zhong Sun. A finite difference scheme for solving the Timoshenko beam equations with boundary feedback. Journal of Computational and Applied Mathematics, 200(2): 606–627, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000392.

Li:2006:OSO

[LLZY06] Qiaoluan Li, Haiyan Liang, Zhenguo Zhang, and Yuanhong Yu. Oscillation of second order self-conjugate differential equation with impulses. Journal of Computational and Applied Mathematics, 197(1):78–88, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500645X.

Lazzaro:2007:EPW

[LM07a] D. Lazzaro and L. B. Montefusco. Edge-preserving wavelet thresholding for image denoising. Journal of Computational and Applied Mathematics, 210(1–2):222– 231, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006649.

Luo:2007:CFS

[LM07b] Zhongxuan Luo and Zhaoliang Meng. Cubature formulas over the n-sphere. Journal of Computational and Applied Mathematics, 202 (2):511–522, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001580.

Liu:2008:RDO

[LM08] Hui Liu and Siliang Ma. R– D optimized tree-structured compression algorithms with discrete directional wavelet transform. Journal of Computational and Applied Mathematics, 219(1):302– 311, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004116.

Lima:2009:ANI

[LM09a] P. M. Lima and L. Morgado. Analytical-numerical investigation of a singular boundary value problem for a generalized Emden–Fowler equation. Journal of Computational and Applied Mathematics, 229(2):480–487, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001684.

Liu:2009:OCS

[LM09b] Haidong Liu and Fanwei Meng. Oscillation criteria for second order linear matrix differential systems with damping. Journal of Computational and Applied Mathematics, 229(1):222–229, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800558X.

Lin:2009:TRD

[LmLgF09] Chang Lin, Mai mai Lin, and Wei gui Feng. Theoretical results of the dust acoustic waves in a magnetized dusty plasma with different dust particles. Journal of Computational and Applied Mathematics, 231(1): 262–267, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000600.

Lebbah:2007:ESF

[LMR07] Yahia Lebbah, Claude Michel, and Michel Rueher. An efficient and safe framework for solving optimization problems. Journal of Computational and Applied Mathematics, 199(2):372–377, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007740.

Lamnii:2008:RCH

[LMS08a] A. Lamnii, H. Mraoui, and D. Sbibih. Recursive computation of Hermite spherical spline interpolants. Journal of Computational and Applied Mathematics, 213 (2):439–453, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000428.

Lyche:2008:QIP

[LMS08b] Tom Lyche, Carla Manni, and Paul Sablonnière. Quasiinterpolation projectors for box splines. Journal of Computational and Applied Mathematics, 221(2):416– 429, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005730.

Li:2008:IGT

[LmY08] Min Li and Xiao ming Yuan. An improved Goldstein’s type method for a class of variant variational inequalities. Journal of Computational and Applied Mathematics, 214 (1):304–312, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001288.

Li:2008:ECB

[LMZ08] Yaqiong Li, Hua Meng, and Qiyuan Zhou. Exponential convergence behavior of shunting inhibitory cellular neural networks with timevarying coefficients. Journal of Computational and Applied Mathematics, 216 (1):164–169, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002269.

Li:2007:SNS

[LmZbL07] Hong Li, Shou ming Zhong, and Hou biao Li. Some new simple stability criteria of linear neutral systems with a single delay. Journal of Computational and Applied Mathematics, 200 (1):441–447, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000264.

Lafitte:2007:HFN

[LN07a] Olivier Lafitte and Youness Noumir. High frequency and numerical Eulerian methods for aeroacoustic problems. Journal of Computational and Applied Mathematics, 204(2):537–548, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003797.

Lohrengel:2007:DGM

[LN07b] Stephanie Lohrengel and Serge Nicaise. A discontinuous Galerkin method on refined meshes for the twodimensional time-harmonic Maxwell equations in composite materials. Journal of Computational and Applied Mathematics, 206(1): 27–54, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003487.

Li:2009:AMS

[LN09] Hengguang Li and Victor Nistor. Analysis of a modified Schrödinger operator in 2D: Regularity, index, and FEM. Journal of Computational and Applied Mathematics, 224(1):320–338, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002112.

Lu:2009:ABM

[LNL09] Lin-Zhang Lu, Michael K. Ng, and Fu-Rong Lin. Approximation BFGS methods for nonlinear image restoration. Journal of Computational and Applied Mathematics, 226(1):84–91, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002240.

Lander:2006:IEF

[LNW06] A. D. Lander, Q. Nie, and F. Y. M. Wan. Internalization and end flux in morphogen gradient formation. Journal of Computational and Applied Mathematics, 190(1– 2):232–251, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002335.

Leentvaar:2008:CTG

[LO08] C. C. W. Leentvaar and C. W. Oosterlee. On coordinate transformation and grid stretching for sparse grid pricing of basket options. Journal of Computational and Applied Mathematics, 222(1):

193–209, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005572.

Lu:2005:ETO

[LOA05a] Haishen Lü, Donal O’Regan, and Ravi P. Agarwal. Existence theorems for the one-dimensional singular p-Laplacian equation with a nonlinear boundary condition. Journal of Computational and Applied Mathematics, 182(1):188–210, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006107.

Lu:2005:ULS

[LOA05b] Haishen Lü, Donal O’Regan, and Ravi P. Agarwal. Upper and lower solutions for the singular p-Laplacian with sign changing nonlinearities and nonlinear boundary data. Journal of Computational and Applied Mathematics, 181(2):442–466, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005965.

Loghmani:2008:ALS

[Log08] G. B. Loghmani. Application of least square method to arbitrary-order problems with separated boundary conditions. Journal of Computational and Applied Mathematics, 222(2):500–510, December 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006267.

Lopez:2007:AEM

[Lóp07a] José L. López. Asymptotic expansions of Mellin convolutions by means of analytic continuation. Journal of Computational and Applied Mathematics, 200(2): 628–636, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000409.

Lopez:2007:NMO

[Lop07b] L. Lopez. Numerical methods for ordinary differential equations on matrix manifolds. Journal of Computational and Applied Mathematics, 210(1–2):232–243, December 31, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006650.

Lorentzen:2005:BON

[Lor05]  L. Lorentzen.  Biography

of Olav Njåstad. Journal of Computational and Applied Mathematics, 179 (1–2):xv–xix, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004376.

Liao:2008:PSN

[LOZ08] Xinyuan Liao, Zigen Ouyang, and Shengfan Zhou. Permanence of species in nonautonomous discrete Lotka– Volterra competitive system with delays and feedback controls. Journal of Computational and Applied Mathematics, 211(1):1–10, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006777.

Lubuma:2006:UCN

[LP06a] Jean M.-S. Lubuma and Kailash C. Patidar. Uniformly convergent non-standard finite difference methods for self-adjoint singular perturbation problems. Journal of Computational and Applied Mathematics, 191 (2):228–238, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005510.

Luo:2006:RSR

[LP06b] Zhongxuan Luo and Xingxuan Peng. A C ^{1}-rational spline in range restricted interpolation of scattered data. Journal of Computational and Applied Mathematics, 194(2):255–266, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004541.

Lacarbonara:2007:SSS

[LP07a] Walter Lacarbonara and Achille Paolone. On solution strategies to Saint-Venant problem. Journal of Computational and Applied Mathematics, 206(1): 473–497, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005000.

Larcher:2007:NOP

[LP07b] Gerhard Larcher and Friedrich Pillichshammer. A note on optimal point distributions in [0, 1)^{s}. Journal of Computational and Applied Mathematics, 206(2): 977–985, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005620.

Lombard:2007:MDE

[LP07c] Bruno Lombard and Joël Piraux. Modeling 1-D elastic P -waves in a fractured rock with hyperbolic jump conditions. Journal of Computational and Applied Mathematics, 204(2):292–305, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600358X.

Li:2008:SIS

[LP08] Song Li and Yali Pan.

Subdivisions with infinitely supported mask. Journal of Computational and Applied Mathematics, 214(1): 288–303, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001252.

Loisel:2009:SAD

[LP09] Stéphane Loisel and Nicolas Privault. Sensitivity analysis and density estimation for finite-time ruin probabilities. Journal of Computational and Applied Mathematics, 230(1):107–120, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005906.

Lopez:2008:P

[LQW08] Marco A. Lopez, Liqun Qi, and Soon-Yi Wu. Preface. Journal of Computational and Applied Mathematics, 217(2):299–300, August 1, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000908.

Lewis:2009:ATR

[LR09] Bryan Lewis and Lothar Reichel. Arnoldi–Tikhonov regularization methods. Journal of Computational and Applied Mathematics, 226 (1):92–102, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002252.

Luzyanina:2005:NBA

[LRB05] Tatyana Luzyanina, Dirk Roose, and Gennady Bocharov. Numerical bifurcation analysis of immunological models with time delays. Journal of Computational and Applied Mathematics, 184(1): 165–176, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000750.

Li:2007:IMC

[LRH07] Tiancheng Li, Ian Robinson, and Michael Hill. The index of merit of k-th-copy integration lattices. Journal of Computational and Applied Mathematics, 205(1): 394–405, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003268.

Loghin:2006:APN

[LRT06] D. Loghin, D. Ruiz, and A. Touhami. Adaptive preconditioners for nonlinear systems of equations. Journal of Computational and Applied Mathematics, 189(1– 2):362–374, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500364X.

Ledoux:2009:EPC

[LRV^{+}09] V. Ledoux, M. Rizea, M. Van Daele, G. Vanden Berghe, and I. Silisteanu. Eigenvalue problem for a coupled channel Schrödinger equation with application to the description of deformed nuclear systems. Journal of Computational and Applied Mathematics, 228(1):197–211, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004652.

Lorentzen:2005:P

[LRW05] Lisa Lorentzen, Frode Rønning, and Haakon Waadeland. Preface. Journal of Computational and Applied Mathematics, 179(1–2):xiii–xiv, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004364.

Lopez:2005:AAS

[LS05] José L. López and Ester Pérez Sinusía. Asymptotic approximations for a singularly perturbed convection–diffusion problem with discontinuous data in a sector. Journal of Computational and Applied Mathematics, 181(1): 1–23, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005461.

Lamba:2006:MSS

[LS06a] H. Lamba and T. Seaman. Mean-square stability properties of an adaptive timestepping SDE solver. Journal of Computational and Applied Mathematics, 194(2): 245–254, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500453X.

Li:2006:PBV

[LS06b] Jianli Li and Jianhua Shen. Periodic boundary value problems for delay differential equations with impulses. Journal of Computational and Applied Mathematics, 193(2):563–573, September 1, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004322.

Llanas:2006:CAI

[LS06c] B. Llanas and F. J. Sainz.

Constructive approximate interpolation by neural networks. Journal of Computational and Applied Mathematics, 188(2):283–308, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002566.

Levesley:2007:PKA

[LS07a] J. Levesley and X. Sun.

Perturbed kernel approximation on homogeneous manifolds. Journal of Computational and Applied Mathematics, 199(1):159–171, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007478.

Li:2007:EPS

[LS07b] Jianli Li and Jianhua Shen.

Existence of positive solution to a singular secondorder initial value problem with impulses. Journal of Computational and Applied Mathematics, 206(2): 699–712, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005164.

Liang:2007:PBV

[LS07c] Ruixi Liang and Jianhua Shen. Periodic boundary value problem for the first order impulsive functional differential equations. Journal of Computational and Applied Mathematics, 202 (2):498–510, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001579.

Li:2008:LCI

[LS08] Chong Li and Weiping Shen. Local convergence of inexact methods under the Hölder condition. Journal of Computational and Applied Mathematics, 222(2): 544–560, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006310.

Li:2009:FNM

[LS09] Song-Hua Li and Ming-Bao Sun. A fast numerical method for a natural boundary integral equation for the Helmholtz equation. Journal of Computational and Applied Mathematics, 230(2): 341–350, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006304.

Shang:2005:FFM

[lSsZ05] You lin Shang and Lian sheng Zhang. A filled function method for finding a global minimizer on global integer optimization. Journal of Computational and Applied Mathematics, 181(1): 200–210, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005710.

Lee:2006:BDP

[LSTS06] Young He Lee, Lena Sherbakov, Jackie Taber, and Junping Shi. Bifurcation diagrams of population models with nonlinear, diffusion. Journal of Computational and Applied Mathematics, 194(2): 357–367, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004991.

Lin:2009:CPL

[LSW09] Yiqin Lin, Xinghua Shi, and Yimin Wei. On computing PageRank via lumping the Google matrix. Journal of Computational and Applied Mathematics, 224(2): 702–708, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002963.

Lee:2005:ROB

[LT05] C. H. Lee and H. T. Tran. Reduced-order-based feedback control of the Kuramoto–Sivashinsky equation. Journal of Computational and Applied Mathematics, 173(1):1–19, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001268.

Ladoucette:2006:RLC

[LT06a] Sophie A. Ladoucette and Jef L. Teugels. Reinsurance of large claims. Journal of Computational and Applied Mathematics, 186(1): 163–190, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001950.

Ling:2006:AMC

[LT06b] Leevan Ling and Manfred R. Trummer. Adaptive multiquadric collocation for boundary layer problems. Journal of Computational and Applied Mathematics, 188(2):265–282, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002554.

Lai:2007:FFO

[LT07] Ming-Chih Lai and Jui-Ming Tseng. A formally fourth-order accurate compact scheme for 3D Poisson equation in cylindrical coordinates. Journal of Computational and Applied Mathematics, 201(1):175–181, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000963.

Liu:2006:APP

[LTC06] Bing Liu, Zhidong Teng, and Lansun Chen. Analysis of a predator–prey model with Holling II functional response concerning impulsive control strategy. Journal of Computational and Applied Mathematics, 193(1): 347–362, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500395X.

Liu:2008:PSK

[LTM08] Xin-Ge Liu, Mei-Lan Tang, and Ralph R. Martin. Periodic solutions for a kind of Liénard equation. Journal of Computational and Applied Mathematics, 219(1): 263–275, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004074.

Li:2007:NCC

[LTW07] Guoyin Li, Chunming Tang, and Zengxin Wei. New conjugacy condition and related new conjugate gradient methods for unconstrained optimization. Journal of Computational and Applied Mathematics, 202(2):523–539, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001592.

Li:2008:HOM

[LTY08] S. J. Li, K. L. Teo, and X. Q. Yang. Higher-order Mond–Weir duality for setvalued optimization. Journal of Computational and Applied Mathematics, 217(2): 339–349, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000945.

Lu:2005:FOM

[Lu05]  Ya Yan Lu.  A fourth-

order Magnus scheme for Helmholtz equation. Journal of Computational and Applied Mathematics, 173(2): 247–258, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001554.

Lu:2007:VIM

[Lu07] Junfeng Lu. Variational iteration method for solving twopoint boundary value problems. Journal of Computational and Applied Mathematics, 207(1):92–95, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600464X.

Lu:2009:ATP

[Lu09a] Lizheng Lu. Approximating tensor product Bézier surfaces with tangent plane continuity. Journal of Computational and Applied Mathematics, 231(1):412– 422, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002064.

Lu:2009:NCD

[Lu09b] Lizheng Lu. A note on constrained degree reduction of polynomials in Bernstein– Bézier form over simplex domain. Journal of Computational and Applied Mathematics, 229(1):324–326, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005591.

Lui:2009:SDE

[Lui09] S. H. Lui. Spectral domain embedding for elliptic PDEs in complex domains. Journal of Computational and Applied Mathematics, 225(2): 541–557, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004263.

Luo:2006:ABS

[Luo06a] Jiaowan Luo. Asymptotic behavior of solutions of second order quasilinear differential equations with delay depending on the unknown function. Journal of Computational and Applied Mathematics, 185(1):133–143, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000531.

Luo:2006:SBS

[Luo06b]  Jiaowan Luo.  Stochasti-

cally bounded solutions of a nonlinear stochastic differential equations. Journal of Computational and Applied Mathematics, 196(1): 87–93, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005194.

Luo:2007:NES

[Luo07] Jiaowan Luo. A note on exponential stability in p-th mean of solutions of stochastic delay differential equations. Journal of Computational and Applied Mathematics, 198(1):143–148, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007119.

Luo:2008:SSP

[Luo08] Jiaowan Luo. Stability of stochastic partial differential equations with infinite delays. Journal of Computational and Applied Mathematics, 222(2):364–371, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006048.

Luther:2007:CRC

[Lut07] Wolfram Luther. The convergence rate of continued fractions representing solutions of a Riccati equation. Journal of Computational and Applied Mathematics, 199(2): 271–276, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007612.

Lievens:2006:IGT

[LV06] S. Lievens and J. Van der Jeugt. Invariance groups of three term transformations for basic hypergeometric series. Journal of Computational and Applied Mathematics, 197(1):1–14, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006291.

Lievens:2007:SGB

[LV07] S. Lievens and J. Van der Jeugt. Symmetry groups of Bailey’s transformations for _{10}ϕ_{9}-series. Journal of Computational and Applied Mathematics, 206(1): 498–519, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005012.

Loghin:2006:BER

[LvGJ06] Daniel Loghin, Martin van Gijzen, and Eline Jonkers. Bounds on the eigenvalue range and on the field of values of non-Hermitian and indefinite finite element matrices. Journal of Computational and Applied Mathematics, 189(1–2):304–323, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003006.

Ledoux:2005:CME

[LVV05] V. Ledoux, M. Van Daele, and G. Vanden Berghe. CP methods and the evaluation of negative energy Coulomb Whittaker functions. Journal of Computational and Applied Mathematics, 183(1): 168–176, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000233.

Laughlin:2005:CTC

[LW05a] J. Mc Laughlin and Nancy J.

Wyshinski. A convergence theorem for continued fractions of the form K_{n=1}^{∞}a_{n}/1. Journal of Computational and Applied Mathematics, 179(1–2):255–262, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004522.

Li:2005:EGA

[LW05b] Wan-Tong Li and Lin-Lin Wang. Existence and global attractivity of positive periodic solutions of functional differential equations with feedback control. Journal of Computational and Applied Mathematics, 180(2): 293–309, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005242.

Lewanowicz:2006:CBT

[LW06a] Stanislaw Lewanowicz and Pawel Woźny. Connections between two-variable Bernstein and Jacobi polynomials on the triangle. Journal of Computational and Applied Mathematics, 197(2): 520–533, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006977.

Li:2006:NNQ

[LW06b] Chong-Jun Li and Ren-Hong Wang. A new 8-node quadrilateral spline finite element. Journal of Computational and Applied Mathematics, 195(1–2):54–65, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004723.

Luo:2006:SDM

[LW06c] Zhongxuan Luo and Renhong Wang. Structure and dimension of multivariate spline space of lower degree on arbitrary triangulation. Journal of Computational and Applied Mathematics, 195(1–2): 113–133, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004796.

Li:2007:TSE

[LW07] Qinghong Li and Xinyuan Wu. A two-step explicit P - stable method of high phaselag order for linear periodic IVPs. Journal of Computational and Applied Mathematics, 200(1):287–296, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000161.

Liu:2008:EGB

[LW08a] Wei-Bing Liu and Xian-Jia Wang. An evolutionary game based particle swarm optimization algorithm. Journal of Computational and Applied Mathematics, 214 (1):30–35, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000799.

Lu:2008:ACI

[LW08b] Lizheng Lu and Guozhao Wang. Application of Chebyshev II–Bernstein basis transformations to degree reduction of Bézier curves. Journal of Computational and Applied Mathematics, 221(1): 52–65, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005249.

Luo:2008:NCC

[LW08c] H. Z. Luo and H. X. Wu. On necessary conditions for a class of nondifferentiable minimax fractional programming. Journal of Computational and Applied Mathematics, 215(1):103–113, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700177X.

Luo:2008:RBP

[LW08d] H. Z. Luo and H. X. Wu. On the relationships between G-preinvex functions and semistrictly Gpreinvex functions. Journal of Computational and Applied Mathematics, 222(2): 372–380, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700605X.

Liang:2006:SRT

[LWC^{+}06] Xue-Zhang Liang, Ren-Hong Wang, Li-Hong Cui, Jie-Lin Zhang, and Ming Zhang. Some researches on trivariate Lagrange interpolation. Journal of Computational and Applied Mathematics, 195(1–2): 192–205, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500484X.

Lee:2009:ALP

[LWH09a] Wen-Chuan Lee, Jong-Wuu Wu, and Ching-Wen Hong. Assessing the lifetime performance index of products with the exponential distribution under progressively Type II right censored samples. Journal of Computational and Applied Mathematics, 231(2): 648–656, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002647.

Lo:2009:PAA

[LWH09b] Keng-Hsin Lo, Kehluh Wang, and Ming-Feng Hsu. Pricing American Asian options with higher moments in the underlying distribution. Journal of Computational and Applied Mathematics, 223(1): 304–313, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000150.

Lin:2009:LBS

[LWLL09] Qun Lin, Yong Hong Wu, Ryan Loxton, and Shaoyong Lai. Linear B-spline finite element method for the improved Boussinesq equation. Journal of Computational and Applied Mathematics, 224(2):658–667, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002914.

Liu:2007:SAN

[LWMT07] Xin-Ge Liu, Min Wu, Ralph Martin, and Mei-Lan Tang. Stability analysis for neutral systems with mixed delays. Journal of Computational and Applied Mathematics, 202(2):478–497, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001567.

Li:2009:RJD

[LWS09] Bo Li, Rong Wu, and Min Song. A renewal jumpdiffusion process with threshold dividend strategy. Journal of Computational and Applied Mathematics, 228 (1):41–55, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004366.

Lai:2008:ETW

[LWW08] Shaoyong Lai, Y. H. Wu, and B. Wiwatanapataphee. On exact travelling wave solutions for two types of nonlinear K(n, n) equations and a generalized KP equation. Journal of Computational and Applied Mathematics, 212(2):291–299, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600731X.

Li:2006:RCP

[LWYT06] S. J. Li, S. Y. Wu, X. Q.

Yang, and K. L. Teo. A relaxed cutting plane method for semi-infinite semi-definite programming. Journal of Computational and Applied Mathematics, 196(2):459– 473, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005972.

Liao:2006:MMD

[LX06] Guojun Liao and Jiaxing Xue. Moving meshes by the deformation method. Journal of Computational and Applied Mathematics, 195(1– 2):83–92, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004747.

Lu:2009:EDP

[LXSH09] Zhaoyang Lu, Wei Xu, Decai Sun, and Weiguo Han. On the expected discounted penalty functions for two classes of risk processes under a threshold dividend strategy. Journal of Computational and Applied Mathematics, 232(2):582–593, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003896.

Liu:2006:STF

[LXUK06] Zeqing Liu, Yuguang Xu, Jeong Sheok Ume, and Shin Min Kang. Solutions to two functional equations arising in dynamic programming. Journal of Computational and Applied Mathematics, 192(2):251–269, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002840.

Ling:2008:DFF

[LXX08] Ai-Fan Ling, Cheng-Xian Xu, and Feng-Min Xu. A discrete filled function algorithm for approximate global solutions of max-cut problems. Journal of Computational and Applied Mathematics, 220(1–2): 643–660, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005067.

Lv:2009:SGF

[LXZG09] Long-Jin Lv, Jian-Bin Xiao, Lin Zhang, and Lei Gao.

Solutions for a generalized fractional anomalous diffusion equation. Journal of Computational and Applied Mathematics, 225(1): 301–308, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001180.

Li:2005:PAF

[LY05a] Yiming Li and Shao-Ming Yu.

A parallel adaptive finite volume method for nanoscale double-gate MOSFETs simulation. Journal of Computational and Applied Mathematics, 175(1):87–99, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002407.

Lu:2005:LBD

[LY05b] Hongtao Lu and Xinzhen Yu.

Local bifurcations in delayed chaos anticontrol systems. Journal of Computational and Applied Mathematics, 181(1):188–199, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400562X.

Lazarov:2007:SDF

[LY07a] Raytcho Lazarov and Xiu Ye. Stabilized discontinuous finite element approximations for Stokes equations. Journal of Computational and Applied Mathematics, 198(1):236–252, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007880.

Liu:2007:RTS

[LY07b] Huipo Liu and Ningning Yan. Recovery type superconvergence and a posteriori error estimates for control problems governed by Stokes equations. Journal of Computational and Applied Mathematics, 209(2): 187–207, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006765.

Li:2009:SSG

[LY09] Hongjie Li and Dong Yue. Synchronization stability of general complex dynamical networks with time-varying delays: a piecewise analysis method. Journal of Computational and Applied Mathematics, 232(2):149–158, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000971.

Liu:2009:NMA

[LYB09] F. Liu, C. Yang, and K. Burrage. Numerical method and analytical technique of the modified anomalous subdiffusion equation with a nonlinear source term. Journal of Computational and Applied Mathematics, 231(1): 160–176, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000508.

Lo:2007:HRN

[LYT07] D. C. Lo, D. L. Young, and C. C. Tsai. High resolution of 2D natural convection in a cavity by the DQ method. Journal of Computational and Applied Mathematics, 203(1):219–236, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001968.

Liu:2008:GSC

[LYW08] Liying Liu, Shengwei Yao, and Zengxin Wei. The global and superlinear convergence of a new nonmonotone MBFGS algorithm on convex objective functions. Journal of Computational and Applied Mathematics, 220(1–2): 422–438, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004633.

Luo:2009:RFD

[LYZ09] Zhendong Luo, Xiaozhong Yang, and Yanjie Zhou. A reduced finite difference scheme based on singular value decomposition and proper orthogonal decomposition for Burgers equation. Journal of Computational and Applied Mathematics, 229 (1):97–107, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005232.

Li:2005:APD

[LZ05] Xiaojun Li and Chengkui Zhong. Attractors for partly dissipative lattice dynamic systems in l^{2} × l^{2}. Journal of Computational and Applied Mathematics, 177 (1):159–174, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004200.

Liu:2006:SNA

[LZ06] Hongyu Liu and Jun Zou. Some new additive Runge– Kutta methods and their applications. Journal of Computational and Applied Mathematics, 190(1–2):74–98, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002220.

Liang:2008:ECMa

[LZ08a] Sihua Liang and Jihui Zhang.

The existence of countably many positive solutions for nonlinear singular m-point boundary value problems. Journal of Computational and Applied Mathematics, 214(1):78–89, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000866.

Liang:2008:ECMb

[LZ08b] Sihua Liang and Jihui Zhang.

The existence of countably many positive solutions for nonlinear singular m-point boundary value problems on the half-line. Journal of Computational and Applied Mathematics, 222(2): 229–243, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005997.

Li:2009:GBN

[LZ09a] Xiaolin Li and Jialin Zhu.

A Galerkin boundary node method and its convergence analysis. Journal of Computational and Applied Mathematics, 230(1): 314–328, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006286.

Liang:2009:ECMa

[LZ09b] Sihua Liang and Jihui Zhang. The existence of countably many positive solutions for nonlinear singular m-point boundary value problems on time scales. Journal of Computational and Applied Mathematics, 223(1): 291–303, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000137.

Liang:2009:ECMb

[LZ09c] Sihua Liang and Jihui Zhang. Existence of countably many positive solutions of n-th-order m-point boundary value problems. Journal of Computational and Applied Mathematics, 224(2): 527–537, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002598.

Liu:2009:NAP

[LZ09d] Yang Liu and Shouwei Zhao. A new approach to practical stability of impulsive functional differential equations in terms of two measures. Journal of Computational and Applied Mathematics, 223(1):449–458, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000356.

Li:2009:OCT

[LZG^{+}09] Qiaoluan Li, Zhenguo Zhang, Fang Guo, Zhiyong Liu, and Haiyan Liang. Oscillatory criteria for third-order difference equation with impulses. Journal of Computational and Applied Mathematics, 225(1):80–86, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800321X.

Liang:2007:FFM

[LZLH07] Y. M. Liang, L. S. Zhang, M. M. Li, and B. S. Han. A filled function method for global optimization. Journal of Computational and Applied Mathematics, 205 (1):16–31, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002536.

Liang:2009:EMP

[LZW09] Sihua Liang, Jihui Zhang, and Zhiyong Wang. The existence of multiple positive solutions for multi-point boundary value problems on the half-line. Journal of Computational and Applied Mathematics, 228(1):10–19, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004287.

Li:2009:HIA

[LZWQ09] Xiang Li, Yang Zhang, Hau-San Wong, and Zhongfeng Qin. A hybrid intelligent algorithm for portfolio selection problem with fuzzy returns. Journal of Computational and Applied Mathematics, 233(2):264–278, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004178.

Li:2009:GES

[LZY09] Jinxian Li, Fengqin Zhang, and Jurang Yan. Global exponential stability of nonautonomous neural networks with time-varying delays and reaction–diffusion terms. Journal of Computational and Applied Mathematics, 233(2):241–247, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004154.

Li:2009:PNI

[LZZ09] Xin-Yue Li, Yuan-Qing Zhang, and Qiu-Lan Zhao. Positive and negative integrable hierarchies, associated conservation laws and Darboux transformation. Journal of Computational and Applied Mathematics, 233(4):1096– 1107, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006311.

Liu:2009:DBS

[LZZD09] Chao Liu, Qingling Zhang, Xue Zhang, and Xiaodong Duan. Dynamical behavior in a stage-structured differential-algebraic preypredator model with discrete time delay and harvesting. Journal of Computational and Applied Mathematics, 231(2):612–625, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002532.

Ma:2006:FEM

[Ma06]  Changfeng Ma.  Finite-

element method for timedependent Maxwell’s equations based on an explicitmagnetic-field scheme. Journal of Computational and Applied Mathematics, 194(2): 409–424, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005121.

Ma:2008:SNN

[Ma08a] Qing-Hua Ma. Some new nonlinear Volterra–Fredholmtype discrete inequalities and their applications. Journal of Computational and Applied Mathematics, 216 (2):451–466, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002853.

Mesloub:2008:CCN

[MA08b] S. Mesloub and H. A. Abdusalam. On a certain class of a nonlinear evolution partial differential equations. Journal of Computational and Applied Mathematics, 217(1):148–156, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003445.

Muratova:2009:MMS

[MA09] Galina V. Muratova and Evgeniya M. Andreeva. Multigrid method for solving convection–diffusion problems with dominant convection. Journal of Computational and Applied Mathematics, 226(1):77–83, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002239.

Magnus:2009:RIS

[Mag09] Alphonse P. Magnus. Rational interpolation to solutions of Riccati difference equations on elliptic lattices. Journal of Computational and Applied Mathematics, 233(3): 793–801, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001277.

Mainge:2008:CTI

[Mai08a] Paul-Emile Maingé. Convergence theorems for inertial KM-type algorithms. Journal of Computational and Applied Mathematics, 219(1): 223–236, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003901.

Mainge:2008:DCP

[Mai08b] Paul-Emile Maingé. Discretization of the Cauchy problem for a fast diffusion equation. Journal of Computational and Applied Mathematics, 213(1):95–110, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000088.

Mao:2007:ESE

[Mao07] Xuerong Mao. Exponential stability of equidistant Euler–Maruyama approximations of stochastic differential delay equations. Journal of Computational and Applied Mathematics, 200 (1):297–316, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000173.

Marciniak:2007:MIM

[Mar07a] Andrzej Marciniak. On multistep interval methods for solving the initial value problem. Journal of Computational and Applied Mathematics, 199(2):229–237, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007557.

Markot:2007:IMV

[Mar07b] Mihály Csaba Markót. Interval methods for verifying structural optimality of circle packing configurations in the unit square. Journal of Computational and Applied Mathematics, 199(2): 353–357, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007715.

Martin:2007:MSS

[Mar07c] P. A. Martin. On the T - matrix for scattering by small obstacles. Journal of Computational and Applied Mathematics, 204(2):219–230, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003529.

Marcozzi:2008:SOC

[Mar08] Michael D. Marcozzi. Stochastic optimal control of ultradiffusion processes with application to dynamic portfolio management. Journal of Computational and Applied Mathematics, 222(1): 112–127, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005523.

Marcozzi:2009:EDG

[Mar09] Michael D. Marcozzi. Extrapolation discontinuous Galerkin method for ultraparabolic equations. Journal of Computational and Applied Mathematics, 224(2): 679–687, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002938.

Mathar:2006:CSE

[Mat06] Richard J. Mathar. Chebyshev series expansion of inverse polynomials. Journal of Computational and Applied Mathematics, 196(2): 596–607, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006230.

Matos:2007:MFP

[Mat07a] Ana C. Matos. Multivariate Frobenius–Padé approximants: Properties and algorithms. Journal of Computational and Applied Mathematics, 202(2):548–572, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001610.

Matsuo:2007:NCS

[Mat07b] Takayasu Matsuo. New conservative schemes with discrete variational derivatives for nonlinear wave equations. Journal of Computational and Applied Mathematics, 203(1):32–56, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001634.

Matsuo:2008:DCG

[Mat08]  Takayasu Matsuo.  Dissi-

pative/conservative Galerkin method using discrete partial derivatives for nonlinear evolution equations. Journal of Computational and Applied Mathematics, 218(2):

506–521, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004293.

Mayer:2007:RSI

[May07] Günter Mayer. On regular and singular interval systems. Journal of Computational and Applied Mathematics, 199(2):220–228, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007545.

Mayo:2008:MRS

[May08] Anita Mayo. Methods for the rapid solution of the pricing PIDEs in exponential and Merton models. Journal of Computational and Applied Mathematics, 222(1): 128–143, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005535.

Mazya:2006:CIC

[Maz06] Vladimir Maz’ya. Conductor inequalities and criteria for Sobolev type twoweight imbeddings. Journal of Computational and Applied Mathematics, 194(1): 94–114, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500405X.

Mazure:2008:URR

[Maz08] Marie-Laurence Mazure. Understanding recurrence relations for Chebyshevian Bsplines via blossoms. Journal of Computational and Applied Mathematics, 219(2): 457–470, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003470.

Moreno-Balcazar:2005:SZS

[MB05] Juan José Moreno-Balcázar. Smallest zeros of some types of orthogonal polynomials: asymptotics. Journal of Computational and Applied Mathematics, 179(1–2):289– 301, July 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004546.

Maes:2006:HRB

[MB06] Jan Maes and Adhemar Bultheel. C^{1} hierarchical Riesz bases of Lagrange type on Powell–Sabin triangulations. Journal of Computational and Applied Mathematics, 196(1):1–19, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005145.

Mandal:2007:ASC

[MB07a] B. N. Mandal and G. H. Bera.

Approximate solution of a class of singular integral equations of second kind. Journal of Computational and Applied Mathematics, 206(1): 189–195, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004249.

Moreno-Balcazar:2007:NZF

[MB07b] Juan J. Moreno-Balcázar. A note on the zeros of Freud– Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 207(2):338–344, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006078.

McMillen:2009:ZCV

[MBA09] T. McMillen, A. Bourget, and A. Agnew. On the zeros of complex Van vleck polynomials. Journal of Computational and Applied Mathematics, 223(2):862–871, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000897.

Martinez:2009:MPE

[MBCV09] A. Martínez, L. Bergamaschi, M. Caliari, and M. Vianello. A massively parallel exponential integrator for advection– diffusion models. Journal of Computational and Applied Mathematics, 231(1): 82–91, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000429.

Milovanovic:2005:GLI

[MC05a] Gradimir V. Milovanović and Aleksandar S. Cvetković. Gauss–Laguerre interval quadrature rule. Journal of Computational and Applied Mathematics, 182(2):433–446, October 15, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006399.

Milovanovic:2005:OPG

[MC05b] Gradimir V. Milovanović and Aleksandar S. Cvetković. Orthogonal polynomials and Gaussian quadrature rules related to oscillatory weight functions. Journal of Computational and Applied Mathematics, 179(1–2):263–287, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042704004534.

Milovanovic:2005:ROS

[MC05c] Gradimir V. Milovanović and Aleksandar S. Cvetković. Remarks on “Orthogonality of some sequences of the rational functions and Müntz polynomials”. Journal of Computational and Applied Mathematics, 173(2):383–388, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002262.

See [MDSR04].

Ma:2007:SNN

[MC07] Qing-Hua Ma and Wing-Sum Cheung. Some new nonlinear difference inequalities and their applications. Journal of Computational and Applied Mathematics, 202 (2):339–351, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001440.

Ma:2008:COS

[MC08] Changfeng Ma and Xiaohong Chen. The convergence of a one-step smoothing Newton method for P_{0}–NCP based on a new smoothing NCPfunction. Journal of Computational and Applied Mathematics, 216(1):1–13, June 15, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001781.

McBain:2009:NSO

[MCA09] G. D. McBain, T. H. Chubb, and S. W. Armfield. Numerical solution of the Orr– Sommerfeld equation using the viscous Green function and split-Gaussian quadrature. Journal of Computational and Applied Mathematics, 224(1):397–404, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800246X.

McCall:2005:GAM

[McC05] John McCall. Genetic algorithms for modelling and optimisation. Journal of Computational and Applied Mathematics, 184(1):205– 222, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000774.

Mustafa:2006:EEB

[MCD06] Ghulam Mustafa, Falai Chen, and Jiansong Deng. Estimating error bounds for binary subdivision curves/surfaces. Journal of Computational and Applied Mathematics, 193(2):596–613, September 1, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004346.

Marchandise:2008:SSS

[MCR08] Emilie Marchandise, Nicolas Chevaugeon, and Jean-François Remacle. Spatial and spectral superconvergence of discontinuous Galerkin method for hyperbolic problems. Journal of Computational and Applied Mathematics, 215 (2):484–494, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600759X.

Masoumi:2008:NMF

[MD08] H. R. Masoumi and G. Degrande. Numerical modeling of free field vibrations due to pile driving using a dynamic soil-structure interaction formulation. Journal of Computational and Applied Mathematics, 215(2):503–511, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007618.

Meiqiang:2009:MPS

[MD09a] Feng Meiqiang and Xie Dongxiu. Multiple positive solutions of multi-point boundary value problem for second-order impulsive differential equations. Journal of Computational and Applied Mathematics, 223(1): 438–448, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000344.

Mohebbi:2009:UCB

[MD09b] Akbar Mohebbi and Mehdi Dehghan. The use of compact boundary value method for the solution of two-dimensional Schrödinger equation. Journal of Computational and Applied Mathematics, 225(1):124–134, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003579.

Mendoza:2006:DAM

[MDB06] O. Barrero Mendoza, B. De Moor, and D. S. Bernstein. Data assimilation for magnetohydrodynamics systems. Journal of Computational and Applied Mathematics, 189(1–2):242–259, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001378.

Macias-Diaz:2009:NTS

[MDMRP09] J. E. Macías-Díaz, I. E. Medina-Ramírez, and A. Puri. Numerical treatment of the spherically symmetric solutions of a generalized Fisher– Kolmogorov–Petrovsky–Piscounov equation. Journal of Computational and Applied Mathematics, 231(2):851–868, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709003045.

Macias-Diaz:2008:EBC

[MDP08] J. E. Macías-Díaz and A. Puri.

An energy-based computational method in the analysis of the transmission of energy in a chain of coupled oscillators. Journal of Computational and Applied Mathematics, 214(2):393–405, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001355.

Marinkovic:2004:OSS

[MDSR04] S. D. Marinković, B. Danković, M. S. Stanković, and P. M. Rajković. Orthogonality of some sequences of the rational functions and the Müntz polynomials. Journal of Computational and Applied Mathematics, 163(2): 419–427, February 15, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/

                                        science/article/pii/S0377042703007647.

See remarks [MC05c].

Mai-Duy:2007:SCM

[MDT07] N. Mai-Duy and R. I. Tanner. A spectral collocation method based on integrated Chebyshev polynomials for two-dimensional biharmonic boundary-value problems. Journal of Computational and Applied Mathematics, 201(1):30–47, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000768.

Manousopoulos:2009:PIF

[MDT09] Polychronis Manousopoulos, Vassileios Drakopoulos, and Theoharis Theoharis. Parameter identification of 1D fractal interpolation functions using bounding volumes. Journal of Computational and Applied Mathematics, 233 (4):1063–1082, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709006190.

Mehr:2005:MAM

[Meh05]  Ramit Mehr.  Modeling

and analysis of the metapopulation dynamics of lymphocyte repertoires. Journal of Computational and Applied Mathematics, 184(1): 223–241, December 1, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000786.

Menendez:2009:RSC

[Men09] M. L. Menéndez. Rao’s statistic for constant and proportional hazard models. Journal of Computational and Applied Mathematics, 224(1): 118–126, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001817.

Merrill:2005:SDE

[Mer05] Stephen J. Merrill. The stochastic dance of early HIV infection. Journal of Computational and Applied Mathematics, 184(1): 242–257, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000798.

Melicher:2008:NSO

[MG08] V. Melicher and V. Gajdosík. A numerical solution of a one-dimensional blood flow model-moving grid approach. Journal of Computational and Applied Mathematics, 215(2):512–520, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600762X.

Matias:2006:MDC

[MGMTO06] J. M. Matías, W. González-Manteiga, J. Taboada, and C. Ordóñez. Managing distribution changes in time series prediction. Journal of Computational and Applied Mathematics, 191(2):206–215, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005492.

McNally:2008:CLP

[MGS08] Jeffrey M. McNally, L. E. Garey, and R. E. Shaw. A communication-less parallel algorithm for tridiagonal Toeplitz systems. Journal of Computational and Applied Mathematics, 212 (2):260–271, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600728X.

Mainardi:2007:BPR

[MGV07] Francesco Mainardi, Rudolf Gorenflo, and Alessandro Vivoli. Beyond the Poisson renewal process: a tutorial survey. Journal of Computational and Applied Mathematics, 205(2):725–735, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003967.

Marcellan:2008:GST

[MH08] F. Marcellán and J. Hernández.

Geronimus spectral transforms and measures on the complex plane. Journal of Computational and Applied Mathematics, 219(2): 441–456, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003196.

Muroya:2009:SCD

[MI09] Yoshiaki Muroya and Emiko Ishiwata. Stability for a class of difference equations. Journal of Computational and Applied Mathematics, 228 (2):561–570, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001386.

Miller:2005:NKM

[Mil05]  Keith Miller.  Nonlinear

Krylov and moving nodes in the method of lines. Journal of Computational and Applied Mathematics, 183(2): 275–287, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000658.

Milgram:2006:CNT

[Mil06] M. Milgram. Comment on “A new two-term relation for the _{3}F_{2} hypergeometric function of unit argument”. Journal of Computational and Applied Mathematics, 196(2): 437–438, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005868.

See [Ext99].

Milgram:2007:CPR

[Mil07] Michael Milgram. Comment on a paper of Rao et al., an entry of Ramanujan and a new _{3}F_{2}(1). Journal of Computational and Applied Mathematics, 201(1):1–2, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000719.

See [RVK05].

Miller:2009:CST

[Mil09] Allen R. Miller. Certain summation and transformation formulas for generalized hypergeometric series. Journal of Computational and Applied Mathematics, 231(2): 964–972, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003252.

Minchev:2005:SNP

[Min05] Emil Minchev. On a system of nonlinear PDE’s for phase transitions with vector order parameter and convective effect. Journal of Computational and Applied Mathematics, 177(2):309–330, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004285.

Meng:2007:SNN

[MJ07] Fan Wei Meng and Dehong Ji. On some new nonlinear discrete inequalities and their applications. Journal of Computational and Applied Mathematics, 208(2): 425–433, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006145.

Masjed-Jamei:2008:BCS

[MJ08a] Mohammad Masjed-Jamei.

A basic class of symmetric orthogonal functions using the extended Sturm–Liouville theorem for symmetric functions. Journal of Computational and Applied Mathematics, 216(1):128–143, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002233.

Masjed-Jamei:2008:CNI

[MJ08b] Mohammad Masjed-Jamei. On constructing new interpolation formulas using linear operators and an operator type of quadrature rules. Journal of Computational and Applied Mathematics, 216(2):307–318, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002488.

Ma:2009:GMM

[MJ09] Jingtang Ma and Yingjun Jiang. On a graded mesh method for a class of weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 231(2): 807–814, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003008.

Markovic:2009:NWL

[MJB09] Darija Marković, Dragan Jukić, and Mirta Bensić. Nonlinear weighted least squares estimation of a threeparameter Weibull density with a nonparametric start. Journal of Computational and Applied Mathematics, 228(1):304–312, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004883.

Ma:2009:NSB

[MJX09] Jingtang Ma, Yingjun Jiang, and Kaili Xiang. Numerical simulation of blowup in nonlocal reaction–diffusion equations using a moving mesh method. Journal of Computational and Applied Mathematics, 230(1):8–21, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005827.

Muroya:2005:GLC

[MK05] Yoshiaki Muroya and Yoshiko Kato. On gopalsamy and Liu’s conjecture for global stability in a population model. Journal of Computational and Applied Mathematics, 181(1):70–82, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005503.

Miyajima:2007:ETS

[MK07] Shinya Miyajima and Masahide Kashiwagi. Existence test for solution of nonlinear systems applying affine arithmetic. Journal of Computational and Applied Mathematics, 199(2):304–309, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500765X.

Margenov:2009:MPF

[MK09] Svetozar Margenov and Nikola Kosturski. MIC(0) preconditioning of 3D FEM problems on unstructured grids: Conforming and non-conforming elements. Journal of Computational and Applied Mathematics, 226(2):288–297, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004081.

Meng:2009:PSL

[ML09] Hua Meng and Fei Long. Periodic solutions for a Liénard type p-Laplacian differential equation. Journal of Computational and Applied Mathematics, 224(2): 696–701, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002951.

Meng:2008:PAP

[MLH08] Yimin Meng, Bingwen Liu, and Lihong Huang. Positive almost periodic solutions for a class of Liénardtype systems with multiple deviating arguments. Journal of Computational and Applied Mathematics, 220(1–2): 615–623, October 15, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004906.

Mai:2009:SFW

[MLL09] Huanhuan Mai, Xiaofeng Liao, and Chuandong Li. A semi-free weighting matrices approach for neutral-type delayed neural networks. Journal of Computational and Applied Mathematics, 225 (1):44–55, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800318X.

Montagner:2006:SFC

[MLOP06] V. F. Montagner, V. J. S.

Leite, R. C. L. F. Oliveira, and P. L. D. Peres. State feedback control of switched linear systems: an LMI approach. Journal of Computational and Applied Mathematics, 194(2):192–206, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004486.

Mo:2007:NTR

[MLY07] Jiangtao Mo, Chunyan Liu, and Shicui Yan. A nonmonotone trust region method based on nonincreasing technique of weighted average of the successive function values. Journal of Computational and Applied Mathematics, 209(1):97–108, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006674.

Mondal:2006:DSP

[MM06] Kajal Kumar Mondal and B. S. Mazumder. On dispersion of settling particles from an elevated source in an open-channel flow. Journal of Computational and Applied Mathematics, 193 (1):22–37, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003778.

Mazya:2007:UAF

[MM07] V. Maz’ya and A. Movchan. Uniform asymptotic formulae for Green’s functions in singularly perturbed domains. Journal of Computational and Applied Mathematics, 208(1):194–206, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006352.

Mursaleen:2009:LSC

[MM09] M. Mursaleen and S. A. Mohiuddine. On lacunary statistical convergence with respect to the intuitionistic fuzzy normed space. Journal of Computational and Applied Mathematics, 233(2): 142–149, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004026.

Marcellan:2007:EMZ

[MMFMG07] F. Marcellán, A. Martínez-Finkelshtein, and P. Martínez-González. Electrostatic models for zeros of polynomials: Old, new, and some open problems. Journal of Computational and Applied Mathematics, 207(2): 258–272, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600611X.

Malek:2008:MCF

[MMM08] Alaeddin Malek and Sayed Hodjatollah Momeni-Masuleh. A mixed collocation-finite difference method for 3D microscopic heat transport problems. Journal of Computational and Applied Mathematics, 217(1):137–147, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003433.

Milagre:2009:MCS

[MMS^{+}09] S. T. Milagre, C. D. Maciel, A. A. Shinoda, M. Hungria, and J. R. B. Almeida. Multidimensional cluster stability analysis from a Brazilian Bradyrhizobium sp. RFLP/ PCR data set. Journal of Computational and Applied Mathematics, 227(2): 308–319, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001131.

Mehri:2005:EPS

[MN05] B. Mehri and M. A. Niksirat. On the existence of periodic solutions for certain differential equations. Journal of Computational and Applied Mathematics, 174(2): 239–249, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400192X.

Minamoto:2007:NMV

[MN07a] Teruya Minamoto and Mitsuhiro T. Nakao. Numerical method for verifying the existence and local uniqueness of a double turning point for a radially symmetric solution of the perturbed Gelfand equation. Journal of Computational and Applied Mathematics, 202(2):177–185, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001208.

Moriya:2007:NSC

[MN07b] K. Moriya and T. Nodera. A new scheme of computing the approximate inverse preconditioner for the reduced linear systems. Journal of Computational and Applied Mathematics, 199(2):345–352, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007703.

Malihoutsaki:2009:INM

[MNG09] E. N. Malihoutsaki, I. A.

Nikas, and T. N. Grapsa. Improved Newton’s method without direct function evaluations. Journal of Computational and Applied Mathematics, 227(1):206–212, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003531.

Molliq:2009:ASF

[MNHA09] R. Yulita Molliq, M. S. M.

Noorani, I. Hashim, and R. R. Ahmad. Approximate solutions of fractional Zakharov– Kuznetsov equations by VIM. Journal of Computational and Applied Mathematics, 233(2):103–108, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002180.

Muhammad:2005:NSI

[MNMS05] Mayinur Muhammad, Ahniyaz Nurmuhammad, Masatake Mori, and Masaaki Sugihara. Numerical solution of integral equations by means of the sinc collocation method based on the double exponential transformation. Journal of Computational and Applied Mathematics, 177 (2):269–286, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400425X.

Mantri:2008:QMB

[MNP08] Pritam S. Mantri, Neela Nataraj, and Amiya K. Pani. A qualocation method for Burgers’ equation. Journal of Computational and Applied Mathematics, 213 (1):1–13, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000027.

Malolepszy:2007:CBS

[MO07a] Tomasz Malolepszy and Wojciech Okrasiński. Conditions for blow-up of solutions of some nonlinear Volterra integral equations. Journal of Computational and Applied Mathematics, 205(2):

744–750, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003980.

Momani:2007:NAD

[MO07b] Shaher Momani and Zaid Odibat. Numerical approach to differential equations of fractional order. Journal of Computational and Applied Mathematics, 207(1): 96–110, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004651.

Momani:2008:NMN

[MO08] Shaher Momani and Zaid Odibat. A novel method for nonlinear fractional partial differential equations: Combination of DTM and generalized Taylor’s formula. Journal of Computational and Applied Mathematics, 220(1– 2):85–95, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004189.

Mendoza:2009:ZLA

[MOC09] C. Díaz Mendoza, R. Orive, and H. Pijeira Cabrera. Zeros and logarithmic asymptotics of Sobolev orthogonal polynomials for exponential weights.

Journal of Computational and Applied Mathematics, 233(3):691–698, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001137.

Mohanty:2005:FVM

[Moh05] R. K. Mohanty. A family of variable mesh methods for the estimates of (d u/d r) and solution of non-linear two point boundary value problems with singularity. Journal of Computational and Applied Mathematics, 182(1): 173–187, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006090.

Mohamad:2007:GES

[Moh07] Sannay Mohamad. Global exponential stability in DC-NNs with distributed delays and unbounded activations. Journal of Computational and Applied Mathematics, 205(1):161–173, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002627.

Mohamad:2008:ESP

[Moh08] Sannay Mohamad. Exponential stability preservation in discrete-time analogues of artificial neural networks with distributed delays. Journal of Computational and Applied Mathematics, 215 (1):270–287, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002026.

Mohsen:2009:FTD

[Moh09] A. Mohsen. On the ijk forms of the truncated LU decomposition. Journal of Computational and Applied Mathematics, 233(2): 582–584, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004208.

Moir:2005:AMS

[Moi05] Nicolette Moir. ARK methods: some recent developments. Journal of Computational and Applied Mathematics, 175(1):101–111, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002420.

Moller:2007:SPE

[Möl07] H. Michael Möller. On square positive extensions and cubature formulas. Journal of Computational and Applied Mathematics, 199(1): 80–88, February 1, 2007.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007405.

Monegato:2007:SNP

[Mon07] G. Monegato. Some new problems in numerical integration. Journal of Computational and Applied Mathematics, 210(1–2):244–253, December 31, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006662.

Monegato:2009:DPA

[Mon09] G. Monegato. Definitions, properties and applications of finite-part integrals. Journal of Computational and Applied Mathematics, 229 (2):425–439, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001635.

Moraca:2007:UBI

[Mor07]  Nenad Morača.  Upper

bounds for the infinity norm of the inverse of SDD and S–SDD matrices. Journal of Computational and Applied Mathematics, 206(2): 666–678, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005139.

Motyl:2006:NMC

[Mot06]  Elzbieta Motyl.  A new

method of calculation of the pressure in the stationary Navier–Stokes equations. Journal of Computational and Applied Mathematics, 189(1–2):207–219, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003699.

Marin:2007:HFA

[MOZ07] A. M. Marín, R. D. Ortíz, and P. Zhevandrov. Highfrequency asymptotics of waves trapped by underwater ridges and submerged cylinders. Journal of Computational and Applied Mathematics, 204(2):356–362, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003633.

Mazzia:2006:TDM

[MP06a] Annamaria Mazzia and Mario Putti. Three-dimensional mixed finite element-finite volume approach for the solution of density-dependent flow in porous media. Journal of Computational and Applied Mathematics, 185(2): 347–359, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001196.

Mazzia:2006:P

[MP06b] Francesca Mazzia and Tiziano Politi. Preface. Journal of Computational and Applied Mathematics, 185(2): 191–195, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500107X.

Mainardi:2007:RFW

[MP07] Francesco Mainardi and Gianni Pagnini. The role of the Fox–Wright functions in fractional sub-diffusion of distributed order. Journal of Computational and Applied Mathematics, 207(2): 245–257, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006017.

Magnus:2008:FRC

[MP08a] Alphonse P. Magnus and Viviane Pierrard. Formulas for recurrence coefficients of orthogonal polynomials related to Lorentzian-like weights. Journal of Computational and Applied Mathematics, 219(2):431–440, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002543.

Maruster:2008:MTI

[MP08b] Stefan Maruster and Cristina Popirlan. On the Manntype iteration and the convex feasibility problem. Journal of Computational and Applied Mathematics, 212 (2):390–396, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007382.

Menendez:2007:MCP

[MPP07] M. L. Menéndez, J. A. Pardo, and L. Pardo. The momentcorrected phi-divergence test statistics for symmetry. Journal of Computational and Applied Mathematics, 202 (2):315–327, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001427.

Menendez:2009:CSC

[MPP09] M. L. Menéndez, L. Pardo, and M. C. Pardo. Confidence sets and coverage probabilities based on preliminary estimators in logistic regression models. Journal of Computational and Applied Mathematics, 224(1):193–203, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001891.

Mishra:2007:GIN

[MPR07] S. K. Mishra, R. P. Pant, and J. S. Rautela. Generalized α-invexity and nondifferentiable minimax fractional programming. Journal of Computational and Applied Mathematics, 206(1): 122–135, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003864.

Mainardi:2005:FFF

[MPS05] Francesco Mainardi, Gianni Pagnini, and R. K. Saxena. Fox H functions in fractional diffusion. Journal of Computational and Applied Mathematics, 178(1– 2):321–331, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003826.

Man:2006:GAF

[MPXK06] Chi-Sing Man, Roberto Paroni, Yu Xiang, and Edward A. Kenik. On the geometric autocorrelation function of polycrystalline materials. Journal of Computational and Applied Mathematics, 190(1–2):200–210, June 1, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002311.

Menendez:2009:OSE

[MPZ09] M. L. Menéndez, L. Pardo, and K. Zografos. Ordering and selecting extreme populations by means of entropies and divergences. Journal of Computational and Applied Mathematics, 232(2): 335–350, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003616.

Munoz:2009:NIM

[MR09] J. F. Muñoz and M. Rueda.

New imputation methods for missing data using quantiles. Journal of Computational and Applied Mathematics, 232(2):305–317, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003574.

Morigi:2006:IMI

[MRSZ06] S. Morigi, L. Reichel, F. Sgallari, and F. Zama. Iterative methods for ill-posed problems and semiconvergent sequences. Journal of Computational and Applied Mathematics, 193(1):157–167, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003857.

Morigi:2007:IML

[MRSZ07] S. Morigi, L. Reichel, F. Sgallari, and F. Zama. An iterative method for linear discrete ill-posed problems with box constraints. Journal of Computational and Applied Mathematics, 198(2): 505–520, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007302.

Messina:2009:CTE

[MRV09] E. Messina, E. Russo, and A. Vecchio. A convolution test equation for double delay integral equations. Journal of Computational and Applied Mathematics, 228 (2):589–599, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001404.

Milovanovic:2005:BEG

[MS05a] Gradimir V. Milovanović and Miodrag M. Spalević. Bounds of the error of Gauss–Turántype quadratures. Journal of Computational and Applied Mathematics, 178(1– 2):333–346, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003838.

Muller:2005:HAN

[MS05b] Markus Müller and Dierk Schleicher. How to add a non-integer number of terms, and how to produce unusual infinite summations. Journal of Computational and Applied Mathematics, 178(1– 2):347–360, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003851.

Milovanovic:2006:QRM

[MS06a] Gradimir V. Milovanović and Miodrag M. Spalević. Quadrature rules with multiple nodes for evaluating integrals with strong singularities. Journal of Computational and Applied Mathematics, 189(1–2):689–702, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003705.

Murat:2006:SPA

[MS06b] M. Murat and D. Szynal.

On some properties of the “apparent length” distribution and a compound distribution. Journal of Computational and Applied Mathematics, 186(1):43–63, February 1, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001883.

Marcellan:2007:SSR

[MS07a] Francisco Marcellán and Ridha Sfaxi. Second structure relation for semiclassical orthogonal polynomials. Journal of Computational and Applied Mathematics, 200(2):537–554, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000343.

Milovanovic:2007:NBE

[MS07b] Gradimir V. Milovanović and Miodrag M. Spalević. A note on the bounds of the error of Gauss–Turán-type quadratures. Journal of Computational and Applied Mathematics, 200(1):276–282, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705008125.

Mohanty:2007:NTL

[MS07c] R. K. Mohanty and Swarn Singh. A new two-level implicit discretization of O(k^{2} + kh^{2} + h^{4}) for the solution of singularly perturbed twospace dimensional non-linear parabolic equations. Journal of Computational and Applied Mathematics, 208(2):

391–403, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006121.

Monaquel:2007:FOT

[MS07d] Shatha Jameel Monaquel and Karl Michael Schmidt. On Mfunctions and operator theory for non-self-adjoint discrete Hamiltonian systems. Journal of Computational and Applied Mathematics, 208(1): 82–101, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006285.

Ma:2008:SCI

[MS08a] Yajun Ma and Jitao Sun. Stability criteria for impulsive systems on time scales. Journal of Computational and Applied Mathematics, 213 (2):400–407, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000398.

Monegato:2008:NET

[MS08b] Giovanni Monegato and Antonio Strozzi. The numerical evaluation of two integral transforms. Journal of Computational and Applied Mathematics, 211(2): 173–180, February 1, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007035.

Mansour:2009:KMS

[MS09a] Toufik Mansour and Chunwei Song. Kernel method and system of functional equations. Journal of Computational and Applied Mathematics, 224(1):133–139, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001830.

Mao:2009:EET

[MS09b] Shipeng Mao and Zhongci Shi. Error estimates of triangular finite elements under a weak angle condition. Journal of Computational and Applied Mathematics, 230(1): 329–331, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006006.

Muller:2009:MFV

[MS09c] Siegfried Müller and Youssef Stiriba. A multilevel finite volume method with multiscale-based grid adaptation for steady compressible flows. Journal of Computational and Applied Mathematics, 227(2):223–233, May 15, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001064.

Merat:2009:IDP

[MSA09] Kaveh Merat, Hassan Salarieh, and Aria Alasty. Implementation of dynamic programming for chaos control in discrete systems. Journal of Computational and Applied Mathematics, 233(2): 531–544, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004658.

Marchuk:2005:AEA

[MSB05] G. I. Marchuk, V. Shutyaev, and G. Bocharov. Adjoint equations and analysis of complex systems: Application to virus infection modelling. Journal of Computational and Applied Mathematics, 184(1): 177–204, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000762.

Miellou:2008:NSC

[MSE08] J. C. Miellou, P. Spiteri, and D. El Baz. A new stopping criterion for linear perturbed asynchronous iterations. Journal of Computational and Applied Mathematics, 219(2):471–483, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005158.

Milovanovic:2008:RTG

[MSP08] Gradimir V. Milovanović, Miodrag M. Spalević, and Miroslav S. Pranić. On the remainder term of Gauss– Radau quadratures for analytic functions. Journal of Computational and Applied Mathematics, 218(2): 281–289, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000556.

Milovanovic:2009:EEG

[MSP09] Gradimir V. Milovanović, Miodrag M. Spalević, and Miroslav S. Pranić. Error estimates for Gaussian quadratures of analytic functions. Journal of Computational and Applied Mathematics, 233(3):802–807, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001289.

Mazroui:2005:RCH

[MST05] A. Mazroui, D. Sbibih, and A. Tijini. A recursive construction of Hermite spline interpolants and applications.

Journal of Computational and Applied Mathematics, 183(1):67–83, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000038.

Maryska:2008:MPF

[MSTT08] Jirí Maryska, Otto Severýn, Miloslav Tauchman, and David Tondr. Modelling of processes in fractured rock using FEM/FVM on multidimensional domains. Journal of Computational and Applied Mathematics, 215 (2):495–502, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007606.

Manouzi:2007:WSF

[MSZ07] Hassan Manouzi, Mohammed Seaid̈, and Mostafa Zahri. Wick-stochastic finite element solution of reaction– diffusion problems. Journal of Computational and Applied Mathematics, 203 (2):516–532, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002317.

Mastroianni:2005:NMG

[MT05] G. Mastroianni and W. Themistoclakis. A numerical method for the generalized airfoil equation based on the de la Vallée Poussin interpolation. Journal of Computational and Applied Mathematics, 180(1):71–105, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004790.

Miyagawa:2008:FSE

[MU08] Yuu Miyagawa and Tsuyoshi Ueta. FEM for Schrödinger equation with Rashba effect. Journal of Computational and Applied Mathematics, 218(1):96–105, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003317.

Muhlbach:2006:EBS

[Müh06a] G. Mühlbach. ECT–B-splines defined by generalized divided differences. Journal of Computational and Applied Mathematics, 187 (1):96–122, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001494.

Muhlbach:2006:OSH

[Müh06b] G. Mühlbach. One sided Hermite interpolation by piecewise different generalized polynomials. Journal of Computational and Applied Mathematics, 196(1): 285–298, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005649.

Murthy:2007:ARA

[Mur07] A. S. Vasudeva Murthy.

An alternate relaxation approximation to conservation laws. Journal of Computational and Applied Mathematics, 203(2):437–443, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002287.

Mustapha:2009:NSS

[Mus09]  Kassem Mustapha.  Nu-

merical solution for a subdiffusion equation with a smooth kernel. Journal of Computational and Applied Mathematics, 231(2): 735–744, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002738.

Markovsky:2005:HPN

[MV05a] Ivan Markovsky and Sabine Van Huffel. High-performance numerical algorithms and software for structured total least squares. Journal of Computational and Applied Mathematics, 180(2): 311–331, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005254.

Matos:2005:SFP

[MV05b] Ana C. Matos and Jeannette Van Iseghem. Simultaneous Frobenius–Padé approximants. Journal of Computational and Applied Mathematics, 176(2):231–258, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003280.

Malengier:2008:PEC

[MV08] B. Malengier and Roger Van Keer. Parameter estimation in convection dominated nonlinear convection–diffusion problems by the relaxation method and the adjoint equation. Journal of Computational and Applied Mathematics, 215(2):477–483, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007588.

Misra:2006:ESP

[MvdMV06] Neeraj Misra, Edward C. van der Meulen, and Karlien Vanden Branden. On estimating the scale parameter of the selected gamma population under the scale invariant squared error loss function. Journal of Computational and Applied Mathematics, 186(1):268–282, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002001.

Mastronardi:2008:FAR

[MVV08] N. Mastronardi, M. Van Barel, and R. Vandebril. A fast algorithm for the recursive calculation of dominant singular subspaces. Journal of Computational and Applied Mathematics, 218(2): 238–246, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000106.

Martin-Vaquero:2006:EFB

[MVVA06] J. Martín-Vaquero and J. Vigo-Aguiar. Exponential fitting BDF algorithms: Explicit and implicit 0-stable methods. Journal of Computational and Applied Mathematics, 192(1):100–113, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003110.

Mastronardi:2006:CEC

[MVVV06] N. Mastronardi, M. Van Barel, E. Van Camp, and R. Vandebril. On computing the eigenvectors of a class of structured matrices. Journal of Computational and Applied Mathematics, 189(1–2):580–591, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001573.

Ma:2007:IAS

[MW07] Zhongtai Ma and Guochun Wen. Iterative approximation of solutions for semilinear parabolic equations system. Journal of Computational and Applied Mathematics, 209(2):167–175, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600673X.

Meek:2008:FBM

[MW08] D. S. Meek and D. J. Walton. The family of biarcs that matches planar, twopoint G^{1} Hermite data. Journal of Computational and Applied Mathematics, 212(1): 31–45, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007102.

Ma:2009:NLM

[MW09a] Cheng Ma and Changyu Wang. A nonsmooth Levenberg–Marquardt method for solving semi-infinite programming problems. Journal of Computational and Applied Mathematics, 230(2): 633–642, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000041.

Ma:2009:ADM

[MW09b] Limin Ma and Zongmin Wu.

Approximation to the k-th derivatives by multiquadric quasi-interpolation method. Journal of Computational and Applied Mathematics, 231(2):925–932, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003227.

Meek:2009:TPH

[MW09c] D. S. Meek and D. J. Walton. A two-point G^{1} Hermite interpolating family of spirals. Journal of Computational and Applied Mathematics, 223(1):97–113, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006711.

Meng:2006:BOS

[MY06] Qiong Meng and Jurang Yan. Bounded oscillation for second-order nonlinear delay differential equation in a critical state. Journal of Computational and Applied Mathematics, 197(1): 204–211, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006539.

Meng:2008:BOS

[MY08] Qiong Meng and Jurang Yan. Bounded oscillation for second-order nonlinear neutral difference equations in critical and non-critical states. Journal of Computational and Applied Mathematics, 211(2):156–172, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007023.

Mao:2005:NMS

[MYY05] Xuerong Mao, Chenggui Yuan, and G. Yin. Numerical method for stationary distribution of stochastic differential equations with Markovian switching. Journal of Computational and Applied Mathematics, 174(1):1–27, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001657.

Milleo:2006:MNS

[MYY06] Ingrid Milléo, J. Yin, and J. Y. Yuan. Minimization of ℓ_{2}-norms of the SOR and MSOR operators. Journal of Computational and Applied Mathematics, 192(2): 431–444, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003729.

Mao:2007:AEM

[MYY07] Xuerong Mao, Chenggui Yuan, and G. Yin. Approximations of Euler–Maruyama type for stochastic differential equations with Markovian switching, under non-Lipschitz conditions. Journal of Computational and Applied Mathematics, 205(2): 936–948, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004146.

Ma:2008:DDR

[MZC08] Shuping Ma, Chenghui Zhang, and Zhaolin Cheng. Delaydependent robust H_{∞} control for uncertain discretetime singular systems with time-delays. Journal of Computational and Applied Mathematics, 217(1):194–211, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003494.

Mo:2006:VSM

[MZW06] Jiangtao Mo, Kecun Zhang, and Zengxin Wei. A variant of SQP method for inequality constrained optimization and its global convergence. Journal of Computational and Applied Mathematics, 197(1): 270–281, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006606.

Nagy:2009:ABC

[Nag09] Bálint Nagy. Analysis of the biological clock of neurospora. Journal of Computational and Applied Mathematics, 226(2):298–305, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004093.

Ning:2007:DSB

[NaLH07] Di Ning, Jun an Lu, and Xiuping Han. Dual synchronization based on two different chaotic systems: Lorenz systems and Rössler systems. Journal of Computational and Applied Mathematics, 206(2):1046–1050, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005772.

Noor:2006:MPP

[NB06] Muhammad Aslam Noor and Abdellah Bnouhachem. Modified proximal-point method for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 197(2): 395–405, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006874.

Nguyen-Ba:2008:HBO

[NBSV08] Truong Nguyen-Ba, Philip W.

Sharp, and Rémi Vaillancourt. Hermite–Birkhoff– Obrechkoff four-stage fourstep ODE solver of order 14 with quantized step size. Journal of Computational and Applied Mathematics, 222(2):608–621, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006358.

Nedzhibov:2008:FMP

[Ned08] Gyurhan H. Nedzhibov. A family of multi-point iterative methods for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 222(2): 244–250, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005924.

Neumaier:2008:CEB

[Neu08] Arnold Neumaier. Certified error bounds for uncertain elliptic equations. Journal of Computational and Applied Mathematics, 218(1): 125–136, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003342.

Nowak:2005:PEA

[NGS05] U. Nowak, A. Grah, and M. Schreier. Parameter estimation and accuracy matching strategies for 2-D reactor models. Journal of Computational and Applied Mathematics, 183(2):301– 311, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000671.

Nakao:2008:GEB

[NH08] Mitsuhiro T. Nakao and Kouji Hashimoto. Guaranteed error bounds for finite element approximations of noncoercive elliptic problems and their applications. Journal of Computational and Applied Mathematics, 218(1): 106–115, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003329.

Nagatou:2007:NVS

[NHN07] K. Nagatou, K. Hashimoto, and M. T. Nakao. Numerical verification of stationary solutions for Navier–Stokes problems. Journal of Computational and Applied Mathematics, 199(2):445–451, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007855.

Nigsch:2007:NST

[Nig07] Martin Nigsch. Numerical studies of time-independent and time-dependent scattering by several elliptical cylinders. Journal of Computational and Applied Mathematics, 204(2):231–241, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003530.

Nadarajah:2008:PPT

[NK08] Saralees Nadarajah and Samuel Kotz. A product Pearson-type VII density distribution. Journal of Computational and Applied Mathematics, 211(1): 103–113, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006753.

Niki:2010:CCT

[NK10] Hiroshi Niki and Toshiyuki Kohno. Comment on ‘A comparison theorem of the SOR iterative method’. Journal of Computational and Applied Mathematics, 234(12): 3507–3510, October 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042710002037.

See [yS05].

Niki:2009:EGM

[NKA09] Hiroshi Niki, Toshiyuki Kohno, and Kuniyoshi Abe. An extended GS method for dense linear systems. Journal of Computational and Applied Mathematics, 231(1): 177–186, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900051X.

Nkemzi:2005:PEA

[Nke05] Boniface Nkemzi. The Poisson equation in axisymmetric domains with conical points. Journal of Computational and Applied Mathematics, 174(2):399–421, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002249.

Niki:2008:PGS

[NKM08] Hiroshi Niki, Toshiyuki Kohno, and Munenori Morimoto. The preconditioned Gauss–Seidel method faster than the SOR method. Journal of Computational and Applied Mathematics, 219(1):59–71, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003639.

See erratum [LC09b].

NGuessan:2005:SCA

[NL05] Assi N’Guessan and Claude Langrand. A Schur complement approach for computing subcovariance matrices arising in a road safety measure modelling. Journal of Computational and Applied Mathematics, 177(2):331–345, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004297.

Niksirat:2006:EPS

[NM06] M. A. Niksirat and B. Mehri.

On the existence of positive solution for secondorder multi-points boundary value problems. Journal of Computational and Applied Mathematics, 193(1): 269–276, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003912.

Nassif:2009:CBS

[NMKS09] Nabil R. Nassif, Noha Makhoul-Karam, and Yeran Soukiassian. Computation of blowing-up solutions for second-order differential equations using re-scaling techniques. Journal of Computational and Applied Mathematics, 227(1):185–195, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003518.

Nurmuhammad:2007:SGM

[NMM07] Ahniyaz Nurmuhammad, Mayinur Muhammad, and Masatake Mori. Sinc–Galerkin method based on the DE transformation for the boundary value problem of fourthorder ODE. Journal of Computational and Applied Mathematics, 206(1):17–26, September 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003475.

Nurmuhammad:2005:DET

[NMMS05] Ahniyaz Nurmuhammad, Mayinur Muhammad, Masatake Mori, and Masaaki Sugihara. Double exponential transformation in the sinc-collocation method for a boundary value problem with fourth-order ordinary differential equation. Journal of Computational and Applied Mathematics, 182(1):32–50, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005990.

Noor:2005:HLI

[Noo05] Muhammad Aslam Noor.

Hemivariational-like inequalities. Journal of Computational and Applied Mathematics, 182(2):316–326, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006211.

Noumi:2007:SFA

[Nou07] Masatoshi Noumi. Special functions arising from discrete Painlevé equations: a survey. Journal of Computational and Applied Mathematics, 202(1):48–55, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001142.

Novati:2008:CRD

[Nov08]  Paolo Novati.  On the

construction of restricteddenominator exponential W - methods. Journal of Computational and Applied Mathematics, 212(1):86–101, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042706007175.

Noschese:2006:ECN

[NP06] S. Noschese and L. Pasquini. Eigenvalue condition numbers: Zero-structured versus traditional. Journal of Computational and Applied Mathematics, 185(1): 174–189, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000567.

Noschese:2007:EPC

[NP07] S. Noschese and L. Pasquini. Eigenvalue patterned condition numbers: Toeplitz and Hankel cases. Journal of Computational and Applied Mathematics, 206(2): 615–624, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005085.

Nawratil:2008:SSF

[NP08] Georg Nawratil and Helmut Pottmann. Subdivision schemes for the fair discretization of the spherical motion group. Journal of Computational and Applied Mathematics, 222(2): 574–591, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006334.

Nie:2009:ESP

[NPTH09] Linfei Nie, Jigen Peng, Zhidong Teng, and Lin Hu. Existence and stability of periodic solution of a Lotka–Volterra predator–prey model with state dependent impulsive effects. Journal of Computational and Applied Mathematics, 224(2):544–555, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002616.

See comment [TSC10].

Ning:2006:ICB

[NQDJ06] Bi Ning, Dai Qinyun, Huang Daren, and Fang Ji. Image coding based on multiband wavelet and adaptive quad-tree partition. Journal of Computational and Applied Mathematics, 195(1– 2):2–7, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500467X.

Naumov:2009:TBH

[NS09a] Maxim Naumov and Ahmed H.

Sameh. A tearing-based hybrid parallel banded linear system solver. Journal of Computational and Applied Mathematics, 226 (2):306–318, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800410X.

Nguyen:2009:FEW

[NS09b] Hoang Nguyen and Rob Stevenson. Finite element wavelets with improved quantitative properties. Journal of Computational and Applied Mathematics, 230(2): 706–727, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000107.

Nilsrakoo:2009:WCT

[NS09c] Weerayuth Nilsrakoo and Satit Saejung. Weak convergence theorems for a countable family of Lipschitzian mappings. Journal of Computational and Applied Mathematics, 230(2): 451–462, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006420.

Nguyen:2007:ATI

[NSC07] Hoang Si Nguyen, Roger B.

Sidje, and Nguyen Huu Cong. Analysis of trigonometric implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 198(1):187–207, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007867.

Nishida:2007:SMP

[NSK07] Tetsushi Nishida, Kokichi Sugihara, and Masato Kimura. Stable marker-particle method for the Voronoi diagram in a flow field. Journal of Computational and Applied Mathematics, 202(2):377–391, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001464.

Nanazawa:2009:MST

[NSK09] Youhei Nanazawa, Hiroshi Suito, and Hideo Kawarada. Mathematical study of tradeoff relations in logistics systems. Journal of Computational and Applied Mathematics, 232(1):122–126, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005517.

Njaastad:2007:DVS

[NSL07] O. Njåstad and J. C. Santos-León. Domain of validity of Szegő quadrature formulas. Journal of Computational and Applied Mathematics, 202(2):440–449, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001518.

Neuberger:2006:CEK

[NSS06] John M. Neuberger, Nándor Sieben, and James W. Swift. Computing eigenfunctions on the Koch snowflake: a new grid and symmetry. Journal of Computational and Applied Mathematics, 191 (1):126–142, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002669.

Niu:2007:VMB

[NSS07] Xiao-Ming Niu, Tetsuya Sakurai, and Hiroshi Sugiura. A verified method for bounding clusters of zeros of analytic functions. Journal of Computational and Applied Mathematics, 199(2): 263–270, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007600.

Nielsen:2008:PMN

[NST08] Bjørn Fredrik Nielsen, Ola Skavhaug, and Aslak Tveito.

Penalty methods for the numerical solution of American multi-asset option problems. Journal of Computational and Applied Mathematics, 222(1):3–16, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005456.

Noutsos:2006:OIC

[NT06] D. Noutsos and M. Tzoumas.

On optimal improvements of classical iterative schemes for Z-matrices. Journal of Computational and Applied Mathematics, 188(1):89–106, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001640.

Nedoma:2008:SHM

[NT08] Jirí Nedoma and Lubos Tomásek. On a solvability of hydro-mechanical problem based on contact problem with visco-plastic friction in Bingham rheology. Journal of Computational and Applied Mathematics, 218(1): 116–124, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003330.

Nesliturk:2006:FEM

[NTS06] A. I. Nesliturk and M. Tezer-Sezgin. Finite element method solution of electrically driven magnetohydrodynamic flow. Journal of Computational and Applied Mathematics, 192(2): 339–352, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003559.

Nerinckx:2008:TMU

[NVD08] Krista Nerinckx, Jan Vierendeels, and Erik Dick. Three Mach-uniform algorithms. Journal of Computational and Applied Mathematics, 215(2):521–527, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007631.

Nakamura:2008:NDS

[NWW08] Gen Nakamura, Shengzhang Wang, and Yanbo Wang. Numerical differentiation for the second order derivatives of functions of two variables. Journal of Computational and Applied Mathematics, 212(2):341–358, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007357.

Ogata:2006:FSMb

[OA06] Hidenori Ogata and Kaname Amano. A fundamental solution method for threedimensional viscous flow problems with obstacles in a periodic array. Journal of Computational and Applied Mathematics, 193(1): 302–318, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003936.

Ogata:2006:FSMa

[Oga06] Hidenori Ogata. A fundamental solution method for three-dimensional Stokes flow problems with obstacles in a planar periodic array. Journal of Computational and Applied Mathematics, 189(1–2):622–634, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003663.

Olyslager:2008:FMM

[OMC08] Femke Olyslager, Lieven Meert, and Kristof Cools. The fast multipole method in electromagnetics applied to the simulation of metamaterials. Journal of Computational and Applied Mathematics, 215(2):528–537, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007643.

Oliveira:2009:WQR

[OMV09] Saulo P. Oliveira, Alexandre L. Madureira, and Frederic Valentin. Weighted quadrature rules for finite element methods. Journal of Computational and Applied Mathematics, 227 (1):93–101, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003324.

Ozaki:2007:MOV

[OOM^{+}07] K. Ozaki, T. Ogita, S. Miyajima, S. Oishi, and S. M. Rump. A method of obtaining verified solutions for linear systems suited for Java. Journal of Computational and Applied Mathematics, 199(2): 337–344, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007697.

Ohmiya:2007:BFT

[OOO^{+}07] M. Ohmiya, H. Ohkura, D. Okaue, D. Saitoh, and T. Shiba. Benjamin–Feir type instability of sine-Gordon equation and spectrum of Lamé equation. Journal of Computational and Applied Mathematics, 207(2): 345–351, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005966.

Ooura:2008:ITQ

[Oou08] Takuya Ooura. An IMT-type quadrature formula with the same asymptotic performance as the DE formula. Journal of Computational and Applied Mathematics, 213(1): 232–239, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000301.

Olmos:2006:PMS

[OS06a] Daniel Olmos and Bernie D.

Shizgal. A pseudospectral method of solution of Fisher’s equation. Journal of Computational and Applied Mathematics, 193(1): 219–242, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003894.

Omelyan:2006:ICF

[OS06b] I. P. Omelyan and V. B.

Solovyan. Improved cubature formulae of high degrees of exactness for the square. Journal of Computational and Applied Mathematics, 188(2):190–204, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002165.

Obermaier:2007:NLL

[OS07a] Josef Obermaier and Ryszard Szwarc. Nonnegative linearization for little q-Laguerre polynomials and Faber basis. Journal of Computational and Applied Mathematics, 199(1):89–94, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007417.

Ohmori:2007:CFE

[OS07b] Katsushi Ohmori and Norikazu Saito. On the convergence of finite element solutions to the interface problem for the Stokes system. Journal of Computational and Applied Mathematics, 198(1): 116–128, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007090.

ORiordan:2007:TPP

[OS07c] E. O’Riordan and G. I. Shishkin. A technique to prove parameter-uniform convergence for a singularly perturbed convection– diffusion equation. Journal of Computational and Applied Mathematics, 206(1): 136–145, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003876.

Ohmori:2009:SRF

[OS09] Katsushi Ohmori and Norikazu Saito. Some remarks on the flux-free finite element method for immiscible twofluid flows. Journal of Computational and Applied Mathematics, 232(1):127– 138, October 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005529.

Osada:2006:AEC

[Osa06] Naoki Osada. Asymptotic error constants of cubically convergent zero finding methods. Journal of Computational and Applied Mathematics, 196(2):347–357, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005686.

Osada:2007:OPF

[Osa07] Naoki Osada. A one parameter family of locally quartically convergent zerofinding methods. Journal of Computational and Applied Mathematics, 205(1): 116–128, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002597.

Osada:2008:CHM

[Osa08] Naoki Osada. Chebyshev– Halley methods for analytic functions. Journal of Computational and Applied Mathematics, 216(2):585–599, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003093.

Oishi:2007:CRM

[OTOR07] Shin’ichi Oishi, Kunio Tanabe, Takeshi Ogita, and Siegfried M. Rump. Convergence of Rump’s method for inverting arbitrarily illconditioned matrices. Journal of Computational and Applied Mathematics, 205(1): 533–544, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003360.

OMalley:2006:DAE

[OW06a] Robert E. O’Malley, Jr. and David B. Williams. Deriving amplitude equations for weakly-nonlinear oscillators and their generalizations. Journal of Computational and Applied Mathematics, 190(1–2):3–21, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002189.

Ou:2006:MFF

[OW06b] Chunhua Ou and Michael J.

Ward. A metastable flamefront interface and Carrier’s problem. Journal of Computational and Applied Mathematics, 190(1– 2):354–375, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002402.

ORegan:2007:SWS

[OYA07] Donal O’Regan, Baoqiang Yan, and Ravi. P. Agarwal. Solutions in weighted spaces of singular boundary value problems on the halfline. Journal of Computational and Applied Mathematics, 205(2):751–763, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003992.

Ou:2009:OBT

[OZL09] Yigui Ou, Qian Zhou, and Haichan Lin. An ODEbased trust region method for unconstrained optimization problems. Journal of Computational and Applied Mathematics, 232(2): 318–326, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003598.

Ouyang:2005:OCO

[OZY05] Zigen Ouyang, Shengfan Zhou, and Fuqi Yin. Oscillation for a class of oddorder delay parabolic differential equations. Journal of Computational and Applied Mathematics, 175(2): 305–319, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002341.

Paris:2005:AAI

[Par05a] R. B. Paris. An asymptotic approximation for incomplete Gauss sums. Journal of Computational and Applied Mathematics, 180(2): 461–477, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400545X.

Paris:2005:KTT

[Par05b] R. B. Paris. A Kummer-type transformation for a _{2}F_{2} hypergeometric function. Journal of Computational and Applied Mathematics, 173(2): 379–382, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002237.

Park:2005:GHU

[Par05c] Chun-Gil Park. Generalized Hyers–Ulam–Rassias stability of n-sesquilinear-quadratic mappings on Banach modules over C^{∗}-algebras. Journal of Computational and Applied Mathematics, 180(2): 279–291, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005230.

Paris:2007:UHEb

[Par07a] R. B. Paris. On the use of Hadamard expansions in hyperasymptotic evaluation of Laplace-type integrals-III: Clusters of saddle points. Journal of Computational and Applied Mathematics, 207(2):273–290, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006066.

Paris:2007:UHEa

[Par07b]  R. B. Paris.  On the

use of Hadamard expansions in hyperasymptotic evaluation of Laplace-type integrals-IV: Poles. Journal of Computational and Applied Mathematics, 206(1): 454–472, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004997.

Paris:2008:AAI

[Par08a] R. B. Paris. An asymptotic approximation for incomplete Gauss sums. II. Journal of Computational and Applied Mathematics, 212(1): 16–30, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007096.

Park:2008:ACM

[Par08b] Ju H. Park. Adaptive control for modified projective synchronization of a fourdimensional chaotic system with uncertain parameters. Journal of Computational and Applied Mathematics, 213(1):288–293, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007242.

Paris:2009:ANE

[Par09a] R. B. Paris. The asymptotics of a new exponential sum. Journal of Computational and Applied Mathematics, 223(1):314–325, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000174.

Paris:2009:AGH

[Par09b] R. B. Paris. The asymptotics of the generalised Hermite– Bell polynomials. Journal of Computational and Applied Mathematics, 232(2): 216–226, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003458.

Paris:2009:HPE

[Par09c] R. B. Paris. High-precision evaluation of the Bessel functions via Hadamard series. Journal of Computational and Applied Mathematics, 224(1):84–100, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001799.

Pillards:2005:TLD

[PC05] Tim Pillards and Ronald Cools. Transforming lowdiscrepancy sequences from a cube to a simplex. Journal of Computational and Applied Mathematics, 174(1): 29–42, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001694.

Perrey-Debain:2006:PWD

[PD06]  E. Perrey-Debain.  Plane

wave decomposition in the unit disc: Convergence estimates and computational aspects. Journal of Computational and Applied Mathematics, 193(1):140–156, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003845.

Pang:2008:ASM

[PDM08] Sulin Pang, Feiqi Deng, and Xuerong Mao. Almost sure and moment exponential stability of Euler–Maruyama discretizations for hybrid stochastic differential equations. Journal of Computational and Applied Mathematics, 213(1):127–141, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000118.

Pedersen:2005:CPS

[Ped05] Henrik L. Pedersen. Canonical products of small order and related pick functions. Journal of Computational and Applied Mathematics, 175(2):245–263, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002316.

Pedersen:2009:LOT

[Ped09] Henrik L. Pedersen. Logarithmic order and type of indeterminate moment problems II. Journal of Computational and Applied Mathematics, 233(3):808–814, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001290.

Petersen:2005:ABS

[Pet05] Vigdis Brevik Petersen. Asymptotic behavior of Szegő polynomials. Journal of Computational and Applied Mathematics, 179(1–2):303–311, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004558.

Petras:2007:PVN

[Pet07a] Knut Petras. Principles of verified numerical integration. Journal of Computational and Applied Mathematics, 199(2):317–328, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007673.

Petronilho:2007:GFV

[Pet07b]  J. Petronilho.  Generic

formulas for the values at the singular points of some special monic classical H_{q,ω}-orthogonal polynomials. Journal of Computational and Applied Mathematics, 205(1):314–324, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003207.

Petronilho:2008:OPU

[Pet08] J. Petronilho. Orthogonal polynomials on the unit circle via a polynomial mapping on the real line. Journal of Computational and Applied Mathematics, 216(1):98–127, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002221.

Petkovic:2009:RHV

[Pet09a] M. S. Petković. A remark on He’s variational iteration technique for solving nonlinear equations. Journal of Computational and Applied Mathematics, 233(4):1187– 1189, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709005895.

See [He07].

Petkovic:2009:SVM

[Pet09b] Miodrag S. Petković. The self-validated method for polynomial zeros of high efficiency. Journal of Computational and Applied Mathematics, 233(4):1175–1186, December 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006384.

Parida:2007:RRN

[PG07] P. K. Parida and D. K.

Gupta. Recurrence relations for a Newton-like method in Banach spaces. Journal of Computational and Applied Mathematics, 206(2): 873–887, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005280.

Peng:2009:PSH

[PH09] Fengfu Peng and Xuli Han.

Parametric splines on a hyperbolic paraboloid. Journal of Computational and Applied Mathematics, 229 (1):183–191, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005554.

Pelovska:2006:NML

[PI06] Galena Pelovska and Mimmo Iannelli. Numerical methods for the Lotka–McKendrick’s equation. Journal of Computational and Applied Mathematics, 197(2):534–557, December 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006989.

Pierce:2008:RHP

[Pie08] Virgil U. Pierce. A Riemann– Hilbert problem for skeworthogonal polynomials. Journal of Computational and Applied Mathematics, 215 (1):230–241, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001999.

Pierantozzi:2009:FED

[Pie09] Teresa Pierantozzi. Fractional evolution Dirac-like equations: Some properties and a discrete von Neumanntype analysis. Journal of Computational and Applied Mathematics, 224(1): 284–295, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002082.

Petkovic:2007:PEB

[PIP07] Miodrag S. Petković, Snezana Ilić, and Ivan Petković. A posteriori error bound methods for the inclusion of polynomial zeros. Journal of Computational and Applied Mathematics, 208(2):

316–330, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005851.

Pironneau:2009:COR

[Pir09] Olivier Pironneau. Calibration of options on a reduced basis. Journal of Computational and Applied Mathematics, 232(1):139–147, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005530.

Pitselis:2009:SSB

[Pit09] Georgios Pitselis. Solvency supervision based on a total balance sheet approach. Journal of Computational and Applied Mathematics, 233(1): 83–96, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004336.

Pechstein:2006:MPI

[PJ06a] Clemens Pechstein and Bert Jüttler. Monotonicitypreserving interproximation of B- H-curves. Journal of Computational and Applied Mathematics, 196(1): 45–57, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005169.

Portero:2006:GPR

[PJ06b] L. Portero and J. C. Jorge. A generalization of Peaceman– Rachford fractional step method. Journal of Computational and Applied Mathematics, 189(1–2):676–688, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003687.

Portero:2008:NCS

[PJ08] L. Portero and J. C. Jorge. A new class of second order linearly implicit fractional step methods. Journal of Computational and Applied Mathematics, 218(2): 603–615, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000277.

Paris:2006:HEP

[PK06] R. B. Paris and D. Kaminski. Hyperasymptotic evaluation of the Pearcey integral via Hadamard expansions. Journal of Computational and Applied Mathematics, 190(1–2):437–452, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002451.

Popova:2007:IOB

[PK07] Evgenija Popova and Walter Krämer. Inner and outer bounds for the solution set of parametric linear systems. Journal of Computational and Applied Mathematics, 199(2):310–316, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007661.

Papatheodorou:2009:NNT

[PK09a] Theodore S. Papatheodorou and Anastasia N. Kandili. Novel numerical techniques based on Fokas transforms, for the solution of initial boundary value problems. Journal of Computational and Applied Mathematics, 227(1):75–82, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003300.

Plubtieng:2009:WCT

[PK09b] Somyot Plubtieng and Poom Kumam. Weak convergence theorem for monotone mappings and a countable family of nonexpansive mappings. Journal of Computational and Applied Mathematics, 224(2):614–621, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002781.

Pashos:2009:SSL

[PKS^{+}09] George Pashos, Michail E. Kavousanakis, Antony N. Spyropoulos, John A. Palyvos, and Andreas G. Boudouvis. Simultaneous solution of large-scale linear systems and eigenvalue problems with a parallel GMRES method. Journal of Computational and Applied Mathematics, 227(1):196–205, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800352X.

Plesniak:2009:MJI

[Ple09] W. Pleśniak. Multivariate Jackson inequality. Journal of Computational and Applied Mathematics, 233(3): 815–820, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001307.

Periaux:2009:FRA

[PLGS09] J. Periaux, D. S. Lee, L. F. Gonzalez, and K. Srinivas. Fast reconstruction of aerodynamic shapes using evolutionary algorithms and virtual Nash strategies in a CFD design environment. Journal of Computational and Applied Mathematics, 232 (1):61–71, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005451.

Pettersson:2008:IRB

[PLMP08] Ulrika Pettersson, Elisabeth Larsson, Gunnar Marcusson, and Jonas Persson. Improved radial basis function methods for multi-dimensional option pricing. Journal of Computational and Applied Mathematics, 222(1):82–93, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700550X.

Liu:2007:MQM

[pLT07] Ya ping Liu and Lü Tao.

Mechanical quadrature methods and their extrapolation for solving first kind Abel integral equations. Journal of Computational and Applied Mathematics, 201 (1):300–313, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001075.

Petkovic:2005:NHO

[PM05] Miodrag S. Petković and Dusan M. Milosević. A new higher-order family of inclusion zero-finding methods. Journal of Computational and Applied Mathematics, 182(2):416–432, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006387.

Phat:2006:SLT

[PN06] Vu N. Phat and Piyapong Niamsup. Stability of linear time-varying delay systems and applications to control problems. Journal of Computational and Applied Mathematics, 194(2): 343–356, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500498X.

Poklukar:2006:NFJ

[Pok06] Darja Rupnik Poklukar. Nonlinear filtering for jumpdiffusions. Journal of Computational and Applied Mathematics, 197(2):558–567, December 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006990.

Philos:2005:BSP

[PP05] Ch. G. Philos and I. K. Purnaras. The behavior of the solutions of periodic linear neutral delay difference equations. Journal of Computational and Applied Mathematics, 175(2):209–230, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002298.

Picq:2008:RTE

[PP08] Martine Picq and Jérôme Pousin. Resolution of the transport equation subject to constraint. Journal of Computational and Applied Mathematics, 218(2): 364–375, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700194X.

Picasso:2009:AAC

[PP09] Marco Picasso and Virabouth Prachittham. An adaptive algorithm for the Crank– Nicolson scheme applied to a time-dependent convection– diffusion problem. Journal of Computational and Applied Mathematics, 233(4):1139– 1154, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006268.

Pelayo:2007:GSS

[PPG07] Fernando L. Pelayo, Maria L.

Pelayo, and Juan L. G. Guirao. Generating the syntactic and semantics graphs for a Markovian process algebra. Journal of Computational and Applied Mathematics, 204(1):38–47, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002378.

Petkovic:2007:PES

[PPR07] Miodrag S. Petković, Ljiljana D. Petković, and Lidija Z. Rancić. Point estimation of simultaneous methods for solving polynomial equations: a survey (II). Journal of Computational and Applied Mathematics, 205 (1):32–52, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002548.

Pavlidis:2005:CNE

[PPV05] N. G. Pavlidis, K. E. Parsopoulos, and M. N. Vrahatis. Computing Nash equilibria through computational intelligence methods. Journal of Computational and Applied Mathematics, 175 (1):113–136, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002444.

Plaza:2005:AAT

[PR05] A. Plaza and M. C. Rivara. Average adjacencies for tetrahedral skeleton-regular partitions. Journal of Computational and Applied Mathematics, 177(1):141–158, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004194.

Polezzi:2007:DVB

[PR07] M. Polezzi and A. Sri Ranga.

On the denominator values and barycentric weights of rational interpolants. Journal of Computational and Applied Mathematics, 200(2): 576–590, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000379.

Pasadas:2008:MAP

[PR08a] M. Pasadas and M. L.

Rodríguez. Multivariate approximation by PDE splines. Journal of Computational and Applied Mathematics, 218(2):556–567, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005808.

Polak:2008:ESS

[PR08b] E. Polak and J. O. Royset. Efficient sample sizes in stochastic nonlinear programming. Journal of Computational and Applied Mathematics, 217(2):301–310, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700091X.

Prentice:2008:RMN

[Pre08a] J. S. C. Prentice. The RKGL method for the numerical solution of initial-value problems. Journal of Computational and Applied Mathematics, 213(2):477–487, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700057X.

Prevost:2008:LMM

[Pré08b] Marc Prévost. Legendre modified moments for Euler’s constant. Journal of Computational and Applied Mathematics, 219(2):484–492, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005134.

Prentice:2009:GEP

[Pre09] J. S. C. Prentice. General error propagation in the RK r GL m method. Journal of Computational and Applied Mathematics, 228 (1):344–354, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005025.

Perez-Rodriguez:2009:IRM

[PRGPS09] S. Perez-Rodriguez, S. Gonzalez-Pinto, and B. P. Sommeijer. An iterated Radau method for time-dependent PDEs. Journal of Computational and Applied Mathematics, 231(1):49–66, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000399.

Papadakis:2009:APO

[PRL09] K. Papadakis, O. Ragos, and C. Litzerinos. Asymmetric periodic orbits in the photogravitational Copenhagen problem. Journal of Computational and Applied Mathematics, 227(1):102–114, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003336.

Petkovic:2008:EHO

[PRP08] Ljiljana D. Petković, Lidija Rancić, and Miodrag S. Petković. An efficient higher order family of root finders. Journal of Computational and Applied Mathematics, 216(1):56–72, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700218X.

Paternoster:2009:P

[PRV09] Beatrice Paternoster, Elvira Russo, and Antonia Vecchio. Preface. Journal of Computational and Applied Mathematics, 228(2):503, June 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001313.

Prevost:2008:ICA

[PRZW08] Marc Prévost, Michela Redivo-Zaglia, and Franck Wielonsky. International Conference on Approximation and Iterative Methods. Journal of Computational and Applied Mathematics, 219(2): 327–328, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700622X.

Psihoyios:2005:FAO

[PS05] G. Psihoyios and T. E. Simos. A fourth algebraic order trigonometrically fitted predictor–corrector scheme for IVPs with oscillating solutions. Journal of Computational and Applied Mathematics, 175(1):137–147, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002535.

Pandey:2007:CFD

[PS07a] R. K. Pandey and Arvind K.

Singh. On the convergence of finite difference methods for weakly regular singular boundary value problems. Journal of Computational and Applied Mathematics, 205(1):469–478, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003311.

Potthast:2007:MRT

[PS07b] Roland Potthast and Jochen Schulz. A multiwave range test for obstacle reconstructions with unknown physical properties. Journal of Computational and Applied Mathematics, 205(1): 53–71, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600255X.

Pelosi:2008:SPH

[PS08] Francesca Pelosi and Paul Sablonnière. Shape-preserving C^{1} Hermite interpolants generated by a Gori–Pitolli subdivision scheme. Journal of Computational and Applied Mathematics, 220(1–2): 686–711, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005092.

Pandey:2009:CFO

[PS09a] R. K. Pandey and Arvind K. Singh. On the convergence of a fourth-order method for a class of singular boundary value problems. Journal of Computational and Applied Mathematics, 224(2): 734–742, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003002.

Petkovic:2009:GMI

[PS09b] Marko D. Petković and Predrag S. Stanimirović. Generalized matrix inversion is not harder than matrix multiplication. Journal of Computational and Applied Mathematics, 230(1):270–282, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006237.

Punzi:2008:MCD

[PSV08] Alessandro Punzi, Alvise Sommariva, and Marco Vianello. Meshless cubature over the disk using thin-plate splines. Journal of Computational and Applied Mathematics, 221(2):430–436, November 15, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005742.

Pedas:2006:SCM

[PT06] Arvet Pedas and Enn Tamme.

Spline collocation method for integro-differential equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 197(1):253–269, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500659X.

Pedas:2008:DGM

[PT08a] Arvet Pedas and Enn Tamme.

Discrete Galerkin method for Fredholm integro-differential equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 213(1): 111–126, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700009X.

Peng:2008:DDR

[PT08b] Chen Peng and Yu-Chu Tian. Delay-dependent robust stability criteria for uncertain systems with interval time-varying delay. Journal of Computational and Applied Mathematics, 214 (2):480–494, May 1, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001501.

Potts:2008:NSN

[PT08c] Daniel Potts and Manfred Tasche. Numerical stability of nonequispaced fast Fourier transforms. Journal of Computational and Applied Mathematics, 222(2): 655–674, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006474.

Parand:2009:RSG

[PT09a] K. Parand and A. Taghavi. Rational scaled generalized Laguerre function collocation method for solving the Blasius equation. Journal of Computational and Applied Mathematics, 233(4): 980–989, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709005883.

Penenko:2009:DAM

[PT09b] Vladimir Penenko and Elena Tsvetova. Discrete-analytical methods for the implementation of variational principles in environmental applications. Journal of Computational and Applied Mathematics, 226(2):319–330, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004111.

Pulova:2009:PPG

[Pul09] N. V. Pulova. A pointwise projected gradient method applied to an optimal control problem. Journal of Computational and Applied Mathematics, 226(2):331–335, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004123.

Postelmans:2005:MLJ

[PV05] Kelly Postelmans and Walter Van Assche. Multiple little q-Jacobi polynomials. Journal of Computational and Applied Mathematics, 178(1–2):361–375, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003929.

Pedas:2009:RSI

[PV09a] Arvet Pedas and Gennadi Vainikko. On the regularity of solutions to integral equations with nonsmooth kernels on a union of open intervals. Journal of Computational and Applied Mathematics, 229(2):440–451, July 15, 2009. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001647.

Pinnau:2009:CFE

[PV09b] René Pinnau and Jorge Mauricio Ruiz V. Convergent finite element discretizations of the density gradient equation for quantum semiconductors. Journal of Computational and Applied Mathematics, 223(2):790–800, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000848.

Pillards:2006:MSD

[PVC06] Tim Pillards, Bart Vandewoestyne, and Ronald Cools. Minimizing the L_{2} and L_{∞} star discrepancies of a single point in the unit hypercube. Journal of Computational and Applied Mathematics, 197(1):282–285, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006618.

Patie:2008:FET

[PW08] P. Patie and C. Winter. First exit time probability for multidimensional diffusions: a PDE-based approach. Journal of Computational and Applied Mathematics, 222(1): 42–53, December 1, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005481.

Peng:2008:SVI

[PZ08] Jian Wen Peng and Dao Li Zhu. A system of variational inclusions with P –η-accretive operators. Journal of Computational and Applied Mathematics, 216(1):198–209, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002282.

Qin:2009:CIA

[QCK09a] Xiaolong Qin, Sun Young Cho, and Shin Min Kang. Convergence of an iterative algorithm for systems of variational inequalities and nonexpansive mappings with applications. Journal of Computational and Applied Mathematics, 233(2): 231–240, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004142.

Qin:2009:CTC

[QCK09b] Xiaolong Qin, Yeol Je Cho, and Shin Min Kang. Convergence theorems of common elements for equilibrium problems and fixed point problems in Banach spaces. Journal of Computational and Applied Mathematics, 225 (1):20–30, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003166.

Qin:2009:SCT

[QCKK09] Xiaolong Qin, Yeol Je Cho, Jung Im Kang, and Shin Min Kang. Strong convergence theorems for an infinite family of nonexpansive mappings in Banach spaces. Journal of Computational and Applied Mathematics, 230(1): 121–127, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005918.

Qian:2006:FOM

[QFX06] Zhi Qian, Chu-Li Fu, and Xiang-Tuan Xiong. Fourthorder modified method for the Cauchy problem for the Laplace equation. Journal of Computational and Applied Mathematics, 192(2): 205–218, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002827.

Qi:2008:CLC

[QG08] Feng Qi and Bai-Ni Guo. A class of logarithmically completely monotonic functions and the best bounds in the second Kershaw’s double inequality. Journal of Computational and Applied Mathematics, 212(2):444–456, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007990.

Qi:2008:NUB

[QGC08] Feng Qi, Senlin Guo, and Shou-Xin Chen. A new upper bound in the second Kershaw’s double inequality and its generalizations. Journal of Computational and Applied Mathematics, 220(1–2): 111–118, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004281.

Qi:2007:CLC

[Qi07] Feng Qi. A class of logarithmically completely monotonic functions and the best bounds in the first Kershaw’s double inequality. Journal of Computational and Applied Mathematics, 206(2): 1007–1014, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005747.

Qi:2008:NLB

[Qi08] Feng Qi. A new lower bound in the second Kershaw’s double inequality. Journal of Computational and Applied Mathematics, 214 (2):610–616, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001586.

Qi:2009:CLC

[Qi09] Feng Qi. A class of logarithmically completely monotonic functions and application to the best bounds in the second Gautschi–Kershaw’s inequality. Journal of Computational and Applied Mathematics, 224(2):538–543, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002604.

Qin:2009:NAD

[Qin09] Jinggang Qin. The new alternating direction implicit difference methods for the wave equations. Journal of Computational and Applied Mathematics, 230(1): 213–223, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006018.

Qiu:2007:WSL

[Qiu07]  Jianxian Qiu.  WENO

schemes with Lax–Wendroff type time discretizations for Hamilton–Jacobi equations. Journal of Computational and Applied Mathematics, 200(2):591–605, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000380.

Qin:2009:PSB

[QLJ09] Zhongfeng Qin, Xiang Li, and Xiaoyu Ji. Portfolio selection based on fuzzy crossentropy. Journal of Computational and Applied Mathematics, 228(1):139–149, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004561.

Quinto:2007:LAE

[Qui07] Eric Todd Quinto. Local algorithms in exterior tomography. Journal of Computational and Applied Mathematics, 199(1):141–148, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007454.

Qamar:2006:AST

[QW06] Shamsul Qamar and Gerald Warnecke. Application of space-time CE/SE method to shallow water magnetohydrodynamic equations. Journal of Computational and Applied Mathematics, 196(1): 132–149, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005224.

Qamar:2008:ANI

[QW08] Shamsul Qamar and Gerald Warnecke. Analytical and numerical investigations of a batch crystallization model. Journal of Computational and Applied Mathematics, 222(2):715–731, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006528.

Qin:2009:MTR

[QWS09] H. H. Qin, T. Wei, and R. Shi. Modified Tikhonov regularization method for the Cauchy problem of the Helmholtz equation. Journal of Computational and Applied Mathematics, 224(1): 39–53, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001520.

Qi:2008:EDK

[QWW08] Liqun Qi, Yiju Wang, and Ed X. Wu. D-eigenvalues of diffusion kurtosis tensors. Journal of Computational and Applied Mathematics, 221(1):150–157, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005316.

Qu:2009:NCT

[QZY09] Shao-Jian Qu, Qing-Pu Zhang, and Yue-Ting Yang. A nonmonotone conic trust region method based on line search for solving unconstrained optimization. Journal of Computational and Applied Mathematics, 224(2): 514–526, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002586.

Qu:2008:NTR

[QZZ08] Shao-Jian Qu, Ke-Cun Zhang, and Jian Zhang. A nonmonotone trust-region method of conic model for unconstrained optimization. Journal of Computational and Applied Mathematics, 220(1–2): 119–128, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700430X.

Reihani:2007:RHF

[RA07] M. H. Reihani and Z. Abadi. Rationalized Haar functions method for solving Fredholm and Volterra integral equations. Journal of Computational and Applied Mathematics, 200(1):12–20, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007922.

Radwan:2005:CHO

[Rad05] Samir F. Radwan. Comparison of higher-order accurate schemes for solving the two-dimensional unsteady Burgers’ equation. Journal of Computational and Applied Mathematics, 174(2): 383–397, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002225.

Rodal:2007:LPD

[RAG07] J. Rodal, I. Area, and E. Godoy. Linear partial difference equations of hypergeometric type: Orthogonal polynomial solutions in two discrete variables. Journal of Computational and Applied Mathematics, 200(2): 722–748, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000690.

Roussos:2005:RER

[RB05] George Roussos and Brad J. C. Baxter. Rapid evaluation of radial basis functions. Journal of Computational and Applied Mathematics, 180(1):51–70, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004789.

Rypl:2006:GCS

[RB06] D. Rypl and Z. Bittnar. Generation of computational surface meshes of STL models. Journal of Computational and Applied Mathematics, 192(1):148–151, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500316X.

Rathinasamy:2008:MSS

[RB08] A. Rathinasamy and K. Balachandran. Mean-square stability of second-order Runge– Kutta methods for multidimensional linear stochastic differential systems. Journal of Computational and Applied Mathematics, 219(1): 170–197, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700386X.

Rihan:2005:FDV

[RCR05] Fathalla A. Rihan, Chris G. Collier, and Ian Roulstone. Four-dimensional variational data assimilation for Doppler radar wind data. Journal of Computational and Applied Mathematics, 176 (1):15–34, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003176.

Rasmussen:2008:GMA

[RD08] Bryan Rasmussen and Luca Dieci. A geometrical method for the approximation of invariant tori. Journal of Computational and Applied Mathematics, 216(2):388–412, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002701.

Reiner:2007:ANI

[RDH07] Robert C. Reiner, Jr., Rabia Djellouli, and Isaac Harari. Analytical and numerical investigation of the performance of the BGT2 condition for low-frequency acoustic scattering problems. Journal of Computational and Applied Mathematics, 204 (2):526–536, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003785.

Read:2007:ASM

[Rea07] W. W. Read. An analytic series method for Laplacian problems with mixed boundary conditions. Journal of Computational and Applied Mathematics, 209(1): 22–32, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006169.

Rezghi:2006:IPN

[RH06] Mansoor Rezghi and S. Mohammad Hosseini. An ILU preconditioner for nonsymmetric positive definite matrices by using the conjugate Gram–Schmidt process. Journal of Computational and Applied Mathematics, 188(1):150–164, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001858.

Rezghi:2009:NVC

[RH09] Mansoor Rezghi and S. Mohammad Hosseini. A new variant of L-curve for Tikhonov regularization. Journal of Computational and Applied Mathematics, 231(2):914–924, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003215.

Rhoads:2005:PTP

[Rho05]  Glenn C. Rhoads.  Pla-

nar tilings by polyominoes, polyhexes, and polyiamonds. Journal of Computational and Applied Mathematics, 174(2):329–353, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002195.

Ronghua:2006:CNS

[RHQ06] Li Ronghua, Meng Hongbing, and Chang Qin. Convergence of numerical solutions to stochastic age-dependent population equations. Journal of Computational and Applied Mathematics, 193(1): 109–120, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003821.

Rakha:2006:CRH

[RI06] Medhat A. Rakha and Adel K. Ibrahim. On the contiguous relations of hypergeometric series. Journal of Computational and Applied Mathematics, 192(2): 396–410, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003584.

Riedel:2006:LOK

[Rie06] Karl O. Riedel. Locally optimal knots and tension parameters for exponential splines. Journal of Computational and Applied Mathematics, 196(1):94–114, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005200.

Rashed:2008:GAN

[RK08a] A. S. Rashed and M. M.

Kassem. Group analysis for natural convection from a vertical plate. Journal of Computational and Applied Mathematics, 222(2): 392–403, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006073.

Ryoo:2008:NCS

[RK08b] C. S. Ryoo and Taekyun Kim. A numerical computation of the structure of the roots of q-Bernoulli polynomials. Journal of Computational and Applied Mathematics, 214(2):319–332, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001264.

Rauh:2007:ROI

[RKAH07] Andreas Rauh, Marco Kletting, Harald Aschemann, and Eberhard P. Hofer. Reduction of overestimation in interval arithmetic simulation of biological wastewater treatment processes. Journal of Computational and Applied Mathematics, 199(2): 207–212, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007521.

Rui:2007:RLS

[RKK07] Hongxing Rui, Seokchan Kim, and Sang Dong Kim. A remark on least-squares mixed element methods for reaction–diffusion problems. Journal of Computational and Applied Mathematics, 202(2):230–236, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001245.

Rui:2009:SLS

[RKK09] Hongxing Rui, Sang Dong Kim, and Seokchan Kim. Split least-squares finite element methods for linear and nonlinear parabolic problems. Journal of Computational and Applied Mathematics, 223(2):938–952, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001155.

Ren:2009:SPB

[RLH09] Yong Ren, Aihong Lin, and Lanying Hu. Stochastic PDIEs and backward doubly stochastic differential equations driven by Lévy processes. Journal of Computational and Applied Mathematics, 223(2):901–907, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001027.

Ren:2008:REU

[RLX08] Yong Ren, Shiping Lu, and Ningmao Xia. Remarks on the existence and uniqueness of the solutions to stochastic functional differential equations with infinite delay. Journal of Computational and Applied Mathematics, 220(1–2): 364–372, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004578.

Roy:2007:FSS

[RM07] A. R. Roy and P. K. Maji.

A fuzzy soft set theoretic approach to decision making problems. Journal of Computational and Applied Mathematics, 203(2):412–418, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002160.

See comment [KGW09].

Rouzine:2005:GFF

[RMKA05] Igor M. Rouzine, Kaja Murali-Krishna, and Rafi Ahmed. Generals die in friendly fire, or modeling immune response to HIV. Journal of Computational and Applied Mathematics, 184(1): 258–274, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000804.

Reilly:2005:SMS

[RO05] M. J. O. Reilly and E. O’Riordan. A Shishkin mesh for a singularly perturbed Riccati equation. Journal of Computational and Applied Mathematics, 182(2):372–387, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006363.

Rump:2007:SFV

[RO07] Siegfried M. Rump and Takeshi Ogita. Super-fast validated solution of linear systems. Journal of Computational and Applied Mathematics, 199(2):199–206, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500751X.

Roberts:2007:RRB

[Rob07] Catherine A. Roberts. Recent results on blow-up and quenching for nonlinear Volterra equations. Journal of Computational and Applied Mathematics, 205(2): 736–743, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003979.

Robenack:2008:CML

[Röb08] Klaus Röbenack. Computation of multiple Lie derivatives by algorithmic differentiation. Journal of Computational and Applied Mathematics, 213(2):454–464, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000441.

Roeger:2008:EFD

[Roe08] Lih-Ing W. Roeger. Exact finite-difference schemes for two-dimensional linear systems with constant coefficients. Journal of Computational and Applied Mathematics, 219(1):102–109, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042707003688.

Roig:2007:OST

[Roi07]  Bernardino Roig.  One-

step Taylor–Galerkin methods for convection–diffusion problems. Journal of Computational and Applied Mathematics, 204(1):95–101, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002421.

Romani:2009:ASS

[Rom09] L. Romani. From approximating subdivision schemes for exponential splines to high-performance interpolating algorithms. Journal of Computational and Applied Mathematics, 224(1): 383–396, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002446.

Roop:2006:CAF

[Roo06]  John Paul Roop.  Com-

putational aspects of FEM approximation of fractional advection dispersion equations on bounded domains in R^{2}. Journal of Computational and Applied Mathematics, 193(1):243–268, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003900.

Rossini:2008:CPB

[Ros08]  Milvia Rossini.  On the

construction of polyharmonic B-splines. Journal of Computational and Applied Mathematics, 221(2): 437–446, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005754.

Rangarajan:2005:LGD

[RP05] S. K. Rangarajan and Sudarshan P. Purushothaman. Lanczos’ generalized derivative for higher orders. Journal of Computational and Applied Mathematics, 177 (2):461–465, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005217.

Rancic:2009:NID

[RP09] Lidija Z. Rancić and Miodrag S. Petković. A note on the improved derivative free root-solvers. Journal of Computational and Applied Mathematics, 223(2): 535–539, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006504.

Roquejoffre:2006:EST

[RR06] Jean-Michel Roquejoffre and Hélène Rouzaud. Existence and stability of time-periodic solutions in a model for spherical flames with timeperiodic heat losses. Journal of Computational and Applied Mathematics, 190(1– 2):376–392, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002414.

Ramm:2005:DPN

[RS05a] Alexander G. Ramm and Alexandra B. Smirnova. On deconvolution problems: numerical aspects. Journal of Computational and Applied Mathematics, 176(2): 445–460, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003498.

Rocha:2005:APO

[RS05b] I. Alvarez Rocha and L. Salto. Asymptotics of polynomials orthogonal with respect to a discrete-complex Sobolev inner product. Journal of Computational and Applied Mathematics, 178(1–2):1–19, June 1, 2005. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003632.

Rodkina:2005:ASA

[RS05c] Alexandra Rodkina and Henri Schurz. Almost sure asymptotic stability of driftimplicit θ-methods for bilinear ordinary stochastic differential equations in R^{1}. Journal of Computational and Applied Mathematics, 180 (1):13–31, August 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004753.

Rosengren:2005:WEM

[RS05d] Hjalmar Rosengren and Michael Schlosser. On Warnaar’s elliptic matrix inversion and Karlsson–Minton-type elliptic hypergeometric series. Journal of Computational and Applied Mathematics, 178(1–2):377–391, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003875.

Ruffing:2005:CBS

[RS05e] Andreas Ruffing and Moritz Simon. Corresponding Banach spaces on time scales. Journal of Computational and Applied Mathematics, 179(1–2):313–326, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400456X.

Reichel:2007:P

[RS07a] Lothar Reichel and Fiorella Sgallari. Preface. Journal of Computational and Applied Mathematics, 198 (2):287, January 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007144.

Rozenblum:2007:FDO

[RS07b] G. Rozenblum and M. Solomyak. On a family of differential operators with the coupling parameter in the boundary condition. Journal of Computational and Applied Mathematics, 208(1):57–71, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006261.

Rapun:2008:MBI

[RS08a] M.-L. Rapún and F.-J. Sayas. Mixed boundary integral methods for Helmholtz transmission problems. Journal of Computational and Applied Mathematics, 214(1): 238–258, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001203.

Reichel:2008:NCI

[RS08b] Lothar Reichel and Hassane Sadok. A new L-curve for ill-posed problems. Journal of Computational and Applied Mathematics, 219(2): 493–508, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000568.

Rafiei:2008:NBF

[RT08] A. Rafiei and F. Toutounian.

New breakdown-free variant of AINV method for nonsymmetric positive definite matrices. Journal of Computational and Applied Mathematics, 219(1):72–80, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003640.

Rafiei:2009:BFV

[RT09] A. Rafiei and F. Toutounian.

Breakdown-free version of ILU factorization for nonsymmetric positive definite matrices. Journal of Computational and Applied Mathematics, 230(2):699–705, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000090.

Rambeerich:2009:ETI

[RTGB09] N. Rambeerich, D. Y. Tangman, A. Gopaul, and M. Bhuruth. Exponential time integration for fast finite element solutions of some financial engineering problems. Journal of Computational and Applied Mathematics, 224(2): 668–678, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002926.

Ruijsenaars:2005:RHF

[Rui05] S. N. M. Ruijsenaars. A relativistic hypergeometric function. Journal of Computational and Applied Mathematics, 178(1–2):393–417, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003887.

Rao:2005:ERHb

[RV05] K. Srinivasa Rao and G. Vanden Berghe. On an entry of Ramanujan in his Notebooks: a nested roots expansion. Journal of Computational and Applied Mathematics, 173(2):371–378, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001906.

Ramos:2005:NSS

[RVA05] Higinio Ramos and Jesús Vigo-Aguiar. A note on stepsize selection in the Störmer– Cowell methods. Journal of Computational and Applied Mathematics, 175(1): 149–159, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400250X.

Ramos:2007:FOR

[RVA07a] Higinio Ramos and Jesús Vigo-Aguiar. A fourthorder Runge–Kutta method based on BDF-type Chebyshev approximations. Journal of Computational and Applied Mathematics, 204 (1):124–136, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002445.

Ramos:2007:VSC

[RVA07b] Higinio Ramos and Jesús Vigo-Aguiar. Variablestepsize Chebyshev-type methods for the integration of second-order I.V.P.’s. Journal of Computational and Applied Mathematics, 204 (1):102–113, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002433.

Ricchiuto:2008:UHO

[RVAD08] M. Ricchiuto, N. Villedieu, R. Abgrall, and H. Deconinck. On uniformly highorder accurate residual distribution schemes for advection– diffusion. Journal of Computational and Applied Mathematics, 215(2):547–556, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007667.

Rao:2005:ERHa

[RVK05] K. Srinivasa Rao, G. Vanden Berghe, and C. Krattenthaler. An entry of Ramanujan on hypergeometric series in his Notebooks. Journal of Computational and Applied Mathematics, 173(2): 239–246, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001542.

See comment [Mil07].

Rauwoens:2008:NSF

[RVM08] Pieter Rauwoens, Jan Vierendeels, and Bart Merci. Numerical study of the flow in a three-dimensional thermally driven cavity. Journal of Computational and Applied Mathematics, 215 (2):538–546, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007655.

Rezounenko:2006:NLP

[RW06] Alexander V. Rezounenko and Jianhong Wu. A nonlocal PDE model for population dynamics with stateselective delay: Local theory and global attractors. Journal of Computational and Applied Mathematics, 190 (1–2):99–113, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002232.

Reginska:2009:WMM

[RW09] Teresa Regińska and Andrzej Wakulicz. Wavelet moment method for the Cauchy problem for the Helmholtz equation. Journal of Computational and Applied Mathematics, 223(1):218–229, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000083.

Wu:2006:BLA

[rWlC06] Tie ru Wu and Hong lu Cheng. Basic lines, axes and geometric modeling on implicit algebraic surfaces. Journal of Computational and Applied Mathematics, 195(1–2):212–219, October 15, 2006. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004863.

Rashidinia:2007:SVI

[RZ07] J. Rashidinia and M. Zarebnia. Solution of a Volterra integral equation by the sinccollocation method. Journal of Computational and Applied Mathematics, 206(2): 801–813, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005231.

Rzadkowski:2009:ETT

[Rza09] Grzegorz Rzadkowski. An extension of trapezoidal type product integration rules. Journal of Computational and Applied Mathematics, 232(2):625–631, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900394X.

Sandstrom:2007:NCZ

[SA07a] Sven-Erik Sandström and Christian Ackrén. Note on the complex zeros of H′_{ν}(x) + iζH_{ν}(x) = 0. Journal of Computational and Applied Mathematics, 201(1):3–7, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000720.

Soliman:2007:NSN

[SA07b] A. A. Soliman and M. A. Abdou. Numerical solutions of nonlinear evolution equations using variational iteration method. Journal of Computational and Applied Mathematics, 207(1):111–120, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004663.

Siddiqi:2008:SSS

[SA08a] Shahid S. Siddiqi and Ghazala Akram. Septic spline solutions of sixth-order boundary value problems. Journal of Computational and Applied Mathematics, 215 (1):288–301, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002038.

Solin:2008:EEM

[SÁ08b] Pavel Solín and José Ávila.

Equidistributed error mesh for problems with exponential boundary layers. Journal of Computational and Applied Mathematics, 218(1): 157–166, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003378.

Salarieh:2009:CSC

[SA09] Hassan Salarieh and Aria Alasty. Control of stochastic chaos using sliding mode method. Journal of Computational and Applied Mathematics, 225(1):135–145, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003580.

Saad:2006:OSG

[Saa06] Rachid Saad. An optimal schedule for Gaussian elimination on an MIMD architecture. Journal of Computational and Applied Mathematics, 185(1):91–106, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000518.

Sablonniere:2008:BSH

[Sab08]  Paul Sablonnière.  B-

splines and Hermite–Padé approximants to the exponential function. Journal of Computational and Applied Mathematics, 219(2): 509–517, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700252X.

Song:2009:CSF

[SACJ09] Xinghua Song, Martin Aigner, Falai Chen, and Bert Jüttler. Circular spline fitting using an evolution process. Journal of Computational and Applied Mathematics, 231(1): 423–433, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002076.

Sezer:2007:TMN

[SAD07] Mehmet Sezer and Aysegül Akyüz-Dascioğlu. A Taylor method for numerical solution of generalized pantograph equations with linear functional argument. Journal of Computational and Applied Mathematics, 200 (1):217–225, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008071.

Soto:2007:FCA

[SAFS07] Jesús Soto, M. Isabel Vigo Aguiar, and Antonio Flores-Sintas. A fuzzy clustering application to precise orbit determination. Journal of Computational and Applied Mathematics, 204(1):137–143, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002457.

Sadek:2009:OSC

[SAHS09] Ibrahim Sadek, Husein Abdul-Hamid, and Julie Scott. Optimal scanning control of flexible structures in twodimensional space. Journal of Computational and Applied Mathematics, 233(2): 389–404, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004361.

Saker:2005:OCS

[Sak05] S. H. Saker. Oscillation criteria of second-order half-linear dynamic equations on time scales. Journal of Computational and Applied Mathematics, 177(2):375–387, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004340.

Saker:2006:OSO

[Sak06] S. H. Saker. Oscillation of second-order nonlinear neutral delay dynamic equations on time scales. Journal of Computational and Applied Mathematics, 187(2): 123–141, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001482.

Sakajo:2009:EDF

[Sak09] Takashi Sakajo. An extension of Draghicescu’s fast treecode algorithm to the vortex method on a sphere. Journal of Computational and Applied Mathematics, 225 (1):158–171, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003609.

Salem:2009:FOP

[Sal09] Hussein A. H. Salem. On the fractional order m-point boundary value problem in reflexive Banach spaces and weak topologies. Journal of Computational and Applied Mathematics, 224(2): 565–572, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800263X.

Shen:2008:AMG

[SaLJ08] Pei-Ping Shen, Xiao ai Li, and Hong-Wei Jiao. Accelerating method of global optimization for signomial geometric programming. Journal of Computational and Applied Mathematics, 214 (1):66–77, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000842.

Sauer:2007:JPB

[Sau07] Tomas Sauer. Jacobi polynomials in Bernstein form. Journal of Computational and Applied Mathematics, 199(1): 149–158, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007466.

Sauer:2008:MRF

[Sau08]  Tomas Sauer.  Multivari-

ate refinable functions, differences and ideals — a simple tutorial. Journal of Computational and Applied Mathematics, 221(2): 447–459, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005766.

Savcenco:2008:CAS

[Sav08] V. Savcenco. Comparison of the asymptotic stability properties for two multirate strategies. Journal of Computational and Applied Mathematics, 220(1–2): 508–524, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004839.

Savcenco:2009:CMR

[Sav09] V. Savcenco. Construction of a multirate RODAS method for stiff ODEs. Journal of Computational and Applied Mathematics, 225(2): 323–337, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003701.

Sabo:2009:BED

[SB09] Kristian Sabo and Mirta Bensić. Border estimation of a disc observed with random errors solved in two steps. Journal of Computational and Applied Mathematics, 229(1):16–26, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005141.

Siklosi:2007:IIS

[SBK07] Malin Siklosi, Begzsuren Batzorig, and Gunilla Kreiss. An investigation of the internal structure of shock profiles for shock capturing schemes. Journal of Computational and Applied Mathematics, 201(1):8–29, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000756.

Sorevik:2009:NST

[SBO09] Tor Sørevik, Tore Birkeland, and Gabriel Oksa. Numerical solution of the 3D time dependent Schrödinger equation in spherical coordinates: Spectral basis and effects of split-operator technique. Journal of Computational and Applied Mathematics, 225(1):56–67, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003191.

Song:2006:SAC

[SC06a] Qiankun Song and Jinde Cao. Stability analysis of Cohen–Grossberg neural network with both timevarying and continuously distributed delays. Journal of Computational and Applied Mathematics, 197(1): 188–203, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006527.

Spaletta:2006:CIB

[SC06b] Giulia Spaletta and Luca Caucci. Constrained iterations for blind deconvolution and convexity issues. Journal of Computational and Applied Mathematics, 197(1): 29–43, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500631X.

Shingareva:2007:FAD

[SC07a] Inna Shingareva and Carlos Lizárraga Celaya. On frequency-amplitude dependences for surface and internal standing waves. Journal of Computational and Applied Mathematics, 200(2): 459–470, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000288.

Song:2007:GES

[SC07b] Qiankun Song and Jinde Cao. Global exponential stability of bidirectional associative memory neural networks with distributed delays. Journal of Computational and Applied Mathematics, 202(2):266–279, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001282.

Sheng:2008:RCG

[SCG08] Xingping Sheng, Guoliang Chen, and Yi Gong. The representation and computation of generalized inverse A_{T,S}^{(2)}. Journal of Computational and Applied Mathematics, 213(1):248–257, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000325.

Schiefermayr:2005:SNP

[Sch05] Klaus Schiefermayr. Some new properties of Jacobi’s theta functions. Journal of Computational and Applied Mathematics, 178(1– 2):419–424, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004613.

Schmidt:2006:ECC

[Sch06a]  K. M. Schmidt.  Excep-

tional coupling constants for the Coulomb–Dirac operator with anomalous magnetic moment. Journal of Computational and Applied Mathematics, 194(1):17–25, September 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003997.

Schoutens:2006:EOU

[Sch06b]  Wim Schoutens.  Exotic

options under Lévy models: an overview. Journal of Computational and Applied Mathematics, 189(1– 2):526–538, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006138.

Schafer:2007:CAP

[Sch07a] Uwe Schäfer. On computerassisted proofs for solutions of linear complementarity problems. Journal of Computational and Applied Mathematics, 199(2):257–262, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007594.

Schwandt:2007:PIN

[Sch07b] Hartmut Schwandt. Parallel interval Newton-like Schwarz methods for almost linear parabolic problems. Journal of Computational and Applied Mathematics, 199(2): 437–444, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007843.

Schaback:2008:LPI

[Sch08]  Robert Schaback.  Limit

problems for interpolation by analytic radial basis functions. Journal of Computational and Applied Mathematics, 212(2):127–149, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007151.

Shen:2009:NOS

[SCM09] Pei-Ping Shen, Yong-Qiang Chen, and Yuan Ma. A nonisolated optimal solution for special reverse convex programming problems. Journal of Computational and Applied Mathematics, 224(1): 219–229, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002033.

Serra-Capizzano:2008:SNA

[SCS08] Stefano Serra-Capizzano and Per Sundqvist. Stability of the notion of approximating class of sequences and applications. Journal of Computational and Applied Mathematics, 219(2):518–536, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700180X.

Scuderi:2009:NSS

[Scu09]  Letizia Scuderi.  A new

smoothing strategy for computing nearly singular integrals in 3D Galerkin BEM. Journal of Computational and Applied Mathematics, 225(2):406–427, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003944.

Sun:2009:PPL

[SCY09] Jia-Chang Sun, Jian-Wen Cao, and Chao Yang. Parallel preconditioners for large scale partial difference equation systems. Journal of Computational and Applied Mathematics, 226(1):125–135, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002380.

Steele:2006:ECS

[SD06] A. G. Steele and R. J. Douglas. Extending chi-squared statistics for key comparisons in metrology. Journal of Computational and Applied Mathematics, 192 (1):51–58, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003079.

Slodicka:2009:NAS

[SD09] Marián Slodicka and Sofiane Dehilis. A numerical approach for a semilinear parabolic equation with a nonlocal boundary condition. Journal of Computational and Applied Mathematics, 231(2):715–724, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002714.

Soleimani-damaneh:2009:SER

[SdJMH09] M. Soleimani-damaneh, G. R.

Jahanshahloo, S. Mehrabian, and M. Hasannasab. Scale elasticity and returns to scale in the presence of alternative solutions. Journal of Computational and Applied Mathematics, 233(2): 127–136, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004002.

Shen:2009:SBD

[SDP09] Pei-Ping Shen, Yun-Peng Duan, and Yong-Gang Pei. A simplicial branch and duality bound algorithm for the sum of convex-convex ratios problem. Journal of Computational and Applied Mathematics, 223(1):145–158, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000046.

Speleers:2006:NSP

[SDV06] Hendrik Speleers, Paul Dierckx, and Stefan Vandewalle. Numerical solution of partial differential equations with Powell–Sabin splines. Journal of Computational and Applied Mathematics, 189(1– 2):643–659, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001056.

Speleers:2007:PSS

[SDV07] Hendrik Speleers, Paul Dierckx, and Stefan Vandewalle. Powell–Sabin splines with boundary conditions for polygonal and nonpolygonal domains. Journal of Computational and Applied Mathematics, 206(1): 55–72, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003815.

Seaid:2007:MNK

[Sea07] Mohammed Seaid̈. Multigrid Newton–Krylov method for radiation in diffusive semitransparent media. Journal of Computational and Applied Mathematics, 203 (2):498–515, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600224X.

Salam:2008:SHT

[SEAA08] A. Salam, A. El Farouk, and E. Al-Aidarous. Symplectic Householder transformations for a QR-like decomposition, a geometric and algebraic approaches. Journal of Computational and Applied Mathematics, 214(2):533–548, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001537.

Senoglu:2007:EPO

[Sen07] Birdal Senoğlu. Estimating parameters in one-way analysis of covariance model with short-tailed symmetric error distributions. Journal of Computational and Applied Mathematics, 201 (1):275–283, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001051.

Sun:2006:LTE

[SF06] Tong Sun and Daria Filippova. Long-time error estimation for semi-linear parabolic equations. Journal of Computational and Applied Mathematics, 185(1):1–18, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000476.

Skiba:2008:CAO

[SF08] Yuri N. Skiba and Denis M. Filatov. Conservative arbitrary order finite difference schemes for shallow-water flows. Journal of Computational and Applied Mathematics, 218(2): 579–591, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006589.

Sifalakis:2008:GDN

[SFFS08] A. G. Sifalakis, A. S. Fokas, S. R. Fulton, and Y. G. Saridakis. The generalized Dirichlet–Neumann map for linear elliptic PDEs and its numerical implementation. Journal of Computational and Applied Mathematics, 219(1):9–34, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003603.

Sifalakis:2009:DIS

[SFPS09] A. G. Sifalakis, S. R. Fulton, E. P. Papadopoulou, and Y. G. Saridakis. Direct and iterative solution of the generalized Dirichlet– Neumann map for elliptic PDEs on square domains. Journal of Computational and Applied Mathematics, 227(1):171–184, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003506.

Shinohara:2007:BVP

[SFSK07] Yasuaki Shinohara, Humiyuki Fujimori, Tomohiro Suzuki, and Mitsunobu Kurihara. On a boundary value problem for delay differential equations of population dynamics and Chebyshev approximation. Journal of Computational and Applied Mathematics, 201(2):348–355, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000872.

Sakkalis:2009:NER

[SFV09] Takis Sakkalis, Rida T. Farouki, and Leonid Vaserstein. Non-existence of rational arc length parameterizations for curves in R^{n}. Journal of Computational and Applied Mathematics, 228 (1):494–497, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800486X.

Sadkane:2005:NLS

[SG05] M. Sadkane and L. Grammont. A note on the Lyapunov stability of periodic discrete-time systems. Journal of Computational and Applied Mathematics, 176 (2):463–466, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003978.

Shi:2008:NTR

[SG08] Zhen-Jun Shi and Jin-Hua Guo. A new trust region method for unconstrained optimization. Journal of Computational and Applied Mathematics, 213 (2):509–520, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000696.

Shi:2009:NFC

[SG09] Zhen-Jun Shi and Jinhua Guo. A new family of conjugate gradient methods. Journal of Computational and Applied Mathematics, 224(1): 444–457, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002513.

Sarfraz:2006:DVU

[SH06] M. Sarfraz and Malik Zawwar Hussain. Data visualization using rational spline interpolation. Journal of Computational and Applied Mathematics, 189(1–2):513–525, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002980.

Sugiura:2009:QRA

[SH09] Hiroshi Sugiura and Takemitsu Hasegawa. Quadrature rule for Abel’s equations: Uniformly approximating fractional derivatives. Journal of Computational and Applied Mathematics, 223(1): 459–468, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000368.

Shardlow:2005:NSS

[Sha05] Tony Shardlow. Numerical simulation of stochastic PDEs for excitable media. Journal of Computational and Applied Mathematics, 175(2):429–446, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003140.

Shampine:2007:DSO

[Sha07] L. F. Shampine. Design of software for ODEs. Journal of Computational and Applied Mathematics, 205(2): 901–911, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004110.

Shampine:2008:VAQ

[Sha08] L. F. Shampine. Vectorized adaptive quadrature in MAT-LAB. Journal of Computational and Applied Mathematics, 211(2):131–140, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600700X.

Shakhno:2009:IAS

[Sha09a] S. M. Shakhno. On an iterative algorithm with superquadratic convergence for solving nonlinear operator equations. Journal of Computational and Applied Mathematics, 231(1): 222–235, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000569.

Shao:2009:APS

[Sha09b] Jianying Shao. An antiperiodic solution for a class of recurrent neural networks. Journal of Computational and Applied Mathematics, 228(1):231–237, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004706.

Shao:2005:NMT

[SHB05] Ji-Cheng Shao, Yozo Hamano, and Michael Bevis. A note on Maxwell’s theory of poles. Journal of Computational and Applied Mathematics, 183(1):101–107, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000166.

Shen:2008:CNS

[SHC08] Shu-Qian Shen, Ting-Zhu Huang, and Guang-Hui Cheng. A condition for the nonsymmetric saddle point matrix being diagonalizable and having real and positive eigenvalues. Journal of Computational and Applied Mathematics, 220(1–2):8–12, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003652.

Shioda:2007:SUL

[Shi07] S. Shioda. Some upper and lower bounds on the coupon collector problem. Journal of Computational and Applied Mathematics, 200(1): 154–167, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008022.

Schmidt:2008:PCN

[SHK^{+}08] Frank Schmidt, Thorsten Hohage, Roland Klose, Achim Schädle, and Lin Zschiedrich. Pole condition: a numerical method for Helmholtztype scattering problems with inhomogeneous exterior domain. Journal of Computational and Applied Mathematics, 218(1):61–69, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003275.

Sickenberger:2008:MSC

[Sic08] Thorsten Sickenberger. Meansquare convergence of stochastic multi-step methods with variable step-size. Journal of Computational and Applied Mathematics, 212(2): 300–319, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007321.

Sidi:2005:ACV

[Sid05] Avram Sidi. Application of class S_{m} variable transformations to numerical integration over surfaces of spheres. Journal of Computational and Applied Mathematics, 184(2): 475–492, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000439.

Sidi:2008:FEC

[Sid08] Avram Sidi. Further extension of a class of periodizing variable transformations for numerical integration. Journal of Computational and Applied Mathematics, 221(1): 132–149, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005304.

Siedentop:2007:SRA

[Sie07a] Heinz Siedentop. The spectrum of relativistic atoms according to Bethe and Salpeter and beyond. Journal of Computational and Applied Mathematics, 208(1): 155–163, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006327.

Siewer:2007:CAN

[Sie07b] Ralf Siewer. A constructive approach to nodal splines. Journal of Computational and Applied Mathematics, 203(1):289–308, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002123.

Siewer:2008:HS

[Sie08] Ralf Siewer. Histopolating splines. Journal of Computational and Applied Mathematics, 220(1–2):661–673, October 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005079.

Simon:2007:CMF

[Sim07a] Barry Simon. CMV matrices: Five years after. Journal of Computational and Applied Mathematics, 208(1): 120–154, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006315.

Sims:2007:RSL

[Sim07b] Robert Sims. Reflectionless Sturm–Liouville equations. Journal of Computational and Applied Mathematics, 208(1):207–225, November 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006364.

Singh:2008:NCG

[Sin08] Vimal Singh. A new criterion for global robust stability of interval delayed neural networks. Journal of Computational and Applied Mathematics, 221(1): 219–225, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005900.

Sinescu:2009:SLR

[Sin09] Vasile Sinescu. Shifted lattice rules based on a general weighted discrepancy for integrals over Euclidean space. Journal of Computational and Applied Mathematics, 232(2):240–251, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003471.

Sivakumar:2009:PIS

[Siv09] B. Sivakumar. A perishable inventory system with retrial demands and a finite population. Journal of Computational and Applied Mathematics, 224(1):29–38, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001519.

Shen:2009:DSE

[SJ09] Shoufeng Shen and Liya Jiang. The Davey–Stewartson equation with sources derived from nonlinear variable separation method. Journal of Computational and Applied Mathematics, 233(2): 585–589, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900421X.

Sanchez:2009:WBC

[SJV^{+}09] Fabrício Lopes Sanchez, Sylvio Barbon Júnior, Lucimar Sasso Vieira, Rodrigo Capobianco Guido, Everthon Silva Fonseca, Paulo Rogério Scalassara, Carlos Dias Maciel, José Carlos Pereira, and Shi-Huang Chen. Wavelet-based cepstrum calculation. Journal of Computational and Applied Mathematics, 227 (2):288–293, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001118.

Shi:2009:FPA

[SJZ09] Zhenwei Shi, Zhiguo Jiang, and Fugen Zhou. A fixedpoint algorithm for blind source separation with nonlinear autocorrelation. Journal of Computational and Applied Mathematics, 223(2): 908–915, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001039.

Shi:2009:BSS

[SJZY09] Zhenwei Shi, Zhiguo Jiang, Fugen Zhou, and Jihao Yin. Blind source separation with nonlinear autocorrelation and non-Gaussianity. Journal of Computational and Applied Mathematics, 229(1):240–247, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005621.

Sakai:2005:NSB

[SK05] Katsuhiro Sakai and Isao Kimura. A numerical scheme based on a solution of nonlinear advection–diffusion equations. Journal of Computational and Applied Mathematics, 173(1):39–55, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704001323.

Spowart:2006:NAE

[SK06] Michael P. Spowart and Edward F. Kuester. A novel asymptotic extraction technique for the efficient evaluation of a class of double Sommerfeld integrals. Journal of Computational and Applied Mathematics, 197(2): 597–611, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007028.

Slemp:2009:IBC

[SK09a] Wesley C. H. Slemp and Rakesh K. Kapania. Imposing boundary conditions in sinc method using highest derivative approximation. Journal of Computational and Applied Mathematics, 230(2): 371–392, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800633X.

Smyrlis:2009:EIM

[SK09b] Yiorgos-Sokratis Smyrlis and Andreas Karageorghis. Efficient implementation of the MFS: the three scenarios. Journal of Computational and Applied Mathematics, 227(1):83–92, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003312.

Schlegel:2009:MRK

[SKAW09] Martin Schlegel, Oswald Knoth, Martin Arnold, and Ralf Wolke. Multirate Runge–Kutta schemes for advection equations. Journal of Computational and Applied Mathematics, 226(2): 345–357, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004147.

Skrzipek:2005:GAP

[Skr05]  M.-R. Skrzipek.  Gener-

alized associated polynomials and functions of second kind. Journal of Computational and Applied Mathematics, 178(1–2):425–436, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003899.

Santos-Leon:2005:SPS

[SL05] J. C. Santos-León. Szegő polynomials and Szegő quadrature for the Fejér kernel. Journal of Computational and Applied Mathematics, 179(1–2):327–341, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004571.

Santos:2006:AFS

[SL06] Felix C. G. Santos and Maria Lencastre. An approach for FEM simulator development. Journal of Computational and Applied Mathematics, 185(2):326–346, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001184.

Sun:2008:NPS

[SL08] Hong-Rui Sun and Wan-Tong Li. On the number of positive solutions of systems of nonlinear dynamic equations on time scales. Journal of Computational and Applied Mathematics, 219(1): 123–133, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003706.

Shieh:2009:NTC

[SL09] Tzong-Hann Shieh and Meng-Rong Li. Numeric treatment of contact discontinuity with multi-gases. Journal of Computational and Applied Mathematics, 230(2): 656–673, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000065.

Shi:2006:NAS

[SLH06] Xiquan Shi, Fengshan Liu, and Minghan Hu. A new asymptotic series for the gamma function. Journal of Computational and Applied Mathematics, 195(1–2): 134–154, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004802.

Shou:2006:MAA

[SLMW06] Huahao Shou, Hongwei Lin, Ralph R. Martin, and Guojin Wang. Modified affine arithmetic in tensor form for trivariate polynomial evaluation and algebraic surface plotting. Journal of Computational and Applied Mathematics, 195(1–2):155– 171, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004814.

Simpson:2006:CDM

[SLN06] Matthew J. Simpson, Kerry A. Landman, and Donald F. Newgreen. Chemotactic and diffusive migration on a nonuniformly growing domain: numerical algorithm development and applications. Journal of Computational and Applied Mathematics, 192(2):282–300, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003298.

Sbarciog:2008:CSB

[SLN08] Mihaela Sbarciog, Mia Loccufier, and Erik Noldus. The computation of stability boundaries in state space for a class of biochemical engineering systems. Journal of Computational and Applied Mathematics, 215 (2):557–567, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007679.

Slodicka:2008:NDT

[Slo08] Marián Slodicka. Nonlinear diffusion in type-II superconductors. Journal of Computational and Applied Mathematics, 215(2):568–576, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007680.

Su:2006:ALP

[SLZ06] Zhixun Su, Ling Li, and Xiaojie Zhou. Arc-length preserving curve deformation based on subdivision. Journal of Computational and Applied Mathematics, 195(1–2): 172–181, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004826.

Sun:2009:PSS

[SLZA09] Yan Sun, Lishan Liu, Jizhou Zhang, and R. P. Agarwal. Positive solutions of singular three-point boundary value problems for second-order differential equations. Journal of Computational and Applied Mathematics, 230(2): 738–750, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000120.

Sawyer:2007:IFV

[SM07] William B. Sawyer and Arthur A. Mirin. The implementation of the finitevolume dynamical core in the community atmosphere model. Journal of Computational and Applied Mathematics, 203(2):387–396, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002238.

Stephens:2009:PUS

[SM09] Meghan Stephens and Niall Madden. A parameteruniform Schwarz method for a coupled system of reaction– diffusion equations. Journal of Computational and Applied Mathematics, 230(2): 360–370, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006328.

Smaoui:2005:BDC

[Sma05] Nejib Smaoui. Boundary and distributed control of the viscous Burgers equation. Journal of Computational and Applied Mathematics, 182(1):91–104, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400603X.

Smietanski:2007:GJB

[Śmi07] Marek J. Śmietański. A generalized Jacobian based Newton method for semismooth block-triangular system of equations. Journal of Computational and Applied Mathematics, 205(1): 305–313, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003190.

Smolarski:2006:DSM

[Smo06] Dennis C. Smolarski. Diagonallystriped matrices and approximate inverse preconditioners. Journal of Computational and Applied Mathematics, 186(2):416–431, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001020.

Song:2007:PDS

[SMT07] Mei Song, Wanbiao Ma, and Yasuhiro Takeuchi. Permanence of a delayed SIR epidemic model with density dependent birth rate. Journal of Computational and Applied Mathematics, 201 (2):389–394, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000926.

Sebastian:2006:SLF

[SNV06] M. V. Sebastián, M. A.

Navascués, and J. R. Valdizán. Surface Laplacian and fractal brain mapping. Journal of Computational and Applied Mathematics, 189(1– 2):132–141, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002992.

Somali:2005:IES

[SO05] S. Somali and V. Oger. Improvement of eigenvalues of Sturm–Liouville problem with t-periodic boundary conditions. Journal of Computational and Applied Mathematics, 180(2):433–441, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005436.

Sohn:2005:ASU

[Soh05] Sung-Ik Sohn. Analytic solutions of unstable interfaces for all density ratios in axisymmetric flows. Journal of Computational and Applied Mathematics, 177(2):367–374, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004327.

Solin:2008:P

[Sol08] Pavel Solín. Preface. Journal of Computational and Applied Mathematics, 218 (1):1, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003202.

Simos:2005:P

[SP05] T. E. Simos and G. Psihoyios.

Preface. Journal of Computational and Applied Mathematics, 175(1):ix, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002547.

Simos:2006:P

[SP06a] T. E. Simos and G. Psihoyios.

Preface. Journal of Computational and Applied Mathematics, 191(2):165, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005455.

Song:2006:PSN

[SP06b] Yongli Song and Yahong Peng. Periodic solutions of a nonautonomous periodic model of population with continuous and discrete time. Journal of Computational and Applied Mathematics, 188(2):256–264, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002542.

Saeidifar:2009:PMF

[SP09a] A. Saeidifar and E. Pasha.

The possibilistic moments of fuzzy numbers and their applications. Journal of Computational and Applied Mathematics, 223(2):1028– 1042, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800143X.

Su:2009:NFT

[SP09b] Ke Su and Dingguo Pu. A nonmonotone filter trust region method for nonlinear constrained optimization. Journal of Computational and Applied Mathematics, 223(1):230–239, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000095.

Spreeuw:2006:TAP

[Spr06] Jaap Spreeuw. Two approximations of the present value distribution of a disability annuity. Journal of Computational and Applied Mathematics, 186(1):217–231, February 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001974.

Roan:2007:SCC

[sR07] Shi shyr Roan. Structure of certain Chebyshev-type polynomials in Onsager’s algebra representation. Journal of Computational and Applied Mathematics, 202 (1):88–104, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001129.

Sarfraz:2006:CPS

[SRB06] M. Sarfraz, M. Riyazuddin, and M. H. Baig. Capturing planar shapes by approximating their outlines. Journal of Computational and Applied Mathematics, 189(1– 2):494–512, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006151.

Sanchez-Ruiz:2005:FIO

[SRD05] Jorge Sánchez-Ruiz and Jesús S. Dehesa. Fisher information of orthogonal hypergeometric polynomials. Journal of Computational and Applied Mathematics, 182(1): 150–164, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006077.

Skalna:2008:SFE

[SRP08] Iwona Skalna, M. V. Rama Rao, and Andrzej Pownuk. Systems of fuzzy equations in structural mechanics. Journal of Computational and Applied Mathematics, 218(1): 149–156, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003366.

Shih:2009:NPB

[SRW09] Y. Shih, C. Rei, and H. Wang. A novel PDE based image restoration: Convectiondiffusion equation for image denoising. Journal of Computational and Applied Mathematics, 231(2): 771–779, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002908.

Sakas:2005:MME

[SS05] D. P. Sakas and T. E. Simos. Multiderivative methods of eighth algebraic order with minimal phase-lag for the numerical solution of the radial Schrödinger equation. Journal of Computational and Applied Mathematics, 175 (1):161–172, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002523.

Shi:2006:CNL

[SS06a] Zhen-Jun Shi and Jie Shen. Convergence of nonmonotone line search method. Journal of Computational and Applied Mathematics, 193(2): 397–412, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004218.

Sofroniou:2006:HSC

[SS06b] M. Sofroniou and G. Spaletta.

Hybrid solvers for composition and splitting methods. Journal of Computational and Applied Mathematics, 185(2):278–291, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001159.

Sutmann:2006:HOC

[SS06c] Godehard Sutmann and Bernhard Steffen. High-order compact solvers for the threedimensional Poisson equation. Journal of Computational and Applied Mathematics, 187(2):142–170, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001500.

Suzuki:2007:EPD

[SS07] Tomohiro Suzuki and Toshio Suzuki. An eigenvalue problem for derogatory matrices. Journal of Computational and Applied Mathematics, 199(2):245–250, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007570.

Smirnov:2008:AUN

[SS08] Evgueni M. Smirnov and Paul E. Smirnov. Application of an unstructured Navier–Stokes code to prediction of adiabatic effectiveness of endwall flush-slot-cooling for a stator vane passage. Journal of Computational and Applied Mathematics, 215(2):577–585, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007692.

Salarieh:2009:MSC

[SS09a] Hassan Salarieh and Mohammad Shahrokhi. Multisynchronization of chaos via linear output feedback strategy. Journal of Computational and Applied Mathematics, 223(2):842–852, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000873.

Sang:2009:PSN

[SS09b] Yanbin Sang and Hua Su. Positive solutions of nonlinear third-order m-point BVP for an increasing homeomorphism and homomorphism with sign-changing nonlinearity. Journal of Computational and Applied Mathematics, 225(1):288–300, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003695.

Sang:2009:SAT

[SS09c] Zhaoyang Sang and Qingying Sun. A self-adaptive trust region method with line search based on a simple subproblem model. Journal of Computational and Applied Mathematics, 232(2):514–522, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003847.

Slevinsky:2009:NFH

[SS09d] Richard M. Slevinsky and Hassan Safouhi. New formulae for higher order derivatives and applications. Journal of Computational and Applied Mathematics, 233(2): 405–419, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004385.

Stahl:2009:EP

[SS09e] Herbert Stahl and Thomas Schmelzer. An extension of the ‘1/9’-problem. Journal of Computational and Applied Mathematics, 233(3): 821–834, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001319.

Swart:2007:NST

[SSMB07] Arno Swart, Gerard L. G. Sleijpen, Leo R. M. Maas, and Jan Brandts. Numerical solution of the twodimensional Poincaré equation. Journal of Computational and Applied Mathematics, 200(1):317–341, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000185.

SanMartin:2009:CFE

[SST09] Jorge San Martín, Loredana Smaranda, and Takéo Takahashi. Convergence of a finite element/ALE method for the Stokes equations in a domain depending on time. Journal of Computational and Applied Mathematics, 230(2): 521–545, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006493.

Shampine:2006:IIS

[SSV06] L. F. Shampine, B. P. Sommeijer, and J. G. Verwer. IRKC: an IMEX solver for stiff diffusion-reaction PDEs. Journal of Computational and Applied Mathematics, 196(2):485–497, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500600X.

Shishkina:2009:STT

[SSW09] Olga Shishkina, Andrei Shishkin, and Claus Wagner. Simulation of turbulent thermal convection in complicated domains. Journal of Computational and Applied Mathematics, 226(2):336–344, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004135.

Sogabe:2009:ECR

[SSZ09] T. Sogabe, M. Sugihara, and S.-L. Zhang. An extension of the conjugate residual method to nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 226 (1):103–113, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002264.

Sanders:2005:TOF

[ST05] J. B. Sanders and N. M.

Temme. On the temporal order of first-passage times in one-dimensional lattice random walks. Journal of Computational and Applied Mathematics, 182(1): 134–149, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006065.

Song:2007:PAP

[ST07] Yihong Song and Hongjiong Tian. Periodic and almost periodic solutions of nonlinear discrete Volterra equations with unbounded delay. Journal of Computational and Applied Mathematics, 205(2):859–870, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004080.

Stadler:2007:PFA

[Sta07] Georg Stadler. Path-following and augmented Lagrangian methods for contact problems in linear elasticity. Journal of Computational and Applied Mathematics, 203 (2):533–547, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002251.

Sun:2009:IGB

[STB09] Yang Sun, Yu-Hui Tao, and Feng-Shan Bai. Incomplete Gröbner basis as a preconditioner for polynomial systems. Journal of Computational and Applied Mathematics, 226(1):2–9, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002161.

Steinbach:2007:PST

[Ste07]  Marc C. Steinbach.  On

PDE solution in transient optimization of gas networks. Journal of Computational and Applied Mathematics, 203(2):345–361, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002263.

Stojanovic:2005:GCM

[Sto05a] Mirjana Stojanović. Global convergence method for singularly perturbed boundary value problems. Journal of Computational and Applied Mathematics, 181(2): 326–335, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005886.

Stoyanov:2005:CAE

[Sto05b] B. J. Stoyanov. Comments on “Asymptotic expansion of a Bessel function integral using hypergeometric functions” by L. J. Landau and N. J. Luswili. Journal of Computational and Applied Mathematics, 176(2):259–262, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003310.

See [LL01].

Stoyanova:2006:ICF

[Sto06] Srebra B. Stoyanova. Invariant cubature formulae of degree 6 for the n-simplex. Journal of Computational and Applied Mathematics, 193(2):446–459, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004243.

Stoll:2008:DPC

[Sto08] Thomas Stoll. Decomposition of perturbed Chebyshev polynomials. Journal of Computational and Applied Mathematics, 214(2):356–370, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700132X.

Sun:2008:ETP

[Sun08] Yongping Sun. Existence of triple positive solutions for a third-order three-point boundary value problem. Journal of Computational and Applied Mathematics, 221(1):194–201, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005791.

Sun:2009:NSR

[Sun09] Tong Sun. Numerical smoothing of Runge–Kutta schemes. Journal of Computational and Applied Mathematics, 233(4):1056–1062, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006207.

Suri:2005:SMF

[Sur05] Manil Suri. Stable hp mixed finite elements based on the Hellinger–Reissner principle. Journal of Computational and Applied Mathematics, 174(2):213–225, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400189X.

Sutmann:2007:CFD

[Sut07] Godehard Sutmann. Compact finite difference schemes of sixth order for the Helmholtz equation. Journal of Computational and Applied Mathematics, 203 (1):15–31, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001622.

Suzuki:2005:POA

[Suz05] Takashi Suzuki. Parallel optimization applied to magnetoencephalography. Journal of Computational and Applied Mathematics, 183(1): 177–190, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000245.

Sima:2007:SNL

[SV07a] Diana M. Sima and Sabine Van Huffel. Separable nonlinear least squares fitting with linear bound constraints and its application in magnetic resonance spectroscopy data quantification. Journal of Computational and Applied Mathematics, 203 (1):264–278, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600210X.

Solin:2007:WDM

[SV07b] Pavel Solín and Tomás Vejchodský. A weak discrete maximum principle for hp-FEM. Journal of Computational and Applied Mathematics, 209(1):54–65, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006194.

Sommariva:2009:GGC

[SV09] Alvise Sommariva and Marco Vianello. Gauss–Green cubature and moment computation over arbitrary geometries. Journal of Computational and Applied Mathematics, 231(2):886–896, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709003173.

Svacek:2008:AFE

[Svá08a] Petr Svácek. Application of finite element method in aeroelasticity. Journal of Computational and Applied Mathematics, 215(2):586–594, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007709.

Svacek:2008:ANN

[Svá08b] Petr Svácek. On approximation of non-Newtonian fluid flow by the finite element method. Journal of Computational and Applied Mathematics, 218(1):167–174, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700338X.

Saadati:2009:QAA

[SVC09] Reza Saadati, S. Mansour Vaezpour, and Yeol J. Cho. Quicksort algorithm: Application of a fixed point theorem in intuitionistic fuzzy quasi-metric spaces at a domain of words. Journal of Computational and Applied Mathematics, 228 (1):219–225, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004688.

Soderlind:2006:ATS

[SW06a] Gustaf Söderlind and Lina Wang. Adaptive timestepping and computational stability. Journal of Computational and Applied Mathematics, 185(2):225–243, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042705001123.

Soderlind:2006:ENO

[SW06b] Gustaf Söderlind and Lina Wang. Evaluating numerical ODE/DAE methods, algorithms and software. Journal of Computational and Applied Mathematics, 185(2): 244–260, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001135.

Shen:2008:GOS

[SW08] Pei-Ping Shen and Chun-Feng Wang. Global optimization for sum of generalized fractional functions. Journal of Computational and Applied Mathematics, 214 (1):1–12, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000763.

Swaminathan:2006:ITC

[Swa06] A. Swaminathan. Inclusion theorems of convolution operators associated with normalized hypergeometric functions. Journal of Computational and Applied Mathematics, 197(1):15–28, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006308.

Sweilam:2007:HWG

[Swe07a] N. H. Sweilam. Harmonic wave generation in non linear thermoelasticity by variational iteration method and Adomian’s method. Journal of Computational and Applied Mathematics, 207 (1):64–72, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004626.

Sweilam:2007:VIM

[Swe07b]  N. H. Sweilam.  Vari-

ational iteration method for solving cubic nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 207(1): 155–163, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004717.

Shao:2009:PSK

[SWYZ09] Jianying Shao, Lijuan Wang, Yuehua Yu, and Jinglei Zhou. Periodic solutions for a kind of Liénard equation with a deviating argument. Journal of Computational and Applied Mathematics, 228 (1):174–181, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004603.

Saucez:2005:AML

[SWZ05] P. Saucez, A. Vande Wouwer, and P. A. Zegeling. Adaptive method of lines solutions for the extended fifth-order Korteweg–de Vries equation. Journal of Computational and Applied Mathematics, 183(2):343–357, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000701.

Song:2009:RAS

[SWZ09] Lina Song, Qi Wang, and Hongqing Zhang. Rational approximation solution of the fractional Sharma–Tasso– Olever equation. Journal of Computational and Applied Mathematics, 224(1): 210–218, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001921.

Shi:2009:CST

[SX09] Zhen-Jun Shi and Zhiwei Xu.

The convergence of subspace trust region methods. Journal of Computational and Applied Mathematics, 231(1): 365–377, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001381.

Sun:2009:AIM

[SY09] Tongjun Sun and Yirang Yuan. An approximation of incompressible miscible displacement in porous media by mixed finite element method and characteristics-mixed finite element method. Journal of Computational and Applied Mathematics, 228 (1):391–411, June 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005050.

Sezer:2008:ASM

[SyS08] Mehmet Sezer, Salih yalçinbas, and Niyazi Sahin. Approximate solution of multipantograph equation with variable coefficients. Journal of Computational and Applied Mathematics, 214 (2):406–416, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001367.

Sogabe:2007:CMS

[SZ07a] Tomohiro Sogabe and Shao-Liang Zhang. A COCR method for solving complex symmetric linear systems. Journal of Computational and Applied Mathematics, 199(2):297–303, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007648.

Sun:2007:IDD

[SZ07b] Bo Sun and Yi Zhao. Impact of dispersion on dynamics of a discrete metapopulation model. Journal of Computational and Applied Mathematics, 200(1):266–275, March 1, 2007. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008113.

Shen:2008:OCF

[SZ08a] Jianhua Shen and Ziran Zou.

Oscillation criteria for firstorder impulsive differential equations with positive and negative coefficients. Journal of Computational and Applied Mathematics, 217 (1):28–37, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003123.

See corrigendum [SZ09b].

Slodicka:2008:TDS

[SZ08b] Marián Slodicka and Viera Zemanová. Time-discretization scheme for quasi-static Maxwell’s equations with a non-linear boundary condition. Journal of Computational and Applied Mathematics, 216 (2):514–522, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002919.

She:2009:CNB

[SZ09a] Zhikun She and Zhiming Zheng. Condition number based complexity estimate for computing local extrema. Journal of Computational and Applied Mathematics, 230(1):233–242, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005943.

Shen:2009:COC

[SZ09b] Jianhua Shen and Ziran Zou. Corrigendum to: “Oscillation criteria for first-order impulsive differential equations with positive and negative coefficients” [J. Comput. Appl. Math. 217 (1) (2008) 28– 37]. Journal of Computational and Applied Mathematics, 224(2):777–778, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002987.

See [SZ08a].

Szafnicki:2005:DEB

[Sza05a] Boleslaw Szafnicki. On the degree elevation of Bernstein polynomial representation. Journal of Computational and Applied Mathematics, 180(2):443–459, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005448.

Szafraniec:2005:OAP

[Sza05b] Franciszek Hugon Szafraniec. Orthogonality of analytic polynomials: a little step further. Journal of Computational and Applied Mathematics, 179(1–2):343–353, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004583.

Szatanik:2008:QSG

[Sza08] Wawrzyniec Szatanik. Quasisolutions for generalized second order differential equations with deviating arguments. Journal of Computational and Applied Mathematics, 216(2):425–434, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002737.

Shidfar:2009:ASC

[SZD09] A. Shidfar, R. Zolfaghari, and J. Damirchi. Application of sinc-collocation method for solving an inverse problem. Journal of Computational and Applied Mathematics, 233(2):545–554, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900466X.

Toutounian:2009:NMC

[TA09] F. Toutounian and A. Ataei.

A new method for computing Moore–Penrose inverse matrices. Journal of Computational and Applied Mathematics, 228(1):412–417, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005062.

Tabata:2007:DBT

[Tab07] Masahisa Tabata. Discrepancy between theory and real computation on the stability of some finite element schemes. Journal of Computational and Applied Mathematics, 199(2): 424–431, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500782X.

Tachikawa:2007:SLP

[Tac07] Masashi Tachikawa. Specific locking in populations dynamics: Symmetry analysis for coupled heteroclinic cycles. Journal of Computational and Applied Mathematics, 201(2):374–380, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000902.

Tang:2006:EAT

[Tan06]  Hui-Chin Tang.  An ex-

haustive analysis of two-term multiple recursive random number generators with efficient multipliers. Journal of Computational and Applied Mathematics, 192(2):

411–416, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003596.

Trong:2009:DBF

[TANT09] Dang Duc Trong, Pham Ngoc Dinh Alain, Phan Thanh Nam, and Truong Trung Tuyen. Determination of the body force of a twodimensional isotropic elastic body. Journal of Computational and Applied Mathematics, 229(1):192–207, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005566.

Tarabia:2009:ARW

[Tar09] Ahmed M. K. Tarabia. Analysis of random walks with an absorbing barrier and chemical rule. Journal of Computational and Applied Mathematics, 225(2):612–620, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004330.

Tatsien:2006:EBC

[Tat06] Li Tatsien. Exact boundary controllability for quasilinear wave equations. Journal of Computational and Applied Mathematics, 190(1– 2):127–135, June 1, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002256.

Taylor:2008:ACF

[Tay08] Mark A. Taylor. Asymmetric cubature formulas for polynomial integration in the triangle and square. Journal of Computational and Applied Mathematics, 218(1): 184–191, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003408.

Tian:2007:SDD

[TBBC07] Tianhai Tian, Kevin Burrage, Pamela M. Burrage, and Margherita Carletti. Stochastic delay differential equations for genetic regulatory networks. Journal of Computational and Applied Mathematics, 205(2):696–707, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003943.

Theofilatos:2009:CES

[TBL09] Konstantinos Theofilatos, Grigorios Beligiannis, and Spiridon Likothanassis. Combining evolutionary and stochastic gradient techniques for system identification. Journal of Computational and Applied Mathematics, 227(1):147–160, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003488.

Takeuchi:2006:PDP

[TCMS06] Yasuhiro Takeuchi, Jing’an Cui, Rinko Miyazaki, and Yasuhisa Saito. Permanence of dispersal population model with time delays. Journal of Computational and Applied Mathematics, 192(2): 417–430, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003602.

Tunga:2006:HHD

[TD06] M. Alper Tunga and Metin Demiralp. Hybrid high dimensional model representation (HHDMR) on the partitioned data. Journal of Computational and Applied Mathematics, 185(1): 107–132, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500052X.

Tatari:2007:CHV

[TD07] Mehdi Tatari and Mehdi Dehghan. On the convergence of He’s variational iteration method. Journal of Computational and Applied Mathematics, 207(1):121–128, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004675.

Temme:2007:AME

[Tem07] N. M. Temme. Analytical methods for an elliptic singular perturbation problem in a circle. Journal of Computational and Applied Mathematics, 207(2):301–322, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006054.

Temir:2009:CTI

[Tem09]  Seyit Temir.  On the

convergence theorems of implicit iteration process for a finite family of Iasymptotically nonexpansive mappings. Journal of Computational and Applied Mathematics, 225(2): 398–405, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003919.

Terwilliger:2005:TLT

[Ter05] Paul Terwilliger. Two linear transformations each tridiagonal with respect to an eigenbasis of the other: comments on the split decomposition. Journal of Computational and Applied Mathematics, 178(1–2):437–452, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003905.

Tapamo:2007:TAM

[TF07] H. Tapamo and A. Frommer.

Two acceleration mechanisms in verified global optimization. Journal of Computational and Applied Mathematics, 199(2):390–396, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007776.

Toth:2007:ECS

[TFC07] B. Tóth, J. Fernández, and T. Csendes. Empirical convergence speed of inclusion functions for facility location problems. Journal of Computational and Applied Mathematics, 199(2): 384–389, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007764.

Tian:2007:FOC

[TG07] Z. F. Tian and Y. B.

Ge. A fourth-order compact ADI method for solving two-dimensional unsteady convection–diffusion problems. Journal of Computational and Applied Mathematics, 198(1):268–286, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007909.

Tao:2009:TDM

[TG09] Youtian Tao and Chuanqing Gu. A two-dimensional matrix Padé-type approximation in the inner product space. Journal of Computational and Applied Mathematics, 231(2):680–695, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002684.

Tangman:2008:FHO

[TGB08a] D. Y. Tangman, A. Gopaul, and M. Bhuruth. A fast highorder finite difference algorithm for pricing American options. Journal of Computational and Applied Mathematics, 222(1):17–29, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005468.

Tangman:2008:NPO

[TGB08b] D. Y. Tangman, A. Gopaul, and M. Bhuruth. Numerical pricing of options using high-order compact finite difference schemes. Journal of Computational and Applied Mathematics, 218(2): 270–280, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700043X.

Tahmasebi:2019:CSC

[TGJ19] Mahdieh Tahmasebi, Azadeh Ghasemifard, and Mohammad Taghi Jahandideh. Comments on “Strong convergence rates for backward Euler on a class of nonlinear jump-diffusion problems” [j. comput. appl. math. 205 (2007) 949–956]. Journal of Computational and Applied Mathematics, 359(??): 69–72, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0377042719301633.

See [HK07b].

Tavazoei:2007:OAB

[TH07] Mohammad Saleh Tavazoei and Mohammad Haeri. An optimization algorithm based on chaotic behavior and fractal nature. Journal of Computational and Applied Mathematics, 206(2):1070– 1081, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005796.

Tian:2008:CGA

[THC08] Gui-Xian Tian, Ting-Zhu Huang, and Shu-Yu Cui. Convergence of generalized AOR iterative method for linear systems with strictly diagonally dominant matrices. Journal of Computational and Applied Mathematics, 213(1):240–247, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000313.

Thi:2005:PSP

[Thi05] N. N. Pham Thi. On positive solutions in a phytoplanktonnutrient model. Journal of Computational and Applied Mathematics, 177(2): 467–473, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005540.

Thianwan:2009:CFP

[Thi09] Sornsak Thianwan. Common fixed points of new iterations for two asymptotically nonexpansive nonself-mappings in a Banach space. Journal of Computational and Applied Mathematics, 224(2): 688–695, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800294X.

Thi:2005:STD

[THS05] N. N. Pham Thi, J. Huisman, and B. P. Sommeijer. Simulation of three-dimensional phytoplankton dynamics: competition in light-limited environments. Journal of Computational and Applied Mathematics, 174(1):57–77, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001712.

Tang:2008:SQC

[TJ08] Chun-Ming Tang and Jin-Bao Jian. A sequential quadratically constrained quadratic programming method with an augmented Lagrangian line search function. Journal of Computational and Applied Mathematics, 220(1–2): 525–547, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004840.

Tiwari:2007:MTP

[TK07] Sudarshan Tiwari and Jörg Kuhnert. Modeling of twophase flows with surface tension by finite pointset method (FPM). Journal of Computational and Applied Mathematics, 203(2):376–386, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002275.

Tian:2007:CRW

[TL07] Jianjun Paul Tian and Zhenqiu Liu. Coalescent random walks on graphs. Journal of Computational and Applied Mathematics, 202 (1):144–154, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001178.

Tapia:2009:APS

[TL09] Juan J. Tapia and P. Gilberto López. Adaptive pseudospectral solution of a diffuse interface model. Journal of Computational and Applied Mathematics, 224(1): 101–117, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001805.

Tong:2008:SIP

[TLQ08] Xiaojiao Tong, Chen Ling, and Liqun Qi. A semi-infinite programming algorithm for solving optimal power flow with transient stability constraints. Journal of Computational and Applied Mathematics, 217(2):432–447, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001008.

Li:2007:IPA

[tLxLlW07] Yao tang Li, Cui xia Li, and Shi liang Wu. Improvements of preconditioned AOR iterative method for L-matrices. Journal of Computational and Applied Mathematics, 206(2):656–665, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005127.

See note [Yun08b].

Twining:2008:CAD

[TM08] Carole J. Twining and Stephen Marsland. Constructing an atlas for the diffeomorphism group of a compact manifold with boundary, with application to the analysis of image registrations. Journal of Computational and Applied Mathematics, 222(2):411–428, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006188.

Triantafyllou:2009:CRB

[TM09] Dimitrios Triantafyllou and Marilena Mitrouli. On the computation of the rank of block bidiagonal Toeplitz matrices. Journal of Computational and Applied Mathematics, 227(1):126–135, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800335X.

Tokarzewski:2009:ESF

[TMG09] Stanislaw Tokarzewski, Alphonse Ph. Magnus, and Jacek Gilewicz. Estimation of a Stieltjes function expanded to Taylor series at complex conjugate points. Journal of Computational and Applied Mathematics, 233(3):835–841, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001320.

Takeuchi:2007:P

[TMMS07] Yasuhiro Takeuchi, R. Miyazaki, S. Morita, and Y. Saito. Preface. Journal of Computational and Applied Mathematics, 201(2):315–316, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000835.

Taboada:2007:CQM

[TMOG07] J. Taboada, J. M. Matías, C. Ordóñez, and P. J. García. Creating a quality map of a slate deposit using support vector machines. Journal of Computational and Applied Mathematics, 204 (1):84–94, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600241X.

Tsolis:2009:ARM

[TND^{+}09] Dimitrios Tsolis, Spiridon Nikolopoulos, Lambros Drossos, Spyros Sioutas, and Theodore Papatheodorou. Applying robust multibit watermarks to digital images. Journal of Computational and Applied Mathematics, 227 (1):213–220, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003543.

Tocino:2005:MSS

[Toc05] A. Tocino. Mean-square stability of second-order Runge– Kutta methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 175(2): 355–367, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002390.

Tocino:2009:SOW

[Toc09] A. Tocino. Simplified order 4.0 weak Taylor schemes for additive noise. Journal of Computational and Applied Mathematics, 231(1): 154–159, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000491.

Toksari:2007:HAF

[Tok07] M. Duran Toksari. A heuristic approach to find the global optimum of function. Journal of Computational and Applied Mathematics, 209(2): 160–166, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006728.

Torney:2005:BAB

[Tor05] David C. Torney. Bayesian analysis of binary sequences. Journal of Computational and Applied Mathematics, 175(2):231–243, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002304.

Torre:2008:GZD

[Tor08] A. Torre. Generalized Zernike or disc polynomials: an application in quantum optics. Journal of Computational and Applied Mathematics, 222(2):622–644, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006450.

Toyonaga:2007:MSN

[Toy07] K. Toyonaga. A method for separating nearly multiple eigenvalues for Hermitian matrix. Journal of Computational and Applied Mathematics, 199(2):432–436, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007831.

Trong:2006:DTD

[TQA06] D. D. Trong, P. H. Quan, and P. N. Dinh Alain. Determination of a two-dimensional heat source: Uniqueness, regularization and error estimate. Journal of Computational and Applied Mathematics, 191(1):50–67, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500261X.

Teng:2006:PNP

[TR06] Zhidong Teng and Mehbuba Rehim. Persistence in nonautonomous predator–prey systems with infinite delays. Journal of Computational and Applied Mathematics, 197(2):302–321, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500662X.

Teofanov:2007:ESP

[TR07] Lj. Teofanov and H.-G. Roos. An elliptic singularly perturbed problem with two parameters. I: Solution decomposition. Journal of Computational and Applied Mathematics, 206(2):1082– 1097, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005802.

Teofanov:2008:ESP

[TR08] Lj. Teofanov and H.-G. Roos. An elliptic singularly perturbed problem with two parameters II: Robust finite element solution. Journal of Computational and Applied Mathematics, 212(2): 374–389, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007370.

Tomar:2009:ECE

[TR09] S. K. Tomar and S. I. Repin. Efficient computable error bounds for discontinuous Galerkin approximations of elliptic problems. Journal of Computational and Applied Mathematics, 226 (2):358–369, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004159.

Tamilselvan:2007:NMS

[TRS07] A. Tamilselvan, N. Ramanujam, and V. Shanthi. A numerical method for singularly perturbed weakly coupled system of two second order ordinary differential equations with discontinuous source term. Journal of Computational and Applied Mathematics, 202(2):203–216, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001221.

Tari:2009:SCT

[TRST09] A. Tari, M. Y. Rahimi, S. Shahmorad, and F. Talati. Solving a class of two-dimensional linear and nonlinear Volterra integral equations by the differential transform method. Journal of Computational and Applied Mathematics, 228 (1):70–76, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800438X.

See comments [Jan09].

Tselios:2005:RKM

[TS05] Kostas Tselios and T. E.

Simos. Runge–Kutta methods with minimal dispersion and dissipation for problems arising from computational acoustics. Journal of Computational and Applied Mathematics, 175(1): 173–181, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002511.

Tang:2006:TTM

[TS06a] Lingyan Tang and Songhe Song. A TVD-type method for 2D scalar Hamilton– Jacobi equations on unstructured meshes. Journal of Computational and Applied Mathematics, 195(1–2): 182–191, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004838.

Tian:2006:REI

[TS06b] Yongge Tian and George P. H. Styan. Rank equalities for idempotent matrices with applications. Journal of Computational and Applied Mathematics, 191(1):77–97, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002633.

Tang:2008:MFV

[TS08] Lingyan Tang and Songhe Song. A multiresolution finite volume scheme for two-dimensional hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 214 (2):583–595, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001562.

Tsallis:2009:CAN

[Tsa09] Constantino Tsallis. Computational applications of nonextensive statistical mechanics. Journal of Computational and Applied Mathematics, 227(1):51–58, May 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003282.

Tezer-Sezgin:2009:BES

[TSB09] M. Tezer-Sezgin and C. Bozkaya.

The boundary element solution of the magnetohydrodynamic flow in an infinite region. Journal of Computational and Applied Mathematics, 225(2):510–521, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004238.

Tian:2010:CES

[TSC10] Yuan Tian, Kaibiao Sun, and Lansun Chen. Comment on “Existence and stability of periodic solution of a Lotka–Volterra predator– prey model with state dependent impulsive effects” [J. Comput. Appl. Math. 224 (2009) 544–555]. Journal of Computational and Applied Mathematics, 234(10):2916– 2923, September 15, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042710001986.

See [NPTH09].

Tsitouras:2006:SRS

[Tsi06] Ch. Tsitouras. Stage reduction on P -stable Numerov type methods of eighth order. Journal of Computational and Applied Mathematics, 191(2):297–305, July 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005571.

Tan:2006:SMM

[TTZ06] Zhijun Tan, Tao Tang, and Zhengru Zhang. A simple moving mesh method for one- and two-dimensional phase-field equations. Journal of Computational and Applied Mathematics, 190(1– 2):252–269, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002347.

Tupper:2007:CSH

[Tup07] P. F. Tupper. Computing statistics for Hamiltonian systems: a case study. Journal of Computational and Applied Mathematics, 205(2): 826–834, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004055.

Turovtsev:2006:ESD

[Tur06] G. V. Turovtsev. An expansion of the solution of Dirichlet boundary value problem for Berger equation. Journal of Computational and Applied Mathematics, 193 (1):1–9, August 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003754.

Turaev:2009:DFS

[Tur09] Mamasali Turaev. Decomposition formulas for Srivastava’s hypergeometric function H_{A} on Saran functions. Journal of Computational and Applied Mathematics, 233(3):842–846, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001332.

Tang:2007:DSS

[TV07] J. M. Tang and C. Vuik. On deflation and singular symmetric positive semidefinite matrices. Journal of Computational and Applied Mathematics, 206(2): 603–614, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005073.

Temme:2009:AER

[TV09a] Nico M. Temme and Vladimir Varlamov. Asymptotic expansions for Riesz fractional derivatives of Airy functions and applications. Journal of Computational and Applied Mathematics, 232(2): 201–215, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003410.

Tsompanopoulou:2009:AIR

[TV09b] P. Tsompanopoulou and E. Vavalis. Analysis of an interface relaxation method for composite elliptic differential equations. Journal of Computational and Applied Mathematics, 226(2):370–387, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004160.

Thangavelu:2007:RTR

[TX07] Sundaram Thangavelu and Yuan Xu. Riesz transform and Riesz potentials for Dunkl transform. Journal of Computational and Applied Mathematics, 199(1): 181–195, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007491.

Tian:2009:NDD

[TXZ09] Junkang Tian, Dongsheng Xu, and Jian Zu. Novel delay-dependent asymptotic stability criteria for neural networks with time-varying delays. Journal of Computational and Applied Mathematics, 228(1):133–138, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800455X.

Tsai:2005:VVP

[TY05] Chiung-Chiou Tsai and Suh-Yuh Yang. On the velocityvorticity-pressure least-squares finite element method for the stationary incompressible Oseen problem. Journal of Computational and Applied Mathematics, 182(1): 211–232, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006120.

Tian:2007:SPS

[TY07a] Lixin Tian and Jiuli Yin. Shock-peakon and shockcompacton solutions for K(p, q) equation by variational iteration method. Journal of Computational and Applied Mathematics, 207 (1):46–52, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004596.

Tsai:2007:ASM

[TY07b] Chiung-Chiou Tsai and Suh-Yuh Yang. Analysis of a splitting method for incompressible inviscid rotational flow problems. Journal of Computational and Applied Mathematics, 200 (1):364–376, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000215.

Tang:2008:DTG

[TYH08] Yinghui Tang, Xi Yun, and Shujuan Huang. Discretetime Geo^{X} /G/1 queue with unreliable server and multiple adaptive delayed vacations. Journal of Computational and Applied Mathematics, 220(1–2):439–455, October 15, 2008. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004645.

Tang:2005:AOI

[TYZ^{+}05] Chen Tang, Haiqing Yan, Hao Zhang, Zhanqing Chen, Ming Liu, and Guimin Zhang. The arbitrary order implicit multistep schemes of exponential fitting and their applications. Journal of Computational and Applied Mathematics, 173(1):155–168, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001487.

Utyuzhnikov:2009:MGW

[UFG09] S. V. Utyuzhnikov, P. Fantini, and M. D. Guenov. A method for generating a well-distributed Pareto set in nonlinear multiobjective optimization. Journal of Computational and Applied Mathematics, 223(2): 820–841, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000861.

Ushijima:2009:FIP

[UHNC09] Teruo Ushijima, Kyousuke Handa, Homare Nishizaki, and Fumihiro Chiba. An FEM identification procedure for a wing section within a class of finite Laurent series. Journal of Computational and Applied Mathematics, 232(1):17–45, October 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005426.

Siraj-ul-Islam:2009:MMN

[uIHA09] Siraj ul Islam, Sirajul Haq, and Arshed Ali. A meshfree method for the numerical solution of the RLW equation. Journal of Computational and Applied Mathematics, 223(2):997–1012, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001416.

Ujevic:2007:IMS

[Uje07] Nenad Ujević. An iterative method for solving nonlinear equations. Journal of Computational and Applied Mathematics, 201(1):208–216, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001002.

Ueno:2009:NSS

[UO09] Toshihide Ueno and Masami Okada. Non-separable splines and numerical computation of evolution equations by the Galerkin methods. Journal of Computational and Applied Mathematics, 223(1): 159–176, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000058.

Vigo-Aguiar:2006:P

[VA06]  J. Vigo-Aguiar.  Preface.

Journal of Computational and Applied Mathematics, 192(1):1, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003018.

Valami:2009:GPE

[Val09] H. Bagherzadeh Valami.

Group performance evaluation, an application of data envelopment analysis. Journal of Computational and Applied Mathematics, 230(2): 485–490, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006456.

Vigo-Aguiar:2005:SEF

[VAMVC05] J. Vigo-Aguiar, J. Martín-Vaquero, and R. Criado. On the stability of exponential fitting BDF algorithms. Journal of Computational and Applied Mathematics, 175 (1):183–194, March 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704002493.

VandeVyver:2005:FEE

[Van05] Hans Van de Vyver. Frequency evaluation for exponentially fitted Runge– Kutta methods. Journal of Computational and Applied Mathematics, 184(2): 442–463, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000415.

VandeVyver:2006:GSE

[Van06a] Hans Van de Vyver. On the generation of P -stable exponentially fitted Runge– Kutta–Nyström methods by exponentially fitted Runge– Kutta methods. Journal of Computational and Applied Mathematics, 188(2): 309–318, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002578.

See erratum [Van07a].

Vigo-Aguiar:2006:ENM

[VAN06b] J. Vigo-Aguiar and S. Natesan. An efficient numerical method for singular perturbation problems. Journal of Computational and Applied Mathematics, 192 (1):132–141, July 15, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003134.

VandeVyver:2007:EGS

[Van07a] Hans Van de Vyver. Erratum to “On the generation of P -stable exponentially fitted Runge–Kutta–Nyström methods by exponentially fitted Runge–Kutta methods”. Journal of Computational and Applied Mathematics, 200(2):778–779, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000987.

See [Van06a].

VandeVyver:2007:PFA

[Van07b] Hans Van de Vyver. Phasefitted and amplification-fitted two-step hybrid methods for y′′ = f(x, y). Journal of Computational and Applied Mathematics, 209(1): 33–53, December 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006182.

VanIseghem:2008:SFT

[Van08] Jeannette Van Iseghem.

Shohat–Favard type theorem for orthogonal series. Journal of Computational and Applied Mathematics, 219(2):

537–550, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002750.

VandeVyver:2009:STS

[Van09] Hans Van de Vyver. Scheifele two-step methods for perturbed oscillators. Journal of Computational and Applied Mathematics, 224(1): 415–432, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002483.

Vigo-Aguiar:2006:VSI

[VAR06] Jesús Vigo-Aguiar and Higinio Ramos. Variable stepsize implementation of multistep methods for y′′ = f(x, y, y′). Journal of Computational and Applied Mathematics, 192(1):114–131, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003122.

Varlamov:2007:SFA

[Var07] Vladimir Varlamov. Special functions arising in the study of semi-linear equations in circular domains. Journal of Computational and Applied Mathematics, 202 (1):105–121, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001191.

Vigo-Aguiar:2007:SID

[VAW07] J. Vigo-Aguiar and B. A.

Wade. Special issue dedicated to Professor shinnosuke oharu on the occasion of his 65th birthday. Journal of Computational and Applied Mathematics, 204(1):1–2, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002329.

VanDeun:2005:COR

[VB05a] J. Van Deun and A. Bultheel.

The computation of orthogonal rational functions on an interval. Journal of Computational and Applied Mathematics, 179(1–2):355–373, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004595.

VanDeun:2005:WSC

[VB05b] J. Van Deun and A. Bultheel.

A weak-star convergence result for orthogonal rational functions. Journal of Computational and Applied Mathematics, 178(1– 2):453–464, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003930.

VandenBerghen:2005:CNP

[VB05c] Frank Vanden Berghen and Hugues Bersini. CONDOR, a new parallel, constrained extension of Powell’s UOBYQA algorithm: Experimental results and comparison with the DFO algorithm. Journal of Computational and Applied Mathematics, 181(1): 157–175, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270400559X.

Vandewoestyne:2006:GPD

[VC06] Bart Vandewoestyne and Ronald Cools. Good permutations for deterministic scrambled Halton sequences in terms of L_{2}-discrepancy. Journal of Computational and Applied Mathematics, 189(1–2):341–361, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003717.

Vanduffel:2008:OAR

[VCD^{+}08] S. Vanduffel, X. Chen, J. Dhaene, M. Goovaerts, L. Henrard, and R. Kaas. Optimal approximations for risk measures of sums of lognormals based on conditional expectations. Journal of Computational and Applied Mathematics, 221(1): 202–218, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700581X.

vanDiejen:2005:AFK

[vD05] J. F. van Diejen. An asymptotic formula for the Koornwinder polynomials. Journal of Computational and Applied Mathematics, 178(1– 2):465–471, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003942.

vandenBoogaart:2007:KBM

[vdBHPS07] K. G. van den Boogaart, R. Hielscher, J. Prestin, and H. Schaeben. Kernel-based methods for inversion of the Radon transform on SO(3) and their applications to texture analysis. Journal of Computational and Applied Mathematics, 199(1): 122–140, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007442.

vandenEshof:2005:RSN

[vdESvG05] Jasper van den Eshof, Gerard L. G. Sleijpen, and Martin B. van Gijzen. Relaxation strategies for nested Krylov methods. Journal of Computational and Applied Mathematics, 177(2):347–365, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004303.

Vanmaele:2006:BPD

[VDL^{+}06] M. Vanmaele, G. Deelstra, J. Liinev, J. Dhaene, and M. J. Goovaerts. Bounds for the price of discrete arithmetic Asian options. Journal of Computational and Applied Mathematics, 185 (1):51–90, January 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000506.

Vierendeels:2008:PSC

[VDV08] J. Vierendeels, K. Dumont, and P. R. Verdonck. A partitioned strongly coupled fluidstructure interaction method to model heart valve dynamics. Journal of Computational and Applied Mathematics, 215(2):602–609, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007722.

vanDoorn:2009:REZ

[vDvFZ09] Erik A. van Doorn, Nicky D. van Foreest, and Alexander I.

Zeifman. Representations for the extreme zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 233(3): 847–851, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001344.

See corrigendum [vDvFZ13].

vanDoorn:2013:CRE

[vDvFZ13] Erik A. van Doorn, Nicky D.

van Foreest, and Alexander I. Zeifman. Corrigendum to “Representations for the extreme zeros of orthogonal polynomials” [J. Comput. Appl. Math. 233 (2009) 847–851]. Journal of Computational and Applied Mathematics, 244(??):155–156, May 15, 2013. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042712005225.

See [vDvFZ09].

Verotta:2005:MEM

[Ver05] Davide Verotta. Models and estimation methods for clinical HIV-1 data. Journal of Computational and Applied Mathematics, 184(1): 275–300, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000816.

Verner:2006:SMT

[Ver06] J. H. Verner. Starting methods for two-step Runge– Kutta methods of stage-order 3 and order 6. Journal of Computational and Applied Mathematics, 185(2): 292–307, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705001160.

Vercher:2008:PFR

[Ver08] Enriqueta Vercher. Portfolios with fuzzy returns: Selection strategies based on semiinfinite programming. Journal of Computational and Applied Mathematics, 217(2): 381–393, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000970.

Vazquez:2009:ESI

[VG09] C. García Vázquez and F. Ortegón Gallego. An elliptic system involving a singular diffusion matrix. Journal of Computational and Applied Mathematics, 229(2):452–461, July 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001659.

Valent:2009:OJP

[VGT09] Galliano Valent, Jacek Gilewicz, and Roland Triay. OPSFA09 (July 2–6, 2007): Proceedings. Journal of Computational and Applied Mathematics, 233(3):591–596, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001009.

Vidunas:2005:TSG

[Vid05] Raimundas Vidunas. Transformations of some Gauss hypergeometric functions. Journal of Computational and Applied Mathematics, 178(1– 2):473–487, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004625.

Villatoro:2009:GOS

[Vil09] Francisco R. Villatoro. Given a one-step numerical scheme, on which ordinary differential equations is it exact? Journal of Computational and Applied Mathematics, 223(2): 1058–1065, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001453.

Vinogradova:2009:CEP

[Vin09] Polina Vinogradova. Convergence estimates of a projection-difference method for an operator-differential equation. Journal of Computational and Applied Mathematics, 231(1):1–10, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709000557.

Vermolen:2009:CAS

[VJ09] F. J. Vermolen and E. Javierre. On the construction of analytic solutions for a diffusionreaction equation with a discontinuous switch mechanism. Journal of Computational and Applied Mathematics, 231(2):983–1003, September 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900329X.

Vasuki:2006:CIR

[VK06] K. R. Vasuki and B. R. Srivatsa Kumar. Certain identities for Ramanujan–Göllnitz– Gordon continued fraction. Journal of Computational and Applied Mathematics, 187(1):87–95, March 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001470.

Voss:2008:SIN

[VK08] David A. Voss and Abdul Q. M. Khaliq. Special issue: Numerical PDE methods in finance: Guest Editors’ foreword. Journal of Computational and Applied Mathematics, 222(1):1–2, December 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005444.

Vlcek:2006:SLM

[VL06] Jan Vlcek and Ladislav Luksan. Shifted limited-memory variable metric methods for large-scale unconstrained optimization. Journal of Computational and Applied Mathematics, 186(2):365– 390, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001007.

Verheyden:2008:ECC

[VLR08] Koen Verheyden, Tatyana Luzyanina, and Dirk Roose. Efficient computation of characteristic roots of delay differential equations using LMS methods. Journal of Computational and Applied Mathematics, 214(1):209–226, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001094.

Vaissmoradi:2009:EAA

[VMMM09] N. Vaissmoradi, A. Malek, and S. H. Momeni-Masuleh. Error analysis and applications of the Fourier–Galerkin Runge–Kutta schemes for high-order stiff PDEs. Journal of Computational and Applied Mathematics, 231(1): 124–133, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000466.

Vandebril:2007:LLA

[VMV07] Raf Vandebril, Nicola Mastronardi, and Marc Van Barel. A Levinson-like algorithm for symmetric strongly nonsingular higher order semiseparable plus band matrices. Journal of Computational and Applied Mathematics, 198(1):75–97, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007077.

Vazquez:2009:MAA

[VN09] Ariel Almendral Vázquez and Bjørn Fredrik Nielsen. The multigrid algorithm applied to a degenerate equation: a convergence analysis. Journal of Computational and Applied Mathematics, 225 (1):251–267, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003671.

Volk:2005:ESI

[Vol05] Wolfgang Volk. Evaluation of the solution of an integralfunctional equation. Journal of Computational and Applied Mathematics, 174(2): 423–436, February 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002274.

Volkmer:2008:AES

[Vol08] Hans Volkmer. Approximation of eigenvalues of some differential equations by zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 213(2):488–500, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000672.

vanOs:2007:CSE

[vOP07] R. G. M. van Os and T. N.

Phillips. The choice of spectral element basis functions in domains with an axis of symmetry. Journal of Computational and Applied Mathematics, 201(1):217–229, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001014.

Vandekerckhove:2008:AAA

[VR08] Christophe Vandekerckhove and Dirk Roose. Accuracy analysis of acceleration schemes for stiff multiscale problems. Journal of Computational and Applied Mathematics, 211(2): 181–200, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007047.

Valdettaro:2007:CRE

[VRBF07] Lorenzo Valdettaro, Michel Rieutord, Thierry Braconnier, and Valérie Frayssé. Convergence and round-off errors in a two-dimensional eigenvalue problem using spectral methods and Arnoldi– Chebyshev algorithm. Journal of Computational and Applied Mathematics, 205(1): 382–393, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003256.

Vandekerckhove:2007:NSA

[VRL07] Christophe Vandekerckhove, Dirk Roose, and Kurt Lust. Numerical stability analysis of an acceleration scheme for step size constrained time integrators. Journal of Computational and Applied Mathematics, 200(2): 761–777, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000744.

Vazquez:2008:GSI

[VRSS08] F. Guerra Vázquez, J.-J.

Rückmann, O. Stein, and G. Still. Generalized semiinfinite programming: a tutorial. Journal of Computational and Applied Mathematics, 217(2):394–419, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000982.

Vejchodsky:2008:SCP

[VS08] Tomás Vejchodský and Pavel Solín. Static condensation, partial orthogonalization of basis functions, and ILU preconditioning in the hp-FEM. Journal of Computational and Applied Mathematics, 218(1):192–200, August 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700341X.

Vegh:2006:HTC

[VT06] V. Vegh and I. W. Turner.

A hybrid technique for computing the power distribution generated in a lossy medium during microwave heating. Journal of Computational and Applied Mathematics, 197(1):122–140, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006473.

Vulanovic:2007:LRT

[Vul07] Relja Vulanović. The layerresolving transformation and mesh generation for quasilinear singular perturbation problems. Journal of Computational and Applied Mathematics, 203(1):177–189, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001932.

Vacus:2005:DSC

[VV05a] Olivier Vacus and Nicolas Vukadinovic. Dynamic susceptibility computations for thin magnetic films. Journal of Computational and Applied Mathematics, 176 (2):263–281, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003322.

Vermolen:2005:SVS

[VV05b] F. J. Vermolen and C. Vuik.

Solution of vector Stefan problems with cross-diffusion. Journal of Computational and Applied Mathematics, 176(1):179–201, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003267.

Vandebril:2006:NNT

[VV06] Raf Vandebril and Marc Van Barel. A note on the nullity theorem. Journal of Computational and Applied Mathematics, 189(1– 2):179–190, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500097X.

VandenBerghe:2007:EFN

[VV07] G. Vanden Berghe and M. Van Daele. Exponentiallyfitted Numerov methods. Journal of Computational and Applied Mathematics, 200(1):140–153, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705008010.

Vandaele:2009:EPU

[VV09a] Nele Vandaele and Michèle Vanmaele. Explicit portfolio for unit-linked life insurance contracts with surrender option. Journal of Computational and Applied Mathematics, 233(1):16–26, November 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001933.

VandenBerghe:2009:TPE

[VV09b] G. Vanden Berghe and M. Van Daele. Trigonometric polynomial or exponential fitting approach? Journal of Computational and Applied Mathematics, 233(4): 969–979, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709005858.

Vaassen:2008:IHO

[VVE08a] J.-M. Vaassen, D. Vigneron, and J.-A. Essers. An implicit high order finite volume scheme for the solution of 3D Navier–Stokes equations with new discretization of diffusive terms. Journal of Computational and Applied Mathematics, 215(2):595–601, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007710.

Vigneron:2008:IFV

[VVE08b] D. Vigneron, J.-M. Vaassen, and J.-A. Essers. An implicit finite volume method for the solution of 3D low Mach number viscous flows using a local preconditioning technique. Journal of Computational and Applied Mathematics, 215(2):610–617, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007734.

Vanberghen:2008:MBS

[VVV08] Yvette Vanberghen, Raf Vandebril, and Marc Van Barel. A QZ-method based on semiseparable matrices. Journal of Computational and Applied Mathematics, 218(2): 482–491, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004104.

Voitcu:2006:CNR

[VW06] Ovidiu Voitcu and Yau Shu Wong. On the construction of a nonlinear recursive predictor. Journal of Computational and Applied Mathematics, 190(1–2):393–407, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002426.

Vinet:2008:GBT

[VZ08] Luc Vinet and Alexei Zhedanov.

Generalized Bochner theorem: Characterization of the Askey–Wilson polynomials. Journal of Computational and Applied Mathematics, 211(1):45–56, January 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006947.

Winkler:2008:STL

[WA08] Joab R. Winkler and John D. Allan. Structured total least norm and approximate GCDs of inexact polynomials. Journal of Computational and Applied Mathematics, 215(1):1–13, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001598.

Waadeland:2005:MCC

[Waa05] Haakon Waadeland. Monotonicity of CF-coefficients in Gauss-fractions. Journal of Computational and Applied Mathematics, 179(1– 2):375–380, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004601.

Wackers:2005:NGD

[Wac05] Jeroen Wackers. A nestedgrid direct Poisson solver for concentrated source terms. Journal of Computational and Applied Mathematics, 180(1):1–12, August 1, 2005.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004649.

Wang:2007:TSS

[Wan07a] Hanquan Wang. A timesplitting spectral method for coupled Gross–Pitaevskii equations with applications to rotating Bose–Einstein condensates. Journal of Computational and Applied Mathematics, 205(1):88–104, August 1, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002573.

Wang:2007:ICO

[Wan07b] Qi-Ru Wang. Interval criteria for oscillation of second-order nonlinear differential equations. Journal of Computational and Applied Mathematics, 205(1):231–238, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002846.

Wang:2007:ESM

[Wan07c] Yuan-Ming Wang. Error and stability of monotone method for numerical solutions of fourth-order semilinear elliptic boundary value problems. Journal of Computational and Applied Mathematics, 200(2):503–519, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000318.

Wang:2008:TSS

[Wan08a] Peng Wang. Three-stage stochastic Runge–Kutta methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 222(2): 324–332, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006012.

Wang:2008:RRM

[Wan08b] Ren-Hong Wang. Recent researches on multivariate spline and piecewise algebraic variety. Journal of Computational and Applied Mathematics, 221(2):460– 471, November 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005778.

Wang:2009:RAB

[Wan09a] Lijuan Wang. RETRACTED: Asymptotic behavior of solutions to a system of differential equations with statedependent delays. Journal of Computational and Applied Mathematics, 228 (1):226–230, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800469X.

Wang:2009:OPU

[Wan09b] Zeng-Qi Wang. Optimization of the parameterized Uzawa preconditioners for saddle point matrices. Journal of Computational and Applied Mathematics, 226 (1):136–154, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002392.

Watson:2006:CHT

[Wat06] G. A. Watson. Computing helmert transformations. Journal of Computational and Applied Mathematics, 197(2):387–394, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006862.

Watson:2007:LMM

[Wat07] G. A. Watson. A Levenberg– Marquardt method for estimating polygonal regions. Journal of Computational and Applied Mathematics, 208(2):331–340, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005863.

Watanabe:2009:CAP

[Wat09]  Yoshitaka Watanabe.  A

computer-assisted proof for the Kolmogorov flows of incompressible viscous fluid. Journal of Computational and Applied Mathematics, 223(2):953–966, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001167.

Wazwaz:2007:CBV

[Waz07a] Abdul-Majid Wazwaz. A comparison between the variational iteration method and Adomian decomposition method. Journal of Computational and Applied Mathematics, 207(1):129–136, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004687.

Wazwaz:2007:VIM

[Waz07b] Abdul-Majid Wazwaz. The variational iteration method for rational solutions for KdV, K(2, 2), Burgers, and cubic Boussinesq equations. Journal of Computational and Applied Mathematics, 207(1):18–23, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004560.

Wang:2009:PIP

[WB09] G. Q. Wang and Y. Q.

Bai. Polynomial interiorpoint algorithms for P_{∗}(κ) horizontal linear complementarity problem. Journal of Computational and Applied Mathematics, 233(2): 248–263, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004166.

Wang:2007:DDE

[WC07] Ying-Ming Wang and Kwai-Sang Chin. Discriminating DEA efficient candidates by considering their least relative total scores. Journal of Computational and Applied Mathematics, 206(1): 209–215, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004262.

Wang:2008:FEP

[WC08] Weichung Wang and Ray-Bing Chen. Finding effective points by surrogate models with overcomplete bases. Journal of Computational and Applied Mathematics, 217(1):110–122, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003184.

Wang:2005:SSL

[WCH05] Yongli Wang, Lifeng Chen, and Guoping He. Sequential systems of linear equations method for general constrained optimization without strict complementarity. Journal of Computational and Applied Mathematics, 182(2): 447–471, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006429.

Wang:2008:ESL

[WCL08] Ji Wang, Jian-Kang Chen, and Shijun Liao. An explicit solution of the large deformation of a cantilever beam under point load at the free tip. Journal of Computational and Applied Mathematics, 212(2):320–330, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007333.

Wang:2009:PAC

[WCW09] Ke Wang, Guoliang Chen, and Yimin Wei. Perturbation analysis for a class of fuzzy linear systems. Journal of Computational and Applied Mathematics, 224(1): 54–65, February 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708001532.

Wang:2009:OMM

[WCY09] Kuo-Hsiung Wang, Wei-Lun Chen, and Dong-Yuh Yang. Optimal management of the machine repair problem with working vacation: Newton’s method. Journal of Computational and Applied Mathematics, 233(2): 449–458, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004518.

Wang:2007:CDM

[WCZ07] Tingchun Wang, Juan Chen, and Luming Zhang. Conservative difference methods for the Klein–Gordon– Zakharov equations. Journal of Computational and Applied Mathematics, 205(1): 430–452, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003281.

Wei:2005:CND

[WD05] Yimin Wei and Huaian Diao. Condition number for the Drazin inverse and the Drazin-inverse solution of singular linear system with their condition numbers. Journal of Computational and Applied Mathematics, 182(2): 270–289, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006156.

Wu:2007:FDE

[WDC07] Chunlin Wu, Jiansong Deng, and Falai Chen. Fast data extrapolating. Journal of Computational and Applied Mathematics, 206(1): 146–157, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003888.

Wyart:2008:SFX

[WDC^{+}08] E. Wyart, M. Duflot, D. Coulon, P. Martiny, T. Pardoen, J.- F. Remacle, and F. Lani. Substructuring FE–XFE approaches applied to threedimensional crack propagation. Journal of Computational and Applied Mathematics, 215(2):626–638, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706007758.

Wei:2009:TPS

[Wei09] Yuming Wei. Triple positive solutions to third order threepoint BVP with increasing homeomorphism and positive homomorphism. Journal of Computational and Applied Mathematics, 231(1): 134–142, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000478.

Wu:2007:DQM

[WeR07] Xionghua Wu and Yu e Ren.

Differential quadrature method based on the highest derivative and its applications. Journal of Computational and Applied Mathematics, 205(1):239–250, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003128.

Wang:2008:UEF

[WF08] Guozhao Wang and Mei’e Fang. Unified and extended form of three types of splines. Journal of Computational and Applied Mathematics, 216(2):498–508, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002890.

Wen:2009:GSC

[WF09] Zijuan Wen and Shengmao Fu. Global solutions to a class of multi-species reaction– diffusion systems with crossdiffusions arising in population dynamics. Journal of Computational and Applied Mathematics, 230 (1):34–43, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005840.

Wang:2007:CPD

[WFL07] Yong Wang, Shu-Cherng Fang, and John E. Lavery. A compressed primal-dual method for generating bivariate cubic L_{1} splines. Journal of Computational and Applied Mathematics, 201 (1):69–87, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000793.

Wang:2009:CPF

[WFW09] Yuncheng Wang, Weiwu Fang, and Tianjiao Wu. A cut-peak function method for global optimization. Journal of Computational and Applied Mathematics, 230(1): 135–142, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005931.

Wahlund:2006:HOD

[WG06] Per Wahlund and Bertil Gustafsson. High-order difference methods for waves in discontinuous media. Journal of Computational and Applied Mathematics, 192 (1):142–147, July 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003158.

Wang:2008:FOC

[WG08] Yuan-Ming Wang and Ben-Yu Guo. Fourth-order compact finite difference method for fourth-order nonlinear elliptic boundary value problems. Journal of Computational and Applied Mathematics, 221(1):76–97, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005262.

Wang:2009:RSE

[WG09] Tingchun Wang and Boling Guo. A robust semi-explicit difference scheme for the Kuramoto–Tsuzuki equation. Journal of Computational and Applied Mathematics, 233(4):878–888, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004531.

Wang:2008:MSM

[WGZ08] Zhong-Qing Wang, Ben-Yu Guo, and Wei Zhang. Mixed spectral method for threedimensional exterior problems using spherical harmonic and generalized Laguerre functions. Journal of Computational and Applied Mathematics, 217(1):277–298, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003585.

Wang:2009:ILM

[WH09a] Wei Wang and Bo Han. An implicit Landweber method for nonlinear ill-posed operator equations. Journal of Computational and Applied Mathematics, 230(2): 607–613, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006559.

Wu:2009:CAW

[WH09b] Shulin Wu and Chengming Huang. Convergence analysis of waveform relaxation methods for neutral differentialfunctional systems. Journal of Computational and Applied Mathematics, 223(1): 263–277, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000113.

Wang:2007:CRP

[WHF07] Xue-Zhong Wang, Ting-Zhu Huang, and Ying-Ding Fu. Comparison results on preconditioned SOR-type iterative method for Z-matrices linear systems. Journal of Computational and Applied Mathematics, 206(2): 726–732, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005188.

Wang:2007:TLE

[WHG07] Hong-Yan Wang, Xing-Biao Hu, and Gegenhasi. 2D Toda lattice equation with selfconsistent sources: Casoratian type solutions, bilinear Bäcklund transformation and Lax pair. Journal of Computational and Applied Mathematics, 202(1):133–143, May 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600118X.

Wang:2008:RKT

[WHL08] W. Wang, B. Han, and L. Li. A Runge–Kutta type modified Landweber method for nonlinear ill-posed operator equations. Journal of Computational and Applied Mathematics, 212(2): 457–468, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706008016.

Wu:2009:MSI

[WHZ09] Shi-Liang Wu, Ting-Zhu Huang, and Xi-Le Zhao. A modified SSOR iterative method for augmented systems. Journal of Computational and Applied Mathematics, 228(1):424–433, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005086.

Wikramaratna:2008:ACR

[Wik08] Roy S. Wikramaratna. The additive congruential random number generator — a special case of a multiple recursive generator. Journal of Computational and Applied Mathematics, 216(2):371–387, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002671.

Wiltshire:2006:TAC

[Wil06] Ron Wiltshire. Two approaches to the calculation of approximate symmetry exemplified using a system of advection–diffusion equations. Journal of Computational and Applied Mathematics, 197(2):287–301, December 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006588.

Willms:2007:PRR

[Wil07]  Allan R. Willms.  Pa-

rameter range reduction for ODE models using cumulative backward differentiation formulas. Journal of Computational and Applied Mathematics, 203(1):87–102, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001762.

Winkler:2006:NCU

[Win06] Joab R. Winkler. The numerical condition of univariate and bivariate degree elevated Bernstein polynomials. Journal of Computational and Applied Mathematics, 191(1):32–49, June 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002608.

Wei:2009:GAP

[WJS^{+}09] Huiming Wei, Yu Jiang, Xinyu Song, G. H. Su, and S. Z. Qiu. Global attractivity and permanence of a SVEIR epidemic model with pulse vaccination and time delay. Journal of Computational and Applied Mathematics, 229(1):302–312, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005761.

Walter:2007:GNP

[WK07a] Éric Walter and Michel Kieffer. Guaranteed nonlinear parameter estimation in knowledge-based models. Journal of Computational and Applied Mathematics, 199(2):277–285, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705007624.

Wu:2007:AMR

[WK07b] Xiaoyong Wu and Xiaowu Ke. Analysis of an M/{D_{n}} /1 retrial queue. Journal of Computational and Applied Mathematics, 200(2): 528–536, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000331.

Wackers:2008:SCM

[WK08] Jeroen Wackers and Barry Koren. A surface capturing method for the efficient computation of steady water waves. Journal of Computational and Applied Mathematics, 215(2):618–625, June 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007746.

Wu:2006:SCMb

[WKL06] Xionghua Wu, Wenbin Kong, and Chen Li. Sinc collocation method with boundary treatment for two-point boundary value problems. Journal of Computational and Applied Mathematics, 196(1): 229–240, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005601.

Wade:2007:SCN

[WKY^{+}07] B. A. Wade, A. Q. M.

Khaliq, M. Yousuf, J. Vigo-Aguiar, and R. Deininger. On smoothing of the Crank– Nicolson scheme and higher order schemes for pricing barrier options. Journal of Computational and Applied Mathematics, 204(1):144–158, July 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002469.

Wang:2006:BQS

[WL06a] Ren-Hong Wang and Chong-Jun Li. Bivariate quartic spline spaces and quasiinterpolation operators. Journal of Computational and Applied Mathematics, 190(1– 2):325–338, June 1, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002384.

Wang:2006:P

[WL06b] Ren-Hong Wang and Xiaonan Luo. Preface. Journal of Computational and Applied Mathematics, 195 (1–2):1, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004668.

Wu:2007:TPS

[WL07] Jun Wu and Yicheng Liu. Two periodic solutions of neutral difference systems depending on two parameters. Journal of Computational and Applied Mathematics, 206(2):713–725, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005176.

Wang:2009:PLM

[WL09a] Yuan-Yuan Wang and Lin-Zhang Lu. Preconditioned Lanczos method for generalized Toeplitz eigenvalue problems. Journal of Computational and Applied Mathematics, 226(1):66–76, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002227.

Wen:2009:FDE

[WL09b] Meilin Wen and Huaishu Li. Fuzzy data envelopment analysis (DEA): Model and ranking method. Journal of Computational and Applied Mathematics, 223(2): 872–878, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000903.

Wang:2009:NPP

[WLF09] Jiaoyan Wang, Qishao Lu, and Zhaosheng Feng. A nonautonomous predator– prey system with stage structure and double time delays. Journal of Computational and Applied Mathematics, 230(1):283–299, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006249.

Wu:2006:SCMa

[WLK06] Xionghua Wu, Chen Li, and Wenbin Kong. A sinc-collocation method with boundary treatment for twodimensional elliptic boundary value problems. Journal of Computational and Applied Mathematics, 196(1): 58–69, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005170.

Wang:2006:RCB

[WLL06] Ruomei Wang, Xiaonan Luo, and Yi Li. Representation and conversion of Bézier surfaces in multivariate B-form. Journal of Computational and Applied Mathematics, 195(1–2):206–211, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004851.

Wang:2007:FCN

[WLL07] Zidong Wang, James Lam, and Xiaohui Liu. Filtering for a class of nonlinear discrete-time stochastic systems with state delays. Journal of Computational and Applied Mathematics, 201 (1):153–163, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600094X.

Wang:2009:RDM

[WLL09] Ying-Ming Wang, Ying Luo, and Liang Liang. Ranking decision making units by imposing a minimum weight restriction in the data envelopment analysis. Journal of Computational and Applied Mathematics, 223(1):

469–484, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800037X.

Widura:2008:OSA

[WLMS08] R. Widura, M. Lehn, K. Muralidhar, and R. Scherer. Operator splitting approach applied to oscillatory flow and heat transfer in a tube. Journal of Computational and Applied Mathematics, 211(2): 115–130, February 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006996.

Wang:2008:NSR

[WLS08] Wan-Sheng Wang, Shou-Fu Li, and Kai Su. Nonlinear stability of Runge–Kutta methods for neutral delay differential equations. Journal of Computational and Applied Mathematics, 214 (1):175–185, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001057.

Wang:2009:NSG

[WLS09] Wan-Sheng Wang, Shou-Fu Li, and Kai Su. Nonlinear stability of general linear methods for neutral delay differential equations. Journal of Computational and Applied Mathematics, 224(2): 592–601, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002768.

Wang:2005:CPN

[WLY05] Changyu Wang, Qian Liu, and Xinmin Yang. Convergence properties of nonmonotone spectral projected gradient methods. Journal of Computational and Applied Mathematics, 182 (1):51–66, October 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704006004.

Walton:2009:ISC

[WM09] D. J. Walton and D. S. Meek. G^{1} interpolation with a single Cornu spiral segment. Journal of Computational and Applied Mathematics, 223(1):86–96, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700670X.

Wagenknecht:2008:SPN

[WMG08] T. Wagenknecht, W. Michiels, and K. Green. Structured pseudospectra for nonlinear eigenvalue problems. Journal of Computational and Applied Mathematics, 212 (2):245–259, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007278.

Wang:2008:IOC

[WML08] Jizhong Wang, Fanwei Meng, and Sanyang Liu. Interval oscillation criteria for second order partial differential equations with delays. Journal of Computational and Applied Mathematics, 212 (2):397–405, March 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007849.

Wegmann:2005:RHP

[WMN05] R. Wegmann, A. H. M.

Murid, and M. M. S. Nasser. The Riemann–Hilbert problem and the generalized Neumann kernel. Journal of Computational and Applied Mathematics, 182(2): 388–415, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006375.

Wegmann:2008:RHP

[WN08] Rudolf Wegmann and Mohamed M. S. Nasser. The Riemann–Hilbert problem and the generalized Neumann kernel on multiply connected regions. Journal of Computational and Applied Mathematics, 214(1):36–57, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000805.

Wang:2009:ASD

[WNZ09] Tingchun Wang, Tao Nie, and Luming Zhang. Analysis of a symplectic difference scheme for a coupled nonlinear Schrödinger system. Journal of Computational and Applied Mathematics, 231(2): 745–759, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709002830.

Wollman:2008:DPM

[WO08] Stephen Wollman and Ercument Ozizmir. A deterministic particle method for the Vlasov–Fokker–Planck equation in one dimension. Journal of Computational and Applied Mathematics, 213 (2):316–365, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000362.

Wodarz:2005:MMI

[Wod05] Dominik Wodarz. Mathematical models of immune effector responses to viral infections: Virus control versus the development of pathology. Journal of Computational and Applied Mathematics, 184(1):301–319, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000828.

Wojdylo:2006:CAL

[Woj06]  John Wojdylo.  On the

coefficients that arise from Laplace’s method. Journal of Computational and Applied Mathematics, 196(1): 241–266, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005613.

Wu:2005:TTB

[WOK05] Y. Wu, L. Ozdamar, and A. Kumar. TRIOPT: a triangulation-based partitioning algorithm for global optimization. Journal of Computational and Applied Mathematics, 177(1):35–53, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003619.

Wong:2005:MFS

[Won05] Patricia J. Y. Wong. Multiple fixed-sign solutions for a system of difference equations with Sturm–Liouville conditions. Journal of Computational and Applied Mathematics, 183(1):108– 132, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000191.

Wong:2008:ESG

[Won08] Patricia J. Y. Wong. Eigenvalues of a system of generalized right focal problems with deviating arguments. Journal of Computational and Applied Mathematics, 218(2): 459–472, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700307X.

Wong:2009:AFE

[Won09a]  Jeff C.-F. Wong.  The

analysis of a finite element method for the three-species Lotka–Volterra competitiondiffusion with Dirichlet boundary conditions. Journal of Computational and Applied Mathematics, 223(1): 421–437, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000332.

Wong:2009:FIP

[Won09b] Manwah Lilian Wong. A formula for inserting point masses. Journal of Computational and Applied Mathematics, 233(3):852–855, December 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709001356.

Wang:2008:QAA

[WPC08] Fengyan Wang, Guoping Pang, and Lansun Chen. Qualitative analysis and applications of a kind of statedependent impulsive differential equations. Journal of Computational and Applied Mathematics, 216(1): 279–296, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002464.

Wright:2008:LAS

[Wri08] R. K. Wright. Local approximation on surfaces with discontinuities, given limited order Fourier coefficients. Journal of Computational and Applied Mathematics, 218(2): 532–542, September 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700516X.

Wouwer:2005:SIM

[WS05] Alain Vande Wouwer and W. E. Schiesser. Special issue on the method of lines: Dedicated to Keith Miller. Journal of Computational and Applied Mathematics, 183(2): 241–244, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000622.

Wang:2008:LDS

[WS08] Xiaoqun Wang and Ian H. Sloan. Low discrepancy sequences in high dimensions: How well are their projections distributed? Journal of Computational and Applied Mathematics, 213 (2):366–386, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000374.

Wang:2009:PAI

[WS09a] Li Wang and Yongzhong Song. Preconditioned AOR iterative methods for Mmatrices. Journal of Computational and Applied Mathematics, 226(1):114–124, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002379.

Weng:2009:PPS

[WS09b] Aizhi Weng and Jitao Sun. Positive periodic solutions of first-order functional differential equations with parameter. Journal of Computational and Applied Mathematics, 229(1):327–332, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005797.

Wang:2009:PSL

[WSM^{+}09] Lijuan Wang, Jianying Shao, Hua Meng, Bing Xiao, and Fei Long. Periodic solutions for Liénard type p-Laplacian equation with two deviating arguments. Journal of Computational and Applied Mathematics, 224(2): 751–758, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003117.

Weiner:2009:SET

[WSPJ09] Rüdiger Weiner, Bernhard A.

Schmitt, Helmut Podhaisky, and Stefan Jebens. Superconvergent explicit twostep peer methods. Journal of Computational and Applied Mathematics, 223(2): 753–764, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000721.

Wouwer:2005:MIU

[WSST05] A. Vande Wouwer, P. Saucez, W. E. Schiesser, and S. Thompson. A MATLAB implementation of upwind finite differences and adaptive grids in the method of lines. Journal of Computational and Applied Mathematics, 183(2): 245–258, November 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000634.

Wang:2009:IBV

[WSZ09] Guotao Wang, Guangxing Song, and Lihong Zhang. Integral boundary value problems for first order integrodifferential equations with deviating arguments. Journal of Computational and Applied Mathematics, 225(2): 602–611, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004329.

Wan:2007:EMF

[WT07] Decheng Wan and Stefan Turek. An efficient multigrid-FEM method for the simulation of solid-liquid two phase flows. Journal of Computational and Applied Mathematics, 203(2):561–580, June 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002305.

Wang:2008:CHP

[WT08] Guo-Jin Wang and Chiew-Lan Tai. On the convergence of hybrid polynomial approximation to higher derivatives of rational curves. Journal of Computational and Applied Mathematics, 214 (1):163–174, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001045.

Wang:2005:OPE

[WTL05] Peiguang Wang, K. L. Teo, and Yanqun Liu. Oscillation properties for even order neutral equations with distributed deviating arguments. Journal of Computational and Applied Mathematics, 182(2):290–303, October 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006168.

Wang:2009:NPA

[WtL09] Hongjuan Wang and Yao tang Li. A new preconditioned AOR iterative method for Lmatrices. Journal of Computational and Applied Mathematics, 229(1):47–53, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005177.

Wu:2008:NMC

[WTZ08] C. Z. Wu, K. L. Teo, and Yi Zhao. Numerical method for a class of optimal control problems subject to nonsmooth functional constraints. Journal of Computational and Applied Mathematics, 217(2):311–325, August 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707000921.

Wu:2007:MHB

[Wu07] Gang Wu. A modified harmonic block Arnoldi algorithm with adaptive shifts for large interior eigenproblems. Journal of Computational and Applied Mathematics, 205(1):343–363, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003232.

Wu:2008:EJC

[Wu08]  Gang Wu.  Eigenvalues

and Jordan canonical form of a successively rank-one updated complex matrix with applications to Google’s PageRank problem. Journal of Computational and Applied Mathematics, 216 (2):364–370, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002579.

Wunsche:2005:GZD

[Wün05] Alfred Wünsche. Generalized Zernike or disc polynomials. Journal of Computational and Applied Mathematics, 174(1):135–163, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001852.

Wei:2007:PPS

[WW07] Fengying Wei and Ke Wang.

Positive periodic solutions of an n-species ecological system with infinite delay. Journal of Computational and Applied Mathematics, 208(2): 362–372, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005905.

Wang:2009:MRA

[WW09] Shuming Wang and Junzo Watada. Modelling redundancy allocation for a fuzzy random parallel-series system. Journal of Computational and Applied Mathematics, 232(2):539–557, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003860.

Wang:2005:FOC

[WWC05] Peiguang Wang, Yonghong Wu, and Lou Caccetta. Forced oscillation of a class of neutral hyperbolic differential equations. Journal of Computational and Applied Mathematics, 177(2):301–308, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004273.

Wang:2007:OCB

[WWC07] Peiguang Wang, Yonghong Wu, and Lou Caccetta. Oscillation criteria for boundary value problems of highorder partial functional differential equations. Journal of Computational and Applied Mathematics, 206(1): 567–577, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600505X.

Wu:2009:ME

[WWD09] Ai-Guo Wu, Hao-Qian Wang, and Guang-Ren Duan. On matrix equations X −AXF = C and X − A\overline{X}F = C. Journal of Computational and Applied Mathematics, 230(2): 690–698, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000089.

Windcliff:2007:HCA

[WWFV07] H. Windcliff, J. Wang, P. A.

Forsyth, and K. R. Vetzal. Hedging with a correlated asset: Solution of a nonlinear pricing PDE. Journal of Computational and Applied Mathematics, 200 (1):86–115, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007983.

Wen:2007:CTB

[WWM07] Rui-Ping Wen, Chuan-Long Wang, and Guo-Yan Meng. Convergence theorems for block splitting iterative methods for linear systems. Journal of Computational and Applied Mathematics, 202 (2):540–547, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001609.

Wu:2008:RTD

[WWP^{+}08] Qing-E Wu, Tuo Wang, Xue-Min Pang, Yong-Xuan Huang, and Ji-Sheng Li. Research on two different mathematical theories on control. Journal of Computational and Applied Mathematics, 222(2):675–689, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006486.

Wang:2009:OPM

[WWP09] Tsung-Yin Wang, Kuo-Hsiung Wang, and Wen Lea Pearn. Optimization of the T policy M/G/1 queue with server breakdowns and general startup times. Journal of Computational and Applied Mathematics, 228 (1):270–278, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004743.

Wang:2008:CIS

[WWW08] Guixia Wang, Zhong Wang, and Hongyou Wu. Computing the indices of Sturm– Liouville eigenvalues for coupled boundary conditions (the EIGENIND–SLP codes). Journal of Computational and Applied Mathematics, 220(1–2):490–507, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004827.

Wang:2009:MPI

[WWWF09] Ying Wang, Weiming Wan, Rui-Sheng Wang, and Enmin Feng. Model, properties and imputation method of missing SNP genotype data utilizing mutual information. Journal of Computational and Applied Mathematics, 229(1):168–174, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005396.

Wu:2009:IMM

[WWZY09] Yu-Jiang Wu, Yang Wang, Min-Li Zeng, and Ai-Li Yang. Implementation of a modified Marder–Weitzner method for solving nonlinear eigenvalue problems. Journal of Computational and Applied Mathematics, 226(1):166–176, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002410.

Wang:2007:MPD

[WY07] Ying-Ming Wang and Jian-Bo Yang. Measuring the performances of decision-making units using interval efficiencies. Journal of Computational and Applied Mathematics, 198(1):253–267, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007892.

Wang:2008:SMP

[WYL08] NingShen Wang, RongHua Yi, and Dian Liu. A solution method to the problem proposed by Wang in voting systems. Journal of Computational and Applied Mathematics, 221(1): 106–113, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005286.

Wang:2009:NFF

[WYL09a] Chengjun Wang, Yongjian Yang, and Jing Li. A new filled function method for unconstrained global optimization. Journal of Computational and Applied Mathematics, 225(1):68–79, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003208.

Wen:2009:SAS

[WYL09b] Liping Wen, Yuexin Yu, and Shoufu Li. Stability and asymptotic stability of θmethods for nonlinear stiff Volterra functional differential equations in Banach spaces. Journal of Computational and Applied Mathematics, 230(2):351–359, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006316.

Wang:2009:RAC

[WYO09] Wentao Wang, Guangxue Yue, and Chunxia Ou. RE-TRACTED: Asymptotic constancy for a differential equation with multiple statedependent delays. Journal of Computational and Applied Mathematics, 233(2): 356–360, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004270.

Wang:2009:BVP

[WYS09] Weibing Wang, Xuxin Yang, and Jianhua Shen. Boundary value problems involving upper and lower solutions in reverse order. Journal of Computational and Applied Mathematics, 230(1):1–7, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005803.

Wang:2008:EPD

[WYW08] Ning-Shen Wang, Rong-Hua Yi, and Wei Wang. Evaluating the performances of decision-making units based on interval efficiencies. Journal of Computational and Applied Mathematics, 216 (2):328–343, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002531.

Wei:2006:BAP

[WZ06] Junjie Wei and Xingfu Zou. Bifurcation analysis of a population model and the resulting SIS epidemic model with delay. Journal of Computational and Applied Mathematics, 197(1):169– 187, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006515.

Wang:2007:EAL

[WZ07] Yong-Jun Wang and Jiang-She Zhang. An efficient algorithm for large scale global optimization of continuous functions. Journal of Computational and Applied Mathematics, 206(2): 1015–1026, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005759.

Wang:2008:MIT

[WZ08] Peiguang Wang and Jing Zhang. Monotone iterative technique for initial-value problems of nonlinear singular discrete systems. Journal of Computational and Applied Mathematics, 221(1): 158–164, November 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005328.

Wang:2009:HOC

[WZ09a] Yuan-Ming Wang and Hong-Bo Zhang. Higher-order compact finite difference method for systems of reaction– diffusion equations. Journal of Computational and Applied Mathematics, 233(2): 502–518, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004592.

Wei:2009:PSC

[WZ09b] Musheng Wei and Jinbing Zhang. On perturbations of some constrained subspaces. Journal of Computational and Applied Mathematics, 226(1):155–165, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002409.

Wang:2009:NBM

[WZD09] Xiaoying Wang, Shugong Zhang, and Tian Dong. Newton basis for multivariate Birkhoff interpolation. Journal of Computational and Applied Mathematics, 228 (1):466–479, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005128.

Wang:2006:MSD

[WZH06] X. P. Wang, S. F. Zhao, and X. F. Han. Micromagnetic simulations of demagnetization curves for single-phase nanocrystalline PrFeB magnet. Journal of Computational and Applied Mathematics, 190(1–2):142–150, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500227X.

Wylie:2006:FIP

[WZS06] Jonathan J. Wylie, Qiang Zhang, and Xiuxin Sun. Fingering instability in particle systems. Journal of Computational and Applied Mathematics, 190(1–2):408–423, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002438.

Xue:2006:FBP

[XC06] Qingshui Xue and Zhenfu Cao. Factoring based proxy signature schemes. Journal of Computational and Applied Mathematics, 195(1–2): 229–241, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004887.

Xiao:2008:BAD

[XC08] Min Xiao and Jinde Cao.

Bifurcation analysis on a discrete-time tabu learning model. Journal of Computational and Applied Mathematics, 220(1–2):725–738, October 15, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005110.

Xu:2009:RFN

[XC09] Aimin Xu and Zhongdi Cen.

A remainder formula of numerical differentiation for the generalized Lagrange interpolation. Journal of Computational and Applied Mathematics, 230(2):418–423, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006389.

Xiong:2008:CFI

[XCZ08] Zhiguang Xiong, Yanping Chen, and Yan Zhang. Convergence of FEM with interpolated coefficients for semilinear hyperbolic equation. Journal of Computational and Applied Mathematics, 214(1):313–317, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001070.

Xiang:2005:PBK

[XDW05] Hua Xiang, Huaian Diao, and Yimin Wei. On perturbation bounds of Kronecker product linear systems and their level-2 condition numbers. Journal of Computational and Applied Mathematics, 183(1): 210–231, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000269.

Xiong:2007:DST

[XF07] Xiang-Tuan Xiong and Chu-Li Fu. Determining surface temperature and heat flux by a wavelet dual least squares method. Journal of Computational and Applied Mathematics, 201(1):198–207, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000999.

Xu:2006:AAP

[XFgS06] Zhiqiang Xu, Renzhong Feng, and Jia guang Sun. Analytic and algebraic properties of canal surfaces. Journal of Computational and Applied Mathematics, 195(1–2): 220–228, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004875.

Xu:2009:SBA

[XGM09] Rui Xu, Qintao Gan, and Zhien Ma. Stability and bifurcation analysis on a ratio-dependent predator– prey model with time delay. Journal of Computational and Applied Mathematics, 230(1):187–203, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005980.

Xiao:2009:CFA

[XGZ09] Zhi Xiao, Ke Gong, and Yan Zou. A combined forecasting approach based on fuzzy soft sets. Journal of Computational and Applied Mathematics, 228(1):326–333, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005001.

Xia:2008:ASN

[XH08] Fu-Quan Xia and Nan-Jing Huang. Algorithm for solving a new class of general mixed variational inequalities in Banach spaces. Journal of Computational and Applied Mathematics, 220(1–2): 632–642, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005055.

Xie:2008:NSO

[XHK^{+}08] Shu-Sen Xie, Sunyeong Heo, Seokchan Kim, Gyungsoo Woo, and Sucheol Yi. Numerical solution of onedimensional Burgers’ equation using reproducing kernel function. Journal of Computational and Applied Mathematics, 214(2):417–434, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001379.

Xu:2007:VIM

[XHW07] Lan Xu, Ji-Huan He, and Abdul-Majid Wazwaz. Variational iteration method — reality, potential, and challenges. Journal of Computational and Applied Mathematics, 207(1):1–2, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706004547.

Xiang:2005:QBT

[Xia05] Shuhuang Xiang. On quadrature of Bessel transformations. Journal of Computational and Applied Mathematics, 177(1):231–239, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004339.

Xiang:2007:EQH

[Xia07a] Shuhuang Xiang. Efficient quadrature for highly oscillatory integrals involving critical points. Journal of Computational and Applied Mathematics, 206(2): 688–698, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005152.

Xiang:2007:FLM

[Xia07b] Shuhuang Xiang. On the Filon and Levin methods for highly oscillatory integral ∫_{a}^{b} f(x)e^{iωg(x)}dx. Journal of Computational and Applied Mathematics, 208(2): 434–439, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005929.

Xie:2008:BQF

[Xie08] Congcong Xie. Best quadrature formula on Sobolev class with Chebyshev weight. Journal of Computational and Applied Mathematics, 214 (2):333–344, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700129X.

Xiong:2007:SCG

[Xio07]  Zhiguang Xiong.  Super-

convergence of the continuous Galerkin finite element method for delay-differential equation with several terms. Journal of Computational and Applied Mathematics, 198(1):160–166, January 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007314.

Xenophontos:2006:VMF

[XKF06] Christos Xenophontos, Jason Kurtz, and Scott Fulton. A pversion MITC finite element method for Reissner–Mindlin plates with curved boundaries. Journal of Computational and Applied Mathematics, 192(2):374–395, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705003572.

Xu:2007:PTO

[XL07] Zhiting Xu and Xiuxiang Liu. Philos-type oscillation criteria for Emden–Fowler neutral delay differential equations. Journal of Computational and Applied Mathematics, 206(2):1116–1126, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005826.

Xu:2005:DDE

[XLHZ05] Shengyuan Xu, James Lam, Daniel W. C. Ho, and Yun Zou. Delay-dependent exponential stability for a class of neural networks with time delays. Journal of Computational and Applied Mathematics, 183(1):16–28, November 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704006442.

Li:2008:SFEb

[xLLhZ08] Ming xia Li, Qun Lin, and Shu hua Zhang. Superconvergence of finite element method for the Signorini problem. Journal of Computational and Applied Mathematics, 222(2): 284–292, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707005973.

Ma:2009:ULS

[xMzY09] De xiang Ma and Xiao zhong Yang. Upper and lower solution method for fourthorder four-point boundary value problems. Journal of Computational and Applied Mathematics, 223(2): 543–551, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707005225.

Xue:2009:PFF

[XSP09] Wenjuan Xue, Chungen Shen, and Dingguo Pu. A penalty-function-free line search SQP method for nonlinear programming. Journal of Computational and Applied Mathematics, 228 (1):313–325, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004998.

Xiao:2009:GCB

[XSW09] Yunhai Xiao, Huijuan Sun, and Zhiguo Wang. A globally convergent BFGS method with nonmonotone line search for non-convex minimization. Journal of Computational and Applied Mathematics, 230(1):95–106, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800589X.

Xiao:2006:CTF

[XTZ06] F. Xiao, X. H. Tang, and X. J. Zhang. Comparison of Taylor finite difference and window finite difference and their application in FDTD. Journal of Computational and Applied Mathematics, 193(2): 516–534, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004292.

Xu:2005:GTO

[Xu05a] Yuan Xu. Generalized translation operator and approximation in several variables. Journal of Computational and Applied Mathematics, 178(1–2):489–512, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003954.

Xu:2005:PIO

[Xu05b] Yuan Xu. Problem 6. Intertwining operators. Journal of Computational and Applied Mathematics, 178(1– 2):535–536, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004066.

Xu:2007:HPE

[Xu07a] Lan Xu. He’s parameterexpanding methods for strongly nonlinear oscillators. Journal of Computational and Applied Mathematics, 207(1): 148–154, October 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004705.

Xu:2007:NTA

[Xu07b]  Xiubin Xu.  Nonlinear

trigonometric approximation and the Dirac delta function. Journal of Computational and Applied Mathematics, 209(2):234–245, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706006881.

Xu:2008:ONS

[Xu08]  Zhiting Xu.  Oscillation

and nonoscillation of second order differential equations with delay depending on the unknown function. Journal of Computational and Applied Mathematics, 214(2):371–380, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707001331.

Xie:2005:PSS

[XW05] Guangming Xie and Long Wang. Periodical stabilization of switched linear systems. Journal of Computational and Applied Mathematics, 181(1):176– 187, September 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005606.

Xu:2006:PSN

[XW06] Rui Xu and Zhiqiang Wang.

Periodic solutions of a nonautonomous predator–prey system with stage structure and time delays. Journal of Computational and Applied Mathematics, 196(1): 70–86, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005182.

Xiang:2007:SMC

[XW07a] Hua Xiang and Yimin Wei.

Structured mixed and componentwise condition numbers of some structured matrices. Journal of Computational and Applied Mathematics, 202(2):217–229, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001233.

Xu:2007:OSO

[XW07b] Zhiting Xu and Peixuan Weng. Oscillation of second order neutral equations with distributed deviating argument. Journal of Computational and Applied Mathematics, 202(2):460–477, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706001531.

Xiang:2008:LIM

[XW08] Shuhuang Xiang and Haiyong Wang. On the Levin iterative method for oscillatory integrals. Journal of Computational and Applied Mathematics, 217(1):38–45, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003135.

Xu:2009:ART

[XW09] Hui-Xia Xu and Guo-Jin Wang. Approximating rational triangular Bézier surfaces by polynomial triangular Bézier surfaces. Journal of Computational and Applied Mathematics, 228 (1):287–295, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004858.

Xu:2009:SLV

[XWT09] Yong Xu, Fuke Wu, and Yimin Tan. Stochastic Lotka– Volterra system with infinite delay. Journal of Computational and Applied Mathematics, 232(2):472–480, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270900380X.

Xie:2008:NMC

[XWZH08] Feng Xie, Jian Wang, Weijiang Zhang, and Ming He. A novel method for a class of parameterized singularly perturbed boundary value problems. Journal of Computational and Applied Mathematics, 213(1):258–267, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000337.

Xu:2009:OCT

[XX09] Youjun Xu and Zhiting Xu.

Oscillation criteria for twodimensional dynamic systems on time scales. Journal of Computational and Applied Mathematics, 225 (1):9–19, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003154.

Xu:2008:PRC

[XXL08] Huiling Xu, Shengyuan Xu, and James Lam. Positive real control for 2-D discrete delayed systems via output feedback controllers. Journal of Computational and Applied Mathematics, 216 (1):87–97, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002208.

Xue:2005:HFH

[XY05] Yonggang Xue and Thomas P.-Y. Yu. Honeycomb and k-fold Hermite subdivision schemes. Journal of Computational and Applied Mathematics, 177(2):401–425, May 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704004637.

Xiao:2008:MSL

[XZ08] Yunhai Xiao and Hongchuan Zhang. Modified subspace limited memory BFGS algorithm for large-scale bound constrained optimization. Journal of Computational and Applied Mathematics, 222(2):429–439, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700619X.

Xie:2008:TMU

[XZL08] Dongxiu Xie, Zhongzhi Zhang, and Zhongyun Liu. Theory and method for updating least-squares finite element model of symmetric generalized centro-symmetric matrices. Journal of Computational and Applied Mathematics, 216(2):484–497, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002889.

Xiong:2009:NDD

[XZL09] Lianglin Xiong, Shouming Zhong, and Duyu Li. Novel delay-dependent asymptotical stability of neutral systems with nonlinear perturbations. Journal of Computational and Applied Mathematics, 232(2):505–513, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003835.

Xiao:2007:AMM

[XZS07] Ying-Xiong Xiao, Ping Zhang, and Shi Shu. An algebraic multigrid method with interpolation reproducing rigid body modes for semi-definite problems in two-dimensional linear elasticity. Journal of Computational and Applied Mathematics, 200(2): 637–652, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706000410.

Xiao:2009:SNM

[XZZ09] Xiantao Xiao, Liwei Zhang, and Jianzhong Zhang. A smoothing Newton method for a type of inverse semidefinite quadratic programming problem. Journal of Computational and Applied Mathematics, 223(1): 485–498, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000381.

Yakhno:2007:AEH

[YA07] V. G. Yakhno and H. K. Akmaz. Anisotropic elastodynamics in a half space: an analytic method for polynomial data. Journal of Computational and Applied Mathematics, 204(2):268–281, July 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003566.

Yilmaz:2008:AVL

[YA08] Yildiz E. Yilmaz and Aysen D. Akkaya. Analysis of variance and linear contrasts in experimental design with generalized secant hyperbolic distribution. Journal of Computational and Applied Mathematics, 216(2):545–553, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002944.

Yamaura:2008:CAF

[Yam08] Yoshihiko Yamaura. A convergence of the approximated free boundary of regularized functionals via Γconvergence. Journal of Computational and Applied Mathematics, 218(2):473– 481, September 1, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003883.

Yang:2005:QFG

[Yan05] Shijun Yang. On a quadrature formula of gori and micchelli. Journal of Computational and Applied Mathematics, 176(1):35–43, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003188.

Yan:2006:HBS

[Yan06a] Xiang-Ping Yan. Hopf bifurcation and stability for a delayed tri-neuron network model. Journal of Computational and Applied Mathematics, 196(2):579– 595, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705006229.

Yang:2006:SPP

[Yan06b] Heping Yang. On a singular perturbation problem with two second-order turning points. Journal of Computational and Applied Mathematics, 190(1–2):287–303, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002360.

Yang:2006:SBV

[Yan06c] Tong Yang. Singular behavior of vacuum states for compressible fluids. Journal of Computational and Applied Mathematics, 190(1– 2):211–231, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002323.

Yang:2006:EPE

[Yan06d] Zuodong Yang. Existence of positive entire solutions for singular and non-singular quasi-linear elliptic equation. Journal of Computational and Applied Mathematics, 197(2):355–364, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006849.

Yao:2009:IMC

[Yao09] Qingliu Yao. An iterative method for a class of quasilinear boundary value problems. Journal of Computational and Applied Mathematics, 230(1):306–313, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006274.

Yaslan:2007:EPS

[Yas07] Ismail Yaslan. Existence of positive solutions for nonlinear three-point problems on time scales. Journal of Computational and Applied Mathematics, 206(2): 888–897, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005383.

Yeung:2005:TFM

[YB05] Man-Chung Yeung and Daniel Boley. Transpose-free multiple Lanczos and its application in Padé approximation. Journal of Computational and Applied Mathematics, 177(1):101–127, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704004091.

Yang:2005:EPS

[YC05] Zhihui Yang and Jinde Cao. Existence of periodic solutions in neutral state-dependent delays equations and models. Journal of Computational and Applied Mathematics, 174(1): 179–199, February 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001888.

Yan:2006:SBA

[YC06] Xiang-Ping Yan and Yan-Dong Chu. Stability and bifurcation analysis for a delayed Lotka–Volterra predator– prey system. Journal of Computational and Applied Mathematics, 196(1):198– 210, November 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005443.

Yulan:2009:NAS

[YCJ09] Wang Yulan, Temuer Chaolu, and Pang Jing. New algorithm for second-order boundary value problems of integro-differential equation. Journal of Computational and Applied Mathematics, 229(1):1–6, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005190.

Chen:2008:MSA

[yCyZ08] Xiao yun Chen and Yan yan Zhan. Multi-scale anomaly detection algorithm based on infrequent pattern of time series. Journal of Computational and Applied Mathematics, 214(1):227–237, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/

science/article/pii/S0377042707001100.

See erratum [yCyZ09].

Chen:2009:EMS

[yCyZ09] Xiao yun Chen and Yan yan Zhan. Erratum to: “Multiscale anomaly detection algorithm based on infrequent pattern of time series” [J. Comput. Appl. Math. 214

(1) (2008) 227–237]. Journal of Computational and Applied Mathematics, 231(2): 1004, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002490.

See [yCyZ08].

Yuan:2008:NPS

[YD08] Yongxin Yuan and Hua Dai.

The nearness problems for symmetric matrix with a submatrix constraint. Journal of Computational and Applied Mathematics, 213(1): 224–231, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000295.

Yuan:2009:GIE

[YD09a] Yong-Xin Yuan and Hua Dai. A generalized inverse eigenvalue problem in structural dynamic model updating. Journal of Computational and Applied Mathematics, 226(1):42–49, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002197.

Yuan:2009:DUD

[YD09b] Yongxin Yuan and Hua Dai. The direct updating of damping and gyroscopic matrices. Journal of Computational and Applied Mathematics, 231(1):255–261, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000594.

Yuan:2006:ASS

[YG06] Chenggui Yuan and William Glover. Approximate solutions of stochastic differential delay equations with Markovian switching. Journal of Computational and Applied Mathematics, 194(2): 207–226, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004498.

Yamamoto:2007:EEF

[YG07] Nobito Yamamoto and Kenta Genma. On error estimation of finite element approximations to the elliptic equations in nonconvex polygonal domains. Journal of Computational and Applied Mathematics, 199(2):286–296, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007636.

Yang:2009:MMP

[YG09a] Aijun Yang and Weigao Ge.

The monotone method for periodic differential equations with the non well-ordered upper and lower solutions. Journal of Computational and Applied Mathematics, 232(2): 632–637, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003951.

Yi:2009:SVF

[YG09b] Lijun Yi and Benqi Guo. Superconvergence of the h–p version of the finite element method in one dimension. Journal of Computational and Applied Mathematics, 233(2):150–164, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900404X.

Yang:2005:DSFa

[YH05] ZhengHong Yang and YongJian Hu. Displacement structures and fast inversion formulas for q-adic Vandermonde-like matrices. Journal of Computational and Applied Mathematics, 176(1):1–14, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003164.

Yoshida:2007:GSD

[YH07] Naoki Yoshida and Tadayuki Hara. Global stability of a delayed SIR epidemic model with density dependent birth and death rates. Journal of Computational and Applied Mathematics, 201(2): 339–347, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000860.

Yin:2009:PGM

[YH09a] Jun-Feng Yin and Ken Hayami. Preconditioned GM-RES methods with incomplete Givens orthogonalization method for large sparse least-squares problems. Journal of Computational and Applied Mathematics, 226 (1):177–186, April 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002422.

Yuan:2009:BEL

[YH09b] Quan Yuan and Zhiqing He. Bounds to eigenvalues of the Laplacian on L-shaped domain by variational methods. Journal of Computational and Applied Mathematics, 233(4):1083–1090, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709006189.

Yuan:2005:SBA

[YHH05] Zhaohui Yuan, Dewen Hu, and Lihong Huang. Stability and bifurcation analysis on a discrete-time neural network. Journal of Computational and Applied Mathematics, 177(1):89–100, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270400408X.

Yi:2009:AIR

[Yi09]  Dokkyun Yi.  An algo-

rithm for image removals and decompositions without inverse matrices. Journal of Computational and Applied Mathematics, 225(2): 428–439, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003956.

Yang:2007:LDC

[YJGL07] Lixing Yang, Xiaoyu Ji, Ziyou Gao, and Keping Li. Logistics distribution centers location problem and algorithm under fuzzy environment. Journal of Computational and Applied Mathematics, 208(2): 303–315, November 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600584X.

Yin:2009:NMP

[YJJ09] G. Yin, Hanqing Jin, and Zhuo Jin. Numerical methods for portfolio selection with bounded constraints. Journal of Computational and Applied Mathematics, 233(2): 564–581, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709005330.

Yaz:2006:LAD

[YJY06] Edwin Engin Yaz, Chung Seop Jeong, and Yvonne Ilke Yaz. An LMI approach to discretetime observer design with stochastic resilience. Journal of Computational and Applied Mathematics, 188 (2):246–255, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002530.

Yuksek:2009:SCD

[YKS09] Kemal Yüksek, Yeliz Koca, and Hasan Sadikoğlu. Solution of counter diffusion problem with position dependent diffusion coefficient by using variational methods. Journal of Computational and Applied Mathematics, 232(2):285–294, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709003550.

Ye:2006:CFD

[YL06] Xintao Ye and Chong Li.

Convergence of the family of the deformed Euler– Halley iterations under the Hölder condition of the second derivative. Journal of Computational and Applied Mathematics, 194(2): 294–308, October 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004759.

Lan:2008:MMS

[yL08] Heng you Lan. Monotone method for a system of nonlinear mixed type implicit impulsive integro-differential equations in Banach spaces. Journal of Computational and Applied Mathematics, 222(2):531–543, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006309.

Yang:2009:RGI

[YL09a] Hu Yang and Deqiang Liu.

The representation of generalized inverse A_{T,S}^{(2,3)} and its applications. Journal of Computational and Applied Mathematics, 224(1): 204–209, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800191X.

Yao:2009:SSB

[YL09b] Huanmin Yao and Yingzhen Lin. Solving singular boundary-value problems of higher even-order. Journal of Computational and Applied Mathematics, 223(2): 703–713, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800068X.

Yang:2009:EEQ

[YLC09] Min Yang, Jiangguo Liu, and Chuanjun Chen. Error estimation of a quadratic finite volume method on right quadrangular prism grids. Journal of Computational and Applied Mathematics, 229(1):274–282, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800575X.

Yang:2009:RKM

[YLN09] Z. W. Yang, M. Z. Liu, and Juan J. Nieto. Runge–Kutta methods for first-order periodic boundary value differential equations with piecewise constant arguments. Journal of Computational and Applied Mathematics, 233(4): 990–1004, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709005871.

Ye:2007:CVC

[YLS07] Xintao Ye, Chong Li, and Weiping Shen. Convergence of the variants of the Chebyshev–Halley iteration family under the Hölder condition of the first derivative. Journal of Computational and Applied Mathematics, 203(1):279–288, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706002111.

Yang:2006:HOS

[YLSX06] Yuyue Yang, Mingjun Li, Shi Shu, and Aiguo Xiao. High order schemes based on upwind schemes with modified coefficients. Journal of Computational and Applied Mathematics, 195(1–2): 242–251, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004899.

Yuan:2009:BTR

[YLW09a] Gonglin Yuan, Xiwen Lu, and Zengxin Wei. BFGS trustregion method for symmetric nonlinear equations. Journal of Computational and Applied Mathematics, 230 (1):44–58, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005852.

Yuan:2009:CGM

[YLW09b] Gonglin Yuan, Xiwen Lu, and Zengxin Wei. A conjugate gradient method with descent direction for unconstrained optimization. Journal of Computational and Applied Mathematics, 233(2): 519–530, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004646.

Yao:2007:ABS

[YM07] Jianli Yao and Fanwei Meng. Asymptotic behavior of solutions of certain higher order nonlinear difference equation. Journal of Computational and Applied Mathematics, 205(1):640–650, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003463.

Yan:2008:SRI

[YM08] Wen-Jing Yan and Yi-Chen Ma. Shape reconstruction of an inverse Stokes problem. Journal of Computational and Applied Mathematics, 216(2):554–562, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002956.

Yao:2008:CCG

[YN08] Yonghong Yao and Muhammad Aslam Noor. On convergence criteria of generalized proximal point algorithms. Journal of Computational and Applied Mathematics, 217(1):46–55, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003147.

Yan:2006:USS

[YOA06] Baoqiang Yan, Donal O’Regan, and Ravi P. Agarwal. Unbounded solutions for singular boundary value problems on the semi-infinite interval: Upper and lower solutions and multiplicity. Journal of Computational and Applied Mathematics, 197(2): 365–386, December 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006850.

Yamatani:2007:DIM

[YOO07] Katsu Yamatani, Takashi Ohe, and Kohzaburo Ohnaka. A direct identification method for current dipoles from electromagnetic fields. Journal of Computational and Applied Mathematics, 201 (1):164–174, April 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000951.

Yabe:2007:LSC

[YOY07] Hiroshi Yabe, Hideho Ogasawara, and Masayuki Yoshino. Local and superlinear convergence of quasi-Newton methods based on modified secant conditions. Journal of Computational and Applied Mathematics, 205(1): 617–632, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600344X.

Yu:2008:NNL

[YP08] Zhensheng Yu and Dingguo Pu. A new nonmonotone line search technique for unconstrained optimization. Journal of Computational and Applied Mathematics, 219(1): 134–144, September 15, 2008.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003718.

Peng:2007:IME

[yPESlZ07] Zhen yun Peng, Salah M. El-Sayed, and Xiang lin Zhang. Iterative methods for the extremal positive definite solution of the matrix equation X + A^{∗}X^{−α}A = Q. Journal of Computational and Applied Mathematics, 200(2): 520–527, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270600032X.

Peng:2005:CJM

[yPlH05] Zhen yun Peng and Xu li Han.

Constructing Jacobi matrices with prescribed ordered defective eigenpairs and a principal submatrix. Journal of Computational and Applied Mathematics, 175(2): 321–333, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704002365.

Peng:2007:RAR

[yPyHZ07] Xiang yang Peng, Xi yan Hu, and Lei Zhang. The reflexive and anti-reflexive solutions of the matrix equation A^{H}XB = C. Journal of Computational and Applied Mathematics, 200(2): 749–760, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000707.

Sun:2005:CTS

[yS05] Li ying Sun. A comparison theorem for the SOR iterative method. Journal of Computational and Applied Mathematics, 181(2): 336–341, September 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704005898.

See comment [NK10].

Yaguchi:2006:HNB

[YS06] Takaharu Yaguchi and Kokichi Sugihara. On the Hedstrom nonreflecting boundary condition for C^{2} solutions. Journal of Computational and Applied Mathematics, 197(1):150–155, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006497.

Yang:2007:PBV

[YS07] Xuxin Yang and Jianhua Shen. Periodic boundary value problems for secondorder impulsive integrodifferential equations. Journal of Computational and Applied Mathematics, 209(2): 176–186, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706006741.

Yin:2009:SOA

[YSY09] G. Yin, Q. S. Song, and H. Yang. Stochastic optimization algorithms for barrier dividend strategies. Journal of Computational and Applied Mathematics, 223(1): 240–262, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000101.

Yi:2008:NCB

[YSYX08] Xuejun Yi, Jianying Shao, Yuehua Yu, and Bing Xiao. New convergence behavior of high-order Hopfield neural networks with timevarying coefficients. Journal of Computational and Applied Mathematics, 219(1): 216–222, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003895.

Yamaguchi:2007:DPS

[YTM07] Masahiro Yamaguchi, Yasuhiro Takeuchi, and Wanbiao Ma. Dynamical properties of a stage structured three-species model with intra-guild predation. Journal of Computational and Applied Mathematics, 201 (2):327–338, April 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000859.

Yu:2006:SSP

[Yu06]  Zhensheng Yu.  Solving

semidefinite programming problems via alternating direction methods. Journal of Computational and Applied Mathematics, 193(2): 437–445, September 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004231.

Yuan:2008:MUM

[Yua08] Yongxin Yuan. A model updating method for undamped structural systems. Journal of Computational and Applied Mathematics, 219(1): 294–301, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004098.

Yuan:2009:IES

[Yua09] Yongxin Yuan. Inverse eigenproblem for R-symmetric matrices and their approximation. Journal of Computational and Applied Mathematics, 233(2):308–314, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709004233.

Yucel:2006:ASL

[Yüc06] Uğur Yücel. Approximations of Sturm–Liouville eigenvalues using differential quadrature (DQ) method. Journal of Computational and Applied Mathematics, 192(2): 310–319, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270500350X.

Yun:2005:CNM

[Yun05a] Jae Heon Yun. Convergence of nonstationary multisplitting methods using ILU factorizations. Journal of Computational and Applied Mathematics, 180(2): 245–263, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005205.

Yun:2005:NSS

[Yun05b] Jae Heon Yun. A note on SOR(ω) splitting of an Mmatrix. Journal of Computational and Applied Mathematics, 176(2):461–462, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003486.

Yun:2008:CSM

[Yun08a] Jae Heon Yun. Convergence of SSOR multisplitting method for an Hmatrix. Journal of Computational and Applied Mathematics, 217(1):252–258, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003548.

Yun:2008:PAM

[Yun08b] Jae Heon Yun. Preconditioned AOR method for Lmatrices. A note on: “Improvements of preconditioned AOR iterative method for Lmatrices” [J. Comput. Appl. Math. 206 (2007), no. 2, 656– 665; MR2333703] by Y. Li, C. Li and S. Wu. Journal of Computational and Applied Mathematics, 220(1– 2):13–16, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003846.

See [tLxLlW07].

Yin:2008:OMO

[YW08] Peng-Yeng Yin and Jing-Yu Wang. Optimal multipleobjective resource allocation using hybrid particle swarm optimization and adaptive resource bounds technique. Journal of Computational and Applied Mathematics, 216(1):73–86, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002191.

Yin:2009:OBS

[YW09a] Chuancun Yin and Chunwei Wang. Optimality of the barrier strategy in de Finetti’s dividend problem for spectrally negative Lévy processes: an alternative approach. Journal of Computational and Applied Mathematics, 233(2): 482–491, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004579.

Yu:2009:ABS

[YW09b] Zhi-Hua Yu and Qi-Ru Wang.

Asymptotic behavior of solutions of third-order nonlinear dynamic equations on time scales. Journal of Computational and Applied Mathematics, 225(2):531–540, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004251.

Yang:2009:ACN

[YWD09] Xiaoyuan Yang, Wei Wang, and Yuanyuan Duan. The approximation of a Crank– Nicolson scheme for the stochastic Navier–Stokes equations. Journal of Computational and Applied Mathematics, 225(1):31–43, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003178.

Yang:2005:DSFb

[YWH05a] ZhengHong Yang, LaiSheng Wang, and YongJian Hu. Displacement structures and fast inversion formulas for confluent polynomial Vandermondelike matrices. Journal of Computational and Applied Mathematics, 177(1): 1–15, May 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003590.

Yang:2005:DSFc

[YWH05b] ZhengHong Yang, LaiSheng Wang, and YongJian Hu. Displacement structures and fast inversion formulas for confluent polynomial Vandermondelike matrices. Journal of Computational and Applied Mathematics, 180(2): 229–243, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005096.

Yu:2007:CCH

[YXF07] Bo Yu, Qing Xu, and Guochen Feng. On the complexity of a combined homotopy interior method for convex programming. Journal of Computational and Applied Mathematics, 200 (1):32–46, March 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007946.

Yin:2008:AEE

[YXJG08] Xiaobo Yin, Hehu Xie, Shanghui Jia, and Shaoqin Gao. Asymptotic expansions and extrapolations of eigenvalues for the Stokes problem by mixed finite element methods. Journal of Computational and Applied Mathematics, 215(1):127–141, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001793.

Yun-Xin:2006:MSU

[YXYJ06] Zhang Yun-Xin and Tan Yong-Ji. Meshless schemes for unsteady Navier–Stokes equations in vorticity formulation using radial basis functions. Journal of Computational and Applied Mathematics, 192(2):328–338, August 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705003547.

Yang:2007:SCF

[YY07a] Min Yang and Yirang Yuan. A symmetric characteristic FVE method with second order accuracy for nonlinear convection diffusion problems. Journal of Computational and Applied Mathematics, 200(2):677–700, March 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706000665.

Yongjian:2007:NDF

[YY07b] Yang Yongjian and Liang Yumei. A new discrete filled function algorithm for discrete global optimization. Journal of Computational and Applied Mathematics, 202(2):280–291, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001294.

Yu:2007:GRS

[YY07c] Wenwu Yu and Lingling Yao. Global robust stability of neural networks with time varying delays. Journal of Computational and Applied Mathematics, 206(2):

679–687, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005140.

Yuan:2008:GDS

[YY08] Junli Yuan and Zuodong Yang. Global dynamics of an SEI model with acute and chronic stages. Journal of Computational and Applied Mathematics, 213 (2):465–476, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000453.

Yan:2009:BIA

[YYF09] Liang Yan, Fenglian Yang, and Chuli Fu. A Bayesian inference approach to identify a Robin coefficient in onedimensional parabolic problems. Journal of Computational and Applied Mathematics, 231(2):840–850, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042709003033.

Yan:2009:PPE

[YZ09a] Ningning Yan and Zhaojie Zhou. A priori and a posteriori error analysis of edge stabilization Galerkin method for the optimal control problem governed by convectiondominated diffusion equation. Journal of Computational and Applied Mathematics, 223(1):198–217, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000071.

Yang:2009:PSA

[YZ09b] Hu Yang and Zhimin Zhang. On a perturbed Sparre Andersen risk model with multilayer dividend strategy. Journal of Computational and Applied Mathematics, 232(2): 612–624, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003938.

Yang:2009:FOS

[YZ09c] Jun Yang and Yu Jing Zhang. Frequent oscillatory solutions of a nonlinear partial difference equation. Journal of Computational and Applied Mathematics, 224(2): 492–499, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708002562.

Yang:2009:FER

[YZ09d] Ying Yang and Aihui Zhou. A finite element recovery approach to Green’s function approximations with applications to electrostatic potential computation. Journal of Computational and Applied Mathematics, 225 (1):202–212, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003634.

Yu:2009:FCG

[YZ09e] Xiaohui Yu and Qiang Zhang.

The fuzzy core in games with fuzzy coalitions. Journal of Computational and Applied Mathematics, 230(1): 173–186, August 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005979.

Yang:2008:MSS

[YZL08] Jun Yang, Shouming Zhong, and Wenpin Luo. Mean square stability analysis of impulsive stochastic differential equations with delays. Journal of Computational and Applied Mathematics, 216(2):474–483, July 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707002877.

Zhang:2009:CBI

[yZXL09] Cheng yi Zhang, Chengxian Xu, and Shuanghua Luo.

Convergence of block iterative methods for linear systems with generalized Hmatrices. Journal of Computational and Applied Mathematics, 229(1):70–84, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005219.

Zhu:2009:SNC

[ZA09] Wenxing Zhu and M. M. Ali. Solving nonlinearly constrained global optimization problem via an auxiliary function method. Journal of Computational and Applied Mathematics, 230(2): 491–503, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006468.

Zadeh:2009:MOO

[Zad09] Kouroush Sadegh Zadeh. Multi-objective optimization in variably saturated fluid flow. Journal of Computational and Applied Mathematics, 223(2):801–819, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800085X.

Zarin:2009:CDF

[Zar09] Helena Zarin. Continuousdiscontinuous finite element method for convection–diffusion problems with characteristic layers. Journal of Computational and Applied Mathematics, 231(2):626– 636, September 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002623.

Zandieh:2009:RPD

[ZAVS09] M. Zandieh, M. Amiri, B. Vahdani, and R. Soltani. A robust parameter design for multi-response problems. Journal of Computational and Applied Mathematics, 230(2):463–476, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708006432.

Zheng:2005:MAE

[ZBB^{+}05] Yanan Zheng, Venkataramanan Balakrishnan, Greg Buzzard, Robert Geahlen, Marietta Harrison, and Ann Rundell. Modeling and analysis of early events in Tlymphocyte antigen-activated intracellular-signaling pathways. Journal of Computational and Applied Mathematics, 184(1):320–341, December 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270500083X.

Zheng:2008:CSC

[ZBL08] Quan Zheng, Rongxia Bai, and Zhongli Liu. The cubic semilocal convergence on two variants of Newton’s method. Journal of Computational and Applied Mathematics, 220(1–2):480–489, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004803.

Zeng:2006:CFD

[ZC06] Xiao-Ming Zeng and X. J. Chen. Computational formula of depth for Catmull– Clark subdivision surfaces. Journal of Computational and Applied Mathematics, 195(1–2):252–262, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004905.

Zhao:2008:EMP

[ZC08a] Yu-Lin Zhao and Hai-Bo Chen. Existence of multiple positive solutions for mpoint boundary value problems in Banach spaces. Journal of Computational and Applied Mathematics, 215 (1):79–90, May 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001756.

Zhu:2008:CSA

[ZC08b] Guoqing Zhu and Shaochun Chen. Convergence and superconvergence analysis of finite element methods on graded meshes for singularly and semisingularly perturbed reaction–diffusion problems. Journal of Computational and Applied Mathematics, 220(1–2):373–393, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700458X.

zuCastell:2007:P

[zCEF07] W. zu Castell, S. Ehrich, and F. Filbir. Preface. Journal of Computational and Applied Mathematics, 199 (1):1, February 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007338.

Zhang:2007:BSN

[ZCH07] Lijun Zhang, Li-Qun Chen, and Xuwen Huo. Bifurcations of smooth and nonsmooth traveling wave solutions in a generalized Degasperis– Procesi equation. Journal of Computational and Applied Mathematics, 205(1): 174–185, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706002731.

Zhang:2006:ULS

[ZCL06] Qin Zhang, Shihua Chen, and Jinhu Lü. Upper and lower solution method for fourthorder four-point boundary value problems. Journal of Computational and Applied Mathematics, 196(2): 387–393, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705005820.

Zhou:2009:NAP

[ZCL09] Yongfang Zhou, Minggen Cui, and Yingzhen Lin. Numerical algorithm for parabolic problems with non-classical conditions. Journal of Computational and Applied Mathematics, 230(2):770–780, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000259.

Zeng:2006:EPS

[ZCS06] Guangzhao Zeng, Lansun Chen, and Lihua Sun. Existence of periodic solution of order one of planar impulsive autonomous system. Journal of Computational and Applied Mathematics, 186(2): 466–481, February 15, 2006.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001275.

Zou:2008:LBC

[ZCZ08] Lan Zou, Xingwu Chen, and Weinian Zhang. Local bifurcations of critical periods for cubic Liénard equations with cubic damping. Journal of Computational and Applied Mathematics, 222(2): 404–410, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006085.

Zhao:2007:CFD

[ZDC07] Jichao Zhao, Matt Davison, and Robert M. Corless. Compact finite difference method for American option pricing. Journal of Computational and Applied Mathematics, 206(1):306–321, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706004511.

Zhu:2009:DDC

[ZF09] Wenxing Zhu and Hong Fan.

A discrete dynamic convexized method for nonlinear integer programming. Journal of Computational and Applied Mathematics, 223(1): 356–373, January 1, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000289.

Zhang:2008:SPS

[ZFG08] Xuemei Zhang, Meiqiang Feng, and Weigao Ge. Symmetric positive solutions for p-Laplacian fourth-order differential equations with integral boundary conditions. Journal of Computational and Applied Mathematics, 222(2):561–573, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006322.

Zlatev:2009:LSS

[ZFS09] Zahari Zlatev, István Faragó, and Peter L. Simon. Large scale scientific computations: Editorial introduction. Journal of Computational and Applied Mathematics, 226 (2):187–189, April 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003981.

Zhang:2009:ESP

[ZG09] Jie Zhang and Zhanji Gui. Existence and stability of periodic solutions of highorder Hopfield neural networks with impulses and delays. Journal of Computational and Applied Mathematics, 224(2):602–613, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800277X.

Zhang:2009:SSB

[ZGH09] Haomin Zhang, Siqing Gan, and Lin Hu. The split-step backward Euler method for linear stochastic delay differential equations. Journal of Computational and Applied Mathematics, 225(2): 558–568, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004275.

Zhang:2008:NCF

[ZGSF08] Hongjuan Zhang, Chonghui Guo, Zhenwei Shi, and Enmin Feng. A new constrained fixed-point algorithm for ordering independent components. Journal of Computational and Applied Mathematics, 220(1–2): 548–558, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004852.

Zaviyalov:2006:ESS

[ZGV06] V. V. Zaviyalov, V. Yu. Gusev, and A. S. Vershinskaya. Exact solutions of system of equations of radiation and energy transfer in 3D spherically nonsymmetrical geometry. Journal of Computational and Applied Mathematics, 189(1–2):635–642, May 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705002967.

Zhou:2008:WLS

[ZH08] Tianhe Zhou and Danfu Han. A weighted least squares method for scattered data fitting. Journal of Computational and Applied Mathematics, 217(1):56–63, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003159.

Zhu:2009:TLG

[ZH09] Liping Zhu and Yinnian He. Two-level Galerkin– Lagrange multipliers method for the stationary Navier– Stokes equations. Journal of Computational and Applied Mathematics, 230(2): 504–512, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270800647X.

Zhang:2005:MAD

[Zha05] Xinhua Zhang. A modification of the Adomian decomposition method for a class of nonlinear singular boundary value problems. Journal of Computational and Applied Mathematics, 180(2): 377–389, August 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704005394.

Zhang:2007:NIT

[Zha07] Shangyou Zhang. Numerical integration with Taylor truncations for the quadrilateral and hexahedral finite elements. Journal of Computational and Applied Mathematics, 205(1):325–342, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706003220.

Zhang:2008:ESN

[Zha08] Qimin Zhang. Exponential stability of numerical solutions to a stochastic agestructured population system with diffusion. Journal of Computational and Applied Mathematics, 220(1– 2):22–33, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004608.

Zhai:2009:PSS

[Zha09a] Chengbo Zhai. Positive solutions for semi-positone threepoint boundary value problems. Journal of Computational and Applied Mathematics, 228(1):279–286, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004755.

Zhang:2009:NLS

[Zha09b] Li Zhang. A new Liu–Storey type nonlinear conjugate gradient method for unconstrained optimization problems. Journal of Computational and Applied Mathematics, 225(1):146–157, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003592.

Zhang:2009:CIS

[ZHG^{+}09] Li-Tao Zhang, Ting-Zhu Huang, Tong-Xiang Gu, Xin-Lan Guo, and Jiang-Hua Yue. Convergent improvement of SSOR multisplitting method for an H-matrix. Journal of Computational and Applied Mathematics, 225(2): 393–397, March 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708003889.

Zhang:2008:GSS

[ZhL08] Guo-Feng Zhang and Qun hua Lu. On generalized symmetric SOR method for augmented systems. Journal of Computational and Applied Mathematics, 219(1): 51–58, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003627.

Zhou:2006:SCI

[Zho06] Fu-Zhao Zhou. The solvability conditions for the inverse eigenvalue problems of reflexive matrices. Journal of Computational and Applied Mathematics, 188(2): 180–189, April 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002153.

Zhou:2008:DRP

[Zho08] Ding-Xuan Zhou. Derivative reproducing properties for kernel methods in learning theory. Journal of Computational and Applied Mathematics, 220(1–2): 456–463, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004657.

Zhou:2009:SCT

[Zho09a]  Ying Zhou.  A smooth-

ing conic trust region filter method for the nonlinear complementarity problem. Journal of Computational and Applied Mathematics, 229(1):248–263, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005645.

Zhou:2009:QBP

[Zho09b] Zhengxin Zhou. On the qualitative behavior of periodic solutions of differential systems. Journal of Computational and Applied Mathematics, 232(2):600–611, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003926.

Zhu:2005:AST

[Zhu05a]  Detong Zhu.  An affine

scaling trust-region algorithm with interior backtracking technique for solving bound-constrained nonlinear systems. Journal of Computational and Applied Mathematics, 184(2): 343–361, December 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705000282.

Zhu:2005:IAS

[Zhu05b] Detong Zhu. An interior affine scaling projective algorithm for nonlinear equality and linear inequality constrained optimization. Journal of Computational and Applied Mathematics, 173(1): 115–148, January 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704001414.

Zhu:2005:ESQ

[Zhu05c] Zhibin Zhu. An efficient sequential quadratic programming algorithm for nonlinear programming. Journal of Computational and Applied Mathematics, 175(2): 447–464, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003152.

Zhu:2006:ASI

[Zhu06] Detong Zhu. Affine scaling inexact generalized Newton algorithm with interior backtracking technique for solving bound-constrained semismooth equations. Journal of Computational and Applied Mathematics, 187(2): 227–252, March 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001548.

Zhuang:2007:AEI

[Zhu07]  Yu Zhuang.  An alter-

nating explicit-implicit domain decomposition method for the parallel solution of parabolic equations. Journal of Computational and Applied Mathematics, 206(1): 549–566, September 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005048.

Zhu:2008:ASI

[Zhu08a] Detong Zhu. Affine scaling interior Levenberg–Marquardt method for bound-constrained semismooth equations under local error bound conditions. Journal of Computational and Applied Mathematics, 219(1):198–215, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707003871.

Zhu:2008:BFS

[Zhu08b] Jinghao Zhu. The bounds of feasible space on constrained nonconvex quadratic programming. Journal of Computational and Applied Mathematics, 213(1): 205–211, March 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270700026X.

Zhu:2008:SEC

[Zhu08c]  Zhibin Zhu.  A sequen-

tial equality constrained quadratic programming algorithm for inequality constrained optimization. Journal of Computational and Applied Mathematics, 212(1): 112–125, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007199.

Zhuang:2008:AOC

[Zhu08d] Rong-Kun Zhuang. Annual oscillation criteria for secondorder nonlinear elliptic differential equations. Journal of Computational and Applied Mathematics, 217 (1):268–276, July 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003561.

Zinkernagel:2005:IIA

[Zin05] Rolf M. Zinkernagel. Immunology and immunity against infection: General rules. Journal of Computational and Applied Mathematics, 184(1):4–9, December 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705000580.

Zhang:2012:CFD

[ZJ12] Li Zhang and Shuyuan Jian. Comments on: “A family of derivative-free conjugate gradient methods for largescale nonlinear systems of equations” [J. Comput. Appl. Math. 224 (2009) 11–19]. Journal of Computational and Applied Mathematics, 236(17):4344–4347, November 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042712001574.

See [CXH09].

Zhang:2008:SNT

[ZJW08] Xinzhen Zhang, Hefeng Jiang, and Yiju Wang. A smoothing Newton-type method for generalized nonlinear complementarity problem. Journal of Computational and Applied Mathematics, 212(1):75–85, February 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706007138.

Zschiedrich:2006:NFE

[ZKSS06] Lin Zschiedrich, Roland Klose, Achim Schädle, and Frank Schmidt. A new finite element realization of the perfectly matched layer method for Helmholtz scattering problems on polygonal domains in two dimensions. Journal of Computational and Applied Mathematics, 188(1):12–32, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001561.

Zhang:2005:HCL

[ZL05] Peng Zhang and Ru-Xun Liu.

Hyperbolic conservation laws with space-dependent fluxes: II. General study of numerical fluxes. Journal of Computational and Applied Mathematics, 176(1):105–129, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003231.

Zhou:2007:EUP

[ZL07] Qiyuan Zhou and Fei Long.

Existence and uniqueness of periodic solutions for a kind of Liénard equation with two deviating arguments. Journal of Computational and Applied Mathematics, 206(2): 1127–1136, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005838.

Zeng:2009:EVP

[ZL09a] J. Zeng and S. J. Li. An Ekeland’s variational principle for set-valued mappings with applications. Journal of Computational and Applied Mathematics, 230(2): 477–484, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708006444.

Zhou:2009:DLB

[ZL09b] Lujun Zhou and Yaqiong Li. A dynamic IS–LM business cycle model with two time delays in capital accumulation equation. Journal of Computational and Applied Mathematics, 228(1):182–187, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708004615.

Zhang:2007:NAL

[ZLA07] H. Zhang, F. Liu, and V. Anh. Numerical approximation of Lévy–Feller diffusion equation and its probability interpretation. Journal of Computational and Applied Mathematics, 206(2):1098– 1115, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005814.

Liu:2006:FMD

[zLCgS06] Xu zheng Liu, Xia Cui, and Jia guang Sun. FDM for multi-dimensional nonlinear coupled system of parabolic and hyperbolic equations.

Journal of Computational and Applied Mathematics, 186(2):432–449, February 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001044.

Zegarra:2009:WBF

[ZLdST09] Javier A. Montoya Zegarra, Neucimar J. Leite, and Ricardo da Silva Torres. Wavelet-based fingerprint image retrieval. Journal of Computational and Applied Mathematics, 227(2):294–307, May 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800112X.

Zhou:2009:WLS

[ZLDW09] Bin Zhou, Zhao-Yan Li, Guang-Ren Duan, and Yong Wang. Weighted least squares solutions to general coupled Sylvester matrix equations. Journal of Computational and Applied Mathematics, 224(2):759–776, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003129.

Zhao:2008:ESI

[ZLL08] Jing Zhao, Zhenbin Liu, and Lishan Liu. The existence of solutions of infinite boundary value problems for first-order impulsive differential systems in Banach spaces. Journal of Computational and Applied Mathematics, 222(2): 524–530, December 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006292.

Zhang:2009:SMS

[ZLL09] Xiangsong Zhang, Sanyang Liu, and Zhenhua Liu. A smoothing method for second order cone complementarity problem. Journal of Computational and Applied Mathematics, 228(1):83–91, June 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708004408.

Zhang:2008:LAB

[ZLLW08] Yunong Zhang, W. E. Leithead, D. J. Leith, and L. Walshe. Log-det approximation based on uniformly distributed seeds and its application to Gaussian process regression. Journal of Computational and Applied Mathematics, 220(1–2): 198–214, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707004360.

Zhang:2009:GSD

[ZLY09] Zhimin Zhang, Shuanming Li, and Hu Yang. The Gerber– Shiu discounted penalty functions for a risk model with two classes of claims. Journal of Computational and Applied Mathematics, 230(2): 643–655, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000053.

Zhang:2009:AAE

[ZLYW09] Yi Zhang, Shaoyong Lai, Jun Yin, and Yonghong Wu. The application of the auxiliary equation technique to a generalized mKdV equation with variable coefficients. Journal of Computational and Applied Mathematics, 223 (1):75–85, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707006693.

Zhang:2009:KQF

[ZLZ09a] Haiyan Zhang, Guojun Li, and Hongluan Zhao. A kind of QP-free feasible method. Journal of Computational and Applied Mathematics, 224(1):230–241, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002045.

Zhang:2009:SDG

[ZLZ09b] Tie Zhang, Jiandong Li, and Shuhua Zhang. Superconvergence of discontinuous Galerkin methods for hyperbolic systems. Journal of Computational and Applied Mathematics, 223(2): 725–734, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000708.

Zhang:2009:RSM

[ZM09a] Zhong-Qiang Zhang and He-Ping Ma. A rational spectral method for the KdV equation on the half line. Journal of Computational and Applied Mathematics, 230(2): 614–625, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000028.

Zheng:2009:TNM

[ZM09b] Bing Zheng and Shu-Xin Miao. Two new modified Gauss–Seidel methods for linear system with M-matrices. Journal of Computational and Applied Mathematics, 233(4):922–930, December 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709004695.

Zandieh:2009:SPA

[ZMAZ09] M. Zandieh and S. Molla-Alizadeh-Zavardehi. Synchronizing production and air transportation scheduling using mathematical programming models. Journal of Computational and Applied Mathematics, 230(2): 546–558, August 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270800650X.

Zhao:2009:NAN

[ZMH09] Zhenyu Zhao, Zehong Meng, and Guoqiang He. A new approach to numerical differentiation. Journal of Computational and Applied Mathematics, 232(2):227–239, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270900346X.

Zampieri:2006:AAW

[ZP06] Elena Zampieri and Luca F.

Pavarino. Approximation of acoustic waves by explicit Newmark’s schemes and spectral element methods. Journal of Computational and Applied Mathematics, 185(2): 308–325, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001172.

Zhang:2005:EPS

[ZS05] Yu Zhang and Jitao Sun. Eventual practical stability of impulsive differential equations with time delay in terms of two measurements. Journal of Computational and Applied Mathematics, 176 (1):223–229, April 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042704003309.

Zhang:2009:SBS

[ZSGF09] Hongjuan Zhang, Zhenwei Shi, Chonghui Guo, and Enmin Feng. Semi-blind source extraction algorithm for fetal electrocardiogram based on generalized autocorrelations and reference signals. Journal of Computational and Applied Mathematics, 223(1): 409–420, January 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708000320.

Zhou:2009:PGM

[ZSWL09] Xiaoxia Zhou, Yongzhong Song, Li Wang, and Qingsheng Liu. Preconditioned GAOR methods for solving weighted linear least squares problems. Journal of Computational and Applied Mathematics, 224(1): 242–249, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002057.

Zeng:2005:SSR

[ZSZ05] X. Y. Zeng, B. Shi, and D. C. Zhang. Stability of solutions for the recursive sequence x_{n+1} = (α−βx_{n})/(γ+ g(x_{n−k})). Journal of Computational and Applied Mathematics, 176(2):283–291, April 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003334.

Zhou:2009:EUP

[ZSZ09] Qiyuan Zhou, Jianying Shao, and Jinglei Zhou. Existence and uniqueness of periodic solutions for a class of generalized Liénard systems with forcing term. Journal of Computational and Applied Mathematics, 224(2): 743–750, February 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003105.

Zhang:2006:RSS

[ZT06a] Zhengru Zhang and Tao Tang. Resolving smallscale structures in Boussinesq convection by adaptive grid methods. Journal of Computational and Applied Mathematics, 195(1–2):

274–291, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004929.

Zhao:2006:BBT

[ZT06b] Qian-Jin Zhao and Jieqing Tan. Block-based Thielelike blending rational interpolation. Journal of Computational and Applied Mathematics, 195(1–2):312–325, October 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004954.

Zhou:2007:EPS

[ZT07] Yinggao Zhou and Xianhua Tang. On existence of periodic solutions of Rayleigh equation of retarded type. Journal of Computational and Applied Mathematics, 203(1):1–5, June 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706001543.

Zhu:2009:SCO

[ZTG09] Jinghao Zhu, Shiming Tao, and David Gao. A study on concave optimization via canonical dual function. Journal of Computational and Applied Mathematics, 224(2): 459–464, February 15, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708002495.

Zeng:2007:ICF

[ZTY07] Lu-Chuan Zeng, Tamaki Tanaka, and Jen-Chih Yao. Iterative construction of fixed points of nonself-mappings in Banach spaces. Journal of Computational and Applied Mathematics, 206(2): 814–825, September 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042706005243.

Zudilin:2005:WPG

[Zud05]  Wadim Zudilin.  Well-

poised generation of Apérylike recursions. Journal of Computational and Applied Mathematics, 178(1– 2):513–521, June 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042704003966.

Zudilin:2007:ATL

[Zud07] Wadim Zudilin. Approximations to -, di- and trilogarithms. Journal of Computational and Applied Mathematics, 202(2):450–459, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704270600152X.

Zudilin:2009:HP

[Zud09] Wadim Zudilin. A hypergeometric problem. Journal of Computational and Applied Mathematics, 233(3): 856–857, December 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709001368.

Zhang:2006:GOS

[ZW06a] Hao Zhang and Shuning Wang. Global optimization of separable objective functions on convex polyhedra via piecewise-linear approximation. Journal of Computational and Applied Mathematics, 197(1): 212–217, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705006540.

Zhu:2006:LIB

[ZW06b] Chun-Gang Zhu and Ren-Hong Wang. Lagrange interpolation by bivariate splines on cross-cut partitions. Journal of Computational and Applied Mathematics, 195(1–2): 326–340, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705004966.

Zhang:2008:LCG

[ZW08] Hao Zhang and Shuning Wang. Linearly constrained global optimization via piecewiselinear approximation. Journal of Computational and Applied Mathematics, 214 (1):111–120, April 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704270700088X.

Zhang:2009:ABG

[ZW09a] Liping Zhang and Soon-Yi Wu. An algorithm based on the generalized D-gap function for equilibrium problems. Journal of Computational and Applied Mathematics, 231(1):403–411, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709002052.

Zhang:2009:SOH

[ZW09b] Yinping Zhang and Qing-Guo Wang. Stationary oscillation for high-order Hopfield neural networks with time delays and impulses. Journal of Computational and Applied Mathematics, 231(1): 473–477, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709000958.

Zhou:2009:CNS

[ZW09c] Shaobo Zhou and Fuke Wu. Convergence of numerical solutions to neutral stochastic delay differential equations with Markovian switching. Journal of Computational and Applied Mathematics, 229(1):85–96, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708005220.

Zhou:2009:MAC

[ZW09d] Xue-Gang Zhou and Kun Wu. A method of acceleration for a class of multiplicative programming problems with exponent. Journal of Computational and Applied Mathematics, 223(2):975–982, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042708001192.

Zhou:2009:GTD

[ZWF09] Qiyuan Zhou, Wentao Wang, and Qiyi Fan. A generalization of the three-dimensional Bernfeld–Haddock conjecture and its proof. Journal of Computational and Applied Mathematics, 233(2): 473–481, November 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709004543.

Zhang:2009:SCS

[ZWY09] Xiaoping Zhang, Jiming Wu, and Dehao Yu. Superconvergence of the composite Simpson’s rule for a certain finitepart integral and its applications. Journal of Computational and Applied Mathematics, 223(2):598–613, January 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000538.

Zhang:2008:SOI

[ZWZ08] Kai Zhang, Jeff C.-F. Wong, and Ran Zhang. Second-order implicit-explicit scheme for the Gray–Scott model. Journal of Computational and Applied Mathematics, 213 (2):559–581, April 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707000726.

Zhao:2006:TSM

[ZX06] Huanxi Zhao and Guoliang Xu. Triangular surface mesh fairing via Gaussian curvature flow. Journal of Computational and Applied Mathematics, 195(1–2): 300–311, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004942.

Zhang:2009:FFM

[ZXZ09] Ying Zhang, Yingtao Xu, and Liansheng Zhang. A filled function method applied to nonsmooth constrained global optimization. Journal of Computational and Applied Mathematics, 232(2): 415–426, October 15, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042709003756.

Zarzo:2007:GRL

[ZYD07] A. Zarzo, R. J. Yáñez, and J. S. Dehesa. General recurrence and ladder relations of hypergeometric-type functions. Journal of Computational and Applied Mathematics, 207(2):166–179, October 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706005991.

Zhou:2009:NMC

[ZYLY09] Chaoying Zhou, Lihua Yang, Yingjun Liu, and Zhihua Yang. A novel method for computing the Hilbert transform with Haar multiresolution approximation. Journal of Computational and Applied Mathematics, 223(2): 585–597, January 15, 2009.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708000526.

Zhang:2006:SGP

[ZZ06a] Li Zhang and Weijun Zhou.

Spectral gradient projection method for solving nonlinear monotone equations. Journal of Computational and Applied Mathematics, 196(2): 478–484, November 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705005996.

Zhao:2006:UAE

[ZZ06b] Yu-Qiu Zhao and Jian-Rong Zhou. Uniform asymptotic expansions of the Pollaczek polynomials. Journal of Computational and Applied Mathematics, 190 (1–2):37–56, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705002207.

Zhu:2006:NVV

[ZZ06c] Xiaolin Zhu and Gongqin Zhu. A note on vector-valued rational interpolation. Journal of Computational and Applied Mathematics, 195(1–2): 341–350, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004978.

Zhang:2007:CIR

[ZZ07] Wanxiong Zhang and Weinian Zhang. Computing iterative roots of polygonal functions. Journal of Computational and Applied Mathematics, 205(1):497–508, August 1, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042706003347.

Zhang:2008:TDH

[ZZ08a] Li Zhang and Weijun Zhou. Two descent hybrid conjugate gradient methods for optimization. Journal of Computational and Applied Mathematics, 216(1):251–264, June 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707002348.

Zhou:2008:IAQ

[ZZ08b] Shuzi Zhou and Zhanyong Zou. An iterative algorithm for a quasivariational inequality system related to HJB equation. Journal of Computational and Applied Mathematics, 219(1): 1–8, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707003597.

Zhang:2009:VSN

[ZZ09a] Jian Zhang and Ke-Cun Zhang. A variant smoothing Newton method for P _{0}- NCP based on a new smoothing function. Journal of Computational and Applied Mathematics, 225(1):1–8, March 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708003142.

Zhang:2009:UGA

[ZZ09b] Xin-Li Zhang and Ke-Cun Zhang. Using genetic algorithm to solve a new multi-period stochastic optimization model. Journal of Computational and Applied Mathematics, 231(1): 114–123, September 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042709000454.

Zhou:2009:GCN

[ZZ09c] Weijun Zhou and Li Zhang.

Global convergence of the nonmonotone MBFGS method for nonconvex unconstrained minimization. Journal of Computational and Applied Mathematics, 223(1):40–47, January 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707006565.

Zhao:2008:TAA

[ZZL08] Caidi Zhao, Shengfan Zhou, and Yongsheng Li. Theorems about the attractor for incompressible non-Newtonian flow driven by external forces that are rapidly oscillating in time but have a smooth average. Journal of Computational and Applied Mathematics, 220(1–2):129–142, October 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042707004311.

Zhang:2006:NST

[ZZZ06] R. Zhang, K. Zhang, and Y. S. Zhou. Numerical study of time-splitting and spacetime adaptive wavelet scheme for Schrödinger equations. Journal of Computational and Applied Mathematics, 195(1–2):263–273, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042705004917.

Zhang:2008:EMB

[ZZZ08] Chun-Xia Zhang, Jiang-She Zhang, and Gai-Ying Zhang. An efficient modified boosting method for solving classification problems. Journal of Computational and Applied Mathematics, 214 (2):381–392, May 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042707001343.

Zhang:2009:MCS

[ZZZ09] Chunrui Zhang, Yazhuo Zhang, and Baodong Zheng. A model in a coupled system of simple neural oscillators with delays. Journal of Computational and Applied Mathematics, 229(1):264–273, July 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042708005657.